A Bibliography of Publications about the Matlab On-Line Matrix Laboratory

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

30 September 2025 Version 4.305

Title word cross-reference

0 = F (t, y(t), y′(t)) [Sha02]. 1 [SRK99]. $100.00 [Mar19]. 10^{435} [Mol95b]. $19.99 [Ano00]. 2 [AV15, AEF^{+}14, BCH06, BBF^{+}19, BBN^{+}22, Can22b, MV22, RV13, Sha08b, ZSW^{+}17]. 2.7 [Mol02b]. $22.33 [Ano96a]. 3 [AV15, MV22, RV13, Zek17, ZSW^{+}17]. $34.95 [Ano99b, Ano99c]. $35.00 [Ano96b]. $39.95 [Ano99a]. 4 [BFSJP^{+}21, Zek17]. $46.36 [Mar19]. $59.95 [Ano95c]. 64 [DH97b]. ^{14} [LO16]. ^{2} [FGCG94]. ^{(}R) [LS04, SM09, Sta05a, Sta05b, Mar14]. α [dVSWAL17]. β [Rum17, Rum23]. COMPl_{e}ib [Hen07]. D [ATvBB17, CZ17]. Γ_{k}^{n}(x) [PF07]. H_{2} [Bur99]. H_{∞} [Bur99, Hen07, KC95]. hp [PR14]. K [BHC^{+}24, Spe95, LLW^{+}23, Rum17, Spe96]. k < m [Rum17]. L_{2} [HAM02]. Λ_{k}^{m,n}(x) [PPS06]. LL^{T} [BSB23]. m [Rum17]. R^{n} [Kat09]. μ [GMT95, GPK05, TACA15]. ODE23 [LS98]. $p [TKD00]. p [Rum23]. T [Wit04].

-Adaptive [PR14]. -Analysis [GPK05]. -bit [DH97b]. -D [Zek17]. -diff [TACA15]. -Means-based [LLW^{+}23]. -norm [HAM02]. -Optimal [ATvBB17]. -Space [Spe95, Spe96]. -Stable [dVSWAL17].

/EUFIT [Ano94h]. /Octave [MBR21].

0 [Ano93b, Sha04, Tay99]. 0-534-93273-8 [Ano93b]. 0-8493-2016-X [Tay99].

1 [Bät20, CV00, Elg01]. 1.0 [RHR^{+}21]. 1.02 [Stu99]. 1.3-$mu [HA95]. 1001 [BKL19]. 1008 [CH20]. 1021 [LCMCD22]. 1030 [BBEM22]. 1040 [PT24]. 10th [Ano94k, IEE96c, Ter94]. 11p [CPUARC20]. 128-bit [DH00, Mol17]. 12th [Ano96t, Ano96s, Bub95]. 13th [D^{+}95, GCP97]. 15th [AO95, W^{+}97]. 16th [WD98]. 17th [KPN^{+}04]. 19 [CK23]. 19.50 [Ano99b, Ano99c]. 1991 [Bar92, Lum91]. 1992 [Lum92]. 1993 [EM94, Sin93]. 1994 [Ano94h, Ano94i, AO95, BS95a, IEE94b, IEE94a, IF94, M^{+}94, Pat94, Ter94]. 1995 [Ano95i, BH95, DG96, FB95a, IEE95b]. 1996 [ACM96, C^{+}97, IGPF96]. 1997 [Ano97h, BLM97, BB97, IEE97a, IEE97b]. 1998 [ACM98]. 19th [Ano96g, Ran96].

2 [Ano97e, CW05b, Elg01, TCE16]. 2.0 [HTNFBS06a, JZW^{+}22, LV16, SBL^{+}10]. 2.1 [HTNFBS06b]. 2000 [Ano97a]. 2003 [ACM03]. 2004 [IEE04, UCL04]. 20th [KZL^{+}20]. 21st [Ber95, Bry96, Bud95, BB97, Jef08, Mal96, BM09]. 23 [Alo06]. 233 [Ano97c]. 233-MHz [Ano97c]. 23rd [KF96, Gra93]. 2445-29 [GSM95]. 24th [Ber95, IEE94e, PJ97]. 25th [Bud95]. 26th [IEE94e, I^{+}96]. 27.00 [Ano99a]. 27th [EP 97]. 28-30 [BLM97]. 28th [IEE95b, Ano93a]. 29th [IEE95b]. 2D [Can22a, Can24, FMS25, HCBAEC23, Mil20, WL94]. 2e [Ano05]. 2nd [Ald96, Ano09, Atk05, Lak97, USE99, Veh07]. 2RRR [MMPOQ^{+}25].

3 [CZ17, MM00, Sha04]. 3-540-44363-0 [Sha04]. 3.0 [Ano97j, Bra97a, Han99, Váz16]. 3.8 [EaoGOBHW14]. 32nd [IEE93b]. 33rd [Ano94i, IEE94d]. 35th [IEE96e]. 39th [C^{+}96]. 3D [FMS25, OBD23]. 3DRSP [BC22]. 3M [CW05d]. 3Ms [CW05b, CW05c, CW05a]. 3rd [Ano95f, Atk05, Coo95, FB95a, Ken95, MO95, Mun12, Ni22].

4.0 [GR93, Spe95, SK95]. 45 [ATvBB17]. 49 [ÖK17]. 4th [Ano93e, IEE94a, IEE95b, OB93].

5 [Ano97f, Ano97g, Ano97k, Bäc97, BB99, Bra97b, HL98, Hau97, Hul99, Kad97, McC98, MM00, PES98, PES99, Pra99, RC98]. 5.0 [Ano97a, Ano97c]. 5.2 [Han99]. 55 [BD20]. 5th [Ano97b].

6 [BB02, EKH02, HL01, HLOR05, Kin01, Kin06, Mar19, Mol00a, Pal01, Res19]. 6.5 [GG04]. 694 [Hig91]. 6th [Ano95s].

7 [Bäc04, EMK04, EKM05, HL05, HLOR05, Pal05b, Pra06, ÜKP05]. 7.1 [BO19]. 7/a [Pra06]. 70-99 [Mun12]. 754 [Rum23]. 756 [Dri96]. 7th [Ano96u, BCD^{+}20, BLM97, Bry96, dOSRS19].

8 [Ano93b, HLR14]. 802.11p [CPUARC20]. 802.16 [Kum14]. 805 [KO00]. 808 [SN01]. 827 [BCR03]. 832 [Dav04b]. 835 [Zen04a]. 836 [DGLN04a]. 837 [ADD04]. 843 [Dri05]. 844 [BPS05]. 845 [MY05]. 847 [KW05]. 849 [Dav05]. 858 [VC06a]. 85th [dOSRS19]. 862 [BK06]. 866 [ERS07]. 868 [Esp07]. 873 [RSS08]. 875 [BY08]. 876 [AS08]. 878 [Jon09]. 884 [DS09]. 899 [Sar10]. 8th [Ano97i, H^{+}96].

90 [Ano97k, Bra97b, DP96b, DP99]. 902 [RBD^{+}10]. ’91 [CF93a]. 916 [ZA11]. 918 [ST12]. ’92 [Van92b]. 927 [CHMN13]. 929 [MG13]. 93 [Mic94b]. 930 [Dav13]. 935 [RKZ^{+}14]. ’94 [Ano94h, Ano94d, Ano94i, Dag94, IF94, M^{+}94, CAC94, IEE94c]. 941 [KT14]. ’95 [Ano95f, Ano95n, BH95, FB95a, Ken95, Mro95b, Hor97, LS05a]. ’96 [ACM96, Ano96e, Ano96v, IEE96c, IEE96b, IGPF96, Sil96]. 960 [DP16]. ’97 [Ano97b, BF97, BLM97, IEE97c, IEE97d]. 973 [DMB17]. 974 [NK17]. 978 [Mar19, Res19]. 978-1-4398-1624-0 [Mun12]. 978-1-4987-7606-6 [Mar19]. 978-1-78548-004-1 [How15]. 978-3-527-40926-6 [Res19]. 981 [Riz17]. 986 [MP17]. ’99 [USE99]. 99-line [YJ24]. 996 [IKK^{+}19]. 998 [AFOP19]. 9th [BF97].

= [BCH06, Oga10].

A&M [Dut16]. Aachen [Ano94h, Ano95f, Ano96e, Ano97b]. AADL [XWZ^{+}22, XWZ^{+}23, ZLW^{+}19]. AAPR [BB97]. AARTC [FB95a]. Ab-initio [XSS20, ZJKS23]. abstract [KSS07]. AC [Att95b, Att95a, Bos01, Bos02, Jac04]. accelerated [Cap13]. Accelerating [DEQOR13, SZCP21]. Acceleration [ASP^{+}18]. accelerators [App19]. Accessing [FR96]. accompany [DIN06, SYTD04]. Accuracy [Hig96, Hig02b, TS21, FBFB04, LNLB19]. Accurate [CH20, SP91, Sha07, TS14, BV21, Jon01, LJ23, PK08]. ACE [IF94, VHM17]. ACEInhibPKPD [VHM17]. achieving [LNLB19]. acid [BSS^{+}23]. acidified [BSS^{+}23]. ACM [ACM97, ACM98]. ACM/IEEE [ACM97, ACM98]. Acoustic [And95]. Acoustics [IEE96f]. Acquisition [FPBO98]. ACSL [VVM93]. Action [RD22]. actions [Wil99]. active [Mea02, SZM^{+}14]. activites [Gro99]. Actuator [UB95, SM95]. AD [SKF05]. Adaptive [AJGO^{+}20, BCHS98, CiA02, CSG98, Din02, FB98, Hay96b, Hay02, IS96, Joh18, KBKS98, PR14, PR06, RBZ96, RSW15, SCC95, SGLBA98, Wel93, AEF^{+}14, Esp07, IP23, MIK00, Say03, Sha08c, Spo02, WSST05]. added [Ano93c]. Adders [Mik23]. Adds [Ano97f]. Adelaide [Ano93d]. ADiGator [WR17]. Adjoint [WB12]. Adjustable [Ste09]. adjustment [Del02]. ADMIT [CV00]. ADMIT-1 [CV00]. ADSL [Bin00]. Adsorption [Do98, Duo98]. Advanced [Bha06, Bur01, C^{+}97, Duf03, Duf10, Duf11, EM94, HDR97, HDR00, Jac04, MM97, MM98, Moh01, MBMW95, Pac00, Rav94a, Shi98a, WT94, WT97, WHT02, WTH03, SLM23, C^{+}97, MO95, Mun12]. Advances [BAY98, BBH^{+}08, FHP^{+}12, IEE93a, BF97, IF94]. Advection [RAW^{+}16]. aerodynamics [BGGL20]. aeroservoelastic [LB99]. aerospace [Mag05, MAB^{+}11, Whi91]. AFDeter [YHC^{+}22]. AFDX [TBH21]. affine [JR99]. Affordable [TS14]. Africon [Van92b, Van92b]. Age [LO16, Sch99]. aid [FSC95]. Aided [BM09, CF91, GR97, IEE96d, MGC94, M^{+}94, UB95, Bar92, BLM97, BM94, Bry96, GG92, HA95, Mro95b, Szy93]. Aids [Saa93]. AIR [HSH12]. Aircraft [Bry94, PFG08, SL03, Sch98, SCB99]. airflow [YHC^{+}22]. airflow-related [YHC^{+}22]. Airplane [Bro94, Man96]. al [EGE95, GSS05]. Albany [IEE95b]. Alberta [Mal96]. Albuquerque [Ano97h]. Alfonso [PY22]. Algebra [Bor18, Bou97, BV18, Dat95, DR96, Dem97, FB95c, Gra04, Gro94, HM88, HZ94, Hil96, HZ96a, HZ96b, Hog07, JRA98, KK01c, Kol93, LE96, Law96a, Law96b, Lay94, Leo94, LHF96, Leo98, LHF03, Lev92, Mor98, ND88, OS06, PBB22, SCL95, Str93, VZDCM25, Wil96a, Ano96a, AB03, Bee05, Ben98b, BQOvdG05, But08, But11, Dav12, GD99, Har05, HK01b, JRA02, KSS07, KH97c, Lay97, Lay03, Lay06, Leo02, MDB01, MY98, PR02, Pen04, Rob95, Smi97, SR90, SB97, Str03, Str06, TB97, Uhl02, Ano95c]. Algebraic [Che93, FK11, Glo98, NPT15, PNT15, ARRY01, HL03a, HSH12, JB06, SHX96, ZL13]. Algorithm [AS15, CMR17, EM14, FGjS15, Gau16a, ICS^{+}18, JKR92, JMD08, Joh18, LLW^{+}23, NPT15, Qur01, TH01, WR17, Zag16, ADD04, BK06, CFPF94, CF95, Cur95, DGLN04a, DGLN04b, HMT13, Kea17, MS14, Mol95c, Mol01, Rob96, Rov90, RB04, RB05, SAKG15, SZCP21, UW12, VC06b, Wit04, Yan17, AFOP19, ADD04, AS08, BK06, BCR03, BY08, BPS05, BBEM22, BKL19, CH20, CHMN13, DGLN04a, Dav04b, Dav05, Dav13, DMB17, DP16, DS09, Dri96, Dri05, ERS07, Esp07, Hig91, IKK^{+}19, Jon09, KW05, KO00, KT14, LCMCD22, MG13, MP17, MY05, NK17, OF92, PT24, RBD^{+}10, RBD^{+}11, RKZ^{+}14, Riz17, RSS08, ST12, Sar10, SN01, VC06a, ZA11, Zen04a]. Algorithmic [FHP^{+}12, New93, WR17]. Algorithmic/Conceptual [New93]. Algorithms [BD20, BL96b, BL96a, CFGG94, DP16, Hig96, Ife05, ICS^{+}18, Jon91, KR22, KBKS98, MBS15, NN94, ÖK17, SBL^{+}10, Wei12, AHS94, BPS02, BQOvdG05, Bob05, BPS99, Bry02, CiA02, Din02, FGCG94, GC99, HBT16, HH04, Hig02b, HSR01, KR23, MG13, MS00, Moo05, MA95, Nab02, Rum01, Sou99, SD96, TS21, Ves98, WS04, FB95a]. Aliasing [Ano95u]. Alignment [MJ01]. Alley [Qur01]. Allocation [Nyh08, Meu05]. along [HG24]. Alpha [KN95]. Alphas [Kad97]. Alternative [Bäc04, ES10, Stu96, MSS^{+}19, Ste08, Ste13]. Amalgamation [Kum14]. AMD [ADD04]. American [Ano95k, Ano97h]. Ames [C^{+}96, IEE04]. AMF [BGPPRW14]. AMF- [BGPPRW14]. AMGKQ [Joh18]. AMPAO [KAR^{+}19]. AMPL [VFG04]. amplifier [Kar05]. amplitude [JAC20]. Amsterdam [Ano95n, KZL^{+}20]. Anaheim [Ano95k]. Analisi [Com90, MR95]. Analizi [Yuk96]. Analog [Amb95, Amb99, ASP^{+}18, Che93, Cou97, Mea02, Su96, cS02, Cou01, Nas01, Paa01]. Analogue [TH01]. Analyse [Sch94]. analyser [LSH95]. Analysing [SM14]. Analysis [Alo06, ÁBZ20, Alv11, Ano88, Ano95e, Ano96r, Ano12, Ans14, Atk05, Att95b, Att95a, AY96, BV95, BDM17, BDM18, Bha05, Bis93, Bow10, Bur10, Car10, Car92, Che94, CF93c, Com90, CSY15a, CSY15b, Cor95, CM67, DH95, DL95, D^{+}95, DH12a, DK99, Fau08, FSO93, Gar96, Góm15, Got95, GPK05, Han06, HK95, Han92, Har02, HM19, HPK18, Irw05, Jer06, KAR^{+}19, KK96, KR22, Koc14, KT10, Kum14, KH94, Kus06, KYN95, LSvdVK09, LSvdV19, Mah00, Mah05, Mar19, MR95, MASV96, MVM97, MBMW95, Nak96, NPT15, NGKM18, New94, Ni22, NP93, Pao99, PN95, PSR16, PNT15, Rah93, RA95, RJ93, RT96, RD22, Rul02, SMS95, SBL^{+}10, Sch94, SR09, SM95, Stu96, SU93, TH01, Tho94, WW94, Whi04, W^{+}97, WD98, XWZ^{+}23, YR19, ZL17]. Analysis [Zla17, AM01, Ajj95, Ano95l, Ano97l, Ano98b, AH04, Att99, Att04, BDR04, BHC^{+}24, Ben98a, Ben04, Beq98, Bha06, Bis97, BAEBAS19, BN01, Boo04, BlD^{+}20, Boy99, Bre03, Bur05, Car98, Car00, CK02, Cat01, CF16, CFN01, CFN02, CM99b, DHS03, Dat04, Dav98, DL01, Den98, Dha03, DGK03, DGK04, DDN02, Do98, Doe98, DHR11, DM06, Duo98, Elg01, ELR02, Fau99, FS23, GW04, GL04, GS12, HBT16, Han94, HG24, HB04, HKD02, Hen94, Hsu99, Irw02, Jes01, JS24, Jon95b, Jon95a, Kar01b, Kar03b, Kar03a, KR23, Kho00, Kin97, Kin93, Kul99, Lea04, LYH^{+}16, LB99, Mar14, MM05, MMS11, MMS17, MF94, McG91, Mid00, Moh01, Moi10, Nak02, Nor04, OBCG19, Paa01, Pal98b, Pao01, Pau00, Pau01, Por23, PSP04, RS02, RHG09, RH00, Rob04, Roq13]. analysis [RL00, SSTD03, SG09, Sch12, SFPO94, STC04, Shi99, Sil96, SP96, Sta03a, Sta03b, SM97, SM05, Sto13, Syd95, Tay99, TR04, TSM93, TtsT04, UCL04, UW12, Van04, Váz16, Wal02, WPK^{+}18, Wet20, Whi00, XZL24, XWZ^{+}22, XCA07, ZL04, ZE95, tHLMN19, LMN18, Yuk96, DIN06]. analysts [Lak97]. Analytic [Dor00, EP94b]. analytical [BTM09, Goc02, GS12, MB94, Ren17]. Analyze [LL92, LL95, WCS92]. Analyzing [SM07, Cat95]. ANDECS [BJG94, FBG94, PGBG94]. ANDECS/MATLAB [FBG94]. ANDURIL [tHLMN19, LMN18]. ANDURYL [tHLMN19]. Angel [Alv11, Ano12, Jer06, Kus02, Kus06, Mar14, Mar19, Ni22, Wik04]. Angewandte [Bät20]. AnIMAGE [VRVAC23]. Animation [Gab98]. anisotropic [RCT20]. ANNIE [Dag94]. annotation [Ozk23]. Annual [Ano94e, Ano95k, Ano96t, Ano96u, Ber95, Bud95, EP 97, Gra93, IEE97c, KF96, Lum91, Lum92, Ter94, Ano94i, Gra94, I^{+}96, Jef08]. anomaly [BSB20]. ANS [JE91]. ANSYS [BR96, HC00, Hat01, Bry96]. Ant [TH01]. Antenna [DM96, EI05, Hau90, Mak02]. Antennas [IEE95a, Lev96, Gro05, VRVAC23, VCK98]. Anterior [RD22]. Anti [BA94]. Anti-Windup [BA94]. Antonio [IEE93b]. Användarhandledning [PES98]. Anwendung [Ben98b]. Anwendungen [von93, Gra04]. anwendungs [Bät20]. anwendungs- [Bät20]. anwendungsorientierte [Beu05]. Anymatrix [HM22]. anymore [Mol15]. aperture [Sou99]. API [FSZD20]. APL [Gid95]. apology [Mol99a]. app [VHM17]. app2web [VFV13]. Applicate [Com92]. Application [Ajj95, Beu05, CKC94, HPZ19, Jon95b, KPS95, LPD^{+}17, Pol92, Riz17, dP96a, Ano94e, Are94, Bät20, BQOvdG05, CZ17, Hsu99, Jon95a, Por23, Tob11]. Application-oriented [Beu05, Hsu99]. Application-Specific [LPD^{+}17]. Applications [AM22, Ano93e, Ano95s, ASP^{+}18, Bha05, BL96b, BL96a, BGGT21, Cha17, Col04, CdFCS98, Cou95, Dat95, DR96, DM90, Gol91, Hil96, IEE94e, IEE95b, IEE97c, IEE97e, KBKS98, KSS07, Kol93, Lay94, Leo94, Leo98, Lu96, MBMW95, Per91, Sch14, VFV13, Wil96a, WT94, WT97, von93, Ali02, B10, BSLK01, BPS02, Bau02, Ben02, Ben98b, BN10, BN11, Bob05, Boo04, Boo98, CCM^{+}03, Cha02c, Cha05c, CAC94, Chi97, CF93b, CF93a, CiA02, CM99a, Cou93, Dav12, DG96, DH97a, FB98, Gil04, Gil05, Gil08, GS08, Gop10, GC99, Har01, HLS08, Jac04, KF96, Kar03b, Kar03a, Kar03c, Kar05, Kat09, KM10, KM11, Kog97, KH97c, KL01, KG05, Lay97, Lay03, Lay06, Leo02, LB99, LCL05, LB00, Lyn04, Lys00a, Lys01, Mac05, Mag05, MAB^{+}11, MSL09, MC04, MM00]. applications [Moo98, MY98, MGT03, PK04, Pen04, RB98, Rec00, RHR^{+}21, Rib00, Riz00, SAE95b, SSTD03, Sem04b, Sta03a, SWrpSetfe02, Str06, SS10, Tho04, TD98, Tra04, TKD00, TW98, VCK98, Wen05, Whi91, WHT02, WTH03, Wun05, vdM96, Com90, Gra04]. Applicazioni [Com90]. Applied [Ano95h, Ano95v, BN00, BV18, Bry02, But11, Cha05b, Com92, Coo14, Dem97, Esf03, Fau99, Fau08, GW04, Gri95, IEE04, Lev03, Mar03, Mar07, MF02, Moo98, Mor00, ND88, Ogu95, OS06, RS02, Rao02, RJ93, Rog03, SH00, Ter94, Ven02, WA96, YCCM05, Ano96t, BH97, But08, Dav04a, KHK05, LSH95, Lys00b, Shi99, Ano95b, Bät20, Veh07]. Applix [Ano96h, Ano96i]. Applixware [Ano96h]. Applying [KC95, MNHH94, Cat01]. Apprendre [MMR97]. Approach [Ano98c, BCK96, Can86, DRR97, GK05, HBC94, HM19, Kat03, KG01, NPP04, New93, Vac95, vdH05, Att02, Att10, BF09, Bro01, CCM^{+}03, CSV96, DCF95, Dor00, FD01, Gaw98, GP98, Gri94, HGM^{+}25, JSM97, KG06, KG12, LH13, Lei02, MSY98, MC02, Mit98, Mit06, Nor05, Rao11, Ros03, SHX96, SB10, TYL^{+}16, TC13, Van97, Van00a, ZM00]. Approaches [DL95, DL01, HT97, TU97]. Approximate [AM22, OBD23, ADD04, DGLN04a, DGLN04b, Mil20]. Approximating [GBM15]. Approximation [KR22, TCD^{+}22, Fas07, Gau06, KR23, PT24, PPS06, PF07, Tre86]. approximations [BPS05, PC13]. approximator [Spo02]. April [Ano94c, Ano97i, Ano97l, BLM97, Ham96, IEE96c, IEE97d, NP96]. AQUINAS [FS23]. Arabie [Lip07]. Arbitrarily [OBD23]. arbitrary [Cur95, VC06a]. Archaeological [BB96]. archaeology [KF96]. Archimedean [GHH20]. architectural [AI21]. architectures [Boo98, FB95a, Jes01, KG05]. ARfit [SN01]. Arguments [BV08, Bat19]. Argyris [DS09]. Arithmetic [Cse99, HP19, Mol17, Rum17, Rum23, Ste09, BMR19, Meu20, Mol96a, Rob94, Rum95]. Arizona [MGC94]. Army [DLA^{+}16]. Arnoldi [Rad96]. ARPIST [LJ23]. Array [Hau90, MJ01, Sou94, JB03, JB06, MIK00, Van02, dP96a]. Arrays [TRD11, FBH17]. Arrival [AS15, CS12]. Art [DTB97]. artificial [CAC94, Dag94]. Ash [OF92]. Asia [Ano95d, IEE93a]. Asia-Pacific [IEE93a]. ASME [Kin93]. Aspect [MLF^{+}12]. aspects [OB93]. assembly [AV15, BCG17, EL16, KA13, RV13]. Asservis [Bou95]. Assessment [FFR^{+}24, Cse92, FJ22, How91]. Asset [Nyh08, Meu05]. assignment [Whi91]. Assistite [CF91]. Associate [Ano95v]. Association [Ano95o, BB97, LYH^{+}16]. asymmetric [FD01]. Asynchronous [SFPO94]. Athens [Lak97]. Atlanta [Ano94e, Bud95, IEE96f]. ATLAST [LHF96, LHF03]. Atoms [Che18, Res19]. ATOSEC5 [BRS94]. ATP [BRS94]. Attenuation [Kum14]. AudExpCreator [NK18]. audio [ZA02]. auditory [NK18]. Aufgaben [BD95, BD00]. August [Ano94a, Ano95f, Ano96s, C^{+}96, Glo98, H^{+}96, IEE94e, IEE94a, IF94, Kin93, KPN^{+}04]. Auslegungskriterien [Bro94]. Auslegungskriterien-Regelungsstrukturen [Bro94]. Austin [IEE94c, USE99]. Australia [Ano93d, Ano93d]. Australian [Ano95o].

Austria [BH95, BB97, Jef08]. Austrian [BB97]. author [Nah02]. Automated [BBB12, Edd09, UW12, ZSW^{+}17]. Automatic [Bir18, BBH^{+}08, Brz97, CMKH03, CF16, DM90, Dja98, Dja00, GCP97, HR91, HWB15, KAR^{+}19, KW95, Kuo95, KG03, MAC08, MJ01, Nei10, PFG08, PAG11, RVV^{+}92, RH92, TCD^{+}22, WR16, BEV06, CV00, For06, HS24, HKF^{+}20, JGGF23, ZC08]. Automatica [Bol94]. automatically [Cam06, Riv01]. automatically-solved [Riv01]. Automating [WS04]. automation [Ano94b, Kin93, Pas04, Bol94]. Automatyki [Brz97]. Automatyzacja [BF95]. Automotive [WM95, WM96, LNLB19]. Autonomous [CBhM12]. autopilot [Jon95b, Jon95a]. autoregressive [SN01]. Auxiliary [Gid95]. available [Mol04a]. Avionics [BBB12]. Aware [SYW^{+}24]. Axially [FKSM97]. AXIOM [Ben98b]. axisymmetric [FS23]. axle [SAE95a]. Azure [TCE21].

B [Mic00]. Backhaul [Kum14].

Backpropagation [VA94]. backscatter [ZE95]. Backstepping [RBZ96]. Balanced [Gaw96]. Baltimore [IEE96a]. band [DG04]. Banks [SN96]. BANSHEE [KMLP^{+}23, PMNWR20]. Barbara [WD98]. base [Rum17, Rum23]. base- [Rum17, Rum23]. Based [AC96, BRPCR94, BDM17, BDM18, B^{+}94, Can86, CGM95, Coo96, DMB17, DDD97, DK99, FFR^{+}24, GV25, GJ20, IoG10, KDAB19, KA09, MLAS21, MBMW95, Nat94, RG21, SDPM04, SU94, WTL00, YKS94, ZL17, Zla17, AS15, Ano96r, Ano97l, BKG05, BBHP^{+}23, BC22, Bra02a, Bra06, BSS^{+}23, BJ02, BEK99, Cap13, CK22, CK23, CDOBA22, Deg20, FGMS21, GI06, GRH^{+}21, GMT96, HBT16, HA95, IZBT21, ICS^{+}18, JS24, JGGF23, Kat09, Kea17, KS01, Kum14, Lei02, LLW^{+}23, LHZT23, McC98, Men01, MVDV97, Mit98, Mit06, NK18, Nyh08, OAKS11, PC13, Por23, RC13, RC16, Ros03, RB04, RB05, RLV11, SSW09, Sem04b, Sem05, SAKG15, SZM^{+}14, Som07, SLM23, TB95, TM97, TC13, UW12, VRVAC23, WEM98, WB16, Wri97, XZL24, YHC^{+}22, ZC08, ZZC^{+}08, ZZG^{+}14, ZLLT23]. based [I^{+}96, RLH03]. Bases [KT10]. Basic [DIN06, Irw02, Irw05, LY97, Pfe95, HAM02, dP96a]. Basics [Kni00, tC04, Ega98]. Basing [Cze95]. Basis [Sar17, Che04, JZW^{+}22]. Baton [NP96]. Bayes [Sto13]. Bayesian [Ano94e, KMLP^{+}23, Row03, ATvBB17, LO16, PMNWR20, She10, Sto13, YR19]. BBCPOP [IKK^{+}19]. BCI [LHZT23]. BDEtools [ADW23]. be [FBG94]. Beach [IEE95a]. beam [SLM23]. beams [HA95, App19]. beginners [HLR01, HLR06a, HLR14]. Behavior [MNBB19, Mol01]. behaviour [AM95]. behavioural [JS24]. Behind [ANV00]. Beispiele [von93]. Beispielorientierte [Hof98]. beispielorientierter [Bät20]. Belgian [AO95]. Belgium [BLM97, FB95a, Sas96]. BEM [AEF^{+}14]. Benchmarks [Mol94]. bending [Roq13]. BER [Kum14]. Berlin [C^{+}97, Sha04, Syd95]. BertiniLab [BNN16]. Bessel [RKZ^{+}14, VC06a, VC06b]. beta [Ano97g]. better [Led01]. Between [MBL^{+}97, LNLB19, MP18, QMS98, dP96a]. Beyond [Chr97, ANV00, CT97, Kni00]. BFDA [YR19]. BFGS [EM14]. Bibliography [Bee94]. BIFTOOL [ELR02]. Bifurcation [KYN95, NGKM18, Ajj95, DGK03, DGK04, ELR02]. bifurcations [BFG^{+}14, GKD05]. big [HL22]. billion [Mol02b]. Binary [ATvBB17, IKK^{+}19, Rum23]. Binomial [Hig02a]. biocontrol [CH23]. Bioengineering [O’B13, KM10, KM11, Vid11]. bioengineers [Sem05]. biofeedback [CK22, CK23]. Biological [VBB18, BDL25, Cun03, EU07, MAB^{+}11, TYK04]. biology [AR03, AR04, SH09]. Biomedical [Bru01, Aka00, Ano97l, Bau02, EBB05, Pat94, Sch12, Sem04b]. Biomimicry [Pas04]. Biomolecular [VPM16]. BioNES [CK22, CK23]. Biosignal [Sem04b]. Bird [BGGL20]. birthday [dOSRS19]. bisection [Che08]. bit [DH97b, DH00, Mol17]. Black [Chr97, CT97, Zla17]. Blacksburg [IEE97d]. BLAS [FSZD20]. BLASFEO [FSZD20]. blind [CiA02]. block [Por23]. Blockchain [Shu17]. blocking [HG24]. Blocks [BBC^{+}94b, BBC^{+}94a, FFR^{+}24]. blockset [CSB03]. Boca [Ano99a, Mar19]. bodies [MP18]. Body [Gab98, BKR97, FCP97]. Bolts [Ano97c]. Boltzmann [Asi10]. Bond [GR97, MVDV97, mP97]. Bondlab [MVDV97]. Book [Alv11, Ano93b, Ano95c, Ano96a, Ano96b, Ano99a, Ano99b, Ano99c, Ano00, Ano09, Ano12, App19, Bar16a, Bar16b, Bow10, Bur10, Car10, Epp11, Han06, How15, Jai09, Jer06, Kus02, Kus06, Lip07, Mar14, Mar19, Mic00, Mun12, Myr17, Ni22, O’B13, Pfe95, Pie05a, Res19, Sha04, Shu01, Ste96, Tay99, Veh07, Wik04, Smi97, Ver09]. Bookware [Pfe95]. Boolean [ADW23]. Boost [AM17]. Bootstrap [ZI04]. Bose [Hoh14]. Boston [Ano93b, Ano95s]. Both [Mat94a]. bottom [ZE95]. Boundary [CHMN13, EP96b, EP04a, BD05, EP00a, KK93, RC13, RC16, YJ24]. bounds [Cha95, FBFB04]. Box [IKK^{+}19]. Brain [RD22, LHZT23]. brain-computer [LHZT23]. Brani [O’B13]. Braunschweig [BCD^{+}20]. breakdown [JS04]. Breast [Shu17]. Brian [Ano00]. Bridge [KN95, Kad95, Can24, Can22a]. Bridging [QMS98, PPL^{+}18]. brief [Mol18a]. Brightness [SS18]. bringing [Ano93c]. Bristol [Tay99]. broader [HL22]. Brushless [ÇA10]. Budapest [JLM96, Zem97]. Buena [Ano96u, IEE94d].

Buffer [DP08, BSS^{+}23]. build [Led01]. Builder [Ano97a]. Building [BBC^{+}94b, BBC^{+}94a, C^{+}96, The93a]. Bulgaria [Zla17]. bulk [FMS25]. bulk-surface [FMS25]. bus [SAE95a]. Business [And05, Kar01a]. BVP [KS01]. BVPs [SKF05]. bvptwp.m [CHMN13].

C [Ano97d, Mei05, Tay99, Ano97a, AC99, BC17b, Cha02c, CA97, EO94, Ife05, Ker95, LPD^{+}20, Led04, LS04, LO16, MNHH94, Mir96, Mir97, MP18, RBC20, dVSWAL17, RS09, SH00, SBL^{+}17, Wal18, WEM98, WWM06]. C# [MP18]. C. [Mei05]. C/C [Led04, dVSWAL17]. C/OpenCL [RBC20]. C6416 [Cha05c]. C6713 [Cha05c]. CA [Ano93e, Ter94, Ano96t, GCP97, IGPF96]. CAA95 [KF96]. cable [WCS92]. CACSD [BLM97, Ano96c, Ano96p, HTCI96, SMS96]. CAD [Kin93, DM96, SHX96]. CAD/CAM/CAE [Kin93]. CAE [Kin93]. Cage [CHT15]. calcium [HG24]. Calcolatore [CF91]. Calcolo [CF91, FT92, MR95]. calculate [CATK11, LHW01]. Calculating [WK95]. Calculation [BKR97, CH20, HG24, Beu05, SZM^{+}14]. Calculations [CH20, FT92, XSS20, Ano95m, Dat13, NSXZ14, Pha99, VPM16, ZJKS23]. Calculus [BS95b, EP94b, EP94a, EP02, HC95, Mac05, Coo01, PK04, MR95]. Calgary [Mal96]. calibrated [FBFB04]. Calibration [LO16, HTNFBS06a, HTNFBS06b, TNBSF04]. Calif [WD98]. California [ACM97, Gra94, Ano95k, IEE94b, IEE95a, Joh95, Kin93]. CAM [Kin93]. Cambridge [KPN^{+}04]. cameras [Ano97c]. Camp [GR97]. Camp-G [GR97]. Campbell [Ano99b, Ano99c]. Can [Ste08]. Canada [Mal96, BEH94, OB93]. Canadian [BEH94, Mal96]. Cancer [Shu17]. CANCSD [Ano96n, NRPH96]. Cannon [Ano99d]. capabilities [Mol00b]. capacity [BSS^{+}23]. capillaries [LLLW06]. CAR [LSvdVK09]. Caratheodory [BV19]. card [SW95]. cardiac [HG24]. Cardiovascular [DS16, NMS^{+}06]. cardiovascular-respiratory [NMS^{+}06]. Carlo [CGRvD15, Mil20, NSXZ14, SAKG15, She08a, She08b]. Carlton [IEE96d]. carried [Ans14]. cascade [SZM^{+}14]. Case [Gou20, MP99, MP00, QCPG96, SYW^{+}24, Asi10, BF09, BGGL20, Mac04, Moe04, RS02, Sch12, TJML92]. cast [GO97]. CaTchDes [BV19]. CAV [BM09]. cavities [Deg20]. ’cbnet [BD20]. CD [Dak06, EI05, PK04]. CD-ROM [PK04]. CDMA [Lan00, VLV00]. CE2014 [MBS15]. Celestial [Ono01]. cell [PSP04]. Center [ACM98, Ano97h, BEH94, IEE94c, IEE96e, Ter94]. centered [Bro01]. Centre [DG96, Van92b]. Centroid [GPR14]. Century [Ber95, Bry96, IEE97d, Bud95, Mal96]. cerebral [dMMLOS20]. Cergy [Ber95]. certain [DH97a]. CF [Tre86]. Chain [Bar97, BBB12, She08a, She08b]. Chains [SMS96, Ano96c, PNGR00]. challenge [Ano95d]. chance [Mol95a]. Change [MON12, EP 97]. Changepoint [WTL00]. channel [FD01, HG24]. channel-blocking [HG24]. channels [Pät02]. Chaos [Gul12]. Chaotic [WK95]. \ [Kah04]. Character [Kar93]. characteristics [BO19, SZM^{+}14]. Characterization [Mar16]. chebfuns [PPT10]. CHEBINT [PC13]. Chebyshev [AT17, DMB17, PC13]. Chelikowsky [Res19]. Chemical [AB96, HPK18, Bre03, CM99a, CS99, EU07, Fog99, WB99, Löw01]. chemically [Bha95]. Chemische [Löw01]. Chemkin [Ano97j, Bra97a]. Chemometrics [Bre03, tC04, HTJ90]. chemometrics/statistics/neural [HTJ90]. Chemotherapy [Shu17]. chercheurs [MM10]. Chicago [IEE98]. Cholesky [Dav05]. Choose [CLMM20]. Chopping [AT17]. Chores [CA97]. Christoffel [Dri96, Dri05]. chromatography [MMD24].

Circles [Var04]. CircStat [Ber09]. Circuit [Att95b, Att95a, DL95, DIN06, Irw05, ICL97, Kar03b, Kar03a, TH01, Att99, Att04, Dav98, DL01, HKD02, Irw02, LB00, Pau00, Pau01, RR02, Sta03b]. Circuits [Ano95p, AS96, C^{+}96, Car00, Con95b, Con95a, Con96, Got95, NR96a, NR96b, Sem05, TH01, Boy99, DS04, Jac04, Kar05, TR04, VCK98]. Circular [Ber09, Nas20]. Císlicové [VS00]. City [D^{+}95, I^{+}96]. civil [Mag05, MAB^{+}11, SCB99]. Clara [Ano93e]. Class [CH20, Kub95]. classes [BK06]. classic [SAE95a]. Classical [LE00, SKA19]. Classification [Rao11, SWS97, vdH05, BHC^{+}24, DHS01, TSA21, ZE95, SYTD04]. Classifying [TKD07, Pac04]. Classroom [KH96, KH97b]. Classwms [BHC^{+}24]. Cleve [Dut16, Hai08, Mol94, Mol95c, Mol95b, Mol95a, Mol95d, Mol96a, Mol96b, Mol97, Mol98a, Mol98b, Mol99b, Mol99a, Mol00a, Mol00b, Mol01, Mol02a, Mol02b, Mol03b, Mol03a, Mol04a, Mol04b, Mol06b, Mol06a, Mol17]. client [Ano97a]. client/server [Ano97a]. Climate [TF02]. Clinical [RSW15]. clone [RGM22]. Closed [SL17]. closet [Ano97a]. Cluster [Alo06, BV95, BKGS02, IoG10, JGGF23]. Clustering [LLW^{+}23, BHC^{+}24, JGGF23]. Clusters [Che18, Res19]. CM [Edl04]. CM-Estimates [Edl04]. CMOS [ASP^{+}18]. CMregr [Edl04]. Co [Res19, XWZ^{+}23, ZLW^{+}19, XWZ^{+}22]. co-modeling [XWZ^{+}22]. Co-modelling [XWZ^{+}23, ZLW^{+}19]. Code [Ano02a, BC17a, CHMN13, Esm14, HYY^{+}15, LZ17, MGW99, WB12, AS12, BQOvdG05, BC22, Bra02b, CR13, CR20, GMT96, HS24, IZBT21, IP23, Mir96, PBI07, Qur05, YJ24]. Codes [BD20, ÖK17, BV19, Esp07, Kok07, Vit11]. Coding [Fis19, PS96a, PS96b, SWS97, TW06, Ano96j, HG24, Moo05, TW02]. coefficients [Cam06, CATK11, MSL09, PPS06, PF07].

Coerror [Gau16a]. coherent [KMBP24]. Cokriging [Mar91]. COLAMD [DGLN04a]. Colin [Ano99b, Ano99c]. Collaborative [OE95]. Collection [Hig89, Hig91, Thi95, YKS94, Bat19, HM22]. College [Lum92]. Collegiate [Lum91, Lum92, Ano96u]. Collocation [PR14]. Colloquium [Ano94c, Ano94l, Ano95b, Ano95h, Ano95t, Ano96f, Ano96m, Ano97l, Ano93c]. colour [WR04]. Columbus [H^{+}96]. column [DGLN04a, DGLN04b]. Combination [CGRvD15, Wit04]. Combinations [JMD08]. Combinatorial [ICS^{+}18, Jon91]. Combined [Jor94, SR94]. Combining [MR11, XWZ^{+}23, XWZ^{+}22]. combustion [MD95]. comes [Ano97a]. Coming [CW05e]. Comm [RC16]. command [Ano95l]. command-line [Ano95l]. Commercial [Zla17]. Communication [Cou95, Cou97, Gar96, Hay01, HYY^{+}15, RD22, ZL17, Ali02, Bla02b, Cou01, Elg01, Har04, JS04, PS98, PS00, PSB04, Tra04, TKD07, ZT02]. Communications [AC96, KK01a, Men95, AHS94, Bat99, FSC95, Gro05, HP02a, Kur00, MZ03, MC04, Nas01, VLV00, WOZ02]. communities [RGM22]. Compact [MP17, Cap13]. Companion [Ife05, Pfe95, App19, Coo01, Gre16]. Comparative [Ans14, BA94, FCP97]. Comparison [GMT96, PPD95, Rus93, Sch14, BGPPRW14, FR18, KS97]. compatible [EL16]. compendium [SG09]. compensation [STF01]. competition [Gra94]. Compilation [BCHS98, BGGT21, ÖK17, RBC20, PAG11]. Compiler [AP01, CA97, De 96, DGGM96, KSF94, LPD^{+}17, LPD^{+}20, RVV^{+}92, BSC^{+}00, BC17b, DJK93]. compilers [H^{+}96]. Compiling [JM94, KSF94, RHB96, AP02, KH14]. Complementarity [IKK^{+}19, FR96]. complete [Boy15, HMT13, Jon09].

Complex [Ano95m, BR96, KPS95, MBS15, Mol98b, Wun05, SSTD03, TYL^{+}16, VRVAC23, Wea97]. complicated [Wea97]. Components [CBhM12]. composer [Qur05]. composite [VK05]. CompPD [LZ15]. COMPRAIL [M^{+}94]. comprehensive [HL96, HL98, HL01, Hay99]. Compression [GJ20, JNR19, Thy10]. Comput [RC16]. Computation [AD14, Bee17, BKGS02, CF91, FBG94, Glo98, HL95, Hig02c, Hig15, Jef08, KT10, The92d, The92a, The92c, The92b, The93a, The93b, The93c, The97, Mol02b, dVSWAL17, Spe95, Spe96, Ano96s, BV21, Bal19, Bat19, BFG^{+}14, BDL25, EA03, Hig04, Kap04, KO00, Lea04, LM91, Mat95a, Mol00b, ZL13]. Computational [Ano96c, Cha17, HM88, IEE94f, Kin93, LC09, MM02, MM08, Mar11, MBS15, Ram97, Saa93, SMS96, SS10, Teo98, Ter94, Van92a, Vog02, War13, WR04, Whi04, Ano96t, BJG94, DJKP07, FD01, Gau06, IEE97a, JSM97, KS05, Man01, MF02, Moe04, Nah08, Nah11, TB95, Tra02, Tra10, VF10, KZL^{+}20, OB93, Ver09, Ano09, Iro15, Jai09, Kus02, Mun13, Wik04]. Computations [Bha05, CCF96, CV88, GV89, GCM25, LV16, Lys03, Nel17, VvBM08, VZDCM25, AH09, BK07a, Bha06, Eat92, Eat97, Eat00, Eat02, Eat05, EBH08, RBC20, RB04, Wat02]. Compute [LSvdVK09, LSvdV19, MV03, BO19, Boy15, Rob96]. Computer [ÅW97a, ATvBB17, BEH94, BM09, BDS97, BDS02, B^{+}94, CF91, Car92, CL96, Cor95, Cse99, Elg01, GG92, GR97, IEE96d, ICS96, Irw05, Khe96, Kin93, LHF96, LHF03, Mal96, MGC94, McC98, M^{+}94, NR96b, Ogu95, Per91, RL00, SB90, SYTD04, Tho94, TW98, UB95, ZL04, AM01, Ano93b, Ben98b, Ber03, Bry96, Cou93, Dav12, DIN06, Don95, KF96, Li99, LHZT23, Man01, Mit98, Mit06, MLB87c, SAE95b, YG99, Bar92, BLM97, BM94, FGMS21, Mro95b, Szy93].

Computer-Aided

[CF91, IEE96d, MGC94, GG92, Szy93]. Computer-Based [B^{+}94, McC98, Mit98]. Computer-controlled [ÅW97a]. Computeralgebra [Ben98b]. Computeralgebra-Systemen [Ben98b]. Computers [Fab95, IEE94a, Kin93, The92e, The92f, Ano94l, EL16, Kep09, Wie94, Kin93, M^{+}94, Ran96]. Computing [ACM97, ACM98, API^{+}19, Ano94h, Ano95h, Ano96e, Ano97b, Ano97k, Ano19, ASP^{+}18, BCHS98, Bra97b, CSY15a, CSY15b, CI01, Cse99, DB08, EP96a, EP96b, EP96c, EP04a, Esm14, FJSD96, GH93, GH95, GH97, Hea97, IEE98, IMM12, Law05, LZ15, MT97, Mol04a, MM96, Sha04, SM09, Sin93, Ste96, Tre15, ZA11, Zag16, Ano95f, BCR03, BSC^{+}00, BKG05, BSB23, CR13, CR20, DJKP07, EP00a, EP00b, EP03, GH04, Gra07, Hel04, H^{+}96, mH12, JSM97, Kat09, Kec01, KS97, LP05, LMAV24, LLZ18, LW03, MC02, Mol04c, Mol04d, Mol08, Möl07, Nas20, PWR13, PIAH12, PK08, QS03, QS06, Qua10, QSG14, TACA15, Tur00, Tur01, VC06b, Van97, Van00a, VF10, Whi07, Zen04a, Zen04b, ZJ01, dMMLOS20, Ano93c, BPS05, VC06a]. con [DACV95, GSS05, SCL95, de 05]. Concept [Shu17, Mol03a]. Concepts [Bha05, She04, AB98, Car00, JS04, MGT03, RR02]. Conceptual [New93, PFG08]. Concise [Bor18, Dav05]. condensates [Hoh14]. Conditions [AFOP19, Cam06, RC13, RC16]. conducteurs [CV96]. cones [Stu99]. Conference [ACM96, ACM97, ACM98, Ald96, Ano93e, Ano94d, Ano95i, Ano95k, Ano95n, Ano95s, Ano95o, Ano96g, Ano96v, Ano97h, BEH94, BF97, BCD^{+}20, Bjø01, BM09, BH95, Bub95, Bud95, Cou93, Dag94, D^{+}95, EP 97, Ham96, IEE93b, IEE94e, IEE94d, IEE94f, IEE95b, IEE96f, IEE96e, IEE97a, IEE97b, IEE97e, IEE04, ICS96, IGPF96, Kin93, KPN^{+}04, KZL^{+}20, Lum91, Lum92, Mal96, M^{+}94, Sil96, Sin93, USE99, W^{+}97, WD98, Ano93d, Ano93a, Ano94a, Ano94k, Ano94i, Ano96s, Ano96u, Ano97i, Ber95, Bry96, Gra93, Gra94, Ham93, IEE97c, I^{+}96, KF96, OB93, Ran96, Sas96, Van92b, BEH94, Mal96, Ter94]. configuration [CF16]. Conformal [AVV97, Nas20]. Congress [Ano94h, Ano96e, Ano97b, IEE94f, Ano95f, Ano95j, BH95, Coo95, GCP97, Lak97, SAE95b]. congruential [DH00]. Conic [PY22]. CONLAB [DJK93, JKR92, Tor02]. connected [Nas20]. Connecticut [IEE97e]. connecting [GGKM09]. Connectivity [Rul02]. connects [Ano95n]. Consensus [LLW^{+}23]. Consistent [HL03a]. Constrained [MON12, SZCP21, NSXZ14]. Constraints [IKK^{+}19, JR99, Mac02]. Construct [AFOP19]. Constructing [AFL^{+}12]. Containers [MS94, SMS96, Ano96c, SMS95]. Contaminated [SWG^{+}94]. Contemporary [AB03, PB01, PS98, PS00, PSB04, SK94, SK95, SK00, Wil96b, Ano93b]. CONTEST [TH09]. Continuation [DGKF12, Ajj95, GGKM09, GKD05]. Continuing [C^{+}96]. Continuous [BT04, EAK01, Ano93c, Chi97, OBCG19, WPK^{+}18, ZTF98]. Contract [WB16]. Contract-based [WB16]. Contributions [MBGV93]. Control [Ada97, Ano93c, Ano94d, Ano95e, Ano95h, Ano96v, Ano97h, ASG94, ARR02, AY97, Bab94, BCK96, Bis93, BP97, Bod98, Bol98, Bou95, Bro94, Bry94, Brz97, CST20, CGM95, CSV94, Che93, CL96, CFGG94, CKC94, DH95, DM90, DDD97, DAC95, DB95, DTB97, FPW90, FPEN94, FC95, Gaw96, GCP97, GGS01, Gra92, Gre94, GPK05, HK95, HBC94, HTCI96, IEE93a, IEE93b, IEE94e, IEE94d, IEE95b, IEE96e, IEE96d, IEE97e, IS96, JRCS95, JCRS96, JH96, JK93, KBKS98, Kir93, KW95, KH94, Kuo95, LL86, LL92, LL95, LY97, Lum02, Lys00a, Lys01, Man96, Mar95b, MGC94, Men95, MMPOQ^{+}25, MBMW95, Ned95, Nis95, Nis04, Oga94a, Oga94b, Oga95, Oga97, PSTO97, PR14, PN95, PH96, Pie96a, POVD96, PSP04, QCPG96, Rah93, RBZ96, RMS93, RLV11, Saa93, SDPM04, SSH94a, SSH94b]. Control [SS96b, SS96c, SCC95, SH93, SKG97, Shi92, Szy93, Tib93, TM97, TBHS94, Vac95, WM95, WM96, WJK02, Zin93, Abo03, Alt95, AHS94, Ano95b, Ano96k, Ano96p, Ano96r, Are94, AY91, Bab98, BC05, BJG94, Bar92, Ben98a, Ben04, Beq03, Bis97, BLM97, BM94, BJ02, Bry02, Bur99, Bur01, Cha02d, CF93b, CF93a, CSG98, Col04, CI01, DHS03, DCF95, DP95, Dat04, Del02, Dja98, Dja00, DB98, DB04, Eng05, EH07, FGCG94, FB95a, FB95b, FPW98, FPEN02, FC00, GL96, Gaw98, Gaw04, GB89, GV94, Gri94, Gup02, Hel04, Hen07, Hoh14, How91, IF94, Kho00, KS01, Kin93, KA02, Kul99, KC95, KG03, Led04, LNC98, LS95, LE00, Mac05, Mac02, Mac04, MD95, Mag02, MC02, Moh01, MO95, Nai03, NM99, Nør00, NMS^{+}06]. control [Oga02, Oga08, Oga10, OG95, Özb00, Pal98b, Pan89, PY98, Pas04, Pat94, PH00, PNGR00, Ram94, RBD^{+}10, RBD^{+}11, Rob96, RA92, Ros03, SFK91, SHX96, SH18, SS00, SEM04a, Shi98a, Shi98b, SP96, Spa03, Spo02, Ste02, SL03, Tew02, UE99, VWHK96, Wel93, WGP95, XCA07, Yan99, YL99, ZD98, Zhu01, ZJ01, ZW93b, ZW93a, Ano95d, CSV96, FC00, SA01, Cou93]. Controllable [TH09]. Controlled [ÇA10, Per91, ÅW97a, RL00]. Controller [AM17, AKB94, KW95, Per93, Ano96d, ASA96, GMT95, Hen07, MD95, PGBG94]. Controllers [KNNM97, TBH21, Ano96f, Bob05, Rob94]. controls [Rob02]. convection [PR02]. Convention [ACM98, Ano97h, BEH94, IEE94c, IEE96e, Ter94, Van92b]. Conventional [DH95]. Convergence [LS98]. conversion [SR94]. Converter [AM17, AM95]. converters [Jes01, OB93]. converting [RB04]. Convex [NN94, MP18]. Coombes [Ano00]. coordinate [YJ24]. Copenhagen [BS95a]. Coping [BV08]. Coprocessor [ASP^{+}18]. copula [Cob21]. Copulas [GHH20]. copy [FBH17]. Corn [SWG^{+}94]. Corner [Mol94, Mol95c, Mol95b, Mol95a, Mol95d, Mol96a, Mol96b, Mol97, Mol98a, Mol98b, Mol99b, Mol99a, Mol00a, Mol00b, Mol01, Mol02a, Mol02b, Mol03b, Mol03a, Mol04a, Mol04b, Mol06b, Mol06a, Mol17]. correct [Liu23]. Correction [SY20, Moo05]. correlated [SR09]. Correspondence [LSvdVK09, LSvdV19, TSM93]. Corrigendum [CK23, RBD^{+}11, RC16]. cortex [dMMLOS20]. Cosimulation [LSPM95]. Cosmology [Gre16, Myr17]. COSSAP [LSPM95]. Cost [Ans14]. Costas [TRD11]. COTS [GRDL^{+}12]. County [ACM98]. Coupled [MBS15, AM95, KMBP24]. Course [AC96, AY96, AY97, EO94, Her96, LQ94, MASV96, PS96a, PS96b, Rah93, SB90, Ano96j, BN01, Cha02d, FSC95, FR95, KHK05, LP05, MY98, Nas01, Por97, SR90, WJK02]. Courses [AY94, AS96, DM90, YA95a, YA95b, Bät20, DACV95, EA95, EM94]. courseware [Bha95]. Coventry [Ano94d, Ano94k]. cover [TF02]. Cözümü [Yuk96]. CPMC [NSXZ14]. CPMC-Lab [NSXZ14]. CPMD [MON12]. CPU [RBC20]. CPU/GPU [RBC20]. Cracow [Mro95b]. Craftsmanship [Mol02a]. CRC [Ano99a, Mar19, Mun12, Tay99]. Create [ES10, Ste08]. Creating [MA96a, MA97a, MA97b, NK18, Ste13]. Creator [Dut16, Hai08]. Criteria [Bro94, Esm14]. Cross [Lu96, ZLLT23]. cross-project [ZLLT23]. crossing [Mol97]. crunching [Nah11]. cryptocurrency [ZLMQ23]. Cryptography [TW06, TW02]. Crystal [YKS94]. crystals [KA13, RDP14]. cubature [CDSV10, CDSV11]. CUDA [ZZG^{+}14]. Culture [DG96]. cumbersome [Ano95l]. Current [PJ97, Boo04, PBB22]. Curriculum [DB93b, TM97, TBHS94]. Cursi [How15]. Cursos [DACV95]. CurveLP [Yan17]. Curves [FK11, Rov10a, Rov10b]. Custom [LPD^{+}20]. CVIK [JGGF23]. Cwiczenia [Brz97]. Cyber [BDM17, BDM18, XWZ^{+}23, ZLW^{+}19, XWZ^{+}22]. Cyber-Physical [BDM17, BDM18, ZLW^{+}19, XWZ^{+}23, XWZ^{+}22]. Cybernetics [IEE97a]. cycle [SR94, SCB99]. cycles [GKD05]. Czech [PJ97, SSV95].

D [Ano93b, BBF^{+}19, Mic00, Mun12, Myr17, Wie94, AV15, AEF^{+}14, BFSJP^{+}21, BPCCM23, BCH06, BBN^{+}22, Can22b, Jes01, MV22, RV13, Sha08b, Zek17, ZSW^{+}17]. D-BEM [AEF^{+}14]. D-STATCOMs [BPCCM23]. DAEs [SRK99]. DAESA [NPT15, PNT15]. DAFX [ZA02]. damping [Joh95]. Dan [RS15a]. d’apprentissage [CGL97]. Darren [Ano99b, Ano99c]. Data [Alo06, ÁBZ17, ÁBZ20, Alv11, Ano12, BB96, Bow10, Bur10, Car10, FY18, FPBO98, Han06, ICL97, JE94, Jer06, KK96, Koc14, KT10, Kus06, LPD^{+}20, Mar19, Mid00, MBMW95, Ni22, Rab20, RHB96, RG21, TBH21, WB12, YR19, YKS94, Ano98b, BLL^{+}15, BAEBAS19, BBF^{+}19, BBN^{+}22, Bre03, Cat95, DdAF^{+}20, HBT16, HL22, JA99, Jon09, JGGF23, Mar14, MM05, MMS11, MMS17, RS02, RHG09, Rid95, SR09, SZCP21, WPK^{+}18, Wet20, XZL24, ZSW^{+}17, dP96a]. database [BSS^{+}23, PGBG94]. databased [BJG94]. Databases [SM14]. dataflow [Cas14]. Daubechies [JNR19]. d’automatique [Riv01]. David [Ano96a, Ano97c]. DC [AM95, ÇA10, Gra93, Tob11]. DC-DC [AM95]. DDE [ELR02]. DDE-BIFTOOL [ELR02]. DDEs [ST01]. deadlock [JS24]. Dearborn [IEE96d]. Deblurring [NPP04]. Decades [Mol06b]. decaying [HS24]. December [Ano96g, IEE93b, IEE94d, IEE96e, IEE96b]. Decision [CFGG94, IEE93b, IEE94d, IEE96e, FGCG94]. decisions [Wil99, tHLMN19, LMN18]. DeCo [CGRvD15]. decomposition [KO00, LYH^{+}16]. decompositions [FHH99, LCL05]. Deconvolution [Rab20]. decreasing [MT84]. dedicated [dOSRS19]. Deep [Fay17]. defect [EH07, ZLLT23]. Defined [MBL^{+}97, JS04]. Deflection [Gab98]. degree [ADD04, DGLN04a, DGLN04b]. del [EGE95]. Delay [ADW23, SM94, BL02, ELR02]. delineation [PNL^{+}21]. Demographic [SM07]. demonstration [SW95]. demonstrations [Nah01]. demystified [McM07, Nas01]. Denmark [BS95a]. Density [BB96, CGRvD15, Rid95, FJ22, MT84]. Department [Lum91]. dependability [Cse92]. dependent [UW12]. Deployment [VFV13]. Deposition [RT96]. Depth [LZ15, CT97]. Derivation [mP97, MMD24]. Derivative [Ano05, SP22, Pri00]. Derivative-Free [SP22]. Derivatives [GJ03, PWR13, Sha07]. DERIVE [Ben98b, Ano95r]. derived [Pan89]. describes [Wea97]. Describing [DZ96]. Design [And95, Ano95e, Ano95r, Ano96d, Ano96r, AKB94, ASA96, Bis93, BP97, Bro94, Che93, CBhM12, DH95, EI05, Esm14, EM94, GMS92, Gre94, GPK05, HDB96, HK95, Her96, IEE96d, JHPK97, KSP11, Kin93, KH94, LL92, LL95, Mah00, Mah05, MGC94, M^{+}94, New93, Nor04, NMS^{+}06, Paa01, PFG08, PH94, Per93, PN95, POVD96, QCPG96, RBZ96, SSH94a, SSH94b, SH93, Shi92, Ste02, TH01, UB95, ZW93b, Zin93, ÅW97a, Ano94l, ARR02, AW97b, BH96, Bar92, Bat99, BPS02, BKG05, Beq03, Bis97, BLM97, BM94, Bur93, BEK99, Cat01, Che99, DHS03, DP95, Dat04, DDN02, Doe98, Dor00, Elg01, FBC00, FR95, GMT95, GL04, GGS01, Gri94, HA95, Hen07, Jan02, Jon95b, Jon95a, Kal97, KK01b, Kul99, LB00, Lum02, Luo95, LTE01, MD95, ME04, Mea02]. design [Nor05, PGBG94, PV99, PSP04, RL00, RB04, RB05, SMB04, Shi98a, Shi98b, Sil96, SA01, SP96, Tew02, TR04, TC13, Tob11, Váz16, WK99, WGP95, XCA07, Yan99, Szy93]. designed [Bro95]. Designers [Stu96]. Designing [Bab94, FKSM97, Mar95b, Oga94a, RSW15, Wea97, NK18]. Designs [ATvBB17, AFL^{+}12, BBB12, BV19, Kao09]. Desmond [IEE95b]. Detecting [Pac04]. Detection [NK17, Van01b, WTL00, BSB20, DP95]. detectors [PBI07]. Determination [Gab98, PPS06, YHC^{+}22]. Deterministic [BL02, Asi10, MKU22]. Detroit [SAE95b]. develop [Ano96h]. Developer [Ano97a]. developers [Led04]. Developing [LS05a, Wal18]. Development [Ano97f, AY91, Bro94, GMT96, KSP11, KNNM97, LW94, MASV96, MBB^{+}09, QCPG96, RA95, Tho94, TM97, Ano97a, JKR92, Pan89, PV99, Rib00]. Developments [GR97, SAE95a, Cse99]. Device [Pie96b, RR02]. Devices [CV96, Boo98, CC02, Dim00, Kar05, RR01, RR02]. Diagnostic [Sza04]. diagnostycznych [BF95]. diagrams [CDS09]. dice [Nah08]. Dichotomy [ST12]. DICOM [FM24, FM24]. Diego [IGPF96, Joh95, Kin93]. dielectrics [Vit11]. Diesel [HHF95]. diff [TACA15]. Difference [MP17, WW94, Gar07, LQT18, MSL09, RC13, RC16]. differences [AS12, WSST05]. Differential [ABB^{+}25, Ban01, Ben05, BC04, BD95, BD00, Bug95a, Bug95b, Coo00, Dav99, EP96a, EP96b, EP96c, EP00a, EP00b, EP03, EP04a, GP98, HLOR05, MR95, Mol03a, NPT15, Pol95, PBA06, PNT15, Riz17, Sch14, SBC04, BF09, BD04, BD05, CBL24, Col05, Coo98, Duf04, EP04b, ELR02, EMMK01, GS07, Goc02, GD99, GS12, HL03a, LS04, LC09, LQT18, Mol96b, Mol97, PIAH12, PT07, PA99, PA04, PSZ08, SG09, Sch12, Sta05b, Ano00]. Differential-Algebraic [NPT15, PNT15, HL03a]. Differentialgleichungen [BD95, BD00, Ben05]. Differentiation [BBH^{+}08, FHP^{+}12, HR91, HWB15, Nei10, PFG08, WR00, WR16, WR17, BEV06, CV00, FGMS21, For06, MG13, RH92]. Differenziale [MR95]. difficult [Mol03a]. DiffMan [EMMK01]. Diffractometer [YKS94]. Diffusion [FRAK15, CATK11, Gar07]. digest [IEE95a]. Digital [Amb95, And99, Ano95r, AC96, Bat99, BC06, Bob05, CGM95, Cav00, Cha05c, Che93, Cou97, Cou01, DDN02, FPW90, FPW98, GW02, GWE04, HBC94, IEE94b, IP97, IP00, IP10, Jac89, Jac96, JH96, JHPK97, Kha05, Kir93, Kum05, KG05, Lei02, Lei11, MMPOQ^{+}25, Mit98, Mit99, Mit06, MGT03, Nah08, NM99, PN95, SSH94a, SSH94b, Sch94, SC97, Ste03, SH11, TS14, Vac95, Amb99, Ano95t, Ano97c, AI21, Bar16b, DM06, EA04, ElA04, Gop10, Hay96a, KL01, Kur00, Lai04, LA01, Luo95, McA04, Nas01, Pac00, Por97, Rob94, SH05, Sch11, SB10, WWM06, ZSW^{+}17, ZA02, Sch94, Wer03, von93]. Digitale [Wer03, von93]. dimension [dMMLOS20]. Dimensional [HPK18, Koc14, Lam95, NGKM18, Rid95, AI21, BV21, Cap13, Kok07, Kok15, MP18, PT24, Rob96, Ano96m]. Dimensions [FGjS15, GBM15, BC22]. dinamico [EGE95]. diode [dP96a]. Direct [KBKS98, DH97a, SZCP21, SP22]. DIRECT-Type [SP22]. DIRECTGO [SP22]. Direction [AS15]. directional [GO97]. directions [Men01, Bro94]. Dirichlet [RC16, RC13]. Discontinuous [ABB^{+}25, FRAK15, JRA^{+}18, OBCG19, RAW^{+}16, RHR^{+}21]. Discovery [MON12]. Discrete [EA04, ElA04, Han92, JNR19, KDAB19, Oga95, PSR16, Qua02, SI00, Sch94, The92h, Chi97, Han94, JL01, Jon95b, Jon95a, OG95, SL17, TSCC05, ZTF98, BC17a]. Discrete-Time [Oga95, Qua02, TSCC05]. discriminant [Van04]. diseases [TYL^{+}16]. diskette [Ano95c]. diskreter [Sch94]. disks [Ano93b, TACA15]. Disney [IEE94d, IEE94f]. DISODE45 [CMR17]. displacement [UW12]. Dispositifs [CV96]. Dissimilar [Has12]. distance [MP18]. Distinction [VA94]. Distributed [BCHS98, CWP98, Cze95, IEE98, Nat94, BSC^{+}00, BKG05, CCM^{+}03, DJK93, JKR92, Mol01]. Distribution [BPCCM23, ICS^{+}18, ÖK17, SBL^{+}10, SU93, TBH21, Wit04]. Distributions [BO19, dVSWAL17, CS12, LHW01, MMD24]. divider [FD01]. Dividing [Gou20]. Division [Ano93c, Ano95h, Ves94]. DM [Ano99a, Ano99b, Ano99c]. DNSLab [VK16]. Do [Kah04, MM96, WW99]. Document [AI21]. documents [AI21, TP03]. does [Ano97a]. Doing [Van00b]. Dolphin [IEE94f]. Domain [DL95, ICL97, KAB97, KPS95, Mac91, USE99, AM17, DL01]. Domain-Specific [USE99]. domains [Nas20]. Doppler [STF01, Ali02]. Dose [WTL00]. Dose-Response [WTL00]. double [JSB20, SM95]. Doubletree [Ano97h, Ter94, WD98]. Doubly [IKK^{+}19, Esp07]. downdating [LLZ18]. Dragonfly [AMR18]. Dresden [Ano96s]. Drive [DRR97, SFPO94]. Driven [CMKH03, PK08]. Drives [ÇA10, IEE97b, LJC93, LJR93, BN99, Bos01, Bos02, ES00, Moh01, OB93, RLH03]. driving [SAE95b]. drug [LYH^{+}16]. drug-pathway [LYH^{+}16]. drugs [HG24, HGM^{+}25]. DSDP5 [BY08]. DSDP5-software [BY08]. DSK [Cha02c, Cha05c, WWM06]. DSL [USE99]. DSP [Ano95t, AB98, Ano96j, BPS02, Bro95, Bur94, CGM95, Cha02c, Ife05, KK01b, Mat91, MSY98, PS96a, PS96b, Qur05, RLV11, SW95]. DSP-Based [CGM95]. DST [RD22]. Dual [DZ96]. Dual-Input [DZ96]. Dubious [MV03]. Dublin [Ken95]. ductile [GO97]. Duelling [Nah00, Nah02]. Duffy [Mun12]. dummies [SM15]. DVD [RS15a]. DVD-ROM [RS15a]. DynABlock 2D [Por23]. Dynamic [Ano05, Ben95b, CF93c, Fis95, FSO93, FPW90, FPEN94, HBC94, Hun98, Lam95, LG94, Mar95b, MNBB19, Ong98, Rab20, RA95, RD22, SKG97, WB12, WL97, CFN01, CFN02, FPW98, FPEN02, Gaj03, Hol04, Kle07, KA11, LH14, Lum02, MKU22, The98, Nør00, Pal98b, Pri00, SFPO94, SZCP21, dP96a]. Dynamical [Lyn04, RBZ96, Sch96, CG00, MC02, Wil99]. Dynamics [AD15, Cun03, GBM15, Gaw98, Gaw04, HR96, Hof98, JK93, Oga98, SLI99a, Yan05, Ajj95, BKR97, Beq98, Bro01, Chu00, CSG98, Dat13, Doe98, EA03, Gra11, Ken95, Moo98, Oga04, Pal05a, Pet96, Por23, Rob02, Sch98, SL17, SH18, SKA19, SEM04a, UE99, WK99, WCS92]. dynamischer [Hof98].

e-book [Mar19]. E.E. [Ano95v]. early [TR04]. Earth [Cha17, Mid00, TGM06, TG10, RG21]. easier [Fos01]. easily [MT84]. Easy [Rul02]. ECG [Ozk23, PNL^{+}21, DS16]. ecg-kit [DS16]. ECGdeli [PNL^{+}21]. Ecole [Pav93]. ecological [Rou98]. ecology [Ber03]. Econometric [CNJ97]. Econometrics [Edd00]. Economic [JLM96, CI01]. Economics [CLMM20, And05, Bra06, MF02]. Ecopath [Kea17]. ecopath matlab [Kea17]. ed [Ni22]. Edge [AV15, McN05]. Edinburgh [CF93a]. Edition [Alv11, Ano12, Atk05, DDW93, HL95, The92g, The92e, The92f, The97, The96, Sch97, Veh07, Ano96a, Ano99a, Ano09, DIN06, Eat00, HL97, The98, Ver09, Mar14]. editions [Mol04a]. Editor [FBG94, Web97]. edn [Mun12]. Eduardo [How15]. educating [Gra94]. Education [Ald96, Ano95k, Ano97i, Bou97, Bud95, C^{+}96, EP 97, IGPF96, KA09, Pie96a, PMA94, TBHS94, EM94, Gra93, Gra94, IF94, I^{+}96, Kin93, Mro95b, OE95, Zal96]. Educational [TQ96, Ano96q]. EE [Nas01]. Effective [Cze95, LNLB19]. Effectiveness [DP96b]. effects [ZA02]. Efficiency [Zag16]. Efficient [AM22, BK07a, BO19, CGRvD15, DP16, Joh18, LYH^{+}16, For06, JZW^{+}22, MBR21, PWR13]. Efficiently [FBH17, DDK14]. effort [LNLB19]. Eigenfunction [Mol03b]. eigenfunctions [BV21]. eigenmodes [SN01]. eigenpairs [BCR03]. Eigenstructure [Whi91]. Eigenvalue [Fab97, Jon01, HMT13, MY05, Rad96]. eigenvalues [BV21, Mol98a]. eigenvector [Mol02b]. eighth [DIN06, NP96]. EIGIFP [MY05]. eigshow [Mol98a]. Einführung [Beu05, BD95, BD00, Hof98, Mei05]. Einstein [Hoh14]. Einstieg [Bät20]. elastic [Car99, RCT20]. Elasticity [ACFK02, BCG17, Kok07, SBL19]. ElasticMatrix [RCT20]. Elastoviscoplastic [CK02]. Electric [Ano95i, AS96, BN99, Cat01, Cha02a, Cha05a, Con95a, FKU03, IEE97b, KSP11, LL03, NR96a, NR96b, BDR04, BEK99, Car00, DS04, ES00, Lys00b, Moh01, Ong98, Pau01]. Electrical [AY94, AY97, BEH94, DRR97, Mal96, SB90, YA95a, YA95b, Ano93d, EA95, Li99, Mag05, MAB^{+}11, Man01, Pha99, Riz00, Sil96, Vin98, YG99, Sim99]. Electroheat [SS96a]. electrohydraulic [SM95]. Electromagnetic [MA96a, MA97a, MA97b, TKD00]. Electromagnetics [LS05c, Ter94, Ano96t, LS05b, VCK98, Wen05]. Electromechanical [FKSM97, Lys00b, Lys05, OB93]. Electronic [BRPCR94, Bla02b, Kar05, Sch99, Boy99, Ken95, Mol04a]. Electronics [Att96, Att99, Att04, FSO93, IEE94a, SO93, Att02, Att10, Bos01, Bos02, Jac04, Nah01]. Electrosoft [Sil96]. Electrostatic [VPM16, WL94]. Elektriske [Sim99]. Element [ACFK02, Bha05, GBM15, KBQ97, Poz05, Wal18, ACF99, Bha06, CK02, DS09, DRS18, FS23, HCBAEC23, Kin93, Kok15, KB97, KB00, LQT18, Li20, OBCG19, Poz14, Sut17, Tho04, VCK98, YJ24]. Elementary [AH04, Atk05, BD04, BD05, EP04b, Gro94, Hil96, LE96, Lev92, Man01, Ono01]. Elementi [Bol94]. Elements [ANM01, Bol94, Fog99, FMS25, Gup02, Hor98, Kat03, Whi07, Zie97, AV15, Car99, Epp11, Joh11, MV22, RV13]. Elfun18 [Bat19]. Elimination [JMD08]. elliptic [Bat19, Boy15, IP23, Mil20]. Ellis [Ano96b]. EMAG [WL94]. Embedded [BGGT21, Led04, NGKM18, Qur05, BC17b, CCM^{+}03, Jan02, Led01]. Embedding [Mat91, MCd^{+}96]. Emergency [HHF95]. emgr [Him23]. emphasis [Den98]. EMpirical [Him23]. Employing [Fis19]. Employment [RT96]. EMTP [LW94, MBL^{+}97]. emulate [Rob94]. enabled [EPJ^{+}05]. Encoding [GJ20]. Encounters [Gul12]. energies [MV22]. Energy [AM22, CLTS20, OB93]. Energy- [AM22]. EnergyPLAN [CLTS20]. Engine [WM95, WM96, BC12, MD95]. Engineer [Kin95, Mag00, Mag05, MAB^{+}11]. Engineering [Ald96, Ano95k, ASG94, AY94, AY97, BEH94, Bod98, Bro01, CBCC96, DIN06, DTB97, DB93b, Ett93, Ett97, HKD02, IEE96a, Inm94, Inm01, Irw05, IGPF96, Kiu05, Lev92, Lev96, LY97, Lys03, Mal96, Mat92a, MBS15, MKP02, MM00, MM96, Ned95, Nis95, Oga94b, Oga97, Pao99, Pao01, Pie05b, PMA94, SB90, Sch05, SKG97, SLI99a, SLI99b, UCL04, Web97, YA95a, YA95b, Yuk96, vdH05, And99, Ano93a, Ano94k, AC99, B10, Bol98, Bry96, Bud95, Bur01, Car00, Cur05, CS99, DACV95, Dag94, DM96, Duf03, Duf10, Duf11, EA03, EBB05, EA95, EM94, Fog99, FR95, GL96, HK01a, HDR97, HDR00, HT97, IEE97d, Irw02, I^{+}96, Jef04, Kin93, Kiu10a, Kiu10b, Led01, Lys00a, Lys01, Mag05, MAB^{+}11, MM97, MM98, Moi10, Mun12, Nis04, Oga02, PPL^{+}18, Pha99]. engineering [Phi00, PV99, Ran96, Rao11, Rib00, Riz00, SSTD03, Sas96, Sch12, SMB04, Sil96, Sta03b, TU97, VF10, Wen05, WJK02, Zie97, Ben98b, Tay99, UCL04]. Engineers [BB95a, BB95b, BB99, Ett96, GP96, GG04, GB03, HV07, Pra96, Rao02, SW94, Beu05, BB02, Bor97, Cha00, Cha02b, Cha04, Cha05b, CC06, CM99a, EU07, Esf03, GS08, Gra94, Hah97, Hah02, Har05, Jaf00, Jef05, Kin98, Kin01, Kin06, Lai04, Li99, Man01, MM10, Moo07, Moo08, Moo09, Moo11, Moo15a, Moo15b, Mun13, Nas01, Nir02, Oga08, Oga10, Pal98a, Pal01, Pal05b, Pra99, Pra02, Pra06, SH00, Vin98, WB99, Wil03, YG99, Mic00]. England [Cou93]. English [Hof98]. Enhanced [TC97a, WL94, DH97a, TS21]. Enhancement [MLAS21]. enhancements [Wel93]. Ensembles [IoG10]. Ensight4Matlab [SBL^{+}17]. EnSight(R) [SBL^{+}17]. entitled [Wie94]. Entwicklungslinien [Bro94]. Entwurf [Bro94]. Envelope [CSY15a, CSY15b]. Envelopment [ÁBZ20]. Environment [Ano05, AW92, BCHS98, BDM17, BDM18, CB98, CNJ97, FKSM97, HTCI96, MM96, Rav94b, SDPM04, SCC95, Ano95l, Ano96p, BJG94, BKG05, Cat95, CB99, Chu00, EPJ^{+}05, GI06, MD95, MDB01, OG95, PGBG94, Pan89, Pri00, SFK91, Sch04, Zha98, Ano96r]. Environmental [CBCC96, SU94, Nir02]. Environmentally [Thi95]. Environments [BRS94, FBG94, Mor98, Che04, DF99]. envlp [CSY15a, CSY15b]. Epidemiological [Alo06]. Equation [ADW23, Mol03b, Asi10, Cap13, CBL24, HL03a, LS04, MBR21, Mai00, Mei05, Mol96b, OAKS11, PSZ08, SG09, Sch12, Tri08, YJ24]. equation-free [MBR21]. Equations [AS08, Ben05, Bir18, BD95, BD00, Bug95a, Bug95b, Can22b, EP96a, EP96b, EP96c, EP04a, Kel95, Mol03a, NPT15, Pol95, PBA06, PNT15, RS08, Sch14, SF14, Ano00, AD14, AD15, BV21, Ban01, BF09, BNN16, BC04, BD04, BD05, CR13, CR20, Col05, Coo00, Coo98, Dav99, Duf04, EP00a, EP00b, EP03, EP04b, ELR02, EMMK01, Gar07, GS07, Goc02, GP98, GD99, GS12, HLOR05, JZW^{+}22, Jon01, Kel03, LVV05, LC09, LQT18, MSL09, Mol97, mP97, PIAH12, PT07, PA99, PA04, SR09, Sta05b, SBC04, YMLW09, ZZC^{+}08, ZZG^{+}14, ZD03]. equidistribution [SSW09]. equilibria [Do98, Duo98]. equilibrium [BFG^{+}14]. Equivalent [ICL97]. equivalents [DPE96]. ER-fMRI [Kao09]. era [EP 97, Mol02a]. Erice [DG96]. Error [HM19, Moo05, BC05, Cam06, Hsu99, Mol95a, RB05]. errors [Liu23]. Esercitazioni [Bol94, CF91]. Esercizi [FT92]. ESM’96 [JLM96]. Essential [Hah97, Hah02, HV07, RS15b, Bar16a, Mic00]. Essentials [Jef04, Kin95, ZD98]. Estimates [Edl04]. Estimating [Bar97, GPR14, RS08, SM07, BSS^{+}23, SR09]. Estimation [AS15, BSL93, BSLK01, BB96, ICS^{+}18, Men95, SBL^{+}10, She08a, She10, SW94, vdH05, ATB05, CBL24, FJ22, Gre96, Kho00, KK93, MIK00, QH01, Rao11, SN01, Spa03, Spo02, Van01b]. Estimators [CSY15a, CSY15b]. ETTH [Ans14]. étudiants [MM10]. EUFIT [Ano94h, Ano95f, Ano96e, Ano97b]. EURISCON’94 [Ano94a]. Europe [Ano95i, Ano95n]. European [Ano94h, Ano94a, Ano95f, Ano96e, Ano97b, Pav93, SSV95]. EUROSIM [BH95, BH95]. Evaluation [BRS94, Gau16a, MJ01, CZ17, HL22, LGML05, MV22, OG95, PF07].

Event [RSW15, Mol97, PK08]. event-driven [PK08]. Events [Tay95]. everything [Ano97a]. evolution [PT07]. exact [Jon09, ZD03]. Exactly [LCMCD22]. Example [KPS95, Bät20, RHR^{+}21]. example-oriented [Bät20]. Examples [GG04, GPK05, PMA94, von93, Bry02, FCP97, Gri95, KA02, Möl07, Qur05, SB10, Ste03, SH11]. Excel [DIN06, Irw05, Sti04]. exchange [dP96a]. exclusion [MMD24]. Executable [KSF94]. execution [PAG11]. Exercises [Bol94, B^{+}94, CF91, Jac96, LHF96, LHF03, Tow16, von93, BH97, Don95, Gau16b, Kin97, McC98, Riv01, Brz97, FT92]. Exeter [Ano96v]. Experience [GRDL^{+}12, SZCP21]. experienced [HLR01, HLR06a, HLR14]. experiences [SR90]. Experiment [Jor94, DPR05]. Experimental [Kao09, MJ01, GMT96]. experimenting [BJG94]. Experiments [FS17, Gek08, HM88, Hil91, MN03, Mol11, Nel17, SB90, GL96, KL01, NK18]. EXPERSYS [CAC94]. EXPERSYS-94 [CAC94]. Expert [Kin93, Pan89, CAC94, tHLMN19, LMN18, TSA21]. EXPINT [BSW07]. explain [Mol98a]. Explained [GB03, LL03]. explicit [BLL^{+}15, Faz10]. Exploiting [LPD^{+}20, RHB96]. exploration [AI21]. Explorations [Bug95b, HC95, BDS97, BDS02, CGI99]. Exploratory [MM05, MMS11, MMS17, Mar14, Alv11, Ano12, Han06, Jer06, Kus06, Mar19, Ni22]. Exploring [LW03, TYL^{+}16]. Exponential [MV03, BSW07, Gau15]. exponentially [HS24]. Exponents [WK95]. Exposition [Kin93]. extend [CLTS20]. Extended [DMB17, OF92]. extensible [HM22]. Extension [BT04, Bro95]. External [The92a, The92b, BEV06, DPE96]. Extraction [CCF02, BO19, HKF^{+}20, NA02]. Extrapolation [Kah02a]. Eye [Jon18, Hun98]. Eye-Tracker [Jon18]. Eyex [Jon18]. Ezulwini [Van92b].

F [Sha04]. Fabrics [ZL17]. Face [KDAB19]. facility [Pan89, SW95]. Factor [RJ93]. factored [BK07a]. Factorization [LCMCD22, Mat92c, Mat94b, BSB23, Dav05]. FACTORIZE [Dav13]. factors [Phi00]. FACTS [Ach04]. Faddeyeva [Zag16, ZA11]. fading [Pät02]. Failure [Kah04, Faz10]. FALCON [DGGM96]. fall [Mar68]. Fame [Mol94]. family [DH97a]. FAQ [Bee05]. Far [Spe95, Spe96]. Far-Field [Spe95, Spe96]. Faraday [CHT15]. Fast [AV15, BCG17, CZ17, Dak05, DP16, EL16, GLJ^{+}91, MS14, MV22, RV13, dVSWAL17, Rum01, Ves98, BK06, BV21, BN10, BN11, CDSV10, CDSV11, Dat13, LHW01, MT84, PC13, Van92a]. Faster [HTNFBS06a, HTNFBS06b, Mol00a, Led01]. FastGAPP [RG21]. Fault [KA09, NE05, DP95, LNLB19]. FCRC [ACM96]. FDEM [DdAF^{+}20]. FDEMtools [DdAF^{+}20]. Fe [USE94]. Feature [NA02]. Features [JHPK97, Kwo93, The93b, KMLP^{+}23, Mol99b]. February [Ano95t, D^{+}95, SAE95b, W^{+}97, WD98]. Federation [GCP97]. Feedback [FPEN94, FPEN02, FC95, FC00, Hel04, PH96, PH00, AW97b, CWP98, Dja98, Dja00, Dor00, LE00, Özb00, PSP04, SP96, Ste02, XCA07, Yan99]. Feedbak [Bou95]. Fejér [DMB17]. FELICITY [Wal18]. FEM [AV15, BCH06, IP23, MPVA25, MV22, RV13]. fermentation [TJML92]. Fermi [DPR05]. ferrofluids [JAC20]. Fess [WD98]. FESTUNG [FRAK15, JRA^{+}18, RAW^{+}16, RHR^{+}21]. fetal [BlD^{+}20]. feval [LH14]. Few [Kwo93, BCR03]. FeynDyn [Dat13]. Feynman [Dat13]. FFTW [Mol00a]. fiber [BN10, BN11]. fibrous [YHC^{+}22]. fiction [Nah11]. FIE’96 [I^{+}96]. Field [OBD23, PH94, Spe95, Spe96, SZM^{+}14]. Field-Oriented [PH94]. Fields [Sim99, JAC20]. Fifth [Lum92]. file [TSA21]. file/Mif [CF16]. Files [KSF94, CF16, Liu23, SBL^{+}17]. Filippov [CMR17, PK08]. Filter [Ano95r, JHPK97, LTE01, SN96, AM95, Bur93, Hay96b, Hay02, Luo95, Mea02, Sim99, WS04]. Filtering [JH96, BH97, Din02, GA01, MIK00, PR06, Say03, SD91, ZM00]. Filters [Jac89, Jac96, DM06, FB98, Paa01, Su96, cS02]. filtry [VS00]. finance [And05, B10, Bra02a, Bra06, HLS08, McN05, MF02, PF10]. Financial [Gru04, GJ03, Hig04]. Finanzderivate [GJ03]. Finanzmathematik [Gru04]. Finding [OF92, Edl04]. Fine [SFK91]. Fine-tuned [SFK91]. Finite [ACFK02, Bar97, Bha05, CCF96, Gar07, GBM15, Kat03, Kin93, KBQ97, MP17, Mol00a, Poz05, VCK98, Wal18, WW94, ACF99, AS12, Bha06, Car99, CK02, DRS18, FS23, Kok15, KB97, KB00, LQT18, Li20, OBCG19, PNGR00, Poz14, RC13, RC16, Tho04, WSST05]. Finite-difference [Gar07, RC13]. FIR [Luo95]. First [Ano97c, Bir18, Rah93, SB90, BN01, Boy15, Cha02d, KHK05, LP05, MSY03, MSY98, MY98]. First-Order [Bir18]. Fishery [Zla17]. Fit [ÖK17]. Fitness [HBT16]. Fitting [McI16, CS12]. Five [MV03]. Fixed [JRCS95, JCRS96, Nyh08, Qur01, Ano96k, Rob94, RB04]. Fixed-Order [JCRS96, Ano96k]. Fixed-Point [Qur01, Rob94, RB04]. FL [Ano99a, Mar19, Ano96u, Ano97i, Ran96]. FLAME [BQOvdG05]. Flap [Ste09]. Flexible [Gaw96, JK93, Rob02, Wei12, Ano96d, ASA96, Ano94b]. Flight [HR96, How91, LB99, Sch98, Bro94, Gre96]. Floating [Gou20, HP19, Mik23, Mol17, Ste09, Meu20, Mol98b, Rob94, RB04, RLV11, Mol96a]. Floating-Point [Gou20, HP19, Mik23, Ste09, Mol98b, Rob94, RB04, RLV11]. FLOATP toolbox [Meu20]. Flood [RH00]. Florida [ACM98, IEE94d, IEE97a, IEE94f, W^{+}97]. Flow [MLAS21, TQ96, AGL07, Ano96q, BLL^{+}15, ERS07, FD01, HKF^{+}20, VK16]. Flugregelung [Bro94]. Flugregelungssystemen [Bro94]. Flugzeugmodell [Bro94]. Fluid [LC98, WB99, Kin93, PR02]. fluids [PWH02, Sha03]. fMRI [Kao09]. FMVol [ST24]. Focus [CH20]. food [BSS^{+}23, Kea17]. force [BFSJP^{+}21]. forced [MSL09]. forecasting [CBL24]. form [mP97]. Formal [BBB12, GRH^{+}21, CDS09]. Format [KT14, SBL^{+}17]. formatting [Liu23]. forms [Bra02b]. Formulas [Yan05]. formulation [GMT96, JRA^{+}18]. formulations [BSS^{+}23]. FORTRAN [Tay99, Ano97j, Bra97a, CFGG94, FGCG94, LHW01, Pao99, Pao01, RVV^{+}92, Ano97k, Bra97b, DP96b, DP99, LP05, LS04, LS05a, Rei93, Zag16]. fortune [Mol94]. Forum [Ano94j]. forward [For06]. Foundation [Bro94, Hay99]. Foundations [GG04, Bar16b, Wil99]. Four [Ano97c, Boy15, GMT95, Zen97]. Fourier [BN01, Van92a, Ano98c, CZ17, CM67, Mol00a, ST24, Sou94, Stu96, Ves98, Wil96b]. Fourth [JE91, Lum91, M^{+}94, Ano96e, Gra94]. FPGA [HNS^{+}01, RLH03, RB04, RB05, TC13]. FPGA-based [RLH03]. fractal [MC02, dMMLOS20]. Fractals [Gul12]. fragments [TSA21, TSA21]. Fragments-Expert [TSA21]. frame [JSB20]. Framework [BW20, DH12a, GV25, Him23, MMPOQ^{+}25, Can24, CDS09, CDOBA22, SR09, ZC08, CF16]. Frameworks [Van92a]. France [AO95, BM09, Ber95, CAC94]. Francisco [GCP97]. Franco [AO95]. Franco-Belgian [AO95]. Fredholm [AS08]. Free [Ano97k, Bra97b, ES10, SP22, MBR21, Ste08, Ste13]. French [Abd94, Bou95, CGL97, CV96, MM10, Riv01].

Frequency

[BO19, BAEBAS19, Ega00, ICL97, KPS95, Mac91, PSR16, Rah93, JAC20, QH01, RH00]. Freshman [FR95]. friendly [Ozk23, TS14, VZDCM25, YG99]. Frontiers [I^{+}96, Gra93, Bud95, EP 97, Gra94]. FTIR [SWG^{+}94]. FTTH [Ans14]. fuel [PSP04]. Full [SS18]. fun [Ano97a]. Function [BD20, Bee17, Che93, DZ96, Gau16a, Mol02a, OBD23, Sar17, TCD^{+}22, Cur95, HBT16, LGML05, PT24, Zek17]. Functional [HSM04, RHG09, YR19, RS02, Bow10, Bur10, Car10]. Functions [Ano95q, BT04, BKGS02, DMB17, GG92, HC92, MAC08, Rid95, WR16, WR17, Zag16, API^{+}19, Bat19, HS24, KS94, MBR21, MT84, MS14, Mol98b, Mol02a, PWR13, RKZ^{+}14, Sha11, SR09, Tre15, VC06a, VC06b, WPK^{+}18, ZA11, Sha12]. Fundamental [Bha05]. Fundamentals [Bay99, Ben95a, BL96b, BL96a, ES00, GC99, KH97a, KH00, Kay93, Kel00, LS05b, Mei01, Moi10, PS96a, PS96b, Pau00, Pau01, Pie96b, SSTD03, Say03, SH05, Sch11, SB10, Tra02, Tra10, Wat02, Wen05, ZM00, von93, Ano96j, Boo98, Cha02a, Cha05a, DG04, JDFV08, Kha05, Lys05, Moe04]. future [Ano95k, Bry96]. Fuzzy [Abo03, Bab94, BV95, Bab98, BD96, ÇA10, HT97, IEE94f, JE94, KW95, KNNM97, PY98, PK04, TU97, YL99, JSM97, Kec01, LL96, Men01, Spo02, WE96, Ano95g, Ano96f]. Fuzzy-Toolbox [Ano95g].

G [Bur10, Car10, DG96, GR97]. GA [Ano94e]. GABLE [MDB01]. GAIL [TCD^{+}22]. Gain [Kah98]. gaining [McN05]. Galerkin [FRAK15, JRA^{+}18, OBCG19, PPS06, PF07, RAW^{+}16, RHR^{+}21]. gallery [Bal19]. Galveston [Pat94]. games [CK22, CK23]. gamma [SAKG15]. gamma-rays [SAKG15]. GAMS [BPCCM23, DF99]. Gander [Ste96]. Gap [QMS98, PPL^{+}18]. Garrett [Ano00].

Gateway [RVV^{+}92, VK16]. Gauss [RBD^{+}11, KS97, Rus93, Gau16a, Joh18, RBD^{+}10, Rov90]. Gaussian [Gri95, LMAV24, PR14]. Gautschi [Tow16]. Gazebo [MMPOQ^{+}25]. GBE [TBH21]. GBP [Ano99a, Ano99b, Ano99c]. GBT [GV94, VWHK96]. geared [BEV06]. GEEQBOX [RS08]. gems [mH12]. gen [Ano95n]. Genaue [ZD03]. General [DCF95, She10]. Generalized [CKC94, Dun99, Fab97, McI16, RS08, MY05, SR09, Ves94]. Generate [MGW99]. generates [Mol01]. Generating [Gou20, KW95, Cam06, MT00, VRVAC23, TP03]. Generation [CMKH03, HYY^{+}15, MAC08, MNBB19, SYW^{+}24, WB12, Can22b, CF16, HBT16, Lan00, RBC20]. GEnerator [CF16, HHF95, PS04, JR99, Kok15]. Generators [CM67, DH97a, DH97b, DH00]. Genetic [CF95, CFPF94, HH04, TYL^{+}16]. genome [TYL^{+}16]. genome-wide [TYL^{+}16]. Gent [BLM97]. Geochemical [RG21]. geodesy [SB97]. geographic [Ano98b]. Geometric [VZDCM25, GI06, HH03, MDB01]. Geometry [EP94b, HZ96a, JAC20]. GeoPDEs [Váz16]. George [Ano96b]. Georgia [Bud95, IEE96f]. geosciences [Zek17]. German [Ben98b, BB95b, Bod98, KK01a, Mei05, Pie05b, Bät20, Ben05, Beu05, BD95, BD00, Gra04, Gru04, GG04, GJ03, Löw01, Sch05, Wer03, ZD03]. Germany [Ano94h, Ano96e, Ano97b, BCD^{+}20, C^{+}97, Ano95f, Ano96s, Glo98, Syd95]. Gersgorin [Var04]. Getting [Pra96, Pra99, Pra02, Pra06]. Gewöhnliche [BD95, BD00]. GHM [GSM95]. GIS [And05]. Givens [GLJ^{+}91]. Glasgow [IEE94e]. GLce [SZCP21]. Gleichungssysteme [Mei05, ZD03]. Glimpse [Mal96]. Gliwice [MO95]. Global [GWA01, SP22, MC97, MA95, SZCP21, HL03b]. Globally [Esp07]. GloptiPoly [HL03b]. GmbH [Res19]. GniCodes [HH03]. GNU [FRAK15, Eat97, Eat00, Eat02, Eat05, EBH08, EaoGOBHW14, FY18, JRA^{+}18, PSR16, Rab20, RAW^{+}16, RHR^{+}21]. Goes [GLJ^{+}91]. Gold [MG18, SBL^{+}17]. Golden [Mol96b]. Goliath [Ano97c]. Golla [GSM95]. Golla-Hughes-McTavish [GSM95]. Golub [MS14]. Good [Mol98a]. Goodness [ÖK17]. Google [Mol02b]. Gournay [CAC94]. Gournay-sur-Marne [CAC94]. government [Mac00]. Governor [HHF95]. GPELab [AD14, AD15]. GPGPU [PIAH12]. GPOPS [RBD^{+}11, PR14, RBD^{+}10]. GPOPS-II [PR14]. GPRlab [XZL24]. GPS [SB97]. GPU [Bro07, Cap13, mH12, KA13, RBC20, ZZG^{+}14]. GPU-accelerated [Cap13]. GPUs [API^{+}19, Dat13, DEQOR13]. Gradient [ZZC^{+}08, Jon09]. Gradient-based [ZZC^{+}08]. Grado [DACV95]. Graduate [AY94, AY96, AY97, YA95a, YA95b, DACV95, EA95]. GRamian [Him23]. Gramm [Mor18]. grammar [Mor18]. Graph [GR97, MVDV97, RD22, Bro01]. graph-centered [Bro01]. Graphic [Nak96, Spe95, Spe96, Nak02]. Graphical [BB96, Coo96, Cra96, Dun99, FBG94, LSvdV19, The93a, Mat94a, XWZ^{+}23, ZLW^{+}19, AH05, Ban01, CF16, FM24, LP05, TSA21, WA96, XWZ^{+}22]. Graphics [Mar95a, Mar99, MH03, Ogu95, Kin93, Mor18, Ano99a]. Graphs [BL96b, BL96a, GCM25, Ano97a, mP97]. grating [HA95]. grating-based [HA95]. Graves [Car10, Bur10]. gravitational [WPK^{+}18, Wet20]. gravitational-wave [WPK^{+}18, Wet20]. Green [FS17, Nel17, Myr17, RS15a]. Greenbelt [Gre96]. Grenoble [BM09]. grid [KW05, SSW09, Som07, EPJ^{+}05, SF14]. Grid-enabled [EPJ^{+}05]. Grids [MSS^{+}19, PT24, WSST05]. Gro2mat [DDK14]. gromacs [DDK14]. Gross [AD14, AD15, CR13, CR20]. ground [CR13, CR20, XZL24]. Grounding [MSS^{+}19]. grown [RDP14]. Growth [Mol06b, Ano95d, BDL25, CI01]. Grundlagen [Bro94, GG04, von93]. GSGPEs [CR13, CR20]. GSGPEs-v1.1 [CR20]. Guaranteed [TCD^{+}22]. GUI [BHC^{+}24, BBHP^{+}23, BSS^{+}23, KA09, Lu96, MVDV97, NK18, Ozk23, SLM23]. GUI-based [BBHP^{+}23, BSS^{+}23, NK18, SLM23]. Guida [CSV94]. guidance [BAY98, Zar97]. Guide [Ano02a, CSV94, DB93a, EI05, Gra92, HL95, HH00, Ife05, Kat03, LL86, LS88, LS93, Lju88, Lju93, The92d, The92a, The92c, The92b, Mat92b, The92g, The93b, The97, The96, Mol80b, Mol81, Mol82, MLB87b, MLB87a, MW93, PHL95, SCC95, Tur00, Tur01, CT97, CGL97, DIN06, HL97, HH05, HH17, HLR01, HLR06b, HLR06a, HLR14, Lay03, LM91, Mag00, Mag05, MAB^{+}11, Mat97, The98, ML90, Nyh08, PES98, Mol18b]. Guided [Rab20]. Guideline [SP91]. GUIs [Ano99a, Mar95a, Mar99, MH03].

H [How15, SCB99]. HACopula [GHH20]. Hagenberg [Jef08]. Hahn [Mic00]. Half [Gau16a, HPZ19]. Half-Range [Gau16a]. halftoning [LA01]. Halifax [BEH94]. Hall [Ano96a]. Hallstatt [BB97]. hamper [Ano95l]. Hampton [Wie94]. Handbook [Ano09, Bee17, CV88, Gek08, Hog07, Ife05, Jai09, Kus02, KS05, MM02, PESMI96, Pha99, Ste10, Wik04, Yan05, Bai05, BPS02, MM08, Nør00, PES99, RC98, Ver09, Ano99b, Ano99c]. Handling [Tay95, How91, Mol97]. hands [Kin01, Kin06]. hands-on [Kin01, Kin06]. Hankel [BC17a]. Hardback [Mar19]. hardware [HNS^{+}01, LGML05, MBBC95, RB05]. harmonic [dMMLOS20]. Harmonics [SF14, AM01]. Harper [Lum92]. Harrogate [Coo95]. Hartford [IEE97e]. Hartley [Ves98]. Hawaii [IEE96c]. HDG [JRA^{+}18, FGjS15]. health [Ozk23]. heart [BlD^{+}20]. heat [Kin93, Mai00]. Heavy [RA95]. Heidelberg [BF97]. held [AO95, C^{+}97, Cou93, Dag94, DG96, EM94]. Heliospheric [HR96]. helps [Mol98a]. Hermite [Gau16a]. Hermitian [BCR03]. Herzliya [Ano96g]. Heteroclinic [DGKF12]. Heterogeneous [BCHS98, BSC^{+}00, PAG11]. heterostructure [RR01, RR02]. Heuristic [SYW^{+}24, CDOBA22]. Hierarchical [GHH20, KT14, KW05]. High [ACM97, ACM98, AS15, AJGO^{+}20, ASP^{+}18, CB98, FJSD96, HL95, HPK18, IEE98, ICL97, Koc14, The92d, The92a, The92c, The92b, The93a, The93b, The93c, The97, MT97, MBB^{+}09, RR01, RR02, dVSWAL17, SCB99, USE94, Cap13, CB99, Eat92, Eat97, Eat00, Eat02, Eat05, EBH08, HCBAEC23, JAC20, KH14, LM91, Mat95a, PT24]. High-Dimensional [HPK18, Koc14, PT24]. high-level [Eat92, Eat97, Eat00, Eat02, Eat05, EBH08]. high-order [Cap13, HCBAEC23]. High-Performance [ASP^{+}18, FJSD96, HL95, The97, The92d, The92a, The92c, The92b, The93a, The93b, The93c, LM91, Mat95a]. High-Precision [dVSWAL17]. High-Productivity [MBB^{+}09]. High-speed [RR01, RR02]. Higher [NP93, FGMS21, IP23]. higher-order [IP23]. highway [Ano95j]. HILBERT [AEF^{+}14, Kah02b]. Hill [Ano96a]. Historic [Mol18b]. history [DPR05, Mol18a, ML20]. HIV [CST20]. hm [MRK20]. hm-toolbox [MRK20]. Hoare [dOSRS19]. HODLR [MRK20]. holography [ZSW^{+}17]. Home [Che04]. Homoclinic [DGKF12, NGKM18]. homogeneous [Asi10, BPS99]. Honolulu [IEE96c]. Hooker [Bur10, Car10]. hopping [SKA19]. Horwood [Ano96b]. Hotel [Ano95d, Ano97h, Bjø01, D^{+}95, IEE93a, IEE94f, IEE95b, IEE96f, IEE96e, IEE97c, IEE97b, IEE97e, Ter94]. Hours [Hau96a, Hau96b, Hau97]. Householder [Tre10]. Hrebícek [Ste96]. HSS [MRK20]. HSTCG [SYW^{+}24]. htucker [KT14]. Hubert [Lip07]. Hughes [GSM95]. Human [Phi00, Fis95, NMS^{+}06]. Hundred [RS15a]. Hungary [Cse99]. Hunt [Ano00]. HVDC [BRS94]. Hyatt [Ano97h, IEE97c, IEE97a, IEE97b]. hybrid [BEK99, MPVA25, TK97, ZLMQ23]. Hybridized [JRA^{+}18]. Hydraulic [VVM93, Ren17]. Hydro [JE94]. Hydrology [Cla95, Hor98]. hydrostatics [Bir03]. Hyper [CDOBA22]. Hyper-Heuristic [CDOBA22]. Hyperbolic [Mat94a, Mol98b, Sha05]. hypercubes [PC13]. hyperplane [DH97a]. Hyphenated [KK96]. Hysteresis [CSB03]. Hytool [Ren17].

IA [IEE04]. IAS [IEE97c, IEE97c]. IASTED [Ham96]. IASTED/ISMM [Ham96]. IbIPP [IZBT21]. IBM [Ano97a]. ICCS [KZL^{+}20]. Ice [MNHH94]. ICIP [IEE94c]. ICIP-94 [IEE94c]. ICMS [BCD^{+}20]. ICPR [KPN^{+}04]. ICSEE [IGPF96]. ideas [Pen04]. identificacion [EGE95]. Identification [Ada97, Ben95b, DPE96, HCV97, ICL97, KDAB19, KPS95, Lju87, Lju88, Lju93, NRHP96, SP91, SWG^{+}94, TJML92, WK03, Abo03, Alt95, Ano96o, BS95a, Gaw04, GB89, GV94, Lju99, Ram94, VWHK96, Zhu01, vdM96, Ano94c, KAB97]. identify [Liu23]. idiots [Nah00, Nah02]. idle [MD95]. IEE [Ano95h]. IEEE [ACM97, ACM98, IEE93b, IEE94e, IEE94d, IEE94f, IEE94b, IEE94a, IEE95b, IEE96f, IEE96e, IEE96d, IEE96b, IEE97c, IEE97a, IEE97b, IEE97e, IEE97d, Ald96, CPUARC20, Kum14, MGC94, Mol96a, Rum23]. IEEE-754 [Rum23]. IEEE/IFAC [MGC94]. IEEE754 [Rum17]. IFAC [Bar92, BS95a, BLM97, FB95a, IF94, Pat94, MGC94]. IFAC/IFIP [FB95a]. IFIP [FB95a]. IFISS [ERS07]. Igniting [ACM03]. II [AD15, Kar03b, KZL^{+}20, PBI07, PR14, RAW^{+}16, SH18, Sha12]. III [JRA^{+}18, SKA19]. IIPBF [RKZ^{+}14]. IIR [DM06]. ile [Yuk96]. Ill [Han92, Han94]. Ill-Posed [Han92, Han94]. Illinois [Lum92, IEE98]. Illustrate [Ano95u]. Illustrated [SDPM04]. IMAC [W^{+}97]. Image [AM22, EO94, GWE04, IZBT21, JNR19, Mar16, NPP04, TS14, TS21, BGGL20, BC06, CiA02, Dha03, GW02, LP13, Mat97, McA04, NA02, PB01, Qur05, Sem04b, SB10, Thy10]. Image-based [IZBT21]. imagery [LHZT23]. Images [Ano95q, KAR^{+}19, MMOP95, HKF^{+}20]. Imaginary [CH20]. Imaging [MLAS21, Shu17, Sou94, Bar16b, HSM04, SBL19, STF01, Sza04]. Impact [WB12, Zem97]. Implement [Hig02a]. Implementation [ACFK02, AHS94, Ano98c, AKB94, EPJ^{+}05, GV25, GSM95, GMS92, Ife05, Joh18, OF92, Rom97, RD22, SBL^{+}10, SO93, Shu17, TBH21, TC97b, ZZG^{+}14, ACF99, AEF^{+}14, Bob05, BPS99, Che08, Din02, DS09, For06, HCBAEC23, Kea17, KMLP^{+}23, Li20, Lum02, MD95, MPVA25, MKU22, Mei05, MR11, PT24, Rad96, Rid95, Rov90, SSW09, SCB99, TC13, VC06b, Váz16, Ves98, Whi91, WS04, WSST05, Yan17, vdM96]. Implementations [ASG94, BC05, DP08, KL01, KG05, Rec00]. Implemented [Rum23, Bün20, MT84]. Implementierung [Mei05]. Implementing [AW92, JMD08, Rei93, AI21, FBH17, Wea97]. Implicit [JH97]. Implicitly [Rad96]. important [Mol03a]. Improved [AW97b]. improvement [FM24]. Improvements [Zag16, Dri05]. in-depth [CT97]. included [Ano93b, Ano95c]. Income [Nyh08]. Incompatibilities [Kwo93]. incomplete [Jon09]. incompressible [ERS07].

Incorporates [Mol00b, Mol00a]. Increase [MLAS21]. Increasing [Mol00b]. independent [BV21, FR18]. Index [SRK99, HL03a, HKF^{+}20, JGGF23]. Index- [SRK99]. index-2 [HL03a]. indices [HG24, HGM^{+}25]. Induction [LJR93, PH94, DP95]. Inductor [SO93]. inductors [AM95]. Industrial [IEE96b, Mac00, MBMW95, Ano96d, ASA96, Gri94, Ran96]. industries [Ano95o]. Industry [IEE97c, Ano95n, Mac00]. inequalities [AVV97]. inequality [JR99]. inertial [GWA01]. infallibility [Mol95c]. infeasible [Yan17]. Infer [VBB18]. inference [Ano94e, JB06]. INFFTM [CZ17]. infinite [HS24, RKZ^{+}14, VC06a, VFG04]. Infinity [FGMS21]. informatics [PJ97]. Information [Ano97i, Sch99, Tót08, YL99, dP96a]. ingegneria [Moo08]. Ingenieria [DACV95]. Ingenieure [Beu05, BB95b, GG04]. Ingenieurmathematik [Ben98b, Sch05]. Ingenieurpraxis [Pie05b]. ingénieurs [MM10]. IngredientDB [BSS^{+}23]. Inherited [Rum17]. inhibition [VHM17]. Initialization [DGKF12, IZBT21, HL03a]. initio [XSS20, ZJKS23]. Innovation [ACM03]. Input [DZ96]. INS [LRD^{+}95]. insects [CH23]. Insight [VF10]. insights [HL22]. Installation [Mol81]. instance [Faz10]. instead [Tre15]. Institute [Ano94i, IEE96b]. Instruction [CW05b, CW05c, LPD^{+}17, WW94]. Instructional [Hil91]. Instructions [LPD^{+}20]. Instructor [SSH94b, SS96b]. instrumentation [BN00]. insurance [Kat09]. Integer [LCMCD22]. Integer-preserving [LCMCD22]. Integers [Gou20]. Integral [AS08, Gau16a, Boy15, Dat13, Mai00, RKZ^{+}14]. Integrals [KR22, Bat19, Gau15, HS24, KR23, VC06a, VC06b]. Integrands [Joh18]. Integrate [Gid95]. Integrated [Ben95b, BDM17, BDM18, Jes01, Att02, Att10, Li99, Nor05, OG95, Rog03, Cou93]. Integrates [Ano88, CdFCS98]. Integrating [MT97, Sha11, Sha12]. Integration [BPCCM23, FPBO98, GRDL^{+}12, MMPOQ^{+}25, TM97, TCD^{+}22, GWA01, HH03, KS05, LJ23, PC13]. integrative [LYH^{+}16, TYL^{+}16]. integrators [BSW07, Cap13]. Intelligence [IEE94f, CAC94, JSM97, SS10, YL99]. Intelligent [Ano94h, Ano96e, Ano97b, ZJ01, Ano94a, Ano95f, HKF^{+}20, LL96, MC02, Dag94]. intensity [ZE95]. interaction [Can22a, Can24]. interactions [Gar07]. Interactive [Ano96r, Ano05, Dak05, DGGM96, DGKF12, Kat03, MASV96, Mol80a, Pao99, Rav94b, Tay99, Van04, Yan05, CT97, Eat92, Eat97, Eat00, Eat02, Eat05, EBH08, GRH^{+}21, LC98, Pao01, Pri00, Rob96, TB95]. intercept [Pac04]. Interconnection [Elg01]. Interface [BPCCM23, Bro07, Coo96, Dun99, Jon18, LSvdV19, The92a, The92b, The93a, SS18, UÇA10, Ano95l, Ano96h, BNN16, CF16, CV00, EK23, FM24, GV08, LHZT23, MC97, TSA21]. interfaces [BLL^{+}15, BQOvdG05, Wet20]. Interfacing [But92, HSR01, Li96, BEV06, KF96]. Interior [NN94, TQ96, Ano96q, BPS99, Yan17, Zha98]. Interior-Point [NN94, BPS99, Yan17, Zha98]. internal [MD95]. International [ACM96, Ald96, Ano93e, Ano94k, Ano94i, Ano95s, Ano96v, Ano97i, BCD^{+}20, BM09, Bry96, Bub95, C^{+}97, Cse99, D^{+}95, DG96, Glo98, Ham93, Ham96, IEE95a, IEE96f, IEE96e, IEE96d, IEE96c, IEE96b, IEE97a, IEE97b, IEE97e, IEE98, IEE04, IGPF96, Jef08, Kin93, KPN^{+}04, KZL^{+}20, Lum91, Lum92, M^{+}94, OB93, Ran96, dOSRS19, SAE95a, SAE95b, Sas96, Sil96, W^{+}97, WD98, Ano96u, H^{+}96, Ken95, Ano93d, Ano94d, Ano95i, Ano95j, GCP97]. Internet [Fos99, OE95]. Interpolation [JE94, Kah02a, CDSV10, CDSV11, Dor00, KW05, PC13]. Interpretation [FY18, LV03, Ren17]. Interpreting [RG21]. interprocess [TKD07]. Intersection [OBD23]. Interval [Har02, Rum95, BDR04, ZL04]. interval-set [ZL04]. INTLAB [Bün20, Rum95]. Introduccion [DACV95]. Introduces [Ano97d, Ano97e]. Introducing [DB93b]. Introduction [Atk05, AC97, AC99, Beu05, Bor18, BD95, BD00, BV18, BH97, BGG98, But08, Coo98, DAC95, DS04, EBB05, Ett96, EKH02, EMK04, EKM05, Ett10, Fis19, FJSD96, Gra11, GK05, Iro15, JRA98, JRA02, JK93, Kao97, KG01, Law05, Li20, Lin99, LS05c, MBM01, Mun13, Nei10, New94, Orf96, Özb00, PS96a, PS96b, Pal98a, Pal01, Pal05b, Pie05a, Poz05, Poz14, Pra96, Sch98, Shi99, Spa03, Sta05a, Sta05b, SM97, Str93, Str03, Tho00, TW02, TW06, Van97, Van00a, Wil96b, AGL07, AR03, AR04, Ano96j, Att09, Att12, Bät20, Ber03, Bra02a, Bra06, CÇ01, Col05, DJKP07, Dow02a, Eng05, Gil04, Gil05, Gil08, GS08, Hig04, Hol04, Kap04, KG06, KG12, Kno00, Kur00, Led04, LQT18, LW03, McA04, Mei05, Men01, Nor04, OD05, Pra99, Pra02]. introduction [Pra06, RB98, Sem05, SH09, Sto13, Tho04, VF10, Wal02, War13, YG99, DACV95]. Introductory [Che18, Hea97, Kol93, KH97c, KFG94, Sch14, WJK02, Tib93, Res19]. Introduttive [Tib93]. Intuitive [Kay05, Kay06, Rus08]. inv [Kah04]. invariants [AVV97]. invasive [CH23]. Inverse [BKL19, Gro99, ATB05, Boy15, Mol98b, Ves94, Vog02]. Inverses [Kah02b]. inversion [BBF^{+}19, BBN^{+}22, DdAF^{+}20]. Inverted [SDPM04]. investigate [RDP14]. Investing [Ano95k]. invitación [GSS05]. Invitation [Sch96]. Iowa [C^{+}96, C^{+}96]. IPPS [IEE96c]. IPscatt [BKL19]. IR [GHN19]. irbleigs [BCR03]. Ireland [Ken95]. iron [GO97]. IRT [She08a, She08b]. IS-95 [Lan00]. ISAP [HPK18]. ISBN [Ano93b, How15, Mar19, Mun12, Res19, Sha04]. ISBN-13 [Mar19]. ISMM [Ham96]. isn’t [Mol95d]. isogeometric [Váz16]. Isoparametric [BCH06]. isotropic [Asi10]. Israel [Ano96g]. ISSAC [Jef08, Glo98]. ISTE [How15]. Italian [Moo08]. Iterated [KR22, KR23]. Iteration [Qur01]. Iterative [BBC^{+}94b, BBC^{+}94a, Kel95, Kel99, NPP04, GHN19, HSH12]. IV [M^{+}94]. IVPs [CMR17]. Izmir [EM94].

J [Ano00, Bur10, Car10, Res19]. Jacobian [Bat19]. Jacqueline [Lip07]. January [Ano95b, Ano96m, IEE96a, IGPF96]. Japan [Ano94b, IF94, Ano94b, IEE96e, IEE96b]. Japan-USA [Ano94b]. Japanese [Ada97, BBC^{+}94a]. Java [Alt12, Ano97a, Ano97j, AC99, Bra97a, Has12, LP05, LS04, PMTL14, SDPM04]. Java/ [SDPM04]. JavaScript [FBH17]. Jeffery [Mar19]. Jeffrey [Alv11, Ano12, Mar14, Ni22]. Jesus [Bar16a]. Jírí [Ste96]. JMASM [ATvBB17, Alo06, BD20, ÖK17]. Joaquim [Veh07]. Joel [Ano95c]. John [Ano96b, Ano00]. Johnson [Epp11]. Joint [MGC94, BKR97]. Jose [ACM97, Gra94]. judgments [tHLMN19, LMN18]. Jul [OB93]. Julia [Bät20, Bät20, Bor18, CLMM20, KR22, KR23, MSS^{+}19]. July [Ald96, Ano94b, Ano94i, Bar92, BCD^{+}20, BS95a, BM09, EM94, IEE93a, IEE98, ICS96, Jef08, Ken95, Lak97, UCL04, Ano95o]. Jun [Wie94]. June [Ano94e, Ano95k, Ano97h, BM09, DG96, FB95a, GCP97, IEE94f, IEE95a, JLM96, KZL^{+}20, MO95, Pav93, SSV95, Syd95, UCL04]. Just [AP01, Mol15]. Just-In-time [AP01].

K. [Ned95]. K6 [Ano97c]. Kalman

[ZM00, BH97, GA01, SD91, WS04]. KCC [LLW^{+}23]. Kernel [BB96, STC04, CCF02]. Kevin [Ano00]. KGaA [Res19]. Kin2 [TCE16]. Kind [AS08, DHR11, Boy15]. Kinect [TCE16, TCE21]. Kinematics [WK99]. kinetics [Do98, Duo98, RDP14]. KinZ [TCE21]. Kirk [Ano93b]. Kit [SGA95, PT24, DS16]. Knife [DLA^{+}16]. Knowledge [CCF02]. Kobe [Ano94b, IEE96e]. Kohn [JZW^{+}22, YMLW09]. Komputerowe [Szy93]. komputrowe [Ano94g]. konferencji [Ano94g]. KPC [CS12]. KPC-toolbox [CS12]. Kramers [Luc05]. Krönig [Luc05]. Kronrod [Joh18]. Krylov [BGPPRW14]. Krylov-methods [BGPPRW14]. KSSOLV [JZW^{+}22, YMLW09]. Kutta [CMR17, Cam06, EH07].

L [Alv11, Ano00, Ano09, Ano12, Han06, Jai09, Jer06, Kus02, Kus06, Mar14, Mar19, Ni22, EM14]. L-BFGS [EM14]. L1C [SS18]. Lab [EO94, NSXZ14]. Laboratories [SDPM04]. Laboratory [Bee94, CFGG94, FSO93, GL96, Jor94, Mol80a, Rum95, SO93, SC97, Bre03, Don95, FGCG94, Kum05, Mit99, Mol88, Mol15, PV10, SI00]. Labs [Law96b, SB96, Ano96a, HZ94, HZ96b]. LabVIEW [Sim99]. Labwork [DL96]. Lær [Hau96a, Hau96b, Hau01]. Lake [Ano96u, IEE94d, I^{+}96]. LALSuite [Wet20]. Language [Ife05, SM09, WW99, Eat92, Eat97, Eat00, Eat02, Eat05, EBH08, JB07, MF94]. Languages [But92, CMKH03, Edw09, PBB22, USE94, USE99, WB12, Cas14, KS97, RGM22, H^{+}96]. Lans [AO95]. LAPACK [Mol00b]. Laplace [Mea02, DL95, DL01, TR04]. LARC [Wie94]. Large [EM14, PSR16, BCR03, BGPPRW14, BFG^{+}14, BD22, FB95b, GHN19, HL03a, Mol95a, Mol99b, MY05, Rad96, RSS08, Zha98]. Large-Scale [EM14, BD22, GHN19, Mol99b, Rad96, RSS08, Zha98]. Largest [Mol02b]. laser [SZM^{+}14, SLM23]. lasers [BN10, BN11, CWP98]. late [Ano97g]. Latent [She10]. Later [MV03]. lattice [Kat09]. lattice-subspaces [Kat09]. law [UW12]. Lawrence [Lip07]. laws [How91]. Layer [GRDL^{+}12, CPUARC20]. layered [CCM^{+}03, RCT20]. layers [BSB20]. LBSA [SLM23]. LCPC [H^{+}96]. Leakage [RT96]. Learn [Dak05, Hau97, Hau96a, Hau96b, Hau01]. Learning [CCF02, CGL97, Dri09, Fay17, Kec01, LMN18, CiA02, Dak06, EP 97, JSM97, MDB01, Möl07, PNGR00, ZLMQ23, ZLLT23, tHLMN19]. Least [BV18, RS08]. Lectures [CCF96, DG96]. Left [LCMCD22, Ves94]. Left-looking [LCMCD22]. left/right [Ves94]. Lehmer [DH97a, DH97b]. Leicester [Cou93]. Leiden [KF96]. Length [GJ20, Zla17]. Length/Weight [Zla17]. LEO [Ali02]. Leonid [Bar16b]. Lessons [Men95]. Letter [Web97]. Leuven [Sas96]. Level [Esm14, JMD08, MP99, MP00, USE94, Boo04, Eat92, Eat97, Eat00, Eat02, Eat05, EBH08, MBR21, SR09, Iro15, Mun13, RS15a]. Levitation [FKSM97]. LIBRA [VH10]. Libraries [PBB22, AI21, Ker95, LHW01, PMTL14, Wet20]. Library [AFL^{+}12, Ano97d, Bee17, BK07b, BCC^{+}17, CMKH03, CBhM12, CA97, DP16, Fay17, NK17, TCD^{+}22, AH09, DEQOR13, KS94, ST24, TC13, VH10, WPK^{+}18, UB95]. libstable [dVSWAL17]. life [HP02b]. lift [Pol92]. light [SLM23, ZSW^{+}17]. Lightweight [Fay17]. Like [HTCI96, Ano96p, Mor98, PMTL14]. likelihood [Jon09, Ano94e]. limit [GKD05, Por23]. Limited [BBEM22]. Limited-memory [BBEM22]. limiting [RAW^{+}16]. Lindfield [Ano96b]. Line [Bee94, MNHH94, Ano95l, YJ24]. lineal [SCL95]. Linear [BBC^{+}94b, BBC^{+}94a, Bor18, Bou97, BV18, Bur99, Che99, DH95, DHS03, Dat95, DR96, Dav98, Dav12, DL95, DL01, Dem97, DAC95, Dun99, FMW07, FB95c, Gaj03, GD99, Gra04, Gro94, Har05, HPZ19, HZ94, Hil96, HZ96a, HZ96b, Hog07, HAM02, JRA98, Kel95, KA02, KK01c, Kol93, LE96, Lat92, Lat05, Law96a, Law96b, Lay94, Lay97, Lay03, Lay06, Leo94, LHF96, Leo98, Leo02, LHF03, Lev92, LCL05, LCMCD22, McI16, Mic94a, Mor98, Nat94, ND88, Oga94a, OS06, Pen04, PBB22, RMS93, SP91, Str93, SB97, Str06, TC97a, TC97b, Wil96a, XCA07, Ano93b, AB03, Bay99, Bee05, BQOvdG05, BCG17, Bry02, But08, But11, Car98, Car00, Dat04, Dav13, Gri95, JRA02, Jon95b, Jon95a, Kok07, KH97c, Luo95, Mei05, MC02, MY98, MA95, OBCG19, Pet96, Smi97, SR90, Str03]. linear [SK94, SK95, SK00, TR04, TB97, Uhl02, Wit04, Yan17, Yan99, Zha98, ZZC^{+}08, ZD03, vdM96, Ano96a]. Linear-Quadratic [DAC95]. Linear/Nonlinear [Nat94]. Linearen [Gra04]. linearer [Mei05, ZD03]. lines [ACF99, CK02, SG09, Som07, Sut17, WSST05]. Link [IoG10, MBL^{+}97]. Link-Based [IoG10]. LinkCluE [IoG10]. Linking [KSF94, RVV^{+}92]. Links [Ano96i]. LINPACK [Mol94]. Liouville [LVV05, LV16]. Lipsman [Ano00]. LiScNLE [PT07]. lisis [GSS05]. LLVM [LH13]. LMI [AFOP19]. load [BDR04]. Loads [MNHH94]. Local [JS24, KYN95]. localization [LNLB19]. Loci [OF92]. lock [Ega98, Ega00]. Locking [RD22]. Locomotive [MFO95]. Logic [ÇA10, KW95, Ano96f, Kec01, Men01, YL99]. Logo [Mol03b]. London [Ano94c, Ano95b, Ano95t, How15, Ano93c, Ano96f, Ano96m, Ano97l]. long [Mol95b]. longitudinal [Kro95]. Look [BCC^{+}17]. looking [LCMCD22]. Looks [Ano97c]. loop [TKD07]. Loss [Kah98]. Lösung [ZD03]. Lösungen [BD95, BD00]. Louis [Dag94]. Louisiana [IEE97c, NP96]. Low [HP19, PR02, Pac04, Pol92]. low-lift [Pol92]. lowest [BC05]. lowest-order [BC05]. LSODE [Rei93]. LSTRS [RSS08]. LTE [Kum14]. lu [Kah04, LCMCD22]. Lumped [Nat94]. Lustre [BGGT21, TSCC05, CCM^{+}03]. Lyapack [DEQOR13]. Lyapunov [WK95]. lymph [ABB^{+}25]. Lyon [Pav93].

M [CF16, HA95, Dak05, KSF94, XSS20, ZJKS23]. M-file [CF16]. M-file/Mif [CF16]. M-Files [KSF94]. M-SPARC [XSS20, ZJKS23]. M-tutor [Dak05]. M2M [MG18]. MA [Ano93b, Ano95c]. Mac [Ano88]. Machine [CKC94, Nor05, JSM97, ZLMQ23]. Machinery [Gab98, Cha02a, Cha05a, FKU03, Nor04, Ong98, WK99]. Machines [IEE97b, Ano93d, Cat01, Gar01, Kec01, Lys00b, Nor04]. Macintosh [Fab95, The92e, Rov90, Tho94]. MACSYMA [Ben98b]. Made [Rul02, Fos01]. Madrid [M^{+}94]. MAGE [CF16]. Magical [Mol93]. Maglev [PGBG94]. Magma [ES10, Ste08, Ste13]. Magnetic [FKSM97, BBN^{+}22, HSM04, JAC20, SZM^{+}14]. Magnetics [Vit11]. Magnetostatic [WL94]. mainly [Mar68]. maitriser [MMR97]. MaJIC [AP01, AP02]. Majorana [DG96]. Making [BLL^{+}15, MDB01, Mol95a]. Malaga [Ano94a]. Malliavin [ST24]. Man [IEE97a]. management [LNC98]. Manifolds [NGKM18, EMMK01]. Manipulation [Hon91, HKF^{+}20, Hon92]. Manipulator [MMPOQ^{+}25]. Manipulators [SS96b, SS96c, Kog97, SS00]. Manual [BL96a, Lin05, Pol95, SSH94b, SS96b, Don95, EBH08, EaoGOBHW14, SYTD04, SBC04]. Manufacture [M^{+}94]. manufacturing [Rib00]. MAPLE [Ben98b, Ste96, GH04, GH93, WEM98, Ano05, BF09, CW05b, CW05c, CW05d, CW05e, ES10, GH95, GH97, GS12, KSS07, LP05, LS04, Möl07, Pri00, SH09, SCL95, Ste08, Ste13]. MAPLE/MATLAB/C [WEM98]. Mapping [Dri96, PBB22, BD22]. mappings [Nas20]. Maps [NGKM18, AVV97, GGKM09]. March [Ano94d, Ano95i, Ano96t, BF97, Joh95, KF96, MGC94, NP96, Pat94, Ran96, SAE95b, Sin93, Ter94]. Marchand [Ano99a]. Marek [Iro15, Mun13]. marine [Are94, BBHP^{+}23, LHW01]. MarineEpi [BBHP^{+}23]. Mark [Wea97]. market [Dow02a, Dow02b, McN05]. Marketing [Ber95]. Markets [Nyh08]. Markov [PNGR00, Bar97, Cse92, She08a, She08b]. Markovian [CS12]. Marne [CAC94]. Marques [Veh07]. Marquis [IEE96f]. Marriott [IEE93b, IEE96f]. MARS [WEM98]. Martinez [Alv11, Ano09, Ano12, Han06, Jai09, Jer06, Kus02, Kus06, Mar19, Ni22, Wik04, Mar14, Han06, Jai09, Jer06, Kus02, Kus06, Wik04]. MartMi [LHZT23]. MartMi-BCI [LHZT23]. Masks2Metrics [MG18]. masonry [Por23]. Mass [M^{+}94, Ben02]. Massachusetts [Ano95s]. masses [BHC^{+}24]. Mastering [AB98, DH98, DH04, HL96, HL98, HL01, HL05, Hul99]. MatCal [LO16]. MATCH [BCHS98]. MATCOM [CdFCS98, Ker95]. MATCONT [DGK03, DGK04]. MatDL [Fay17]. MatDSP [CGM95]. Mateda [SBL^{+}10]. Mateda-2.0 [SBL^{+}10]. Matematica [MR95]. matematik [Bäc97]. MatEMTP [MA96a, MA97a, MA97b]. material [GMT96]. materials [GSM95, JAC20, Joh95, Luc05, TW98, VK05, YHC^{+}22]. Materialy [Ano94g]. Math [Ano97d, AH09, CA97, Fos99, Fos91, MR95]. MATHCAD [Ben98b, Ben05, Thi95, Jaf00, Möl07, Ben05]. MathCW [Bee17]. MATHEMATICA [Ben98b, Tay99, Bha05, Bha06, CW05b, CW05c, CW05d, CW05e, ES10, LP05, LTE01, MAC08, MM97, MM98, Möl07, Pao99, Pao01, Ste08, Ste13]. Mathematical [AR03, AR04, Ano95v, Ano98a, BF09, Bee17, BCD^{+}20, Bro94, CGI99, ES10, EM94, Gek08, Gop10, Hai08, Moe04, MS00, SH09, WR16, WR17, C^{+}97, Eng05, Hen94, JR99, Kno00, Man01, Mol88, Mol02a, Moo05, Nah11, PWR13, EM94]. Mathematical-Function [Bee17]. Mathematics [Bäc95, Bäc97, Bät20, CHT15, Coo14, DG96, Gru04, Jef05, Kar01a, Lum91, Lum92, Mat92a, Ogu95, Sch05, Whi04, WT94, WT97, Ano96u, Bäc00, Bäc04, Ben98b, Dav04a, Duf03, Duf10, Duf11, Esf03, HDR97, HDR00, Hig04, Jef04, JL01, Mac00, MM97, MM98, Mun12, Rog03, WHT02, WTH03]. Mathematik [Bät20]. mathematisches [Bro94]. MatHH [CDOBA22]. maths [Ano95m]. MathWorks [Ano96i, Ano97d, Ano97e, Ano98b, Mol03b, Mol06b, Ano96h]. MATLAB [Ano95b, Ano12, App19, Bar16a, Bät20, Ben98b, BB95b, Bjø01, CSV94, CK23, Epp11, FRAK15, Löw01, Mar14, Mei05, Mic00, MO95, O’B13, Pie05a, RBD^{+}11, Res19, Tay99, Tib93, Tow16, Veh07, WEM98, ZW93a, Beu05, PT24, RC16, Sha08a, AM17, Ada97, AS15, ADW23, AP02, Alt12, ÁBZ17, ÁBZ20, AB98, AS12, AHS94, AB96, And95, And05, AV15, Ano88, Ano94f, Ano95e, Ano95l, Ano95p, Ano95q, Ano95r, Ano95u, Ano95v, Ano96j, Ano96k, Ano96n, Ano96o, Ano96p, Ano97f, Ano97d, Ano98c, Ano98a, Ano99d, Ano02b, Ano19, Ans14, AC96, ANV00, Asu02, ASA96, AI21, AY91, ASG94, Atk05, Att09, Att12, Att95b, Att95a, Att96, Att99, Att02, Att04, Att10, AEF^{+}14, AY94, AS96, AY96, AY97, AW92, AMR18, Bab94, BV95, Bäc95, Bäc97]. MATLAB [Bäc00, Bäc04, BK06, BK07a, B10, BCR03, BC05, BP96, BD20, BCHS98, BSC^{+}00, Bar97, BKR97, BF09, BFSJP^{+}21, BPCCM23, BNN16, Bat19, Bät20, BT04, BD96, BB96, BKGS02, BKG05, BHC^{+}24, BGGL20, BW20, Ben95b, Ber09, BSW07, Bha95, Bha05, Bha06, BN10, BN11, BMR19, BB95a, BB95b, BB99, BB02, BCG17, Bis93, BPB96, Bis97, BC17b, BK07b, BAEBAS19, Bod98, Bol94, Bor18, Bor97, BV19, BlD^{+}20, Bou97, Boy99, BFM89, Bra02a, Bra06, BPS99, Bro07, BH97, BBEM22, Brz97, BDS97, BDS02, BEV06, BV08, BDL25, Bün20, BKL19, B^{+}94, But08, But11, CPUARC20, CLTS20, CH23, CR13, CZ17, CR20, CF91, CST20, CS12, CGRvD15, CHMN13, CSB03, Cat95, Cat01, CSV94, CSV96, CK22, Cha00, Cha02b, Cha04, Cha05b, CATK11, Cha02d, Cha17, Che18, CL96]. MATLAB [CGI99, Che08, Chi97, CF93b, CF93a, CF95, CFGG94, CBL24, Chu00, CGL97, Cob21, CV00, Col05, Con95b, Con95a, Con96, CM99a, CdFCS98, CSY15a, CSY15b, Coo00, Coo96, Coo98, Coo01, Cor95, Cor96, Cra96, CI01, CNJ97, CFF^{+}91, Cur95, Cze95, DACV95, DHS03, Dak06, DJKP07, Dat13, DR96, DDW93, Dav99, Dav04a, DS05, Dav11, Dav13, DM90, De 96, DGGM96, DP96b, DP99, DGKF12, DDD97, Deg20, DdAF^{+}20, DPE96, DB93a, Den98, DGK03, DDK14, DIN06, DF99, Dja98, Dja00, DHR11, DH12a, DM06, Don95, Dri96, Dri05, Dri09, DH12b, Duf03, Duf10, Duf11, DK99, Dun99, Dut16, EO94, EP94a, Edw09, EK23, EAK01, EA04, ElA04, EI05, EA95, EMMK01, ES10, EM14, Esm14, EGE95, Ett93, Ett96, Ett97, EMK04, EKM05, Ett10, Fab95, Fab97]. MATLAB [FR18, FJ22, Fas07, Fau99, Fau08, Fay17, Faz10, FD01, FGCG94, FR96, FS23, FBG94, Fis95, FPBO98, FCP97, FBH17, For06, FC95, FC00, FM24, FMS25, FGjS15, GL96, GV25, GH93, GH95, GH97, GH04, Gar96, GP96, Gar01, Gar07, GR93, GRH^{+}21, Gau16b, GHN19, Gek08, GLJ^{+}91, GG92, GGKM09, GMT96, GSM95, Gid95, Gil04, Gil05, Gil08, GS08, GMS92, GS07, GL04, GD99, Góm15, GWE04, GCM25, GHH20, GB89, Got95, GKD05, Gra92, Gra04, GV94, Gre16, GA01, GS12, Gri95, Gro05, Gru04, GG04, GPK05, GV08, GK05, GO97, GJ03, GPR14, GB03, HTJ90, HC95, Hah97, Hah02, HV07, HNS^{+}01, HHF95, HL03a, HK95, HL96, HL97, HL98, HL01, HL05, HSH12, Har05, Har02, HDR97, HDR00, tHLMN19, Har01]. MATLAB [HG24, HGM^{+}25, HC00, Hat01, Hau96a, Hau97, Hau01, Hau90, HCV97, HR96, Her01, Her96, HCBAEC23, Hig89, Hig91, Hig93, HH00, Hig02c, Hig02a, HH05, HH17, HM22, HR91, Hil91, HC92, HZ94, HZ96a, HZ96b, HT97, HTCI96, HSR01, HKF^{+}20, HLP96, HL22, Hon92, Hor97, How91, HPK18, HAM02, HAM06, Hul99, Hun98, HLR01, HLOR05, HLR06a, HLR14, HLS08, IoG10, IZBT21, Ife05, IP97, IP00, Jac96, JRCS95, JCRS96, JH96, JHPK97, JH97, Jaf00, JSB20, JRA^{+}18, JR99, JZW^{+}22, JAC20, Joh11, Joh18, JB03, JB07, Jon91, Jon95b, Jon95a, Jon18, JDFV08, JD13, KN95, Kad95, Kad97, Kah04, KSP11, KH97a, KH00, KK01a, Kao09, Kar01b, Kar03b, Kar03a, Kar03c, Kar05, KR22, Kat03, Kay05, Kay06, Kep09, Ker95, KG01, KG06, KG12, KS01, Kin95, Kin97, Kin01]. MATLAB [Kin06, KM10, KM11, KAB97, KA02, Kiu10a, Kla92, KA11, KK01c, KSS07, KW05, KH96, KH97b, Kni00, Jiř97, KK97, KA09, KSF94, KPS95, KS94, KMLP^{+}23, KSL93, Kro95, KK93, KNNM97, KS97, KT10, KH14, Kum14, KC95, KH94, KFG94, KYN95, KBQ97, KB97, KB00, Kwo93, LS98, Lam95, LSPM95, LH14, Lan00, LPD^{+}17, LPD^{+}20, LL86, Law96b, Led04, LVV05, Lei02, Lei11, LL92, LW94, LL95, LMN18, Lev92, Li96, LZ17, Li20, LQ94, LLLW06, LLW^{+}23, LP97, LW03, LS88, LM91, LS93, LZ15, Liu23, Lju88, Lju93, LS05b, LSvdVK09, LSvdV19, Löw01, LSH95, Lu96, LHW01, LJR93, Luo95, LE00, Lus09, LTE01, Lyn04, Lys03, MMD24, Mac91, Mag02, Mag00, Mag05, MAB^{+}11, ME04, Mah05, MA96a, MA97a, MA97b, MBL^{+}97, MPVA25, Man01, Man96, MSL09].

MATLAB [MKU22, Mar95a, Mar99, MH03, Mar92, MBBC95, Mar03, Mar07, MNHH94, MGW99, MM02, MM05, MMS11, MMS17, MLF^{+}12, MRK20, The90, The92d, The92a, The92c, The92b, Mat92b, The92g, The93a, The93b, The93c, Mat95a, Mat95b, Mat95c, The98, Mat94a, MF94, MF99, MF04, Mat92c, Mat94b, McA04, McC98, McG91, McI16, McM07, Mei05, MCd^{+}96, MP99, MP00, MMPOQ^{+}25, Mic94a, Mic94b, MVDV97, Mit99, MA96b, MMR97, MM00, Mol80a, Mol80b, Mol81, Mol82, MLB87c, Mol88, ML90, Mol93, Mol94, Mol95d, Mol96b, Mol00a, Mol00b, Mol02a, Mol04a, Mol04b, Mol04c, Mol04d, Mol06b, Mol08, Mol11, Mol18a, Mol18b, ML20, MY05, MR11, Moo07, Moo09, Moo11, Moo15a, Moo15b, MVM97, Mor98, MV22, MSS^{+}19, Mro95a, MM96, MBMW95, Nah01, Nak96, Nak02, Nas20, Nei10]. MATLAB [New93, NRPH96, NRHP96, NK17, NMS^{+}06, OF92, OBCG19, OBD23, Oga94a, Oga94b, Oga08, Ogu95, OG95, ÖK17, Ong98, OE95, OD05, OAKS11, Ozk23, PF10, PGBG94, PS96a, PS96b, PV10, Pal98a, Pal01, Pal05b, PH94, Pan89, Pao99, Pao01, PMTL14, PMNWR20, mP97, PESMI96, PES98, PES99, PWR13, PR14, PHL95, PPD95, PIAH12, PA11, PL95, PS04, Pfe95, Pie96a, Pie05b, PNL^{+}21, Pol95, PA99, PA04, PB01, PPS06, PC13, Por23, PR06, Poz05, Poz14, PAG11, Pra96, Pra99, Pra02, Pra06, PS98, PS00, PSB04, PSR16, QCPG96, QS03, QS06, QMS98, Qur05, RV13, Rah93, RDP14, RCT20, RHB96, RHG09, RA95, RBD^{+}10, Rao11, RS08, Rav94a, Rav94b, Rec00, RC98, RBC20, RGM22, Ren17, RVV^{+}92, RAW^{+}16, RHR^{+}21, RT96, RH92, Rid95, RG21]. MATLAB [RBZ96, Rob96, Rob04, RA92, Rod92, RS15b, RSS08, RD22, Rov90, Rov10b, RB04, RB05, dVSWAL17, Rum95, Rus08, Rus93, RSW15, Saa93, ST12, SMS95, SMS96, SHX96, SBL^{+}10, SB90, Sch97, SI00, Sch12, SH00, SH05, Sch05, Sch04, SY20, Sem04b, Sem05, SR97, SRK99, ST01, SGT03, Sha05, Sha07, Sha08b, Sha11, SM09, She08a, She08b, She10, Sig92, Sig94, ST98, SD02, SM14, SA01, SM15, SCL^{+}18, SZM^{+}14, SD91, Smi97, SBL19, SGA95, SCB99, Sou99, Spe95, Spe96, SWS97, SWG^{+}94, SD96, Ste03, SH11, Ste13, SB96, SYTD04, SZCP21, SP22, SK94, SK95, SK00, Stu99, SS10, Sut17, SU94, TJML92, Tay95, TC97a, TK97, TC97b, TH09, TRD11, TSA21, TB95, TCE16, Tew02, TS14, TS21, Thi95, Tho94, Thy10]. MATLAB [Tib93, Tob11, TTT99, TCD^{+}22, TGM06, Tre86, TT^{+}96, TMC^{+}99, Tre00, TW98, TBHS94, ÜKP05, UÇA10, UW12, Van97, Van00a, VF10, Van04, VFV13, VA94, Ven02, VH04, VH10, VWHK96, VHM17, Ves94, Ves98, Vid11, VRVAC23, Vit11, VK05, WTL00, Wan15, War13, WD96, WM95, WM96, WR00, WR16, WR17, WWM06, WL94, Wer03, WR04, WW94, WGP95, Whi91, Whi04, WB16, WCS92, WT94, WHT02, WTH03, WE96, WSST05, WK95, XZL24, XCA07, YCCM05, Yan05, YMLW09, Yan17, YR19, YHC^{+}22, YA95a, YA95b, YJ24, Yuk96, Zek17, ZL04, Zen97, ZC08, Zha98, ZZC^{+}08, ZSW^{+}17, ZLMQ23, ZW93b, ZLLT23, dMMLOS20, von93, Rob95, Kir93, Lin05, Ano97c, AKB94, Cse92, DL96, MC97, Rei93, AGL07, ATvBB17, AB13, ACF99, ACFK02, AP01, Alo06, API^{+}19, Alt95, Ano95g]. Matlab [Ano95h, Ano95m, Ano96c, Ano96d, Ano96h, Ano96l, Ano96q, Ano96r, Ano97a, Ano97g, Ano97k, Ano97j, Ano05, AD14, AD15, Asi10, AS08, AH09, AC99, AW97b, BC17a, BH96, BR96, Bal19, BJG94, BCH06, BGPPRW14, Bee94, Bee05, Ben05, Beu05, BBHP^{+}23, BFG^{+}14, BC22, Bla02a, BC06, BO19, BP97, BBF^{+}19, BBN^{+}22, Bra97b, Bra97a, Bra02b, BSS^{+}23, Bro95, Bug95b, BF95, Bur93, Bur94, But92, BEK99, BD22, CDSV10, CDSV11, CMR17, Cam06, Can22a, Can22b, Can24, Cap13, CGM95, CK02, CH20, CCF02, CBCC96, CB98, CB99, CF16, CBhM12, CFPF94, CG00, CW05b, CW05c, CW05d, CW05e, Cho02, CA97, CLMM20, CFR19, Coo14, CDOBA22, ÇA10, Dak05, DZ96, DLA^{+}16, DP16, DS16, DGK04, DS09, DRR97, DRS18, Edd09, Edl04, ERS07, EL16, EPJ^{+}05]. Matlab [Esp07, EKH02, FMW07, FHH99, Fis19, FSC95, Fos01, FKSM97, FK11, Gab98, GI06, GRDL^{+}12, GBM15, Gau05, Gau06, GSS05, Gop10, GR97, Gre94, Hai08, HH03, HL95, Han92, Han94, Han99, HTNFBS06a, HTNFBS06b, Han07, Has12, HS24, HA95, HL03b, Hen07, Hof98, HT12, Hoh14, Hon91, How95, HWB15, IP10, IP23, ICS^{+}18, Irw05, ICL97, IMM12, JNR19, JM94, Jon09, JGGF23, KAR^{+}19, Kah98, Kah02b, Kah02a, KR23, Kat09, Kea17, KMBP24, Kin98, Kiu05, KW95, Kog97, Kok07, Kok15, KA13, KT14, Kub95, Kun04, LP13, LMAV24, Law05, LV16, LLZ18, LCL05, LHZT23, LS05c, LRD^{+}95, LO16, LJC93, MAC08, MS94, MBR21, Mah00, Mak02, MM97, MM98, MDB01, Mar91, Mar16, MM10, The92e, The92f, Mat97, The97, MASV96, MT97, MS14]. Matlab [Meu20, Mid00, MG18, Mil20, MJ01, Mir96, Mir97, MMOP95, MB94, MON12, MLB87b, MLB87a, Möl07, MLAS21, MP18, Moo08, Mor18, MA95, NPP04, Nat94, NPT15, NSXZ14, NK18, NE05, Nyh08, Oga10, Ono01, PFG08, PSTO97, PY22, PR02, PBI07, PL00, Per93, Pet96, PT07, PT24, PMA94, Pol92, PF07, POVD96, Pri00, PNT15, Qua10, QSG14, Rad96, RZR12, RC13, Riv01, Rom97, Roq13, Rov10a, Rum01, SFK91, SDPM04, SS96a, Sar10, Sar17, SSW09, SG09, Sch14, Sch11, SL17, SH18, SKA19, SN01, SH93, SR95, SR02, Sha02, SKF05, Sha08c, Sha09, Sha12, SAKG15, SH09, SR09, Sim99, SR90, SB10, Som07, SCL95, SS15, Ste08, Ste09, Sti04, SW95, SBC04, SLM23, SM94, TP03, TCE21, TACA15, TSM93, TM97]. Matlab [TNBSF04, TQ96, Tót08, TG10, Tri08, UCL04, VC06b, Van00b, Váz16, VZDCM25, VPM16, VK16, Wal18, Wea97, Web97, WW99, Wei12, WB12, WT97, Wit04, Wri97, XSS20, YKS94, Zen04a, Zen04b, ZL13, ZJKS23, ZZG^{+}14, Zin93, Zla17, dP96a, de 05, vdH05, Alv11, Bol94, Bur10, Han06, How15, Iro15, Jai09, Jer06, LS04, Lip07, Mar19, Mun12, Mun13, Myr17, Ned95, Ni22, RKZ^{+}14, Sha04, Sta05a, Sta05b, Ste96, Ver09, Wik04, Ano96b, Bar16b, Car10, Hig95, RS15a, Shu01, Bow10]. Matlab-aided [HA95]. MATLAB-based [CK23, RC16, DK99, WTL00, BKG05, Bra02a, Bra06, CK22, Lei02, MVDV97, OAKS11, RB05, Sem04b, Sem05, SZM^{+}14, TB95, UW12, VRVAC23, YHC^{+}22, MLAS21, SDPM04, Zla17, Ano96r, BC22, BEK99, Cap13, CDOBA22, GI06, JGGF23, Kat09, Kea17, LHZT23, RC13, SAKG15, Som07]. MATLAB-derived [Pan89]. MATLAB-Implementierung [Mei05]. MATLAB-Like [HTCI96, Ano96p, Mor98, PMTL14]. MATLAB-MPI [BKGS02]. Matlab-Simulation [ZJKS23, XSS20]. MATLAB-software [KK93]. MATLAB-style [WB16]. MATLAB-to-C [LPD^{+}20, Mir97]. Matlab/C [Wal18]. MATLAB/GNU [FRAK15, JRA^{+}18, PSR16, RAW^{+}16, RHR^{+}21]. MATLAB/GUI [KA09]. MATLAB/INTLAB [Bün20]. MATLAB/Octave [AMR18, FR18, PC13, GV08, CDSV10, CDSV11]. MATLAB/power [CSB03]. MATLAB/PSI [PPD95]. MATLAB/SIMULINK [FCP97, HHF95, Hun98, Jon95a, KNNM97, MCd^{+}96, Ong98, RA95, WM95, WM96, KSP11, ZC08, ZLMQ23, Jon95b, AB13, CBhM12, ÇA10, DRR97, UCL04]. MatlabHTM [BSB20]. Matlabie [Brz97]. MatlabMPI [KA04, Kep05].

MATLAB(R) [Ano96a, Ano99b, Ano99c, JB06, Mar11, Rab20, SS18, ST24, BTM09, EU07, HLR06b, Kle07, LC09, MM08, SBL^{+}17, Whi07, WLB^{+}09, Ano09, Kus02, Kus06, Ano99a, Ano93b]. MATlab^{T M} [Ano95c, TH01, Ano00]. MATMOL [Som07]. MatNEC [Coo96]. Matrices [BV18, Bug95a, FSZD20, GMS92, Hig89, Hig91, Kah02b, Kla92, Lip07, Mar92, MRK20, VvBM08, AV15, BPS05, HAM06, KS94, LP97, RV13, RZR12]. Matrix [Bar97, Bee94, CV88, Cze95, FBG94, GV25, Gol91, GV89, Got95, Har01, Hig93, Hig95, Hig15, HPZ19, HM88, Hon91, IMM12, KS97, MGW99, Mol80a, Mol02b, MV03, Rob95, ST12, TH09, UÇA10, VvBM08, WR00, API^{+}19, BCR03, BMR19, GRH^{+}21, HM22, HK01b, HSR01, Hon92, KO00, LYH^{+}16, Mol00b, Mol15, PIAH12, Van97, Van00a, Wat02, Whi07, WB16, ZZC^{+}08, Hig02c, Ano95c]. matrix-vector [Van97]. MATS [KT10]. MATSLISE [LVV05, LV16]. matter [SAKG15]. mattorabu [Oga10]. MATVines [Cob21]. MATX [KSF94]. May [ACM96, Ano95n, Ber95, Bry96, BB97, FB95a, Gre96, Ham93, IEE96f, IEE97b, Mal96, Mro95b, Zem97]. McSAF [DH12a]. McTavish [GSM95]. MD [Gre96, IEE96a]. mean [UW12]. mean-square [UW12]. Means [LLW^{+}23, Alt95, Mil20]. MeasureBrArt [Lin05]. measurement [Dow02a, Liu23, Whi00]. Measurements [ATvBB17, Ano99d, FBC00, PBI07]. Measures [KT10]. Measuring [Dow02b]. Mechanical [Ben98a, Ben04, Rao04, SMB04, ARR02, FBC00, Kel00, Mag05, MAB^{+}11]. Mechanics [Che18, Gek08, Ono01, PWH02, Sha03, VK05, WT94, WT97, Cur05, Gre96, Hul99, Kin93, Lev03, LC98, Mac05, MKP02, Res19, SLI99a, SLI99b, TW98, WB99, WHT02, WTH03]. mechanics/estimation [Gre96]. mechanisms [Nor04]. Mechatronic [SU94, Moo98]. Mechatronics [Nec02, PV99, Lys00b]. med [Bäc97, Sim99]. MEDIA [Mro95b, Ald96, AGL07, App19, Gre16, RCT20]. mediated [Bha95]. Medical [Dha03, Shu17]. medicine [HP02b]. Medium [BKL19]. Meeting [AO95, IEE97c, JE91, Ano94e, Ano95d, IEE96a, Jef08, SAE95a]. meets [Ano97c]. megaflops [Mol94]. Melbourne [Ald96, Ano95d, Ano95o]. memory [BSB20, BBEM22, DJK93, JKR92]. Memristor [ASP^{+}18]. Memristor-CMOS [ASP^{+}18]. MENHIR [CB98, CB99]. Mesh [PS04, Kok15]. Meshed [BPCCM23]. Meshfree [Fas07]. meshless [Mil20]. metaheuristic [HBT16]. metastatic [BDL25]. meteorology [Hol04]. Method [ACFK02, AW92, GPR14, Hig02a, KBQ97, Lam95, Riz17, SF14, WR16, Dav04b, FRAK15, GSM95, HS24, HCBAEC23, JRA^{+}18, Kat09, Kel03, Kok15, KB97, KB00, Li20, MT84, MT00, Mil20, PWR13, PK08, PPS06, RBD^{+}10, RAW^{+}16, SSW09, SG09, Som07, Sut17, Tho04, VCK98, WSST05, YJ24, Rad96, RBD^{+}11]. methodologies [ZJ01]. Methodology [BP97, GV25, FGMS21, Jan02]. Methods [ABB^{+}25, AJGO^{+}20, BBC^{+}94b, BBC^{+}94a, BBEM22, Che93, Com90, Com92, Coo14, Cou93, DB08, Dav04a, Gar94, GBM15, Gek08, GK05, IEE04, JH97, Kel95, KG01, Kiu05, Mat92a, NPP04, Pao99, PR14, PL95, Poz05, Rao02, Ste10, TQ96, Tre00, Wal18, Whi04, Ajj95, ARRY01, Ano96b, Ano96q, Ban01, BGPPRW14, BO19, BTM09, Bor97, Bra02a, Bra06, BM94, Bur99, Cha05b, CC06, CM99a, CM99b, CS99, Dat04, Duf04, EH07, FB03, Fas07, Gar00, Gau06, GHN19, GS08, Goc02, GS12, HSH12, HAM02, Jon01, KF96, Kel99, KG06, KG12, KM10, KM11, Kiu10a, Kiu10b, KA13, KC95, KS05, LP13, LQT18, LW03, MF99, MF04, Mei05, Mol99b, MS00, Moo05, OBCG19, OD05, Pao01, PL00, Poz14, Ram97, RS02, Rec00, RHR^{+}21, Rob04, SH00, STC04]. methods [Shu01, Tay99, VH04, Vog02, WA96, YCCM05, Zha98]. metod [Mro95a]. Metodi [Com90, Com92]. Metrics [MLF^{+}12]. metrology [C^{+}97]. Metropolitan [Ano94i]. Meulman [Lip07]. Mexico [Ano97h, USE94]. MFEM [BC05]. MHz [Ano97c]. MI [SAE95b]. Miami [Ran96]. Michigan [IEE96d]. micompm [FR18]. micro [Bal19, Lys05]. micro-electromechanical [Lys05]. micro-resonators [Bal19]. microchips [SGLBA98]. microcomputers [Rov90]. microengineering [Lys05]. Micromagnetic [CF16]. Micromechatronics [GL04]. MICRONDE [MMOP95]. microscopes [HL22]. microscopy [BFSJP^{+}21, ZSW^{+}17]. microstrip [VRVAC23]. Microwave [Lev96, VCK98]. Midwest [C^{+}96]. mielno [Ano94g]. Mif [CF16]. Milovy [PJ97]. Milwaukee [Ano95j, IEE97b]. MIMO [AW97b]. Mineral [MLAS21]. MINImal [Ano95c, Asi10, Rob95]. minimization [Rov90]. minimum [ADD04, DGLN04a, DGLN04b]. mirror [HA95]. MIRT [She10]. missile [BAY98, Jon95b, Jon95a, Zar97]. Missouri [Dag94]. MIST [KFG94]. MIX10 [KH14]. Mixed [JRCS95, JCRS96, Ano96k, FR96, RC13, RC16]. Mixed-Norm [JRCS95, JCRS96, Ano96k]. MNPBEM [HT12]. Mobile [Pät02, FSC95, HP02a, VLV00]. Modal [D^{+}95, W^{+}97, WD98, Gaw98, Mag02]. MODCONS [WGP95]. Mode [Kah04, Zin93, BH96, For06]. Model [BW20, BDM17, BDM18, Bro94, Can86, CKC94, DDD97, HCV97, Rab20, RA95, Ros03, SO93, She08b, SYW^{+}24, Abo03, Ano97l, BSB20, BJ02, Bün20, CPUARC20, CH23, Cha95, Hun98, Mol96a, SH18, VHM17, Wri97, ZLMQ23]. Model-Based [BDM17, BDM18, Can86, Ros03, BJ02].

Modeling [AB13, ADW23, Ano95v, Ano98a, BP96, BR96, BD20, Cla95, CF93c, CFN01, CFN02, DF99, EP96a, EP96b, EP96c, EP04a, GR97, HHF95, HGM^{+}25, HB04, HP02b, How15, Khe96, Kir93, Kul99, Lam95, LG94, LRD^{+}95, MBL^{+}97, MVDV97, Nir02, Pal98b, RLH03, Rov10a, Rov10b, SP91, SS96b, SS96c, SKG97, SCL^{+}18, SS15, TK97, Wal18, Wea97, WM95, WM96, WL97, ZL17, ZE95, Bab98, Beq98, Beq03, BC04, Bru01, BEK99, CSB03, CDS09, Dav99, Doe98, EP00a, EP00b, EP03, Elg01, Gar07, GL04, Hay96a, JA99, Kro95, Mac00, Mak02, MIK00, Moe04, Moh01, PF10, PSP04, RR02, SBC04, TtsT04, Whi07, XWZ^{+}22, Pat94]. modelled [PR02]. Modelli [Com90]. Modelling [BV95, CH23, CBCC96, GG92, MC02, SS00, SR94, SU94, Tho94, AM01, Ach04, Are94, BH96, BF09, ERS07, EM94, Mor00, Nør00, TJML92, XWZ^{+}23, ZLW^{+}19, Ham93, Ham96, JLM96, Pav93, SSV95]. modelo [EGE95]. Models [Bug95a, Chr97, Com90, Dun99, FFR^{+}24, Góm15, HYY^{+}15, MNBB19, McI16, PA11, SS96a, She08a, She10, SM07, Whi04, AR03, AR04, And05, Ano93c, BN10, BN11, BSS^{+}23, CBL24, CT97, CI01, DP08, GRH^{+}21, HBT16, JS24, Kec01, LNLB19, Mic94a, RL00, Row03, SG09, SN01]. moderate [JAC20]. Modern [Bis93, Bis97, Bos01, Bos02, Cou95, DH95, DB95, DB98, DB04, FS23, GL96, HK01b, LA01, Oga97, Oga02, Ogu95, Shi92, Shi98b, Tew02, ANM01, Ben02, Mei05, MN03, MBM01, Shi98a]. moderne [Mei05]. modes [Bal19]. Modified [Riz17]. Modular [DRR97, LH13]. Modulated [LJR93]. modulation [Bin00, Van01b, ZT02]. modulators [SLM23]. Module [AB13, AM95]. Modules [MASV96, Rob94]. molecular [MMD24]. Molecules [Che18, Res19]. Moler [Dut16, Hai08]. Moment [SF14]. Monitor [MNHH94]. Monitoring [Shu17]. Monotonicity [Mik23]. Monte [NSXZ14, CGRvD15, Mil20, SAKG15, She08a, She08b]. Monterey [Ano96t, Ter94]. Montfort [Cou93]. Montréal [OB93]. MooAFEM [IP23]. Moosh [DLA^{+}16]. MORLAB [BW20]. morphological [BlD^{+}20]. Morphometrics [MBGV93, MG18]. Motif [MON12]. Motion [IEE93a, Can22b, STF01]. motions [PK08]. Motor [ÇA10, Kir93, LJR93, PH94, RD22, DP95, LHZT23, SFPO94]. movement [Hun98]. moving [SSW09, Som07]. moving-grid [Som07]. MPBEC [VPM16]. MPCC [MR11]. MPEC [DF99]. MPI [BKGS02, Whi04]. MR2404036 [Ver09]. MS [The92f]. MS-DOS [The92f]. MSCPDPLab [ZLLT23]. MSS [EM14]. Mu [Man96]. Mu-Synthesis [Man96]. Mühendislik [Yuk96]. Multi [Ald96, ABB^{+}25, Kao09, Mik23, PGBG94, She08b, AMR18, BKR97, Cap13, SR09, WGP95, ZLLT23, Ano96m]. multi-dimensional [Cap13, Ano96m]. multi-level [SR09]. Multi-Media [Ald96]. Multi-Objective [Kao09, PGBG94, WGP95]. Multi-Physics [ABB^{+}25]. multi-platform [AMR18]. multi-segment [BKR97]. multi-source [ZLLT23]. Multi-Term [Mik23]. Multi-Unidimensional [She08b]. multibody [Moo98]. multicarrier [Bin00]. multichannel [BAEBAS19]. Multicomplex [CH20, CH20]. multicomputer [DJK93]. multicomputers [JKR92]. Multiconference [Pav93, SSV95]. multicore [EL16, Kep09]. Multidimensional [JNR19, AH09, KMBP24]. Multifaceted [KAR^{+}19]. multifrontal [Dav04b]. multigrid [ZZG^{+}14]. Multilayers [DLA^{+}16]. multilinear [KW05]. Multimaterial [GV25]. MultiMATLAB [MT97, TT^{+}96, TMC^{+}99]. Multimedia [IGPF96, Man03, MSY98, Mro95b]. multimodal [CK22, CK23]. multinode [Kep09]. Multiple [KT10, LSvdV19, PR14, TMC^{+}99, LMAV24, RBD^{+}10, RBD^{+}11, TACA15, TT^{+}96]. Multiple-Phase [PR14, RBD^{+}10, RBD^{+}11]. MultipleCar [LSvdV19]. multiplicities [Zen04a, Zen04b]. Multipliers [AM22, DH97b, DH00]. multiply [Nas20]. Multiprecision [Ano19, Sch04]. Multirate [Har04, QCPG96]. Multisensor [BAEBAS19]. Multistep [AJGO^{+}20]. multitask [DP08]. Multithreaded [HYY^{+}15]. Multithreading [IMM12]. Multivariable [CKC94, SP96, Zhu01, Zin93, Coo01, SMS95, ZW93a, BFM89, Mac91]. Multivariate [CSY15a, CSY15b, Joh18, Koc14, Row03, Cur95, FR18, PC13, SN01, Wan15]. MultRoot [Zen04a]. MuPAD [Ben98b, Ben98b]. MUPen2DTool [BBN^{+}22]. Music [D^{+}95]. mutation [HBT16]. mutually [AM95]. Myex [Jon18]. Mystery [Mol93].

N [Rid95]. N-Dimensional [Rid95]. Nachrichtentechnik [KK01a]. NAClab [ZL13]. Nano [Lys05, KA13]. Nano- [Lys05]. nano-crystals [KA13]. Nanocrystals [Che18, Res19]. nanoparticles [HT12]. Nashville [D^{+}95]. NASTRAN [GSM95]. National [Ano95o, IEE94b]. native [BCG17]. native-MATLAB [BCG17]. NATO [EM94]. Natural [RJ93]. Naturwissenschaftler [Beu05]. Nauczaaie [Mro95a]. nauczaniu [BF95]. naukowo [MM96]. naukowo-technicznych [MM96]. Navigation [LRD^{+}95, BSLK01, GWA01, Rog03]. NC [SAE95a]. NDES [Ken95]. Near [BV19, CW05e]. Near-Optimal [BV19]. Nearest [BHC^{+}24]. NEC [Lu96]. negative [MSL09]. Neighbors [BHC^{+}24]. Neoadjuvant [Shu17]. neocortical [BSB20]. NES [CK23, CK22]. nested [TKD07]. nested-loop [TKD07]. Net [Ano97a]. Netherlands [KZL^{+}20, Ano95n, KF96]. NETLAB [Nab02]. Network [AM22, BL96b, BL96a, BF95, DB93a, DL96, HDB96, RD22, SS96a, Sta03a, SWG^{+}94, VA94, LWN02, TC13, TKD07, ZZC^{+}08]. Networking [ACM97, ACM98, Che04, MZ03]. Networks [BD20, BPCCM23, HPK18, Kar93, KDAB19, NRHP96, PMNWR20, SU93, VA94, VBB18, Abd94, Ach04, Ano96o, BD22, Dag94, Elg01, HTJ90, HB04, Hay99, Kec01, MBBC95, McN05, Nør00, KMLP^{+}23]. Neumann [RC16, RC13]. Neural [AM22, BD20, BL96b, BL96a, BF95, DDD97, DB93a, DL96, EA03, HDB96, Hay99, Kar93, KDAB19, LL96, McN05, NRHP96, Nør00, SS96a, SWG^{+}94, VA94, Ano96o, Dag94, HTJ90, HT97, Kec01, MBBC95, Spo02, TC13, TU97, ZZC^{+}08, CCF02, Abd94]. NeuralWorks [SWS97]. Neuro [JSM97, KNNM97, LL96]. Neuro-Fuzzy [KNNM97, JSM97, LL96]. neurobiological [EA03]. neurocomputing [HK01a]. neurones [Abd94]. neuroscience [Tra02, Tra10, Wil99]. neuroscientists [WLB^{+}09]. neutral [MSL09]. Newport [IEE95a]. News [Ano97k, Ano97j, Bra97b, Bra97a, DDK14]. Newton [Kel03, LP13, Rov90]. NGSC [GJ20]. Nicely [GG92]. Nine [Hig02a]. Nineteen [MV03]. NJ [Ano96a]. NLSEmagic [Cap13]. NMR [BBF^{+}19]. NNCTRL [Ano96n, NRPH96]. NNSYSID [Ano96o, NRHP96]. No [Ano95m, UCL04, Oga10]. Nobody [Kah98]. Nodal [MV22, RV13]. Noise [DK99, ZT02, Sas96]. Non [Gri95, PMNWR20, BAEBAS19, MC02, OBCG19, Pet96, KMLP^{+}23]. Non-Gaussian [Gri95]. non-linear [MC02, OBCG19, Pet96]. Non-Parametric [PMNWR20, KMLP^{+}23]. non-stationary [BAEBAS19]. Nonlinear [Aka00, Asi10, AY96, Cap13, DG96, GG92, Kel95, MVM97, Nat94, PR14, SO93, TC97a, TC97b, WTL00, Ano97e, AY91, Col04, FCP97, HL03a, IP23, Kel03, Ken95, MV22, PT07, Por23, Spo02, WK03, Yan99, DG96, Lak97]. Nonnegative [IKK^{+}19, LP13]. nonparametric [FJ22]. Nonsymmetric [PY22]. Nonuniform [Mar01]. Nordic [Bjø01]. Norfolk [Sin93]. Norm [GV25, JRCS95, JCRS96, Ano96k, HAM02]. Norm-Based [GV25]. Normal [Mol01, ÖK17, She08a]. Normale [Pav93]. normality [Wan15]. normally [Mol01]. Northern [Lum92]. Norwegian [Hau96a, Hau96b, Hau01, Sim99]. Notation [Cze95]. Note [Tib93]. Notes [The93c, Tib93, Mat95a]. Nov [AO95, Gra93]. Novel [AB13, HBT16, UW12, ZLMQ23]. Novell [Ano97a]. November [ACM97, ACM98, ACM03, Ano93c, Ano96f, Ano96u, Bud95, Dag94, EP 97, Gra94, IEE94c, I^{+}96, Lum91, Lum92, PJ97, SAE95a]. NPDF [Rid95]. NPTool [LP13]. Nuclear [Ano96g, BBN^{+}22]. nucleation [RDP14]. Number [BV08, CH20, CM67, Nah11, DH97a, DH97b, DH00, PR02]. Number-crunching [Nah11]. Numbers [Gou20, Mol95a, Mar68, Mol01, Tre15]. Numeric [CF91, DR96, HL95, JHPK97, The92d, The92a, The92c, The92b, The93a, The93b, The93c, The97, LP05, LM91, Mat95a]. Numerica [Com90]. Numerical [Atk05, BFG^{+}14, BTM09, Bor18, Bor97, Bra02a, Bra06, Bur05, CST20, Car92, CHMN13, CC06, CM99a, DB08, Dat95, Dat04, DLA^{+}16, Dem97, EU07, ELR02, FB03, Fau08, Gar94, Gar00, GGKM09, GS08, GKD05, Gra07, GK05, Hig96, Kar01b, KG01, KM10, KM11, Kiu05, Kiu10a, Kiu10b, Lea04, LQT18, Mat92a, Mat94a, MF99, MF04, Mol02a, Mol04a, Mol04c, Mol08, Nak96, Nak02, NGKM18, PL95, PL00, Pie05a, Rao02, Rec00, Ste10, TB97, AS12, Ano95l, AH04, Ban01, Bee05, BDL25, Cha05b, CS99, Dat13, DPR05, DGK03, DGK04, Eat92, Eat97, Eat00, Eat02, Eat05, EBH08, Fau99, GW04, Goc02, GS12, HH03, HCBAEC23, Hig02b, KG06, KG12, LVV05, LLZ18, LJ23, LW03, MF94, Moi10, Mol03a, Mol04d, OD05, Roq13, SH00, SL17, SH18, SKA19]. numerical [Sha07, Sta05a, Sta05b, WA96, WCS92, YCCM05, ZL13, Com90, Com92, FT92, Mei05, Ano96b]. numerically [Tre15]. Numerici [Com92]. Numerico [CF91, FT92, GSS05]. Numerics [Cse99, Van00b, AGL07]. Numerik [Mei05]. numerycznych [Mro95a]. Nuts [Ano97c].

O [Car10]. OAGM [BB97]. OAGM/AAPR [BB97]. Object [DP16, NPP04, Nei10, ARR02, Dav13, DRS18, EMMK01, IP23, CF16]. Object-Oriented [DP16, NPP04, Nei10, ARR02, Dav13, EMMK01]. Objective [Kao09, PGBG94, WGP95]. Objectively [Mol99a]. objectives [JR99]. Objects [MS94, MCd^{+}96]. obliczen [MM96]. observables [BDL25]. observations [FR18]. Observatory [HR96]. Observer [DP95]. Obtain [Mat94a]. occasion [dOSRS19]. ocean [ZE95]. OCT [HKF^{+}20]. OCT-A [HKF^{+}20]. OctApps [WPK^{+}18]. Octave [Eat92, Eat00, EBH08, FRAK15, Ano96p, AMR18, Bee05, BW20, CDSV10, CDSV11, CST20, DLA^{+}16, Eat97, Eat00, Eat02, Eat05, Edd00, FR18, FY18, GHH20, GV08, HTCI96, JRA^{+}18, Joh18, MBR21, PC13, PSR16, QS06, Qua10, QSG14, Rab20, RAW^{+}16, RHR^{+}21, RS15b, SS18, SM14, Váz16, WPK^{+}18, Wet20, Bar16a]. Octave/MATLAB [CST20, Joh18, DLA^{+}16]. Octave/MATLAB(R) [SS18]. OCTBEC [Hoh14]. October [Ano93e, Ano95d, Ano95s, Bjø01, IEE94b, IEE97c, IEE97a, IEE97e, dOSRS19, USE94, USE99]. Ocular [KAR^{+}19]. ODE [Ano97j, ANV00, BGPPRW14, Bra97a, Rei93, SR95, SR97, SR02]. ODE45 [Faz10]. ODEs [Bir18, Bün20, DGK03, DGK04, Mol96b, SGT03, Sha09]. oFEM [DRS18]. off [Ano95j, LNLB19, RB05]. off-highway [Ano95j]. Offers [Stu96]. Ogata [Ned95]. Ogive [She08a]. ogolnopolskiej [Ano94g]. Ohio [H^{+}96]. Old [Ano95d, YKS94]. Older [RD22]. On-Line [Bee94]. on-stack [LH13]. Onditas [DACV95]. One [Lam95, NGKM18, BV21, OAKS11, RS15a]. One-Dimensional [NGKM18, BV21]. one-way [OAKS11]. onto [Nas20]. OOP [Wei12]. OOPS [Mol99a]. Open [LO16, SH18, SM14, TCD^{+}22, BGGL20, Dat13, PNL^{+}21, Ren17, Ste08, Ste13, TACA15, VHM17]. Open-Source [TCD^{+}22, SM14, BGGL20, TACA15, VHM17]. OpenCL [RBC20]. OpenGJK [MP18]. OpenMP [Ano97j, Bra97a]. OpenPIV [BGGL20]. OpenPIV-MATLAB [BGGL20]. OpenQSEI [SBL19]. Operating [Gab98]. Operation [M^{+}94]. Operational [Lin05, BC12]. operations [Ben02]. Operativa [CSV94]. Operative [CSV94]. Operator [WR16, WR17, FRAK15, RAW^{+}16]. Operators [BT04, Ves94]. Optical [Ano98c, Wil96b, Bal19, Deg20, HTNFBS06a, HTNFBS06b, Luc05, PB01, TNBSF04]. Optics [DLA^{+}16, Möl07, Wil96b, Boo98, DM96, Tót08]. Optimal [ATvBB17, AFL^{+}12, Ano96q, AY97, BPCCM23, BV19, CST20, Kao09, LS95, Nai03, PR14, SD91, TQ96, Bry02, Bur99, CI01, DCF95, Gri94, Hoh14, Jon95b, Jon95a, RBD^{+}10, RBD^{+}11, Rob96, SH18]. Optimally [Mol99b]. optimisation [Alt95, LWN02]. Optimization [EM14, GV25, ICS^{+}18, IKK^{+}19, MBS15, MW93, PY22, Qur01, SP22, TCD^{+}22, But11, CÇ01, CFF^{+}91, Cur95, DCF95, DP08, DG96, FB95b, HL03b, HAM02, IZBT21, JB03, Kel99, MC97, MA95, PF10, Pas04, Por23, SZM^{+}14, Spa03, SZCP21, Stu99, Ven02, Mol99b]. optimization-based [Por23]. Optimizations [HYY^{+}15]. Optimized [HWB15, MGW99]. Optimizing [FSZD20, LH14, LGML05, Wei12]. Optimum [Van02]. Option [Chr97, Hig02a, CT97, Hig04]. oracle [LNLB19]. Orange [ACM98]. oraz [BF95]. Orbital [Cur05]. Orbits [DGKF12, NGKM18, GGKM09]. Order [AJGO^{+}20, BW20, Bir18, JRCS95, JCRS96, NP93, SF14, Wat93, AS12, Ano96k, BC05, Cam06, Cap13, FGMS21, HCBAEC23, IP23, Mol02b]. ordering [ADD04, DGLN04a, DGLN04b]. Ordinal [JA99]. Ordinary [LS04, Pol95, PA99, PA04, Sch14, CBL24, Mol96b, Mol03a, Sta05b, BD95, BD00]. Oregon [Lum91, ICS96]. orientated [RL00]. Oriented [CF16, DP16, NPP04, Nei10, OBD23, PH94, ARR02, Bät20, Beu05, Dav13, DRS18, EMMK01, Hsu99, IP23]. Origins [Mol04b]. Orlando [ACM98, Ano97i, IEE94f, IEE97a, W^{+}97]. Orleans [IEE97c]. ORT [GJ20]. Orthogonal [DMB17, Gau05, Gau06, Gau16b, Gau15, LMAV24, Tow16]. Osborn [Ano00]. OSCAR [Deg20]. Oscillations [MA96b, Ano96l, MSL09, Rog00]. oscillatory [Sha11, Sha12]. Oscilloscope [Rul02]. Oslo [Bjø01]. Ostend [FB95a]. Other [But92, M^{+}94, Mac05, Mac91, Nah00, Nah02]. Outcomes [RSW15]. Outlier [NK17]. OutlierLib [NK17]. Output [FS17, Nel17, AM95, DDK14, SZM^{+}14]. overloaded [For06, PWR13]. Overloading [WR16, WR17]. Overview [RHR^{+}21, Dav04a]. Ox [KS97].

P [Bar16b, Veh07, Lam95]. P-Method [Lam95]. P.D.E. [Mat94a]. P2Q2Iso2D [BCH06]. PA [EP 97, Bry96, Ham93]. Pacific [Ano95d, IEE93a, IEE93a]. pack [App19, Gre16]. Package [ADW23, BA94, Bou97, DL96, Esm14, FSO93, GCM25, Han92, HPK18, IoG10, KAR^{+}19, LLW^{+}23, LZ15, LSvdVK09, PY22, RSW15, SBL^{+}10, She08b, Ste09, SM07, SU93, XSS20, ZJKS23, BSW07, BO19, BAEBAS19, BEK99, Cam06, Cob21, CFR19, CI01, Dav05, Deg20, DdAF^{+}20, DPE96, DGK03, DGK04, DDK14, DRS18, Edl04, FMS25, GHN19, GS07, Han94, HTNFBS06a, HTNFBS06b, HSH12, HKF^{+}20, JSB20, LMAV24, LVV05, LLZ18, NSXZ14, OG95, PT07, SL17, SH18, SKA19, SN01, SFPO94, SBL19, TTT99, Tót08, Tri08, Wan15, Zen04a, Zen04b]. Packages [Ano97k, Bra97b, Jon95b, Jon95a]. packing [BC22]. Padua [CDSV10, CDSV11]. Padua2DM [CDSV10, CDSV11]. PageRank [Mol02b]. pages [Ano95c, Ano96b, Ano99a, Ano99b, Ano99c, Ano00, Ano95c]. pakietu [Mro95a]. Palace [IEE94d]. Pan [IEE93a]. Panel [ÁBZ17]. Pao [Tay99]. Paper [CKC94, Ano93b]. paperback [Res19]. Papers [AO95, Cse99, Bar92]. para [DACV95]. parabolic [OAKS11]. paradigm [AI21]. paradigms [SS10]. PARADISE [Ano96r]. Parallel [BKGS02, BK07b, CGRvD15, Cho02, HLP96, HI97, IEE96c, IMM12, Kad97, Kep05, Kep09, Lus09, MBS15, MT97, MMPOQ^{+}25, Mor98, PHL95, dVSWAL17, SM09, Sin93, AMR18, HSR01, H^{+}96, Mol95d, PPD95]. Parallelism [LPD^{+}20, RHB96]. parallelization [SZCP21]. Parallelized [MJ01]. Parameter [ATB05, CBL24, Cze95, vdH05, KK93, Rao11]. parameters [SN01, SZM^{+}14, YHC^{+}22]. Parametric [BCK96, PMNWR20, Ano96r, KMLP^{+}23]. Parchar [Mar16]. Park [IEE94b]. Parker [WD98]. Parser [AFOP19]. Part [KZL^{+}20, CW05b, Elg01, FRAK15, JRA^{+}18, RAW^{+}16]. Part2Track [JSB20]. Partial [Goc02, PSZ08, Sch12, Sch14, Col05, Coo98, Duf04, GS12, LS04, LC09, PT07, SG09, Sta05b]. Particle [TS14, TS21, BGGL20, CATK11, UW12, ZSW^{+}17, JSB20]. Particles [Mar16]. Partitioning [UÇA10]. Passive [Joh95]. past [KF96]. Pasta [DPR05]. Pastes [MLAS21]. Path [NSXZ14]. pathogen [BBHP^{+}23]. pathway [LYH^{+}16]. PaToH [UÇA10]. Patrick [Ano99a]. Pattern [BB97, DHS01, KPN^{+}04, SWG^{+}94, Dav04b, MC97, Nab02, STC04, BB97, SYTD04]. Patterns [Spe95, Spe96, JS24, NE05, TYL^{+}16]. PC [MLB87b, SR90, ZW93b]. PC-Matlab [MLB87b, SR90, ZW93b]. PCs [AatupcA03]. PD [AFL^{+}12]. PDE [JH97, PPS06, PF07]. PDE-s [PPS06]. PDEs [FMS25, IP23, Sha05]. pedagogical [DPR05]. penalized [LYH^{+}16, McI16]. Penalties [McI16]. PENalty [BBF^{+}19]. Pencils [ST12]. Pendulum [SDPM04]. penetrating [XZL24]. Pennsylvania [ACM96, Ham96]. Penny [Ano96b]. Pentium [Mol95a]. Pentiums [Ano97c]. Performance [ACM97, ACM98, Ans14, ASP^{+}18, BRS94, CB98, Esm14, FJSD96, FSZD20, HL95, IEE98, Kum14, Mar95b, The97, MT97, TtsT04, CB99, KH14, LM91, The92d, The92a, The92c, The92b, The93a, The93b, The93c, Mat95a]. perm mateda [ICS^{+}18]. Permutation [ICS^{+}18]. Permutation-based [ICS^{+}18]. Personal [The92f]. Perspective [Mol15, BC04, Paa01, SBC04]. perspectives [DPR05]. perturbation [AS12]. Peters [Ano95c]. PETOOL [OAKS11]. Petrological [RG21]. pH [BSS^{+}23, LHW01]. pharmacodynamic [VHM17]. pharmacokinetic [VHM17]. pharmacokinetic/pharmacodynamic [VHM17]. Phase [Ega98, PR14, RD22, Wat93, CS12, Ega00, LJR93, RBD^{+}10, RBD^{+}11]. Phase-lock [Ega98]. Phase-Locking [RD22]. phase-type [CS12]. Phasor [DL95, DL01]. PHClab [GV08]. PHCpack [GV08]. Phenomena [BPB96, Tho00, TYK04]. Phenomenon [Ano95u]. Philadelphia [ACM96]. Phipps [Lip07]. Phoenix [ACM03]. Phoneme [SWS97]. Photoelectron [Fis19]. photogrammetry [MBM01]. photomechanics [Asu02]. Photonics [Iro15, Mun13, War13]. Photovoltaic [AB13, CBhM12]. PHY [GJ20]. PHY-NGSC-Based [GJ20]. Phys [RC16]. Physical [BDM17, BDM18, Bro94, ZLW^{+}19, CPUARC20, DM96, DH97a, Hor98, XWZ^{+}22, XWZ^{+}23]. Physics [ABB^{+}25, CV96, CC02, Gar94, RS15a, ANM01, B10, Gar00, MN03, Nah11]. physikalische [Bro94]. PhySim [CPUARC20]. PhySim-11p [CPUARC20]. Physiological [Kho00, Tho94, WK03]. PhysioNet [SM14]. Physique [CV96]. PID [BA94, Kir93, Per93]. Piecewise [CMR17, HC92, PPT10, KW05]. Piecewise-smooth [PPT10]. Pioneer [Hai08]. Pipeline [RD22]. Pitaevskii [AD14, AD15, CR13, CR20]. Pittsburgh [Bry96, EP 97, Ham93, Ham96]. PIVlab [TS14, TS21]. PK [AFL^{+}12]. PK/PD [AFL^{+}12]. PkStaMp [AFL^{+}12]. Planar [MMPOQ^{+}25]. Plane [OBD23, Spe95, Spe96, Wat93, JZW^{+}22]. Plane-Wave [Spe95, Spe96, JZW^{+}22]. planes [Mar68]. planning [CLTS20, LWN02]. plant [Bre03, LSH95]. plasmonic [HT12]. plates [Roq13]. Platform [AJGO^{+}20, BKGS02, Lu96, RLV11, AMR18, BO19, LHZT23, RDP14]. Platforms [Has12]. play [CK22, CK23]. PlgCirMap [Nas20]. plotting [Mor18]. plug [CK22, CK23]. plug-and-play [CK22, CK23]. plus [Ano97a]. pMATLAB [BK07b, MBB^{+}09]. Pneumatic [UB95].

Point [CHMN13, Gou20, HP19, Mik23, MON12, Mol17, NN94, Qur01, Ste09, TQ96, Ano96q, BPS99, KK93, Meu20, Mol98b, Rob94, RB04, RLV11, Yan17, Zha98]. Points [OBD23, CDSV10, CDSV11, Mol96a]. Poisson [RC16, RC13, YJ24]. Poland [Bub95, MO95, Mro95b]. Polarization [SS18]. polarized [HA95]. Polioptymalizacja [Ano94g]. Polish [Brz97, MM96]. Polmat [AH09]. polygonal [Nas20]. Polynomial [IKK^{+}19, NN94, BNN16, Gri94, Jon01, KS94, Zen04a, Zen04b, DP16]. Polynomials [DP16, Gau15, Tow16, Gau05, Gau06, Gau16b, HL03b, LMAV24]. Polytopal [ABB^{+}25]. popbio [SM07]. Popular [Has12, ÖK17]. Popular-but-Seemingly-Dissimilar [Has12]. Population [AFL^{+}12, Ber03]. porous [AGL07, GMT96]. port [Bai05]. Portable [Bee17]. portfolio [Kat09]. Portland [ICS96, Lum91]. Porto [dOSRS19]. Portopia [IEE96e]. Portraits [Wat93]. Portugal [dOSRS19]. Posed [Han92, Han94]. positioning [GWA01]. positive [MSL09]. Post [IZBT21, Lu96, MBBC95]. post- [MBBC95]. Post-Processing [IZBT21, Lu96]. Posterior [RD22]. Posterior-to- [RD22]. posteriori [BC05]. postgraduate [Mun13]. postprint [BS95a, FB95a]. postprocessing [Sar10]. pour [MM10]. Power [AM01, BRPCR94, FSO93, IEE94a, IEE96a, IEE04, JE94, KA09, KYN95, LSH95, MASV96, Rog00, SO93, TQ96, VVM93, Ach04, Ajj95, Ano93a, Ano96q, BH96, BDR04, Bos01, Bos02, CSB03, Cha02a, Cse92, DPE96, Dja98, Dja00, PSP04, UW12, Van01a, Ano95n, Ano96s]. Power-gen [Ano95n]. powerful [WW99]. powerplant [Ano95j]. Pp [Mun12, Ano93b, How15, Mar19, Res19]. PQser [CFR19]. Practical [AC96, Bäc95, Bäc00, HH04, Ife05, Lai04, SP91, AH05, Att09, Att12, Bau02, CSV96, DM96, Din02, Nah08, OG95, Ros03, SB10, ZM00, Bäc97]. practice [Ano96f, GA01, Mar01, PJ97, Qua02]. practices [PPL^{+}18]. practicing [Nas01]. practising [Mun13]. practitioner [Nør00]. Prague [SSV95]. Praktisk [Bäc97]. Prandtl [PR02]. PRBS [Alt95]. PRBS-signals [Alt95]. Pre [Lu96]. Pre- [Lu96]. precise [HTNFBS06a, HTNFBS06b]. Precision [CCF96, HP19, HPZ19, Mol17, dVSWAL17, Rum17, Rum23, Ste09, Meu20, TNBSF04]. Precision- [Rum17, Rum23]. Predator [GBM15, Gar07]. predator-prey [Gar07]. prediction [ZLMQ23, ZLLT23]. Predictive [CKC94, DDD97, Mac02, SCC95, McN05, Ros03]. predictors [SMS95]. preface [Nah02]. preliminaries [Sta05a, Pie05a]. PREMER [VBB18]. Prentice [Ano96a]. Prepackaged [Fos91]. PrepAnnECG [Ozk23]. preprints [Ano95d]. preprocessing [Ozk23]. presentacion [DACV95]. presentation [DACV95]. presented [Cse99, DG96]. Preservice [Bou97]. preserving [LCMCD22]. Press [Ano99a, How15, Mar19, Mun12, Tay99]. Prey [GBM15, Gar07]. price [ZLMQ23]. Pricing [Ano05, Chr97, Pri00, CT97]. primal [MPVA25]. Primer [Rou98, Sig92, Sig94, ST98, Asu02, BGG98, Car99, DS05, Dav11, JD13, PR06, SD02]. prin [Bla02a]. principle [SSW09]. Principles [BSL93, Ben02, Cou95, HK01a, Riz00, Tay94, Tra04, ZT02, BN10, BN11, Jac04, Pee01, PSP04, Qua02]. print [Mol04a]. Prioritization [MNBB19]. Pro [AH05, LM91, The90, MLB87c, ML90, MLB87a]. PRO-MATLAB [LM91, The90, MLB87c, ML90, MLB87a]. Probabilistic [CM99b, HM19, IEE04]. Probability [Chi97, Kay05, Li99, MC04, Pee01, Pfe95, Rid95, Rus08, SW94, SWrpSetfe02, Vin98, Wil03, YG99, AC97, Dav12, FJ22, Kay06, Nah00, Nah02, Nah08, Pac04, Zie97, Beu05]. Problem [BKL19, CS99, Ett93, GG92, PBB22, Ano95m, Att09, Att12, EPJ^{+}05, Ett97, GI06, LP05, RC13, RC16, Sti04]. problem-solving [GI06]. problema [EGE95]. problemas [de 05]. Problems [ABB^{+}25, BD95, BD00, CHMN13, Con95a, EP96b, EP04a, Fab97, FC95, GH93, GH95, GH97, Han92, HI97, ICS^{+}18, IKK^{+}19, MBS15, Ned95, Oga94b, PR14, Pfe95, Riz17, Ste96, AS12, Ano96m, ATB05, BD05, Bro95, DCF95, EP00a, FR96, FC00, GH04, GHN19, Gro99, HMT13, Han94, Hen07, KA02, KK93, Mil20, Mol99b, MY05, Nah08, Nah11, Rad96, RBD^{+}10, RBD^{+}11, Som07, TACA15, Vog02, WCS92]. procedures [Cou93, SR09]. Proceedings [ACM97, ACM98, Ano95s, Cou93, EM94, Glo98, Hor97, IEE93a, KZL^{+}20, Mic94b, Pav93, Zem97, ACM96, Ano94i, Ano95k, Ano95o, BEH94, Bjø01, BLM97, BH95, Bud95, BB97, Dag94, EP 97, Gra94, Ham96, IEE94c, IEE96f, I^{+}96, JE91, Ken95, Kin93, Mal96, MGC94, Ter94, Van92b, Ano93e, Ano94h, Ano96e, Ano97h, Ano97b, BCD^{+}20, BM09, Bub95, C^{+}96, D^{+}95, IEE93b, IEE94e, IEE94d, IEE94f, IEE95b, IEE96e, IEE96d, IEE96c, IEE96b, IEE97e, IEE97d, IEE98, ICS96, Jef08, KPN^{+}04, Lum91, Lum92, NP96, Sin93, USE94, USE99, W^{+}97, WD98, H^{+}96, PJ97, dOSRS19]. Process [Alt95, Beq98, Beq03, Cha02d, Cla03, Kin93, LNC98, Mar95b, MCd^{+}96, Ram94, SEM04a, Del02, Ram97, RL00, SBL^{+}17, TJML92, TKD07, Wea97, Zhu01]. process-orientated [RL00]. Processes [Gri95, Kao97, Kay05, Rus08, SW94, AC97, CS12, Chi97, Cse92, JD13, Kay06, Kro95, MC04, Vin98, Wil03, YG99]. Processing [AM22, Amb95, Ano93e, Ano95s, Ano95q, BAEBAS19, Bro95, B^{+}94, Can86, DS16, EO94, Fab95, FY18, How95, IEE94b, IEE96f, IEE96c, IZBT21, Jac89, Jac96, Kad97, KSL93, Kub95, Lei11, LS88, LS93, Mar16, Men95, MB94, MLAS21, Orf96, SS18, Sch94, SM14, Sin93, SC97, The92h, Wer03, von93, Aka00, Amb99, Ano95t, Ano97l, BC06, Boo98, Bru01, Cav00, Cha05c, Chi00, CiA02, Den98, DDN02, EA04, ElA04, GI06, GW02, GWE04, Gop10, Har04, Hay96a, IP97, IP00, IP10, Kal97, Kay93, Kha05, Kum05, KL01, Lai04, Lei02, Lu96, LTE01, MIK00, Mat97, McA04, McC98, MSY03, Mea02, MC04, Mit98, Mit99, Mit06, MS00, MGT03, NM99, NA02, Paa01, PB01, Por97, Qua02, Qur05, SH05, Sch11, Sem04b, SB10, Sou99, SWrpSetfe02, SD96]. processing [Ste03, SH11, TB95, Teo98, VLV00, Van02, WOZ02, WWM06, Wel93, Wri97, XZL24, ZI04]. Processor [Gid95, HSR01, MBBC95]. Processors [LPD^{+}17, TMC^{+}99, KG05, PAG11, TT^{+}96]. product [PV99, Rib00, VC06a]. Productivity [MBB^{+}09]. Products [Ano97k, Bra97b, Ano97j, Bra97a, RKZ^{+}14, VC06b]. professionals [Ozk23]. Professor [Mol06a]. Program [ATvBB17, EI05, GR97, Hon91, KSF94, MA96a, MA97a, MA97b, MLF^{+}12, RG21, ST12, VPM16, Are94, Asi10, BCR03, BQOvdG05, BSS^{+}23, CATK11, Dat13, Hon92, Liu23, MY05, RDP14, Sha08b, SZM^{+}14, TNBSF04, Tót08, UW12, ZSW^{+}17, Bjø01]. Programmer [Wea97]. Programmi [Bol94]. Programmierbeispiele [GG04]. Programming [Bee17, CA97, CNJ97, Edw09, GG04, Her01, IMM12, Kep05, Lus09, MSS^{+}19, Nei10, NN94, PR14, Alt12, AH05, Ano95l, Att09, Att12, AC99, Bai05, BY08, BPS99, Cas14, Cha00, Cha02b, Cha04, FMW07, Kap04, KS97, Kun04, LLLW06, Luo95, MKU22, MF94, MA95, OD05, dOSRS19, Tho04, TTT99, VFG04, Ven02, Yan17]. Programs [Bol94, BV08, Góm15, MGW99, Pao99, RHB96, SB90, AY91, DP99, FCP97, HH03, HLS08, JR99, Kah02a, Mir96, Pao01, PAG11, Tay99, Tre86, TKD07, WB16, Zha98]. Progress [Ter94, Mol04d, Ano96t]. project [Smi97, ZLLT23, ES10, Tor02]. Projection [LZ15, LP13]. projectowania [Ano94g]. Projects [EP96a, GP96, DJKP07]. projektowaniu [Szy93]. Propagation [Ano98c, BP96, BPB96, IEE95a, GMT96, OAKS11, RCT20, Vit11]. properties [YHC^{+}22]. property [Are94]. propulsion [Are94]. protocols [Che04]. Prototyping [CGM95, BK06]. Provably [LJ23]. proving [GRH^{+}21, JS24]. Proximity [Lip07, HAM06]. przedmiotow [BF95]. Przyklad [BF95]. PSE [KS01]. pseudorandom [DH97b, DH00]. Pseudospectra [BKGS02, BKG05]. pseudospectral [RBD^{+}10, RBD^{+}11]. PSI [PPD95]. PSPICE [Att02, Att10, Sim99, AY94, DIN06, Irw05]. Psychophysi [NK18]. Publication [CW05e]. Publications [Bee94]. Pulse [BP96, BPB96, LJR93]. puzzlers [Nah00, Nah02]. PVS [BDM17, BDM18]. PVS-Simulink [BDM18]. PWM [LJC93]. PWS [Ano93b]. PyBanshee [KMLP^{+}23]. pyelonephritis [HGM^{+}25]. Python [CLMM20, Kiu10b, KMLP^{+}23, MSS^{+}19, TCE21, Wet20].

Q6965 [UCL04]. QDT [KMBP24]. QFT [AW92, Nat94, SFK91]. QFT-Based [Nat94]. QGLAB [GCM25]. qls [SR09]. QPECgen [JR99]. QR [Mat92c, Mat94b, Mol95c]. QSPR [HG24, HGM^{+}25]. Quadratic [DAC95, Jiř97, HMT13, JR99, Jon95b, Jon95a, Luo95]. Quadrature [DMB17, Gau16a, Joh18, PR14, Esp07, Gau06, HS24, LMAV24, Sha08b, Sha08c]. Quadruple [Mol17]. quadrupoles [Mai00]. qualities [How91]. Quality [AM22, Del02]. Quality-Efficient [AM22]. QuantDiffForecast [CBL24].

Quantification [How15, SY20, SS15, PT24]. quantified [CBL24]. Quantitative [Yan99, AW97b, KF96]. quantization [RB05]. Quantum [Che18, GCM25, CZ17, Dat13, Hoh14, KMBP24, Lev03, Res19, SL17, SH18, SKA19, SZM^{+}14, Tót08, SKA19]. Quantum-classical [SKA19]. Quarteroni [Sha04]. quartic [Jiř97]. Quasi [BMR19, RS08, SBL19]. Quasi-Least [RS08]. Quasi-Toeplitz [BMR19]. quasiconformal [AVV97]. quasimatrix [Tre10]. QUBIT4MATLAB [Tót08]. Québec [IEE94a, OB93]. queries [MP18]. Queuing [IMM12]. Quick [Pra96, Mat95b, Pra99, Pra02, Pra06]. QuickBasic [Tay99, Pao99, Pao01].

R [Alv11, Ano93b, Ano96a, Ano99d, Ano00, Ano12, Ben95b, Bow10, Bur10, Car10, Han06, Jai09, Jer06, KH96, Kus02, Kus06, Mar14, Mar19, Ni22, Res19, Wie94, Wik04, EK23, LHW01, Mar07, RHG09, dVSWAL17, Sch14, SM07, VFV13, Veh07]. R. [Ano09]. R/MATLAB [VFV13]. r2009 [MM10]. Radar [LM04, Mah00, Mah05, Wei12, ME04, Pac04, STF01, Sou99, XZL24]. Radial [BPCCM23, Sar17]. Radiative [SF14]. Radio [LWN02, DG04, HP02a, JS04, Nah01]. Radionuclide [RT96]. radiowave [OAKS11]. Radisson [Bjø01]. rail [Chu00]. Railway [M^{+}94]. Railways [Ano95i, M^{+}94]. Rainey [Lum92]. Ramsay [Bur10, Car10]. Random [CM67, Gou20, Her96, Jaf00, JD13, Kay05, Mar68, Mol95b, New94, Rus08, SW94, The92h, AC97, BC22, BH97, Chi97, DH97a, Gri95, JDFV08, Kay06, Li99, MT00, Mil20, MC04, Mol01, Pee01, Vin98, Wil03, Wit04]. Randomized [RSW15]. Range [Gau16a, STF01, VC06a]. Range-Doppler [STF01]. rank [BPS05, FHH99, LLZ18]. rank-revealing [FHH99]. ranking [Cas14].

RankRev [LLZ18]. Rapid [CGM95, OBCG19, SS18, Kat09, YHC^{+}22]. RARtool [RSW15]. rate [BlD^{+}20]. Rational [DMB17]. Raton [Ano99a, Mar19]. Raviart [BC05]. ray [ANM01, Fis19, LLLW06]. rays [SAKG15]. RC [Ano95p, Con95a, Con96]. Rcall [EK23]. RCR [Nel17]. re [I^{+}96]. re-engineering [I^{+}96]. Reaction [Löw01, Fog99, Gar07]. reaction-diffusion [Gar07]. read [DDK14, SBL^{+}17]. Reaktionstechnik [Löw01]. Real [AKB94, KL01, POVD96, WM95, WM96, WWM06, XSS20, ZJKS23, ARR02, Bat99, Bat19, Bra02b, FB95a, Jan02, Li96, LHZT23, Zal96]. Real-space [XSS20, ZJKS23]. Real-Time [AKB94, POVD96, WM95, WM96, KL01, WWM06, ARR02, FB95a, Jan02, Li96, LHZT23, Zal96]. realistic [Deg20, FR96]. Realization [ÇA10]. realized [Rob94]. recapture [Cat95]. receivers [Pac00]. recipes [TGM06, TG10]. Recognition [Kar93, KPN^{+}04, SWG^{+}94, BB97, Kin97, MC97]. recognitions [Nab02]. reconfigurable [BSC^{+}00]. reconnections [CZ17]. reconstruction [FBFB04, HSH12]. record [IEE97c, IEE97b]. Recovery [MLAS21, Cat95]. Redfern [Ano99b, Ano99c]. REDUCE [PR02]. reduced [BPS05]. reduced-rank [BPS05]. Reduction [BW20, SH18]. Reference [The92c, EaoGOBHW14, HL96, HL98, HL01, Mat95b, SCB99]. reference-H [SCB99]. Refinement [CCF02]. Reformulation [Bir18]. Reg [UCL04]. Regelungsstrukturen [Bro94]. Regelungstechnik [Bod98]. Regency [Ano97h, IEE97c, IEE97b]. Region [EM14, BBEM22, RSS08, SZM^{+}14]. Regression [BV19, Edl04, WTL00]. regulacji [Szy93]. Regular [ST12]. Regularization [Han92, Han07, GHN19, MR11, RSS08, Han94, Han99]. related [BO19, BM94, WCS92, YHC^{+}22]. relation [Mac91]. relational [PK04]. relations [Luc05]. Relationship [Zla17]. Relationships [WTL00]. Relaxation [IKK^{+}19, BBF^{+}19, BBN^{+}22]. Relay [MA96b, Ano96l]. Release [The93c, Mat95a]. releases [Ano98b]. Relevant [Thi95]. Reliability [Cha95]. Reliable [Cse99, Wei12, EH07, Esp07, MP18]. rely [Mol02a]. Remark [Zag16]. Remarks [ACF99]. Remote [SDPM04, Shu17, JE91]. remotely [Rid95]. Renamer [FM24]. renaming [FM24]. Repeated [ATvBB17, FS17, Gau16a, Nel17]. replacement [LH13]. Replicated [Nel17]. Report [Nel17, Li96]. Representation [BB96, Lip07, EA03, HAM06, TKD07]. representations [SL17]. Representing [BQOvdG05]. Republic [PJ97, SSV95]. required [BLL^{+}15]. Research [EM94, Li96, Luc05]. Researchers [Lei11, MM10]. réseaux [Abd94]. Reshaping [WB12]. residual [KS01]. resilience [BD22]. ResMapper [BD22]. résolus [Riv01]. Resolution [AS15]. resolved [JSB20]. resonance [BBN^{+}22, HSM04, SGLBA98]. resonant [BH96]. resonators [Bal19]. Resort [IEE94d, W^{+}97, WD98]. resource [Bor97, SBC04]. Resources [Ano02b, CdFCS98, Fos01, RB05]. respect [Gau15]. respiratory [Fis95, NMS^{+}06]. Response [CW05a, HHF95, RSW15, WTL00, Web97]. Response-Adaptive [RSW15]. Responses [Shu17]. responsiveness [AP02]. Restarted [Rad96]. restoration [LP13]. Restructuring [DGGM96]. resueltos [de 05]. resultant [WEM98]. resultant-based [WEM98]. Results [Nel17, CF16, GMT96, LS98]. Retargetable [LPD^{+}20]. reusable [MKU22]. Reusing [FS17, Nel17]. revealing [FHH99]. Reveals [RD22]. Reverse [BGGT21]. Review [Alv11, Ano93b, Ano95c, Ano95m, Ano96a, Ano96b, Ano99a, Ano99b, Ano99c, Ano00, Ano09, Ano12, App19, Bar16a, Bar16b, Bow10, Bur10, Car10, Epp11, GR93, Han06, How15, Jai09, Jer06, Kus02, Kus06, Lip07, Mar14, Mar19, Mic00, Mun12, Myr17, Ni22, O’B13, Pie05a, Res19, RS15a, Sha04, Tay99, Ter94, Veh07, Wik04, Ano96t, DDW93, Kwo93, Shu01, Tow16, Ver09]. Reviews [Ano97j, Bra97a, CW05c, CW05e, Ste96, Mir96]. Revisiting [Mol98b]. RF [LB00]. Riccati [PIAH12]. Richard [Epp11]. Riemann [FK11]. right [Ves94]. Rigid [Gab98, BKR97]. Rigorous [FBFB04, Tay95]. ring [Cat95, Liu23]. ring-recovery [Cat95]. Ripples [JL01]. Risk [Meu05, Dow02a, Dow02b, PF10]. Ritz [IEE96d]. Ritz-Carlton [IEE96d]. River [Ano96a]. Rivercenter [IEE93b]. Rivières [IEE94a]. RL [Ano95p, Con95a, Con96]. RLC [Con95b]. RMatlab [VFV13]. RMatlab-app2web [VFV13]. Road [Can24, BD22]. Robbin [Ano95c]. Robot [Ano95o, SS96b, SS96c, SU94, Ano96d, ASA96, GB89, Kog97, Rob02, SS00]. Robotic [UB95]. Robotics [Cor95, Cor96, Hon92, JE91, Kin93, LQ94, Ano94a, Hon91, JE91]. Robots [Ano95o]. Robust [AFOP19, Ano96r, BCK96, EH07, GMT95, Gre94, Gri94, GPK05, IS96, LB99, LHW01, Mac04, Mag02, OBD23, PFG08, Col04, Hen07, VH04, ZD98]. Robustness [WD96]. rocking [Por23]. ROF [Kum14]. Rogel [Bar16a]. Rogel-Salazar [Bar16a]. Role [Car92, Wie94]. ROM [RS15a, PK04]. Ronald [Ano00]. Root [OF92]. RooTri [OBD23]. roots [Zen04a, Zen04b]. Roppongi [IEE96b]. ROS [MMPOQ^{+}25]. ROS/Gazebo [MMPOQ^{+}25]. Rosemont [Lum92]. Rostock [Glo98]. rotating [JAC20]. rotational [CATK11]. Rotations [LSvdVK09]. Rouge [NP96]. Rounding [HM19]. Roundoff [DK99]. routine [LS98]. Routines [BB96, RVV^{+}92, VZDCM25, LS04]. Royal [Van92b]. Rubens [How15]. rule [Men01, Sto13]. rule-based [Men01]. Rules [DMB17, Fis19, KW95, LMAV24]. Run [GJ20]. Runge [CMR17, Cam06, EH07]. running [OG95]. Runtime [HWB15]. rwl read [Liu23].

S [Bur10, Iro15, Mun13, AM17, FGCG94, PPS06]. S-domain [AM17]. S. [Car10]. S02YSCODE [CFGG94]. Sá [Veh07]. Saddle [Ano96a]. Sage [ES10, Ste13]. Salazar [Bar16a]. Salem [SAE95a]. Saleri [Sha04]. Sales [JE94]. Salt [I^{+}96]. Sampaio [How15]. Samples [Zla17]. sampling [MT84, Mar01]. San [ACM97, GCP97, Gra94, IEE93b, IGPF96, Joh95, Kin93]. Santa [Ano93e, USE94, WD98]. SAS [Bjø01]. satellite [Ali02]. saver [Mir97]. SC-SR1 [BBEM22]. SC2003 [ACM03]. SC97 [ACM97, ACM97]. SC98 [ACM98, ACM98]. SCADE [CCM^{+}03]. SCADE/lustre [CCM^{+}03]. Scalability [Ans14]. ScalaLab [PMTL14]. ScaLAPACK [QMS98, RHB96]. Scalar [OBD23]. Scale [BO19, EM14, BD22, GHN19, Mol99b, Rad96, RSS08, Set05, Zha98]. scaled [YJ24]. scaling [HAM02]. SCAN [Cse99]. SCAN-98 [Cse99]. Scandinavia [Bjø01]. Scattering [BKL19, TKD00, TACA15, VCK98]. Scenario [Kum14]. Scheman [C^{+}96]. Scheme [GJ20]. Schemes [MP17, Cap13, Gar07]. Scholes [Chr97, CT97]. School [DG96, Hor97, Mic94b, MO95]. Schrödinger [BV21, Cap13, LVV05, Tri08]. Schur [Bra02b]. Schwarz [Dri96, Dri05]. Science [Ano95v, Bub95, IEE96b, KZL^{+}20, Mat92a, Web97, Dav12, HK01a, Kar01a, Mac00, Nah01, Nah11, SSTD03, Tót08, VF10, WR04]. Sciences [Cha17, O’B13, RJ93, HP02b, Mid00, TGM06, TG10, Vid11, Com92, JLM96]. Scientific [BF97, Cla03, Cse99, DB08, DG96, FJSD96, GH93, GH95, GH97, Hea97, Law05, Lys03, MM96, QS03, QS06, Qua10, QSG14, Sin93, Ste96, Ano95l, DJKP07, GH04, Kap04, LP05, Lea04, LW03, PMTL14, Tur00, Tur01, Van97, Van00a, Ano97j, Bra97a, Sha04]. Scientists [Ett96, GP96, HV07, Mic00, Pra96, Rao02, RG21, Beu05, Bor97, Cha05b, GS08, Hah97, Hah02, Har05, Jef05, Mun13, Nir02, Pra99, Pra02, Pra06]. Scienze [Com92]. SCOM [CI01]. Scope [Iro15, Mun13, RS15a]. Scotland [IEE94e]. SCOUT [SY20]. script [HL22]. Scripting [WW99, PMTL14]. scripts [CF16]. SCTBEM [YJ24]. SDELab [GS07]. SDPHA [BPS99]. SDPT3 [TTT99]. SDS [FCP97]. Sea [Zla17, Boo04]. Search [Esm14, Spa03, ZC08]. search-based [ZC08]. SEC2MWD [MMD24]. Second [Alv11, Ano94h, Ano12, AS08, SF14, Wat93, AS12, Ano96a, Ano99a, IEE97c, Mar14, Ver09]. Secrets [Boo04, Alt12]. Securities [Ano05, Pri00]. Sedimentological [RG21]. SeDuMi [HL03b, Stu99]. Seemingly [Has12]. segment [BKR97]. Seigyo [Oga10]. Seismic [FY18, BD22]. selected [Bar92]. Self [BP97, PNGR00, Bob05, KA13]. self-assembly [KA13]. Self-learning [PNGR00]. Self-Tuning [BP97, Bob05]. semantics [BC12, FBH17]. Semi [CV96, HS24, VFG04]. Semi-conducteurs [CV96]. semi-infinite [HS24, VFG04]. Semiconductor [CV96, Pie96b, CWP98, CC02, Dim00]. semidefinite [BY08, BPS99, TTT99]. semidiscrete [KO00]. Seminar [Mro95b, PJ97]. Semiseparable [VvBM08]. sensed [Rid95]. Sensitivity [CH20, HPK18, WD96, DM06]. Sep [Ano93d]. separation [Row03]. Sept [Cou93]. September [Ano93e, Ano93a, Ano94h, Ano94k, Ano94l, Ano95j, Ano96e, Ano96v, Ano97b, BEH94, BH95, Bub95, C^{+}97, Coo95, IEE95b, IEE96d, IEE04, M^{+}94, Sas96, Van92b]. Sequence [FFR^{+}24, BSB20]. Sequential [CGRvD15, JMD08]. serial [Bai05]. seriation [CFR19]. Series [AT17, BD20, KT10, Wil96b, WK95, CZ17, ZLMQ23]. serious [Ano97a]. Server [HI97, Mor98, Ano97a]. Service [TBH21]. servos [Tob11]. Session [SAE95b, Ano94i]. Set [Atk05, GG92, Góm15, LPD^{+}17, JZW^{+}22, ZL04]. Seventh [IEE98]. sFr [Ano99a, Ano99b, Ano99c]. Sham [JZW^{+}22, YMLW09]. shape [JB06]. Shapes [Gab98, VRVAC23]. shaping [SLM23]. shells [FS23]. Sheraton [D^{+}95, IEE97e, W^{+}97]. Shift [RD22]. Ship [Bir03]. Short [Alv11, Che08, O’B13, Veh07, ACF99]. SIAM [Sin93]. sic [RBD^{+}10]. SICE [Ano94i, Ano94i]. Signal [Amb95, Ano93e, Ano95s, BAEBAS19, Boo98, Bro95, B^{+}94, Can86, Car98, Che94, DS16, Gar96, How95, IEE94b, IEE96f, Jac89, Jac96, Kal97, KSL93, Kub95, Lei11, LS88, LS93, MSY03, Men95, Orf96, Sch94, SC97, SD96, The92h, Wer03, Whi00, von93, Aka00, AM95, Amb99, Ano95t, Ano97l, BC06, BO19, Bru01, Cav00, Cha05c, CiA02, CM99b, Den98, DDN02, EA04, ElA04, Gop10, Har04, Hay96a, IP97, IP00, IP10, Kay93, Kha05, Kin97, Kum05, KL01, KG05, Lai04, Lan00, Lei02, LTE01, MIK00, McC98, McG91, Mea02, MC04, Mit98, Mit99, Mit06, MS00, MGT03, NM99, Paa01, Pee01, Por97, Qua02, Row03, SH05, Sch11, Shi99, Sou99, SWrpSetfe02, Ste03, SH11, Teo98, WWM06, Wel93, ZI04, Fab95, SY20]. Signale [Sch94]. Signals [HV02, HV03, Her96, Jac91, Kar03c, Lat92, MMOP95, OWN97, PV10, PP95, PPR03, Rob04, Sch94, Tay94, The92h, ZTF98, Alt95, Ano97l, Bau02, BH97, BDS97, BDS02, EAK01, Gaj03, Jaf00, JD13, KH97a, KH00, Lat05, LV03, LM04, Li99, Lin99, Man03, MB94, Sem05, She04, SM05, SB96]. Signalverarbeitung [Sch94, Wer03, von93]. Signature [BLL^{+}15]. Silicon [HA95]. Simple [GBM15, OBD23, PS04, DS09, WW99, YHC^{+}22]. Simplify [CA97]. simulate [BBHP^{+}23, Deg20, PK08]. Simulated [DL96]. Simulating [DRR97, HP19]. Simulation [AS15, Ano95v, BRPCR94, BP96, Ben95a, BRS94, CPUARC20, Chu00, Cor95, Cze95, FS17, Fis19, FSO93, FFR^{+}24, HP02a, HBC94, Hof98, ICS96, Irw05, IGPF96, Jor94, KR22, Khe96, Kir93, Kle07, KA11, KA09, Led01, LRD^{+}95, LJC93, LJR93, MVM97, Nel17, PF10, Pet96, SS96a, Thi95, VVM93, Wal18, WL97, XSS20, Zen97, Ach04, And05, Ano93c, Ano94l, Ano97e, Beq98, Beq03, BC12, BH95, Bry96, BEK99, CH23, Can22a, Can24, Cou93, DIN06, Doe98, Fis95, FCP97, Gra11, Gri95, Ham93, Ham96, HC00, Hat01, HT12, HP02b, HLS08, JLM96, Jes01, KR23, KMBP24, Kho00, KA13, LLLW06, MBR21, The98, MFO95, MC02, Moe04, OG95, Ong98, Pav93, PPD95, Ram97, RLH03, SAE95b, SFPO94, SSV95, SCB99, Spa03, SL03, Syd95, TK97, Tra04, VK16]. Simulation [ZJKS23, IEE97a]. Simulation-Based [FFR^{+}24]. Simulations [Gar01, HR96, Hun98, MSS^{+}19, WD96, AD15, Ber03, ME04, SKA19]. simulative [ZZC^{+}08]. Simulator [Wei12, GO97]. SIMULINK [CSV94, MO95, Hau96b, Hof98, PSB04, RLV11, SYW^{+}24, Ano95e, AY96, Bab94, Bis97, BA94, BRS94, CSV94, DH98, FCP97, Gra07, HHF95, Hun98, Jon95b, Jon95a, KNNM97, Löw01, Mat95c, The96, MCd^{+}96, MM00, Mol96b, Ong98, RA95, Rav94a, Rob94, SO93, SA01, SU93, Tew02, UB95, WM95, WM96, Wel93, AB13, Ano92, Ano97e, Are94, BRPCR94, BH96, BBB12, BLL^{+}15, BDM17, BDM18, BC12, BGGT21, BCC^{+}17, Brz97, CCM^{+}03, CBCC96, CSV96, CDS09, CBhM12, CKC94, ÇA10, DH04, DP08, DRR97, DB93b, FGMS21, FSO93, FFR^{+}24, Gar01, Gra11, HBT16, Hau96b, Hof98, HYY^{+}15, HLOR05, HLR14, JE94, JS24, Jor94, KSP11, KA11, LNLB19, MM10, The98, MFO95, MNBB19, Moh01, PPL^{+}18, Pet96, Pol92, RLH03, Riv01, RS09, SCC95, SR94, SFPO94, SRK99, TBH21, TC13]. Simulink [TSCC05, UCL04, VVM93, Wat93, XWZ^{+}22, XWZ^{+}23, ZC08, ZLW^{+}19, ZL17, ZLMQ23]. Simulink-based [RLV11, FGMS21]. Simulink/Stateflow [XWZ^{+}22, XWZ^{+}23, ZLW^{+}19]. Simulink(R) [Kle07]. Simulinku [Brz97]. Simultaneous [FB95b, Joh18, Rib00, LMAV24]. Sinc [Ste10]. Singapore [IEE93a]. Single [YKS94, CATK11, UW12]. Singular [LZ17, AS12, Mol98a]. SIPAMPL [VFG04]. SIPG [BCG17]. SISOTOOL [AM17]. sistema [EGE95]. sistemas [de 05]. Sistemlerinin [Yuk96]. Sixth [IEE94b, Sin93]. size [MMD24]. Sketching [HC92]. skills [Mol02a]. Slater [Fis19]. Sliding [Zin93, PK08]. slope [RAW^{+}16]. Slow [GLJ^{+}91]. Small [AM95, CH20, FSZD20, KSP11, Mol95a]. Small-signal [AM95]. Smart [Gro05, Joh95]. SMI [HCV97]. Smith [SMS95]. Smooth [CMR17, PPT10]. SMOS [SS18]. SMPS [BH96]. smt [RZR12]. Snippet [Esm14, LZ17]. Societies [Ano96g]. Society [Ano95k, IEE95a, IEE96a, Zem97, Ano97i]. SOFSEM’97 [PJ97]. Soft [Ano94h, Ano96e, Ano97b, Ano95f, JSM97, Kec01, MC02, ZJ01]. Softstat [BF97]. Software [Ano94j, Ano96r, Ano97k, AJGO^{+}20, BSL93, Bee17, BCD^{+}20, BPB96, Bra97b, BBEM22, But92, Car92, CFGG94, DDK14, DL96, Edd09, ES10, EM14, Fos01, GRDL^{+}12, GR97, Hai08, HL95, Hen94, JMD08, MRK20, The92d, The92a, The92c, The92b, The93a, The93b, The93c, The97, Mir96, MW93, MBB^{+}09, MBMW95, NPT15, PS96a, PS96b, PPL^{+}18, PR14, RSW15, Sil96, Tho94, TM97, WTL00, Web97, YKS94, Ano96j, Ano97e, Ano97j, AI21, ARR02, BF97, BGGL20, BY08, BO19, Bra97a, BEV06, BDL25, DPE96, DM96, Edl04, FGMS21, FGCG94, FSC95, Gau06, HP02a, JS04, Kro95, KK93, LP13, Led04, LS05a, Li96, Li99, LM91, Mat95a, Meu20, Mil20, PK04, RDP14, RBD^{+}10, RBD^{+}11, RSS08, SGA95, SLM23, TTT99, VH04, XZL24, Sil96]. software-defined [JS04]. SoftwareX [CK23]. Softwate [YKS94]. SOHO [HR96]. Sojourn [Cse92]. Solar [HR96]. solid [FCP97]. solidification [GO97]. Solka [Mar14, Mar19, Alv11, Ano12, Ni22]. solucion [EGE95]. Solution [BBC^{+}94b, BBC^{+}94a, CHMN13, Han92, Sha09, Yuk96, AS12, Ano94e, DF99, FGMS21, HMT13, Han94, Jon01, LVV05, LQT18, Mil20, Mol03a, RDP14, ZD03]. Solutions [BL96a, BD95, BD00, JH97, Mat94a, SSH94b, SS96b, Yan05, Ano96m, AD14, BH97, Gau16b, Gri95, MP18, Nah08, Ren17, Tow16]. Solve [Riz17, AD14, AD15, MR11, SKF05, Som07]. solved [DJKP07, KA02, Riv01]. Solver [CMR17, WL94, Dav13, Faz10, KS01, Rei93, RC13, RC16, WEM98, ZZG^{+}14]. solvers [Mol96b]. Solving [Ano95p, AS08, Con95b, Con95a, Con96, Ett93, Fab97, GH93, GH95, GH97, GH04, HPZ19, Kel03, LCMCD22, Ned95, Oga94b, PR14, SRK99, ST01, Sha02, SGT03, Sha05, Ste96, Zha98, Att09, Att12, BNN16, Bro95, Bün20, CS99, DCF95, Duf04, EMMK01, EPJ^{+}05, Ett97, GI06, GS07, JZW^{+}22, LP05, Mai00, MY05, PIAH12, PPS06, PF07, Rad96, RBD^{+}10, RBD^{+}11, Sti04, YMLW09, YJ24, ZZC^{+}08]. SOM [TC13]. Some [FCP97, ÖK17, Nas01, PT07, SR90]. sonar [ZE95]. Sons [Ano00]. Soon [CW05e].

Sophomore [MASV96]. sorting [Bra02b]. sound [SW95]. Source [BD20, HWB15, LO16, MP99, MP00, TCD^{+}22, WR16, WR17, BGGL20, PNL^{+}21, Ren17, Row03, SM14, Ste08, Ste13, TACA15, VHM17, ZLLT23]. Source-Level [MP99, MP00]. Southampton [Ano94l]. Southeastcon [IEE97d]. Southeastern [NP96]. Souza [How15]. SP [OE95]. Space [Bar97, Che93, Góm15, HCV97, PA11, SCC95, Spe95, Spe96, Vac95, VLV00, Bay99, FB95b, MP18, XSS20, ZJKS23]. Space-time [VLV00]. Spacecraft [Bry94]. Spain [M^{+}94, Ano94a]. SPARC [XSS20, ZJKS23]. Sparse [GMS92, IKK^{+}19, LCMCD22, Mat92c, Mat94b, PR14, SCL^{+}18, BK07a, BCR03, BPS05, Dav05, EL16, HL03a, KW05, LP97, PT24, PT24]. SpaSM [SCL^{+}18]. Spatial [Hsu99, SLM23]. Speaking [Mol99b, Mol99a]. specdicho [ST12]. special [Mol02a]. specialist [Ken95]. Specific [GG92, LPD^{+}17, She10, USE99]. specifically [Bro95]. Specification [Jan02]. Spectra [NP93, Spe95, Spe96, SWG^{+}94]. Spectral [New94, Poz05, ST12, Shu01, SM05, Tre00, CFR19, MIK00, Poz14, SM97]. Spectram [Rab20]. spectrophotometer [dP96a]. Spectroscopic [Rab20]. Spectroscopy [Fis19, KMBP24]. spectrum [Set05, Cas14]. Speech [Chi00, IEE96f, Kin97, Kub95, PS96a, PS96b, SWS97, Ano96j, Qua02]. Speed [GPR14, Kir93, AP02, MD95, Mol00b, RLH03, RR01, RR02, SCB99]. Speeding [Cla03, PIAH12]. SPEX [LCMCD22]. sphere [BC22]. Spherical [SF14, Bal19, LJ23, dMMLOS20]. Spikes [Wil99]. spin [JAC20]. spin-up [JAC20]. Spinterp [KW05]. SpinUpFlowDescriptor [JAC20]. Spiral [VA94]. Spline [KK97]. splines [Jiř97]. split [OAKS11]. split-step [OAKS11]. spreadsheets [Kar01b]. Springer [Ano99b, Ano99c, Sha04]. Springer-Verlag [Ano99b, Ano99c, Sha04]. SPSS [Veh07, Mar03, Mar07]. SQP [MR11]. square [UW12]. Squares [BV18, RS08]. Squeezing [HPZ19]. SR1 [BBEM22]. srodowisko [MM96]. srodowisku [Mro95a]. SSMMATLAB [Góm15]. SSNT [PBI07]. St [Dag94]. Stability [Hig96, RA95, TC97a, TC97b, Bir03, DPE96, Hig02b, LB99, Rob96]. Stable [dVSWAL17, Spo02, LJ23]. stack [LH13]. Stafford [Ano93a]. Staggered [SF14]. Stampacchia [DG96]. Standalone [CdFCS98]. Standard [BCC^{+}17, MG18, Ano97j, Bra97a, Mol96a]. Standardization [TRD11]. StaRMAP [SF14]. Started [Pra96, Pra99, Pra02, Pra06]. Stata [SR09]. STATCOMs [BPCCM23]. State [Bar97, C^{+}96, Che93, Góm15, HCV97, Lum91, PA11, PBB22, SYW^{+}24, Tay95, Vac95, vdH05, Bay99, CR13, CR20, mP97, Rao11]. State-Aware [SYW^{+}24]. State-Space [Che93, Vac95]. Stateflow [MM10, Riv01, XWZ^{+}22, XWZ^{+}23, ZLW^{+}19]. Static [DH12a, JB03, Por23, SBL19]. Statics [SLI99b]. stationary [AD14, BAEBAS19]. STATISTICA [Veh07, Bla02a, Mar03, Mar07]. Statistical [Cla95, Com92, Del02, Góm15, Hay96a, KFG94, MIK00, SCL^{+}18, The92h, BF97, Kay93, KM10, KM11, KS97, LS05a, Shi99, VH04]. Statistici [Com92]. Statisticians [AO95]. Statistics [Ano09, Ber09, Cha17, Jai09, Kus02, Mar07, MM02, Mar11, Veh07, Vid11, Wik04, AO95, Beu05, HTJ90, Li99, Mar03, MM08, Row03, Ver09, Wil03, Zem97, Zie97, Zem97, O’B13]. Statistik [Beu05]. STDI [KAB97]. steering [GMT95, SAE95a, Zen97]. Step [CH20, MS94, CT97, Hau01, OAKS11]. step-by-step [CT97]. stepping [KA13]. stepwise [Van04]. stereo [FBFB04]. sterowania [BF95]. Stiff [Mol03a, BGPPRW14]. Stiffness [Mol03a, BCG17]. Still [Thy10]. Stimulus [RD22]. Stochastic [CFGG94, How15, HPK18, HLS08, Kao97, KR22, SS15, AD15, BL02, FGCG94, GS07, KR23, Kul99, Mor00, Spa03, YG99]. stochastics [Hig04]. Stoichiometry [AB96]. Stokes [FD01, ZZG^{+}14]. storage [JB03]. storm [Boo04]. Strain [Yan05]. Strategic [Nyh08, Zar97]. Strategies [BA94, Lum02]. Strategy [SYW^{+}24, MR11]. Strathclyde [IEE94e]. streamlined [Nas01]. Street [Ano96i]. strength [Cha95]. Stress [Yan05, Cha95]. stress-strength [Cha95]. Striving [Mol95c]. strongly [ZZG^{+}14]. Structural [Lev96, Lip07, NPT15, PNT15, Yan05, FB95b, HAM06, LCL05, SM95, YHC^{+}22, KAB97]. Structure [IEE96b, LV03, MGW99, ZW93b, ZW93a]. structured [RZR12]. Structures [Bro94, DL96, Gaw96, JK93, Ano94e, Bha06, Car99, CSG98, FB95b, Gaw98, Gaw04, Joh95, Por23, SZCP21]. Struktury [BF95]. Strum [Ano93b]. Stuck [Ano00]. Student [DDW93, HL95, The92g, The92e, The92f, The97, The96, MVM97, Sch97, HL97, The98, Wit04]. Students [Ano95v, CBCC96, Lei11, Lev92, Cur05, MM10, Nas01]. Studiengänge [Bät20]. Studies [AFL^{+}12, BA94, BRS94, Cra96, SWG^{+}94, BF09, CLTS20, DPE96, Mac04, Moe04, RS02, Sch12]. studio [Qur05]. Study [Ans14, DIN06, Gou20, Lay03, MA96b, QCPG96, Ano96l, DH97b, DH00, PGBG94, TJML92]. Sturm [LVV05, LV16]. style [Joh11, WB16, Epp11]. Subband [Ano95q]. Subband/Transform [Ano95q]. Subject [Mol98b]. Subjects [RD22]. sublattices [Kat09]. Subproblems [EM14, RSS08]. subset [BC17b]. Subspace [AS15, vdM96]. subspaces [Kat09]. subtle [Mol03a]. Such [WB12]. SUGAR [VZDCM25]. Suite [MP17, Riz17, SR95, SR02, WR00, ANV00, MG13, SR97]. summary [Gop10]. Summer [Hor97, ICS96, Mic94b, MO95]. sump [Pol92]. Sun [Van92b, The90, ML90]. Supercomputing [ACM96]. Superconducting [FKSM97]. superfast [Dak06]. Supérieure [Pav93]. SuperNEC [Coo96]. Supervision [MCd^{+}96]. supplement [HT97]. Supplementary [Pie96a]. supply [BH96, Cou93]. Support [BBB12, DH97a, Kec01, Ker95, Vid11, O’B13]. Supported [TBHS94]. Supports [RG21]. suppression [DH97a]. surface [FMS25, SKA19]. Surfaces [FK11, GI06, Rov10a, Rov10b]. surges [Boo04]. Survey [Hea97, TRD11, Cho02]. Surveying [RGM22]. Suspended [Mar16]. suspension [PGBG94, SAE95a]. SVD [Mol06a]. svt [LZ17]. Swansea [Bar92]. Swazi [Van92b]. Swaziland [Van92b]. Swedish [Bäc97, PES98]. SWIGLAL [Wet20]. Swiss [DLA^{+}16]. switch [Elg01]. switched [BH96]. Symbolic [ARRY01, Ano97k, Bra97b, Con95b, DR96, DM06, Glo98, JHPK97, Jef08, RBZ96, Roq13, SHX96, VZDCM25, LP05, mP97, PR02, AH09]. Symmetric [FKSM97, MT84, MS14, MY05, Stu99]. symplectic [BSB23]. Symposium [Ano94e, BLM97, C^{+}96, Cse99, Glo98, IEE95a, IEE96d, IEE96c, IEE98, Jef08, MGC94, NP96, Pat94, dOSRS19, Syd95, USE94, Ano94b, Bar92, BS95a, Gre96, IF94]. Synchronization [RD22]. Synchronous [BCC^{+}17]. synergism [LL96]. synergistic [PAG11]. Synthesis [GPK05, JRCS95, JCRS96, Nat94, Ano96k, Chi00, Ega00, GMT95, HNS^{+}01, Kin97, Nor04, Man96]. Synthetic [Sou99]. SYSID’94 [BS95a]. System [Ada97, Ano95e, Bau02, Ben95b, BRS94, Che93, Che94, CBhM12, DH95, DM90, Den98, Doe98, FC95, FC00, Gra92, HK95, HWB15, IEE96d, Kir93, KA09, KPS95, KH94, LL86, Lju87, Lju88, Lju93, Lju99, MGC94, MASV96, NP96, NRHP96, Oga04, Pal05a, PN95, SDPM04, SSH94a, SSH94b, Sch94, SH93, Shi92, SA01, TC97a, TC97b, Tib93, UE99, VVM93, YKS94, Ajj95, Ano93c, Ano96o, Are94, BM94, Bro01, BEV06, Car98, CSB03, Cha02a, Che99, CG00, CM99b, DHS03, Dat13, Dav13, DPE96, DDN02, Dor00, Fis95, GL96, GGS01, Hun98, JB06, JB07, KK01b, Kul99, MBR21, The98, MFO95, NMS^{+}06, Pan89, Pet96, Pol92, Rob96, Rog00, SHX96, Shi98a, Shi98b, SM95, WGP95, Zhu01, BS95a]. Systeme [Hof98]. Systemen [Ben98b]. Systèmes [Bou95]. systemow [BF95]. Systems [AFOP19, AY96, BRPCR94, Bab94, BR96, BCHS98, BBC^{+}94b, BBC^{+}94a, BDM17, BDM18, Bir18, Bis93, Bou95, BGGT21, Bub95, C^{+}96, CGM95, CL96, CFGG94, CF93c, Cou95, Cou97, Cze95, DB95, DRR97, DB93b, FSO93, FPW90, FPEN94, FKSM97, GPK05, HBC94, HPZ19, HTCI96, Hof98, IEE94f, IEE96b, IEE97a, IEE04, Jac91, Khe96, Kuo95, KYN95, Lam95, Lat92, LL92, LL95, LY97, LG94, LRD^{+}95, LCMCD22, Mah00, Mah05, Man96, Mar95b, MA96b, MVM97, M^{+}94, Nat94, Nis95, Oga94a, Oga95, Oga98, PH94, Per91, PP95, PH96, Pie96a, Rod92, RMS93, Saa93, SS96a, Sch96, SP91, SKG97, Shu17, Szy93, Tay94, Thi95, Tho94, UB95, Wat93, WL97, XWZ^{+}23, Yuk96, ZLW^{+}19, Zin93, ÅW97a, AM01, Ali02, Ano93b, Ano94a, Ano94c, Ano96m, Ano96l, Ano96p, Ano96s, AY91]. systems [AH09, ARR02, BSC^{+}00, BDR04, Bar92, BNN16, Bay99, BD96, BGPPRW14, Ben98b, BFG^{+}14, Bis97, BC17b, Bla02b, BL02, BLM97, BDS97, BDS02, CR13, CR20, CAC94, Chi97, CFN01, CFN02, Col04, Cou01, Cou93, Cun03, Dag94, Dat04, DB98, DB04, EAK01, EA04, ElA04, EA03, FGCG94, FCP97, FPW98, FPEN02, Gaj03, GMT95, GP98, GWA01, Gri94, Gup02, HL03a, Har04, Hay01, HV02, HV03, Hel04, JE91, Jan02, Kal97, KH97a, KH00, Kar03c, KMBP24, Kho00, Kin93, KA02, Kle07, KA11, KG03, Lat05, Led01, Led04, LV03, LL96, Lin99, LCL05, LHW01, Lum02, Lyn04, Lys00a, Lys00b, Lys01, Lys05, Mac04, MAB^{+}11, ME04, Man03, Mei05, MC02, Men01, Moo98, Nai03, Nis04, Nør00, OG95, OWN97, PV10, Pal98b, Pat94, PH00, PPR03, PK08, PS98]. systems [PS00, PSB04, PSP04, Rob04, RA92, Rog03, RL00, SAE95b, SI00, SL17, SH18, Sem05, SFPO94, She04, Spo02, Ste02, SB96, SK94, SK95, SK00, TK97, Tra04, TYK04, Van01a, Wea97, WK03, WE96, XWZ^{+}22, Yan99, Zal96, ZD03, ZTF98, ZT02, ZJ01, ZW93b, ZW93a, vdM96, Ano94k, JE91, Sch94, Syd95, Ken95]. Szeged [Cse99].

Tactical [Zar97]. Tailings [MLAS21]. taken [Zla17]. Talbot [Riz17]. Tale [Mol95a]. tame [Oga10]. Taming [DH12b, Nah11]. Tape [CW05d]. Targeting [LPD^{+}17, BC17b, TC13]. Task [RHB96]. Taylor [Bün20]. Tchakaloff [BV19]. TDIST [Wit04]. Teach [Ano95v, Web97, Bha95]. Teacher [Ano97i, Bou97]. Teaching [Ano95r, ASG94, Att95b, Att95a, Att96, Bur93, Bur94, Car92, CL96, EP 97, JH96, JHPK97, Kla92, Lev92, LQ94, McG91, POVD96, RA92, Rod92, Ano95t, JDFV08, Mol04d]. Tech [IEE97d]. Technical [Ano95v, Mol15, Sti04, Bät20]. technicians [Lai04]. technicznych [MM96]. Techniques [Ano94h, Ano95h, Ano96e, Ano97b, BSL93, BJ02, De 96, DP99, Irw05, KK96, Löw01, POVD96, Wil96b, Ano95b, Ano95f, Ano97l, BPS02, DCF95, DM96, DIN06, EU07, GMT95, LH14, Pac00, PPD95, Som07, Spo02, ZI04]. technische [Bät20]. Technology [Ano93e, Ano94i, Ano95s, Ano97i, Ano97k, Ano97j, Bra97b, Bra97a, KK01a, Kum14, Lum91, Lum92, Stu96, Kar01a, LL03, SAE95a, Sta03b, SBC04, ZL04, Ano96u, I^{+}96]. Technology-based [I^{+}96]. Telecommunication [JS04].

Telecommunications [Nas01, HB04]. Telemedicine [Shu17]. Telescoping [CMKH03]. Temperature [SS18]. Templates [BBC^{+}94b, BBC^{+}94a, FHH99]. Tennessee [D^{+}95]. tensor [BK06]. Tensors [KT14, BK07a]. Teoría [de 05]. Term [Mik23]. Terms [Ans14, CH20, Kum14]. terrain [OAKS11]. ters [Ano95c]. Test [FFR^{+}24, Hig89, Hig91, MNHH94, MNBB19, SYW^{+}24, TH09, BGGL20, GHN19, HBT16, LB99, LNLB19, Som07, Wan15, Hig93, Hig95]. Testing [Edd09, FFR^{+}24, Fos01, SM95, Whi00, ZC08]. Tests [ÖK17, Ren17]. Tetragamma [Mol02a]. Texas [IEE93b, Pat94, USE99, Dut16, IEE94c]. text [LC98]. textbook [Iro15, Mol04a, Mun13]. textgraph [Li99]. TFMLAB [BFSJP^{+}21]. TFSAP [BO19]. Their [ÖK17, Ano94e, LCL05, Ves98]. theme [Mal96]. Theorem [TC97b, GRH^{+}21, JS24]. Theoretical [Bar16b]. Theories [TKD00, dOSRS19]. Theory [Brz97, DG96, FBC00, Gol91, HBC94, KBKS98, Lju87, Men95, NP96, Per91, PNT15, Rod92, Shi92, SW94, TD98, TW06, ÅW97a, Bau02, BAY98, Che99, Eng05, FB98, GC99, Gre96, GA01, Gri95, Har01, Hay96b, Hay02, Kal97, Kno00, Lju99, LB00, Lys00a, Lys01, Mac05, Mac04, Mar01, MC02, Özb00, PK04, PJ97, RB98, Rob96, Rou98, SGLBA98, Shi98a, Shi98b, SS10, Tho04, TW02, Van01b, vdM96]. there [Mol95a, Mol95d, Mol97]. Thermal [Mai00]. Thermath [MAC08]. things [WW99]. Third [C^{+}97, IEE94e, IEE94f, Sil96, Van92b]. Thirty [IEE97c]. Thirty-second [IEE97c]. Thomas [BC05]. thoughts [Mol95b]. Three [BC05, FGjS15, LJR93, AI21, BC22, MP18, Som07]. three-dimensional [AI21, MP18]. Three-phase [LJR93]. Thresholding [LZ17]. TIA [Lan00]. tidal [Boo04]. tide [Boo04]. timbre [Set05].

Time [AKB94, BD20, BO19, BAEBAS19, DL95, KAB97, KA13, KT10, Oga95, POVD96, PSR16, RSW15, WM95, WM96, WK95, AP01, ARR02, BV21, BL02, Chi97, DL01, FB95a, Jan02, JSB20, KL01, Li96, LHZT23, Mir97, Mol95b, Qua02, TSCC05, UW12, VLV00, WWM06, ZLMQ23, SM94, Zal96]. time-dependent [UW12]. Time-Frequency [BO19, BAEBAS19, PSR16]. time-independent [BV21]. time-saver [Mir97]. Time-Scale [BO19]. Time-stepping [KA13]. Time-to-Event [RSW15]. timed [JS24]. Timer [Hau96a, Hau96b]. times [Cse92]. TMS320C30 [AHS94, SC97]. TMS320C55X [KL01]. TMS320C6000 [KK01b, Qur05]. TMS320C6x [Cha02c]. TMS320Cx [WWM06]. Tobii [Jon18]. today [Ber95]. Toeplitz [BMR19]. Tokyo [Ano94i, IEE96b, IF94]. Tolerance [Boy99]. Tony [dOSRS19]. Tool [BBB12, Ben95b, CL96, Cor95, Gre94, KA09, KNNM97, Lu96, MLAS21, NPT15, PS96a, PS96b, PNT15, RVV^{+}92, SWG^{+}94, TQ96, Van00b, VBB18, WTL00, WK95, ZL17, Ano95l, Ano96j, Ano96q, Ano98b, BAEBAS19, BBF^{+}19, BBN^{+}22, BD22, CLTS20, Can22a, Can22b, Can24, CK22, CK23, CF16, FJ22, Lin05, MA95, NK18, OAKS11, TB95, VRVAC23, YHC^{+}22, dMMLOS20]. Toolbox [AFOP19, ÁBZ17, ÁBZ20, Ano97j, Ano19, BV95, Ber09, BP97, Bra97a, Bro95, BF95, BKL19, CLTS20, CGM95, CGRvD15, CCF02, CSV94, Con95b, CSY15a, CSY15b, Cor95, DDD97, DS16, DB93a, Dri96, FC95, GHH20, Gra92, GPK05, HCV97, Hig15, HTCI96, HLP96, ICS^{+}18, JRCS95, JCRS96, Jon91, KR22, KAB97, KPS95, KSL93, KT14, KFG94, LL86, LV16, LS88, LS93, Lju88, Lju93, LRD^{+}95, LSvdV19, Lu96, McI16, MG18, MON12, Nat94, NRHP96, PH94, PHL95, PA11, Per93, PSR16, Rab20, RS08, RD22, SS96a, SM14, SCL^{+}18, SP22, TH01, TH09, TRD11, Tib93, Wal18, WR17, YR19, Ano96k, Ano96o, Ano96p, Ano96r, AD14, AD15, AMR18, Bal19, BFSJP^{+}21, BD96, BBHP^{+}23, BMR19, BlD^{+}20, Bün20, CFPF94, CF95, CBL24, CV00, Cor96, CFF^{+}91]. toolbox [Dri05, ERS07, EMMK01, FR18, FS23, GV94, HTJ90, tHLMN19, Hen07, HT12, Hoh14, HL22, Hon92, How91, JRA^{+}18, JZW^{+}22, JAC20, JGGF23, KR23, KMBP24, KK97, KMLP^{+}23, LMN18, LSH95, MMD24, MBR21, Mag02, MBBC95, Mat97, Nas20, PMNWR20, PNL^{+}21, PC13, RCT20, RKZ^{+}14, RZR12, Ren17, RAW^{+}16, RHR^{+}21, Rum95, SHX96, Som07, Stu99, TSA21, TCE16, TACA15, WGP95, WE96, YMLW09, ZL13, ZLLT23, Ano95g, AH09, BFM89, CSV94, CSV96, Fab95, FC00, Hig93, Hig95, Hig02c, Hon91, Mac91, MMOP95, Mol99b, NK18, Sar17, SY20, SA01, SM94, Tib93, VA94, ZW93a, CS12, FRAK15, MRK20]. Toolboxes [Yan05, Chi00, CF93b, CF93a, Mac91, Mat95c]. Toolchest [Wea97]. Toolkit [Kin95, KT10, MA96b, NRPH96, Ano96l, Ano96n, CT97, LCL05, Sar10, TCE21]. Tools [Ano97f, FGjS15, GHN19, GR97, HK95, Han92, Han99, KH94, NR96b, Pie96a, QCPG96, Rah93, TC97a, TC97b, BM94, Bry96, CG00, C^{+}97, FHH99, FR95, Man01, Nir02, Han94, Han07, HSH12]. Top [TBH21, Cas14]. Topical [JE91]. Topics [Pfe95, Rav94a, Bha06, BM94]. TopModel [Rom97]. topographical [Zek17]. topological [HG24, HGM^{+}25]. Topology [GV25, IZBT21]. TopoZeko [Zek17]. total [MFO95]. Tour [Mol93]. Towers [AI21]. Track [Can22a, PBI07]. Tracker [Jon18]. Tracking [BSL93, JSB20, Wei12, BSLK01, QH01, ZSW^{+}17]. traction [BFSJP^{+}21, SFPO94]. Trade [BEH94, LNLB19]. trade-off [LNLB19]. trading [RB05]. Train [Can22a]. Traits [She10]. trajectories [SKA19]. Trajectory [MMPOQ^{+}25]. Transducer [Ano94l]. Transducers [And95]. Transfer [BKGS02, Che93, GV25, SF14, Ben02, Kin93, ZLLT23]. Transfer-Function [Che93]. Transform [Ano95q, DL95, Duf04, JNR19, KDAB19, Sta03b, Uhl02, Wil96b, tC04, DL01, JL01, Rob04, Ves98, BC17a, Van92a]. Transformation [HWB15, RD22, WR16, WR17, YJ24]. Transformations [MP99, MP00]. Transformer [ICL97]. Transforming [BV08]. Transforms [Jac91, Mol00a, PP95, PSR16, BGG98, Mai00, Mea02, PPR03, RB98]. Transient [HHF95, SU93, AM95]. Transients [MA96a, MA97a, MA97b, Van01a]. Transit [M^{+}94]. Transition [Bar97, Rab20]. Translating [Mir96, TSCC05]. Translation [RS09, Zag16, DP99]. translational [CATK11]. Translator [DP96b, JB07, Ker95]. Transmission [BRS94, CST20, MNHH94, And99, BBHP^{+}23, Che04, Cse92]. transmitted [JS04]. Transport [TYK04, Bha95, SAKG15, SCB99, Tho00]. Transpose [IMM12]. transputer [Coo95, Kad95]. Traveling [GS12]. treatment [Hsu99]. tree [Liu23]. tree-ring [Liu23]. trees [Cam06]. Trends [MBS15, PJ97, Boo04]. TRIAC [PBI07]. Trials [RSW15]. triangles [LJ23]. triangularization [Tre10]. trig [Mol98b]. trigonometric [API^{+}19]. Trigonometry [Mol98b]. trinn [Hau01]. Trois [IEE94a]. Trois-Rivières [IEE94a]. truck [SAE95a]. truncation [Cam06]. Trust [BBEM22, EM14, RSS08]. Trust-Region [EM14, BBEM22, RSS08]. TTA [CCM^{+}03]. TTB [Can22a]. TTB-2D [Can22a]. Tucker [KT14]. Tucson [MGC94]. tumor [BDL25]. tuned [SFK91]. Tuning [BP97, Set05, Bob05]. turbines [KC95]. Turbocharging [KH96, KH97b]. turbulent [VK16]. Turkey [EM94]. Turkish [Yuk96]. turning [Kro95]. tutor [Dak05, Dak06]. Tutorial [Dak05, Hau90, CBL24, HL96, HL98, HL01, Kin01, Kin06, Lei02, Mar92, Sto13]. Tutorials [TM97, PSTO97]. tweezercalib [HTNFBS06a, HTNFBS06b]. tweezers [HTNFBS06a, HTNFBS06b, TNBSF04]. twelve [DJKP07]. Twenty [MV03, NP96, Gra94]. Twenty-eighth [NP96]. Twenty-Five [MV03]. twenty-fourth [Gra94]. Twin [MMPOQ^{+}25, VA94]. Two [CHMN13, Esm14, GRDL^{+}12, GBM15, Has12, Lam95, Mol95a, Mol06b, YA95a, YA95b, AM95, EA95, Kok07, Kok15, KK93, OAKS11, RKZ^{+}14, Rob96]. two-dimensional [Kok07, Kok15, Rob96]. Two-Layer [GRDL^{+}12]. Two-Level [Esm14]. two-module [AM95]. two-point [KK93]. two-way [OAKS11]. Type [CMR17, CMKH03, SP22, CS12]. Type-Driven [CMKH03]. Typed [WB12].

Ubungen [von93]. UK [Bar92, Ano93a, Ano94d, Ano94k, Ano94c, Ano94l, Ano95b, Ano95t, Ano96v, IEE94e, KPN^{+}04]. UKACC [Ano96v]. ukladow [BF95, Szy93]. Ulam [DPR05]. Ultra [BN10, BN11, DG04]. Ultra-fast [BN10, BN11]. Ultrasonic [Ano99d]. Ultrasound [Shu17, FM24, Sza04, FM24]. UMFPACK [Dav04b]. UMTS [LWN02]. UmUTracker [ZSW^{+}17]. Uncertain [AFOP19, Men01, Ano94c]. uncertainties [Ben98a, Ben04]. Uncertainty [MBS15, SY20, SS15, BDR04, CBL24, PT24, tHLMN19, LMN18, How15]. Undergraduate [AY94, AS96, DM90, RS15a, TM97, WW94, Iro15, Mun13]. Undergraduates [DL96, Gro99]. Understanding [DG04, Dim00, KK01c]. Undocumented [Alt12]. Unidimensional [She08b, HAM02]. Unified [XWZ^{+}22, XWZ^{+}23, ZLW^{+}19, Bro01, Hsu99]. unifies [Mol96a]. Uniform [CM67, BBF^{+}19]. Unifying [ES10, Has12, dOSRS19]. unimodal [MT84]. unit [Are94]. United [Ano95i]. Univariate [BD20]. Universal [MM96]. Université [IEE94a]. Universities [Ano93a, Ano95t, Zem97]. University [Ald96, Ano94d, Ano96v, C^{+}96, Cou93, Glo98, IEE94e, IEE96b, JLM96, Lum91, Lum92, Mal96, Dut16]. Uniwersalne [MM96]. UNIX [FY18, The92d, The92a, The92b, The93c]. unmixing [Row03]. unruly [Nah11]. unsymmetric [Dav04b]. unsymmetric-pattern [Dav04b]. Update [tHLMN19]. updated [HLOR05, HLR14, KMLP^{+}23]. Updates [DDK14]. updating [LLZ18]. updating/downdating [LLZ18]. Upen2DTool [BBF^{+}19]. Upper [Ano96a, BB97]. Upward [Kwo93]. upwind [WSST05]. US$34.50 [Ano93b]. USA [ACM96, ACM98, Ano94b, H^{+}96, IEE97a, IEE04, Pat94, ACM97, Ano93e, Ano94e, Ano95j, Ano95s, Dag94, Ham96, IEE93b, IEE94d, IEE96f, IEE96d, IEE97b, IEE97e, IEE97d, SAE95a, USE94, USE99]. Usage [MSS^{+}19, Ogu95, RHR^{+}21]. USD [How15]. Use [BF97, CL96, JDFV08, KAB97, KSL93, Ano97l, BD96, BFM89, CK22, CK23, DB93a, Den98, GB89, Gra92, LL86, LS88, LS93, Lju88, Lju93, Mag02, MA95, SGA95, WE96, Wri97]. Used [FBG94]. USENIX [USE94]. User [DB93a, Dun99, EI05, Gra92, HL95, LL86, LS88, LS93, Lju87, Lju88, Lju93, LSvdV19, MBL^{+}97, The92d, Mat92b, The92g, The93a, The97, The96, Mol80b, Mol82, MLB87a, Mol18b, PHL95, SCC95, TS14, VZDCM25, FM24, HL97, LM91, Lju99, Mat97, The98, ML90, Nyh08, Ozk23, TSA21, PES98]. User-Defined [MBL^{+}97]. User-friendly [TS14, VZDCM25]. Users [MLB87b, HLR01, HLR06a, HLR14, RGM22]. uses [KO00, Wea97]. Using [Ada97, AB96, Ano95c, Ano95e, Ano95h, Ano95p, Ano95r, Ano95u, Ano95v, Ano98a, Ano99d, ASG94, AY97, Bäc95, Bäc97, BR96, BPCCM23, Bee17, BKGS02, Bis93, Brz97, B^{+}94, Bur94, CBCC96, CSV96, CG00, Con95b, Con96, ÇA10, DR96, EO94, EI05, Fab97, FSO93, FC95, GH93, GH95, Gar96, GBM15, Got95, GV94, HC95, HHF95, Hau90, HR96, Her96, Hil91, How95, ICL97, JH96, JHPK97, JH97, Kao09, Kay05, Kir93, KK01c, KW95, Kro95, Kub95, KBQ97, Lam95, Law05, Lei11, LL92, LL95, LP97, LJC93, Lys03, MAC08, Mah00, Mah05, Mar07, Mat94a, MF94, MMPOQ^{+}25, MVM97, New93, NR96b, Ono01, PFG08, PR14, PL95, Pfe95, POVD96, Poz05, Rah93, RA95, Rob96, RS15a, Saa93, SMS96, SS96a, SH93, SKF05, Shu17, Som07]. Using [SC97, SWS97, Ste96, SM07, Stu99, VVM93, VA94, Veh07, VWHK96, Wat93, WM95, WM96, WR17, WCS92, WT94, WT97, ZLW^{+}19, dP96a, vdH05, AM17, AGL07, AB13, AM95, Alt95, AB98, AHS94, And95, Ano93b, Ano94l, Ano95b, Ano96b, Ano96c, Ano96d, ASA96, Att99, Att04, Bäc00, Bäc04, BH96, Bar16b, BKR97, BF09, BHC^{+}24, Bis97, BC06, Boo98, Boy99, BDS97, BDS02, Bug95b, Bur93, But11, CH23, Cap13, Cha02c, Chi97, CBL24, Con95a, DCF95, DPE96, DM96, Dja98, Dja00, DM06, DEQOR13, EP94a, EU07, ELR02, FGMS21, Fau99, Fau08, FD01, Fis95, FC00, GH97, GH04, GMT95, Gar01, GRH^{+}21, GSM95, GS08, GD99, GWE04, Gra11, GA01, GPR14, Har05, HDR97, HGM^{+}25, HC00, Hat01, HAM02, Hul99, Hun98, IP97, IP00, JKR92]. using [Jaf00, JNR19, Jor94, KSP11, KH97a, KH00, Kar01b, Kay06, KK01b, Kin97, KC95, KYN95, KB97, KB00, Lan00, LP05, Led04, LW94, LC09, LW03, LJR93, Luo95, LTE01, Lyn04, Man01, Mar03, MFO95, MF99, MF04, McC98, MC02, Mid00, MB94, Mit99, Moh01, Möl07, MV22, NM99, NMS^{+}06, Ong98, OE95, PGBG94, mP97, PR02, PIAH12, PL00, Pet96, PA99, PA04, Poz14, PS98, PS00, PSB04, RDP14, RBD^{+}10, RBD^{+}11, Rao11, Rav94a, Rei93, Ren17, RLH03, Rob95, Rob04, RA92, Rod92, Roq13, Rov90, Rus08, SMS95, SI00, SH00, SH05, Sch11, SR94, SFPO94, SD91, Sta05b, SW95, SB96, SK94, SK95, SK00, SS10, SU93, TJML92, TW98, UCL04, Van97, Van00a, WD96, WR04, WHT02, WTH03, YCCM05, ZJ01, ZW93b]. Utah [I^{+}96]. utilisation [EPJ^{+}05]. Utilization [HWB15, MGW99]. Utilizing [AS96, AY96, EA95, YA95a, YA95b]. UTP [dOSRS19]. UTV [FHH99].

V [Ano05, Pri00]. V.4 [HDR97, IP97]. v1.1 [CR20]. V3.0 [Tót08]. V4.3 [Dav04b]. VA [Sin93, Wie94]. Validated [Cse99]. validating [CDS09]. Validation [RS09]. validity [JGGF23]. Valley [Van92b]. Valuation [Hig02a, Hig04]. Value [CHMN13, EP96b, EP04a, LZ17, Ano93c, BD05, EP00a, KK93]. values [Mol98a]. Variable [BV08, IEE96b, Meu20, OAKS11, RLH03, SL17, ZZG^{+}14, ZW93b, ZW93a]. variable-speed [RLH03]. variables [JDFV08, MT00, Pee01, Wit04, Wun05]. variants [MS14]. variational [JR99]. variations [Mac05]. VARMA [Jon09]. VAX [LM91, MLB87c]. VAX/VMS [MLB87c]. VBA [PF10]. VBI [Can24]. VBI-2D [Can24]. VCH [Res19]. VDSL [Bin00]. ve [Yuk96]. Vector [Cze95, Kat09, Kec01, Van97, Van00a]. Vectorization [Ano02a]. Vectorized [Kok07, MPVA25, Sha08c, Sha09]. Vectorizing [LPD^{+}17]. Vectors [BV18]. VEE [AH05]. Vehicle [GPR14, KSP11, RA95, BEK99, Can24, Chu00, LL03, SAE95b, SAE95b]. vehicle-bridge [Can24]. vehicles [Can22b, Zen97]. Vehicular [TBH21]. Velocimetry [TS14, TS21, BGGL20, JSB20]. VEMcomp [FMS25]. venue [Ano96v]. VEqMon2D [Can22b]. Verfahren [Mei05]. vergence [Hun98]. Verification [BBB12, BM09, XWZ^{+}23, ARRY01, Can24, GRH^{+}21, Rum01, WB16, XWZ^{+}22, ZZC^{+}08]. Verlag [Ano99b, Ano99c, Res19, Sha04]. versatile [Rov90, ZSW^{+}17]. Version [Con95b, EI05, GR93, HL95, Han99, Hig95, Hig02c, Him23, The93c, The97, MVM97, ZJKS23, Eat00, HL97, HTNFBS06a, HTNFBS06b, Han07, KMLP^{+}23, The93b, Mat95a, Mat97, MLB87c, TTT99, VWHK96]. Versions [CFGG94, FGCG94]. Very [USE94, Mol95b, Mol95a]. vessel [HKF^{+}20]. VHLL [USE94]. via [Ano96q, JS04, Kum14, LYH^{+}16, LCMCD22, RBC20, RBZ96, SR09, TYL^{+}16, TQ96, VA94, WR16, WR17, ZLMQ23]. Viable [ES10, Ste08, Ste13]. Vibration [HC00, Hat01, How95, Inm94, Ben98a, Ben04, Gri95, Inm01, Sas96, TD98]. Vibrations [Mol98a, New94, Kel00, Mei01, Rao04]. Victoria [Ano95d]. Vidakovic [O’B13]. Video [AC96, GJ20, WOZ02, Thy10]. Vienna [BH95]. Villard [AO95]. vine [Cob21]. Virginia [IEE97d]. Virtual [FMS25, BN00, HCBAEC23, HSR01, Sut17]. viscoelastic [GSM95]. viscosity [ZZG^{+}14]. Vision [Ano97a]. Visits [Dut16]. VisSim [JH96]. Vista [Ano96u, IEE94d]. visual [AI21, dP96a]. Visualisation [TH01]. Visualisations [RS15a]. Visualization [Ano99d, BPB96, EI05, HL95, JH97, The92d, The92a, The92c, The92b, The93a, The93b, The93c, The97, Nak96, Spe95, Spe96, Ano95l, Mat95a, Mol88, Nak02, WA96, Zek17]. VMAP [WW94]. VMS [MLB87c]. Voigt [Zag16, ZA11]. Vol [Cou93]. Volatility [ST24]. Volume [Sch94, Bal19, BS95a, BLM97, FB95a]. Vömel [Mei05]. vortex [CZ17]. vs [Ano97c]. VSC [ZW93a]. VSP [FY18]. VSS [IEE96b]. VXI [FPBO98].

W [Ano95c, Ano09, Han06, Jai09, BF95, Mro95a, Szy93, Brz97]. Wahrscheinlichkeitsrechnung [Beu05]. walk [Mil20]. Wall [Ano96i]. Walt [IEE94d, IEE94f]. Walter [Ste96]. warfare [Sch99]. Wartak [Iro15, Mun13]. Warwick [Ano94d]. Washington [Gra93]. WaspSim [CH23]. Water [MLAS21, SU93, BHC^{+}24, Cou93]. Wave [Ano98c, Mol03b, Spe95, Spe96, GMT96, GS12, JZW^{+}22, RCT20, Vit11, WPK^{+}18, Wet20]. Wavefront [BPB96]. Wavelet [JNR19, KDAB19, MMOP95, New94, PSR16, RB98, Stu96, TB95, BGG98, tC04, JL01, MG13, PPS06, PF07, Wal02]. wavelet-Galerkin [PPS06, PF07]. Wavelets [DACV95, SN96, BN01, BGG98, GC99, Teo98, AO95]. WavePacket [SL17, SH18, SKA19]. waves [Kno00, TKD00]. way [OAKS11, WW99]. Ways [Hig02a, MV03, Boy15]. weak [GMT96]. Weakly [WB12]. Weakly-Typed [WB12]. Wear [Ano99d]. web [Kea17, Ano97k, Bra97b, KH00, TM97, VFV13, PSTO97]. Web-based [TM97]. Wednesday [Ano93c]. Weight [Zla17, MMD24, MS14]. Weinheim [Res19]. Wellesley [Ano95c]. Welsch [MS14]. Wendy [Alv11, Ano12, Jer06, Kus02, Kus06, Mar19, Ni22, Wik04, Mar14]. WGMode [Bal19]. wheel [GMT95, Zen97]. where [Ano95n]. whispering [Bal19]. WI [Ano95j]. Wide [PSTO97, DG04, TYL^{+}16]. Width [LJR93]. Wiley [Ano00, Res19]. Wiley-VCH [Res19]. William [Lum92]. Wilson [Web97]. WiMAX [Kum14]. WinBUGS [O’B13, Vid11]. Wind [MNHH94, KC95]. Windows [Bai05, FSC95]. Windup [BA94]. Winston [SAE95a]. Winston-Salem [SAE95a]. Winter [IEE96a]. wired [Che04]. wired/wireless [Che04]. Wireless [MZ03, Che04, Gro05, Tra04]. Wisconsin [IEE97b]. WITec [HL22]. within [AH09, FR96, OG95]. without [CW05d]. WITio [HL22]. Women [Shu17]. Work [Mol04d]. Workbench [Nah01]. workflow [FM24]. WorkPlace [Ano97j, Bra97a]. Workshop [BB97, C^{+}97, DG96, EM94, IEE93a, IEE94b, IEE94a, IEE96b, POVD96, FB95a, H^{+}96, Ken95, Zem97, Wie94]. workstation [TB95]. Workstations [The90, The93c, The92d, The92a, The92b, ML90]. World [BEH94, GCP97, IEE94f, Lak97, Mol02b, TH01, W^{+}97, ARR02, Bat99, Gra94, Coo95, PSTO97]. Wrapping [GRDL^{+}12]. write [SBL^{+}17]. Written [But92, IZBT21]. Wrocław [Bub95]. wspomaganie [Ano94g, Szy93]. wykorzystania [BF95].

X [ANM01, Fis19, LLLW06, Tay99]. X-ray [LLLW06, ANM01, Fis19]. X10 [KH14]. xMAS [ZL17]. xMAS-Based [ZL17]. xtmultcorr [SR09]. xxv [Mar19].

Y. [Tay99]. YADPF [MKU22]. Yaroslavsky [Bar16b]. Years [MV03, Mol95b]. York [Ano96b, Ano99b, Ano99c, Ano00, IEE95b]. Yosemite [IEE94b]. YSCODE [FGCG94].

Z [AM17, Brz97, Edw09]. Z-Domain [AM17]. Zero [Mol97]. ziggurat [MT00, Mol01]. Zitarelli [Ano96a].

References

Austerlitz:2003:DAT

[AatupcA03] Howard Austerlitz and Howard Data acquisition techniques using personal computers Austerlitz. Data acquisition techniques using PCs. Academic Press, New York, NY, USA, second edition, 2003. ISBN 0-12-068377-6. xii + 416 pp. LCCN TK7888.3 .A8718 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els031/2002107710.
html;   http://www.loc.
gov/catdir/toc/els031/
2002107710.html.

Andersen:1996:CSU

[AB96] P. K. Andersen and G. Bjedov. Chemical stoichiometry using MATLAB. In Iskander et al. [I^{+}96], pages 612–615. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Ambardar:1998:MDC

[AB98] Ashok Ambardar and Craig Borghesani. Mastering DSP concepts using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-534976-1. x + 72 pp. LCCN TK5102.9.A4524 1998.

Anton:2003:CLA

[AB03] Howard Anton and Robert C.

Busby. Contemporary linear algebra. Wiley, New York, NY, USA, 2003. ISBN 0-471-16362-7. xviii + 594 + 40 pp. LCCN QA184.2 .A58 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;

                                        http://www.loc.gov/catdir/
                                        toc/wiley032/2003266717.
                                        html.

Acakpovi:2013:NPM

[AB13] Amevi Acakpovi and Essel Ben Hagan. Novel photovoltaic module modeling using Matlab/Simulink. International Journal of Computer Applications, 83(??): 27–32, December 2013. CO-DEN ???? ISSN 0975-

                                        8887. URL https://www.
                                        ijcaonline.org/archives/
                                        volume83/number16/14535-
                                        2978/.

Antonietti:2025:LDP

[ABB^{+}25] Paola F. Antonietti, Stefano Bonetti, Michele Botti, Mattia Corti, Ivan Fumagalli, and Ilario Mazzieri. lymph: Discontinuous polytopal methods for multiphysics differential problems. ACM Transactions on Mathematical Software, 51(1):3:1–3:??, March 2025. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Abdi:1994:RN

[Abd94] Hervé Abdi. Les réseaux de neurones. (French) [Neural networks]. Sciences et technologies de la connaissance. Presses universitaires de Grenoble, Grenoble, France, 1994. ISBN 2-7061-0554-2. iv + 269 pp. LCCN QA76.87 .A33 1994. 160F.

Abonyi:2003:FMI

[Abo03]  Janos Abonyi.  Fuzzy

model identification for control. Birkhäuser Boston Inc., Cambridge, MA, USA, 2003. ISBN 0-8176-4238-2 , 3-7643-4238-2. x + 273 pp. LCCN TJ213 .A222 2003.

Alvarez:2017:PDT

[ÁBZ17] Inmaculada C. Álvarez, Javier Barbero, and José L. Zofío. A panel data toolbox for MATLAB. Journal of Statistical Software, 76(??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v076i06;
https://www.jstatsoft.
org/index.php/jss/article/
view/v076i06/v76i06.pdf.

Alvarez:2020:DEA

[ÁBZ20] Inmaculada C. Álvarez, Javier Barbero, and José L. Zofío. A data envelopment analysis toolbox for MAT-LAB. Journal of Statistical Software, 95(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v095i03;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v095i03/
v95i03.pdf.

Arnold:1996:PCD

[AC96] John F. Arnold and Michael C.

Cavenor. A practical course in digital video communications based on MATLAB. IEEE Transactions on Education, 39(2):127–136, May 1996. CODEN IEEDAB. ISSN 0018-9359.

Aunon:1997:IPR

[AC97] Jorge Aunon and V. Chandrasekar. Introduction to probability and random processes. Communications and signal processing. Mc-Graw-Hill, New York, NY, USA, 1997. ISBN 0-07-001563-5. xv + 532 pp. LCCN QA273.A86 1996.

Austin:1999:IEP

[AC99] Mark Austin and David Chancogne. Introduction to engineering programming: in C, Matlab and Java. Wiley, New York, NY, USA, 1999. ISBN 0-471-00116-3. xv + 656 pp. LCCN QA76.73.C15 A98 1999. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley034/98008481.
html;   http://www.loc.
gov/catdir/toc/onix05/
98008481.html.

Alberty:1999:RAL

[ACF99] Jochen Alberty, Carsten Carstensen, and Stefan A. Funken. Remarks around 50 lines of Matlab: short finite element implementation. Numerical Algorithms, 20(2-3):117–137, 1999. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Alberty:2002:MIF

[ACFK02] J. Alberty, C. Carstensen, S. A. Funken, and R. Klose. Matlab implementation of the finite element method in elasticity. Computing, 69(3): 239–263, November 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Acha:2004:FMS

[Ach04] Enrique Acha, editor. FACTS:

modelling and simulation in power networks. Wiley, New York, NY, USA, 2004. ISBN 0-470-85271-2. xv + 403 pp. LCCN TK3148 .F33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley041/2004299463.
html;   http://www.loc.
gov/catdir/toc/wiley041/
2004299463.html.

ACM:1996:FCP

[ACM96] ACM, editor. FCRC ’96:

Conference proceedings of the 1996 International Conference on Supercomputing: Philadelphia, Pennsylvania, USA, May 25–28, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number: 415961.

ACM:1997:SHP

[ACM97] ACM, editor. SC’97: High Performance Networking and Computing: Proceedings of the 1997 ACM/IEEE SC97 Conference: November 15–21, 1997, San Jose, California, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-89791-985-8. LCCN ???? URL http://

                                        www.supercomp.org/sc97/
                                        proceedings/.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

ACM:1998:SHP

[ACM98] ACM, editor. SC’98: High Performance Networking and Computing: Proceedings of the 1998 ACM/IEEE SC98 Conference: Orange County Convention Center, Orlando, Florida, USA, November 7–13, 1998. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN ???? LCCN ???? URL http://

                                        www.supercomp.org/sc98/
                                        papers/.

ACM:2003:SII

[ACM03] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

Antoine:2014:GMT

[AD14] Xavier Antoine and Romain Duboscq. GPELab, a Matlab toolbox to solve Gross–Pitaevskii equations I: Computation of stationary solutions. Computer Physics Communications, 185(11):2969–2991, November 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465514002318.

Antoine:2015:GMT

[AD15] Xavier Antoine and Romain Duboscq. GPELab, a Matlab toolbox to solve Gross– Pitaevskii equations II: Dynamics and stochastic simulations. Computer Physics Communications, 193(??): 95–117, August 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465515001113.

Adachi:1997:SIC

[Ada97] Shuichi Adachi. System Identification for Control Using MATLAB (Japanese). Tokyo Denki Publishing, Tokyo, Japan, 1997. ISBN 4-254-11073-1. ???? pp. LCCN ????

Amestoy:2004:AAA

[ADD04] Patrick R. Amestoy, Timothy A. Davis, and Iain S. Duff. Algorithm 837: AMD, an approximate minimum degree ordering algorithm. ACM Transactions on Mathematical Software, 30(3):381–388, September 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Akman:2023:BMP

[ADW23] Ozgur E. Akman, Kevin Doherty, and Benjamin J. Wareham. BDEtools: a MATLAB package for Boolean delay equation modeling. Journal of Computational Biology, 30(1): 52–69, January 2023. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (electronic). URL https://

                                        www.liebertpub.com/doi/
                                        abs/10.1089/cmb.2021.0658;
                                         https://www.liebertpub.
                                        com/doi/reader/10.1089/
                                        cmb.2021.0658.

Aurada:2014:HMI

[AEF^{+}14] Markus Aurada, Michael Ebner, Michael Feischl, Samuel Ferraz-Leite, Thomas Führer, Petra Goldenits, Michael Karkulik, Markus Mayr, and Dirk Praetorius.

HILBERT — a MATLAB implementation of adaptive 2D-BEM. Numerical Algorithms, 67(1):1–32, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/s11075-013-9771-
2.pdf.

Aliev:2012:PLC

[AFL^{+}12] Alexander Aliev, Valerii Fedorov, Sergei Leonov, Brian McHugh, and Mindy Magee. PkStaMp library for constructing optimal population designs for PK/ PD studies. Communications in Statistics: Simulation and Computation, 41 (6):717–729, 2012. CODEN CSSCDB. ISSN 0361-0918.

Agulhari:2019:ARL

[AFOP19] Cristiano M. Agulhari, Alexandre Felipe, Ricardo C. L. F. Oliveira, and Pedro L. D. Peres. Algorithm 998: The robust LMI parser — a toolbox to construct LMI conditions for uncertain systems. ACM Transactions on Mathematical Software, 45(3):36:1– 36:25, August 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3323925.

Aarnes:2007:INF

[AGL07] Jørg E. Aarnes, Tore Gimse, and Knut-Andreas Lie. An introduction to the numerics of flow in porous media using Matlab. In Geometric modelling, numerical simulation, and optimization: applied mathematics at SINTEF, pages 265–306. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007.

Atkinson:2004:ENA

[AH04] Kendall E. Atkinson and Weimin Han. Elementary numerical analysis. Wiley, New York, NY, USA, third edition, 2004. ISBN 0-471-43337-3. xvi + 560 pp. LCCN QA297 .A83 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        toc/wiley032/2003053836.
                                        html.

Angus:2005:VPP

[AH05] Robert B. (Robert Brownell) Angus and Thomas E. Hulbert. VEE Pro: practical graphical programming. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-85233-870-9. various pp. LCCN QA76.65 .A54 2005.

Augusta:2009:PLN

[AH09] P. Augusta and Z. Hurak. Polmat library now within Symbolic Math Toolbox for Matlab in multidimensional systems computations. In IEEE, editor, International Workshop on Multidimensional (nD) Systems, 2009. nDS 2009, Aristotle University of Thessaloniki, Thessaloniki, Greece, June 29 2009–July 1 2009, pages 1–3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-2797-5. LCCN ???? IEEE catalog number CFP09612.

Anakwa:1994:ICC

[AHS94] Winfred K. N. Anakwa, Chris Hoadley, and Thomas L. Stewart. Implementation of communications and control algorithms on TMS320C30 using MATLAB. Computers in education journal, 4 (2):41–47, April–June 1994. CODEN CEJOE7. ISSN 1069-3769.

Atanasiu:2021:DTM

[AI21] Vlad Atanasiu and Rolf Ingold. Document Towers: a MATLAB software implementing a three-dimensional architectural paradigm for the visual exploration of digital documents and libraries. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021000297.

Arevalo:2020:SPA

[AJGO^{+}20] Carmen Arévalo, Erik Jonsson-Glans, Josefine Olander, Monica Selva Soto, and Gustaf Söderlind. A software platform for adaptive high order multistep methods. ACM Transactions on Mathematical Software, 46(1):2:1–2:17, March 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3372159.

Ajjarapu:1995:ABC

[Ajj95] Venkataramana Ajjarapu. Application of bifurcation and continuation methods for the analysis of power system dynamics. IEEE Conference on Control Applications — Proceedings, pages 52–56, 1995. CODEN ICOAE8. ISSN 1085-1992.

Akay:2000:NBS

[Aka00] Metin Akay, editor. Nonlinear biomedical signal processing. IEEE Press series on biomedical engineering. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7803-6011-7 (vol. 1), 0-7803-6012-5 (vol. 2). ???? pp. LCCN R857.S47 N66 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley043/00027777.
html;   http://www.loc.
gov/catdir/description/
wiley036/00027777.html;
 http://www.loc.gov/catdir/
toc/onix07/00027777.html.

Ariaans:1994:DRI

[AKB94] L. J. J. M. Ariaans, M. Keulers, and J. W. J. J. Beckers. Design and realtime implementation of a controller in MatLab. In Anonymous [Ano94i], pages 883–886. ISBN 0-7803-2024-7. LCCN TK 7870 K37 1994.

Aldeen:1996:IIC

[Ald96] M. (Mohammad) Aldeen, editor. IEEE International Conference on Multi-Media Engineering Education (2nd: July 1996: University of Melbourne). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3173-7, 0-7803-3174-5. LCCN T65.3.I35 1996.

Ali:2002:DAL

[Ali02] Irfan Ali, editor. Doppler applications in LEO satellite communication systems, volume 656 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 2002. ISBN 0-7923-7616-1. xii + 113 pp. LCCN TK5104 .D57 2002.

Alonso:2006:JCA

[Alo06] Andrés M. Alonso. JMASM 23: Cluster analysis in epidemiological data (Matlab). Journal of Modern Applied Statistical Methods, 5(1):273–280, May 2006. CODEN ???? ISSN 1538-9472. URL https:

                                        //jmasm.com/index.php/
                                        jmasm/article/view/243.

Altena:1995:PCO

[Alt95] G. Altena. Process control optimisation by means of process identification using PRBS-signals and Matlab. In Anonymous [Ano95n], pages 403–412. ISBN ???? LCCN ???? Seven volumes.

Altman:2012:USM

[Alt12] Yair M. Altman. Undocumented secrets of MATLAB– Java programming. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-6904-9 (e-book), 1-4398-6903-0 (hardback), 1-4398-6903-0. xxi + 663 + 16 pp. LCCN QA297 .A544 2012.

AlvarezP:2011:SBR

[Alv11] Adolfo Alvarez P. Short book review: Exploratory Data Analysis with MAT-LAB, Second Edition, by Wendy L. Martinez, Angel R. Martinez, Jeffrey L. Solka. International Statistical Review = Revue Internationale de Statistique, 79(3):492, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:/

/www.jstor.org/stable/
41305073.

Al-Mothafar:1995:STB

[AM95] M. R. D. Al-Mothafar.

Small-signal and transient behaviour of a two-module DC-DC converter using mutually coupled output filter inductors. International Journal of Electronics Theoretical & Experimental, 79(6):917–932, December 1995. CODEN IJELA2. ISSN 0020-7217.

Acha:2001:PSH

[AM01] Enrique Acha and Manuel Madrigal. Power systems harmonics: computer modelling and analysis. Wiley, New York, NY, USA, 2001. ISBN 0-471-52175-2. xvi + 365 pp. LCCN TK3226 .A24 2001.

ALgamluoli:2017:CBC

[AM17] Ammar ALgamluoli and Muatafa Falah Mahmood. Controller for boost converter in S-domain and Zdomain by using SISO-TOOL MATLAB. International Journal of Computer Applications, 174(??):3–7, September 2017. CODEN ???? ISSN 0975-8887. URL

                                        https://www.ijcaonline.
                                        org/archives/volume174/
                                        number9/28433-2017915244/

.

Ahmadinejad:2022:EQE

[AM22] Mohammad Ahmadinejad and Mohammad Hossein Moaiyeri. Energy- and quality-efficient approximate multipliers for neural network and image processing applications. IEEE Transactions on Emerging Topics in Computing, 10 (2):1105–1116, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ambardar:1995:ADS

[Amb95] Ashok Ambardar. Analog and Digital Signal Processing. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1995. ISBN 0-534-94086-2. xx + 700 pp. LCCN TK 5102.9 A45 1995.

Ambardar:1999:ADS

[Amb99] Ashok Ambardar. Analog and digital signal processing. Brooks/Cole, Pacific Grove, CA, USA, second edition, 1999. ISBN 0-534-95409-X. xvii + 807 pp. LCCN TK5102.9 .A45 1999.

Azzini:2018:DMP

[AMR18] Ivano Azzini, Ronal Muresano, and Marco Ratto. Dragonfly: a multi-platform parallel toolbox for MAT-LAB/Octave. Computer Languages, Systems and Structures, 52(??):21–42, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1477842417300660.

Anderson:1995:DAT

[And95] G. Anderson. Design of acoustic transducers using MATLAB. In SAE [SAE95a], pages 6–16. ISBN 1-56091-716-4. LCCN TL255.N393 1995.

Anderson:1999:DTE

[And99] John B. Anderson. Digital transmission engineering. IEEE series on mobile and digital communication. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7803-3457-4 , 0-13-082961-7 (Prentice Hall). xx + 369 pp. LCCN TK5103.7 .A534 1999.

Anderson:2005:BEF

[And05] Patrick L. Anderson. Business economics and finance with MATLAB, GIS and simulation models. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-348-0. xxii + 472 pp. LCCN HD30.22 .A53 2005.

Als-Nielsen:2001:EMX

[ANM01] Jens Als-Nielsen and Des McMorrow. Elements of modern X-ray physics. Wiley, New York, NY, USA, 2001. ISBN 0-471-49857-2 (hardcover), 0-471-49858-0 (paperback). xi + 318 pp. LCCN QC481 .A47 2001. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley035/00063342.
                                        html;   http://www.loc.
                                        gov/catdir/toc/onix07/
                                        00063342.html.

Anonymous:1988:MIA

[Ano88] Anonymous. MATLAB integrates analysis on Mac. Computers in Physics, 2(6): 96–??, November 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822803.

Anonymous:1992:S

[Ano92] Anonymous. Simulink 1.2. Macworld, 9(12):268–??, December 1992. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1993:UPE

[Ano93a] Anonymous, editor. 28th Universities power engineering conference: September 1993, Stafford, UK, volume 2. Staffordshire University, Stafford, UK, 1993. ISBN ???? LCCN ???? Two volumes.

Anonymous:1993:BRCc

[Ano93b] Anonymous. Book review:

Contemporary linear systems using MATLAB(R) (computer disks included): R. D. Strum and D. E. Kirk, PWS, Boston, MA, 1994. 673 pp. (paper), US$34.50, ISBN 0-534-93273-8. Mathematics and Computers in Simulation, 35(6):527, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/0378475493900767.

Anonymous:1993:CCD

[Ano93c] Anonymous, editor. Computing and Control Division colloquium on bringing added value to continuous system simulation models: Wednesday, 3 November 1993, London, number 201 in Colloquium Digest

— IEE. IEE, London, UK, 1993. ISBN ???? LCCN ????

Anonymous:1993:ICE

[Ano93d] Anonymous, editor. International conference on electrical machines in Australia

— 1993 Sep: Adelaide; Australia. University of South Australia, Adelaide, South Australia, Australia, 1993. ISBN ???? LCCN ????

Anonymous:1993:PIC

[Ano93e] Anonymous, editor. Proceedings of the 4th International Conference on Signal Processing Applications and Technology: September 28–October 1, 1993, Santa Clara, CA, USA. DSP Associates, Newton, MA, USA, 1993. ISBN ???? LCCN ???? Two volumes.

Anonymous:1994:EER

[Ano94a] Anonymous, editor. EU-RISCON’94: European robotics and intelligent systems conference — August 1994, Malaga, Spain, volume 2. University of Bristol, Bristol, UK, 1994. ISBN ???? LCCN ???? Three volumes.

Anonymous:1994:FAJ

[Ano94b] Anonymous, editor. Flexible automation: Japan-USA symposium — July 1994, Kobe, Japan, volume 3. Institute of Systems, Control and Information Engineers, Kyoto, Japan, 1994. ISBN ???? LCCN ???? Three volumes.

Anonymous:1994:IUS

[Ano94c] Anonymous, editor. Identification of uncertain systems: Colloquium, April 1994, London, UK, number 105 in Colloquium Digest — IEE. IEE, London, UK, 1994. ISBN ???? LCCN ????

Anonymous:1994:ICC

[Ano94d] Anonymous, editor. International Conference on Control ’94: 21–24 March 1994: University of Warwick, Coventry, UK, number 389 in IEE Conference Publication. IEE, London, UK, 1994. ISBN ???? ISSN 0537-9989. LCCN ???? Two volumes.

Anonymous:1994:SAM

[Ano94e] Anonymous, editor. Likelihood, Bayesian inference and their application to the solution of new structures: Symposium: Annual meeting — June 1994, Atlanta, GA, USA, volume 22. American Crystallographic Association, Pittsburgh, PA, USA, 1994. ISBN ???? ISSN 0569-4221. LCCN ????

Anonymous:1994:M

[Ano94f]  Anonymous.  MATLAB

4.1. Macworld, 11(11):81– ??, November 1994. CO-DEN MACWEA. ISSN 0741-8647.

Anonymous:1994:PKW

[Ano94g] Anonymous, editor. Polioptymalizacja i komputrowe wspomaganie projectowania mielno: Materialy 12 ogolnopolskiej konferencji, volume 18 of Zeszyty Naukowe — Wyzsza Szkola Inzynierska w Koszalinie Wydzial Mechaniczny. Wyzsza Szkola Inzynierska w Koszalinie Wydzial Mechaniczny, Koszalin, Poland, 1994. ISBN ???? ISSN 0860-0325. LCCN ????

Anonymous:1994:ECI

[Ano94h] Anonymous, editor. Proceedings /EUFIT ’94. Second European Congress on Intelligent Techniques and Soft Computing, Aachen, Germany, September 20–23, 1994, volume 1. Verl. der Augustinus-Buchh., Aachen, Germany, 1994. ISBN 3-86073-286-2. LCCN ???? Three volumes.

Anonymous:1994:SPS

[Ano94i] Anonymous, editor. SICE ’94, proceedings of the 33rd SICE annual conference. International session: Tokyo Metropolitan Institute of Technology, July 26–28, 1994. Society of Instrument and Control Engineers, Tokyo, Japan, 1994. ISBN 0-7803-2024-7. LCCN TK 7870 K37 1994.

Anonymous:1994:SF

[Ano94j] Anonymous. Software forum. Design news, 49(17): 150–??, September 1994. CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1994:ICS

[Ano94k] Anonymous, editor. Systems engineering: 10th International conference — September 1994, Coventry, UK, volume 1. Coventry University, Coventry, UK, 1994. ISBN 0-905949-23-4. LCCN ???? Two volumes.

Anonymous:1994:TDS

[Ano94l] Anonymous, editor. Transducer design and simulation using computers: Colloquium — September 1994, Southampton, UK. USITT, Southampton, UK, 1994. ISBN ???? LCCN ????

Anonymous:1995:

[Ano95a]  Anonymous. ????  In

Mościński and Ogonowski [MO95], page ?? ISBN 0-13-309667-X. LCCN QA402.3.A33 1995.

Anonymous:1995:ACT

[Ano95b] Anonymous, editor. Applied control techniques using MATLAB: Colloquium — January 1995, London, UK, number 14 in IEE Colloquium Digest. IEE, London, UK, 1995. ISBN ???? LCCN ????

Anonymous:1995:BRMh

[Ano95c] Anonymous. Book review:

Matrix algebra: Using MIN-Imal MATlab^{T M}: By Joel W. Robbin. A. K. Peters, Wellesley, MA. (1995). 544 pages. $59.95 (diskette included)ters, Wellesley, MA. (1995). 544 pages. Computers and Mathematics with Applications, 29(11): 110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122195902740.

Anonymous:1995:CMC

[Ano95d] Anonymous, editor. Control 95: meeting the challenge of Asia Pacific growth: preprints: 23–25 October 1995, Old Melbourne Hotel, Melbourne, Victoria, volume 95/09 of National Conference Publication — Institution of Engineers Australia NCP 1995. Institution of Engineers, Australia, Barton, A.C.T., Australia, 1995. ISBN 0-85825-631-

2. ISSN 0313-6922. LCCN ????

Anonymous:1995:CSD

[Ano95e] Anonymous. Control system design and analysis using MATLAB and SIMULINK. In Investing in the future: 1995 Annual Conference proceedings, American Society for Engineering Education, June 25–28, 1995, Anaheim, California [Ano95k], pages 2518–2522. ISBN ???? LCCN ???? Two volumes.

Anonymous:1995:EEC

[Ano95f] Anonymous, editor. EU-FIT ’95: 3rd European congress on intelligent techniques and soft computing — August 1995, Aachen, Germany, volume 3 of EU-FIT — European Congress — 1995. Elite Foundation, Aachen, Germany, 1995. ISBN 3-930911-67-1 (??invalid ISBN checksum??). LCCN ???? Three volumes.

Anonymous:1995:FM

[Ano95g] Anonymous. Fuzzy-Toolbox für Matlab. Elektronik, 44 (7):76–??, ???? 1995. CO-DEN EKRKAR. ISSN 0013-5658.

Anonymous:1995:ICC

[Ano95h] Anonymous, editor. IEE Computing and Control Division Colloquium on Applied Control Techniques Using Matlab, number 014 in IEE Colloquium (Digest). IEE, London, UK, 1995. CODEN DCILDN. ISBN ???? ISSN 0963-3308. LCCN ????

Anonymous:1995:ICM

[Ano95i] Anonymous, editor. International Conference on Electric Railways in a United Europe: 27–30 March 1995, number 405 in IEE Conference Publication. IEE, London, UK, 1995. ISBN 0-85296-631-8. ISSN 0537-9989. LCCN TF1055 .I57 1995.

Anonymous:1995:IOP

[Ano95j] Anonymous, editor. International off-highway and powerplant congress, September 1995, Milwaukee, WI, USA, number 952144 in Papers — Society of Automotive Engineers New York. Society of Automotive Engineers International, New York, NY, USA, 1995. ISBN ???? ISSN 0148-7191. LCCN ????

Anonymous:1995:IFA

[Ano95k] Anonymous, editor. Investing in the future: 1995 Annual Conference proceedings, American Society for Engineering Education, June 25–28, 1995, Anaheim, California. American Society for Engineering Education, Washington, DC, USA, 1995. ISBN ???? LCCN ???? Two volumes.

Anonymous:1995:MCP

[Ano95l]  Anonymous.  MATLAB:

a cumbersome programming environment and commandline interface hamper this tool for numerical analysis and scientific visualization. MacUser, 11(10):68– ??, ???? 1995. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1995:MRC

[Ano95m] Anonymous. Matlab review: Complex maths calculations? no problem for Matlab. Electronics world + wireless world, ??(3):197– ??, March 1995. CODEN EWWWE6. ISSN 0959-8332.

Anonymous:1995:PGE

[Ano95n] Anonymous, editor. Powergen Europe ’95: where the industry connects: Conference — May 1995, Amsterdam, The Netherlands, volume 6. Power Gen Europe, Utrecht, The Netherlands, 1995. ISBN ???? LCCN ???? Seven volumes.

Anonymous:1995:RAI

[Ano95o] Anonymous, editor. Robots for Australian industries: proceedings of the 1995 National Conference of the Australian Robot Association, Melbourne, 5–7 July. Australian Robot Association Inc., Sydney, NSW, Australia, 1995. ISBN 0-9587583-0-1. LCCN ????

Anonymous:1995:SRR

[Ano95p] Anonymous. Solving RL and RC circuits using MAT-LAB. In Investing in the future: 1995 Annual Conference proceedings, American Society for Engineering Education, June 25–28, 1995, Anaheim, California [Ano95k], pages 2241–2246. ISBN ???? LCCN ???? Two volumes.

Anonymous:1995:STM

[Ano95q]  Anonymous.  Subband/

transform MATLAB functions for processing images. NASA tech briefs, 19(11): 58–??, ???? 1995. CODEN NSTBAT. ISSN 0145-319X.

Anonymous:1995:TDF

[Ano95r] Anonymous. Teaching digital filter design using MAT-LAB and Derive. In Investing in the future: 1995 Annual Conference proceedings, American Society for Engineering Education, June 25–28, 1995, Anaheim, California [Ano95k], pages 618–621. ISBN ???? LCCN ???? Two volumes.

Anonymous:1995:PIC

[Ano95s] Anonymous, editor. The Proceedings of the 6th International Conference on Signal Processing Applications & Technology: October 24– 26, 1995, Boston, Massachusetts, USA, volume 2. DSP Associates, Waltham, MA, USA, 1995. ISBN ???? LCCN ????

Anonymous:1995:TDS

[Ano95t] Anonymous, editor. The teaching of digital signal processing (DSP) in universities: Colloquium — February 1995, London, UK, number 35 in IEE Colloquium Digest. IEE, London, UK, 1995. ISBN ???? LCCN ????

Anonymous:1995:UMI

[Ano95u] Anonymous. Using MAT-LAB to illustrate the phenomenon of aliasing. In Investing in the future: 1995 Annual Conference proceedings, American Society for Engineering Education, June 25–28, 1995, Anaheim, California [Ano95k], pages 612–617. ISBN ???? LCCN ???? Two volumes.

Anonymous:1995:UMT

[Ano95v] Anonymous. Using MAT-LAB to teach mathematical modeling and simulation to associate in applied science in E.E. technical students. In Investing in the future: 1995 Annual Conference proceedings, American Society for Engineering Education, June 25–28, 1995, Anaheim, California [Ano95k], pages 2311–2315. ISBN ???? LCCN ???? Two volumes.

Anonymous:1996:BRLh

[Ano96a] Anonymous. Book review:

Linear algebra labs with MATLAB(R) (second edition): By David R. Hill and David E. Zitarelli. Prentice Hall, Upper Saddle River, NJ. (1996). $22.33. Computers and Mathematics with Applications, 32(6):133, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812219690231X.

Anonymous:1996:BRNb

[Ano96b] Anonymous. Book review: Numerical methods using Matlab: By John Penny and George Lindfield. Ellis Horwood, New York. (1995). 328 pages. $35.00. Computers and Mathematics with Applications, 31(7): 147, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122196904733.

Anonymous:1996:CCC

[Ano96c]  Anonymous.  Computa-

tional chains for CACSD using Matlab containers. Proceedings of the IEEE International Symposium on Computer-Aided Control System Design, pages 392– 397, 1996. IEEE catalog number 96TH8136.

Anonymous:1996:DIF

[Ano96d] Anonymous. Design of an industrial flexible robot controller using Matlab. Computers & industrial engineering, 31(1-2):131–134, October 1996. CODEN CINDDL. ISSN 0360-8352 (print), 1879-0550 (electronic).

Anonymous:1996:EFE

[Ano96e] Anonymous, editor. EU-FIT ’96: fourth European Congress on Intelligent Techniques and Soft Computing, Aachen, Germany, September 2–5, 1996: Proceedings, volume 3. ELITE-Foundation, Aachen, Germany, 1996. ISBN 3-89653-187-5. LCCN ????

Anonymous:1996:FLC

[Ano96f] Anonymous, editor. Fuzzy logic controllers in practice: Colloquium — November 1996, London, number 200 in Colloquium Digest — IEE 1996. IEE, London, UK, 1996. ISBN ???? ISSN 0963-3308. LCCN ????

Anonymous:1996:INS

[Ano96g] Anonymous, editor. Israel Nuclear Societies: Conference; 19th — December 1996, Herzliya, Israel, Atomic Energy Commission USA -Reports- 1996; INISmf-15507. ????, ????, 1996. ISBN ???? LCCN ????

Anonymous:1996:MAD

[Ano96h] Anonymous. Mathworks, Applix develop interface for Matlab, Applixware. Silicon Graphics World, 6(6): 14, June 1996. ISSN 1057-7041.

Anonymous:1996:MLA

[Ano96i] Anonymous. MathWorks links to Applix on Wall Street. SunExpert Magazine, 7(10):16, October 1996. ISSN 1053-9239. Describes a product, Applix Link, that will allow the user to access real-time Wall Street financial data feeds, process that information with Matlab, and display the results in an Applixware spreadsheet.

Anonymous:1996:MST

[Ano96j]  Anonymous.  MATLAB

software tool for the introduction of speech coding fundamentals in a DSP course. Proceedings of the International Conference on Acoustics, Speech, and Signal Processing, 2:1133–1136, 1996. CODEN IPRODJ. ISSN 0736-7791. IEEE catalog number 96CB35903.

Anonymous:1996:MTF

[Ano96k]  Anonymous.  MATLAB

toolbox for fixed-order, mixed-norm control synthesis. IEEE Control Systems Magazine, 16(5):36–44, October 1996. CODEN IS-MAD7. ISSN 0272-1708. Anonymous:1996:MTS

[Ano96l] Anonymous. Matlab toolkit for the study of oscillations in relay systems. Proceedings of the IEEE International Symposium on Computer-Aided Control System Design, pages 286– 291, 1996. IEEE catalog number 96TH8136. Anonymous:1996:MSP

[Ano96m] Anonymous, editor. Multidimensional systems: problems and solutions: Colloquium — January 10, 1996, London, number 3 in Colloquium Digest — IEE 1996. IEE, London, UK, 1996. CODEN DCILDN. ISBN ???? ISSN 0963-3308. LCCN ????

Anonymous:1996:NCT

[Ano96n] Anonymous. NNCTRL — a CANCSD toolkit for MAT-LAB. Proceedings of the IEEE International Symposium on Computer-Aided Control System Design, pages 368–373, 1996. IEEE catalog number 96TH8136.

Anonymous:1996:NTM

[Ano96o]  Anonymous.  NNSYSID

toolbox — a MATLAB toolbox for system identification with neural networks. Proceedings of the IEEE International Symposium on Computer-Aided Control System Design, pages 374– 379, 1996. IEEE catalog number 96TH8136.

Anonymous:1996:OCS

[Ano96p] Anonymous. Octave control systems toolbox: a MATLAB-like CACSD environment. Proceedings of the IEEE International Symposium on Computer-Aided Control System Design, pages 386–391, 1996. IEEE catalog number 96TH8136.

Anonymous:1996:OPF

[Ano96q] Anonymous. Optimal power flow via interior point methods: an educational tool in Matlab. Canadian Conference on Electrical and Computer Engineering, 2: 996–999, 1996. CODEN CCCEFV. ISSN 0840-7789. IEEE catalog number 96TH8157.

Anonymous:1996:PPR

[Ano96r] Anonymous. PARADISE — PArametric Robust Analysis and Design Interactive Software Environment: a Matlab-based robust control toolbox. Proceedings of the IEEE International Symposium on Computer-Aided Control System Design, pages 380–385, 1996. IEEE catalog number 96TH8136.

Anonymous:1996:PSC

[Ano96s] Anonymous, editor. Power systems computation conference: 12th — August 1996, Dresden, Germany, volume 2 of POWER SYS-TEMS COMPUTATION CONFERENCE 1996. Power Systems Computation Conference, ????, 1996. ISBN ???? LCCN ????

Anonymous:1996:PAC

[Ano96t] Anonymous, editor. Progress in applied computational electromagnetics: Annual review; 12th — March 1996, Monterey, CA, ANNUAL REVIEW OF PROGRESS IN APPLIED COMPU-TATIONAL ELECTRO-MAGNETICS 1996; CONF 12/V2. Naval Postgraduate School, Monterey, CA, USA, 1996. CODEN CPCEFK. ISBN ???? LCCN ????

Anonymous:1996:TCM

[Ano96u] Anonymous, editor. Technology in collegiate mathematics: Annual international conference; 7th — November 1994, Lake Buena Vista, FL, PROCEED-INGS OF THE ANNUAL INTERNATIONAL CON-FERENCE ON TECH-NOLOGY IN COLLE-GIATE MATHEMATICS 1994; 7th. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-87020-7. LCCN ????

Anonymous:1996:UIC

[Ano96v] Anonymous, editor. UKACC International Conference on Control ’96, 2–5 September 1996, venue, University of Exeter, UK, volume 427 of Conference publication. IEE, London, UK, 1996. ISBN 0-85296-668-7, 0-85296-666-0. ISSN 0537-9989. LCCN TJ212.2 .U32 1996; TK5.I4 no.427. Two volumes.

Anonymous:1997:CBC

[Ano97a] Anonymous. C++ Builder comes out of the closet, IBM does everything Java, a new vision for client/server development, Novell’s Developer Net 2000, plus Matlab 5.0: serious fun with graphs. .EXE: the software developers’ magazine, 11(8): 7–??, ???? 1997. CODEN EXEEE5. ISSN 0268-6872.

Anonymous:1997:EEC

[Ano97b] Anonymous, editor. EU-FIT ’97: 5th European Congress on Intelligent Techniques and Soft Computing, Aachen, Germany, September 8–11, 1997: Proceedings, volume 2 of EU-FIT -EUROPEAN CONGRESS-1997. ELITE-Foundation, Aachen, Germany, 1997. ISBN 3-89653-200-6. LCCN ????

Anonymous:1997:FLD

[Ano97c] Anonymous. First looks — David meets Goliath: K6 vs. 233-MHz Pentiums. four digital cameras. MatLab 5.0 nuts and bolts. PC Magazine, 16(12):45–??, ???? 1997. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1997:MIM

[Ano97d] Anonymous. The Math-Works introduces MATLAB C++ math library. C/C++ Users Journal, 15(9):97, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:MIS

[Ano97e] Anonymous. MathWorks introduces Simulink 2 nonlinear simulation software. Silicon Graphics World, 7(4): 15, April 1997. ISSN 1057-7041. Anonymous:1997:MAD

[Ano97f] Anonymous. MATLAB 5 adds development tools. Research & Development, 39 (8):44, July 1997. CODEN REDEEA. ISSN 0746-9179. Anonymous:1997:MLB

[Ano97g]  Anonymous.  Matlab 5

(late beta). Personal computer world, 20(2):83–??, ???? 1997. CODEN PC-WODU. ISSN 0142-0232. Anonymous:1997:ACC

[Ano97h] Anonymous, editor. Proceedings of the 1997 American Control Conference: Albuquerque Convention Center, Hyatt Regency Hotel and Doubletree Hotel, Albuquerque, New Mexico, June 4 – June 6, 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-3833-2, 0-7803-3832-4, 0-7803-3834-0, 0-7803-3835-9. ISSN 0743-1619. LCCN TJ212 .A57 1997. Five volumes.

Anonymous:1997:SIT

[Ano97i] Anonymous, editor. Society for Information Technology and Teacher Education: International conference; 8th — April 1997, Orlando, FL, volume 1 of Technology and Teacher Education Annual. Association for the Advancement of Computing in Education, Charlottesville, VA, USA, 1997. ISBN 1-880094-25-8. LCCN ????

Anonymous:1997:TNR

[Ano97j] Anonymous. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–??, October/ December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://dlib.
computer.org/cs/books/
cs1997/pdf/c4075.pdf.

Anonymous:1997:TNF

[Ano97k] Anonymous. Technology news: Fortran 90 news; free software; symbolic computing packages; Matlab 5; Web products. IEEE Computational Science & Engineering, 4(1):87–??, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://dlib.
computer.org/cs/books/
cs1997/pdf/c1087.pdf.

Anonymous:1997:UMB

[Ano97l] Anonymous, editor. The use of model based signal processing techniques in the analysis of biomedical signals: Colloquium — April 1997, London, number 9 in Colloquium Digest — IEE 1997. IEE, London, UK, 1997. ISBN ???? ISSN 0963-3308. LCCN ???? Anonymous:1998:MMU

[Ano98a] Anonymous. Mathematical modeling using MAT-LAB. Technical report ADa358 796, Naval Postgraduate School, Monterey, CA, USA, 1998. 132 pp.

Anonymous:1998:MRG

[Ano98b] Anonymous. MathWorks releases geographic data analysis tool. SunServer, 12(1): 8, January 1998. ISSN 1091-

4986. URL http://www.
crcpress.com.

Anonymous:1998:MIF

[Ano98c] Anonymous. MATLAB implementation of a Fourier approach to optical wave propagation. Technical report AD-a356 153, Naval Postgraduate School, Monterey, CA, USA, 1998. 129 pp.

Anonymous:1999:BRGp

[Ano99a]  Anonymous.  Book re-

view: Graphics and GUIs with Matlab(R), second edition: By Patrick Marchand. CRC Press, Boca Raton, FL. (1999). 445 pages. $39.95, DM 70.00, öS 511.90, sFr 64.00, GBP 27.00. Computers and Mathematics with Applications, 38(11–12):289, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812219991249X.

Anonymous:1999:BRMc

[Ano99b]  Anonymous.  Book re-

view: The Matlab(R) 5 handbook: By Darren Redfern and Colin Campbell. Springer-Verlag, New York. (1998). 488 pages. $34.95, DM 69.00, öS 504.00, sFr 63.00, GBP 19.50. Computers and Mathematics with Applications, 37(3):134, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199904017.

Anonymous:1999:BRMh

[Ano99c]  Anonymous.  Book re-

view: The Matlab(R) 5 handbook: By Darren Redfern and Colin Campbell. Springer-Verlag, New York. (1998). 488 pages. $34.95, DM 69.00, öS 504.00, sFr 63.00, GBP 19.50. Computers and Mathematics with Applications, 37(6):

134, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199902651.

Anonymous:1999:VCW

[Ano99d] Anonymous. Visualization of cannon wear using ultrasonic measurements and MATLAB(R). Technical report AD-a361 642, Army Armament Research Development and Engineering Center, Watervliet, NY, Benet Weapons Laboratory, 1999. 18 pp.

Anonymous:2000:BRDa

[Ano00] Anonymous. Book review:

Differential equations with MATLAB^{T M}: By Kevin R. Coombes, Brian R. Hunt, Ronald L. Lipsman, John E. Osborn and Garrett J. Stuck. John Wiley & Sons, New York. (2000). 256 pages. $19.99. Computers and Mathematics with Applications, 39(3–4):262, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900336.

Anonymous:2002:CVG

[Ano02a] Anonymous. Code vectorization guide. Technical note 1109 (Revision 2.0), The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, October

                                        15, 2002. URL http://www.
                                        mathworks.com/support/
                                        tech-notes/1100/1109.shtml.

Anonymous:2002:MR

[Ano02b] Anonymous. MATLAB resources. Technical report, Department of Mathematics, University of Haifa, Haifa, Israel, 2002. URL

                                        http://math.haifa.ac.il/
                                        matlab.html. This is an

extensive list of pointers to MATLAB material on the World-Wide Web.

Anonymous:2005:PDS

[Ano05] Anonymous. Pricing Derivative Securities with Matlab 2e: An Interactive, Dynamic Environment with Maple V and Matlab. Academic Press, New York, NY, USA, 2005. ISBN 0-12-565651-3. ???? pp. LCCN ????

Anonymous:2009:BRCa

[Ano09]  Anonymous.  Book re-

view: Computational Statistics Handbook with MAT-LAB(R), 2nd edition by Martinez, W. L. and Martinez, A. R. Biometrics, 65(1):338, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2012:BRBh

[Ano12] Anonymous. Book review: Exploratory Data Analysis With MATLAB (Second Edition) by Wendy L. Martinez; Angel R. Martinez; Jeffrey L. Solka. Technometrics, 54(3):330, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714912.

Anonymous:2019:MCT

[Ano19] Anonymous. Multiprecision computing toolbox for MATLAB. Web site, September 26, 2019. URL

https://www.advanpix.com/

.

Ansari:2014:PAE

[Ans14] Nadia Mustaqim Ansari.

Performance analysis of ETTH and FTTH in terms of cost and scalability: a comparative study carried out on MATLAB. International Journal of Computer Applications, 98(??): 1–5, July 2014. CODEN ???? ISSN 0975-8887. URL

https://www.ijcaonline.
org/archives/volume98/
number11/17224-7546/.

Ashino:2000:BBM

[ANV00] R. Ashino, M. Nagase, and R. Vaillancourt. Behind and beyond the MATLAB ODE suite. Computers and Mathematics with Applications, 40(4-5):491–512, 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Antoniadis:1995:WSP

[AO95] Anestis Antoniadis and Georges Oppenheim, editors. Wavelets and statistics: Papers from the 15th Franco-Belgian Meeting of Statisticians, held in Villard de Lans, France, Nov. 16–18, 1994, volume 103 of Lecture Notes in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94564-4. ISSN 0930-0325. LCCN QA403.3 .W3857 1995.

Almasi:2001:MJM

[AP01] George Almasi and David A. Padua. MaJIC: a Matlab just-in-time compiler. Lecture Notes in Computer Science, 2017:68–81, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2017/20170068.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2017/20170068.
                                        pdf.

Almasi:2002:MCM

[AP02] George Almási and David Padua. MaJIC: compiling MATLAB for speed and responsiveness. ACM SIG-PLAN Notices, 37(5):294– 303, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alonso:2019:CMT

[API^{+}19] Pedro Alonso, Jesús Peinado, Javier Ibáñez, Jorge Sastre, and Emilio Defez. Computing matrix trigonometric functions with GPUs through Matlab. The Journal of Supercomputing, 75 (3):1227–1240, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Appleby:2019:BRB

[App19] Rob Appleby. Book review:

Beams and accelerators with MATLAB, (with companion media pack). Contemporary Physics, 60(2):209, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Allman:2003:MMB

[AR03] Elizabeth Spencer Allman and John A. (John Anthony) Rhodes. Mathematical models in biology: an introduction. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-81980-6, 0-521-52586-1 (paperback). xiii + 370 pp.

Allman:2004:MMB

[AR04] Elizabeth Spencer Allman and John A. (John Anthony) Rhodes. Mathematical models in biology: an introduction. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-81980-6 (hb.), 0-521-52586-1 (paperback). xiii + 370 pp. LCCN QH323.5 .A44 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam032/2003043929.
html;   http://www.loc.
gov/catdir/toc/cam031/
2003043929.html.

Arendt:1994:MMP

[Are94] R. Arendt. The modelling of marine propulsion unit control system property with an application of program Simulink. In Anonymous [Ano94g], pages 7–18. ISBN ???? ISSN 0860-0325. LCCN ????

Auslander:2002:CSM

[ARR02] D. M. (David Martin) Auslander, J. R. Ridgely, and J. D. Ringgenberg. Control software for mechanical systems: object-oriented design in a real-time world. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-786302-0. xvi + 339 pp.

Alefeld:2001:SAM

[ARRY01] Götz Alefeld, Jiři Rohn, Siegfried Rump, and Tetsuro Yamamoto, editors.

Symbolic algebraic methods and verification methods. Springer mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-211-83593-8. ix + 266 pp. LCCN QA76.9.M35 S92 2001. US$69.95. URL http:

//www.springer-ny.com/
detail.tpl?cart=10209516271260963&
isbn=3211835938.

Azemi:1996:UMU

[AS96] A. Azemi and C. Stook.

Utilizing MATLAB in undergraduate electric circuits courses. In Iskander et al. [I^{+}96], pages 599–

602. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Atkinson:2008:ASF

[AS08] Kendall E. Atkinson and Lawrence F. Shampine. Algorithm 876: Solving Fredholm integral equations of the second kind in Matlab. ACM Transactions on Mathematical Software, 34 (4):21:1–21:20, July 2008. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Amodio:2012:FDM

[AS12] Pierluigi Amodio and Giuseppina Settanni. A finite differences MATLAB code for the numerical solution of second order singular perturbation problems. Journal of Computational and Applied Mathematics, 236(16):3869–3879, October 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042712001719.

Agrawal:2015:MSS

[AS15] Pallavi Agrawal and Madhu Shandilya. MATLAB simulation of subspace based high resolution direction of arrival estimation algorithm. International Journal of Computer Applications, 130(??):22–27, November 2015. CODEN ???? ISSN 0975-8887. URL

https://www.ijcaonline.
org/archives/volume130/
number15/23286-2015907175/

.

Ata:1996:DIF

[ASA96] Atef A. Ata, Ali R. Shahin, and Shihab S. Asfour. Design of an industrial flexible robot controller using MATLAB. In Randhawa [Ran96], pages 131–

134. ISBN ???? ISSN 0360-8352 (print), 1879-0550 (electronic). LCCN ????

Atherton:1994:TCE

[ASG94] D. P. Atherton, O. B.

Sorensen, and A. Goucem. Teaching control engineering using implementations of MATLAB. In Ichikawa and Furuta [IF94], pages 273–276. ISBN 0-08-042230-

6. LCCN TJ212.2 .A393 1995.

Asinari:2010:NBE

[Asi10]  Pietro Asinari.  Nonlin-

ear Boltzmann equation for the homogeneous isotropic case: Minimal deterministic Matlab program. Computer Physics Communications, 181(10):1776–1788, October 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465510002420.

Athreyas:2018:MCA

[ASP^{+}18] Nihar Athreyas, Wenhao Song, Blair Perot, Qiangfei Xia, Abbie Mathew, Jai Gupta, Dev Gupta, and J. Joshua Yang. Memristor-CMOS analog coprocessor for acceleration of highperformance computing applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):38:1–38:??, October 2018. CODEN ???? ISSN 1550-4832.

Asundi:2002:MPP

[Asu02]  Anand Asundi.  MAT-

LAB for photomechanics: a primer. Elsevier, Amsterdam, The Netherlands, 2002. ISBN 0-08-044050-9. viii + 189 pp. LCCN TA418.12 .A88 2002.

Aurentz:2017:CCS

[AT17] Jared L. Aurentz and Lloyd N. Trefethen. Chopping a Chebyshev series. ACM Transactions on Mathematical Software, 43 (4):33:1–33:21, March 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Aster:2005:PEI

[ATB05] Richard C. Aster, Clifford H. Thurber, and Brian Borchers. Parameter estimation and inverse problems, volume 90 of International geophysics series. Elsevier Academic Press, Amsterdam, The Netherlands, 2005. ISBN 0-12-065604-3. xii + 301 pp. LCCN QA276.8 .A88 2005.

Atkinson:2005:ENA

[Atk05]  Kendall Atkinson.  Ele-

mentary Numerical Analysis (3rd Edition) with MAT-LAB: An Introduction (2nd Edition) Set. Wiley, New York, NY, USA, 2005. ISBN 0-471-74257-0. ???? pp. LCCN ????

Attia:1995:TACb

[Att95a] J. O. Attia. Teaching AC circuit analysis with MAT-LAB. In Bergadaa [Ber95], pages 2C6.9–2C6.12. CO-DEN PFECDR. ISBN 2-911209-00-1. ISSN 0190-5848. LCCN ????

Attia:1995:TACa

[Att95b] John Okeyere Attia. Teaching AC circuit analysis with MATLAB. Frontiers in Education Conference, 1(????): 333–336, ???? 1995. CO-DEN PFECDR. ISSN 0190-5848. IEEE catalog number 95CB35867.

Attia:1996:TEM

[Att96] J. Okyere Attia. Teaching electronics with MATLAB. In Iskander et al. [I^{+}96], pages 609–611. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Attia:1999:ECA

[Att99] John Okyere Attia. Electronics and circuit analysis using MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-1176-4. 378 pp. LCCN TK7835 .A88 1999.

Attia:2002:PME

[Att02] John Okyere Attia. PSPICE and MATLAB for electronics: an integrated approach. VLSI circuits series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2002. ISBN 0-8493-1263-9. 338 (est.) pp. LCCN TK7874.75 .A88 2002.

Attia:2004:ECA

[Att04] John Okyere Attia. Electronics and circuit analysis using MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2004. ISBN 0-8493-1892-0. 393 (est.) pp. LCCN TK7835 .A88 2004.

Attaway:2009:MPI

[Att09] Stormy Attaway. MATLAB: a practical introduction to programming and problem solving. Butterworth-Heinemann, Amsterdam, The Netherlands, 2009. ISBN 0-7506-8762-2 (paperback). xx + 452 pp. LCCN QA297 .A87 2009.

Attia:2010:PME

[Att10] John Okyere Attia. PSPICE and MATLAB for electronics: an integrated approach. VLSI circuits series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2010. ISBN 1-4200-8658-8. ???? pp. LCCN TK7874.75 .A88 2010.

Attaway:2012:MPI

[Att12] Stormy Attaway. MATLAB: a practical introduction to programming and problem solving. Butterworth-Heinemann, Waltham, MA, USA, second edition, 2012. ISBN 0-12-385081-9. xx + 518 pp. LCCN QA297 .A87 2012.

Abebe:2017:JCP

[ATvBB17] Haftom Temesgen Abebe, Frans E. S. Tan, Gerard J. P. van Breukelen, and Martijn P. F. Berger. JMASM 45: a computer program for Bayesian D-optimal binary repeated measurements designs (Matlab). Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/
jmasm/article/view/920.

Anjam:2015:FMA

[AV15] I. Anjam and J. Valdman. Fast MATLAB assembly of FEM matrices in 2D and 3D: Edge elements. Applied Mathematics and Computation, 267(??):252–263, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300315004191.

Anderson:1997:CII

[AVV97] Glen D. (Glen Douglas) Anderson, M. K. (Mavina Krishna) Vamanamurthy, and Matti Vuorinen. Conformal invariants, inequalities, and quasiconformal maps. Canadian Mathematical Society series of monographs and advanced texts. Wiley, New York, NY, USA, 1997. ISBN 0-471-59486-5 (cloth). xxvii + 505 pp. LCCN QA360 .A64 1997. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley033/97013356.
                                        html;   http://www.loc.
                                        gov/catdir/toc/wiley022/
                                        97013356.html.

Azvine:1992:IQM

[AW92] B. Azvine and R. J. Wynne. Implementing the QFT method in the MATLAB environment. In Barker [Bar92], pages 335–340. ISBN 0-08-041269-6. LCCN TJ213 .C57 1992.

Aastrom:1997:CCS

[ÅW97a] Karl J. (Karl Johan) Åström and Björn Wittenmark. Computer-controlled systems: theory and design. Prentice Hall information and system sciences series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1997. ISBN 0-13-314899-8. xiv + 557 pp. LCCN TJ213 .A78 1997.

                                        URL ftp://ftp.control.
                                        llth.se/pub/books/ccs.

Azvine:1997:IMQ

[AW97b] B. Azvine and R. J. Wynne.

Improved MIMO quantitative feedback design in Matlab. In Gertler et al. [GCP97], pages 321–326. ISBN 0-08-042926-2 (plenary vol.), 0-08-042909-2 (vol. A), 0-08-042910-6 (vol. B), 0-08-042911-4 (vol. C), 0-08-042912-2 (vol. D), 0-08-042913-0 (vol. E), 0-08-042914-9 (vol. F), 0-08-042915-7 (vol. G), 0-08-042916-5 (vol. H), 0-08-042917-3 (vol. I), 0-08-042918-1 (vol. J), 0-08-042919-X (vol. K), 0-08-042920-3 (vol. L), 0-08-042921-1 (vol. M), 0-08-042922-X (vol. N), 0-08-042923-8 (vol. O), 0-08-042924-6 (vol. P), 0-08-042925-4 (vol. Q). LCCN ????

Atherton:1991:DPN

[AY91] D. P. Atherton and T. C.

Yang. Development of programs for nonlinear control systems in MATLAB. IEE Conference Publication, 2 (332):1240–1245, 1991. CO-DEN IECPB4. ISSN 0537-9987 (??invalid ISSN checksum??).

Azemi:1994:PMU

[AY94] Asad Azemi and Edwin Engin Yaz. PSpice and MAT-LAB in undergraduate and graduate electrical engineering courses. Frontiers in Education Conference, ???? (????):456–459, ???? 1994.

CODEN PFECDR. ISSN 0190-5848.

Azemi:1996:USM

[AY96] A. Azemi and E. Engin Yaz. Utilizing SIMULINK and MATLAB in a graduate nonlinear systems analysis course. In Iskander et al. [I^{+}96], pages 595– 598. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Azemi:1997:UMG

[AY97] A. Azemi and E. Engin Yaz. Using MATLAB in a graduate electrical engineering optimal control course. In EP Innovations [EP 97], pages 13–17. ISBN 0-7803-4087-6, 0-7803-4086-8, 0-7803-4088-4, 0-7803-4089-2. ISSN 0190-5848. LCCN T62 .F76 1997. Three volumes.

Burrus:1994:CBE

[B^{+}94] C. Sidney Burrus et al. Computer-Based Exercises for Signal Processing Using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-219825-8, 0-13-364845-1. xii + 404 pp. LCCN TK5102.9 .C57 1994.

BaezLopez:2010:MAE

[B10] David Báez López. MAT-LAB with applications to engineering, physics and finance. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-0697-7 (hardcover). xiv + 412 pp. LCCN QA297 .B28 2010.

Bohn:1994:SPC

[BA94] C. Bohn and D. P. Atherton. A SIMULINK package for comparative studies of PID anti-windup strategies. In Mattsson et al. [MGC94], pages 447–452. ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Babuska:1994:DFC

[Bab94]  R. Babuska.  Designing

fuzzy control systems with MATLAB and SIMULINK. In Anonymous [Ano94h], pages 269–273. ISBN 3-86073-286-2. LCCN ???? Three volumes.

Babuska:1998:FMC

[Bab98] Robert Babuška. Fuzzy modeling for control, volume ISIT 12 of International series in intelligent technologies. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1998. ISBN 0-7923-8154-8. xii + 260 pp. LCCN TJ217.5 .B33 1998.

Backstrom:1995:PMU

[Bäc95] Gunnar Bäckström. Practical Mathematics Using MATLAB. Studentlitteratur and Chartwell-Bratt, Lund, Sweden and Bromley, Kent, UK, 1995. ISBN 91-44-49231-6 (Studentlitteratur), 0-86238-397-8 (Chartwell Bratt). 165 pp. LCCN ????

Backstrom:1997:PMM

[Bäc97] Gunnar Bäckström. Praktisk matematik med MAT-LAB 5 (Swedish) [Practical Mathematics Using MAT-LAB 5]. Studentlitteratur, Lund, Sweden, 1997. ISBN 91-44-00544-X (Studentlitteratur), 0-86238-491-5 (Chartwell Bratt), 91-44-00244-6 (Studentlitteratur Swedish version). ???? pp. LCCN QA76.95 .B3 1997.

Backstrom:2000:PMU

[Bäc00] Gunnar Bäckström. Practical mathematics using MATLAB. Studentlitteratur, Lund, Sweden, second edition, 2000. ISBN 91-44-01552-6. 235 pp. LCCN QA76.95 .B3 2000.

Backstrom:2004:AMU

[Bäc04] Gunnar Bäckström. Alternative mathematics using MATLAB 7. GB Publishing, Malmö, Sweden, 2004. ISBN ???? 239 pp. LCCN ????

Boashash:2019:MTF

[BAEBAS19] Boualem Boashash, Abdeldjalil Aïssa-El-Bey, and Mohammad F. Al-Sa’d. Multisensor time-frequency signal processing MAT-LAB package: an analysis tool for multichannel non-stationary data. SoftwareX, 8(??):53–58, ???? 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711017300687.

Bai:2005:WSP

[Bai05] Ying Bai. The Windows serial port programming handbook. Auerbach Publications, Boca Raton, FL, USA, 2005. ISBN 0-8493-2213-8. 801 (est.) pp. LCCN TK7887.5 .B35 2005.

Balac:2019:WMT

[Bal19] Stéphane Balac. WGMode:

a Matlab toolbox for whispering gallery modes volume computation in spherical optical micro-resonators. Computer Physics Communications, 243(??):121–134, October 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465519301511.

Banks:2001:DEG

[Ban01] Bernard W. Banks. Differential equations with graphical and numerical methods. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-084376-8. xiii + 462 pp. LCCN QA371 .B26 2001.

Barker:1992:SJS

[Bar92] H. A. Barker, editor. Computer aided design in control systems: selected papers from the IFAC symposium, Swansea, UK, 15–17 July 1991, number 1 in IFAC Symposia Series. Pergamon Press, New York, NY, USA, 1992. ISBN 0-08-041269-6. LCCN TJ213 .C57 1992.

Baran:1997:ETM

[Bar97] S. Baran. Estimating the transition matrix of a finite state space Markov chain with MATLAB. In Zempleni [Zem97], pages 29–34. ISBN 963-463-082-0. LCCN ????

Barber:2016:BRE

[Bar16a] D. C. Barber. Book review: Essential MATLAB and Octave, by Jesus Rogel-Salazar. Contemporary Physics, 57(3):455–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barber:2016:BRT

[Bar16b] D. C. Barber. Book review: Theoretical foundations of digital imaging using Matlab, by Leonid P. Yaroslavsky. Contemporary Physics, 57(4):608–609, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bateman:1999:DCD

[Bat99] Andrew Bateman. Digital communications: design for the real world. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-34301-

0. xxv + 221 pp. LCCN TK5103.7 .B377 1999.

Batista:2019:ECM

[Bat19] Milan Batista. Elfun18 — a collection of MATLAB functions for the computation of elliptic integrals and Jacobian elliptic functions of real arguments. SoftwareX, 10(??):Article 100245, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711018302796.

Battig:2020:AMM

[Bät20]  Daniel Bättig.  Ange-

wandte Mathematik 1 mit MATLAB und Julia: Ein anwendungs- und beispielorientierter Einstieg für technische Studiengänge. (German) [Applied Mathematics 1 with MATLAB and Julia: an application and example-oriented introduction to technical courses]. Springer Vieweg, Berlin and Heidelberg, Germany, 2020. ISBN 3-662-60951-7 (print), 3-662-60952-5 (ePub). xiii + 254 pp. LCCN ¿¿¿¿.

Baura:2002:STP

[Bau02] Gail D. Baura. System theory and practical applications of biomedical signals. IEEE Press series in biomedical engineering. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-471-23653-5. xxvii + 440 pp. LCCN R857.S47 B38 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley044/2002512546.
html;   http://www.loc.
gov/catdir/description/
wiley037/2002512546.html;
 http://www.loc.gov/catdir/
toc/wiley031/2002512546.
html.

Ben-Asher:1998:AMG

[BAY98] Joseph Z. Ben-Asher and Isaac Yaesh. Advances in missile guidance theory, volume 180 of Progress in astronautics and aeronautics. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024–2518, 1998. ISBN 1-56347-275-9. x + 200 pp. LCCN TL507 .P75.

Bay:1999:FLS

[Bay99] John S. Bay. Fundamentals of linear state space systems. WCB/McGraw-Hill, Boston, MA, USA, 1999. ISBN 0-256-24639-4. xviii + 571 pp. LCCN QA402 .B39 1999. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/mh023/98025177.
html;   http://www.loc.
gov/catdir/toc/mh022/98025177.
html.

Biran:1995:ME

[BB95a] Adrian Biran and Moshe M. G. Breiner. MATLAB for Engineers. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-56524-

2. xiv + 668 pp. LCCN TA345.B485 1995.

Biran:1995:MI

[BB95b] Adrian B. Biran and Moshe M. G. Breiner. MAT-LAB für Ingenieure (German) [English: MATLAB for Engineers]. Addison-Wesley, Reading, MA, USA, 1995. ISBN 3-89319-856-3. ???? pp. LCCN ????

Beardah:1996:MRK

[BB96] C. C. Beardah and M. J.

Baxter. MATLAB routines for kernel density estimation and the graphical representation of archaeological data. In Kamermans and Fennema [KF96], pages 179–184. ISBN 90-73368-10-3. ISSN 0169-7447. LCCN GN700 .A62 v.28 v.1-2 (1996).

Burger:1997:PRP

[BB97] Wilhelm Burger and Mark Burge, editors. Pattern recognition 1997: proceedings of the 21st Workshop of the Austrian Association for Pattern Recognition (OAGM/AAPR), Hallstatt, Upper Austria, May 26–27, 1997, number 103 in Schriftenreihe — Österreichischen Computer Gesellschaft 1997. Oldenbourg-Verlag, München, Wien, Oldenbourg, 1997. ISBN 3-486-24494-9 (München), 3-7029-0436-0 (Wien), 3-85403-103-3 (Österreichischen Computer Gesellschaft). LCCN TA1650.O88 1997.

Biran:1999:ME

[BB99] Adrian Biran and Moshe Breiner. MATLAB 5 for Engineers. Addison-Wes-ley, Reading, MA, USA, second edition, 1999. ISBN 0-201-36043-8. xvi + 688 pp. LCCN QA297.B52 1999.

URL http://www.awl-he.
com/titles/11606.html.

Rev. ed. of: MATLAB for engineers (1995).

Biran:2002:ME

[BB02] Adrian Biran and Moshe Breiner. MATLAB 6 for engineers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-033631-9 (paperback). xx + 777 pp. LCCN QA297 .B522 2002.

Barnat:2012:TCS

[BBB12] Jiri Barnat, Jan Beran, and Lubos Brim. Tool chain to support automated formal verification of avionics Simulink designs. Lecture Notes in Computer Science, 7437:78–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-32469-7_
6/.

Barrett:1994:TSLb

[BBC^{+}94a] Richard Barrett, Michael Berry, Tony F. Chan, James W. Demmel, June Donato, Jack Dongarra, Victor Eijkhout, Roldan Pozo, Charles Romine, and Henk van der Vorst. Templates for the Solution of Linear Systems: Building Blocks for Iterative Methods (Japanese). Asakura Shoten, Tokyo, Japan, 1994. ISBN 4-254-11401-X. ???? pp. LCCN ????

Barrett:1994:TSLa

[BBC^{+}94b] Richard Barrett, Mike Berry, Tony Chan, Jim Demmel, June Donato, Jack Dongarra, Victor Eijkhout, Roldan Pozo, Chuck Romine, and Henk van der Vorst. Templates for the Solution of Linear Systems: Building Blocks for Iterative Methods. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-328-5. xiii + 112 pp. LCCN QA297.8 .T45 1994.

Brust:2022:ASS

[BBEM22] Johannes Brust, Oleg Burdakov, Jennifer Erway, and Roummel Marcia. Algorithm 1030: SC-SR1: MAT-LAB software for limitedmemory SR1 trust-region methods. ACM Transactions on Mathematical Software, 48(4):48:1–48:??, December 2022. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3550269.

Bortolotti:2019:UUP

[BBF^{+}19] V. Bortolotti, L. Brizi, P. Fantazzini, G. Landi, and F. Zama. Upen2DTool: a Uniform PENalty Matlab tool for inversion of 2D NMR relaxation data. SoftwareX, 10(??):Article 100302, July/December 2019. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711019302006.

Bischof:2008:AAD

[BBH^{+}08] Christian H. Bischof, H. Martin Bücker, Paul Hovland, Uwe Naumann, and Jean Utke, editors. Advances in Automatic Differentiation, volume 64 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNCSA6. ISBN 3-540-68935-4 (print), 3-540-68942-7 (e-book). ISSN 1439-7358. LCCN QA304 .I58 2008. URL http://

link.springer.com/book/
10.1007/978-3-540-68942-
3;   http://www.springerlink.
com/content/978-3-540-
68942-3.

Bidegain:2023:MGB

[BBHP^{+}23] Gorka Bidegain, Tal Ben-Horin, Eric N. Powell, John M. Klinck, and Eileen E. Hofmann. MarineEpi: a GUI-based Matlab toolbox to simulate marine pathogen transmission. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000535.

Bortolotti:2022:MNM

[BBN^{+}22] Villiam Bortolotti, Leonardo Brizi, Anastasiia Nagmutdinova, Fabiana Zama, and Germana Landi. MUPen2DTool: a new Matlab tool for 2D nuclear magnetic resonance relaxation data inversion. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001583.

Borrelli:2004:DEM

[BC04] Robert L. Borrelli and Courtney S. Coleman. Differential equations: a modeling perspective. Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-43332-2. xii + 718 pp. LCCN QA371 .B74 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley041/2004533274.
                                        html;   http://www.loc.
                                        gov/catdir/toc/wiley041/
                                        2004533274.html.

Bahriawati:2005:TMI

[BC05] C. Bahriawati and C. Carstensen. Three MATLAB implementations of the lowest-order Raviart–Thomas MFEM with a posteriori error control. Computational Methods in Applied Mathematics, 5(4):333–361 (electronic), 2005. ISSN 1609-4840.

Blanchet:2006:DSI

[BC06] Gerard Blanchet and Maurice Charbit. Digital signal and image processing using Matlab. Digital signal and image processing series. ISTE Ltd., London, UK and Newport Beach, CA, USA, 2006. ISBN 1-905209-13-4. 763 pp. LCCN TK5102.9 .B545 2006. Translation of: Signaux et images sous Matlab.

Bouissou:2012:OSS

[BC12] Olivier Bouissou and Alexandre Chapoutot. An operational semantics for Simulink’s simulation engine. ACM SIGPLAN Notices, 47(5):129–138, May 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCTES ’12 proceedings.

Baddour:2017:MCD

[BC17a] Natalie Baddour and Ugo Chouinard. Matlab code for the Discrete Hankel Transform. Journal of Open Research Software, 5(1): 4–??, January 11, 2017. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.82/.

Bispo:2017:MSC

[BC17b] João Bispo and João M. P.

Cardoso. A MATLAB subset to C compiler targeting embedded systems. Software—Practice and Experience, 47(2):249–272, February 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Black:2022:MBR

[BC22] Travis J. Black and Alexei F.

Cheviakov. 3DRSP: Matlabbased random sphere packing code in three dimensions. SoftwareX, 18(??): ??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711022000437.

Bourke:2017:SLS

[BCC^{+}17] Timothy Bourke, François Carcenac, Jean-Louis Colaço, Bruno Pagano, Cédric Pasteur, and Marc Pouzet. A synchronous look at the Simulink standard library. ACM Transactions on Embedded Computing Systems, 16(5s):176:1–176:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bigatti:2020:MSI

[BCD^{+}20] Anna Maria Bigatti, Jacques Carette, James H. Davenport, Michael Joswig, and Timo de Wolff, editors. Mathematical Software — ICMS 2020: 7th International Conference, Braunschweig, Germany, July 13–16, 2020, Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020.

Bird:2017:FNM

[BCG17] R. E. Bird, W. M. Coombs, and S. Giani. Fast native-MATLAB stiffness assembly for SIPG linear elasticity. Computers and Mathematics with Applications, 74(12):3209–3230, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812211730514X.

Bartels:2006:PIF

[BCH06] S. Bartels, C. Carstensen, and A. Hecht. P2Q2Iso2D = 2D isoparametric FEM in Matlab. Journal of Computational and Applied Mathematics, 192(2):219–250, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042705002839.

Banerjee:1998:MMC

[BCHS98] Prithviraj Banerjee, Alok Choudhary, Scott Hauck, and Nagaraj Shenoy. MATCH: a MATLAB compilation environment for distributed heterogeneous adaptive computing systems. Technical report, Northwestern University, Evanston, IL, USA,

1998. URL http://www.
accelchip.com/;   http:
//www.ece.northwestern.
edu/cpdc/Match/Match.html.

From the Web page: “An exclusive license of the Match Compiler has been transferred to a startup company called AccelChip, Inc. (formerly called Mach Design Systems), located in Schaumburg, Illinois.”.

Bhattacharyya:1996:RCP

[BCK96] S. P. Bhattacharyya, H. Chapellat, and L. H. Keel. Robust Control: The Parametric Approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-781576-X. xiv + 648 pp. LCCN TJ213.B534 1995.

Baglama:2003:AIM

[BCR03] J. Baglama, D. Calvetti, and L. Reichel. Algorithm 827: irbleigs: A MAT-LAB program for computing a few eigenpairs of a large sparse Hermitian matrix. ACM Transactions on Mathematical Software, 29(3):337–348, September 2003. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Boyce:1995:GDE

[BD95] William E. Boyce and Richard C. DiPrima. Gewöhnliche Differentialgleichungen: Einführung, Aufgaben, Lösungen. (German) [Ordinary Differential Equations: Introduction, Problems, and Solutions]. Spektrum, Akad. Verl., Heidelberg, Germany; Berlin, Germany; Oxford, UK, 1995. ISBN 3-86025-151-1. xv + 699 pp. LCCN ???? Translated from English by Horst Welter.

Beale:1996:FST

[BD96] Mark H. Beale and Howard B. Demuth. Fuzzy systems toolbox for use with MAT-LAB. PWS productivity tools. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, student, Mac version edition, 1996. ISBN 0-534-95258-5 (disk), 0-534-94579-1 (manual). 1 computer disk pp. LCCN QA297.

Boyce:2000:GDE

[BD00] William E. Boyce and Richard C. DiPrima. Gewöhnliche Differentialgleichungen: Einführung, Aufgaben, Lösungen. (German) [Ordinary Differential Equations: Introduction, Problems, and Solutions]. Spektrum, Akad. Verl., Heidelberg, Germany; Berlin, Germany; Oxford, UK, 2000. ISBN 3-8274-0597-1. xv + 699 pp. LCCN ???? Translated from English by Horst Welter.

Boyce:2004:EDE

[BD04] William E. Boyce and Richard C. DiPrima. Elementary differential equations. Wiley, New York, NY, USA, eighth edition, 2004. ISBN 0-471-43339-X. xviii + 622 pp. LCCN QA371 .B77 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley047/2004043373.
html;   http://www.loc.
gov/catdir/description/

                                        wiley041/2004043373.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley041/2004043373.
                                        html.

Boyce:2005:EDE

[BD05] William E. Boyce and Richard C. DiPrima. Elementary differential equations and boundary value problems. Wiley, New York, NY, USA, eighth edition, 2005. ISBN 0-471-43338-1. xviii + 790 pp. LCCN QA371 .B773 2005. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley047/2003070545.
                                        html;   http://www.loc.
                                        gov/catdir/description/
                                        wiley041/2003070545.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley041/2003070545.
                                        html.

Bal:2020:JMA

[BD20] Cagatay Bal and Serdar Demir. JMASM 55: MAT-LAB algorithms and source codes of ’cbnet’ function for univariate time series modeling with neural networks (MATLAB). Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https:

                                        //jmasm.com/index.php/
                                        jmasm/article/view/1100.

Byun:2022:RMT

[BD22] Ji-Eun Byun and Dina D’Ayala. ResMapper: Matlab tool for seismic resilience mapping of largescale road networks. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001674.

Bulai:2025:NMS

[BDL25] Iulia Martina Bulai, Maria Carmela De Bonis, and Concetta Laurita. A new MATLAB software for numerical computation of biological observables for metastatic tumor growth. Mathematics and Computers in Simulation, 234(??):31–49, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0378475425000503.

Bernardeschi:2017:PSI

[BDM17] C. Bernardeschi, A. Domenici, and P. Masci. A PVS– Simulink integrated environment for model-based analysis of cyber-physical systems. IEEE Transactions on Software Engineering, PP(99):1, ???? 2017. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=
7900400.

Bernardeschi:2018:PSI

[BDM18] Cinzia Bernardeschi, Andrea Domenici, and Paolo Masci. A PVS-Simulink integrated environment for model-based analysis of cyber-physical systems. IEEE Transactions on Software Engineering, 44(6):512–533, ???? 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL https:

                                        //ieeexplore.ieee.org/
                                        document/7900400/.

Barboza:2004:TIA

[BDR04] L. V. Barboza, G. P. Dimuro, and R. H. S. Reiser. Towards interval analysis of the load uncertainty in power electric systems. In IEEE [IEE04], pages 538– 541. ISBN 0-9761319-1-9 (paperback). LCCN TK1005 .I547 2004.

Buck:1997:CES

[BDS97] John R. Buck, Michael M. Daniel, and Andrew Singer. Computer explorations in signals and systems using MATLAB. Prentice Hall signal processing series MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-732868-0. 207 pp. LCCN TK5102.9.B83 1997.

Buck:2002:CES

[BDS02] John R. Buck, Michael M.

Daniel, and Andrew Singer. Computer explorations in signals and systems using MATLAB. Prentice-Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2002. ISBN 0-13-042155-3. xii + 207 pp. LCCN TK5102.9 .B83 2002.

Beebe:1994:BPMb

[Bee94] Nelson H. F. Beebe. A bibliography of publications about the Matlab on-line matrix laboratory. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 7, 1994. 18

pp. URL https://www.
math.utah.edu/pub/tex/
bib/index-table-m.html#

matlab. This report is updated frequently.

Beebe:2005:MON

[Bee05] Nelson H. F. Beebe. Matlab and Octave (numerical linear algebra) FAQ. World-Wide Web frequently-asked question document., September 2005. URL https:/

/www.math.utah.edu/faq/

matlab/. This report is updated frequently.

Beebe:2017:MFC

[Bee17] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (e-book). xxxvi + 1114 pp. LCCN QA75.5-76.95. URL http:

                                        //www.springer.com/us/
                                        book/9783319641096.

Baird:1994:CPC

[BEH94] C. R. Baird and M. E. El-Hawary, editors. Conference proceedings: 1994 Canadian Conference on Electrical and Computer Engineering, September 25– 28, 1994, World Trade and Convention Center, Halifax, Canada, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-2416-1, 0-7803-2417-X. LCCN TK 7801 C36 1994.

Butler:1999:MMS

[BEK99] Karen L. Butler, Mehrdad Ehsani, and Preyas Kamath. A Matlab-based modeling and simulation package for electric and hybrid electric vehicle design. IEEE transactions on vehicular technology, 48(6):1770– 1778, November 1999.

Bennett:1995:SF

[Ben95a] B. S. Bennett. Simulation Fundamentals. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-813262-3. xvii + 329 pp. LCCN QA76.9.C65 .B46 1995.

Benson:1995:IDS

[Ben95b] D. Benson. An integrated dynamic system identification tool for MATLAB[R]. In Anonymous [Ano95s], pages 1305–1309. ISBN ???? LCCN ????

Benaroya:1998:MVA

[Ben98a] Haym Benaroya. Mechanical vibration: analysis, uncertainties, and control. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-948373-X. xvi + 511 pp. LCCN QA935 .B36 1998.

Benker:1998:ICS

[Ben98b] Hans Benker. Ingenieurmathematik mit Computeralgebra-Systemen. AXIOM, DE-RIVE, MACSYMA, MAPLE, MATHCAD, MATHEMAT-ICA, MATLAB und Mu-PAD in der Anwendung. (German) Engineering mathematics with computer algebra systems. The applications: AXIOM, DERIVE, MACSYMA, MAPLE, MATH-CAD, MATHEMATICA, MATLAB UND MuPAD. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1998. xiii + 439 pp.

Benitez:2002:PMA

[Ben02]  Jaime Benítez.  Princi-

ples and modern applications of mass transfer operations. Wiley-Interscience, New York, NY, USA, 2002. ISBN 0-471-20344-0. xxi + 499 pp. LCCN TP156.M3 B44 2002. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley0310/
                                        2002284580.html;   http:
                                        //www.loc.gov/catdir/toc/
                                        wiley023/2002284580.html.

Benaroya:2004:MVA

[Ben04] Haym Benaroya. Mechanical vibration: analysis, uncertainties, and control, volume 181 of Mechanical engineering. Marcel Dekker, New York, NY, USA, second edition, 2004. ISBN 0-8247-5380-1 (hardcover). xvii + 712 pp. LCCN QA935 .B36 2004.

Benker:2005:DMM

[Ben05] H. Benker. Differentialgleichungen mit Mathcad und Matlab. (German) [Differential Equations with Mathcad and Matlab]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-23440-3. 270 (est.) pp. LCCN ???? EUR 39.95; SFR 68.00.

Bequette:1998:PDM

[Beq98] B. Wayne Bequette. Process dynamics: modeling, analysis, and simulation. Prentice Hall international series in the physical and chemical engineering sciences. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-206889-3. xviii + 621 pp. LCCN TP155.7 .B45 1998.

Bequette:2003:PCM

[Beq03] Wayne B. Bequette. Process control: modeling, design, and simulation. Prentice-Hall international series in the physical and chemical engineering sciences. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-353640-8 (paper). xxix + 769 pp. LCCN TS156.8 .B46 2003.

Bergadaa:1995:MTC

[Ber95] M. Bergadaa, editor. Marketing today and for the 21st Century: 24th Annual conference — May 1995, Cergy, France, number 24 in Proceedings of the Annual Conference — European Marketing Academy //V1. 1995, Ecole Superieure des Sciences Economiques et Commerciales, 1995. CODEN PFECDR. ISBN 2-911209-00-1. ISSN 0190-5848. LCCN ????

Bernstein:2003:PEI

[Ber03] Ruth Bernstein. Population ecology: an introduction to computer simulations. Wiley, New York, NY, USA, 2003. ISBN 0-470-85148-1 (paper). viii + 158 pp. LCCN QH352 .B458 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley039/2002155483.
html;   http://www.loc.
gov/catdir/toc/wiley031/
2002155483.html.

Berens:2009:CMT

[Ber09] Philipp Berens. CircStat: a MATLAB toolbox for circular statistics. Journal of Statistical Software, 31(10): ??, September 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v31/i10.

Beucher:2005:WSM

[Beu05] Ottmar Beucher. Wahrscheinlichkeitsrechnung und Statistik mit Matlab: anwendungsorientierte Einführung für Ingenieure und Naturwissenschaftler. (German) [Probability calculation and statistics with Matlab: Applicationoriented Introduction for engineers and scientists]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-23416-0.

509 (est.) pp. LCCN ???? EUR 39.95; SFR 68.00.

Bucker:2006:SIM

[BEV06] H. Martin Bücker, Atya Elsheikh, and Andre Vehreschild. A system for interfacing MATLAB with external software geared toward automatic differentiation. In Mathematical software—ICMS 2006, volume 4151 of Lecture Notes in Comput. Sci., pages 373–384. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Bukowski:1995:PWN

[BF95] L. Bukowski and J. Feliks. Przyklad wykorzystania neural network toolbox — Matlab w nauczaniu przedmiotow “Automatyzacja systemow diagnostycznych” oraz “Struktury ukladow sterowania”. In Mrozek [Mro95b], pages 1–

6. ISBN 83-86547-04-9. LCCN ????

Bandilla:1997:SAS

[BF97] W. Bandilla and F. Faulbaum, editors. Softstat ’97: advances in statistical software 6: the 9th Conference on the Scientific Use of Statistical Software, March 3–6, 1997, Heidelberg, volume 6 of Advances in Statistical Software 1997. Lucius and Lucius, Stuttgart, Germany, 1997. ISBN 3-8282-0032-X. LCCN QA276.4.K66 1997.

Barnes:2009:MMC

[BF09] Belinda Barnes and Glenn Fulford. Mathematical modelling with case studies: a differential equations approach using Maple and MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2009. ISBN 1-4200-8348-1 (hardcover). ix + 349 pp. LCCN QA371.3 .B37 2009. URL http://

                                        www.loc.gov/catdir/toc/
                                        fy0904/2008049372.html.

Bindel:2014:NCB

[BFG^{+}14] D. Bindel, M. Friedman, W. Govaerts, J. Hughes, and Yu. A. Kuznetsov. Numerical computation of bifurcations in large equilibrium systems in Matlab. Journal of Computational and Applied Mathematics, 261(??):232–248, May 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005785.

Boyle:1989:MTU

[BFM89] John M. Boyle, Matthew P. Ford, and Jan M. Maciejowski. Multivariable Toolbox for use with MAT-LAB. IEEE Control Systems Magazine, 9(1):59–65, January 1989. CODEN IS-MAD7. ISSN 0730-6598.

Barrasa-Fano:2021:TMT

[BFSJP^{+}21] Jorge Barrasa-Fano, Apeksha Shapeti, Álvaro Jorge-Peñas, Mojtaba Barzegari, José Antonio Sanz-Herrera, and Hans Van Oosterwyck. TFMLAB: a MAT-LAB toolbox for 4D traction force microscopy. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000625.

Burrus:1998:IWW

[BGG98] C. Sidney Burrus, Ramesh A.

Gopinath, and Haitao Guo. Introduction to wavelets and wavelet transforms: a primer. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-489600-9. xiv + 268 pp. LCCN QA403.3 .B87 1998. With additional material and programs by Jan E. Odegard and Ivan W. Selesnick.

Ben-Gida:2020:OMO

[BGGL20] Hadar Ben-Gida, Roi Gurka, and Alex Liberzon. OpenPIV-MATLAB — an opensource software for particle image velocimetry; test case: Birds’ aerodynamics. SoftwareX, 12(??):Article 100585, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711020302983.

Bourbouh:2021:LSR

[BGGT21] Hamza Bourbouh, Pierre-Loïc Garoche, Christophe Garion, and Xavier Thirioux. From Lustre to Simulink: Reverse compilation for embedded systems applications. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3):31:1– 31:20, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3461668.

Beck:2014:CAK

[BGPPRW14] S. Beck, S. González-Pinto, S. Pérez-Rodríguez, and R. Weiner. A comparison of AMF- and Krylov-methods in Matlab for large stiff ODE systems. Journal of Computational and Applied Mathematics, 262(??):292– 303, May 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042713005153.

Breitenecker:1995:ESC

[BH95] Felix Breitenecker and Irmgard Husinsky, editors. EUROSIM ’95: simulation congress: proceedings of the EUROSIM Conference, EUROSIM ’95, Vienna, Austria, 11–15 September 1995. North-Holland Publishing Co., Amsterdam, The Netherlands, 1995. ISBN 0-444-82241-0. LCCN QA76.9.C65E966 1995.

Baha:1996:DMR

[BH96] B. Baha and D. C. Hamill.

The design and modelling of resonant switched mode power supply (SMPS) using Simulink and Matlab. In Silvester [Sil96], pages 35–44. ISBN 1-85312-395-1, 1-85312-385-1 (??invalid ISBN checksum??). LCCN TK5.I59 1996.

Brown:1997:IRS

[BH97] Robert Grover Brown and Patrick Y. C. Hwang. Introduction to random signals and applied Kalman filtering: with MATLAB exercises and solutions. Wiley, New York, NY, USA, third edition, 1997. ISBN 0-471-12839-2 (cloth). xi + 484 pp. LCCN TK5102.9.B75 1997.

Bhatnagar:1995:MCT

[Bha95] Tanmaya Shubham Bhatnagar. MATLAB courseware to teach chemically mediated transport. Thesis (M. Eng.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1995. 127 pp.

Bhatti:2005:FFE

[Bha05] M. Asghar Bhatti. Fundamental Finite Element Analysis and Applications: with Mathematica and MAT-LAB Computations: Fundamental Concepts. Wiley, New York, NY, USA, 2005. ISBN 0-471-64808-6. xx + 700 pp. LCCN TA646 .B56 2005. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip0414/2004002270.
html.

Bhatti:2006:ATF

[Bha06] M. Asghar Bhatti. Advanced topics in finite element analysis of structures: with Mathematica and MATLAB computations. Wiley, New York, NY, USA, 2006. ISBN 0-471-64807-8 (cloth). xvi + 590 pp. LCCN TA647 .B494 2006. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip058/2005005179.
html.

Belattmania:2024:CMG

[BHC^{+}24] Ayoub Belattmania, Mounir Hakkou, Taoufiq Chtioui, Abdelhaq Aangri, Assia Abdenour, El mehdi Latni, Abdessalam Benharra, Abdelkrim EL Arrim, Azdine Dahaoui, Ahmed Raissouni, Lamiae Khali Issa, and Lhoussaine Ed-daoudy.

Classwms: a MATLAB GUI for water masses classification using clustering analysis and K nearest neighbors. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711024001122.

Bingham:2000:AVM

[Bin00] John A. C. Bingham.

ADSL, VDSL, and multicarrier modulation. Wiley series in telecommunications and signal processing. Wiley, New York, NY, USA, 2000. ISBN 0-471-29099-8. xvii + 289 pp. LCCN TK5103.7 .B535 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley041/99015963.
html;   http://www.loc.
gov/catdir/description/
wiley032/99015963.html;
http://www.loc.gov/catdir/
toc/onix02/99015963.html.

Biran:2003:SHS

[Bir03] Adrian Biran. Ship hydrostatics and stability. Butterworth-Heinemann, Oxford, UK, 2003. ISBN 0-7506-4988-7. xvi + 344 pp.

Birkisson:2018:ARO

[Bir18]  Ásgeir Birkisson.  Au-

tomatic reformulation of ODEs to systems of firstorder equations. ACM Transactions on Mathematical Software, 44(3):31:1– 31:18, April 2018. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3159443.

Bishop:1993:MCS

[Bis93] Robert H. Bishop. Modern Control Systems Analysis and Design Using MAT-LAB. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-59657-1. viii + 161 pp. LCCN TJ213 .B5357 1993. Companion text to: Modern control systems. 6th ed. / Richard C. Dorf.

Bishop:1997:MCS

[Bis97] Robert H. Bishop. Modern control systems analysis and design using MATLAB and SIMULINK. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-49846-4. xii + 251 pp. LCCN TJ216 .B53 1997.

Brosilow:2002:TMB

[BJ02] Coleman Brosilow and Babu Joseph. Techniques of model-based control. Prentice-Hall international series in the physical and chemical engineering sciences. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-028078-X. xxi + 680 pp. LCCN TS156.8 .B755 2002.

Bals:1994:AED

[BJG94] J. Bals, H.-D. Joos, and G. Grubel. ANDECS– Matlab environment for databased computational control experimenting. In Mattsson et al. [MGC94], pages 233–238. ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Bjornaraa:2001:PPN

[Bjø01] Tore I. Bjørnarå, editor.

Program and proceedings for the Nordic MATLAB Conference 2001: October 17– 18, 2001, Radisson SAS Scandinavia Hotel, Oslo. Comsol, Trondheim, Norway, 2001. ISBN 82-995955-0-9. LCCN ????

Bader:2006:AMT

[BK06] Brett W. Bader and Tamara G.

Kolda. Algorithm 862: MATLAB tensor classes for fast algorithm prototyping. ACM Transactions on Mathematical Software, 32(4):635–653, December 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bader:2007:EMC

[BK07a] Brett W. Bader and Tamara G.

Kolda. Efficient MATLAB computations with sparse and factored tensors. SIAM Journal on Scientific Computing, 30(1):205–231, 2007. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Bliss:2007:PPM

[BK07b] N. Travinin Bliss and J. Kepner. pMATLAB parallel MATLAB library. The International Journal of High Performance Computing Applications, 21(3):336– 359, August 2007. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/21/
3/336.full.pdf+html.

Bekas:2005:DDM

[BKG05] C. Bekas, E. Kokiopoulou, and E. Gallopoulos. The design of a distributed MATLAB-based environment for computing pseudospectra. Future Generation Computer Systems, 21(6):930–941, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bekas:2002:PCP

[BKGS02] Constantine Bekas, Efrosini Kokiopoulou, Efstratios Gallopoulos, and Valeria Simoncini. Parallel computation of pseudospectra using transfer functions on a MATLAB-MPI cluster platform. Lecture Notes in Computer Science, 2474: 199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/
service/series/0558/bibs/
2474/24740199.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740199.pdf.

Burgel:2019:AIM

[BKL19] Florian Bürgel, Kamil S.

Kazimierski, and Armin Lechleiter. Algorithm 1001: IPscatt — a MATLAB toolbox for the inverse medium problem in scattering. ACM Transactions on Mathematical Software, 45(4):45:1– 45:??, December 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3328525.

Barker:1997:CMR

[BKR97] T. M. Barker, C. Kirtley, and J. Ratanapinunchai. Calculation of multisegment rigid body joint dynamics using MATLAB. Proceedings of the Institution of Mechanical Engineers. Part H, Journal of engineering in medicine, 211 (6):483–??, ???? 1997. ISSN 0954-4119.

Bose:1996:SMN

[BL96a] N. K. Bose and P. Liang.

Solutions Manual for Neural Network Fundamentals with Graphs, Algorithms, and Applications. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-006619-1. ???? pp. LCCN ???? This manual is for the book [BL96b].

Bose:1996:NNF

[BL96b] N. K. (Nirmal K.) Bose and Ping Liang. Neural Network Fundamentals with Graphs, Algorithms, and Applications. McGraw-Hill series in electrical and computer engineering. Communications and signal processing. Mc-Graw-Hill, New York, NY, USA, 1996. ISBN 0-07-006618-3. xxxiii + 478 pp. LCCN QA76.87 .B68 1996. An instructor’s manual is available [BL96a].

Boukas:2002:DST

[BL02] El-Kébir Boukas and Zi-Kuan Liu. Deterministic and stochastic time delay systems. Control engineering. Birkhäuser Verlag, Basel, Switzerland, 2002. ISBN 0-8176-4245-5. xvi + 423 pp. LCCN TJ216 .B65 2002.

Blaga:2002:SPM

[Bla02a] Petru Blaga. Statistică … prin Matlab. Presa Universitară Clujeană, Cluj-Napoca, 2002. ISBN 973-610-096-0. x + 391 pp.

Blake:2002:ECS

[Bla02b] Roy Blake. Electronic communication systems. Delmar/Thomson Learning, Albany, NY, USA, second edition, 2002. ISBN 0-7668-2684-8. xx + 985 pp. LCCN TK5101 .B558 2002.

Boudet:2020:FHR

[BlD^{+}20] S. Boudet, A. Houzé l’Aulnoit, R. Demailly, A. Delgranche, L. Peyrodie, R. Beuscart, and D. Houzé de l’Aulnoit. A fetal heart rate morphological analysis toolbox for MATLAB. SoftwareX, 11(??):Article 100428, January/June 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018302498.

Bender:2015:SRM

[BLL^{+}15] Marc Bender, Karen Laurin, Mark Lawford, Vera Pantelic, Alexandre Korobkine, Jeff Ong, Bennett Mackenzie, Monika Bialy, and Steven Postma. Signature required: Making Simulink data flow and interfaces explicit. Science of Computer Programming, 113 (part 1)(??):29–50, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642315001392.

Boullart:1997:CAC

[BLM97] L. Boullart, M. Loccufier, and S. E. Mattsson, editors. Computer aided control systems design, CACSD ’97: a proceedings volume from the 7th IFAC Symposium, Gent, Belgium, 28-30 April 1997. Pergamon Press, New York, NY, USA, 1997. ISBN 0-08-042383-3. LCCN TJ212.2 .C33 1997.

Brdys:1994:CAC

[BM94] Mieczyslaw A. Brdys and Krzysztof Malinowski, editors. Computer aided control system design: methods, tools, and related topics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994. ISBN 981-02-1391-3. LCCN TJ213 .C568 1994.

Bouajjani:2009:CAV

[BM09] Ahmed Bouajjani and Oded Maler, editors. Computer Aided Verification: 21st International Conference, CAV 2009, Grenoble, France, June 26–July 2, 2009, Proceedings, volume 5643 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-02657-5 (paperback), 3-642-02658-3. LCCN QA76.76.V47 .C38 2009.

Bini:2019:QTM

[BMR19] Dario A. Bini, Stefano Massei, and Leonardo Robol. Quasi-Toeplitz matrix arithmetic: a MATLAB toolbox. Numerical Algorithms, 81(2):741–769, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Boldea:1999:ED

[BN99] I. Boldea and S. A. Nasar.

Electric drives. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-2521-8. 411 (est.) pp. LCCN TK4058 .B64 1999.

Baican:2000:AVI

[BN00] R. (Roman) Baican and D. S. (Dan S.) Necsulescu. Applied virtual instrumentation. WIT, Southampton, UK, 2000. ISBN 1-85312-800-7. 277 (est.) pp. LCCN TK7878.4 .B3 2000. US$85.00.

Boggess:2001:FCW

[BN01] Albert Boggess and Francis J. Narcowich. A first course in wavelets with Fourier analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-022809-5. xix + 283 pp. LCCN QA403.3 .B64 2001.

Binh:2010:UFF

[BN10] Le Nguyen Binh and Nam Quoc Ngo. Ultrafast fiber lasers: principles and applications with MAT-LAB models. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-1128-8. ???? pp. LCCN TA1800 .B56 2010.

Binh:2011:UFF

[BN11] Le Nguyen Binh and Nam Quoc Ngo. Ultrafast fiber lasers: principles and applications with MATLAB models. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-1130-X (electronic book), 1-4398-1128-8. xviii + 419 pp. LCCN TA1800 .B56 2011.

Bates:2016:BMI

[BNN16] Daniel J. Bates, Andrew J. Newell, and Matthew Niemerg. BertiniLab: A MATLAB interface for solving systems of polynomial equations. Numerical Algorithms, 71(1):229–244, January 2016. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11075-015-0014-6.

Boashash:2019:ESP

[BO19] Boualem Boashash and Samir Ouelha. Efficient software platform TFSAP 7.1 and Matlab package to compute time-frequency distributions and related time-scale methods with extraction of signal characteristics. SoftwareX, 8 (??):48–52, ???? 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711017300353.

Bobal:2005:DST

[Bob05] V. (Vladimír) Bobál, editor. Digital self-tuning controllers: algorithms, implementation and applications. Advanced textbooks in control and signal processing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-85233-980-2 (paperback). xvi + 317 pp. LCCN TJ217 .D54 2005.

Bode:1998:MRG

[Bod98] Helmut Bode. MATLAB in der Regelungstechnik. (German) [MATLAB in Control Engineering]. Teubner, Stuttgart, Germany; Leipzig, Germany, 1998. ISBN 3-519-06252-6. xiii + 356 pp. LCCN ????

Bolzern:1994:PMP

[Bol94] Paolo Bolzern. Programmi MATLAB per Esercitazioni di Elementi di Automatica [English: MATLAB Programs for Exercises in Elements of Automation]. Masson, Masson, France, 1994. ISBN 88-214-0678-4. ???? pp. LCCN ????

Bolton:1998:CE

[Bol98] W. (William) Bolton. Control engineering. Longman, Harlow, UK, second edition, 1998. ISBN 0-582-32773-3. xii + 397 pp.

Boone:1998:SPU

[Boo98] Bradley G. (Bradley Gilbert) Boone. Signal processing using optics: fundamentals, devices, architectures, and applications. The Johns Hopkins University/Applied Physics Laboratory series in science and engineering. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-508424-1 (cloth). xxi + 394 pp. LCCN TA1632 .B66 1998.

Boon:2004:STT

[Boo04] John D. Boon. Secrets of the tide: tide and tidal current analysis and applications, storm surges and sea level trends. Horwood Publishers, Chichester, UK, 2004. ISBN 1-904275-17-6. xi + 212 pp. LCCN GC303 .B66 2004.

Borse:1997:NMM

[Bor97] G. J. (Garold J.) Borse. Numerical methods with MAT-LAB: a resource for scientists and engineers. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1997. ISBN 0-534-93822-1. xvi + 640 pp. LCCN QA297.B647 1997.

Bornemann:2018:NLA

[Bor18] Folkmar Bornemann. Numerical Linear Algebra: a Concise Introduction with MATLAB and Julia. Springer Undergraduate Mathematics Series. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-319-74221-3, 3-319-74222-1 (e-book). ISSN 1615-2085. x + 153 pp. LCCN QA184-205; QA297-299.4.

Bose:2001:MPE

[Bos01] Bimal K. Bose. Modern power electronics and AC drives. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-016743-6. xxi + 711 pp.

Bose:2002:MPE

[Bos02] Bimal K. Bose. Modern power electronics and AC drives. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-016743-6. xxi + 711 pp. LCCN TK2781 .B67 2002; TK 2781 .B67 2002X ENGI.

Boukas:1995:SAF

[Bou95] El-Kebir Boukas. Systèmes Asservis (French) [Feedbak Control Systems]. Editions de L’École Polytechnique de Montréal, Montréal, PQ, Canada, 1995. ISBN 2-553-00430-3. ???? pp. LCCN ????

Bouniaev:1997:LAM

[Bou97] M. M. Bouniaev. Linear algebra with MATLAB package in preservice teacher education. In Anonymous [Ano97i], pages 588–592. ISBN 1-880094-25-8. LCCN ????

Bowman:2010:BRB

[Bow10] Adrian Bowman. Book review: Functional Data Analysis with R and MAT-LAB. Journal of Statistical Software, 34(BR-3):??, April 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v34/b03.

Boyd:1999:TAE

[Boy99] Robert (Robert R.) Boyd. Tolerance analysis of electronic circuits using MAT-LAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-2276-6. x + 147 pp. LCCN TK7867 .B65 1999.

Boyd:2015:FWC

[Boy15] John P. Boyd. Four ways to compute the inverse of the complete elliptic integral of the first kind. Computer Physics Communications, 196(??):13–18, November 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465515001733.

Baiocchi:1996:PPM

[BP96] O. R. Baiocchi and S. A.

Pradels. Pulse propagation: Modeling and simulation with MATLAB. In Ingalls et al. [ICS96], pages 530–534. ISBN 1-56555-098-

6. ISSN 0094-7474. LCCN ????

Bobal:1997:STC

[BP97] V. Bobal and R. Prokop.

Self-tuning control Matlab toolbox — methodology and design. In Boullart et al. [BLM97], pages 127–132. ISBN 0-08-042383-3. LCCN TJ212.2 .C33 1997.

Bishop:1996:VPP

[BPB96] W. J. Bishop, S. A. Pradels, and O. R. Baiocchi. Visualization of pulse propagation phenomena with MATLAB and wavefront software. In Iskander et al. [I^{+}96], pages 246–249. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Barreto-Parra:2023:OID

[BPCCM23] German Francisco Barreto-Parra, Brandon Cortés-Caicedo, and Oscar Danilo Montoya. Optimal integration of D-STATCOMs in radial and meshed distribution networks using a MATLAB–GAMS interface. Algorithms (Basel), 16(3), March 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/
16/3/138.

Brixius:1999:SMI

[BPS99] Nathan Brixius, Florian A. Potra, and Rongqin Sheng. SDPHA: a MATLAB implementation of homogeneous interior-point algorithms for semidefinite programming. Optimization Methods and Software, 11/12(1-4):583– 596, 1999. CODEN OM-SOE2. ISSN 1055-6788. Interior point methods.

Bateman:2002:DHA

[BPS02] Andrew Bateman and Iain Paterson-Stephens. The DSP handbook: algorithms, applications and design techniques. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-201-39851-6. xiii + 664 + 16 pp. LCCN TK5102.5 .B38 2002.

Berry:2005:ACS

[BPS05] Michael W. Berry, Shakhina A. Pulatova, and G. W. Stewart. Algorithm 844: Computing sparse reduced-rank approximations to sparse matrices. ACM Transactions on Mathematical Software, 31(2):252–269, June 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bientinesi:2005:RLA

[BQOvdG05] Paolo Bientinesi, Enrique S.

Quintana-Ortí, and Robert A. van de Geijn. Representing linear algebra algorithms in code: the FLAME application program interfaces. ACM Transactions on Mathematical Software, 31(1):27–59, March 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Baker:1996:MCS

[BR96] J. R. Baker and K. E.

Rouch. Modeling of complex systems using ANSYS and Matlab. In Bryan [Bry96], pages 2.221–2.234. ISBN ???? LCCN ????

Bramley:1997:TNR

[Bra97a] Randall Bramley. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific Work-Place 3.0. IEEE Computational Science & Engineering, 4(4):75–78, October/ December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://dlib.
computer.org/cs/books/
cs1997/pdf/c4075.pdf.

Bramley:1997:TNF

[Bra97b] Randall Bramley. Technology news: Fortran 90 news; free software; symbolic computing packages; Matlab 5; Web products. IEEE Computational Science & Engineering, 4(1):87, 89–90, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://dlib.
computer.org/cs/books/
cs1997/pdf/c1087.pdf;
http://www.cs.berkeley.
edu/~xiaoye;   http://
www.cs.cornell.edu/home/
vavasis/qmg-home.html;
 http://www.globus.org/
sage; http://www.macsyma.
com/;   http://www.mathworks.
com/; http://www.netlib.
org/benchmark/linpackjava;
 http://www.netlib.org/
scalapack;   http://www.
ucmp.berkeley.edu/subway/
phylogen.html;   http:
//www.vni.com/products/
wpd/jnl/jnl_1_0.html;
http://www.wolfram.com/
look/cse.

Brandimarte:2002:NMF

[Bra02a] Paolo Brandimarte. Numerical methods in finance: a MATLAB-based introduction. Wiley series in probability and statistics. Wiley, New York, NY, USA, 2002. ISBN 0-471-39686-9 (cloth). xv + 403 pp. LCCN HG176.5 .B73 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley042/2001026767.
html;   http://www.loc.
gov/catdir/description/
wiley035/2001026767.html;
http://www.loc.gov/catdir/
toc/onix06/2001026767.
html.

Brandts:2002:MCS

[Bra02b] J. H. Brandts. Matlab code for sorting real Schur forms. Numerical linear algebra with applications, 9 (3):249–261, 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brandimarte:2006:NMF

[Bra06] Paolo Brandimarte. Numerical methods in finance and economics: a MATLABbased introduction. Statistics in practice. Wiley-Interscience, New York, NY, USA, second edition, 2006. ISBN 0-471-74503-0 (cloth). xxiv + 669 pp. LCCN HG176.5 .B73

2006. URL http://www.
loc.gov/catdir/enhancements/
fy0826/2006045787-b.html;
http://www.loc.gov/catdir/
enhancements/fy0826/2006045787-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0826/2006045787-t.html.

Brereton:2003:CDA

[Bre03] Richard G. Brereton. Chemometrics: data analysis for the laboratory and chemical plant. Wiley, New York, NY, USA, 2003. ISBN 0-471-48977-8 (hardback), 0-470-84911-8 (paperback). xiv + 489 pp. LCCN QD75.4.S8 B74 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley036/2002027212.
html;   http://www.loc.
gov/catdir/toc/wiley031/
2002027212.html.

Brockhaus:1994:FPG

[Bro94]  R. Brockhaus.  Flu-

gregelung: physikalische Grundlagen, mathematisches Flugzeugmodell, Auslegungskriterien-Regelungs-strukturen, Entwurf von Flugregelungssystemen, Entwicklungslinien [English: Flight Control: Physical Foundation, Mathematical Airplane Model, Design Criteria and Control Structures, Development Directions]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-55416-5. xxii + 820 pp. LCCN TL678.B67 1994.

Brown:1995:SPT

[Bro95] A. Brown. Signal processing in a toolbox: An extension to Matlab specifically designed for solving DSP problems. Electronics world + wireless world, ??(12):1040– ??, December 1995. CO-DEN EWWWE6. ISSN 0959-8332.

Brown:2001:ESD

[Bro01] Forbes T. Brown. Engineering system dynamics: a unified graph-centered approach, volume 8 of Control engineering. Marcel Dekker, New York, NY, USA, 2001. ISBN 0-8247-0616-1. xx + 1006 pp. LCCN TA168 .B675 2001.

Brodtkorb:2007:MIG

[Bro07] André Rigland Brodtkorb.

A MATLAB interface to the GPU. Master’s thesis, University of Oslo, Department of Informatics, Oslo, Norway, May 1, 2007. xiv + 108 pp. URL http:

//babrodtk.at.ifi.uio.
no/files/publications/
brodtkorb_msc_thesis.pdf.

Ba-Razzouk:1994:SBS

[BRPCR94] A. Ba-Razzouk, A. Pittet, A. Cheriti, and V. Rajagopalan. Simulink based simulation of power electronic systems. In IEEE [IEE94a], pages 105–108. ISBN 0-7803-2091-3, 0-7803-2092-1. LCCN TK7881.15 .I35 1994.

Brillon:1994:PES

[BRS94] D. Brillon, V. Rajagopalan, and V. K. Sood. Performance evaluation of simulation studies of a HVDC transmission system in ATP, ATOSEC5 and SIMULINK environments. In IEEE [IEE94a], pages 96–104. ISBN 0-7803-2091-3, 0-7803-2092-1. LCCN TK7881.15 .I35 1994.

Bruce:2001:BSP

[Bru01] Eugene N. Bruce. Biomedical signal processing and signal modeling. Wiley series in telecommunications and signal processing; Wiley-Interscience publication. Wiley, New York, NY, USA, 2001. ISBN 0-471-34540-7 (cloth). xiv + 520 pp. LCCN R857.S47 B78 2001; R 857 .S47B78 2001X GERSTM.

Bryson:1994:CSA

[Bry94] Arthur E. (Arthur Earl) Bryson, Jr. Control of Spacecraft and Aircraft. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-08782-2. xxiii + 378 pp. LCCN TL3260 .B78 1994. US$49.50. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/prin021/92029102.
html;   http://www.loc.
gov/catdir/toc/prin031/
92029102.html.

Bryan:1996:AFS

[Bry96] W. J. Bryan, editor. Ansys: The future of simulation tools, computer aided engineering in the 21st century:

International conference; 7th — May 1996, Pittsburgh; PA, volume 2 of AN-SYS CONFERENCE PRO-CEEDINGS 1996. Swansea Analysis Systems, Houston, TX, USA, 1996. ISBN ???? LCCN ????

Bryson:2002:ALO

[Bry02] Arthur E. (Arthur Earl)

Bryson. Applied linear optimal control: examples and algorithms. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-81285-2 (hardback), 0-521-01231-7 (paperback). xxi + 362 pp. LCCN TJ220 .B78 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam022/2001052553.
html;   http://www.loc.
gov/catdir/samples/cam031/
2001052553.html;   http:
//www.loc.gov/catdir/toc/
cam024/2001052553.html.

Brzozka:1997:CZA

[Brz97] J. Brzozka. Cwiczenia Z Automatyki W Matlabie I Simulinku (Polish) [Exercises for Automatic Control Theory Using MATLAB and Simulink]. MIKOM, Warszawa, Poland, 1997. ISBN 83-87102-25-3. ???? pp. LCCN ????

Blanke:1995:SIS

[BS95a] M. Blanke and T. Söderström, editors. System identification (SYSID’94): a postprint volume from the IFAC symposium, Copenhagen, Denmark, 4–6 July 1994, volume 3 of SYSID — Symposium. Pergamon Press, New York, NY, USA, 1995. ISBN 0-08-042225-X. LCCN QA402.S957 1995. Three volumes.

Bradley:1995:C

[BS95b] Gerald L. Bradley and Karl J. Smith. Calculus. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-178617-2. ???? pp. LCCN QA303.B88218 1995. A supplemental lab manual is available [HC95].

Bautista:2020:MSM

[BSB20] Ilia Bautista, Sudeep Sarkar, and Sanjukta Bhanja. MatlabHTM: a sequence memory model of neocortical layers for anomaly detection. SoftwareX, 11(??):Article 100491, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711018302863.

Bujok:2023:CSE

[BSB23] Maksymilian Bujok, Alicja Smoktunowicz, and Grzegorz Borowik. On computing the symplectic LL^{T} factorization. Numerical Algorithms, 93(3):1401– 1416, July 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

https://link.springer.
com/article/10.1007/s11075-
022-01472-y.

Banerjee:2000:MCD

[BSC^{+}00] P. Banerjee, N. Shenoy, A. Choudhary, S. Hauck, C. Bachmann, M. Haldar, P. Joisha, A. Jones, A. Kanhare, A. Nayak, S. Periyacheri, M. Walkden, and D. Zaretsky. A MATLAB compiler for distributed, heterogeneous, reconfigurable computing systems. In IEEE, editor, 2000 IEEE Symposium on Field-Programmable Custom Computing Machines, Napa Valley, California, 17–19 April 2000, pages 39– 48. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN ???? LCCN ????

Bar-Shalom:1993:ETP

[BSL93] Yaakov Bar-Shalom and Xiao Rong Li. Estimation and Tracking: Principles, Techniques, and Software. Artech House Inc., Boston, MA, USA, 1993. ISBN 0-89006-643-4. xxii + 511 pp. LCCN TK6580.B32 1993.

Bar-Shalom:2001:EAT

[BSLK01] Yaakov Bar-Shalom, Xiao-Rong Li, and Thiagalingam Kirubarajan. Estimation with applications to tracking and navigation. Wiley, New York, NY, USA, 2001. ISBN 0-471-41655-X (cloth). xxiii + 558 pp. LCCN TJ214.5 .B37 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley043/2001022366.
html;   http://www.loc.
gov/catdir/description/
wiley035/2001022366.html;
 http://www.loc.gov/catdir/
toc/onix07/2001022366.
html.

Breidt:2023:IGB

[BSS^{+}23] Fred Breidt, Caitlin R. Skinner, Mileah Shriner, Mollie Ruinsky, Seo Young Yang, Robert P. Wine, and Lynette Johnston.

IngredientDB: a GUI-based

Matlab database program for estimating the pH of acid or acidified food formulations from buffer capacity models. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002418.

Berland:2007:EMP

[BSW07] Håvard Berland, Bård Skaflestad, and Will M. Wright. EXPINT — a MATLAB package for exponential integrators. ACM Transactions on Mathematical Software, 33(1):4:1– 4:17, March 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Battles:2004:EMC

[BT04] Zachary Battles and Lloyd N.

Trefethen. An extension of MATLAB to continuous functions and operators. SIAM Journal on Scientific Computing, 25(5):1743– 1770, September 2004. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43012.

Bober:2009:NAM

[BTM09] William Bober, Chi-Tay Tsai, and Oren Masory. Numerical and analytical methods with MATLAB(R). CRC Series in Computational Mechanics and Applied Analysis. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-9356-8. viii + 452 pp.

Bubnicki:1995:PIC

[Bub95] Zdzisław Bubnicki, editor. Proceedings of the 12th International Conference on Systems Science: 12–15 September 1995, Wrocław, Poland, volume 1(12) of Proceedings of the International Conference on Systems Science. Oficyna Wydawnicza Politechniki Wrocławskiej, Wrocław, Poland, 1995.

ISBN 83-7085-152-5. LCCN TA168.I526 1995. Three volumes.

Budny:1995:FEA

[Bud95] Dan Budny, editor. Frontiers in Education 1995: 25th Annual Conference: proceedings, November 1– 4, 1995, Atlanta, Georgia: engineering education for the 21st century. ETP/Harrison, New York, NY, USA, 1995. ISBN 0-7803-3023-4, 0-7803-3022-6, 0-7803-3024-2, 0-7803-3025-0. ISSN 0190-5848. LCCN T62 .F76 1995. Two volumes.

Bugl:1995:DEM

[Bug95a]  Paul Bugl.  Differential

Equations: Matrices and Models. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-02-316540-5. various pp. LCCN QA372.B874 1995.

Bugl:1995:EDE

[Bug95b] Paul Bugl. Explorations in Differential Equations using Matlab 4.0. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-374760-3. 147 pp. LCCN QA371.5.D37B84 1995.

Bunger:2020:TMT

[Bün20] Florian Bünger. A Taylor model toolbox for solving ODEs implemented in MATLAB/INTLAB. Journal of Computational and Applied Mathematics, 368 (??):Article 112511, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042719305163.

Burrus:1993:TFD

[Bur93] C. S. Burrus. Teaching filter design using Matlab. Proceedings of the International Conference on Acoustics, Speech, and Signal Processing, 1:I–20–I–22, 1993. CODEN IPRODJ. ISBN 0-7803-0946-4. ISSN 0736-7791. IEEE catalog number 93CH3252-4.

Burrus:1994:TDU

[Bur94] C. S. Burrus. Teaching DSP using Matlab. In IEEE [IEE94b], pages 127–

130. ISBN 0-7803-1948-6, 0-7803-1949-4. LCCN TK5102.9 .I328 1994.

Burl:1999:LOC

[Bur99] Jeffrey B. Burl. Linear optimal control: H_{2} and H_{∞} methods. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 1999. ISBN 0-201-80868-4, 0-201-80869-2 (Solutions Manual). xxi + 462 pp. LCCN TJ220 .B87

1999. URL http://www.
ece.mtu.edu/ee/faculty/
burl/book.htm;   http:/
/www.ece.mtu.edu/ee/faculty/
burl/errata_frame.htm.

Burns:2001:ACE

[Bur01]  Roland S. Burns.  Ad-

vanced control engineering. Butterworth-Heinemann, Oxford, UK, 2001. ISBN 0-7506-5100-8. xiii + 450 pp. LCCN TJ213 .B87 2001. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/els051/2004351720.
                                        html;   http://www.loc.
                                        gov/catdir/toc/els051/
                                        2004351720.html.

Burden:2005:NA

[Bur05] Richard L. Burden. Numerical analysis. Brooks/Cole, Pacific Grove, CA, USA, eighth edition, 2005. ISBN 0-534-39200-8. ???? pp.

Burr:2010:BRB

[Bur10] Tom Burr. Book review: Functional Data Analysis With R and MATLAB by J. Ramsay; G. Hooker; S. Graves. Technometrics, 52 (4):466, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

                                        http://www.jstor.org/stable/
                                        40997262.

Butler:1992:IMS

[But92] B. P. Butler. Interfacing Matlab and software written in other languages. NPL Report DITC 200/92, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, May 1992. ???? pp.

Butt:2008:IAL

[But08] Rizwan Butt. Introduction to applied linear algebra with MATLAB, volume 7 of Sigma Series in Applied Mathematics. Heldermann Verlag, Lemgo, Germany, 2008. ISBN 3-88538-407-8. xiv + 517 pp.

Butt:2011:ALA

[But11] Rizwan Butt. Applied linear algebra and optimization using MATLAB. Mercury Learning and Information, Dulles, VA, USA, 2011. ISBN 1-936420-04-X. ???? pp. LCCN ????

Babuska:1995:FMC

[BV95] R. Babuska and H. B. Verbruggen. Fuzzy modelling and cluster analysis toolbox for MATLAB. In Anonymous [Ano95f], pages 1479– 1483. ISBN 3-930911-67-1 (??invalid ISBN checksum??). LCCN ???? Three volumes.

Bucker:2008:CVN

[BV08] H. Martin Bücker and Andre Vehreschild. Coping with a variable number of arguments when transforming MATLAB programs. In Bischof et al. [BBH^{+}08], pages 211–222. CODEN LNCSA6. ISBN 3-540-68935-4 (print), 3-540-68942-7 (e-book). ISSN 1439-7358. LCCN QA304 .I58 2008. URL http://

link.springer.com/content/
pdf/10.1007/978-3-540-
68942-3_19.

Boyd:2018:IAL

[BV18] Stephen P. In. Boyd and Lieven Vandenberghe. Introduction to Applied Linear Algebra: Vectors, Matrices, and Least Squares. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-108-69394-6, 1-316-51896-5 (hardcover). xii + 463 pp. LCCN QA184.2 .B69 2018.

Bos:2019:CMC

[BV19] Len Bos and Marco Vianello. CaTchDes: MATLAB codes for Caratheodory–Tchakaloff near-optimal regression designs. SoftwareX, 10 (??):Article 100349, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711019302274.

Baeyens:2021:FAC

[BV21] Toon Baeyens and Marnix Van Daele. The fast and accurate computation of eigenvalues and eigenfunctions of time-independent one-dimensional Schrödinger equations. Computer Physics Communications, 258(??):Article 107568, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S001046552030271X.

Benner:2020:MMO

[BW20] Peter Benner and Steffen W. R. Werner. MORLAB

— a model order reduction framework in MATLAB and Octave. In Bigatti et al. [BCD^{+}20], pages 432–441.

Benson:2008:ADS

[BY08] Steven J. Benson and Yinyu Ye. Algorithm 875: DSDP5software for semidefinite programming. ACM Transactions on Mathematical Software, 34(3):16:1–16:20, May 2008. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cameron:1996:PMS

[C^{+}96] Greg Cameron et al., editors. Proceedings of the 39th Midwest Symposium on Circuits and Systems: August 18–21, 1996, Scheman Continuing Education Building, Iowa State University, Ames, Iowa, volume 39. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3637-2, 0-7803-3636-4, 0-7803-3638-0. LCCN TK3226 .M553 1996. Three volumes. IEEE catalog number: 96CH35995.

Ciarlini:1997:AMT

[C^{+}97] P. Ciarlini et al., editors. Advanced mathematical tools in metrology: Third International Workshop on Advanced Mathematical Tools in Metrology held in Berlin, Germany, in September 1996, volume 45 of Series on advances in mathematics for applied sciences. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1997. ISBN 981-02-2918-6. LCCN QA465.A283 1997.

Coisson:1997:MCC

[CA97] R. Coïsson and G. Allodi. Matlab compiler and C math library simplify programming chores. Computers in Physics, 11(1):89– ??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/
10.1063/1.4822521.

Cunkas:2010:RFL

[ÇA10] Mehmet Çunkas and Omer Aydoğdu. Realization of fuzzy logic controlled brushless DC motor drives using Matlab/Simulink. Mathematical and Computational Applications, 15(2):218– 229, August 2010. CO-DEN ???? ISSN 2297-8747.

URL https://www.mdpi.
com/2297-8747/15/2/218.

Chen:1994:AIC

[CAC94] J. G. Chen, F. G. Attia, and Donald L. Crabtree, editors. EXPERSYS-94: expert systems applications and artificial intelligence, Gournaysur-Marne, France. IITT International, Gournay-sur-Marne, France, 1994. ISBN 2-907669-30-3. LCCN ????

Cameron:2006:MPA

[Cam06] Frank Cameron. A Matlab package for automatically generating Runge– Kutta trees, order conditions, and truncation error coefficients. ACM Transactions on Mathematical Software, 32(2):274–298, June 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Candy:1986:SPM

[Can86]  James V. Candy.  Sig-

nal Processing: The Model-Based Approach. McGraw-Hill, New York, NY, USA, 1986. ISBN 0-07-009725-9. xv + 240 pp. LCCN TK5102.5 .C27 1986.

Cantero:2022:TTT

[Can22a] Daniel Cantero. TTB-2D:

Train–Track–Bridge interaction simulation tool for Matlab. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001716.

Cantero:2022:VEM

[Can22b] Daniel Cantero. VEqMon2D — equations of motion generation tool of 2D vehicles with Matlab. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022000693.

Cantero:2024:VRV

[Can24] Daniel Cantero. VBI-2D — road vehicle-bridge interaction simulation tool and verification framework for Matlab. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711024000967.

Caplan:2013:NNS

[Cap13] R. M. Caplan. NLSEmagic: Nonlinear Schrödinger equation multi-dimensional Matlabbased GPU-accelerated integrators using compact high-order schemes. Computer Physics Communications, 184(4):1250–1271, April 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465512004067.

Carroll:1992:RCS

[Car92] John Carroll. The role of computer software in numerical analysis teaching. ACM SIGNUM Newsletter, 27(2):2–31, April 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Carlson:1998:SLS

[Car98] Gordon E. Carlson. Signal and linear system analysis. Wiley, New York, NY, USA, second edition, 1998. ISBN 0-471-12465-6 (cloth). xvi + 752 pp. LCCN TK5102.5.C313 1998.

Carroll:1999:PFE

[Car99] W. F. (William F.) Carroll.

A primer for finite elements in elastic structures. Wiley, New York, NY, USA, 1999. ISBN 0-471-28345-2 (cloth). xv + 494 pp. LCCN TA653 .C37 1999. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley047/98013435.
html;   http://www.loc.
gov/catdir/description/
wiley032/98013435.html;
 http://www.loc.gov/catdir/
toc/onix02/98013435.html.

Carlson:2000:CEC

[Car00] A. Bruce Carlson. Circuits: engineering concepts and analysis of linear electric circuits. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37097-7. xx + 840 pp. LCCN TK454 .C357 2000.

Cardot:2010:BRF

[Car10] Hervé Cardot. Book review: Functional Data Analysis with R and Matlab, by Ramsay, J. O., Hooker, G., and Graves, S. Biometrics, 66(4):1312–1313, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cass:2014:TPL

[Cas14] Stephen Cass. The top 10 programming languages: Spectrum’s 2014 ranking [dataflow]. IEEE Spectrum, 51(7):68, July 2014. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Catchpole:1995:MEA

[Cat95] E. A. Catchpole. MAT-LAB: An environment for analyzing ring-recovery and recapture data. Journal of Applied Statistics, 22(5): 801–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Cathey:2001:EMA

[Cat01] Jimmie J. Cathey. Electric machines: analysis and design applying MATLAB. McGraw-Hill series in electrical and computer engineering. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-242370-6. xiv + 530 pp. LCCN TK2182 .C36 2001.

Charsooghi:2011:MPC

[CATK11] Mohammad A. Charsooghi, Ehsan A. Akhlaghi, Sharareh Tavaddod, and H. R. Khalesifard. A MAT-LAB program to calculate translational and rotational diffusion coefficients of a single particle. Computer Physics Communications, 182(2):400–408, February 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465510003620.

Cavicchi:2000:DSP

[Cav00] Thomas J. Cavicchi. Digital signal processing. Wiley, New York, NY, USA, 2000. ISBN 0-471-12472-9. xv + 793 pp. LCCN TK5102.9 .C37 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley031/99022866.
html;   http://www.loc.
gov/catdir/toc/onix01/
99022866.html.

Chauveau:1998:MEH

[CB98] S. Chauveau and F. Bodin.

MENHIR: An environment for high performance Matlab. Lecture Notes in Computer Science, 1511:27–??, 1998. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Chauveau:1999:MEH

[CB99] Stéphane Chauveau and François Bodin. Menhir: An environment for high performance Matlab. Scientific Programming, 7(3–4): 303–312, ???? 1999. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=53f7mftrrm4r73yyrqau%
                                        26referrer=parent%26backto=
                                        issue%2C9%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Cathers:1996:MEE

[CBCC96] B. Cathers, M. Boyd, E. Craig, and M. Chadwick. Modelling for environmental engineering students using Matlab and Simulink. In Aldeen [Ald96], pages 261–268. ISBN 0-7803-3173-7, 0-7803-3174-5. LCCN T65.3.I35 1996.

Chermitti:2012:DLC

[CBhM12] Ali Chermitti, Omar Bouklihacene, and Samir Mouhadjer. Design of a library of components for autonomous photovoltaic system under Matlab/Simulink. International Journal of Computer Applications, 53(??):13–19, September 2012. CODEN ???? ISSN 0975-8887. URL

https://www.ijcaonline.
org/archives/volume53/
number14/8488-2434/.

Chowell:2024:PEF

[CBL24] Gerardo Chowell, Amanda Bleichrodt, and Ruiyan Luo. Parameter estimation and forecasting with quantified uncertainty for ordinary differential equation models

using QuantDiffForecast:

a MATLAB toolbox and tutorial. Statistics in Medicine, 43(9):1826–1848, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chong:2001:IO

[CÇ01] Edwin Kah Pin Chong and Stanislaw H. Çzak. An introduction to optimization. Wiley-Interscience series in discrete mathematics and optimization. Wiley, New York, NY, USA, second edition, 2001. ISBN 0-471-39126-3. xv + 476 pp. LCCN QA402.5 .C476 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley044/2001026801.
html;   http://www.loc.
gov/catdir/description/
wiley037/2001026801.html;
 http://www.loc.gov/catdir/
toc/onix06/2001026801.
html.

Colinge:2002:PSD

[CC02] Jean-Pierre Colinge and C. A. (Cynthia A.) Colinge. Physics of semiconductor devices. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 2002. ISBN 1-4020-7018-7. xiii + 436 pp. LCCN QC611 .C787 2002; QC 611 .C787 2002X GERSTM. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        toc/fy031/2002025492.html.

Chapra:2006:NME

[CC06] Steven C. Chapra and Raymond P. Canale. Numerical methods for engineers. McGraw-Hill Higher Education, Boston, MA, USA, fifth edition, 2006. ISBN 0-07-291873-X (hard copy). xviii + 926 pp.

Chaitin-Chatelin:1996:LFP

[CCF96] Françoise Chaitin-Chatelin and Valérie Frayssé. Lectures on Finite Precision Computations. Software, environments, tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-358-7. xv + 235 pp. LCCN QA297 .C417 1996.

Castellano:2002:KMT

[CCF02] G. Castellano, C. Castiello, and A. M. Fanelli. KER-NEL: a Matlab toolbox for knowledge extraction and refinement by NEural learning. Lecture Notes in Computer Science, 2329: 970–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer-ny.com/
link/service/series/0558/
bibs/2329/23290970.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2329/23290970.
pdf.

Caspi:2003:SSL

[CCM^{+}03] Paul Caspi, Adrian Curic, Aude Maignan, Christos Sofronis, Stavros Tripakis, and Peter Niebert. From Simulink to SCADE/lustre to TTA: a layered approach for distributed embedded applications. ACM SIG-PLAN Notices, 38(7):153– 162, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Consularo:1998:MIM

[CdFCS98] Luis Augusto Consularo, Luciano da Fontoura Costa, and Donald L. Shirer. MAT-COM integrates MATLAB resources into standalone applications. Computers in Physics, 12(5):460–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

                                        aip.scitation.org/doi/
                                        10.1063/1.168722.

Cruz-Duarte:2022:MMB

[CDOBA22] Jorge M. Cruz-Duarte, José C. Ortiz-Bayliss, and Ivan Amaya. MatHH: a Matlab-based Hyper-Heuristic framework. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022000413.

Chen:2009:FFM

[CDS09] Chunqing Chen, Jin Song Dong, and Jun Sun. A formal framework for modeling and validating Simulink diagrams. Formal Aspects of Computing, 21(5):451– 483, October 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00165-009-0108-9.

Caliari:2010:PFI

[CDSV10] Marco Caliari, Stefano De Marchi, Alvise Sommariva, and Marco Vianello. Padua2DM: fast interpolation and cubature at the Padua points in Matlab/ Octave. Numerical Algorithms, 54(??):??, ???? 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

                                        /www.springerlink.com/

openurl.asp?genre=article&
issn=1017-1398&volume=
0&issue=0&spage=??.

Caliari:2011:PFI

[CDSV11] Marco Caliari, Stefano De Marchi, Alvise Sommariva, and Marco Vianello. Padua2DM: fast interpolation and cubature at the Padua points in Matlab/ Octave. Numerical Algorithms, 56(1):45–60, January 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=1017-1398&volume=
56&issue=1&spage=45.

Calio:1991:MEC

[CF91] Franca Caliò and Marco Frontini. MATLAB Esercitazioni di Calcolo Numerico Assistite da Calcolatore [English: MATLAB Exercises in Computer-Aided Numeric Computation]. clup CittàStudi, Milano, Italy, 1991. ISBN 88-251-0002-7. LCCN ???? Chipperfield:1993:SMT

[CF93a] A. J. Chipperfield and P. J.

Fleming, editors. MAT-LAB toolboxes and applications for control: Control ’91, Edinburgh, volume 48 of IEE Control Engineering Series. Peter Peregrinus Ltd, Michael Faraday House, Six Hills Way, Stevenage, Herts SG1 2AY, UK, 1993. ISBN 0-86341-290-4. ISSN 0262-1797. LCCN TJ213 .M38 1993. US$75.00.

Chipperfield:1993:MTA

[CF93b] A. J. (Andrew J.) Chipperfield and P. J. (Peter J.) Fleming, editors. MAT-LAB toolboxes and applications for control, volume 48 of IEE control engineering series. Peter Peregrinus on behalf of The Institution of Electrical Engineers, London, UK, 1993. ISBN 0-86341-290-4. xi + 237 pp. LCCN TJ213 .M38 1993; TJ213.M37.

Close:1993:MAD

[CF93c] Charles Close and Dean Frederick. Modeling and Analysis of Dynamic Systems. Houghton-Mifflin and Wiley, Boston and New York, NY, USA, second edition, 1993. ISBN 0-395-55114-5 (Houghton-Mifflin), 0-471-12517-2 (Wiley). xv + 681 pp. LCCN QA402.C53 1993.

Chipperfield:1995:MGA

[CF95] A. J. Chipperfield and P. J. Fleming. MATLAB genetic algorithm toolbox. IEE Colloquium (Digest), 014: 10/1–10/4, 1995. CODEN DCILDN. ISSN 0963-3308.

Checinski:2016:MMF

[CF16] Jakub Checiński and Marek Frankowski. MAGE (Mfile/Mif Automatic GEnerator): a graphical interface tool for automatic generation of Object Oriented Micromagnetic Framework configuration files and Matlab scripts for results analysis. Computer Physics Communications, 207(??):487– 498, October 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465516302016.

Crummey:1991:OTM

[CFF^{+}91] T. P. Crummey, R. Farshadnia, P. J. Fleming, A. C. W. Grace, and S. D. Hancock. An optimization toolbox for MATLAB. IEE Conference Publication, 2 (332):744–749, 1991. CO-DEN IECPB4. ISSN 0537-9987 (??invalid ISSN checksum??).

Choi:1994:SSL

[CFGG94] J. Choi, E. Fernandez-Gaucherand, and D. Gerhart. S02YSCODE: a software laboratory for stochastic systems control and decision algorithms, FORTRAN and MATLAB versions. In IEEE [IEE94d], pages 2528– 2533. ISBN 0-7803-1969-9, 0-7803-1968-0, 0-7803-1970-

2. ISSN 0888-3610. LCCN TJ 217 I11c 1994. Four volumes.

Close:2001:MAD

[CFN01] Charles M. Close, Dean K. Frederick, and Jonathan C. Newell. Modeling and analysis of dynamic systems. Wiley, New York, NY, USA, third edition, 2001. ISBN 0-471-39442-4, 0-471-45296-3 (International ed.). xiv + 576 pp.

Close:2002:MAD

[CFN02] Charles M. Close, Dean K. Frederick, and Jonathan C. Newell. Modeling and analysis of dynamic systems. Wiley, New York, NY, USA, third edition, 2002. ISBN 0-471-39442-4 (cloth). xiv + 576 pp. LCCN QA402 .C53 2002; QA 402 .C53 2002X MATH. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley035/2001033010.
                                        html;   http://www.loc.
                                        gov/catdir/toc/onix06/
                                        2001033010.html.

Chipperfield:1994:GAT

[CFPF94] A. J. Chipperfield, P. J. Fleming, H. Pohlheim, and C. M. Fonseca. A genetic algorithm toolbox for Matlab. In Anonymous [Ano94k], pages 200–207. ISBN 0-905949-23-4. LCCN ???? Two volumes.

Concas:2019:PMP

[CFR19] Anna Concas, Caterina Fenu, and Giuseppe Rodriguez. PQser: a Matlab package for spectral seriation. Numerical Algorithms, 80(3):879–902, March 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Choe:2000:UDS

[CG00] Won Gyu Choe and J. Guckenheimer. Using dynamical system tools in Matlab. In Numerical methods for bifurcation problems and large-scale dynamical systems (Minneapolis, MN, 1997), volume 119 of IMA Vol. Math. Appl., pages 85–113. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000.

Chen:1999:MEM

[CGI99] Ke Chen, P. J. Giblin, and Alan Irving. Mathematical explorations with MAT-LAB. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-63078-9 (hardback), 0-521-63920-4 (paperback). xiv + 306 pp. LCCN QA76.95 .C44 1999. US$24.95. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam029/99019617.
html;   http://www.loc.

                                        gov/catdir/toc/cam026/
                                        99019617.html.

Cloutier:1997:MGD

[CGL97] Guy Cloutier, Richard Gourdeau, and Daniel Leblanc. MATLAB: guide d’apprentissage. (French) [MATLAB: Learning Guide]. Cours […]; no 5167 Cours … (Ecole polytechnique (Montréal, Québec)); no 5167. École polytechnique de Montréal, Montréal, PQ, Canada, fourth edition, 1997. ISBN 2-553-00652-7. iii + 84 pp. LCCN ????

Carabelli:1995:MDM

[CGM95] S. Carabelli, C. Greco, and F. Mannino. MatDSP: a DSP-based Matlab toolbox for rapid prototyping of digital control systems. In Fleming and Boullart [FB95a], pages 293–298. ISBN 0-08-042599-2. LCCN TJ217.7.I35 1995. Two volumes.

Casarin:2015:PSM

[CGRvD15] Roberto Casarin, Stefano Grassi, Francesco Ravazzolo, and Herman K. van Dijk. Parallel sequential Monte Carlo for efficient density combination: The DeCo MAT-LAB toolbox. Journal of Statistical Software, 68(3): ??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/index.php/

jss/article/view/v068i03;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v068i03/v68i03.pdf.

Casado:2020:AMN

[CH20] Jose Maria Varas Casado and Rob Hewson. Algorithm 1008: Multicomplex number class for Matlab, with a focus on the accurate calculation of small imaginary terms for multicomplex step sensitivity calculations. ACM Transactions on Mathematical Software, 46(2):18:1–18:26, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3378542.

Cacho:2023:MBI

[CH23] Oscar J. Cacho and Susan M. Hester. Modelling biocontrol of invasive insects using Wasp-Sim: a MATLAB simulation model. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000171.

Chang:1995:RBS

[Cha95] Dong Shang Chang. Reliability bounds for the stressstrength model. Computers & industrial engineering, 29(1-4):15–19, September 1995. CODEN CINDDL.

ISSN 0360-8352 (print), 1879-0550 (electronic).

Chapman:2000:MPE

[Cha00] Stephen J. Chapman. MAT-LAB programming for engineers. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-95151-1. xvii + 444 pp. LCCN QA297.C418 2000.

Chapman:2002:EMP

[Cha02a] Stephen J. Chapman. Electric machinery and power system fundamentals. McGraw-Hill series in electrical and computer engineering. McGraw-Hill, New York, NY, USA, 2002. ISBN 0-07-229135-4. xix + 673 pp. LCCN TK2000 .C455 2002. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/mh021/00066824.
                                        html;   http://www.loc.
                                        gov/catdir/toc/mh021/00066824.
                                        html.

Chapman:2002:MPE

[Cha02b] Stephen J. Chapman. MAT-LAB programming for engineers. BookWare companion series. Brooks/Cole, Pacific Grove, CA, USA, second edition, 2002. ISBN 0-534-39056-0. xviii + 478 pp. LCCN QA297 .C418 2002.

Chassaing:2002:DAU

[Cha02c] Rulph Chassaing. DSP applications using C and the TMS320C6x DSK. Topics in digital signal processing. Wiley, New York, NY, USA, 2002. ISBN 0-471-20754-3 (cloth). xix + 335 pp. LCCN TK5102.9 .C47423 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley044/2001046889.
html;   http://www.loc.
gov/catdir/description/
wiley036/2001046889.html;
 http://www.loc.gov/catdir/
toc/wiley021/2001046889.
html.

Chau:2002:PCF

[Cha02d] Pao C. Chau. Process control: a first course with MATLAB. Cambridge series in chemical engineering. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-80760-3, 0-521-00255-9 (paperback). xiii + 314 pp. LCCN TP155.75 .C42 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam022/2001052567.
html;   http://www.loc.
gov/catdir/samples/cam033/
2001052567.html;   http:
//www.loc.gov/catdir/toc/
cam025/2001052567.html.

Chapman:2004:MPE

[Cha04] Stephen J. Chapman. MAT-LAB programming for engineers. Thomson, London, UK, third edition, 2004. ISBN 0-534-42417-1. xviii + 540 pp. LCCN ????

Chapman:2005:EMF

[Cha05a] Stephen J. Chapman. Electric machinery fundamentals. McGraw-Hill Higher Education, New York, NY, USA, fourth edition, 2005. ISBN 0-07-246523-9. xxii + 746 pp. LCCN TK2000 .C46 2005. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/mh041/2003065174.
                                        html;   http://www.loc.
                                        gov/catdir/toc/mh041/2003065174.
                                        html.

Chapra:2005:ANM

[Cha05b] Steven C. Chapra. Applied numerical methods with MATLAB for engineers and scientists. McGraw-Hill Higher Education, Boston, MA, USA, 2005. ISBN 0-07-239265-7 , 0-07-297677-2 (set). xvi + 384 pp. LCCN QA297 .C4185 2005.

Chassaing:2005:DSP

[Cha05c] Rulph Chassaing. Digital signal processing and applications with the C6713 and C6416 DSK. Topics in digital signal processing. Wiley-Interscience, New York, NY, USA, 2005. ISBN 0-471-69007-4. xxi + 518 pp. LCCN TK5102.9 .C47422 2005. URL ftp://

                                        uiarchive.cso.uiuc.edu/

pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley042/2004050924.
html;   http://www.loc.
gov/catdir/toc/wiley041/
2004050924.html.

Chave:2017:CSE

[Cha17] Alan Dana Chave. Computational Statistics in the Earth Sciences: with Applications in MATLAB. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-09600-6 (hardcover), 1-108-51494-4, 1-316-15610-9 (e-book). xiii + 451 pp. LCCN QE26.3 .C43 2017.

Chen:1993:ADC

[Che93] Chi-Tsong Chen. Analog and Digital Control System Design: Transfer-Function, State-Space, and Algebraic Methods. Saunders College Publishing, Ft. Worth, TX, USA, 1993. ISBN 0-03-094070-2. xx + 600 pp. LCCN TJ213.C475 1993.

Chen:1994:SSA

[Che94] Chi-Tsong Chen. System and Signal Analysis. Saunders College Publishing, Ft. Worth, TX, USA, 1994. ISBN 0-03-097709-6. xx + 705 pp. LCCN QA402 .C4423 1994.

Chen:1999:LST

[Che99] Chi-Tsong Chen. Linear system theory and design.

The Oxford series in electrical and computer engineering. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, third edition, 1999. ISBN 0-19-511777-8 (cloth), 0-19-511778-6 (paper). xiii + 334 pp. LCCN QA402 .C44 1999.

Chen:2004:HNB

[Che04] Walter Y. Chen. Home networking basis: transmission environments and wired/wireless protocols. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-016511-5. xxii + 544 pp. LCCN TK5105.75 .C47 2003.

Chen:2008:SIB

[Che08] Long Chen. Short implementation of bisection in MATLAB. In Recent advances in computational sciences, pages 318–332. World Sci. Publ., Hackensack, NJ, 2008.

Chelikowsky:2018:IQM

[Che18] James R. Chelikowsky. Introductory Quantum Mechanics with MATLAB For Atoms, Molecules, Clusters, and Nanocrystals. Wiley-VCH, Weinheim, Germany, 2018. ISBN 3-527-40926-2 (paperback), 3-527-65501-8 (e-book), 3-527-65499-2 (Mobipocket ebook), 3-527-65500-X (ePub ebook), 3-527-65501-8 (PDF ebook). 240 pp. LCCN QC174.17.D37 .C445 2019. URL http://

swbplus.bsz-bw.de/bsz506870707cov.
htm;   http://www.wiley-
vch.de/publish/dt/books/
ISBN978-3-527-40926-6/.

Childers:1997:PRP

[Chi97] Donald G. Childers. Probability and random processes: using MATLAB with applications to continuous and discrete time systems. Irwin, Chicago, IL, USA, 1997. ISBN 0-256-13361-

1. xiii + 433 pp. LCCN QA273.19.E4C47 1997.

Childers:2000:SPS

[Chi00] Donald G. Childers. Speech processing and synthesis toolboxes. Wiley, New York, NY, USA, 2000. ISBN 0-471-34959-3 (cloth/CD-ROM). ix + 483 pp. LCCN TK7882.S65 C485 2000.

URL ftp://uiarchive.
cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.
loc.gov/catdir/description/
wiley032/99038270.html;
http://www.loc.gov/catdir/
toc/onix03/99038270.html.

Cash:2013:AMC

[CHMN13] J. R. Cash, D. Hollevoet, F. Mazzia, and A. M. Nagy. Algorithm 927: The MAT-LAB code bvptwp.m for the numerical solution of two point boundary value problems. ACM Transactions on Mathematical Software, 39(2):15:1–15:12, February 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Choy:2002:PMS

[Cho02] Ron Choy. Parallel Matlab survey. World-Wide Web document, 2002. URL

http://supertech.lcs.mit.
edu/~cly/survey.html#PTTO.

This is an extensive survey of “parallel functionalities in Matlab”, with links to many Internet sites.

Chriss:1997:BSBa

[Chr97] Neil A. Chriss. Black– Scholes and Beyond: Option Pricing Models. Irwin/McGraw Hill, Boston, MA, USA, 1997. ISBN 0-7863-1025-1. viii + 496 pp. LCCN HG6024.A3 C495 1997. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/mh022/96017361.
html;   http://www.loc.
gov/catdir/toc/mh022/96017361.
html.

Chapman:2015:MFC

[CHT15] S. Jonathan Chapman, David P. Hewett, and Lloyd N. Trefethen. Mathematics of the Faraday cage. SIAM Review, 57(3):398– 417, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chudzikiewicz:2000:SRV

[Chu00] Andrzej Chudzikiewicz. Simulation of rail vehicle dynamics in MATLAB environment. Vehicle system dynamics, 33(2):107– 119, February 2000. ISSN 0042-3114.

Craven:2001:COC

[CI01] B. D. Craven and S. M. N.

Islam. Computing optimal control on MATLAB— the SCOM package and economic growth models. In Optimization and related topics (Ballarat/Melbourne, 1999), volume 47 of Appl. Optim., pages 61–70. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 2001.

Cichocki:2002:ABS

[CiA02] Andrzej Cichocki and Shun ichi Amari. Adaptive blind signal and image processing: learning algorithms and applications. Wiley, New York, NY, USA, 2002. ISBN 0-471-60791-6. xxxi + 554 pp. LCCN TK5102.9 .C488 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley034/2001055943.
html;   http://www.loc.
gov/catdir/toc/wiley021/
2001055943.html.

Carstensen:2002:EFE

[CK02] C. Carstensen and R. Klose. Elastoviscoplastic finite element analysis in 100 lines of Matlab. J. Numer. Math., 10(3):157–192, 2002. ISSN 1570-2820.

Chand:2022:BPP

[CK22] Kulbhushan Chand and Arun Khosla. BioNES: a plug-and-play MATLABbased tool to use NES games for multimodal biofeedback. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S235271102200111X.

Chand:2023:CBP

[CK23] Kulbhushan Chand and Arun Khosla. Corrigendum to “BioNES: a plug-and-play MATLAB-based tool to use NES games for multimodal biofeedback” [SoftwareX 19 (2022) 101184]. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S235271102200214X.

See [CK22].

Chow:1994:AMG

[CKC94] C.-M. Chow, A. G. Kuznetsov, and D. W. Clarke. Application of multivariable generalized predictive control to the Simulink model of a paper machine. In IEEE [IEE94e], pages 1675– 1684. ISBN 0-7803-1873-0, 0-7803-1872-2, 0-7803-1874-9. LCCN TJ212.2.I3247 1994. Three volumes.

Chen:1996:UMC

[CL96] B. Chen and S. Lampe. The use of MATLAB as a computer tool for teaching control systems. In Iskander et al. [IGPF96], pages 181– 184. ISBN 1-56555-084-6. ISSN 0735-9276. LCCN T65.5.C65 I583 1996.

Clarke:1995:SMH

[Cla95] Robin T. Clarke. Statistical Modeling in Hydrology. Wiley, New York, NY, USA, 1995. ISBN 0-471-95016-5. xii + 412 pp. LCCN GB656.2.M33 C58 1994.

Clanton:2003:SSP

[Cla03] Sam Clanton. Speeding up the scientific process. Linux Journal, 110:56–60, June 2003. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Coleman:2020:MPJ

[CLMM20] Chase Coleman, Spencer Lyon, Lilia Maliar, and Serguei Maliar. Matlab, Python, Julia: What to choose in economics? Computational Economics, page ??, 2020. CODEN CNOMEL. URL http://

link.springer.com/article/

10.1007/s10614-020-09983-
3.

Cabrera:2020:MTE

[CLTS20] Pedro Cabrera, Henrik Lund, Jakob Zinck Thellufsen, and Peter Sorknæs. The MATLAB Toolbox for EnergyPLAN: a tool to extend energy planning studies. Science of Computer Programming, 191(??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642320300162.

Coveyou:1967:FAU

[CM67] R. R. Coveyou and R. D. MacPherson. Fourier analysis of uniform random number generators. Journal of the ACM, 14(1):100–119, January 1967. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Constantinides:1999:NMC

[CM99a] A. Constantinides and Navid Mostoufi. Numerical methods for chemical engineers with MATLAB applications. Prentice-Hall international series in the physical and chemical engineering sciences. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-013851-7. xxv + 560 pp. LCCN QA297 .C6494 1999.

Cooper:1999:PMS

[CM99b] George R. Cooper and Clare D. McGillem. Probabilistic methods of signal and system analysis. The Oxford series in electrical and computer engineering. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, third edition, 1999. ISBN 0-19-512354-9 (cloth), 0-19-512355-7 (Solutions manual). xiii + 480 pp. LCCN TK5102.9 .C59 1999; TK 454.2 .C664 1999X ENGI.

Chauhan:2003:ATD

[CMKH03] Arun Chauhan, Cheryl Mc-Cosh, Ken Kennedy, and Richard Hanson. Automatic type-driven library generation for telescoping languages. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

/www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10692#1;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap296.pdf.

Calvo:2017:ADM

[CMR17] Manuel Calvo, Juan I. Montijano, and Luis Rández. Algorithm 968: DISODE45: A Matlab Runge–Kutta solver for piecewise smooth IVPs of Filippov type. ACM Transactions on Mathematical Software, 43(3):25:1– 25:??, January 2017. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

                                        //dl.acm.org/citation.
                                        cfm?id=2907054.

Cribari-Neto:1997:MEP

[CNJ97] F. Cribari-Neto and M. J. Jensen. MATLAB as an econometric programming environment. Journal of Applied Econometrics, 12 (6):735–??, ???? 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coblenz:2021:MVC

[Cob21] Maximilian Coblenz. MATVines: a vine copula package for MATLAB. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711021000455.

Colgren:2004:ARC

[Col04] Richard D. (Richard Dean) Colgren. Applications of robust control to nonlinear systems, volume 205 of Progress in astronautics and aeronautics. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024–2518, 2004. ISBN 1-56347-666-5. xv + 175 pp. LCCN TL507 .P75 vol. 205.

Coleman:2005:IPD

[Col05] Matthew P. Coleman. An introduction to partial differential equations with MATLAB, volume 4 of Chapman and Hall/CRC applied mathematics and nonlinear science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-373-1. xii + 671 pp. LCCN QA371.35 .C66 2005.

Comincioli:1990:ANM

[Com90] Valeriano Comincioli. Analisi Numerica Metodi Modelli Applicazioni [English: Numerical Analysis: Methods, Models, and Applications]. McGraw-Hill, New York, NY, USA, 1990. ISBN 88-386-0646-3. LCCN ????

Comincioli:1992:MNS

[Com92] Valeriano Comincioli. Metodi Numerici e Statistici per le Scienze Applicate [English: Numerical and Statistical Methods for Applied Sciences]. C.E.A. Casa Editrice Ambrosiana, Milano, Italy, 1992. ISBN 88-408-0757-8. LCCN ????

Conrad:1995:SRR

[Con95a] William R. Conrad. Solving RL and RC electric circuits problems using MAT-LAB. In ASEE, editor, Investing in the future: 1995 Annual Conference proceedings, American Society for Engineering Education, June 25–28, 1995, Anaheim, California, pages 2241–2246. American Society for Engineering Education, Washington, DC, USA, 1995. ISBN ???? LCCN T62.A2 1995.

Conrad:1995:SRC

[Con95b] William R. Conrad. Solving RLC circuits using MAT-LAB version 4 using the symbolic toolbox. Computers in education journal, 5 (4):45–??, ???? 1995. CO-DEN CEJOE7. ISSN 1069-3769.

Conrad:1996:SRR

[Con96] William R. Conrad. Solving RL and RC circuits using MATLAB. Computers in education journal, 6(1): 12–??, ???? 1996. CODEN CEJOE7. ISSN 1069-3769.

Cook:1995:WTC

[Coo95] B. M. Cook, editor. World transputer congress: 3rd — September 1995, Harrogate, volume 46 of CONCUR-RENT SYSTEMS ENGI-NEERING SERIES 1995. IOS Press, Amsterdam, The Netherlands, 1995. ISBN 90-5199-235-1, 4-274-90062-2. ISSN 1383-7575. LCCN ????

Cooper:1996:MMB

[Coo96] R. M. Cooper. MatNEC: a MATLAB based graphical interface to SuperNEC. In Anonymous [Ano96t], pages 687–695 (or 687–693??). CODEN CPCEFK. ISBN ???? LCCN ????

Cooper:1998:IPD

[Coo98] Jeffery Cooper. Introduction to partial differential equations with MATLAB. Applied and Numerical Harmonic Analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 1998. ISBN 0-8176-3967-5, 3-7643-3967-

5. xv + 540 pp. LCCN QA371.35 .C66 1998.

Coombes:2000:DEM

[Coo00] Kevin Robert Coombes.

Differential equations with MATLAB. Wiley, New York, NY, USA, 2000. ISBN 0-471-32227-X. v + 256 pp. LCCN QA371.5.D37 D55 2000.

Cooper:2001:MCM

[Coo01] Jeffery Cooper. A MAT-LAB companion for multivariable calculus. Harcourt/ Academic Press, San Diego, CA, USA, 2001. ISBN 0-12-187625-X. xvi + 294 pp. LCCN QA303.5.D37 C66 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els031/00106079.
html;   http://www.loc.
gov/catdir/toc/els031/
00106079.html.

Cook:2014:AMM

[Coo14] Brandon G. Cook. Applied mathematics: Methods and Matlab. Computing in Science and Engineering, 16 (4):6–7, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Corke:1995:CTS

[Cor95] P. I. Corke. A computer tool for simulation and analysis: the robotics toolbox for MATLAB. In Anonymous [Ano95o], pages 319– 330. ISBN 0-9587583-0-1. LCCN ????

Corke:1996:RTM

[Cor96] Peter I. Corke. Robotics toolbox for MATLAB. IEEE Robotics & Automation Magazine, 3(1):24–32, March 1996. CODEN IRAMEB. ISSN 1070-9932.

Coulbeck:1993:ICA

[Cou93] Bryan Coulbeck, editor. Integrated computer applications in water supply: Proceedings of a Conference held at De Montfort University, Leicester, England, Sept. 7–8, 1993. Vol 1: Methods and procedures for systems simulation and control, number 6 in Mechanical Engineering Research Studies Engineering Control Series. Research Studies Press, Taunton, Somerset, England, 1993. ISBN 0-86380-154-4. LCCN TD353.I524 1993.

Couch:1995:MCS

[Cou95] Leon W. Couch II. Modern Communication Systems: Principles and Applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-02-325286-3. xxv + 598 pp. LCCN TK5101.A3C68 1995.

Couch:1997:DAC

[Cou97] Leon W. Couch, II. Digital and Analog Communication Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fifth edition, 1997. ISBN 0-13-522583-3. xxv + 742 pp. LCCN TK5101 .C69 1997. URL http://

www.eel.ufl.edu/~lcouc.

Couch:2001:DAC

[Cou01] Leon W. Couch. Digital and analog communication systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, sixth edition, 2001. ISBN 0-13-081223-4. xxv + 758 pp. LCCN TK5101 .C69 2001.

Cabezas:2020:PSM

[CPUARC20] Xavier Alejandro Flores Cabezas, Martha Cecilia Paredes Paredes, Luis Felipe Urquiza-Aguiar, and Diego Javier Reinoso-Chisaguano. PhySim-11p: Simulation model for IEEE 802.11p physical layer in MATLAB.

SoftwareX, 12(??):Article 100580, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711020302934.

Caliari:2013:GMC

[CR13] Marco Caliari and Stefan Rainer. GSGPEs: a MAT-LAB code for computing the ground state of systems of Gross–Pitaevskii equations. Computer Physics Communications, 184(3): 812–823, March 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465512003402.

Caliari:2020:GVM

[CR20] Marco Caliari and Stefan Rainer. GSGPEsv1.1: a MATLAB code for computing the ground state of systems of Gross– Pitaevskii equations. Computer Physics Communications, 247(??):Article 106968, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465519303170.

Crane:1996:GSM

[Cra96] C. M. Crane. Graphical studies in MATLAB. In Anonymous [Ano96u], pages 97–101. ISBN 0-201-87020-7. LCCN ????

Cutlip:1999:PSC

[CS99] Michael B. Cutlip and Mordechai Shacham. Problem solving in chemical engineering with numerical methods. Prentice-Hall international series in the physical and chemical engineering sciences. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-862566-2. xxxii + 458 pp. LCCN TP168 .C88 1999.

Su:2002:AF

[cS02] Kendall L. (Kendall Ling chiao) Su. Analog filters. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, second edition, 2002. ISBN 1-4020-7033-0. xv + 406 pp. LCCN TK7872.F5 S7997 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://site.ebrary.com/
lib/ucsc/Doc?id=10067273.

Casale:2012:KTF

[CS12] Giuliano Casale and Evgenia Smirni. KPC-toolbox: fitting Markovian arrival processes and phase-type distributions with MAT-LAB. ACM SIGMET-RICS Performance Evaluation Review, 39(4):47, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casoria:2003:HMM

[CSB03] Silvano Casoria, Gilbert Sybille, and Patrice Brunelle. Hysteresis modeling in the MATLAB/power system blockset. Mathematics and Computers in Simulation, 63(3–5):237–248, 2003. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Csenki:1992:STM

[Cse92]  Attila Csenki.  Sojourn

times in Markov processes for power transmission dependability assessment with MatLab. Microelectronics and Reliability, 32(7):945– 960, July 1992. CODEN MCRLAS. ISSN 0026-2714 (print), 1872-941X (electronic).

Csendes:1999:DRC

[Cse99] Tibor Csendes, editor. Developments in Reliable Computing: Papers presented at the International Symposium on Scientific Computing, Computer Arithmetic, and Validated Numerics, SCAN-98, in Szeged, Hungary, volume 5(3) of Reliable Computing = Nadezhnye vychisleniia. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1999. ISBN 0-7923-6057-5. LCCN QA76.9.E94 D48 1999.

Clark:1998:ASD

[CSG98] Robert L. Clark, William R.

Saunders, and Gary P. Gibbs. Adaptive structures: dynamics and control. Wiley, New York, NY, USA, 1998. ISBN 0-471-12262-9 (cloth). xvii + 467 pp. LCCN TA654.9 .C54 1998. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley042/97020308.
html;   http://www.loc.
gov/catdir/description/
wiley034/97020308.html;
http://www.loc.gov/catdir/
toc/onix05/97020308.html.

Campos:2020:NOC

[CST20] Carlos Campos, Cristiana J.

Silva, and Delfim F. M. Torres. Numerical optimal control of HIV transmission in Octave/MATLAB. Mathematical and Computational Applications, 25(1): ??, March 2020. CODEN ???? ISSN 2297-8747.

URL https://www.mdpi.
com/2297-8747/25/1/1.

Cavallo:1994:GOM

[CSV94] Alberto Cavallo, Roberto Setola, and Francesco Vasca. Guida Operativa a MAT-LAB, SIMULINK e Control Toolbox [English: Operative Guide to MATLAB, SIMULINK, and the Control Toolbox]. Liguori Editore, Napoli, Italy, 1994. ISBN 88-207-2474-X. ix + 502 pp. LCCN ????

Cavallo:1996:UMS

[CSV96] Alberto Cavallo, Roberto Setola, and Francesco Vasca. Using MATLAB, Simulink, and Control Toolbox: a practical approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-261058-2 (paperback). ???? pp. LCCN QA297.C3713 1996.

Cook:2015:PEM

[CSY15a] Dennis Cook, Zhihua Su, and Yi Yang. envlp: A MATLAB toolbox for computing envelope estimators in multivariate analysis. Journal of Statistical Software, 62(8):??, January 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v62/i08.

Cook:2015:PMT

[CSY15b] R. Dennis Cook, Zhihua Su, and Yi Yang. envlp: A MATLAB toolbox for computing envelope estimators in multivariate analysis. Journal of Statistical Software, 62(8):??, January 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v62/i08.

Chriss:1997:BSBb

[CT97] Neil A. Chriss and The MathWorks. The Black– Scholes and beyond interactive toolkit: a step-by-step guide to in-depth option pricing models. Irwin/McGraw Hill, Boston, MA, USA, 1997. ISBN 0-7863-1026-X. viii + 152 pp. LCCN HG6024.A3 C494 1997.

Cundiff:2003:DBS

[Cun03] John S. Cundiff. Dynamics of biological systems. American Society of Agricultural Engineers, St Joseph, MI, USA, 2003. ISBN 1-892769-30-1. ???? pp.

Curtis:1995:MAO

[Cur95] M. A. Curtis. A MATLAB algorithm for optimization of an arbitrary multivariate function. Data handling in science and technology, 15 (??):445–??, ???? 1995. CO-DEN DHSTEV. ISSN 0922-3487.

Curtis:2005:OME

[Cur05] Howard D. Curtis. Orbital mechanics for engineering students. Elsevier aerospace engineering series. Elsevier Butterworth-Heinemann, Oxford, UK, 2005. ISBN 0-7506-6169-0 (hardcover). xv + 673 pp. LCCN TL1050 .C87 2005.

Coleman:1988:HMC

[CV88] Thomas F. Coleman and Charles F. Van Loan. Handbook for Matrix Computations, volume 4 of Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1988. ISBN 0-89871-227-0 (paperback), 1-61197-104-7 (e-book). vii + 264 pp. LCCN QA188 .C651 1988. US$34.00.

Colinge:1996:PDS

[CV96] Jean-Pierre Colinge and Fernand Van de Wiele. Physique des Dispositifs Semi-conducteurs (French) (Physics of Semiconductor Devices). De Boeck Université, Bruxelles, Belgium, 1996. ISBN 2-8041-2107-0. 388 pp. LCCN ????

Coleman:2000:AAD

[CV00] Thomas F. Coleman and Arun Verma. ADMIT-1: automatic differentiation and MATLAB interface toolbox. ACM Transactions on Mathematical Software, 26(1):150– 175, March 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toms/2000-26-
1/p150-coleman/;   http:
//www.acm.org/pubs/citations/
journals/toms/2000-26-

1/p150-coleman/p150-coleman.
pdf.

Chonacky:2005:R

[CW05a] N. Chonacky and D. Winch.

3Ms: a response. Computing in Science and Engineering, 7(5):7–9, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//ieeexplore.ieee.org/
iel5/5992/32219/01501733.
pdf?isnumber=32219&prod=
JNL&arnumber=1501733&arSt=
+7&ared=+9&arAuthor=+Chonacky%
2C+N.%3B++Winch%2C+D.;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
32219&arnumber=1501733&
count=14&index=1.

Chonacky:2005:IPM

[CW05b] N. Chonacky and D. Winch.

3Ms for instruction, Part 2: Maple, Mathematica, and Matlab. Computing in Science and Engineering, 7(4):14–23, July/August 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//ieeexplore.ieee.org/
iel5/5992/31456/01463131.
pdf?isnumber=31456&prod=
JNL&arnumber=1463131&arSt=
+14&ared=+23&arAuthor=
Chonacky%2C+N.%3B+Winch%
2C+D.;   http://ieeexplore.
ieee.org/xpls/abs_all.
jsp?isnumber=31456&arnumber=

                                        1463131&count=14&index=
                                        1.

Chonacky:2005:IRM

[CW05c] Norman Chonacky and David Winch. 3Ms for instruction: Reviews of Maple, Mathematica, and Matlab. Computing in Science and Engineering, 7(3): 7–13, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        csdl.computer.org/comp/
                                        mags/cs/2005/03/c3007.
                                        pdf;   http://csdl.computer.
                                        org/comp/mags/cs/2005/
                                        03/c3007abs.htm.

Chonacky:2005:MMM

[CW05d] Norman Chonacky and David Winch. Maple, Mathematica, and Matlab: The 3M’s without the tape. Computing in Science and Engineering, 7(1):8–16, January/ February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/01/c1008.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1008.
                                        htm.

Chonacky:2005:RMM

[CW05e] Norman Chonacky and David Winch. Reviews of Maple, Mathematica, and Matlab: Coming soon to a publication near you. Computing in Science and Engineering, 7(2):9–10, March/ April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/02/c2009.pdf;
http://csdl.computer.org/
dl/mags/cs/2005/02/c2009.
htm.

Carroll:1998:DFS

[CWP98] John E. Carroll, James Whiteaway, and Dick Plumb. Distributed feedback semiconductor lasers, volume 10 of IEE Circuits, devices and systems series. IEE, London, UK, 1998. ISBN 0-85296-917-1 (IEE), 0-8194-2660-1 (SPIE). xxv + 412 pp. LCCN TA1700 .C367 1998.

Caliari:2017:IFE

[CZ17] Marco Caliari and Simone Zuccher. INFFTM: Fast evaluation of 3D Fourier series in MATLAB with an application to quantum vortex reconnections. Computer Physics Communications, 213(??):197– 207, April 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465516303691.

Czeczot:1995:EMS

[Cze95] J. Czeczot. Effective MAT-LAB simulation of distributed parameter systems basing on the matrix and vector notation. In Snorek et al. [SSV95], pages 143– 147. ISBN 1-56555-080-3. LCCN ????

DeMichele:1995:PIM

[D^{+}95] Dominick J. DeMichele et al., editors. Proceedings of the 13th International Modal Analysis Conference: February 13– 16, 1995, Sheraton Music City Hotel, Nashville, Tennessee, volume 2460 of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photo-optical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1995. ISBN 0-912053-48-8. ISSN 1046-6770. LCCN TS510.S63 v.2460; TA 654.15 I58 1995. Two volumes.

Dorato:1995:LCI

[DAC95] Peter Dorato, Chaouki Abdallah, and Vito Cercone. Linear-Quadratic Control: An Introduction. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-02-329962-2. xiv + 205 pp. LCCN TJ 216 D65 1995.

DAttellis:1995:IOP

[DACV95] Carlos E. D’Attellis, Marta T. Anaya, Maria I. Cavallaro, and Francisco F. Villaverde. Introduccion a las Onditas:

Una presentacion para Cursos de Grado de Ingenieria con MATLAB [English: Introduction to Wavelets: A presentation for graduate engineering courses with MATLAB]. Neuva Libreria, S.R.L., ????, 1995. ISBN 950-9088-77-3. ???? pp. LCCN ????

Dagli:1994:IES

[Dag94] Cihan H. Dagli, editor. Intelligent engineering systems through artificial neural networks: proceedings of the Artificial Neural Networks in Engineering (AN-NIE ’94) Conference, held November 13–16, 1994, in St. Louis, Missouri, USA, ASME Press Series on International Advances in Design Productivity. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1994. ISBN 0-7918-0045-8. LCCN ???? Three volumes.

Daku:2005:LMF

[Dak05] Brian Daku. Learn Matlab Fast: The M-tutor Interactive Tutorial. Wiley, New York, NY, USA, 2005. ISBN 0-471-72879-9. ???? pp. LCCN ????

Daku:2006:MTC

[Dak06] Brian Daku. MATLAB tutor CD: learning MAT-LAB superfast! Wiley, New York, NY, USA, 2006.

ISBN 0-471-72879-9 (CD-ROM), 0-471-27469-0 (paperback). xxi + 162 pp. LCCN ????

Datta:1995:NLA

[Dat95] Biswa Nath Datta. Numerical Linear Algebra and Applications. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-17466-3. xxii + 680 pp. LCCN QA184 .D37 1995.

Datta:2004:NML

[Dat04] Biswa Nath Datta. Numerical methods for linear control systems: design and analysis. Elsevier Academic Press, Amsterdam, The Netherlands, 2004. ISBN 0-12-203590-9. xxxviii + 695 pp. LCCN QA402.3 .D368 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/els041/2003058331.
                                        html;   http://www.loc.
                                        gov/catdir/toc/els041/
                                        2003058331.html.

Dattani:2013:FMP

[Dat13] Nikesh S. Dattani. Feyn-Dyn: a MATLAB program for fast numerical Feynman integral calculations for open quantum system dynamics on GPUs. Computer Physics Communications, 184(12):2828–2833, December 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465513002294.

Davis:1998:LCA

[Dav98] Artice M. Davis. Linear circuit analysis. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1998. ISBN 0-534-95095-7. xix + 1108 pp. LCCN TK454 .D414 1998.

Davis:1999:DEM

[Dav99] Paul W. Davis. Differential equations: modeling with MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-736539-X. xvi + 685 pp. LCCN QA371.5.D37 D38 1999.

Davis:2004:MAM

[Dav04a] Jon H. Davis. Methods of applied mathematics with a MATLAB overview. Applied and Numerical Harmonic Analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 2004. ISBN 0-8176-4331-1 , 3-7643-4331-1. xii + 721 pp. LCCN QA37.3 .D39 2004.

Davis:2004:AUV

[Dav04b] Timothy A. Davis. Algorithm 832: UMFPACK V4.3—an unsymmetricpattern multifrontal method. ACM Transactions on Mathematical Software, 30(2):

196–199, June 2004. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Davis:2005:ACS

[Dav05] Timothy A. Davis. Algorithm 849: a concise sparse Cholesky factorization package. ACM Transactions on Mathematical Software, 31(4):587–591, December 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Davis:2011:MP

[Dav11] Timothy A. Davis. MAT-LAB primer. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, eighth edition, 2011. ISBN 1-4398-2862-8. xvi + 232 pp. LCCN QA297 .D38 2011.

Davis:2012:LAP

[Dav12] Ernest Davis. Linear algebra and probability for computer science applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4665-0155-3 (hardcover). xviii + 413 pp. LCCN QA76.9.M35 D38 2012.

Davis:2013:AFO

[Dav13] Timothy A. Davis. Algorithm 930: FACTOR-IZE: an object-oriented linear system solver for MAT-LAB. ACM Transactions on Mathematical Software, 39(4):28:1–28:18, July 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Demuth:1993:NNT

[DB93a] Howard Demuth and Mark Beale. Neural Network Toolbox: For use with MATLAB: User’s Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1993. ISBN ???? LCCN ????

Dwan:1993:ISS

[DB93b] T. E. Dwan and T. E.

Bechert. Introducing Simulink into a systems engineering curriculum. In Grayson [Gra93], pages 627–631. ISBN 0-7803-1482-4, 0-7803-1483-2, 0-7803-1484-0. ISSN 0190-5848. LCCN ????

Dorf:1995:MCS

[DB95] Richard C. Dorf and Robert H.

Bishop. Modern Control Systems. Addison-Wesley, Reading, MA, USA, seventh edition, 1995. ISBN 0-201-50174-0. xix + 807 pp. LCCN TJ216 .D67 1994.

Dorf:1998:MCS

[DB98] Richard C. Dorf and Robert H.

Bishop. Modern control systems. Addison-Wesley, Reading, MA, USA, eighth edition, 1998. ISBN 0-201-30864-9. xxiv + 855 pp. LCCN TJ216.D67 1998.

Dorf:2004:MCS

[DB04] Richard C. Dorf and Robert H. Bishop. Modern control systems. Pearson Education, Upper Saddle River, NJ, USA, tenth edition, 2004. ISBN 0-13-145733-0 (hardcover). ???? pp. LCCN TJ216 .D67 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip0415/2004004466.
html.

Dahlquist:2008:NMS

[DB08] Germund Dahlquist and Åke Björck. Numerical Methods in Scientific Computing, volume 1. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2008. ISBN 0-89871-644-6. xxviii + 717 pp. LCCN QA297 .D335 2008. US$109.00. URL

http://www.ec-securehost.
com/SIAM/OT103.html;
http://www.loc.gov/catdir/
enhancements/fy0834/2007061806-
b.html;   http://www.loc.
gov/catdir/enhancements/
fy0834/2007061806-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0834/2007061806-
t.html.

Dakev:1995:GAS

[DCF95] N. V. Dakev, A. J. Chipperfield, and P. J. Fleming. General approach for solving optimal control problems using optimization techniques. Proceedings of the IEEE International Conference on Systems, Man and Cybernetics, 5:4503–4508, 1995. CO-DEN PICYE3. ISSN 0884-3627. IEEE catalog number 95CB35767.

Deidda:2020:FMP

[DdAF^{+}20] G. P. Deidda, P. Díaz de Alba, C. Fenu, G. Lovicu, and G. Rodriguez. FDEMtools: a MATLAB package for FDEM data inversion. Numerical Algorithms, 84(4):1313–1327, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11075-019-00843-
2.

Declercq:1997:NMB

[DDD97] F. Declercq, P. De Smedt, and R. De Keyser. A neural model based predictive control toolbox for MATLAB. In Boullart et al. [BLM97], pages 305–310. ISBN 0-08-042383-3. LCCN TJ212.2 .C33 1997.

Dien:2014:SNU

[DDK14] Hung Dien, Charlotte M.

Deane, and Bernhard Knapp. Software news and updates: Gro2mat: a package to efficiently read gromacs output in MATLAB. Journal of Computational Chemistry, 35(20):1528–1531, July 30, 2014. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Diniz:2002:DSP

[DDN02] Paulo Sergio Ramirez Diniz, Eduardo A. B. (Eduardo António Barros) Da Silva, and Sergio L. (Sergio Lima) Netto. Digital signal processing: system analysis and design. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-78175-2. xix + 612 pp. LCCN TK5102.9 .D63 2002. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/cam021/2001025447.
                                        html;   http://www.loc.
                                        gov/catdir/toc/cam027/
                                        2001025447.html.

Davis:1993:SEM

[DDW93] Joel Davis, Trevian Dray, and André Weideman. The student edition of MAT-LAB (review). Notices of the American Mathematical Society, 40(4):329–332, April 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https:/

                                        /www.ams.org/journals/
                                        notices/199304/199304FullIssue.
                                        pdf.

DeRose:1996:CTM

[De 96] L. De Rose. Compiler Techniques for MATLAB. Ph.D. thesis, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, 1996. ??–?? pp.

delaPenaEsteban:2005:TSP

[de 05] Francisco David de la Peña Esteban. Teoría de sistemas: problemas resueltos con Matlab. Vision Net, Madrid, Spain, 2005. ISBN 84-9821-030-5. ii + 109 pp. LCCN ????

Degallaix:2020:OMB

[Deg20] Jérôme Degallaix. OSCAR:

a MATLAB based package to simulate realistic optical cavities. SoftwareX, 12 (??):Article 100587, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711020303009.

DelCastillo:2002:SPA

[Del02] Enrique Del Castillo. Statistical process adjustment for quality control. Wiley series in probability and statistics. Wiley, New York, NY, USA, 2002. ISBN 0-471-43574-0 (cloth). xviii + 357 pp. LCCN TS156 .D388 2002; TS 156 .D388 2002X ENGI.

Demmel:1997:ANL

[Dem97] James W. Demmel. Applied Numerical Linear Algebra. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-389-7 (paperback). xi + 419 pp. LCCN QA184 .D455

1997. URL http://www.
siam.org/books/demmel/
demmel_class.

Denbigh:1998:SAS

[Den98] Philip Denbigh. System analysis and signal processing: with emphasis on the use of MATLAB. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-17860-5. xiv + 513 pp. LCCN TK5102.9.D46 1998.

Dufrechu:2013:ALL

[DEQOR13] Ernesto Dufrechu, Pablo Ezzatti, Enrique S. Quintana-Ortí, and Alfredo Remón. Accelerating the Lyapack library using GPUs. The Journal of Supercomputing, 65(3):1114–1124, September 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-013-0889-8.

Dirkse:1999:MSE

[DF99] Steven P. Dirkse and Michael C. Ferris. Modeling and solution environments for MPEC: GAMS & MATLAB. In Reformulation: nonsmooth, piecewise smooth, semismooth and smoothing methods (Lausanne, 1997), volume 22 of Appl. Optim., pages 127– 147. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1999.

DiPillo:1996:NOA

[DG96] G. Di Pillo and F. Giannessi, editors. Nonlinear optimization and applications: Lectures presented at the Workshop on Nonlinear Optimization: Theory and Applications, held in Erice at the “G. Stampacchia” School of Mathematics of the “E. Majorana” International Centre for Scientific Culture, June 13–21, 1995. Plenum Press, New York, NY, USA; London, UK, 1996. ISBN 0-306-45316-9. LCCN QA402.5.N645 1996.

DiBenedetto:2004:UUW

[DG04] Maria-Gabriella Di Benedetto and Guerino Giancola. Understanding ultra wide band radio fundamentals. Prentice Hall communications engineering and emerging technologies series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-148003-0. xx + 505 pp. LCCN TK5103.4 .D52 2004.

DeRose:1996:FMI

[DGGM96] L. De Rose, K. Gallivan, E. Gallopoulos, and B. Marsolf. FALCON: a MAT-LAB interactive restructuring compiler. In Huang et al. [H^{+}96], pages 269–288. CO-DEN LNCSD9. ISBN 3-540-60765-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1995.

Dhooge:2003:MMP

[DGK03] A. Dhooge, W. Govaerts, and Yu. A. Kuznetsov. MATCONT: A MATLAB package for numerical bifurcation analysis of ODEs. ACM Transactions on Mathematical Software, 29(2): 141–164, June 2003. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dhooge:2004:MMP

[DGK04] A. Dhooge, W. Govaerts, and Yu. A. Kuznetsov. MATCONT: a Matlab package for numerical bifurcation analysis of ODEs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):21–22, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

DeWitte:2012:IIC

[DGKF12] Virginie De Witte, Willy Govaerts, Yuri A. Kuznetsov, and Mark Friedman. Interactive initialization and continuation of homoclinic and heteroclinic orbits in MAT-LAB. ACM Transactions on Mathematical Software, 38 (3):18:1–18:34, April 2012. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Davis:2004:ACC

[DGLN04a] Timothy A. Davis, John R.

Gilbert, Stefan I. Larimore, and Esmond G. Ng. Algorithm 836: COLAMD, a column approximate minimum degree ordering algorithm. ACM Transactions on Mathematical Software, 30(3):377–380, September 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Davis:2004:CAM

[DGLN04b] Timothy A. Davis, John R.

Gilbert, Stefan I. Larimore, and Esmond G. Ng. A column approximate minimum degree ordering algorithm. ACM Transactions on Mathematical Software, 30(3):353–376, September 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

DAzzo:1995:LCS

[DH95] John D’Azzo and Constantine H. Houpis. Linear Control System Analysis and Design: Conventional and Modern. McGraw-Hill, New York, NY, USA, fourth edition, 1995. ISBN 0-07-016321-9. xviii + 763 pp. LCCN TJ 213 D33 1995.

Dyadkin:1997:FEL

[DH97a] Iosif G. Dyadkin and Kenneth G. Hamilton. A family of enhanced Lehmer random number generators, with hyperplane suppression, and direct support for certain physical applications. Computer Physics Communications, 107(1–3):258–280, December 22, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
cpc.cs.qub.ac.uk/cpc/;
 http://www.cpc.cs.qub.
ac.uk/cpc/cgi-bin/list_
summary.pl?CatNumber=ADGW.

Dyadkin:1997:SBM

[DH97b] Iosif G. Dyadkin and Kenneth G. Hamilton. A study of 64-bit multipliers for Lehmer pseudorandom number generators. Computer Physics Communications, 103(2–3):103–130, ???? 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Dabney:1998:MS

[DH98] James Dabney and Thomas L. Harman. Mastering SIMULINK 2. MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-243767-8. xvii + 345 pp. LCCN QA76.9.C65D32 1998.

Dyadkin:2000:SBM

[DH00] Iosif G. Dyadkin and Kenneth G. Hamilton. A study of 128-bit multipliers for congruential pseudorandom number generators. Computer Physics Communications, 125(1–3):239– 258, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://cpc.
cs.qub.ac.uk/summaries/
ADLK;   http://www.elsevier.
com/gej-ng//10/15/40/55/
25/42/abstract.html.

Dabney:2004:MS

[DH04] James Dabney and Thomas L.

Harman. Mastering Simulink. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-142477-7 (paperback). xix + 376 pp. LCCN QA76.9.C65 D34 2004.

Doherty:2012:MSA

[DH12a] Jesse Doherty and Laurie Hendren. McSAF: a static analysis framework for MATLAB. Lecture Notes in Computer Science, 7313: 132–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-31057-7_
7/.

Dubrau:2012:TM

[DH12b] Anton Willy Dubrau and Laurie Jane Hendren. Taming MATLAB. ACM SIG-PLAN Notices, 47(10):503– 522, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dhawan:2003:MIA

[Dha03] Atam P. Dhawan. Medical image analysis. IEEE Press series in biomedical engineering. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-471-45131-2. xv + 315 pp. LCCN RC78.7.D53 D48 2003.

Doherty:2011:KAM

[DHR11] Jesse Doherty, Laurie Hendren, and Soroush Radpour. Kind analysis for MATLAB. ACM SIGPLAN Notices, 46 (10):99–118, October 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Duda:2001:PC

[DHS01] Richard O. Duda, Peter E. (Peter Elliot) Hart, and David G. Stork. Pattern classification. Wiley, New York, NY, USA, second edition, 2001. ISBN 0-471-05669-3. xx + 654 pp. LCCN Q327 .D83 2001. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley041/99029981.
                                        html;   http://www.loc.

gov/catdir/description/
wiley032/99029981.html;
 http://www.loc.gov/catdir/
toc/onix03/99029981.html.

DAzzo:2003:LCS

[DHS03] John Joachim D’Azzo, Constantine H. Houpis, and Stuart N. Sheldon. Linear control system analysis and design with MATLAB, volume 14 of Control engineering series. Marcel Dekker, New York, NY, USA, fifth edition, 2003. ISBN 0-8247-4038-6. xvi + 839 pp. LCCN TJ213 .D33 2003.

Dimitrijev:2000:USD

[Dim00] Sima Dimitrijev. Understanding semiconductor devices. The Oxford series in electrical and computer engineering. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-513186-X (cloth). xviii + 574 pp. LCCN TK7871.85 .D547 2000.

Diniz:2002:AFA

[Din02] Paulo Sergio Ramirez Diniz. Adaptive filtering: algorithms and practical implementation, volume SECS 694 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, second edition, 2002. ISBN 1-4020-7125-6. xix + 568 pp. LCCN TK5102.9 .D6 2002; TK 5102.9 .D6 2002X ENGI.

Dillard:2006:SGA

[DIN06] William Dillard, J. David Irwin, and R. M. Nelms. Study guide to accompany Basic Engineering Circuit Analysis, eighth edition: with computer simulation techniques for Excel, MAT-LAB, and PSpice. Wiley, New York, NY, USA, 2006. ISBN 0-471-73106-4. 140 pp. LCCN TK454 .I78 2006.

Djaferis:1998:ACP

[Dja98] Theodore Euclid Djaferis. Automatic control: the power of feedback using MATLAB. The PWS BookWare companion series. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1998. ISBN 0-534-95051-5. xii + 144 pp. LCCN TJ213.D544 1998.

Djaferis:2000:ACP

[Dja00] Theodore Euclid Djaferis. Automatic control: the power of feedback using MATLAB. BookWare companion series. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37171-X (paperback). xii + 144 pp. LCCN TJ216 .D53 2000.

Drakenberg:1993:CCD

[DJK93] P. Drakenberg, P. Jacobsen, and B. Kåström. A CONLAB compiler for a distributed memory multicomputer. In Sincovec [Sin93], pages 814–821. ISBN 0-89871-315-3. LCCN QA76.58.S55 1993. Two volumes.

Danaila:2007:ISC

[DJKP07] Ionut Danaila, Pascal Joly, Sidi Mahmoud Kaber, and Marie Postel, editors. An introduction to scientific computing: twelve computational projects solved with MATLAB. Springer, New York, NY, USA, 2007. ISBN 0-387-30889-X (hardcover). xv + 294 pp. LCCN QA297 .I585

2007. URL http://www.
loc.gov/catdir/enhancements/
fy0824/2006931780-d.html;
http://www.loc.gov/catdir/
toc/fy0712/2006931780.
html.

Dunay:1999:MBA

[DK99] Rezső Dunay and István Kollár. MATLAB-based analysis of roundoff noise. In Csendes [Cse99], pages 373–382. ISBN 0-7923-6057-

5. LCCN QA76.9.E94 D48 1999.

DeCarlo:1995:LCA

[DL95] Raymond DeCarlo and Pen-Min Lin. Linear Circuit Analysis: Time Domain, Phasor, and Laplace Transform Approaches. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-473869-1 (two volumes), 0-13-043134-6 (vol. 1), 0-13-043142-7 (vol. 2). various pp. LCCN TK454.D44 1995.

Dumitras:1996:NNS

[DL96] A. Dumitras and V. Lazarescu. Neural network structures simulated with MatLab. A labwork software package for undergraduates. In Anonymous [Ano96e], pages 2269–2273. ISBN 3-89653-187-5. LCCN ????

DeCarlo:2001:LCA

[DL01] Raymond A. DeCarlo and Pen-Min Lin. Linear circuit analysis: time domain, phasor, and Laplace transform approaches. The Oxford series in electrical and computer engineering. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2001. ISBN 0-19-513666-7. xvi + 1008 pp. LCCN TK454 .D43 2001.

Defrance:2016:PMN

[DLA^{+}16] Josselin Defrance, Caroline Lemaître, Rabih Ajib, Jessica Benedicto, Emilien Mallet, Rémi Pollès, Jean-Pierre Plumey, Martine Mihailovic, Emmanuel Centeno, Cristian Ciracì, David Smith, and Antoine Moreau. Moosh: A numerical Swiss Army knife for the optics of multilayers in Octave/Matlab. Journal of Open Research Software, 4(1):e13–??, April 22, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.100/.

DeMoyer:1990:AMU

[DM90] Robert DeMoyer, Jr. and E. Eugene Mitchell, Jr. Applications of MATLAB in undergraduate automatic control system courses. CoED, 10(4):2–4, October– December 1990. CODEN CWLJDP. ISSN 0736-8607.

Diaz:1996:AEU

[DM96] Leo Diaz and Thomas A.

Milligan. Antenna engineering using physical optics: practical CAD techniques and software. Artech House antenna library. Artech House Inc., Boston, MA, USA, 1996. ISBN 0-89006-732-5. xiii + 343 pp. LCCN TK7871.6 .D48 1996.

Dolecek:2006:SSA

[DM06] G. Jovanovic Dolecek and S. K. Mitra. Symbolic sensitivity analysis of IIR digital filters using MAT-LAB. International Journal of Control, 79(11):1331– 1339, 2006. CODEN IJ-COAZ. ISSN 0020-7179.

Deckers:2017:AER

[DMB17] Karl Deckers, Ahlem Mougaida, and Hédi Belhadjsalah. Algorithm 973: Extended rational Fejér quadrature rules based on Chebyshev orthogonal rational functions. ACM Transactions on Mathematical Software, 43 (4):37:1–37:29, March 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

deMiras:2020:MTC

[dMMLOS20] Juan Ruiz de Miras, Guillermo Martínez-Lledó, William Orwig, and Jorge Sepulcre. A MATLAB tool for computing the spherical harmonic fractal dimension of the cerebral cortex. Computer Physics Communications, 254(??):Article 107381, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465520301648.

Do:1998:AAE

[Do98] Duong D. Do. Adsorption analysis: equilibria and kinetics, volume 2 of Series on chemical engineering. Imperial College Press, London, UK, 1998. ISBN 1-86094-130-3, 1-86094-137-0 (paperback). xxi + 892 pp.

Doebelin:1998:SDM

[Doe98] Ernest O. Doebelin. System dynamics: modeling, analysis, simulation, design. Marcel Dekker, New York, NY, USA, 1998. ISBN 0-8247-0126-7. xiv + 755 pp. LCCN TA168 .D588 1998.

Donnelly:1995:MMC

[Don95] Karen Donnelly. MAT-LAB manual: computer laboratory exercises. Saunders College Publishing, Ft. Worth, TX, USA, 1995. ISBN 0-03-094896-7. v + 196 + 16 pp. LCCN ????

Dorato:2000:AFS

[Dor00]  Peter Dorato.  Analytic

feedback system design: an interpolation approach. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-36917-0 (paperback). ix + 109 pp. LCCN TJ216 .D65 2000.

Ribeiro:2019:UTP

[dOSRS19] Pedro (Pedro Fernando de Oliveira Salazar) Ribeiro and Augusto Sampaio, editors. Unifying theories of programming: 7th International Symposium, UTP 2019, dedicated to Tony Hoare on the occasion of his 85th birthday, Porto, Portugal, October 8, 2019: proceedings, volume 11885 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2019. ISBN 3-030-31037-X (paperback), 3-030-31038-8 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .U86

                                        2019. URL http://link.
                                        springer.com/.

Dowd:2002:IMR

[Dow02a] Kevin Dowd. An introduction to market risk measurement. Wiley finance. Wiley, New York, NY, USA, 2002. ISBN 0-470-84748-4. xviii + 284 pp.

Dowd:2002:MMR

[Dow02b] Kevin Dowd. Measuring market risk. Wiley finance series. Wiley, New York, NY, USA, 2002. ISBN 0-471-52174-4. xix + 370 pp. LCCN HG6024.3 .D683 2002. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley046/2002071367.
                                        html;   http://www.loc.
                                        gov/catdir/description/
                                        wiley0310/2002071367.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley023/2002071367.
                                        html.

Daley:1995:ODC

[DP95] S. Daley and R. J. Patton. Observer design for control and fault detection of an induction motor. In Anonymous [Ano95b], pages 5–?? ISBN ???? LCCN ????

deCerqueira:1996:UDD

[dP96a] E. O. de Cerqueira and R. J. Poppi. Using dynamic data exchange to exchange information between Visual Basic and Matlab. application to a diode array spectrophotometer. Trends in analytical chemistry: TrAC, 15(10):500–??, ???? 1996. ISSN 0165-9936.

DeRose:1996:MFT

[DP96b] Luiz De Rose and David Padua. A MATLAB to Fortran 90 translator and its effectiveness. In ACM [ACM96], pages 309–316. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number: 415961.

DeRose:1999:TTM

[DP99] Luiz De Rose and David Padua. Techniques for the translation of MAT-LAB programs into Fortran 90. ACM Transactions on Programming Languages and Systems, 21(2):286– 323, March 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toplas/1999-21-
2/p286-de_rose/.

DiNatale:2008:BOM

[DP08] Marco Di Natale and Valerio Pappalardo. Buffer optimization in multitask implementations of Simulink models. ACM Transactions on Embedded Computing Systems, 7(3):23:1– 23:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Delgado:2016:APO

[DP16] Jorge Delgado and Juan Manuel Peña. Algorithm 960: POLYNOMIAL: an objectoriented Matlab library of fast and efficient algorithms for polynomials. ACM Transactions on Mathematical Software, 42(3):23:1– 23:19, May 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Demmig:1996:IES

[DPE96] S. Demmig, H. Pundt, and I. Erlich. Identification of external system equivalents for power system stability studies using the software package MATLAB. In Anonymous [Ano96s], pages 1055–1061. ISBN ???? LCCN ????

Dauxois:2005:FPU

[DPR05] Thierry Dauxois, Michel Peyrard, and Stefano Ruffo. The Fermi–Pasta–Ulam ‘numerical experiment’: history and pedagogical perspectives. European Journal of Physics, 26(5):S3–S11, 2005. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL

                                        https://arxiv.org/pdf/
                                        nlin/0501053.pdf.

Davenport:1996:NSL

[DR96] J. W. Davenport and L. F.

Roberts. Numeric and symbolic linear algebra applications using MATLAB. In Anonymous [Ano96u], pages 107–110. ISBN 0-201-87020-7. LCCN ????

Driscoll:1996:AMT

[Dri96] Tobin A. Driscoll. Algorithm 756: a MAT-LAB toolbox for Schwarz– Christoffel mapping. ACM Transactions on Mathematical Software, 22(2):168– 186, June 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:

//www.acm.org/pubs/toc/
Abstracts/0098-3500/229475.
html.

Driscoll:2005:AIS

[Dri05] Tobin A. Driscoll. Algorithm 843: Improvements to the Schwarz–Christoffel toolbox for MATLAB. ACM Transactions on Mathematical Software, 31(2):239– 251, June 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Driscoll:2009:LM

[Dri09] Tobin A. Driscoll. Learning MATLAB. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2009. ISBN 0-89871-683-7. xiv + 97 pp.

Doumbia:1997:MAS

[DRR97] M. L. Doumbia, G. Roy, and V. Rajagopalan. Modular approach for simulating electrical drive systems with Matlab/Simulink. In IEEE [IEE97b], pages TC3– 3–?? ISBN 0-7803-3946-0, 0-7803-3947-9. LCCN TK4058.I4 1997. IEEE catalog number: 97TH8282.

Dudzinski:2018:OOO

[DRS18] Michael Dudzinski, Marco Rozgić, and Marcus Stiemer. oFEM: An object oriented finite element package for Matlab. Applied Mathematics and Computation, 334(??):117–140, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031730824X.

Dorf:2004:IEC

[DS04] Richard C. Dorf and James A. Svoboda. Introduction to electric circuits. Wiley, New York, NY, USA, sixth edition, 2004. ISBN 0-471-44795-1, 0-471-45233-5 (WIE), 0-471-46344-2. xxii + 809 pp. LCCN TK454 .D67 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley039/2003270579.
html;   http://www.loc.

gov/catdir/toc/wiley032/
2003270579.html.

Davis:2005:MP

[DS05] Timothy A. Davis and Kermit Sigmon. MATLAB primer. Chapman and Hall/CRC, Boca Raton, FL, USA, seventh edition, 2005. ISBN 1-58488-523-8. xii + 215 pp. LCCN QA297 .D38 2005.

Dominguez:2009:ASM

[DS09] Víctor Domínguez and Francisco-Javier Sayas. Algorithm 884: a simple Matlab implementation of the Argyris element. ACM Transactions on Mathematical Software, 35(2):Art. 16, 11, 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Demski:2016:PEK

[DS16] Andrés Demski and Mariano Llamedo Soria. ecg-kit: a Matlab toolbox for cardiovascular signal processing. Journal of Open Research Software, 4(1): e8–??, April 04, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.86/.

Dutton:1997:ACE

[DTB97] Ken Dutton, Steve Thompson, and Bill Barraclough.

The Art of Control Engineering. Addison-Wes-ley, Reading, MA, USA, 1997. ISBN 0-201-17545-2 (paperback). xviii + 813 pp. LCCN TJ213 .D87 1997. URL ftp:/

/ftp.mathworks.com/pub/
books/dutton.

Duffy:2003:AEM

[Duf03] Dean G. Duffy. Advanced engineering mathematics with MATLAB. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2003. ISBN 1-58488-349-9. xvi + 818 pp. LCCN TA345 .D84 2003.

Duffy:2004:TMS

[Duf04] Dean G. Duffy. Transform methods for solving partial differential equations. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2004. ISBN 1-58488-451-7. xvii + 708 pp. LCCN QA403.5 .D84 2004.

Duffy:2010:AEM

[Duf10]  Dean G. Duffy.  Ad-

vanced engineering mathematics with MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2010. ISBN 1-4398-1624-7. ???? pp. LCCN TA345 .D84 2010.

Duffy:2011:AEM

[Duf11]  Dean G. Duffy.  Ad-

vanced engineering mathematics with MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2011. ISBN 1-4398-1624-7 (hardback). 1079 pp. LCCN TA345 .D84 2011.

Dunn:1999:GUI

[Dun99] Peter Dunn. A graphical user interface to generalized linear models in MAT-LAB. Journal of Statistical Software, 4(4):1–25, 1999. CODEN JSSOBK. ISSN

???? URL http://www.
jstatsoft.org/v04/i04;
http://www.jstatsoft.
org/v04/i04/guiglm.pdf;
http://www.jstatsoft.
org/v04/i04/updates.

Duong:1998:AAE

[Duo98] D. Do Duong. Adsorption analysis: equilibria and kinetics, volume 2 of Series on chemical engineering. Imperial College Press, London, UK, 1998. ISBN 1-86094-130-3, 1-86094-137-0 (paperback). xxi + 892 pp. LCCN TP156.A35 D86 1998.

Dutta:2016:MCC

[Dut16] Sourav Dutta. MATLAB creator Cleve Moler visits Texas A&M University. SIAM News, 49(6):??, July/ August 2016. URL https:/

/sinews.siam.org/DetailsPage/
TabId/900/ArtMID/2243/
ArticleID/783/MATLAB-Creator-

Cleve-Moler-Visits-Texas-
AM-University.aspx.

Royuela-del-Val:2017:LFP

[dVSWAL17] Javier Royuela del Val, Federico Simmross-Wattenberg, and Carlos Alberola-López. libstable: Fast, parallel, and high-precision computation of α-stable distributions in R, C/C++, and MATLAB. Journal of Statistical Software, 78 (??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v078i01;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v078i01/v78i01.pdf.

Davari:1996:DDF

[DZ96] Asad Davari and Yun Zhang. Dual-input describing function with Matlab. In Naraghi-Pour [NP96], pages 198–201. ISBN 0-8186-7352-4, 0-8186-7350-4 (??invalid ISBN checksum??). ISSN 0094-2898. LCCN TA 168 S727p 1996.

EnginYaz:1995:UMT

[EA95] Edwin Engin Yaz and Asad Azemi. Utilizing MATLAB in two graduate electrical engineering courses. Frontiers in Education Conference, 1:325–328, 1995. CO-DEN PFECDR. ISSN 0190-5848. IEEE catalog number 95CB35867.

Eliasmith:2003:NEC

[EA03] Chris Eliasmith and C. H.

(Charles H.) Anderson. Neural engineering: computation, representation, and dynamics in neurobiological systems. Computational neuroscience. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-05071-4 (hardcover). xii + 356 pp. LCCN QP363.3 .E454 2003.

El-Ali:2004:DSD

[EA04] Taan Said El-Ali. Discrete systems and digital signal processing with MAT-LAB. Electrical Engineering Textbook Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2004. ISBN 0-8493-1093-8. xviii + 667 pp. LCCN TK5102.9 .E35 2004.

El-Ali:2001:CSS

[EAK01] Taan S. El-Ali and Mohammad A. Karim. Continuous signals and systems with MATLAB. Electrical engineering textbook series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-0321-4. 532 (est.) pp. LCCN TK5102.9 .E34 2001.

Eaton:2014:GOR

[EaoGOBHW14] John W. (John Wesley)

Eaton, David (Co author of GNU Octave) Bateman, Søren Hauberg, and Rik Wehbring. The GNU Octave 3.8 reference manual. Samurai Media Limited, London, USA, third edition, 2014. ISBN 988-13277-3-3 (paperback: vol. 1), 988-13277-4-1 (paperback: vol. 2). xii + 890 (two volumes) pp. LCCN QA 76.95 .E1 G68 2014.

Eaton:1992:OHL

[Eat92] John W. Eaton. Octave: a high-level interactive language for numerical computations, February 1992.

URL ftp://ftp.gnu.org/

octave/. Originally developed at the University of Texas in Austin, and later contributed to the GNU Project of the Free Software Foundation. Octave is Matlab-like, but is not a full reimplementation.

Eaton:1997:GOH

[Eat97] John W. Eaton. GNU Octave: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, 1997. ISBN 0-9541617-2-6. viii + 311 pp. LCCN ????

Eaton:2000:GOH

[Eat00] John W. (John Wesley) Eaton. GNU Octave: a high-level interactive language for numerical computations: edition 3 for Octave version 2.0.13. Network Theory Ltd., Bristol, UK, 2000. ISBN 0-9541617-2-6 (paperback). viii + 311 pp. LCCN ????

Eaton:2002:GOH

[Eat02] John W. Eaton. GNU Octave: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, third edition, 2002. ISBN 0-9541617-2-6. viii + 311 pp. LCCN ???? For Octave version 2.0.13, February 1997.

Eaton:2005:GOH

[Eat05] John W. Eaton. GNU Octave: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, 2005. ISBN 0-9541617-2-6. viii + 311 pp. LCCN

???? URL http://www.
gbv.de/dms/ilmenau/toc/
520168526eaton.PDF. For

Octave version 2.0.17 (stable).

Enderle:2005:IBE

[EBB05] John D. (John Denis) Enderle, Joseph D. Bronzino, and Susan M. Blanchard, editors. Introduction to biomedical engineering. Academic Press series in biomedical engineering. Elsevier Academic Press, Amsterdam, The Netherlands, second edition, 2005. ISBN 0-12-238662-0. xxi + 1118 pp. LCCN R856 .I47 2005. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        toc/ecip055/2004030223.
                                        html.

Eaton:2008:GOM

[EBH08] John W. Eaton, David Bateman, and Søren Hauberg. GNU Octave manual: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, 2008. ISBN 0-9546120-6-X. x + 555 pp. LCCN ???? US$39.95. Version 3 for Octave version 3.0.2, August 2008.

Eddelbuttel:2000:EO

[Edd00] Dirk Eddelbüttel. Econometrics with Octave. Journal of Applied Econometrics, 15(5):531–542, September– October 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eddins:2009:AST

[Edd09] Steven L. Eddins. Automated software testing for Matlab. Computing in Science and Engineering, 11(6):48–55, November/December 2009. CO-DEN CSENFA. ISSN 0740-7475 (print), 1558-1918 (electronic).

Edlund:2004:CMS

[Edl04] Ove Edlund. CMregr — a Matlab software package for finding CM-estimates for regression. Journal of Statistical Software, 10 (3):1–11, 2004. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=92&url=v10/i03/
CMregr_rev2.pdf&ct=1.

Edwards:2009:ZPL

[Edw09] Kathryn Edwards. The A–Z of programming languages: MATLAB. ComputerWorld, December 9, 2009. CO-DEN CMPWAB. ISSN 0010-4841. URL http:/

/www.computerworld.com.
au/article/329191/a-z_
programming_languages_
matlab/.

Egan:1998:PLB

[Ega98] William F. Egan. Phase-lock basics. Wiley, New York, NY, USA, 1998. ISBN 0-471-24261-6 (cloth). xxx + 484 pp. LCCN TK7872.P38 E33 1998; TK 7872 .P38E33 1998X ENGI.

Egan:2000:FSP

[Ega00] William F. Egan. Frequency synthesis by phase lock. Wiley, New York, NY, USA, second edition, 2000. ISBN 0-471-32104-4. xxviii + 597 pp. LCCN TK7872.F73 E32 2000. URL ftp://

uiarchive.cso.uiuc.edu/

                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley047/99025863.
                                        html;   http://www.loc.
                                        gov/catdir/description/
                                        wiley031/99025863.html;
                                         http://www.loc.gov/catdir/
                                        toc/onix01/99025863.html.

Espinosa:1995:MSA

[EGE95] F. Espinosa, J. J. Garcia, and M. A. Espinosa. MAT-LAB: solucion al problema de identificacion del modelo de un sistema dinamico. Revista española de electronica, ??(486):60–??, ???? 1995. CODEN RVEEBT. ISSN 0482-6396.

Enright:2007:RRD

[EH07] W. H. Enright and Wayne B. Hayes. Robust and reliable defect control for Runge– Kutta methods. ACM Transactions on Mathematical Software, 33(1):1:1– 1:19, March 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Elsherbeni:2005:ADV

[EI05] Atef Z. Elsherbeni and Matthew J. Inman. Antenna Design and Visualization Using MATLAB: Version 1.2 — Program and User Guide on CD. Scitech Publishing Inc., Raleigh, NC, USA, 2005. ISBN 1-891121-39-1. ???? pp. LCCN ????

Egert:2023:RRI

[EK23] Janine Egert and Clemens Kreutz. Rcall: an R interface for MATLAB. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001947.

Etter:2002:IM

[EKH02] D. M. Etter, David C.

Kuncicky, and Douglas W. Hull. Introduction to Matlab 6. ESource–the Prentice Hall engineering source. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-032845-6. xviii + 142 pp. LCCN TA345 .E8725 2002.

Etter:2005:IM

[EKM05] D. M. Etter, David C. Kuncicky, and Holly Moore. Introduction to MATLAB 7. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-147492-8. xix + 283 pp. LCCN TA345 .E8726 2005 UCB.

Engblom:2016:FMC

[EL16] Stefan Engblom and Dimitar Lukarski. Fast Matlab compatible sparse assembly on multicore computers. Parallel Computing, 56(??): 1–17, August 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167819116300102.

ElAli:2004:DSD

[ElA04] Taan S. ElAli. Discrete systems and digital signal processing with MATLAB. Electrical Engineering Textbook Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2004. ISBN 0-8493-1093-8. xviii + 667 pp.

Elguibaly:2001:CCN

[Elg01]  Fayez Elguibaly.  Com-

puter communication networks: analysis and design. Part 1: ????; Part 2: Interconnection networks and switch design and modeling: Interconnection networks and switch design and modeling. NorthStar Digital Design, Victoria, BC, Canada, 2001. ISBN 0-9686116-2-1 (both volumes). xxvi + 404 + 405–621 pp.

Engelborghs:2002:NBA

[ELR02] K. Engelborghs, T. Luzyanina, and D. Roose. Numerical bifurcation analysis of delay differential equations using DDE-BIFTOOL. ACM Transactions on Mathematical Software, 28(1):1–21, March 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Esroy:1994:NAR

[EM94] Ysar Esroy and Alfredo O.

Moscardini, editors. Mathematical modelling courses for engineering education: Proceedings of the NATO Advanced Research Workshop on the Design of Mathematical Modelling Courses for Engineering Education, held in Izmir, Turkey, July 12–16, 1993, volume 132 of NATO ASI Series F Computer and Systems Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58010-7, 0-387-58010-7. ISSN 0258-1248. LCCN QA401 .M3943 1994.

Erway:2014:AMM

[EM14] Jennifer B. Erway and Roummel F. Marcia. Algorithm 943: MSS: MAT-LAB software for L-BFGS trust-region subproblems for large-scale optimization. ACM Transactions on Mathematical Software, 40 (4):28:1–28:12, June 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Etter:2004:IM

[EMK04] Delores M. Etter, Holly Moore, and David C. Kuncicky. Introduction to MAT-LAB 7. ESource–the Prentice Hall engineering source.

Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-147492-8. xix + 283 pp. LCCN ????

Engo:2001:DOO

[EMMK01] Kenth Engø, Arne Marthinsen, and Hans Z. Munthe-Kaas. DiffMan: an objectoriented MATLAB toolbox for solving differential equations on manifolds. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):323–347, 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Special issue: Themes in geometric integration.

Engelberg:2005:MIC

[Eng05]  Shlomo Engelberg.  A

mathematical introduction to control theory, volume 2 of Series in electrical and computer engineering. Imperial College Press, London, UK, 2005. ISBN 1-86094-570-8. xvii + 350 pp. LCCN QA402.3 .E644 2005.

Eddins:1994:UMC

[EO94] Steven L. Eddins and Michael T. Orchard. Using MATLAB and C in an image processing lab course. In IEEE [IEE94c], pages 515– 519. ISBN 0-8186-6952-7, 0-8186-6950-0, 0-8186-6951-9. LCCN TK8315.I22 1994. Three volumes.

EP:1997:FEA

[EP 97] EP Innovations, editor.

Frontiers in Education 1997: 27th Annual Conference: proceedings, November 5–8, 1997, Pittsburgh, PA: Teaching and learning in an era of change, volume 2 (or 1??). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4087-6, 0-7803-4086-8, 0-7803-4088-4, 0-7803-4089-2. ISSN 0190-5848. LCCN T62 .F76 1997. Three volumes.

Edwards:1994:CUM

[EP94a] C. Henry Edwards, Jr. and David E. Penney. Calculus using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-303256-6. ???? pp. LCCN QA 303 E223 1994.

Edwards:1994:CAG

[EP94b] C. Henry Edwards, Jr. and David E. Penney. Calculus with Analytic Geometry. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 1994. ISBN 0-13-457912-7, 0-13-463993-6. various pp. LCCN QA 303 E223 1994.

Edwards:1996:CPD

[EP96a] C. H. Edwards, Jr. and David E. Penney. Computing Projects for Differential Equations: Computing and Modeling. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-504465-0. 304 pp. LCCN

???? URL http://www.
prenhall.com/books/esm_
0-13-504465-0.html. Lab

supplement to [EP96b].

Edwards:1996:DEB

[EP96b] C. H. Edwards, Jr. and David E. Penney. Differential Equations and Boundary Value Problems: Computing and Modeling. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-382094-7. 718 pp. LCCN QA371.E28 1996. URL http://

www.prenhall.com/books/
esm_0-13-382094-7.html.

See also lab supplement [EP96a].

Edwards:1996:DEC

[EP96c] C. H. Edwards, Jr. and David E. Penney. Differential Equations: Computing and Modeling. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-382102-1. ???? pp. LCCN QA371.E282 1996.

Edwards:2000:DEB

[EP00a] C. H. (Charles Henry) Edwards and David E. Penney. Differential equations and boundary value problems: computing and modeling. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2000. ISBN 0-13-079770-7 (U.S.), 0-13-083444-0 (International ed.). xii + 781 pp. LCCN QA371 .E28 2000.

Edwards:2000:DEC

[EP00b] C. H. (Charles Henry) Edwards and David E. Penney. Differential equations: computing and modeling. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2000. ISBN 0-13-079779-0. xi + 545 + 3 pp. LCCN QA371 .E29 2000.

Edwards:2002:C

[EP02] C. H. (Charles Henry) Edwards and David E. Penney. Calculus. Prentice-Hall, Upper Saddle River, NJ 07458, USA, sixth edition, 2002. ISBN 0-13-092071-1, 0-13-095006-8 (International edition), 0-13-092082-7 (CD-ROM). ???? pp. LCCN QA303.2 .E42 2002.

Edwards:2003:DE

[EP03] C. H. (Charles Henry) Edwards and David E. Penney. Differential equations: computing and modeling. Pearson Education, Upper Saddle River, NJ, USA, third edition, 2003. ISBN 0-13-067337-4. various pp. LCCN QA371 .E29 2003.

Edwards:2004:DEC

[EP04a] C. H. (Charles Henry) Edwards and David E. Penney. Differential Equations and Boundary Value Problems: Computing and Modeling. Pearson Education, Upper Saddle River, NJ, USA, 2004. ISBN 0-13-065245-8. xv + 787 + 7 pp. LCCN QA371 .E28 2004.

Edwards:2004:EDE

[EP04b] C. H. (Charles Henry) Edwards and David E. Penney. Elementary differential equations. Pearson Education, Upper Saddle River, NJ, USA, fifth edition, 2004. ISBN 0-13-145773-X. xi + 613 + 6 pp. LCCN QA371 .E33 2004b.

Eres:2005:IUG

[EPJ^{+}05] M. Hakki Eres, Graeme E. Pound, Zhouan Jiao, Jasmin L. Wason, Fenglian Xu, Andy J. Keane, and Simon J. Cox. Implementation and utilisation of a Grid-enabled problem solving environment in Matlab. Future Generation Computer Systems, 21(6):920– 929, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Epps:2011:BRE

[Epp11] Bob Epps. Book review: The elements of MATLAB style by Richard K. Johnson. ACM SIGSOFT Software Engineering Notes, 36 (3):33–34, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Elman:2007:AIM

[ERS07] Howard C. Elman, Alison Ramage, and David J. Silvester. Algorithm 866: IFISS, a Matlab toolbox for modelling incompressible flow. ACM Transactions on Mathematical Software, 33(2):1–18, June 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

El-Sharkawi:2000:FED

[ES00] Mohamed A. El-Sharkawi.

Fundamentals of electric drives. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-95222-4. xii + 314 pp. LCCN TK4058 .E39 2000.

Erocal:2010:SPU

[ES10] Burçin Eröcal and William Stein. The Sage Project: Unifying free mathematical software to create a viable alternative to Magma, Maple, Mathematica and MATLAB. Lecture Notes in Computer Science, 6327: 12–27, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.

                                        springer.com/content/pdf/
                                        10.1007/978-3-642-15582-
                                        6_4.pdf.

Esfandiari:2003:AME

[Esf03] Ramin S. Esfandiari. Applied mathematics for engineers. Atlantis Publishers, Irvine, CA, USA, third edition, 2003. ISBN 0-9729990-0-0. 959 (est.) pp. LCCN *.

Esmailzadeh:2014:CSM

[Esm14] Nabaz Esmailzadeh. Code snippet: A MATLAB package for computing two-level search design performance criteria. Journal of Statistical Software, 56(CS-1): ??, January 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v56/c01.

Espelid:2007:AGD

[Esp07] Terje O. Espelid. Algorithm 868: Globally doubly adaptive quadrature— reliable Matlab codes. ACM Transactions on Mathematical Software, 33(3):21:1– 21:21, August 2007. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Etter:1993:EPS

[Ett93] D. M. Etter. Engineering Problem Solving with MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-280470-0 (includes computer disk). xxi + 434 pp. LCCN TA331 .E88 1993.

Etter:1996:IME

[Ett96] D. M. Etter. Introduction to MATLAB for Engineers and Scientists. Prentice Hall modular series for engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-519703-1. xi + 145 pp. LCCN TA345.E873 1996.

Etter:1997:EPS

[Ett97] Delores M. Etter. Engineering problem solving with MATLAB. MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1997. ISBN 0-13-397688-2. xx + 329 pp. LCCN TA345.E547 1997.

Etter:2010:IM

[Ett10] Delores M. Etter. Introduction to MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2010. ISBN 0-13-608123-1. ???? pp. LCCN TA345 .E8724 2010.

Elnashaie:2007:NTC

[EU07] Said Elnashaie and Frank Uhlig. Numerical techniques for chemical and biological engineers using MAT-LAB(R). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-34433-0. xvi + 590 pp. A simple bifurcation approach, With the assistance of Chadia Affane, With 1 CD-ROM (Windows, Macintosh and UNIX/Linux).

Faber:1995:MSP

[Fab95] Erica Faber. MATLAB 4.1 with Signal Processing Toolbox for Macintosh computers. Computer Music Journal, 19(2):112–??, ???? 1995. CODEN CMUJDY. ISSN 0148-9267 (print), 1531-5169 (electronic).

Faber:1997:SGE

[Fab97] K. Faber. On solving generalized eigenvalue problems using MATLAB. Journal of Chemometrics, 11(1):87– ??, ???? 1997. CODEN JOCHEU. ISSN 0886-9383. Fasshauer:2007:MAM

[Fas07] Gregory E. Fasshauer. Meshfree approximation methods with MATLAB, volume 6 of Interdisciplinary Mathematical Sciences. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2007. ISBN 981-270-634-8. xviii + 500 pp. With 1 CD-ROM (Windows, Macintosh and UNIX).

Fausett:1999:ANA

[Fau99] Laurene V. Fausett. Applied numerical analysis using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-319849-9. xii + 596 pp. LCCN QA297 .F38 1999.

Fausett:2008:ANA

[Fau08] Laurene V. Fausett. Applied Numerical Analysis using MATLAB. Pearson Prentice Hall, Upper Saddle River, NJ, USA, second edition, 2008. ISBN 0-13-239728-5. xiii + 673 pp. LCCN QA297 .F38 2008.

Fayek:2017:PML

[Fay17] Haytham M. Fayek. MatDL:

A lightweight deep learning library in MATLAB. Journal of Open Source Software, 2(19):413:1–413:2, November 2017. CO-DEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/
10.21105/joss.00413.

Fazio:2010:IFM

[Faz10] R. Fazio. An instance of failure for the MATLAB explicit ODE45 solver. In S. I. Ao, L. Gelman, D. W. L. Hukins, A. Hunter, and A. M. Korsunsky, editors, Proceedings World Congress on Engineering 2010, held in London, 30 June–2 July, volume III, pages 1935– 1940. IAENG, Kwun Tong, Hong Kong, 2010.

Fleming:1995:AAR

[FB95a] P. J. Fleming and L. Boullart, editors. Algorithms and architectures for real-time control, AARTC ’95: a postprint volume from the 3rd IFAC/IFIP workshop, Ostend, Belgium, 31 May– 2 June 1995, volume 1 of Algorithms and Architectures for Real Time Control. Pergamon Press, New York, NY, USA, 1995. ISBN 0-08-042599-2. LCCN TJ217.7.I35 1995. Two volumes.

Fonseca:1995:SSC

[FB95b] I. M. Fonseca and P. M. Bainum. Simultaneous structural and control optimization of large space structures. Applied mechanics reviews, 48(11/2): S175–S180, November 1995. CODEN AMREAD. ISSN 0003-6900.

Fraleigh:1995:LA

[FB95c] John B. Fraleigh and Raymond A. Beauregard. Linear Algebra. Addison-Wes-ley, Reading, MA, USA, third edition, 1995. ISBN 0-201-52675-1. xii + 538 + 57 pp. LCCN QA184.F73 1995.

Farhang-Boroujeny:1998:AFT

[FB98] B. Farhang-Boroujeny. Adaptive filters: theory and applications. Wiley, New York, NY, USA, 1998. ISBN 0-471-98337-3. xv + 529 pp. LCCN TK7872.F5 F37 1998. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;

http://www.loc.gov/catdir/
description/wiley0310/
98008783.html;   http:/
/www.loc.gov/catdir/toc/
onix03/98008783.html.

Faires:2003:NM

[FB03] J. Douglas Faires and Richard L. Burden. Numerical methods. Brooks/Cole– Thomson Learning, Pacific Grove, CA, USA, third edition, 2003. ISBN 0-534-40761-7. xii + 622 pp. LCCN QA297 .F35 2003.

Figliola:2000:TDM

[FBC00] R. S. Figliola, Donald E.

Beasley, and Timothy A. Conover. Theory and design for mechanical measurements. Wiley, New York, NY, USA, third edition, 2000. ISBN 0-471-35083-4. xiv + 536 pp. LCCN T50 .F54 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/onix05/00036794.html.

Farenzena:2004:RAB

[FBFB04] M. Farenzena, A. Busti, A. Fusiello, and A. Benedetti. Rigorous accuracy bounds for calibrated stereo reconstruction. In Kittler et al. [KPN^{+}04], pages 288–292. ISBN ???? ISSN 1051-4651. LCCN ???? IEEE Computer Society Order Number P2128.

Finsterwalder:1994:GME

[FBG94] R. Finsterwalder, E. G. Berger, and G. Grübel. A graphical matrix editor to be used in ANDECS/ MATLAB computation environments. In Mattsson et al. [MGC94], pages 165– 170. ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Foley-Bourgon:2017:EIC

[FBH17] Vincent Foley-Bourgon and Laurie Hendren. Efficiently implementing the copy semantics of MATLAB’s arrays in JavaScript. ACM SIGPLAN Notices, 52(2): 72–83, February 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frederick:1995:FCP

[FC95] Dean K. Frederick and Joe H. Chow. Feedback Control Problems Using MAT-LAB and the Control System Toolbox. Tom Robbins’ BookWare companion series. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1995. ISBN 0-534-93798-5. xx + 219 pp. LCCN TJ216.F73 1995. US$30.95.

Frederick:2000:FCP

[FC00] Dean K. Frederick and J. H.

(Joe H.) Chow. Feedback control problems: using MATLAB and the Control System Toolbox. BookWare companion series. Brooks/ Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37172-8. xxi + 225 pp. LCCN TJ216 .F73 2000.

Florin:1997:SEC

[FCP97] I. Florin, G. Constantin, and V. Platagea. Some examples of comparative simulation of nonlinear solid body systems with MAT-LAB/SIMULINK and SDS simulation programs. In Lakshmikantham [Lak97], pages 1969–1976. ISBN 0-08-042032-X, 0-08-042036-2. ISSN 0362-546X (print), 1873-5215 (electronic). LCCN ????

Fehribach:2001:SFA

[FD01] Joseph D. Fehribach and Anthony M. J. Davis. Stokes flow around an asymmetric channel divider; a computational approach using MAT-LAB. J. Engrg. Math., 39 (1-4):207–220, 2001. CO-DEN JLEMAU. ISSN 0022-0833. Special issue on practical asymptotics.

Formica:2024:SBT

[FFR^{+}24] Federico Formica, Tony Fan, Akshay Rajhans, Vera Pantelic, Mark Lawford, and Claudio Menghi. Simulation-based testing of Simulink models with test sequence and test assessment blocks. IEEE Transactions on Software Engineering, 50(2):239–257, February 2024. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Fernandez-Gaucherand:1994:SSS

[FGCG94] Emmanuel Fernandez-Gaucherand, Jongsup Choi, and Dan Gerhart. S^{2}YSCODE: stochastic systems control and decision algorithms software laboratory, FORTRAN & MATLAB versions. In Mattsson et al. [MGC94], pages 179–186. ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Fu:2015:AMT

[FGjS15] Zhixing Fu, Luis F. Gatica, and Francisco javier Sayas. Algorithm 949: MATLAB tools for HDG in three dimensions. ACM Transactions on Mathematical Software, 41(3):20:1– 20:21, June 2015. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Falcone:2021:SBS

[FGMS21] Alberto Falcone, Alfredo Garro, Marat S. Mukhametzhanov, and Yaroslav D. Sergeyev. A Simulinkbased software solution using the Infinity Computer methodology for higher order differentiation. Applied Mathematics and Computation, 409 (??):Article 125606, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305610.

Fierro:1999:UTM

[FHH99] Ricardo D. Fierro, Per Christian Hansen, and Peter Søren Kirk Hansen. UTV tools: Matlab templates for rank-revealing UTV decompositions. Numerical Algorithms, 20(2-3):165–194, 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Forth:2012:RAA

[FHP^{+}12] Shaun Forth, Paul Hovland, Eric Phipps, Jean Utke, and Andrea Walther, editors. Recent Advances in Algorithmic Differentiation, volume 87 of Lecture Notes in Computational Science and Engineering. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CO-DEN LNCSA6. ISBN 3-642-30022-7 (print), 3-642-30023-5 (e-book). ISSN 1439-7358. LCCN ????

URL http://link.springer.
com/book/10.1007/978-3-
642-30023-3;   http://
www.springerlink.com/content/
978-3-642-30023-3. Pro-

ceedings of the Sixth International Conference on Automatic Differentiation (AD2012) held July 23–27, 2012, in Fort Collins, Colorado, USA.

Fisher:1995:DSH

[Fis95] Brice Fisher. Dynamic simulation of the human respiratory system using MAT-LAB. Thesis (M.S.), College of Engineering. Department of Engineering Science. University of Toledo., Toledo, OH, USA, 1995. vi + 65 pp.

Fisher:2019:ICM

[Fis19] Aidan A. E. Fisher. An introduction to coding with Matlab: Simulation of Xray photoelectron spectroscopy by employing Slater’s rules. Journal of Chemical Education, 96 (7):1502–1505, June 2019. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Farmer:2022:MTP

[FJ22] Jenny Farmer and Donald J. Jacobs. MATLAB tool for probability density assessment and nonparametric estimation. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022000231.

Fosdick:1996:IHP

[FJSD96] Lloyd D. Fosdick, Elizabeth R. Jessup, Carolyn J. C. Schauble, and Gitta Domik. An Introduction to High-Performance Scientific Computing. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-06181-3 (hardcover). xxiii + 760 pp. LCCN QA76 .A594 1996. US$70.00.

Frauendiener:2011:ACR

[FK11] Jörg Frauendiener and Christian Klein. Algebraic curves and Riemann surfaces in Matlab. Lecture Notes in Mathematics, 2013: 125–162, 2011. CODEN LN-MAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-17413-1_
3/.

Frantsuz:1997:DAS

[FKSM97] E. Frantsuz, V. Khavinson, F. Shiota, and Y. Miki. Designing axially symmetric electromechanical systems of superconducting magnetic levitation in Matlab environment. In Ciarlini et al. [C^{+}97], pages 254–256. ISBN 981-02-2918-6. LCCN QA465.A283 1997.

Fitzgerald:2003:EM

[FKU03] A. E. (Arthur Eugene)

Fitzgerald, Charles Kingsley, and Stephen D. Umans. Electric machinery. McGraw-Hill series in electrical engineering. Power and energy. McGraw-Hill, New York, NY, USA, sixth edition, 2003. ISBN 0-07-366009-4, 0-07-112193-5. xv + 688 pp. LCCN TK2181 .F5 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/mh024/2002070988.
html;   http://www.loc.
gov/catdir/toc/mh023/2002070988.
html.

Freeborn:2024:UDR

[FM24] Todd J. Freeborn and Ja-

cob A. Mota. Ultrasound
DICOM Renamer: a MAT-

LAB graphical user interface for workflow improvement for DICOM ultrasound renaming. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711024001146.

Frittelli:2025:VVE

[FMS25] Massimo Frittelli, Anotida Madzvamuse, and Ivonne Sgura. VEMcomp: a virtual elements MATLAB package for bulk-surface PDEs in 2D and 3D. Numerical Algorithms, 99(3):1365– 1392, July 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

https://link.springer.
com/article/10.1007/s11075-
024-01919-4.

Ferris:2007:LPM

[FMW07] Michael C. Ferris, Olvi L.

Mangasarian, and Stephen J. Wright. Linear programming with Matlab, volume 7 of MPS/SIAM Series on Optimization. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2007. ISBN 0-89871-643-8. xii + 266 pp. LCCN QA402.5 .F425 2007.

Fogler:1999:ECR

[Fog99] H. Scott Fogler. Elements of chemical reaction engineering. Prentice-Hall international series in the physical and chemical engineering sciences. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 1999. ISBN 0-13-531708-8 (cloth). xxx + 967 pp. LCCN TP157 .F65 1999.

Forth:2006:EOI

[For06] Shaun A. Forth. An efficient overloaded implementation of forward mode automatic differentiation in MATLAB. ACM Transactions on Mathematical Software, 32(2):195–222, June 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Foster:1991:PM

[Fos91] Kenneth R. Foster. Prepackaged math. IEEE Spectrum, 28(11):44–50, November 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1999:MI

[Fos99] Kenneth R. Foster. Math on the Internet. IEEE Spectrum, 36(4):36–40, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2001:MME

[Fos01] Kenneth R. Foster. Matlab made easier and testing with Matlab [software resources]. IEEE Spectrum, 38(2):87– 89, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fladung:1998:IVD

[FPBO98] W. A. Fladung, A. W. Phillips, D. L. Brown, and N. Olsen. The integration of VXI data acquisition into MATLAB. In Wicks and DeMichele [WD98], pages 1089–1093. ISBN 0-912053-59-3. ISSN 1046-6770. LCCN TA654.15 .I57 1998. Two volumes.

Franklin:1994:FCD

[FPEN94] Gene F. Franklin, J. David Powell, and Abbas Emami-Naeini. Feedback Control of Dynamic Systems. Addison-Wesley, Reading, MA, USA, third edition, 1994. ISBN 0-201-52747-2. xx + 778 pp. LCCN TJ216 .F723 1994.

Franklin:2002:FCD

[FPEN02] Gene F. Franklin, J. David Powell, and Abbas Emami-Naeini. Feedback control of dynamic systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2002. ISBN 0-13-032393-4. xvii + 910 pp. LCCN TJ216 .F723 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy032/2002279119.html.

Franklin:1990:DCD

[FPW90] Gene F. Franklin, J. David Powell, and Michael L. Workman. Digital Control of Dynamic Systems. Addison-Wesley, Reading, MA, USA, second edition, 1990. ISBN 0-201-11938-2. xxi + 837 pp. LCCN TJ223.M53 F73 1990.

Franklin:1998:DCD

[FPW98] Gene F. Franklin, J. David Powell, and Michael L. Workman. Digital control of dynamic systems. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-82054-4. xxiii + 742 pp. LCCN TJ223.M53 F73 1998.

Freeman:1995:FDE

[FR95] James J. Freeman and Sal Rositano. Freshman design & engineering tools course. Frontiers in Education Conference, 1:33–35, 1995. CO-DEN PFECDR. ISSN 0190-5848. IEEE catalog number 95CB35867.

Ferris:1996:ARM

[FR96] M. C. Ferris and T. F. Rutherford. Accessing realistic mixed complementarity problems within MAT-LAB. In Di Pillo and Giannessi [DG96], pages 141– 154. ISBN 0-306-45316-9. LCCN QA402.5.N645 1996.

Fachada:2018:PMM

[FR18] Nuno Fachada and Agostinho C.

                                        Rosa.  micompm:  A

MATLAB/Octave toolbox for multivariate independent comparison of observations. Journal of Open Source Software, 3 (23):430:1, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00430.

Frank:2015:FMG

[FRAK15] Florian Frank, Balthasar Reuter, Vadym Aizinger, and Peter Knabner. FES-TUNG: A MATLAB/GNU Octave toolbox for the discontinuous Galerkin method, Part I: Diffusion operator.

Computers and Mathematics with Applications, 70(1): 11–46, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122115001820.

Feng:2017:GSR

[FS17] Mingbin Feng and Jeremy Staum. Green simulation: Reusing the output of repeated experiments. ACM Transactions on Modeling and Computer Simulation, 27(4):23:1–23:??, December 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See replication report [Nel17].

Filippidis:2023:MAA

[FS23] Achilleas Filippidis and Adam J. Sadowski. Modern analysis of axisymmetric shells with AQUINAS: a MATLAB finite element toolbox. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711023001309.

Fontan:1995:MWS

[FSC95] F. P. Fontan, A. Seoane, and M. V. Castro. Matlab for Windows software aid in a mobile communications course. International Journal of Electrical Engineering Education, 32(4):341– 349, October 1995. CODEN IJEEAF. ISSN 0020-7209.

Flinders:1993:PES

[FSO93] F. Flinders, S. Senini, and W. Oghanna. Power electronics simulation laboratory using “Simulink” dynamic systems analysis package. In Olivier and Bouchard [OB93], pages 643–650. ISBN ???? LCCN ????

Frison:2020:BAB

[FSZD20] Gianluca Frison, Tommaso Sartor, Andrea Zanelli, and Moritz Diehl. The BLAS API of BLASFEO: Optimizing performance for small matrices. ACM Transactions on Mathematical Software, 46(2):15:1–15:36, June 2020. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3378671.

Frontini:1992:CNN

[FT92] Marco Frontini and Aldo Tagliani. Calcolo Numerico— Esercizi [English: Numerical Calculations—Exercises]. clup CittàStudi, Milano, Italy, 1992. ISBN 88-251-7056-4. LCCN ????

Farfour:2018:SUG

[FY18] Mohammed Farfour and Wang Jung Yoon. Seismic UNIX and GNU Octave for VSP Data Processing and Interpretation, pages 73–92. Wiley, New York, NY, USA, 2018. ISBN 1-119-22751-8.

Grewal:2001:KFT

[GA01] Mohinder S. Grewal and Angus P. Andrews. Kalman filtering: theory and practice using MATLAB. Wiley, New York, NY, USA, second edition, 2001. ISBN 0-471-39254-5. 401 pp. LCCN QA402.3 .G697 2001.

Gaberson:1998:DAM

[Gab98] H. A. Gaberson. Determination and animation of machinery rigid body operating deflection shapes with Matlab. In Wicks and DeMichele [WD98], pages 1526–1532. ISBN 0-912053-59-3. ISSN 1046-6770. LCCN TA654.15 .I57 1998. Two volumes.

Gajic:2003:LDS

[Gaj03] Zoran Gajic. Linear dynamic systems and signals. Prentice Hall / Pearson Education, Upper Saddle River, NJ, USA, 2003. ISBN 0-201-61854-0. xxi + 646 pp. LCCN TK5102.9 .G355 2003.

Garcia:1994:NMP

[Gar94] Alejandro Garcia. Numerical Methods for Physics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-151986-7. xiii + 368 pp. LCCN QC20.G37

Garcia:1996:CSA

[Gar96] T. R. Garcia. Communication signal analysis using MATLAB. In Iskander et al. [I^{+}96], pages 616– 619. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Garcia:2000:NMP

[Gar00] Alejandro L. Garcia. Numerical methods for physics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2000. ISBN 0-13-906744-2. vii + 423 pp. LCCN QC20 .G37 2000.

Gardner:2001:SMU

[Gar01] John F. (John Francis) Gardner. Simulations of machines using MATLAB and Simulink. Bookware companion series. Wadsworth, Pacific Grove, CA, USA, 2001. ISBN 0-534-95279-8. xii + 137 pp. LCCN TJ153 .G27 2001.

Garvie:2007:FDS

[Gar07] Marcus R. Garvie. Finitedifference schemes for reactiondiffusion equations modeling predator-prey interactions in MATLAB. Bulletin of Mathematical Biology, 69 (3):931–956, 2007. CODEN BMTBAP. ISSN 0092-8240.

Gautschi:2005:OPM

[Gau05] Walter Gautschi. Orthogonal polynomials (in Matlab). Journal of Computational and Applied Mathematics, 178(1–2):215–234, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Gautschi:2006:OPQ

[Gau06] Walter Gautschi. Orthogonal polynomials, quadrature, and approximation: computational methods and software (in Matlab). In Orthogonal polynomials and special functions, volume 1883 of Lecture Notes in Math., pages 1–77. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Gautschi:2015:POR

[Gau15] Walter Gautschi. Polynomials orthogonal with respect to exponential integrals. Numerical Algorithms, 70(1):215–226, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11075-014-9943-8.

Gautschi:2016:AER

[Gau16a] Walter Gautschi. Algorithm 957: Evaluation of the repeated integral of the coerror function by half-range Gauss–Hermite quadrature. ACM Transactions on Mathematical Software, 42(1):9:1–9:10, February 2016. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:2016:OPM

[Gau16b] Walter Gautschi. Orthogonal polynomials in MAT-LAB: exercises and solutions, volume 26 of Software, environments, and tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2016. ISBN 1-61197-429-1 (paperback), 1-61197-430-5. ix + 335 pp. LCCN QA404.5 .G3564 2016.

Gawronski:1996:BCF

[Gaw96] Wodek Gawronski. Balanced Control of Flexible Structures, volume 211 of Lecture notes in control and information sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-76017-2. xiv + 262 pp. LCCN TA660.F53G39 1996. Includes an appendix of MATLAB functions for control system design methods.

Gawronski:1998:DCS

[Gaw98] Wodek Gawronski. Dynamics and control of structures: a modal approach. Mechanical engineering series. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98527-1 (hardcover). xx + 231 pp. LCCN TA654.15 .G39 1998.

Gawronski:2004:DIC

[Gaw04] Wodek K. Gawronski. Dynamics identification and control of structures. Mechanical engineering series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-40649-2. xxii + 396 pp. LCCN TA654 .G36 2004.

Gorez:1989:UMR

[GB89] R. Gorez and Ph. Buyze.

On the use of MATLAB in robot identification and control. IFAC Proceedings Series, 5:133–135, 1989. CO-DEN IPSEET. ISSN 0741-1146.

Gustafsson:2003:MEE

[GB03] Fredrik Gustafsson and Niclas Bergman. MAT-LAB for Engineers Explained. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 1-85233-697-8. x + 218 pp. LCCN TA345 .G87 2003.

Garvie:2015:SFE

[GBM15] Marcus R. Garvie, John Burkardt, and Jeff Morgan. Simple finite element methods for approximating predator–prey dynamics in two dimensions using Matlab. Bulletin of Mathematical Biology, 77(3):548– 578, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11538-015-0062-z;
 http://link.springer.
com/content/pdf/10.1007/
s11538-015-0062-z.pdf.

Goswami:1999:FWT

[GC99] Jaideva C. Goswami and Andrew K. Chan. Fundamentals of wavelets: theory, algorithms, and applications. Wiley series in microwave and optical engineering. Wiley, New York, NY, USA, 1999. ISBN 0-471-19748-3 (cloth). xvi + 306 pp. LCCN TK5102.9 .G69 1999; TK 5102.9 .G69 1999X ENGI.

Goodman:2025:QMP

[GCM25] Roy H. Goodman, Grace Conte, and Jeremy L. Marzuola. QGLAB: a MAT-LAB package for computations on quantum graphs. SIAM Journal on Scientific Computing, 47(2):B428– B453, ???? 2025. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL https://

epubs.siam.org/doi/10.
1137/23M1627729.

Gertler:1997:IFA

[GCP97] J. J. Gertler, J. B. Cruz, and M. Peshkin, editors. International Federation of Automatic Control: World congress; 13th — June 1996, San Francisco, CA, Proceedings of the World Congress — International Federation of Automatic Control 1996; Conf 13; Vol H. Pergamon Press, New York, NY, USA, 1997. ISBN 0-08-042926-2 (plenary vol.), 0-08-042909-2 (vol. A), 0-08-042910-6 (vol. B), 0-08-042911-4 (vol. C), 0-08-042912-2 (vol. D), 0-08-042913-0 (vol. E), 0-08-042914-9 (vol. F), 0-08-042915-7 (vol. G), 0-08-042916-5 (vol. H), 0-08-042917-3 (vol. I), 0-08-042918-1 (vol. J), 0-08-042919-X (vol. K), 0-08-042920-3 (vol. L), 0-08-042921-1 (vol. M), 0-08-042922-X (vol. N), 0-08-042923-8 (vol. O), 0-08-042924-6 (vol. P), 0-08-042925-4 (vol. Q). LCCN ????

Golubitsky:1999:LAD

[GD99] Martin Golubitsky and Michael Dellnitz. Linear algebra and differential equations using MATLAB. Brooks/Cole, Pacific Grove, CA, USA, 1999. ISBN 0-534-35425-4. xiv + 704 pp. LCCN QA185.D37 G65 1999. Includes CD-ROM.

Gekeler:2008:MMM

[Gek08] Eckart W. Gekeler. Mathematical Methods for Mechanics: a Handbook with MATLAB Experiments. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 3-540-69279-7. 300 (est.) pp. LCCN ????

Geromel:1992:CNN

[GG92] J. C. Geromel and G. O. Guardabassi. Computeraided nicely nonlinear modelling: a set of problem — specific MATLAB functions. In Barker [Bar92], pages 221–226. ISBN 0-08-041269-6. LCCN TJ213 .C57 1992.

Grupp:2004:MIG

[GG04] Frieder Grupp and Florian Grupp. MATLAB 6.5 für Ingenieure: Grundlagen und Programmierbeispiele. (German) [MATLAB 6.5 for Engineers: Foundations and Programming Examples]. Oldenbourg-Verlag, München, Wien, Oldenbourg, third edition, 2004. ISBN 3-486-27376-0. viii + 99 pp. LCCN ????

Ghaziani:2009:NCC

[GGKM09] R. Khoshsiar Ghaziani, W. Govaerts, Yu. A.

Kuznetsov, and H. G. E. Meijer. Numerical continuation of connecting orbits of maps in MATLAB. Journal of Difference Equations and Applications, 15(8-9): 849–875, 2009. CODEN JDEAEA. ISSN 1023-6198 (print), 1563-5120 (electronic).

Goodwin:2001:CSD

[GGS01] Graham C. (Graham Clifford) Goodwin, Stefan F. Graebe, and Mario E. Salgado. Control system design. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-958653-9. ???? pp. LCCN TJ213 .G585 2000.

Gander:1993:SPS

[GH93] Walter Gander and Jiři Hřebíček. Solving Problems in Scientific Computing Using MAPLE and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57329-1, 3-540-57329-1. xiii + 268 pp. LCCN Q183.9 .G36 1993.

Gander:1995:SPS

[GH95] Walter Gander and Jiři Hřebíček. Solving Problems in Scientific Computing Using Maple and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1995. ISBN 3-540-58746-2, 0-387-58746-2. xv + 315 pp. LCCN Q183.9 .G36 1995.

Gander:1997:SPS

[GH97] Walter Gander and Jiři Hřebíček. Solving Problems in Scientific Computing using Maple and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1997. ISBN 3-540-61793-0 (paperback). xvii + 408 pp. LCCN Q183.9 .G36 1997. US$49.95.

Gander:2004:SPS

[GH04] Walter Gander and Jiří Hřebíček. Solving problems in scientific computing using Maple and MAT-LAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2004. ISBN 3-540-21127-6. xxii + 476 pp. LCCN Q183.9 .G36 2004.

Gorecki:2020:HAC

[GHH20] Jan Górecki, Marius Hofert, and Martin Holena. Hierarchical Archimedean copulas for MATLAB and Octave: The HACopula toolbox. Journal of Statistical Software, 93(??):??, ???? 2020. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/

view/v093i10;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v093i10/
v93i10.pdf.

Gazzola:2019:ITM

[GHN19] Silvia Gazzola, Per Christian Hansen, and James G. Nagy. IR tools: a MATLAB package of iterative regularization methods and largescale test problems. Numerical Algorithms, 81(3):773– 811, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Galvez:2006:MBP

[GI06] A. Gálvez and A. Iglesias. Matlab-based problemsolving environment for geometric processing of surfaces. In Mathematical software—ICMS 2006, volume 4151 of Lecture Notes in Comput. Sci., pages 35–

46. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Gidh:1995:API

[Gid95] K. Gidh. An auxiliary processor to integrate MAT-LAB with APL. APL Quote Quad, 26(1):8–??, ???? 1995. CODEN APLQD9. ISSN 0163-6006.

Gilat:2004:MIA

[Gil04] Amos Gilat. MATLAB:

an introduction with applications. Wiley, New York, NY, USA, 2004. ISBN 0-471-43997-5 (paperback). viii + 296 pp. LCCN QA297 .G48 2004 UCR.

Gilat:2005:MIA

[Gil05] Amos Gilat. MATLAB: an introduction with applications. Wiley, New York, NY, USA, second edition, 2005. ISBN 0-471-69420-7. viii + 343 pp. LCCN QA297 .G48 2005.

Gilat:2008:MIA

[Gil08] Amos Gilat. MATLAB: an introduction with applications. Wiley, New York, NY, USA, third edition, 2008. ISBN 0-470-10877-0. x + 374 pp. LCCN ????

Gunther:2003:FMG

[GJ03] Michael Günther and Ansgar Jüngel. Finanzderivate mit MATLAB. (German) [Financial Derivatives with MATLAB]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 2003. ISBN 3-528-03204-9. xi + 302 pp. LCCN ????

Girishwaingankar:2020:PNB

[GJ20] Poorva Girishwaingankar and Sangeeta Milind Joshi. The PHY-NGSC-based ORT run length encoding scheme for video compression. International Journal of Image and Graphics (IJIG), 20 (02):??, April 2020. CO-DEN ???? ISSN 0219-4678. URL https:

//www.worldscientific.
com/doi/10.1142/S0219467820500072.

Guenther:2005:INM

[GK05] Ronald B. Guenther and Abdelwahab Kharab. An Introduction to Numerical Methods: a MATLAB Approach. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-557-2. ???? pp. LCCN ????

Govaerts:2005:NCB

[GKD05] Willy Govaerts, Yuri A.

Kuznetsov, and Annick Dhooge. Numerical continuation of bifurcations of limit cycles in MATLAB. SIAM Journal on Scientific Computing, 27(1):231–252, 2005. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Gajic:1996:MCS

[GL96] Zoran Gajić and M. Lelić.

Modern control system engineering: with MAT-LAB Laboratory experiments. Prentice Hall international series in systems and control engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-134116-2. ???? pp. LCCN TJ213.G26 1996.

Giurgiutiu:2004:MMA

[GL04] Victor Giurgiutiu and Sergey Edward Lyshevski. Micromechatronics: modeling, analysis, and design with MATLAB. Nano- and microscience, engineering, technology, and medicine series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2004. ISBN 0-8493-1593-X. 828 (est.) pp. LCCN TJ163.12 .G58 2004.

Genz:1991:FGG

[GLJ^{+}91] Alan Genz, Zongli Lin, Charles Jones, Dali Luo, and Thorsten Prenzel. Fast Givens goes slow in MAT-LAB. ACM SIGNUM Newsletter, 26(2):11–16, April 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gloor:1998:IPI

[Glo98] Oliver Gloor, editor. IS-SAC 98: Proceedings of the 1998 International Symposium on Symbolic and Algebraic Computation, August 13–15, 1998, University of Rostock, Germany. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-002-3. LCCN ????

Gilbert:1992:SMM

[GMS92] John R. Gilbert, Cleve Moler, and Robert Schreiber. Sparse matrices in MAT-LAB: Design and implementation. SIAM Journal on Matrix Analysis and Applications, 13(1):333–356, January 1992. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp

from math.liu.se in pub/
references.

Gao:1995:RCD

[GMT95] Xiaoming Gao, Brian D.

McVey, and Robert L. Tokar. Robust controller design of four wheel steering systems using μ synthesis techniques. Proceedings of the IEEE Conference on Decision and Control, 1:875–882, 1995. CO-DEN PCDCDZ. ISSN 0191-2216. IEEE catalog number 95CH3580-3.

Gialamas:1996:DMC

[GMT96] T. P. Gialamas, D. A.

Manolas, and D. T. Tsahalis. Development of a MATLAB code based on a weak formulation of wave propagation through a porous material. comparison with experimental results. In Sas [Sas96], pages 775–786. ISBN 90-73802-60-1. LCCN ????

Gundersen:1997:MSD

[GO97] Øyvind Gundersen and Morten I. Onsøien. MAT-LAB simulator for directional solidification of ductile cast iron. SINTEF, Materials Technology, Trondheim, Norway, 1997. ISBN 82-14-00471-3. 22 + 41 pp.

Gockenbach:2002:PDE

[Goc02] Mark S. Gockenbach. Partial differential equations: analytical and numerical methods. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2002. ISBN 0-89871-518-0. xxii + 614 pp. LCCN QA377 .G63 2002.

Goldberg:1991:MTA

[Gol91] Jack L. Goldberg. Matrix Theory with Applications. McGraw-Hill, New York, NY, USA, 1991. ISBN 0-07-557200-1. xxi + 520 pp. LCCN QA188 .G645 1991.

Gomez:2015:SSM

[Góm15] Victor Gómez. SSMMATLAB: A set of MATLAB programs for the statistical analysis of state space models. Journal of Statistical Software, 66(9):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v066i09;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v066i09/v66i09.pdf.

Gopi:2010:MSD

[Gop10] E S. Gopi. Mathematical summary for digital signal processing applications with Matlab. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3746-2. ???? pp. LCCN ????

Gottling:1995:MAC

[Got95] James G. Gottling. Matrix Analysis of Circuits Using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-127044-3. viii + 130 pp. LCCN TK7867.G633 1995.

Goualard:2020:GRF

[Gou20] Frédéric Goualard. Generating random floatingpoint numbers by dividing integers: a case study. In Krzhizhanovskaya et al. [KZL^{+}20], pages 15–28. ISBN 3-030-50416-6, 3-030-50417-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). URL

https://link.springer.
com/book/10.1007/978-3-
030-50417-5.

Garcia:1996:MPS

[GP96] Alejandro L. Garcia and Cecile Penland. MATLAB Projects for Scientists and Engineers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-460171-8. ???? pp. LCCN ???? URL ftp:/

/ftp.mathworks.com/pub/
books/penland/.

Goldberg:1998:DES

[GP98] Jack L. (Jack Leonard) Goldberg and Merle C. Potter. Differential equations: a systems approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-211319-8. xvii + 476 pp. LCCN QA372 .G598 1998.

Gu:2005:RCS

[GPK05] Da-Wei Gu, Petko Petkov, and Mihail M. Konstantinov. Robust Control Systems with MATLAB: Design Examples with the μ-Analysis and Synthesis Toolbox. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-85233-983-7. xiv + 389 pp. LCCN TJ217.2 .G8 2005. Includes CD-ROM.

Gupta:2014:ESV

[GPR14] Pratishtha Gupta, G N Purohit, and Manisha Rathore. Estimating speed of vehicle using centroid method in MATLAB. International Journal of Computer Applications, 102(??): 1–8, September 2014. CO-DEN ???? ISSN 0975-

                                        8887. URL https://www.
                                        ijcaonline.org/archives/
                                        volume102/number14/17880-
                                        8451/.

Gatto:1993:RMV

[GR93] Marino Gatto and Andrea Rizzoli. Review of MAT-LAB, Version 4.0. Natural Resource Modeling, 7(1):85– 88, Winter 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Granda:1997:NDB

[GR97] J. J. Granda and J. Reus.

New developments in bond graph modeling software tools: The computer aided modeling program Camp-G and Matlab. In IEEE [IEE97a], pages 1542–1547. ISBN 0-7803-4054-X, 0-7803-4053-1, 0-7803-4055-8. ISSN 1062-922X. LCCN Q300 .I485 1997. Five volumes.

Grace:1992:CST

[Gra92] Andrew Grace. Control System Toolbox for use with MATLAB: User’s Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, July 1992. ISBN ???? various pp. LCCN QA188 .C66 1992.

Grayson:1993:ACF

[Gra93] L. P. Grayson, editor. 23rd Annual conference on frontiers in education — 1993 Nov: Washington; DC, Frontiers in Education Conference 23rd. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1482-4, 0-7803-1483-2, 0-7803-1484-0. ISSN 0190-5848. LCCN ????

Grayson:1994:FEP

[Gra94] Lawrence P. Grayson, editor. Frontiers in education: proceedings, twentyfourth annual conference, November 2–6, 1994, San Jose, California: educating engineers for world competition. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-2414-5, 0-7803-2413-7, 0-7803-2415-3. ISSN 0190-5848. LCCN T 62 F76 1994.

Gramlich:2004:ALA

[Gra04] Günter Gramlich. Anwendungen der Linearen Algebra mit MATLAB. (German) [Applications of Linear Algebra with MATLAB]. Carl Hanser, München, Germany, 2004. ISBN 3-446-22655-9. 179 (est.) pp. LCCN ????

Gran:2007:NCS

[Gra07]  Richard J. Gran.  Nu-

merical computing with SIMULINK. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2007. ISBN 0-89871-637-3. ???? pp. LCCN QA297 .G676 2007.

Gray:2011:ISD

[Gra11] Michael A. Gray. Introduction to the simulation of dynamics using Simulink. Computational science series. Taylor and Francis, Boca Raton, FL, USA, 2011. ISBN 1-4398-1897-5 (hardcover). xvi + 308 pp. LCCN TA352 .G73 2011.

Garcia-Rosello:2012:TLW

[GRDL^{+}12] Emilio Garcia-Rosello, Jacinto G. Dacosta, Maria J. Lado, Arturo J. Mendez, and Jose Garcia Perez-Schofield. Two-layer wrapping for COTS software integration: An experience with Matlab. IEEE Software, 29(4):76–82, July/ August 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gregory:1994:MRC

[Gre94] I. Gregory. Matlab as a robust control design tool. In Wieseman [Wie94], pages 485–496. ISBN ???? ISSN 0191-7811. LCCN ????

Greatorex:1996:FME

[Gre96] S. A. Greatorex, editor. Flight mechanics/ estimation theory symposium: – May 1996, Greenbelt, MD, number 3333 in NASA CONFERENCE PUBLICATION 1996. National Aeronautics and Space Administration, Washington, DC, USA, 1996. ISBN ???? ISSN 0191-7811. LCCN ????

Green:2016:CMC

[Gre16] Dan Green. Cosmology with MATLAB: with companion media pack. World Scientific Publishing Co. Pte. Ltd., Singapore, 2016. ISBN 981-310-839-8 (hardcover), 981-310-840-1 (paperback). xi + 250 pp. LCCN QB981 .G74 2016.

Gauhar:2021:FVM

[GRH^{+}21] Ayesha Gauhar, Adnan Rashid, Osman Hasan, João Bispo, and João M. P. Cardoso. Formal verification of matrix based MATLAB models using interactive theorem proving. PeerJ Computer Science, 7:e440:1–e440:21, March 2021. ISSN 2376-5992.

Grimble:1994:RIC

[Gri94] Michael J. Grimble. Robust industrial control: optimal design approach for polynomial systems. Prentice-Hall international series in systems and control engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-655283-8. xvi + 597 pp. LCCN TJ213 .G765 1994.

Grigoriu:1995:ANP

[Gri95] Mircea Grigoriu. Applied Non-Gaussian Processes: examples, theory, simulation, linear random vibration, and MATLAB solutions. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-367095-3. xii + 442 pp. LCCN QA274.G76 1995.

Grossman:1994:ELA

[Gro94] Stanley I. Grossman. Elementary Linear Algebra. Saunders College Publishing, Ft. Worth, TX, USA, fifth edition, 1994. ISBN 0-03-097354-6. xx + 634 pp. LCCN QA184.G76 1994.

Groetsch:1999:IPA

[Gro99]  C. W. Groetsch.  In-

verse problems: activites for undergraduates. Classroom resource materials. Mathematical Association of America, Washington, DC, USA, 1999. ISBN 0-88385-716-2. xii + 222 pp. LCCN QA371 .G73 1999. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam0210/99062793.
html;   http://www.loc.
gov/catdir/toc/cam027/
99062793.html.

Gross:2005:SAW

[Gro05] Frank B. Gross. Smart antennas for wireless communications: with MAT-LAB. McGraw-Hill, New York, NY, USA, 2005. ISBN 0-07-144789-X (hardback). xiv + 270 pp. LCCN TK7871.67.A33 G767 2005.

Grundmann:2004:FMG

[Gru04] Wolfgang Grundmann. Finanzmathematik mit MAT-LAB. (German) [Financial Mathematics with MAT-LAB]. Teubner, Stuttgart, Germany; Leipzig, Germany, 2004. ISBN 3-519-00450-X. 226 pp. LCCN ????

Gilsing:2007:SPS

[GS07] Hagen Gilsing and Tony Shardlow. SDELab: a package for solving stochastic differential equations in MATLAB. Journal of Computational and Applied Mathematics, 205(2): 1002–1018, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Gilat:2008:NME

[GS08] Amos Gilat and Vish Subramaniam. Numerical methods for engineers and scientists: an introduction with applications using MAT-LAB. Wiley, New York, NY, USA, second edition, 2008. ISBN 0-471-73440-3. xx + 459 pp. LCCN QA297 .G49

                                        2008. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0835/2008299899-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0835/2008299899-
                                        d.html;   http://www.loc.

gov/catdir/enhancements/
fy0835/2008299899-t.html.

Griffiths:2012:TWA

[GS12] Graham W. Griffiths and W. E. Schiesser. Traveling wave analysis of partial differential equations: numerical and analytical methods with MATLAB and Maple. Academic Press, New York, NY, USA, 2012. ISBN 0-12-384652-8 (hardcover). xiii + 447 pp. LCCN QA374 .G75 2012.

Gibson:1995:IGG

[GSM95] Warren C. Gibson, Christian A. Smith, and Donald J. McTavish. Implementation of the Golla-Hughes-McTavish (GHM) method for viscoelastic materials using MATLAB and NAS-TRAN [2445-29]. In Johnson [Joh95], pages 312– 323. ISBN 0-8194-1794-7. ISSN 0361-0748. LCCN TS510.S63 v.2445; TA355 .S528 1995.

Gonzalez:2005:IAL

[GSS05] Narciso García González, Mariano Lozano Sánchez, and Mere Macià Soler. Una invitación al an?lisis numérico con Matlab. Popular Libros, Albacete, Spain, 2005. ISBN 84-932789-9-8. vi + 279 pp. LCCN ????

Gulick:2012:ECF

[Gul12] Denny Gulick. Encounters with Chaos and Fractals. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2012. ISBN 1-58488-517-3 (hardcover). xvi + 371 pp. LCCN Q172.5.C45 G85 2012.

Gupta:2002:ECS

[Gup02] Sudhir (Sudhir K.) Gupta. Elements of control systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-011989-X. xvi + 573 pp. LCCN TJ213 .G883 2002.

Golub:1989:MC

[GV89] Gene H. Golub and Charles F. Van Loan. Matrix Computations, volume 3 of Johns Hopkins Series in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, second edition, 1989. ISBN 0-8018-3772-3 (hardcover), 0-8018-3739-1 (paperback). xix + 642 pp. LCCN QA188 .G65 1989. US$14.50.

Graves:1994:UMT

[GV94] C. H. Graves and S. M. Veres. Using MATLAB toolbox “GBT” in identification and control. In Anonymous [Ano94c], pages 11/1–11/5. ISBN ???? LCCN ????

Guan:2008:PMO

[GV08] Yun Guan and Jan Verschelde. PHClab: a MAT-LAB/Octave interface to PHCpack. In Software for algebraic geometry, volume 148 of IMA Vol. Math. Appl., pages 15– 32. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008.

Galuppini:2025:TMN

[GV25] Giacomo Galuppini and Paolo Venini. A transfer matrix norm-based framework for multimaterial topology optimization: Methodology and MATLAB implementation. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/
18/5/282.

Gonzalez:2002:DIP

[GW02] Rafael C. Gonzalez and Richard E. (Richard Eugene) Woods. Digital image processing. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2002. ISBN 0-201-18075-8. xx + 793 pp. LCCN TA1632 .G66 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy031/2002276271.html.

Gerald:2004:ANA

[GW04] Curtis F. Gerald and Patrick O. Wheatley. Applied numerical analysis. Pearson/Addison-Wesley, Boston, MA, USA, seventh edition, 2004. ISBN 0-321-13304-8. x + 609 pp. LCCN QA297 .G47 2004.

Grewal:2001:GPS

[GWA01] Mohinder S. Grewal, Lawrence Randolph Weill, and Angus P. Andrews. Global positioning systems, inertial navigation, and integration. Wiley, New York, NY, USA, 2001. ISBN 0-471-35032-X (cloth). xix + 392 pp. LCCN G109.5 .G74 2001. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://prod.library.utoronto.
                                        ca/datalib/datart/datalib/
                                        mixed_media/1000/1919/.

Gonzalez:2004:DIP

[GWE04] Rafael C. Gonzalez, Richard E. (Richard Eugene) Woods, and Steven L. Eddins. Digital Image processing using MATLAB. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-008519-7. xiv + 609 pp. LCCN TA1637 .G66 2004. US$120.00.

Huang:1996:LCP

[H^{+}96] C.-H. (Chua-Huang) Huang et al., editors. Languages and compilers for parallel computing: 8th international workshop, LCPC 95, Columbus, Ohio, USA, August 10–12, 1995: proceedings, volume 1033 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CO-DEN LNCSD9. ISBN 3-540-60765-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1995.

Hava:1995:SGM

[HA95] S. Hava and M. Auslender.

Silicon grating-based mirror for 1.3-$mu@m polarized beams: Matlab-aided design. Applied Optics, 34(6): 1053–1058, February 1995. CODEN APOPAI. ISSN 0003-6935.

Hahn:1997:EMS

[Hah97] Brian D. Hahn. Essential MATLAB for scientists and engineers. Arnold, London, UK, 1997. ISBN 0-340-69144-1 (Arnold), 0-470-25013-5 (Wiley). x + 265 pp. LCCN TA347.D4 H34 1997.

Hahn:2002:EMS

[Hah02] Brian D. Hahn. Essential MATLAB for scientists and engineers. Butterworth-Heinemann, Oxford, UK, second edition, 2002. ISBN 0-7506-5240-3. xvi + 298 pp. LCCN TA345 .H314 2002.

Haigh:2008:CMM

[Hai08]  Thomas Haigh.  Cleve

Moler: Mathematical software pioneer and creator of Matlab. IEEE Annals of the History of Computing, 30(1):87–91, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hamza:1993:MSI

[Ham93] M. H. Hamza, editor. Modelling and simulation: International conference — May 1993, Pittsburgh, PA. IASTED, Anaheim, CA, USA, 1993. ISBN 0-88986-173-0. ISSN 1021-8181. LCCN ????

Hamza:1996:MSP

[Ham96] M. H. Hamza, editor. Modelling and simulation: proceedings of the IASTED/ ISMM International Conference, April 25–27, 1996, Pittsburgh, Pennsylvania, USA, MODELLING AND SIMULATION -IASTED INTERNATIONAL SYMPOSIUM-1021- 8181 1996. IASTED, ACTA Press, Anaheim, CA, USA, 1996. ISBN 0-88986-201-X. LCCN QA76.9.C65I28 1996.

Hubert:2002:LUS

[HAM02] L. J. Hubert, P. Arabie, and J. J. Meulman. Linear unidimensional scaling in the L_{2}-norm: basic optimization methods using MAT-LAB. J. Classification, 19 (2):303–328, 2002. ISSN 0176-4268.

Hubert:2006:SRP

[HAM06] Lawrence J. Hubert, Phipps Arabie, and Jacqueline Meulman. The structural representation of proximity matrices with MAT-LAB. ASA-SIAM series on statistics and applied probability. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2006. ISBN 0-89871-607-1 (paperback). xvi + 214 pp. LCCN QA195 .H83 2006 UCB.

Hansen:1992:RTM

[Han92] Per Christian Hansen. Regularization tools. A Matlab package for analysis and solution of discrete ill-posed problems. Report UNIC-92-03, UNI·C, Technical University of Denmark, DK-2800 Lyngby, Denmark, June 1992. ???? pp.

Hansen:1994:RTM

[Han94] Per Christian Hansen. REG-ULARIZATION TOOLS: A Matlab package for analysis and solution of discrete ill-posed problems. Numerical Algorithms, 6(1-2):1– 35, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hansen:1999:RTV

[Han99] Per Christian Hansen. Regularization Tools Version 3.0 for Matlab 5.2. Numerical Algorithms, 20(2–3):195– 196, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.
com/content/getfile/5058/
18/8/abstract.htm;   http:
//ipsapp007.kluweronline.
com/content/getfile/5058/
18/8/fulltext.pdf.

Hannachi:2006:BRB

[Han06] A. Hannachi. Book review: Exploratory Data Analysis with MATLAB, by W. L. Martinez; A. R. Martinez. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):390– 391, March 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
3559689.

Hansen:2007:RTV

[Han07] Per Christian Hansen. Regularization Tools version 4.0 for Matlab 7.3. Numerical Algorithms, 46(2): 189–194, 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hartfiel:2001:MTA

[Har01] Darald J. Hartfiel. Matrix theory and applications with MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 1-58488-108-9. ix + 371 pp. LCCN QA188 .H37 2001. US$79.95, UK£29.99.

Hargreaves:2002:IAM

[Har02] Gareth I. Hargreaves. Interval analysis in MAT-LAB. Numerical analysis report 416, Manchester Centre for Computational Mathematics, Manchester, England, December 2002. 49 pp. URL http:

//www.maths.man.ac.uk/~
nareports/narep416.pdf.

Harris:2004:MSP

[Har04] Fred (Fredric J.) Harris.

Multirate signal processing for communication systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-146511-2. xiii + 478 pp. LCCN TK5103.7 .H38 2004.

Hardy:2005:LAE

[Har05] Kenneth Hardy. Linear algebra for engineers and scientists using MATLAB. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-010988-6 (international edition), 0-13-906728-0. xv + 480 + 6 pp. LCCN QA184.2 .H37 2005.

Hasbun:2012:UTP

[Has12] Javier E. Hasbun. Unifying two popular-but-seemingly-dissimilar platforms: Matlab and Java. Computing in Science and Engineering, 14(3):6–7, May/June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hatch:2001:VSU

[Hat01] Michael R. Hatch. Vibration simulation using MATLAB and ANSYS. Chapman and Hall/CRC, Boca Raton, FL, USA, 2001. ISBN 1-58488-205-0. 654 (est.) pp. LCCN TJ177 .H38 2001.

Haupt:1990:AAT

[Hau90] Randy L. Haupt. An antenna array tutorial using MATLAB. CoED, 10(4): 59–61, October–December 1990. CODEN CWLJDP. ISSN 0736-8607.

Haugen:1996:LMP

[Hau96a] Finn Haugen. Lær MAT-LAB: på 5 Timer! (Norwegian) [Learn MATLAB: in 5 Hours!]. Haugen, ??, Norway, 1996. ISBN 82-91748-00-4. ???? pp. LCCN ????

Haugen:1996:LSP

[Hau96b] Finn Haugen. Lær Simulink: på 3 Timer! (Norwegian) [Learn Simulink in 3 hours!]. Haugen, ??, Norway, 1996. ISBN 82-91748-01-2. ???? pp. LCCN ????

Haugen:1997:LMH

[Hau97] Finn Haugen. Learn MAT-LAB 5 in 6 hours! Tech Teach, ????, 1997. ISBN 82-91748-02-0. ???? pp. LCCN ????

Haugen:2001:RMT

[Hau01] Finn Haugen. Lær MAT-LAB trinn for trinn. (Norwegian) [Learn Matlab step by step]. TechTeach, Skien, Norway, 2001. ISBN 82-91748-07-1. 205 pp.

Hayes:1996:SDS

[Hay96a] M. H. (Monson H.) Hayes.

Statistical digital signal processing and modeling. Wiley, New York, NY, USA, 1996. ISBN 0-471-59431-8 (cloth). xv + 608 pp. LCCN TK5102.9 .H39 1996.

Haykin:1996:AFT

[Hay96b] Simon S. Haykin. Adaptive filter theory. Prentice-Hall information and system sciences series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1996. ISBN 0-13-322760-X (book), 0-13-393539-6 (instructor’s manual). xvii + 989 pp. LCCN TK7872.F5H368 1996.

Haykin:1999:NNC

[Hay99] Simon S. Haykin. Neural networks: a comprehensive foundation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-273350-1. xxi + 842 pp. LCCN QA76.87 .H39 1999.

Haykin:2001:CS

[Hay01] Simon S. Haykin. Communication systems. Wiley, New York, NY, USA, fourth edition, 2001. ISBN 0-471-17869-1 (cloth). xviii + 816 pp. LCCN TK5101 .H37 2001.

Haykin:2002:AFT

[Hay02] Simon S. Haykin. Adaptive filter theory. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2002. ISBN 0-13-090126-1. xvi + 920 pp. LCCN TK7872.F5 H368 2002.

Hayes:2004:MAT

[HB04] Jeremiah F. Hayes and Thimma V. J. Ganesh Babu. Modeling and analysis of telecommunications networks. Wiley-Intersci-ence, New York, NY, USA, 2004. ISBN 0-471-34845-7 (Cloth hc.). xx + 393 pp. LCCN TK5101 .H39 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley046/2003020806.
                                        html;   http://www.loc.
                                        gov/catdir/description/
                                        wiley041/2003020806.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley041/2003020806.
                                        html.

Hartley:1994:DSD

[HBC94] Tom T. Hartley, Guy O. Beale, and Stephen P. Chicatelli. Digital Simulation of Dynamic Systems: a Control Theory Approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-219957-2. xiii + 393 pp. LCCN QA 76.9 C65 H36 1994.

Hanh:2016:NFF

[HBT16] Le Thi My Hanh, Nguyen Thanh Binh, and Khuat Thanh Tung. A novel fitness function of metaheuristic algorithms for test data generation for Simulink models based on mutation analysis. The Journal of Systems and Software, 120(??):17– 30, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0164121216301017.

Hill:1992:SPF

[HC92] David R. Hill and Gene B.

Chase. Sketching piecewise functions in MAT-LAB. Collegiate Microcomputer, X(3):160–??, August 1992. CODEN CMICDL. ISSN 0731-4213.

Hagin:1995:CEU

[HC95] Frank Hagin and Jack Cohen. Calculus Explorations Using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-328600-2. ???? pp. LCCN ???? This is a supplemental lab manual for [BS95b].

Hatch:2000:VSU

[HC00] Michael R. Hatch and CR-CnetBASE (Online service). Vibration simulation using MATLAB and ANSYS. Chapman and Hall/CRC, Boca Raton, FL, USA, 2000. 654 pp. LCCN TJ177 .H38 2001.

Herrera:2023:NIH

[HCBAEC23] César Herrera, Ricardo Corrales-Barquero, Jorge Arroyo-Esquivel, and Juan G. Calvo. A numerical implementation for the highorder 2D virtual element method in MATLAB. Numerical Algorithms, 92(3): 1707–1721, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s11075-
                                        022-01361-4.

Haverkamp:1997:STM

[HCV97] Bert Haverkamp, Chun Tung Chou, and Michel Vemaegen. SMI toolbox: a MAT-LAB toolbox for state space model identification. Journal A. Revue A. Tijdschrift A. Zeitschrift A, 38(3):34– ??, ???? 1997. CODEN JR-NAAD. ISSN 0771-1107.

Hagan:1996:NND

[HDB96] Martin T. Hagan, Howard B. Demuth, and Mark Beale.

Neural Network Design. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1996. ISBN 0-534-94332-2. various pp. LCCN QA76.87.H34 1996.

Harman:1997:AEM

[HDR97] Thomas L. Harman, James Dabney, and Norman Richert. Advanced engineering mathematics using MATLAB V.4. The PWS BookWare companion series. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1997. ISBN 0-534-94350-0. xxxix + 643 pp. LCCN TA345.H32 1997.

Harman:2000:AEM

[HDR00] Thomas L. Harman, James Dabney, and Norman Richert. Advanced engineering mathematics with MATLAB. Brooks/Cole, Pacific Grove, CA, USA, second edition, 2000. ISBN 0-534-37164-7. xxxiv + 750 pp. LCCN TA345 .H324 2000.

Heath:1997:SCI

[Hea97] Michael T. Heath. Scientific Computing: An Introductory Survey. McGraw-Hill, New York, NY, USA, 1997. ISBN 0-07-027684-6. xx + 408 pp. LCCN Q183.9.H4 1997.

Hellerstein:2004:FCC

[Hel04] Joseph Hellerstein, editor. Feedback control of computing systems. Wiley-In-terscience, New York, NY, USA, 2004. ISBN 0-471-26637-X (cloth). xx + 429 pp. LCCN TJ216 .F44 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley047/2004040490.
;   http://www.loc.gov/
catdir/bios/wiley047/2004040490.
html;   http://www.loc.
gov/catdir/description/
wiley041/2004040490.;
http://www.loc.gov/catdir/
description/wiley041/2004040490.
html;   http://www.loc.
gov/catdir/toc/wiley041/
2004040490.;   http://
www.loc.gov/catdir/toc/
wiley041/2004040490.html;
UCLA.

Hench:1994:SMA

[Hen94] D. L. Hench. Software for mathematical analysis. WESCON Conference Record, pages 258–263, 1994. CODEN WCREDI. ISSN 1044-6036, 0083-8837.

Henrion:2007:CDP

[Hen07]  Didier Henrion.  H_{∞}

controller design on the COMPl_{e}ib problems with the robust control toolbox for Matlab. International Journal of Tomography & Statistics, 6(S07):69– 73, 2007. ISSN 0972-9976.

Herrera:1996:DCR

[Her96] A. Herrera. Design of a course of random signals using MATLAB. In Iskander et al. [I^{+}96], pages 1219–1222. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Herniter:2001:PM

[Her01] Marc E. Herniter. Programming in MATLAB. Brooks/ Cole-Thomson Learning, Pacific Grove, CA, USA, 2001. ISBN 0-534-36880-8 (paperback). xx + 486 pp. LCCN QA297 .H43 2001.

Hasani:2024:CTI

[HG24] Mehri Hasani and Masoud Ghods. Calculation of topological indices along with MATLAB coding in QSPR analysis of calcium channel-blocking cardiac drugs. Journal of Mathematical Chemistry, 62(10): 2456–2477, November 2024. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (electronic). URL

https://link.springer.
com/article/10.1007/s10910-
023-01570-9.

Hasani:2025:MQP

[HGM^{+}25] Mehri Hasani, Masoud Ghods, Sourav Mondal, Muhammad Kamran Siddiqui, and Imran Zulfiqar Cheema. Modeling QSPR for pyelonephritis drugs: a topological indices approach using MATLAB. The Journal of Supercomputing, 81 (3):??, February 2025. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s11227-
                                        025-06967-8.

Higham:2000:MG

[HH00] Desmond J. Higham and Nicholas J. Higham. MAT-LAB Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-469-9. xxii + 283 pp. LCCN QA297 .H5217 2000.

Hairer:2003:GMP

[HH03] Ernst Hairer and Martin Hairer. GniCodes— Matlab programs for geometric numerical integration. In Frontiers in numerical analysis (Durham, 2002), Universitext, pages 199–240. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003.

Haupt:2004:PGA

[HH04] Randy L. Haupt and Sue Ellen Haupt. Practical genetic algorithms. Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-45565-2. xvii + 253 pp. LCCN QA402.5 .H387 2004; QA402.5 HAU.

URL ftp://uiarchive.
cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.
loc.gov/catdir/bios/wiley047/
2004043360.html;   http:
//www.loc.gov/catdir/description/
wiley041/2004043360.html;
 http://www.loc.gov/catdir/
toc/wiley041/2004043360.
html.

Higham:2005:MG

[HH05] D. J. (Desmond J.) Higham and Nicholas J. Higham. MATLAB guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2005. ISBN 0-89871-578-

4. xxiii + 382 pp. LCCN QA297 .H5217 2005.

Higham:2017:MG

[HH17] Desmond J. Higham and Nicholas J. Higham. MAT-LAB guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2017. ISBN 1-61197-465-8. xxvi + 476 pp.

Haller:1995:MED

[HHF95] C. L. Haller, D. L. Harris, and R. T. Fink. Modeling of emergency diesel generator and governor transient response using MATLAB/ SIMULINK. In Anonymous [Ano95j], pages 1–?? ISBN ???? ISSN 0148-7191. LCCN ????

Husbands:1997:PPS

[HI97] Parry Husbands and Charles L. Isbell, Jr. The parallel problems server. Technical report, Learning & Vision Group, Artificial Intelligence Laboratory, Massachusetts Institute Of Technology, Cambridge, MA, USA, October 3, 1997. 2 pp. URL www.

                                        ai.mit.edu/projects/lv/
                                        Abstracts/zz-husbell.pdf.

Higham:1989:CTM

[Hig89] Nicholas J. Higham. A collection of test matrices in MATLAB. Numerical Analysis Report 172, University of Manchester, Manchester M13 9PL, England, July 1989.

Higham:1991:ACT

[Hig91] Nicholas J. Higham. Algorithm 694: a collection of test matrices in MAT-LAB. ACM Transactions on Mathematical Software, 17(3):289–305, September 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/
                                        Abstracts/0098-3500/116805.
                                        html.

Higham:1993:TMT

[Hig93] Nicholas J. Higham. The Test Matrix Toolbox for MATLAB. Numerical Analysis Report 237, University of Manchester, Manchester M13 9PL, England, December 1993. URL ftp://vtx.

ma.man.ac.uk/pub/narep/
narep237.ps.Z.

Higham:1995:TMT

[Hig95] Nicholas J. Higham. The Test Matrix Toolbox for Matlab, version 3.0. Numerical Analysis Report 276, Manchester Centre for Computational Mathematics, Manchester, England, September 1995. 70

pp. URL http://p2chpd-
cluster.univ-lyon1.fr/
P2CHPD/matlab/pdf_doc/
otherdocs/testmatrix.pdf.

Supersedes Numerical Analysis Report No. 237. In preparation.

Higham:1996:ASN

[Hig96] Nicholas J. Higham. Accuracy and Stability of Numerical Algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-355-2 (paperback). xxviii + 688 pp. LCCN QA297.H53 1996. US$39.00. Typeset with L^{A}TEX2e.

Higham:2002:NWI

[Hig02a] Desmond J. Higham. Nine ways to implement the binomial method for option valuation in MATLAB. SIAM Review, 44(4):661– 677 (electronic) (2003), December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39326.

Higham:2002:ASN

[Hig02b] Nicholas J. Higham. Accuracy and stability of numerical algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2002. ISBN 0-89871-521-0. xxx + 680 pp. LCCN QA297 .H53 2002.

Higham:2002:MCT

[Hig02c] Nicholas J. Higham. The Matrix Computation Toolbox for MATLAB (version 1.0). Numerical Analysis Report 410, Manchester Centre for Computational Mathematics, Manchester, England, August 2002. 19 pp.

Higham:2004:IFO

[Hig04] D. J. (Desmond J.) Higham. An introduction to financial option valuation: mathematics, stochastics, and computation. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-83884-3, 0-521-54757-1 (paperback). ???? pp. LCCN HG6024.A3 H532 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam041/2003069572.
html;   http://www.loc.
gov/catdir/toc/cam041/
2003069572.html.

Higham:2015:MCT

[Hig15] Nicholas J. Higham. Matrix computation toolbox. Web site., 2015.

URL http://www.ma.man.
ac.uk/~higham/mctoolbox.

Hill:1991:IEU

[Hil91] D. R. Hill. Instructional experiments using MATLAB. In Lum [Lum91], pages 147–

161. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Hill:1996:ELA

[Hil96] Richard O. Hill, Jr. Elementary Linear Algebra with Applications. Saunders College Publishing, Ft. Worth, TX, USA, third edition, 1996. ISBN 0-03-010347-

9. xvii + 445 pp. LCCN QA184 .H55 1996.

Himpe:2023:EEG

[Him23] Christian Himpe. emgr — EMpirical GRamian framework version 5.99. ACM Transactions on Mathematical Software, 49(3):31:1– 31:??, September 2023. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
10.1145/3609860.

Hanselman:1995:MTC

[HK95] Duane C. Hanselman and Benjamin C. Kuo. MAT-LAB Tools for Control System Analysis and Design. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1995. ISBN 0-13-202293-1 (PC version), 0-13-202574-4 (Macintosh version). viii + 207 pp. LCCN TJ216.K818 1995.

Ham:2001:PNS

[HK01a] Fredric M. Ham and Ivica Kostanic. Principles of neurocomputing for science and engineering. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-025966-6. xxx + 642 pp. LCCN QA76.87 .H352 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/mh021/00042710.
html;   http://www.loc.
gov/catdir/toc/mh021/00042710.
html.

Hill:2001:MMA

[HK01b] David R. (David Ross) Hill and Bernard Kolman. Modern matrix algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-948852-9. xv + 511 pp. LCCN QA188 .H55 2000.

Hayt:2002:ECA

[HKD02] William Hart Hayt, Jack E.

(Jack Ellsworth) Kemmerly, and Steven M. Durbin. Engineering circuit analysis. McGraw-Hill, New York, NY, USA, sixth edition, 2002. ISBN 0-07-228364-5 , 0-07-112227-3 (international edition). xviii + 781 pp. LCCN TK454 .H4 2002.

Hojati:2020:MPA

[HKF^{+}20] Sahar Hojati, Rahele Kafieh, Parisa Fardafshari, Masoud Aghsaei Fard, and Hatef Fouladi. A MAT-LAB package for automatic extraction of flow index in OCT-A images by intelligent vessel manipulation. SoftwareX, 12(??):Article 100510, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711020300364.

Hanselman:1995:SEM

[HL95] Duane C. Hanselman and Bruce Littlefield. The Student Edition of Matlab: High-Performance Numeric Computation and Visualization Software: Version 4 User’s Guide. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-184979-4, 0-13-184995-6 (Windows version), 0-13-184987-5 (Macintosh version), 0-13-459207-7 (Macintosh/Power Macintosh version). xii + 833 pp. LCCN QA297 .S8427 1995.

Hanselman:1996:MMC

[HL96] Duane Hanselman and Bruce Littlefield. Mastering MATLAB: a comprehensive tutorial and reference. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-191594-0. xii + 542 pp. LCCN QA297.H295 1996.

Hanselman:1997:SEM

[HL97] Duane C. Hanselman and Bruce Littlefield. The student edition of MATLAB: version 5, user’s guide. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-272550-9 (paperback user’s guide), 0-13-272485-5 (user’s guide), 0-13-272477-4 (set), 0-13-022598-3 (CD-ROM). xxxix + 429 pp. LCCN QA297 .S8436 1997.

Hanselman:1998:MMC

[HL98] Duane C. Hanselman and Bruce Littlefield. Mastering MATLAB 5: a comprehensive tutorial and reference. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-858366-8. xviii + 638 pp. LCCN QA297.H296 1998.

Hanselman:2001:MMC

[HL01] Duane C. Hanselman and Bruce Littlefield. Mastering MATLAB 6: a comprehensive tutorial and reference. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-019468-9. xviii + 814 pp. LCCN QA297 .H293 2001.

Hanke:2003:CIN

[HL03a] Michael Hanke and René Lamour. Consistent initialization for nonlinear index-2 differential-algebraic equation: large sparse systems in MATLAB. Numerical Algorithms, 32(1):67–85, 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Henrion:2003:GGO

[HL03b] Didier Henrion and Jean-Bernard Lasserre. GloptiPoly: Global optimization over polynomials with Matlab and SeDuMi. ACM Transactions on Mathematical Software, 29(2):165– 194, June 2003. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hanselman:2005:MM

[HL05] Duane C. Hanselman and Bruce Littlefield. Mastering MATLAB 7. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-143018-1 (paperback), 0-13-185714-2. xi + 852 pp. LCCN QA297 .H2963 2005.

Holmi:2022:WMD

[HL22] Joonas Tapani Holmi and Harri Lipsanen. WITio: a MATLAB data evaluation toolbox to script broader insights into big data from WITec microscopes. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271102200019X.

Hunt:2005:DEM

[HLOR05] Brian R. Hunt, Ronald L. Lipsman, John E. Osborn, and Jonathan M. Rosenberg, editors. Differential equations with MAT-LAB: updated for MATLAB 7 and Simulink 6. Wiley, New York, NY, USA, second edition, 2005. ISBN 0-471-71812-2 (paperback). x + 295 pp. LCCN QA371.5.D37 D55 2005.

Hollingsworth:1996:PTM

[HLP96] J. Hollingsworth, K. Liu, and V. Paúl Pauca. Parallel toolbox for MAT-LAB. Technical report, Wake Forest University, Winston-Salem, NC, USA, 1996. URL http://

www.mthcsc.wfu.edu/pt/
pt.html. From http:/
/supertech.lcs.mit.edu/
~cly/survey.html#PTTO:

“The webpage has been scrapped. We are trying to contact the authors to obtain more information about the project.”.

Hunt:2001:GMB

[HLR01] Brian R. Hunt, Ronald L.

Lipsman, and Jonathan M. Rosenberg. A guide to MATLAB: for beginners and experienced users. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80380-2 (hardcover), 0-521-00859-X (paperback). xvii + 327 pp. LCCN QA297 .H86

2001. URL http://www.
cambridge.org/uk/catalogue/
email.asp?isbn=052100859X.

With Kevin R. Coombes, John E. Osborn, and Garrett J. Stuck.

Hunt:2006:GMB

[HLR06a] Brian R. Hunt, Ronald J.

Lipsman, and J. (Jonathan) Rosenberg. A guide to MATLAB: for beginners and experienced users. Cambridge University Press, Cambridge, UK, second edition, 2006. ISBN 0-521-85068-1, 0-521-61565-8 (paperback). xv + 311 pp. LCCN QA297 .H86 2006 UCD.

Hunt:2006:GM

[HLR06b] Brian R. Hunt, Ronald L.

Lipsman, and Jonathan M. Rosenberg. A guide to MATLAB(R). Cambridge University Press, Cambridge, UK, second edition, 2006. ISBN 0-521-61565-8. xvi + 311 pp. With Kevin R. Coombes, John E. Osborn, and Garrett J. Stuck, For beginners and experienced users, Updated for MATLAB 7 and Simulink 6.

Hunt:2014:GMB

[HLR14] Brian R. Hunt, Ronald L. Lipsman, and Jonathan M. (Jonathan Micah) Rosenberg. A guide to MAT-LAB: for beginners and experienced users: updated for MATLAB 8 and Simulink 8. Cambridge University Press, Cambridge, UK, third edition, 2014. ISBN 1-107-66222-2 (paperback). ???? pp. LCCN QA297 .H86 2014.

Huynh:2008:SSA

[HLS08] Huu Tue Huynh, Van Son Lai, and Issouf Soumaré. Stochastic simulation and applications in finance with MATLAB programs. The Wiley finance series. Wiley, New York, NY, USA, 2008. ISBN 0-470-72538-9 (hardcover). xvi + 338 pp. LCCN ???? US$60.00.

Hill:1988:ECM

[HM88] David R. Hill and Cleve B. Moler (consulting editor). Experiments in Computational Matrix Algebra. Random House, New York, NY, USA, 1988. ISBN 0-394-35678-0 (paperback). xii + 446 pp. LCCN QA188 .H55 1987.

Higham:2019:NAP

[HM19] Nicholas J. Higham and Theo Mary. A new approach to probabilistic rounding error analysis. SIAM Journal on Scientific Computing, 41(5):A2815–A2835, ???? 2019. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Higham:2022:AEM

[HM22] Nicholas J. Higham and Mantas Mikaitis. Anymatrix: an extensible MATLAB matrix collection. Numerical Algorithms, 90(3):1175– 1196, July 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

https://link.springer.
com/article/10.1007/s11075-
021-01226-2.

Hammarling:2013:ACS

[HMT13] Sven Hammarling, Christopher J. Munro, and Françoise Tisseur. An algorithm for the complete solution of quadratic eigenvalue problems. ACM Transactions on Mathematical Software, 39 (3):18:1–18:19, April 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Haldar:2001:FHS

[HNS^{+}01] M. Haldar, A. Nayak, N. Shenoy, A. Choudhary, and P. Banerjee. FPGA hardware synthesis from MATLAB. In IEEE, editor, Fourteenth International Conference on VLSI Design, Bangalore, India, 3–7 January 2001, pages 299–304. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN ???? LCCN ????

Hoffmann:1998:MSB

[Hof98] Josef Hoffmann. Matlab and Simulink: Beispielorientierte Einführung in die Simulation dynamischer Systeme (English: Simulation of Dynamics Systems with Matlab and Simulink). Studentlitteratur, Lund, Sweden, 1998. ISBN 3-8273-1077-6. ???? pp. LCCN ????

Hogben:2007:HLA

[Hog07] Leslie Hogben, editor. Handbook of Linear Algebra. Discrete Mathematics and its Applications (Boca Raton). Chapman and Hall/CRC, Boca Raton, FL, USA, 2007. ISBN 1-58488-510-6 (hardcover), 1-4200-1057-3 (e-book). xxx + 1370 pp. LCCN QA184.2 .H36 2007. URL

http://www.crcnetbase.

com/isbn/9781420010572;
http://www.crcnetbase.
com/isbn/9781584885108;
http://www.loc.gov/catdir/
enhancements/fy0647/2006045491-

d.html. Associate editors: Richard Brualdi, Anne Greenbaum and Roy Mathias.

Hohenester:2014:OMT

[Hoh14] Ulrich Hohenester. OCT-BEC — a Matlab toolbox for optimal quantum control of Bose–Einstein condensates. Computer Physics Communications, 185(1): 194–216, January 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465513003172.

Holton:2004:IDM

[Hol04] James R. Holton. An introduction to dynamic meteorology. Elsevier Academic Press, Amsterdam, The Netherlands, fourth edition, 2004. ISBN 0-12-354015-1. ???? pp. LCCN QC880 .H65 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els051/2004044072.
html;   http://www.loc.
gov/catdir/toc/els051/
2004044072.html.

Honey:1991:RTM

[Hon91] W. E. Honey. Robotics Toolbox for the Matlab matrix manipulation program. In Jamshidi and Eicker [JE91], pages 101– 110. ISBN ???? LCCN TJ210.3 .A58 1991.

Honey:1992:RTM

[Hon92] William Edward Honey. Robotics toolbox for the MATLAB matrix manipulation program. Master’s thesis, University of New Mexico, Albuquerque, NM, USA, 1992. viii + 80 pp.

Horova:1997:SSM

[Hor97] Ivana Horová, editor. Summer School MATLAB 94, 95 Proceedings, volume 5 of Folia Facultatis Scientiarium Naturalium Universitatis Masarykianae Brunensis. Mathematica. Masaryk University, Brno, 1997. ISBN 80-210-1517-9. iv + 209 pp. Held in Velké Karlovice August 27–31, 1994, and in Čeřínek, August 28–September 1, 1995.

Hornberger:1998:EPH

[Hor98] George M. Hornberger, editor. Elements of physical hydrology. The Johns Hopkins University Press, Baltimore, MD, USA, 1998. ISBN 0-8018-5856-9 , 0-8018-5857-7 (paperback). viii + 302 pp. LCCN GB661.2 .E44 1998. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/jhu051/97042533.html;

http://www.loc.gov/catdir/
description/jhu052/97042533.
html;   http://www.loc.
gov/catdir/toc/jhu051/
97042533.html.

Howitt:1991:MTH

[How91] J. Howitt. MATLAB toolbox for handling qualities assessment of flight control laws. IEE Conference Publication, 2(332):1251–1256, 1991. CODEN IECPB4. ISSN 0537-9987 (??invalid ISSN checksum??).

Howard:1995:VSP

[How95] I. Howard. Vibration signal processing using Matlab. Acoustics Australia, 23(1): 9–??, ???? 1995. CODEN ACAUEC. ISSN 0814-6039.

Howard:2015:BRU

[How15] James P. Howard II. Book review: Uncertainty Quantification and Stochastic Modeling with MATLAB, Eduardo Souza de Cursi and Rubens Sampaio, ISTE Press, London, 2015. ISBN 978-1-78548-004-1. 442 pp. USD 180 (H). Journal of Statistical Software, 67(BR-7):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v067b07;
http://www.jstatsoft.
org/index.php/jss/article/
view/v067b07/v67b07.pdf;
http://www.sciencedirect.
com/science/book/9781785480058.

Harada:2002:SSR

[HP02a] Hiroshi Harada and Ramjee Prasad. Simulation and software radio for mobile communications. The Artech House universal personal communications series. Artech House Inc., Boston, MA, USA, 2002. ISBN 1-58053-044-3. xx + 467 pp. LCCN TK5103.2 .H368 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy031/2002019677.html.

Hoppensteadt:2002:MSM

[HP02b] F. C. Hoppensteadt and Charles S. Peskin. Modeling and simulation in medicine and the life sciences, volume 10 of Texts in applied mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2002. ISBN 0-387-95072-9. xiv + 354 pp. LCCN QH323.5 .H67 2002.

Higham:2019:SLP

[HP19] Nicholas J. Higham and Srikara Pranesh. Simulating low precision floating-point arithmetic. SIAM Journal on Scientific Computing, 41 (5):C585–C602, ???? 2019. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Hu:2018:IMP

[HPK18] Weilong Hu, Yannis Pantazis, and Markos A. Katsoulakis. ISAP — MAT-LAB package for sensitivity analysis of high-dimensional stochastic chemical networks. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085i03;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v085i03/
v85i03.pdf.

Higham:2019:SMH

[HPZ19] Nicholas J. Higham, Srikara Pranesh, and Mawussi Zounon. Squeezing a matrix into half precision, with an application to solving linear systems. SIAM Journal on Scientific Computing, 41(4): A2536–A2551, ???? 2019. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Hill:1991:ADM

[HR91] David R. Hill and Lawrence C.

Rich. Automatic differentiation in MATLAB. Applied Numerical Mathematics: Transactions of IMACS, 1991. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). To appear.

Headrick:1996:SHO

[HR96] R. Headrick and J. Rowe. Solar and heliospheric observatory (SOHO) flight dynamics simulations using MATLAB. In Greatorex [Gre96], pages 111–124. ISBN ???? ISSN 0191-7811. LCCN ????

Hasegawa:2024:AQM

[HS24] Takemitsu Hasegawa and Hiroshi Sugiura. An automatic quadrature method for semi-infinite integrals of exponentially decaying functions and its Matlab code. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003941.

Hansen:2012:ATM

[HSH12] Per Christian Hansen and Maria Saxild-Hansen. AIR Tools — a MATLAB package of algebraic iterative reconstruction methods. Journal of Computational and Applied Mathematics, 236(8):2167–2178, February 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042711005188.

Huettel:2004:FMR

[HSM04] Scott A. Huettel, Allen W.

Song, and Gregory Mc-Carthy. Functional magnetic resonance imaging. Sinauer Associates, Publishers, Sunderland, MA, USA, 2004. ISBN 0-87893-288-7. xviii + 492 pp. LCCN RC386.6.M34 H84 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip0413/2004000500.
html.

Hoffbeck:2001:IMP

[HSR01] Joseph P. Hoffbeck, Mansoor Sarwar, and Eric J. Rix. Interfacing MAT-LAB with a parallel virtual processor for matrix algorithms. The Journal of Systems and Software, 56(1): 77–80, February 1, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/57/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/29/11/57/24/30/article.
pdf.

Hsu:1999:SEA

[Hsu99] David Y. Hsu. Spatial error analysis: a unified, application-oriented treatment. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7803-3453-1. xvii + 217 pp. LCCN TK153 .H78 1999. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley043/98012430.
html;   http://www.loc.
gov/catdir/description/
wiley036/98012430.html;
 http://www.loc.gov/catdir/
toc/onix07/98012430.html.

Hines:1997:MSF

[HT97] J. Wesley Hines and Lefteri H. Tsoukalas. MAT-LAB supplement to Fuzzy and neural approaches in engineering. Adaptive and learning systems for signal processing, communications, and control. Wiley, New York, NY, USA, 1997. ISBN 0-471-19247-

3. xii + 210 pp. LCCN QA76.87.H56 1997.

Hohenester:2012:MMT

[HT12] Ulrich Hohenester and Andreas Trügler. MNPBEM

— a Matlab toolbox for the simulation of plasmonic nanoparticles. Computer Physics Communications, 183(2):370–381, February 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465511003274.

Hodel:1996:OCS

[HTCI96] A. S. Hodel, R. B. Tenison, D. C. Clem, and J. E. Ingram. The Octave control systems toolbox: a MATLAB-like CACSD environment. In IEEE [IEE96d], pages 386–

391. ISBN 0-7803-3032-3, 0-7803-3033-1. LCCN TJ212.2.I32495 1996. IEEE catalog number 96TH8136.

Haario:1990:CSN

[HTJ90] H. Haario, V. M. Taavitsainen, and P. A. Jokinen. A chemometrics/statistics/ neural networks toolbox for MATLAB. Data handling in science and technology, 6(??):133–150, ???? 1990. CODEN DHSTEV. ISSN 0922-3487.

Hansen:2006:TFVa

[HTNFBS06a] Poul Martin Hansen, Iva Marija Tolić-Nørrelykke, Henrik Flyvbjerg, and Kirstine Berg-Sørensen. tweezercalib 2.0: Faster version of Matlab package for precise calibration of optical tweezers. Computer Physics Communications, 174(6):518–520, March 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465505006247.

Hansen:2006:TFVb

[HTNFBS06b] Poul Martin Hansen, Iva Marija Tolic-Nørrelykke, Henrik Flyvbjerg, and Kirstine Berg-Sørensen. tweezercalib 2.1: Faster version of Matlab package for precise calibration of optical tweezers. Computer Physics Communications, 175(8):572–573, October 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S001046550600289X.

Hull:1999:MMU

[Hul99] Douglas W. Hull. Mastering mechanics I using MAT-LAB 5. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-864034-3. xiv + 336 pp. LCCN TA345 .H85 1999.

Hung:1998:DMV

[Hun98] G. K. Hung. Dynamic model of the vergence eye movement system: Simulations using MATLAB/ SIMULINK. Computer Methods and Programs in Biomedicine, 55(1):59–??, ???? 1998. CODEN CMP-BEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Haykin:2002:SS

[HV02] Simon S. Haykin and Barry Van Veen. Signals and systems. Wiley, New York, NY, USA, second edition, 2002. ISBN 0-471-16474-7 (cloth), 0-471-37851-8 (WIE). xvii + 802 pp. LCCN TK5102.5 .H37 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley0310/
2002027040.html;   http:
//www.loc.gov/catdir/toc/
wiley023/2002027040.html.

Haykin:2003:SS

[HV03] Simon S. Haykin and Barry Van Veen. Signals and systems. Wiley, New York, NY, USA, second edition, 2003. ISBN 0-471-16474-7 (domestic), 0-471-37851-8 (WIE). xvii + 802 pp.

Hahn:2007:EME

[HV07] Brian Hahn and Dan Valentine. Essential MATLAB for Engineers and Scientists. Newnes, London, UK, third edition, 2007. ISBN 0-7506-8417-8 (paperback). 448 (est.) pp. LCCN ????

Huck:2015:STO

[HWB15] Alexander Hück, Johannes Willkomm, and Christian Bischof. Source transformation for the optimized utilization of the Matlab runtime system for automatic differentiation. In Mehl et al. [MBS15], pages 115–131. ISBN 3-319-22996-6, 3-319-22997-4 (ebook). LCCN QA71-90;

TA329. URL http://link.
springer.com/chapter/10.

                                        1007/978-3-319-22997-3_
                                        7/.

Huang:2015:COM

[HYY^{+}15] Kai Huang, Min Yu, Rongjie Yan, Xiaomeng Zhang, Xiaolang Yan, Lisane Brisolara, Ahmed Amine Jerraya, and Jiong Feng. Communication optimizations for multithreaded code generation from Simulink models. ACM Transactions on Embedded Computing Systems, 14(3):59:1– 59:??, May 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hill:1994:LAL

[HZ94] David R. Hill and David E. Zitarelli. Linear Algebra LABS with MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-02-354811-8. various pp. LCCN ????

Hill:1996:LAG

[HZ96a] D. R. Hill and D. E. Zitarelli. Linear algebra + geometry + MATLAB. In Anonymous [Ano96u], pages 214–218. ISBN 0-201-87020-7. LCCN ????

Hill:1996:LAL

[HZ96b] David R. Hill and David E. Zitarelli. Linear Algebra LABS with MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1996. ISBN 0-13-505439-7. LCCN ????

Iskander:1996:TBR

[I^{+}96] Magdy F. Iskander et al., editors. Technology-based re-engineering engineering education: proceedings of Frontiers in Education FIE’96 26th annual conference, November 6–9, 1996, Salt Lake City, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Islam:1997:IHF

[ICL97] S. Mofizul Islam, K. M.

Coates, and G. Ledwich. Identification of high frequency transformer equivalent circuit using Matlab from frequency domain data. In IEEE [IEE97c], pages 357–364. ISBN 0-7803-4068-X, 0-7803-4067-1, 0-7803-4069-8, 0-7803-4070-1. ISSN 0197-2618. LCCN TK5 .I19 1997. IEEE catalog number 97CH36096 (softbound), 97CB36096 (hardcover).

Ingalls:1996:PSC

[ICS96] V. Wayne Ingalls, Joseph Cynamon, and Annie V.

Saylor, editors. Proceedings of the 1996 Summer Computer Simulation Conference: July 21–25, 1996, Portland, Oregon. Society for Computer Simulation, San Diego, CA, USA, 1996. ISBN 1-56555-098-6. ISSN 0094-7474. LCCN ????

Irurozki:2018:APM

[ICS^{+}18] Ekhine Irurozki, Josu Ceberio, Josean Santamaria, Roberto Santana, and Alexander Mendiburu. Algorithm 989: perm mateda: a Matlab toolbox of estimation of distribution algorithms for permutationbased combinatorial optimization problems. ACM Transactions on Mathematical Software, 44(4):47:1– 47:13, August 2018. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3206429.

IEEE:1993:AWA

[IEE93a] IEEE, editor. Asia-Pacific Workshop on Advances in Motion Control: Proceedings, July 15–16, 1993, the Pan Pacific Hotel, Singapore. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1223-6. LCCN TJ212.2 .A75 1993.

IEEE:1993:PIC

[IEE93b] IEEE, editor. Proceedings of the 32nd IEEE Conference on Decision and Control, December 15–17, 1993, Marriott Rivercenter, San Antonio, Texas, USA, volume 3 of IEEE Conference on Decision and Control. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1299-6, 0-7803-1298-8, 0-7803-1300-3. ISSN 0888-3610. LCCN TJ 217 I11c. Four volumes.

IEEE:1994:WAT

[IEE94a] IEEE, editor. 1994 IEEE 4th Workshop on Computers in Power Electronics: Université du Québec à Trois-Rivières, August 7– 10, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-2091-3, 0-7803-2092-1. LCCN TK7881.15 .I35 1994.

IEEE:1994:SID

[IEE94b] IEEE, editor. 1994 Sixth IEEE Digital Signal Processing Workshop: October 2–5, 1995, Yosemite National Park, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1948-6, 0-7803-1949-4. LCCN TK5102.9 .I328 1994.

IEEE:1994:IPN

[IEE94c] IEEE, editor. ICIP-94: proceedings, November 13– 16, 1994, Austin Convention Center, Austin, Texas, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6952-7, 0-8186-6950-0, 0-8186-6951-9. LCCN TK8315.I22 1994. Three volumes.

IEEE:1994:PIC

[IEE94d] IEEE, editor. Proceedings of the 33rd IEEE Conference on Decision and Control, December 14–16, 1994, Buena Vista Palace at Walt Disney Resort, Lake Buena Vista, Florida, USA, volume 3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1969-9, 0-7803-1968-0, 0-7803-1970-2. ISSN 0888-3610. LCCN TJ 217 I11c 1994. Four volumes.

IEEE:1994:CAG

[IEE94e] IEEE, editor. Proceedings of the Third IEEE Conference on Control Applications, August 24th–26th, 1994, the University of Strathclyde, Glasgow, Scotland, UK, volume 3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1873-0, 0-7803-1872-2, 0-7803-1874-9. LCCN TJ212.2.I3247 1994. Three volumes.

IEEE:1994:PTI

[IEE94f] IEEE, editor. Proceedings of the Third IEEE Conference on Fuzzy Systems: IEEE World Congress on Computational Intelligence, June 26–June 29, 1994, Walt Disney World Dolphin Hotel, Orlando, Florida, volume 3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1896-X, 0-7803-1897-8, 0-7803-1898-6. LCCN TJ212.2.I3249 1994. Three volumes.

IEEE:1995:IAP

[IEE95a] IEEE, editor. IEEE Antennas and Propagation Society International Symposium: 1995 digest, June 18–June 23, 1995, Newport Beach, California, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2720-9. LCCN TK 7871.6 A2 1995. Four volumes. IEEE catalog number: 95CH35814.

IEEE:1995:PIC

[IEE95b] IEEE, editor. Proceedings of the 4th IEEE Conference on Control Applications: the Desmond Hotel, Albany, New York, September 28th– 29th, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2551-6, 0-7803-2550-8. LCCN TJ212.2.I3247 1995.

IEEE:1996:IPE

[IEE96a] IEEE, editor. IEEE Power Engineering Society: Winter meeting — January 1996, Baltimore, MD, volume 12(1) of IEEE TRANS-ACTIONS ON POWER DELIVERY PWRD 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? ISSN 0885-8977 (print), 1937-4208 (electronic). LCCN ????

IEEE:1996:PIIc

[IEE96b]  IEEE, editor.  Proceed-

ings, 1996 IEEE International Workshop on Variable Structure Systems, VSS ’96: December 5–6, 1996, Institute of Industrial Science, University of Tokyo, Roppongi, Tokyo, Japan, volume 19 of Seiken Symposium. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3718-2. LCCN TJ216 .I28 1996.

IEEE:1996:PIIb

[IEE96c] IEEE, editor. Proceedings of IPPS ’96: the 10th International Parallel Processing Symposium, April 15– 19, 1996, Honolulu Hawaii. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7255-2, 0-8186-7257-9. LCCN QA 76.58 I56 1996.

IEEE:1996:PIIa

[IEE96d]  IEEE, editor.  Proceed-

ings of the 1996 IEEE International Symposium on Computer-Aided Control System Design, September 15–18, 1996, Ritz-Carlton, Dearborn, Dearborn, Michigan, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3032-3, 0-7803-3033-1. LCCN TJ212.2.I32495 1996. IEEE catalog number 96TH8136.

IEEE:1996:PIC

[IEE96e] IEEE, editor. Proceedings of the 35th IEEE Conference on Decision and Control, December 11–13, 1996, Portopia Hotel and International Convention Center, Kobe, Japan, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3591-0, 0-7803-3590-2, 0-7803-3592-9, 0-7803-3593-7. ISSN 0191-2216. LCCN TJ 217 I11c 1996. Four volumes.

IEEE:1996:IIC

[IEE96f] IEEE, editor. The 1996 IEEE International Conference on Acoustics, Speech, and Signal Processing conference proceedings, May 7– 10, 1996, Marriott Marquis Hotel, Atlanta, Georgia, USA, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3193-1, 0-7803-3192-3, 0-7803-3194-X, 0-7803-3195-8. ISSN 0749-8411. LCCN TK 7882 S65 I16 1996. Six volumes.

IEEE:1997:IIC

[IEE97a] IEEE, editor. 1997 IEEE International Conference on Systems, Man, and Cybernetics: Hyatt Orlando, Orlando, Florida, USA, October 12–15, 1997: computational cybernetics and simulation, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4054-X, 0-7803-4053-1, 0-7803-4055-8. ISSN 1062-922X. LCCN Q300 .I485 1997. Five volumes.

IEEE:1997:IIE

[IEE97b] IEEE, editor. 1997 IEEE International Electric Machines and Drives Conference record: May 18–21, 1997, Hyatt Regency Hotel, Milwaukee, Wisconsin, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-3946-0, 0-7803-3947-9. LCCN TK4058.I4 1997. IEEE catalog number: 97TH8282.

IEEE:1997:ICR

[IEE97c] IEEE, editor. IAS ’97:

conference record of the 1997 IEEE Industry Applications Conference: Thirtysecond IAS Annual Meeting, October 5–9, 1997, Hyatt Regency Hotel, New Orleans, Louisiana, Conference Record- IEEE Industry Applications Society Annual Meeting 1997/v1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4068-X, 0-7803-4067-1, 0-7803-4069-8, 0-7803-4070-1. ISSN 0197-2618. LCCN TK5 .I19 1997. IEEE catalog number 97CH36096 (softbound), 97CB36096 (hardcover).

IEEE:1997:PIS

[IEE97d] IEEE, editor. Proceedings, IEEE Southeastcon ’97: engineering the new century, April 12–14, 1997, Virginia Tech, Blacksburg, Virginia, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-3845-6, 0-7803-3844-8, 0-7803-3846-4, 0-7803-3847-2. LCCN TK7801.I33 1997. IEEE Catalog number 97CH36044.

IEEE:1997:PII

[IEE97e] IEEE, editor. Proceedings of the 1997 IEEE International Conference on Control Applications: October 5–7, 1997, Sheraton Hartford Hotel, Hartford, Connecticut, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-3877-4, 0-7803-3876-6, 0-7803-3878-2. ISSN 1085-1992. LCCN TJ212.2.C565 1997. IEEE catalog number: 97CH36055.

IEEE:1998:PIM

[IEE98]  IEEE, editor.  Proceed-

ings: the Seventh IEEE International Symposium on High Performance Distributed Computing, July 28–31, 1998, Chicago, Illinois. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8579-4, 0-8186-8581-6 (microfiche). LCCN QA76.9.D5 I157 1998. IEEE Order Plan Catalog Number 98TB100244. IEEE Computer Society Press order number PR08579.

IEEE:2004:ICP

[IEE04] IEEE, editor. 2004 International Conference on Probabilistic Methods Applied to Power Systems, Ames, IA, USA, 12–16 September 2004. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-9761319-1-9 (paperback). LCCN TK1005 .I547 2004.

Ichikawa:1994:ACE

[IF94] A. Ichikawa and K. Furuta, editors. Advances in control education 1994 (ACE ’94): IFAC symposium, Tokyo, Japan, 1–2 August 1994, volume 3 of Advances in Control Education. Pergamon Press, New York, NY, USA, 1994. ISBN 0-08-042230-6. LCCN TJ212.2 .A393 1995.

Ifeachor:2005:DCH

[Ife05] Emmanuel Ifeachor. DSP (Companion Handbook): a Practical Guide for MAT-LAB and C Language Implementation of DSP Algorithms. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-043302-0. ???? pp. LCCN ????

Iskander:1996:ICS

[IGPF96] Magdy F. Iskander, Clark C. Guest, Jerry Place, and Suzanne Fitzgerald, editors. 1996 International Conference on Simulation and Multimedia in Engineering Education (ICSEE ’96), January 1996, San Diego, CA, volume 28(1) of Simulation series. Society for Computer Simulation, San Diego, CA, USA, 1996. ISBN 1-56555-084-6. ISSN 0735-9276. LCCN T65.5.C65 I583 1996.

Ito:2019:ABS

[IKK^{+}19] Naoki Ito, Sunyoung Kim, Masakazu Kojima, Akiko Takeda, and Kim-Chuan Toh. Algorithm 996: BBCPOP: A sparse doubly nonnegative relaxation of polynomial optimization problems with binary, box, and complementarity constraints. ACM Transactions on Mathematical Software, 45(3):34:1–34:16, July 2019. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3309988.

Ismail:2012:CMT

[IMM12] I. A. Ismail, G. S. Mokaddis, and Mariam K. Metry. Computing a matrix transpose of multithreading for queuing parallel in Matlab programming. International Journal of Computer Applications, 49(??): 41–47, July 2012. CODEN ???? ISSN 0975-8887. URL

https://www.ijcaonline.
org/archives/volume49/
number5/7627-0694/.

Inman:1994:EV

[Inm94] Daniel J. Inman. Engineering Vibration. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-951773-1. xv + 560 pp. LCCN TA355.I48 1994.

Inman:2001:EV

[Inm01] D. J. Inman. Engineering vibration. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2001. ISBN 0-13-726142-X. xiv + 621 pp. LCCN TA355 .I519 2001.

Iam-on:2010:LMP

[IoG10] Natthakan Iam-on and Simon Garrett. LinkCluE: a MATLAB package for linkbased cluster ensembles. Journal of Statistical Software, 36(9):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v36/
i09.

Ingle:1997:DSP

[IP97] Vinay K. Ingle and John G.

Proakis. Digital signal processing using MATLAB V.4. The PWS BookWare companion series. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1997. ISBN 0-534-93805-1. xii + 420 pp. LCCN TK5102.9.I535 1997.

Ingle:2000:DSP

[IP00] Vinay K. Ingle and John G. Proakis. Digital signal processing using MATLAB. BookWare companion series. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37174-4. xiv + 418 pp. LCCN TK5102.9 .I534 2000.

Ingle:2010:DSP

[IP10] Vinay K. Ingle and John G. Proakis. Digital signal processing with Matlab. Cengage Learning, Mason, OH, USA, third edition, 2010. ISBN 1-111-42737-2. ???? pp. LCCN ????

Innerberger:2023:MOO

[IP23] Michael Innerberger and Dirk Praetorius. MooAFEM: an object oriented Matlab code for higher-order adaptive FEM for (nonlinear) elliptic PDEs. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007998.

Ironside:2015:CPI

[Iro15] Charlie Ironside. Computational Photonics: An Introduction with MATLAB, by Marek S. Wartak, Scope: textbook. Level: undergraduate. Contemporary Physics, 56(1):88, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Irwin:2002:BEC

[Irw02] J. David Irwin. Basic engineering circuit analysis. Wiley, New York, NY, USA, seventh edition, 2002. ISBN 0-471-40740-2 (cloth). xxiii + 670 pp. LCCN TK454 .I78 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley035/2001023737.
html;   http://www.loc.
gov/catdir/toc/onix07/
2001023737.html.

Irwin:2005:BEC

[Irw05] J. David Irwin. Basic Engineering Circuit Analysis, with Computer Simulation Techniques For Excel, Matlab, and Pspice. Wiley, New York, NY, USA, 2005. ISBN 0-471-73106-4. ???? pp. LCCN ????

Ioannou:1996:RAC

[IS96] Petros A. (Petros A.) Ioannou and Jing Sun. Robust Adaptive Control. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-439100-4. xvii + 825 pp. LCCN TJ217 .I66 1996.

Ibhadode:2021:ITO

[IZBT21] Osezua Ibhadode, Zhidong Zhang, Ali Bonakdar, and Ehsan Toyserkani. IbIPP for topology optimization

— an image-based initialization and post-processing code written in MATLAB. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000467.

Johnson:1999:ODM

[JA99] Valen E. Johnson and James H. Albert. Ordinal data modeling. Statistics for social science and public policy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98718-5 (hardcover). x + 258 pp. LCCN HA29 .J588 1999. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://site.ebrary.com/
lib/ucsc/Doc?id=5006008.

Jackson:1989:DFS

[Jac89] Leland Jackson. Digital Filters and Signal Processing. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, second edition, 1989. ISBN 0-89838-276-9. xv + 410 pp. LCCN TK 7872 F5 J33 1989.

Jackson:1991:SST

[Jac91] Leland B. Jackson. Signals, Systems, and Transforms. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-09589-0. xiii + 482 pp. LCCN TK5102.5.J33 1991.

Jackson:1996:DFS

[Jac96] Leland B. Jackson. Digital Filters and Signal Processing, with MATLAB Exercises. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, third edition, 1996. ISBN 0-7923-9559-X. xiii + 502 pp. LCCN TK7872.F5 J33 1996.

Jacob:2004:AAC

[Jac04] J. Michael Jacob. Advanced AC circuits and electronics: principles and applications. Herrick and Jacob series. Thomson/Delmar Learning, Clifton Park, NY, USA, 2004. ISBN 0-7668-2330-X (hardcover). xvii + 455 pp. LCCN TK454.15.A48 J33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip042/2003007979.
html.

Jimenez:2020:SMT

[JAC20] Cristian Jimenez, Iván Amaya, and Rodrigo Cor-

rea. SpinUpFlowDescriptor:

a MATLAB toolbox for ferrofluids materials under moderate and high amplitude and frequency of magnetic rotating fields in a spin-up geometry. SoftwareX, 12(??):Article 100567, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711020302636.

Jaffe:2000:RSE

[Jaf00] Richard C. Jaffe. Random signals for engineers using MATLAB and Math-CAD. AIP series in modern acoustics and signal processing. American Institute of Physics, Woodbury, NY, USA, 2000. ISBN 0-387-98956-0. xv + 374 pp. LCCN TK5102.9 .J34 2000.

Jain:2009:BRB

[Jai09] Varinder Jain. Book review: Computational Statistics Handbook with MAT-LAB, by W. L. Martinez; A. R. Martinez. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):942– 943, October 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
20622568.

Janka:2002:SDM

[Jan02] Randall S. Janka. Specification and design methodology for real-time embedded systems. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 2002. ISBN 0-7923-7626-9. xxxi + 221 pp. LCCN TK7895.E42 J36 2002.

Joisha:2003:SAS

[JB03] Pramod G. Joisha and Prithviraj Banerjee. Static array storage optimization in MATLAB. ACM SIG-PLAN Notices, 38(5):258– 268, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joisha:2006:AAS

[JB06] Pramod G. Joisha and Prithviraj Banerjee. An algebraic array shape inference system for MAT-LAB(R). ACM Transactions on Programming Languages and Systems, 28(5): 848–907, September 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Joisha:2007:TSM

[JB07] Pramod G. Joisha and Prithviraj Banerjee. A translator system for the MATLAB language. Software—Practice and Experience, 37(5):535–578, April 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jacques:1996:MTF

[JCRS96] D. R. Jacques, R. A. Canfield, D. B. Ridgely, and M. S. Spillman. A MAT-LAB toolbox for fixed-order, mixed-norm control synthesis. IEEE Control Systems Magazine, 16(5):36–??, ???? 1996. CODEN ISMAD7. ISSN 0272-1708.

Jovanovic-Dolecek:2013:RSP

[JD13] Gordana Jovanovic-Dolecek. Random signals and processes primer with MAT-LAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-2385-6 (hardcover), 1-4614-2386-4 (e-book). xviii + 528 pp. LCCN QA273 .D65 2013; QA273 .J68 2013.

Jovanovic-Dolecek:2008:UMT

[JDFV08] Gordana Jovanovic-Dolecek and Alfonso Fernandez-Vazquez. Use of MATLAB in teaching the fundamentals of random variables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):46–51, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jamshidi:1991:TMF

[JE91] Mohammad Jamshidi and Patrick J. Eicker, editors.

Robotics and remote systems: proceedings of the Fourth ANS Topical Meeting on Robotics and Remote Systems. Dept. of Energy, Albuquerque Operations; Sandia National Laboratories; CAD Laboratory for Systems/Robotics, UNM, Albuquerque, NM, USA, 1991. ISBN ???? LCCN TJ210.3 .A58 1991.

Jantzen:1994:FIH

[JE94] J. Jantzen and B. Eliasson. Fuzzy interpolation of hydro power sales data in Simulink. In IEEE [IEE94f], pages 1857–1860. ISBN 0-7803-1896-X, 0-7803-1897-8, 0-7803-1898-6. LCCN TJ212.2.I3249 1994. Three volumes.

Jeffrey:2004:EEM

[Jef04] Alan Jeffrey. Essentials of engineering mathematics. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2004. ISBN 1-58488-489-4. vii + 882 pp. LCCN TA333 .J44 2004.

Jeffrey:2005:MES

[Jef05] Alan Jeffrey. Mathematics for engineers and scientists. Chapman and Hall/CRC, Boca Raton, FL, USA, sixth edition, 2005. ISBN 1-58488-488-6. xvi + 994 pp. LCCN QA37.3 .J44 2005.

Jeffrey:2008:PAM

[Jef08] David Jeffrey, editor. Proceedings of the 21st annual meeting of the International Symposium on Symbolic Computation, ISSAC 2008, July 20–23, 2008, Hagenberg, Austria. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-59593-904-0. LCCN ????

Jersky:2006:BRB

[Jer06] Brian Jersky. Book review: Exploratory Data Analysis with MATLAB, by Wendy L. Martinez; Angel R. Martinez. Journal of the American Statistical Association, 101(473): 399–400, March 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        30047483.

Jespers:2001:ICD

[Jes01] Paul G. Jespers. Integrated converters: D to A and A to D architectures, analysis and simulation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-856446-5. 259 (est.) pp. LCCN TK7887.6 .J47 2001.

Jose-Garcia:2023:CMB

[JGGF23] Adán José-García and Wilfrido Gómez-Flores. CVIK: a Matlab-based cluster validity index toolbox for automatic data clustering. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000559.

Jacquot:1996:TDC

[JH96] R. G. Jacquot and J. C.

Hamann. Teaching digital control and filtering using MATLAB and VisSim. In Iskander et al. [I^{+}96], pages 591–594. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-

4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Jacquot:1997:VPS

[JH97] Raymond G. Jacquot and J. C. Hamann. Visualization of PDE solutions using implicit methods and MAT-LAB. Computers in education journal, 7(3):2–5, July/ September 1997. CODEN CEJOE7. ISSN 1069-3769.

Jacquot:1997:TDF

[JHPK97] Raymond G. Jacquot, J. C.

Hamann, J. W. Pierre, and R. F. Kubichek. Teaching digital filter design using symbolic and numeric features of MATLAB. Computers in education journal, 7(1):8–11, January/March 1997. CODEN CEJOE7. ISSN 1069-3769.

Kobza:1997:QQS

[Jiř97] JiříKobza. Quadratic and quartic splines in MAT-LAB. In Summer Schools MATLAB 94, 95 Proceedings (Velké Karlovice, 1994; Čeřínek, 1995), volume 5 of Folia Fac. Sci. Natur. Univ. Masaryk. Brun. Math., pages 47–65. Masaryk Univ., Brno, 1997.

Junkins:1993:IDC

[JK93] John L. Junkins and Youdan Kim. Introduction to Dynamics and Control of Flexible Structures. AIAA education series. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024–2518, 1993. ISBN 1-56347-054-3. 452 (est.) pp. LCCN TL910 .J86 1993.

Jacobson:1992:ADD

[JKR92] P. Jacobson, B. Kågström, and M. Ränner. Algorithm development for distributed memory multicomputers using CONLAB. Scientific Programming, 1(??): 185–203, 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jensen:2001:RMD

[JL01] A. (Arne) Jensen and A. (Anders) La Cour-Harbo. Ripples in mathematics: the discrete wavelet transform. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-41662-5 (softcover). ix + 246 pp. LCCN QA403.3 .J46 2001.

Javor:1996:MSE

[JLM96] A. Javor, A. Lehmann, and I. Molnar, editors. Modelling and simulation 1996: ESM’96: June 2–6, 1996, Budapest University of Economic Sciences. Society for Computer Simulation, San Diego, CA, USA, 1996. ISBN 1-56555-097-8. LCCN ????

Johnson:1994:CM

[JM94] Stephen C. Johnson and Cleve Moler. Compiling Matlab. In USENIX Association [USE94], pages 119– 128. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

Johnson:2008:SIS

[JMD08] Kjell Johnson, Abhyuday Mandal, and Tan Ding. Software for implementing the sequential elimination of level combinations algorithm. Journal of Statistical Software, 25(6):1– 13, March 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v25/i06.

Jan:2019:MIC

[JNR19] Nazir Jan, Nasruminallah, and Rehmatullah. Multidimensional image compression through discrete wavelet transform in Matlab using Daubechies wavelet. International Journal of Computer Applications, 178(??):13–16, May 2019. CODEN ???? ISSN 0975-8887. URL

https://www.ijcaonline.
org/archives/volume178/
number8/30549-2019918778/

.

Johnson:1995:SSM

[Joh95] Conor D. Johnson, editor. Smart structures and materials 1995. Passive damping: 1–2 March 1995, San Diego, California, volume 2445 of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1995. ISBN 0-8194-1794-7. ISSN 0361-0748. LCCN TS510.S63 v.2445; TA355 .S528 1995.

Johnson:2011:EMS

[Joh11] Richard K. Johnson. The elements of MATLAB style. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-73258-1. ???? pp. LCCN QA76.73.M296 J64 2011.

Johnson:2018:AAE

[Joh18] Robert W. Johnson. Algorithm 988: AMGKQ: An efficient implementation of adaptive multivariate Gauss–Kronrod quadrature for simultaneous integrands in Octave/MATLAB. ACM Transactions on Mathematical Software, 44(3):36:1– 36:19, April 2018. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3157735.

Jones:1991:MTC

[Jon91] Karin Elizabeth Jones. A MATLAB toolbox of combinatorial algorithms. Master’s thesis, University of California, Santa Barbara, Santa Barbara, CA, USA, 1991. xi + 179 pp.

Jones:1995:AMS

[Jon95a] L. C. Jones. The application of the MATLAB/ SIMULINK analysis packages to the design of a discrete linear quadratic optimal missile autopilot. In Anonymous [Ano95b], pages 3–?? ISBN ???? LCCN ????

Jones:1995:AMA

[Jon95b] Llyr Campbell Jones. Application of the MAT-LAB/SIMULINK analysis packages to the design of a discrete linear quadratic optimal missile autopilot. IEE Colloquium (Digest), 014: 3/1–3/6, 1995. CODEN DCILDN. ISSN 0963-3308.

Jonsson:2001:EMA

[Jon01] Gudbjorn Freyr Jonsson. Eigenvalue methods for accurate solution of polynomial equations. PhD thesis, Cornell University, Ithaca, NY, USA,

2001.  117 pp.  URL

http://wwwlib.umi.com/
dissertations/fullcit/
9988142;   http://wwwlib.
umi.com/dissertations/
preview/9988142.

Jonasson:2009:AEV

[Jon09] Kristjan Jonasson. Algorithm 878: exact VARMA likelihood and its gradient for complete and incomplete data with Matlab. ACM Transactions on Mathematical Software, 35(1):Art. 6, 11, 2009. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Jones:2018:PMM

[Jon18] Pete Richard Jones. Myex: A MATLAB interface for the Tobii Eyex eye-tracker. Journal of Open Research Software, 6(1):16–??, April 20, 2018. CODEN ???? ISSN 2049-9647. URL

https://openresearchsoftware.
metajnl.com/articles/10.
5334/jors.196/.

Jorgl:1994:CSL

[Jor94]  H. P. Jorgl.  A com-

bined simulation and laboratory experiment using Simulink. In Ichikawa and Furuta [IF94], pages 241– 244. ISBN 0-08-042230-6. LCCN TJ212.2 .A393 1995.

Jiang:1999:QMG

[JR99] Houyuan Jiang and Daniel Ralph. QPECgen, a MAT-LAB generator for mathematical programs with quadratic objectives and affine variational inequality constraints. Computational Optimization and Applications, 13(1-3):25–59, 1999. CODEN CPPPEF. ISSN 0926-6003. Computational optimization—a tribute to Olvi Mangasarian, Part II.

Johnson:1998:ILA

[JRA98] Lee W. Johnson, R. Dean Riess, and Jimmy T. Arnold. Introduction to Linear Algebra. Addison-Wes-ley, Reading, MA, USA, fourth edition, 1998. ISBN 0-201-82416-7. various pp. LCCN QA184.J63 1997. Includes MATLAB projects at the end of each chapter.

Johnson:2002:ILA

[JRA02] Lee W. Johnson, R. Dean (Ronald Dean) Riess, and Jimmy T. (Jimmy Thomas) Arnold. Introduction to linear algebra. Addison-Wes-ley, Reading, MA, USA, fifth edition, 2002. ISBN 0-201-65859-3. ???? pp. LCCN QA184.2 .J63 2002.

Jaust:2018:FAG

[JRA^{+}18] Alexander Jaust, Balthasar Reuter, Vadym Aizinger, Jochen Schütz, and Peter Knabner. FESTUNG: a MATLAB/GNU Octave toolbox for the discontinuous Galerkin method. Part III: Hybridized discontinuous Galerkin (HDG) formulation. Computers and Mathematics with Applications, 75(12):4505–4533, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122118301895.

Jacques:1995:MTF

[JRCS95] David R. Jacques, D. Brett Ridgely, Robert A. Canfield, and Mark S. Spillman. A MATLAB toolbox for fixed order, mixednorm control synthesis. In IEEE [IEE95b], pages 470–

475. ISBN 0-7803-2551-6, 0-7803-2550-8. LCCN TJ212.2.I3247 1995.

Johnson:2004:TBC

[JS04] C. Richard Johnson and William A. Sethares. Telecommunication breakdown: concepts of communication transmitted via softwaredefined radio. Pearson Education, Upper Saddle River, NJ, USA, 2004. ISBN 0-13-143047-5. xiv + 379 pp. LCCN TK5103.48 .J65 2004.

Jesus:2024:LDA

[JS24] Joabe Jesus and Augusto Sampaio. Local deadlock analysis of Simulink models based on timed behavioural patterns and theorem proving. Science of Computer Programming, 236(??):??, September 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642324000364.

Janke:2020:PMP

[JSB20] T. Janke, R. Schwarze, and

K. Bauer. Part2Track:

a MATLAB package for double frame and time resolved Particle Tracking Velocimetry. SoftwareX, 11 (??):Article 100413, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711019303395.

Jang:1997:NFS

[JSM97] Jyh-Shing Roger Jang, Chuen-Tsai Sun, and Eiji Mizutani. Neuro-fuzzy and soft computing: a computational approach to learning and machine intelligence. MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-261066-3. xxvi + 614 pp. LCCN QA76.9.S63 J36 1997.

Jiao:2022:KEM

[JZW^{+}22] Shizhe Jiao, Zhenlin Zhang, Kai Wu, Lingyun Wan, Huanhuan Ma, Jielan Li, Sheng Chen, Xinming Qin, Jie Liu, Zijing Ding, Jinlong Yang, Yingzhou Li, Wei Hu, Lin Lin, and Chao Yang. KSSOLV 2.0: an efficient MAT-LAB toolbox for solving the Kohn–Sham equations with plane-wave basis set. Computer Physics Communications, 279(??):Article 108424, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465522001436.

Kisacanin:2002:LCS

[KA02] Branislav Kisačanin and Gyan C. Agarwal. Linear control systems: with solved problems and MAT-LAB examples. Kluwer Academic… Plenum Publishers, New York, NY, USA, 2002. ISBN 0-306-46743-7. xiii + 381 pp. LCCN TJ220 .K57 2002.

Kepner:2004:M

[KA04] Jeremy Kepner and Stan Ahalt. MatlabMPI. Journal of Parallel and Distributed Computing, 64(8):997–1005, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Koc:2009:MGB

[KA09] Savas Koç and Zafer Aydoğmus.

A MATLAB/GUI based fault simulation tool for power system education. Mathematical and Computational Applications, 14(3): 207–217, December 2009. CODEN ???? ISSN 2297-

8747. URL https://www.
mdpi.com/2297-8747/14/
3/207.

Klee:2011:SDS

[KA11] Harold Klee and Randal Allen. Simulation of dynamic systems with MAT-LAB and Simulink. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2011. ISBN 1-4398-3673-6. ???? pp. LCCN QA76.9.C65 K585 2011.

Korzec:2013:TSM

[KA13] M. D. Korzec and T. Ahnert. Time-stepping methods for the simulation of the self-assembly of nanocrystals in Matlab on a GPU. Journal of Computational Physics, 251(??):396– 413, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0021999113004063.

Kirkegaard:1997:STD

[KAB97] P. H. Kirkegaard, P. Andersen, and R. Brincker. Structural Time Domain Identification (STDI) toolbox for use with MATLAB. In Wicks et al. [W^{+}97], pages 896–900. ISBN 0-912053-53-

4. ISSN 1046-6770. LCCN TS510.S63 v.3089. Two volumes.

Kadlec:1995:MTB

[Kad95] J. Kadlec. MATLAB Transputer bridge. In Blanke and Söderström [BS95a], page 1627. ISBN 0-08-042225-X. LCCN QA402.S957 1995. Three volumes.

Kadlec:1997:PPA

[Kad97] J. Kadlec. Parallel processing on Alphas under MATLAB 5. In Plasil and Jeffery [PJ97], pages 440–447. ISBN 3-540-63774-5 (softcover). LCCN QA76.751.S62 1997.

Kahan:1998:MLN

[Kah98] W. Kahan. Matlab’s loss is nobody’s gain. Technical report, Department of Mathematics and Department of Electrical Engineering and Computer Science, University of California, Berkeley, Berkeley, CA, USA, August 23, 1998. 35 pp. URL http:

                                        //www.cs.berkeley.edu/
                                        ~wkahan/MxMulEps.pdf;

http://www.cs.berkeley.
edu/~wkahan/MxMulEps.ps.

Kahan:2002:MPI

[Kah02a] W. Kahan. Matlab programs for interpolation and extrapolation. World-Wide Web document, February 11, 2002. URL http:/

/www.cs.berkeley.edu/~
wkahan/Math128/XPOLATE.

Lecture notes for Math 128.

Kahan:2002:MIH

[Kah02b] W. Kahan. Matlab’s inverses of Hilbert matrices. World-Wide Web document, May 2, 2002. URL

http://www.cs.berkeley.
edu/~wkahan/MathH110/HilbMats.

pdf. Lecture notes for Math H110.

Kahan:2004:FMD

[Kah04] W. Kahan. Failure mode:

Do MATLAB’s lu(...),

inv(...), / and \ have a

failure mode? World-Wide Web document, February 17, 2004. URL http:/

/www.cs.berkeley.edu/~
wkahan/Math128/FailMode.

pdf. Lecture notes for Math 128.

Kalouptsidis:1997:SPS

[Kal97] N. Kalouptsidis. Signal processing systems: theory and design. Wiley series in telecommunications and signal processing. Wiley, New York, NY, USA, 1997. ISBN 0-471-11220-8. xxii + 840 pp. LCCN TK5102.9 .K34 1996. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley047/96012447.
                                        html;   http://www.loc.
                                        gov/catdir/description/
                                        wiley032/96012447.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley022/96012447.
                                        html.

Kao:1997:ISP

[Kao97] Edward P. C. Kao. An Introduction to Stochastic Processes. Duxbury Press, North Scituate, MA, USA, 1997. ISBN 0-534-25518-3. x + 438 pp. LCCN QA274 .K355 1997.

Kao:2009:MOO

[Kao09] Ming-Hung Kao. Multiobjective optimal experimental designs for ER-fMRI using MATLAB. Journal of Statistical Software, 30 (11):??, June 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v30/i11.

Kaplan:2004:ISC

[Kap04] Daniel Kaplan. An introduction to scientific computation and programming. Brooks/Cole–Thomson Learning, Belmont, CA, 2004. ISBN 0-534-38913-9. xiii + 546 pp. LCCN QA76.6 .K37 2004.

Karnofsky:1993:NNC

[Kar93] Ken Karnofsky. Neural networks and character recognition. Dr. Dobb’s Journal of Software Tools, 18(6):96, 98–100, 102–103, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Karris:2001:MBS

[Kar01a] Steven T. Karris. Mathematics for business, science, and technology. Orchard Publications, Fremont, CA, USA, 2001. ISBN 0-9709511-0-8. various pp. LCCN QA37.3 .K37 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/ipg051/2002284744.
html;   http://www.loc.
gov/catdir/description/
ipg051/2002284744.html.

Karris:2001:NAU

[Kar01b] Steven T. Karris. Numerical analysis using MAT-LAB and spreadsheets. Orchard Publications, Fremont, CA, USA, 2001. ISBN 0-9709511-1-6. various pp. LCCN QA297 .K37 2001.

Karris:2003:CAM

[Kar03a] Steven T. Karris. Circuit analysis I: with MATLAB applications. Orchard Publications, Fremont, CA, USA, 2003. ISBN 0-9709511-2-4 (paperback). various pp. LCCN TK7867 .K27 2003.

Karris:2003:CAI

[Kar03b] Steven T. Karris. Circuit analysis II: with MATLAB applications. Orchard Publications, Fremont, CA, USA, 2003. ISBN 0-9709511-5-9. ???? pp. LCCN TK454 .K38 2003.

Karris:2003:SSM

[Kar03c] Steven T. Karris. Signals and systems: with MAT-LAB applications. Orchard Publications, Fremont, CA, USA, 2003. ISBN 0-9709511-6-7. ???? pp. LCCN ????

Karris:2005:EDA

[Kar05] Steven T. Karris. Electronic devices and amplifier circuits with MATLAB applications. Orchard Publications, Fremont, CA, USA, 2005. vii + 629 pp. LCCN TK7871.2 .E54 2005.

Kafieh:2019:AMM

[KAR^{+}19] Rahele Kafieh, Zahra Amini, Hossein Rabbani, Bahareh Kaviani Baghbaderani, Bahareh Salafian, Fatemeh Mazaheri, and Marzieh Mokhtari. Automatic multifaceted Matlab package for analysis of ocular images (AMPAO). SoftwareX, 10 (??):Article 100339, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711019301852.

Kattan:2003:MGF

[Kat03] Peter Issa Kattan. MAT-LAB Guide to Finite Elements: An Interactive Approach. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-43874-2. x + 385 pp. LCCN TA347.F5 K38 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy034/2002030558.html.

Includes CD-ROM.

Katsikis:2009:MBR

[Kat09] Vasilios N. Katsikis. A Matlab-based rapid method for computing lattice-subspaces and vector sublattices of R^{n}: applications in portfolio insurance. Applied Mathematics and Computation, 215(3):961–972, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kay:1993:FSS

[Kay93] Steven M. Kay. Fundamentals of statistical signal processing. Prentice Hall signal processing series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-345711-7 (vol. 1), 0-13-504135-X (vol. 2). ???? pp. LCCN TK5102.5 .K379 1993.

Kay:2005:IPR

[Kay05] Steven Kay. Intuitive Probability and Random Processes Using MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-24157-4. ???? pp. LCCN ????

Kay:2006:IPR

[Kay06]  Steven M. Kay.  Intu-

itive probability and random processes using MAT-LAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-24157-4, 0-387-24158-2 (e-book). xviii + 833 pp. LCCN QA273 .K326 2006 UCB.

Kwon:1997:FEMb

[KB97] Young W. Kwon and Hyochoong Bang. The finite element method using MATLAB. CRC mechanical engineering series. Advanced topics in mechanical engineering series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1997. ISBN 0-8493-9653-0. 519 pp. LCCN TA347.F5K86 1997. US$79.95.

Kwon:2000:FEM

[KB00] Young W. Kwon and Hyochoong Bang. The finite element method using MATLAB. CRC Mechanical Engineering Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2000. ISBN 0-8493-0096-7. xiv + 607 pp. LCCN TA347.F5 K86 2000. With 1 CD-ROM (Windows 95, 98 or NT4.0).

Kaufman:1998:DAC

[KBKS98] Howard Kaufman, Izhak Bar-Kana, and Kenneth Sobel. Direct Adaptive Control Algorithms: Theory and Applications. Communications and control engineering series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1998. ISBN 0-387-94884-8. xxvi + 424 pp. LCCN TJ217 .K38 1998. URL ftp:/

/ftp.mathworks.com/pub/
books/kaufman/.

Kwon:1997:FEMa

[KBQ97] Young W. Kwon, Hyochoong Bang, and M. S. Qatu. Finite element method using MATLAB. Applied mechanics reviews, 50(3):B25, ???? 1997. CO-DEN AMREAD. ISSN 0003-6900.

Kundu:1995:ACM

[KC95] N. Kundu and C. M. Crane.

Applying H_{∞} control methods to wind turbines using MATLAB. International journal of ambient energy, 16(3):131–138, July 1995. CODEN IJAEDW. ISSN 0143-0750.

Khadilkar:2019:FIB

[KDAB19] Samrat P. Khadilkar, Sunil R. Das, Mansour H. Assaf, and Satyendra N. Biswas. Face identification based on discrete wavelet transform and neural networks. International Journal of Image and Graphics (IJIG), 19(4): ??, October 2019. ISSN 0219-4678. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0219467819500220.

Kearney:2017:PEM

[Kea17] Kelly A. Kearney. ecopath matlab: A Matlab-based implementation of the Ecopath food web algorithm. Journal of Open Source Software, 2 (9):64:1–64:2, January 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00064.

Kecman:2001:LSC

[Kec01] V. (Vojislav) Kecman. Learning and soft computing: support vector machines, neural networks, and fuzzy logic models. Complex adaptive systems. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-11255-8. xxxii + 541 pp. LCCN QA76.9.S63 K43 2001; QA 76.9 .S63K43 2001X ENGI. URL ftp://

                                        uiarchive.cso.uiuc.edu/

pub/etext/gutenberg/;
http://link.library.utoronto.
ca/eir/EIRdetail.cfm?Resources_
_ID=24153&T=F.

Kelley:1995:IML

[Kel95] C. T. Kelley. Iterative Methods for Linear and Nonlinear Equations, volume 16 of Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-352-8 (paperback). xiii + 165 pp. LCCN QA297.8 .K45 1995. URL http://

www.siam.org.

Kelley:1999:IMO

[Kel99] C. T. Kelley. Iterative methods for optimization. Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1999. ISBN 0-89871-433-8 (paperback). xv + 180 pp. LCCN QA402.5 .K44 1999.

Kelly:2000:FMV

[Kel00] S. Graham Kelly. Fundamentals of mechanical vibrations. McGraw-Hill, New York, NY, USA, second edition, 2000. ISBN 0-07-230092-2. xi + 629 pp. LCCN QA935 .K38 2000.

Kelley:2003:SNE

[Kel03] C. T. Kelley. Solving nonlinear equations with Newton’s method. Fundamentals of algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2003. ISBN 0-89871-546-6 (paperback). xiii + 104 pp. LCCN QA297.8 .K455 2003.

Kennedy:1995:NDI

[Ken95] Michael Peter Kennedy, editor. NDES ’95, Dublin Ireland 28–29 July, 1995: proceedings of the 3rd international specialist workshop on nonlinear dynamics of electronic systems. University College Dublin, Dublin, Ireland, 1995. ISBN 1-898473-24-2. LCCN ????

Kepner:2005:PPM

[Kep05]  Jeremy Kepner.  Par-

allel programming with MatlabMPI. World-Wide Web site., 2005. URL

                                        http://www.ll.mit.edu/
                                        MatlabMPI/.

Kepner:2009:PMM

[Kep09]  Jeremy Kepner.  Paral-

lel MATLAB for multicore and multinode computers, volume 21 of Software, Environments, and Tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2009. ISBN 0-89871-673-X. xxvi + 253 pp.

Keren:1995:MMC

[Ker95] Y. Keren. MATCOM: a MATLAB to C++ translator and support libraries.

Technical report, Israel Institute of Technology, Haifa, Israel, 1995.

Kamermans:1996:IPC

[KF96] H. Kamermans and K. Fennema, editors. Interfacing the past: computer applications and quantitative methods in archaeology CAA95: Annual conference: 23rd March 1995, Leiden, The Netherlands, volume 28 of Analecta Praehistorica Leidensia. Institute of Prehistory, University of Leiden, Leiden, The Netherlands, 1996. ISBN 90-73368-10-3. ISSN 0169-7447. LCCN GN700 .A62 v.28 v.1-2 (1996).

Kurz:1994:MMI

[KFG94] Mary Elizabeth Kurz and Emmanuel Fernandez-Gaucherand. MIST: MATLAB introductory statistical toolbox. Frontiers in Education Conference, ????(????):397–404, ???? 1994. CODEN PFECDR. ISSN 0190-5848.

Kharab:2001:INM

[KG01] Abdelwahab Kharab and Ronald B. Guenther. An Introduction to Numerical Methods: a MATLAB Approach. Chapman and Hall/CRC, Boca Raton, FL, USA, 2001. ISBN 1-58488-281-6. 440 (est.) pp. LCCN QA297 .K52 2001. US$69.95, UK£29.99.

Kuo:2003:ACS

[KG03] Benjamin C. Kuo and M. Farid Golnaraghi. Automatic control systems. Wiley, New York, NY, USA, eighth edition, 2003. ISBN 0-471-13476-7, 0-471-38148-9 (International ed.). xiii + 609 pp. LCCN TJ213 .K8354 2003. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley035/2002513209.
                                        html;   http://www.loc.
                                        gov/catdir/toc/wiley031/
                                        2002513209.html.

Kuo:2005:DSP

[KG05] Sen M. (Sen-Maw) Kuo and Woon-Seng Gan. Digital signal processors: architectures, implementations, and applications. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-035214-4. xviii + 601 pp. LCCN TK5102.9 K84 2005.

Kharab:2006:INM

[KG06] Abdelwahab Kharab and Ronald B. Guenther. An introduction to numerical methods: a MATLAB approach. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2006. ISBN 1-58488-557-2. 608 pp. LCCN QA297 .K52 2006. Includes CD-ROM.

Kharab:2012:INM

[KG12] Abdelwahab Kharab and Ronald B. Guenther. An introduction to numerical methods: a MATLAB approach. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2012. ISBN 1-4398-6899-9 (hardback), 1-4398-6900-6 (ebook). 14 + 567 pp. LCCN QA297 .K52 2012.

Kuo:1994:MTC

[KH94] Benjamin C. Kuo and Duane C. Hanselman. MAT-LAB Tools for Control System Analysis and Design. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-034646-2 (IBM PC version), 0-13-103045-0 (Apple Macintosh version). viii + 227 pp. LCCN TJ216 .K816b 1994.

Kmiecik:1996:TMC

[KH96] C. Kmiecik and R. Hartnett. Turbocharging MAT-LAB[R] for the classroom. In Iskander et al. [I^{+}96], pages 620–623. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Kamen:1997:FSS

[KH97a] Edward W. Kamen and Bonnie S. Heck. Fundamentals of signals and systems using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-02-361942-2. xiii + 688 pp. LCCN TK5102.9.K35 1997.

Kmiecik:1997:TMC

[KH97b] Chris Kmiecik and Richard J. Hartnett. Turbocharging MATLAB for the classroom. Technical report 02-97, U.S. Coast Guard Academy, Center for Advanced Studies, New London, CT, USA, March 1997. 4 pp.

Kolman:1997:ILA

[KH97c] Bernard Kolman and David R. Hill. Introductory linear algebra with applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, sixth edition, 1997. ISBN 0-13-266313-9. xvii + 608 + 70 + 8 pp. LCCN QA184.K67 1997.

Kamen:2000:FSS

[KH00] Edward W. Kamen and Bonnie S. Heck. Fundamentals of signals and systems using the Web and MAT-LAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2000. ISBN 0-13-017293-6. xiii + 722 pp. LCCN TK5102.9 .K3523 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy022/00267308.html.

Kumar:2014:MCM

[KH14] Vineet Kumar and Laurie Hendren. MIX10: compiling MATLAB to X10 for high performance. ACM SIG-PLAN Notices, 49(10):617– 636, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Khan:2005:DSP

[Kha05] Ashfaq A. Khan. Digital signal processing fundamentals. Da Vinci Engineering Press, Hingham, MA, USA, 2005. ISBN 1-58450-281-9. xxi + 391 pp. LCCN TK5102.9 .K3365 2005. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip0422/2004020495.
html.

Kheir:1996:SMC

[Khe96] Naim A. Kheir, editor. Systems Modeling and Computer Simulation, volume 94 of Electrical engineering and electronics. Marcel Dekker, New York, NY, USA, second edition, 1996. ISBN 0-8247-9421-4. xxii + 729 pp. LCCN QA76.9.C65 S975 1996.

Kolman:2005:ILA

[KHK05] Bernard Kolman, David R. (David Ross) Hill, and Bernard Kolman, Introductory linear algebra with applications. Introductory linear algebra: an applied first course. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, eighth edition, 2005. ISBN 0-13-143740-2. various pp. LCCN QA184.2 .K65 2005.

Khoo:2000:PCS

[Kho00] Michael C. K. Khoo. Physiological control systems: analysis, simulation, and estimation. IEEE Press series on biomedical engineering. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7803-3408-6. xvii + 319 pp. LCCN QP33.6.M36 K48 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley043/99027107.
html;   http://www.loc.
gov/catdir/description/
wiley036/99027107.html;
 http://www.loc.gov/catdir/
toc/onix07/99027107.html.

Kinoglu:1993:AIC

[Kin93] Fatih Kinoglu, editor. Computers in engineering, 1993: CAD/CAM/CAE, Computers in engineering education, Computational fluid mechanics, Computer graphics, Computational heat transfer, Design automation, Expert systems, Finite element analysis, Process control, Robotics: proceedings of the 1993 ASME International Computers in Engineering Conference and Exposition, August 8– 12,San Diego, California. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1993. ISBN 0-7918-1169-7. LCCN TA345.A86 1993.

King:1995:ETM

[Kin95] Joe King. The Engineer’s Toolkit: MATLAB Essentials. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1995. ISBN 0-8053-6442-0. LCCN ????

King:1997:SSA

[Kin97] R. W. King. Speech signal analysis, synthesis and recognition exercises using MATLAB. International Journal of Electrical Engineering Education, 34(2): 161–??, ???? 1997. CODEN IJEEAF. ISSN 0020-7209.

King:1998:ME

[Kin98] Joe King. Matlab for engineers. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-35094-7. iii + 124 pp. LCCN TA345 .K56 1998.

King:2001:MEH

[Kin01] Joe King. MATLAB 6 for engineers: hands-on tutorial. R. T. Edwards, Philadelphia, PA, USA, 2001. ISBN 1-930217-06-4. x + 308 pp. LCCN TA345 .K56 2001.

King:2006:MEH

[Kin06] Joe King. MATLAB 6 for engineers: hands-on tutorial. R. T. Edwards, Philadelphia, PA, USA, 2006. ISBN 1-930217-06-4. ???? pp. LCCN TA345 .K56 2006.

Kirk:1993:MSD

[Kir93] James R. Kirk. Modeling and simulation of a digital PID motor speed control system using MatLAB. CoED, 3(4):9–16, October/ December 1993. CODEN CWLJDP. ISSN 0736-8607. Kiusalaas:2005:NME

[Kiu05] Jaan Kiusalaas. Numerical Methods in Engineering with Matlab. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-85288-9 (hardback). viii + 426 pp. LCCN TA345 .K58 2005.

Kiusalaas:2010:NMEa

[Kiu10a] Jaan Kiusalaas. Numerical methods in engineering with MATLAB. Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-19133-5 (hardback). x + 431 pp. LCCN TA345 .K58 2010. Kiusalaas:2010:NMEb

[Kiu10b] Jaan Kiusalaas. Numerical methods in engineering with Python. Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-19132-7 (hardback). x + 422 pp. LCCN TA345 .K584 2010.

Kroller:1993:MSP

[KK93] M. Kroller and K. Kunisch.

MATLAB-software for parameter estimation in twopoint boundary value problems. In Identification and control in systems governed by partial differential equations (South Hadley, MA, 1992), pages 59–68. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993.

Karjalainen:1996:DAH

[KK96] E. J. Karjalainen and U. P.

Karjalainen. Data Analysis for Hyphenated Techniques. Elsevier, Amsterdam, The Netherlands, 1996. ISBN 0-444-82237-2. 475 pp. LCCN QD95.5.S72K37 1996. A CD-ROM containing MAT-LAB M-files and data files is available from the publisher.

Kobza:1997:STM

[KK97] J. Kobza and R. Kučera.

Spline toolbox in MAT-LAB. In Summer Schools MATLAB 94, 95 Proceedings (Velké Karlovice, 1994; Čeřínek, 1995), volume 5 of Folia Fac. Sci. Natur. Univ. Masaryk. Brun. Math., pages 25–46. Masaryk Univ., Brno, 1997.

Kammeyer:2001:MNG

[KK01a] Karl-Dirk Kammeyer and Volker Kühn. MATLAB in der Nachrichtentechnik. (German) [MATLAB in Communications Technology]. Schlembach, ????, 2001. ISBN 3-935340-05-2. xv + 560 pp. LCCN ????

Kehtarnavaz:2001:DSD

[KK01b] Nasser Kehtarnavaz and Mansour Keramat. DSP system design: using the TMS320C6000. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-091031-7. xii + 273 pp.

Kleinfeld:2001:ULA

[KK01c] Erwin Kleinfeld and Margaret Kleinfeld. Understanding Linear Algebra Using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-060945-5. vii + 210 pp. LCCN QA184.2 .K54 2001.

Kuo:2001:RTD

[KL01] Sen M. (Sen-Maw) Kuo and Bob H. Lee. Real-time digital signal processing: implementations, applications, and experiments with the TMS320C55X. Wiley, New York, NY, USA, 2001. ISBN 0-470-84137-0. xvi + 496 pp. LCCN TK5102.9 .K86 2001. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;

http://www.loc.gov/catdir/
description/wiley0310/
2001026651.;   http://
www.loc.gov/catdir/toc/
wiley021/2001026651.; UCLA.

Klatt:1992:TAM

[Kla92] G. Klatt. Teaching about matrices with MATLAB. In Lum [Lum92], pages 846–

853. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Klee:2007:SDS

[Kle07] Harold Klee. Simulation of dynamic systems with MAT-LAB(R) and Simulink(R). CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2007. ISBN 1-4200-4418-4. xii + 762 pp. With a foreword by Chris Bauer and Chris Schwarz.

King:2010:NSM

[KM10] Michael R. King and Nipa A. Mody. Numerical and statistical methods for bioengineering: applications in MATLAB. Cambridge texts in biomedical engineering. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-87158-1. ???? pp. LCCN R857.M34 K56 2010. URL http:

//assets.cambridge.org/
97805218/71587/cover/9780521871587.
jpg.

King:2011:NSM

[KM11] Michael R. King and Nipa A. Mody. Numerical and statistical methods for bioengineering: applications in MATLAB. Cambridge Texts in Biomedical Engineering. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-87158-1 (hardback). xii + 581 pp. LCCN R857.M34 K56 2011; R857.M34 K56X 2011 (LC). Applications in MATLAB.

Kenneweg:2024:QMT

[KMBP24] Tristan Kenneweg, Stefan Mueller, Tobias Brixner, and Walter Pfeiffer. QDT

— a Matlab toolbox for the simulation of coupled quantum systems and coherent multidimensional spectroscopy. Computer Physics Communications, 296(??):Article 109031, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465523003764.

Koot:2023:PVP

[KMLP^{+}23] Paul Koot, Miguel Angel Mendoza-Lugo, Dominik Paprotny, Oswaldo Morales-Nápoles, Elisa Ragno, and Daniël T. H. Worm. PyBanshee version (1.0): a Python implementation of the MATLAB toolbox BANSHEE for Non-Parametric Bayesian Networks with updated features. SoftwareX, 21(??):

??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711022001972.

Kadlec:1995:ABM

[KN95] J. Kadlec and N. Nakhaee.

Alpha bridge for MATLAB

4. In Cook [Coo95], pages 175–189. ISBN 90-5199-235-1, 4-274-90062-2. ISSN 1383-7575. LCCN ????

Knight:2000:BMB

[Kni00] Andrew (Andrew James)

Knight. Basics of MATLAB and beyond. Chapman and Hall/CRC, Boca Raton, FL, USA, 2000. ISBN 0-8493-2039-9. xii + 202 pp. LCCN TA345 .K63 2000.

Kruse:1997:NDT

[KNNM97] R. Kruse, D. Nauck, A. Nuernberger, and L. Merz. A neuro-fuzzy development tool for fuzzy controllers under MATLAB/SIMULINK. In Anonymous [Ano97b], pages 1029–1033. ISBN 3-89653-200-6. LCCN ????

Knobel:2000:IMT

[Kno00] Roger Knobel. An introduction to the mathematical theory of waves, volume 3 of Student mathematical library, 1520-9121 ; IAS/Park City mathematical subseries. American Mathematical Society, Providence, RI, USA, 2000.

ISBN 0-8218-2039-7 (softcover). xiv + 196 pp. LCCN QA927 .K693 2000.

Kolda:2000:ACU

[KO00] Tamara G. Kolda and Dianne P. O’Leary. Algorithm 805: computation and uses of the semidiscrete matrix decomposition. ACM Transactions on Mathematical Software, 26(3):415–435, September 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Koch:2014:AMH

[Koc14]  Inge Koch.  Analysis

of Multivariate and High-Dimensional Data. Cambridge series in statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-88793-3 (hardcover). xxv + 504 pp. LCCN QA278 .K5935 2014. URL http:

//assets.cambridge.org/
97805218/87939/cover/9780521887939.
jpg.

Kogketsof:1997:MAR

[Kog97] Georgios Kogketsof. Matlab applications to robot manipulators. Thesis (M.S.E.E.), Department of Electrical Engineering, University of Toledo, Toledo, OH, 1997. x + 182 pp.

Koko:2007:VMC

[Kok07] Jonas Koko. Vectorized Matlab codes for linear twodimensional elasticity. Scientific Programming, 15(3): 157–172, ???? 2007. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Koko:2015:MMG

[Kok15]  Jonas Koko.  A Mat-

lab mesh generator for the two-dimensional finite element method. Applied Mathematics and Computation, 250(??):650–664, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300314015148.

Kolman:1993:ILA

[Kol93] Bernard Kolman. Introductory Linear Algebra with Applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fifth edition, 1993. ISBN 0-02-366032-

5. xx + 619 pp. LCCN QA184.K67 1993.

Kittler:2004:PIC

[KPN^{+}04] Josef Kittler, Maria Petrou, Mark S. Nixon, Edwin R. Hancock, et al., editors. Proceedings of the 17th International Conference on Pattern Recognition: ICPR 2004: August 23–26, 2004, Cambridge UK, volume 4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2128-2. ISSN 1051-4651. LCCN Q327; Internet. IEEE Computer Society Order Number P2128.

Kollar:1995:FDS

[KPS95] I. Kollar, R. Pintelon, and J. Schoukens. Frequency domain system identification toolbox for MAT-LAB: a complex application example. In Blanke and Söderström [BS95a], pages 1619–1624. ISBN 0-08-042225-X. LCCN QA402.S957 1995. Three volumes.

Kastner:2022:AAA

[KR22] Felix Kastner and Andreas Rößler. An analysis of approximation algorithms for iterated stochastic integrals and a Julia and MATLAB simulation toolbox. arXiv.org, January 20,

2022. URL http://arxiv.
org/abs/2201.08424.

Kastner:2023:AAA

[KR23] Felix Kastner and Andreas Rößler. An analysis of approximation algorithms for iterated stochastic integrals and a Julia and Matlab simulation toolbox. Numerical Algorithms, 93(1): 27–66, May 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

https://link.springer.
com/article/10.1007/s11075-
022-01401-z.

Krol:1995:UMS

[Kro95] O. S. Krol. Using the MATLAB software in modeling the longitudinal turning processes. Russian engineering research, 15(4):80– ??, ???? 1995. ISSN 1068-798X.

Konkova:1994:LMF

[KS94] T. Ya. Kon′kova and V. N.

Simonova. The library of MATLAB functions for polynomial matrices. Zapiski naucnyh seminarov Leningradskogo otdelenija ordena Lenina Matematiceskogo instituta im. V.A. Steklova Akademii nauk SSSR, 219(Chisl. Metody i Voprosy Organ. Vychisl. 10):53–80, 221–222, 1994. CODEN ZNSLAF. ISSN 0373-2703.

Kuesters:1997:MPL

[KS97] U. Kuesters and J. P.

Steffen. Matrix programming languages for statistical computing: a comparison of GAUSS, MATLAB and Ox. In Bandilla and Faulbaum [BF97], pages 349–362. ISBN 3-8282-0032-X. LCCN QA276.4.K66

Kierzenka:2001:BSB

[KS01] Jacek Kierzenka and Lawrence F. Shampine. A BVP solver based on residual control and the MATLAB PSE. ACM Transactions on Mathematical Software, 27(3):299–316, September 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kythe:2005:HCM

[KS05] Prem K. Kythe and Michael R. Schäferkotter. Handbook of computational methods for integration. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-428-2. xxii + 598 pp. LCCN QA299.3 .K98 2005.

Koga:1994:CMM

[KSF94] Masanobu Koga, Mitsuji Sampei, and Katsuhisa Furuta. A compiler of MAT-LAB to MATX: Compiling and linking of M-files to an executable program. In Mattsson et al. [MGC94], pages 137–142. ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Krauss:1993:SPT

[KSL93] Thomas P. Krauss, Loren Shure, and John Little. Signal Processing Toolbox for Use with MATLAB. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1993.

ISBN ???? various pp. LCCN ????

Kaloko:2011:DDS

[KSP11] Bambang Sri Kaloko, Soebagio, and Mauridhi Hery Purnomo. Design and development of small electric vehicle using MAT-LAB/Simulink. International Journal of Computer Applications, 24(??):19–23, June 2011. CODEN ???? ISSN 0975-8887. URL

https://www.ijcaonline.
org/archives/volume24/
number6/2960-3940/.

Klima:2007:AAA

[KSS07] Richard E. Klima, Neil P.

Sigmon, and Ernest L. Stitzinger. Applications of abstract algebra with Maple and MATLAB. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2007. ISBN 1-58488-610-2, 1-4200-1119-7. xii + 505 pp. LCCN QA162 .K65 2007. Includes CD-ROM.

Kugiumtzis:2010:MAT

[KT10] Dimitris Kugiumtzis and Alkiviadis Tsimpiris. Measures of analysis of time series (MATS): a MATLAB toolkit for computation of multiple measures on time series data bases. Journal of Statistical Software, 33 (5):??, February 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v33/i05.

Kressner:2014:AHM

[KT14] Daniel Kressner and Christine Tobler. Algorithm 941: htucker — a Matlab toolbox for tensors in hierarchical Tucker format. ACM Transactions on Mathematical Software, 40(3):22:1– 22:22, April 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kubichek:1995:UMS

[Kub95] R. F. Kubichek. Using Matlab in a speech signal processing class. Computers in education journal, 5(3):2– ??, ???? 1995. CODEN CE-JOE7. ISSN 1069-3769.

Kulkarni:1999:MAD

[Kul99] Vidyadhar G. Kulkarni. Modeling, analysis, design, and control of stochastic system. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98725-8 (hardcover). xiv + 374 pp. LCCN QA274 .K84 1999.

Kumar:2005:DSP

[Kum05] B. Preetham Kumar. Digital signal processing laboratory. Taylor and Francis, Boca Raton, FL, USA, 2005. ISBN 0-8493-2784-9. 257 (est.) pp. LCCN TK5102.9 .K835 2005.

Kumar:2014:PAL

[Kum14] Hitesh Kumar. Performance analysis of LTE and IEEE 802.16 WiMAX in terms of attenuation and BER via MATLAB; scenario based ROF amalgamation as backhaul technology. International Journal of Computer Applications, 97(??): 32–36, July 2014. CODEN ???? ISSN 0975-8887. URL

https://www.ijcaonline.
org/archives/volume97/
number20/17127-7787/.

Kuncicky:2004:MP

[Kun04] David C. Kuncicky. Matlab programming. ESource– the Prentice Hall engineering source. Pearson Education, Upper Saddle River, NJ, USA, 2004. ISBN 0-13-035127-X. xviii + 267 pp. LCCN QA297 .K768 2004.

Kuo:1995:ACS

[Kuo95] Benjamin C. Kuo. Automatic Control Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, seventh edition, 1995. ISBN 0-13-304759-8. xxii + 897 pp. LCCN TJ213 .K8354 1995.

Kurzweil:2000:IDC

[Kur00] Jack Kurzweil. An introduction to digital communications. Wiley, New York, NY, USA, 2000. ISBN 0-471-15772-4. xv + 535 pp. LCCN TK5103.7 .K87 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley034/99027789.
html;   http://www.loc.
gov/catdir/toc/onix05/
99027789.html.

Kushler:2002:BRB

[Kus02] Robert H. Kushler. Book review: Computational Statistics Handbook with MAT-LAB(R) by Wendy L. Martinez; Angel R. Martinez. Technometrics, 44(4):405– 406, November 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
1271563.

Kushler:2006:BRB

[Kus06] Robert H. Kushler. Book review: Exploratory Data Analysis with MATLAB(R) by Wendy L. Martinez; Angel R. Martinez. Technometrics, 48(2):311, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/
25471179.

Kluska:1995:AGF

[KW95] J. Kluska and K. Wiktorowicz. Automatic generating of fuzzy control rules for a fuzzy logic controller using Matlab. In Bubnicki [Bub95], pages 434– 443. ISBN 83-7085-152-5. LCCN TA168.I526 1995. Three volumes.

Klimke:2005:ASP

[KW05] Andreas Klimke and Barbara Wohlmuth. Algorithm 847: Spinterp: piecewise multilinear hierarchical sparse grid interpolation in MATLAB. ACM Transactions on Mathematical Software, 31(4):561–579, December 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kwong:1993:MNF

[Kwo93] M. K. Kwong. MATLAB 4.0: New features, a few upward incompatibilities (review). SIAM News, 26 (2):??, March 1993. ISSN 0036-1437. URL ftp:/

/info.mcs.anl.gov/pub/
tech_reports/reports/P343.
ps.Z.

Kwatny:1995:LBA

[KYN95] Harry G. Kwatny, Xiao-Ming Yu, and Chika Nwankpa. Local bifurcation analysis of power systems using MAT-LAB. In IEEE [IEE95b], pages 57–62. ISBN 0-7803-2551-6, 0-7803-2550-8. LCCN TJ212.2.I3247 1995.

Krzhizhanovskaya:2020:CSI

[KZL^{+}20] Valeria V. Krzhizhanovskaya, Gábor Závodszky, Michael H. Lees, Jack J. Dongarra, Peter M. A. Sloot, Sérgio Brissos, and João Teixeira, editors. Computational Science — ICCS 2020 20th International Conference, Amsterdam, The Netherlands, June 3–5, 2020, Proceedings, Part II, volume 12138 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020. ISBN 3-030-50416-6, 3-030-50417-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). URL

https://link.springer.
com/book/10.1007/978-3-
030-50417-5.

Lau:2001:MDH

[LA01] Daniel L. Lau and Gonzalo R. Arce. Modern digital halftoning, volume 8 of Signal processing and communications. Marcel Dekker, New York, NY, USA, 2001. ISBN 0-8247-0456-8. xiii + 431 + 16 pp. LCCN TA1637 .L38 2001.

Lai:2004:PDS

[Lai04]  Edmund Lai.  Practical

digital signal processing for engineers and technicians. Newnes, London, UK, 2004. ISBN 0-7506-5798-7. ix + 289 pp. LCCN TK5102.5 .L24 2004.

Lakshmikantham:1997:NAW

[Lak97] V. Lakshmikantham, editor. Nonlinear analysts:

World congress; 2nd — July 1996, Athens, volume 30(4) of Nonlinear Analysis Theory Methods and Applications. Elsevier Science Publishers, Amsterdam, The Netherlands, 1997. ISBN 0-08-042032-X, 0-08-042036-2. ISSN 0362-546X (print), 1873-5215 (electronic). LCCN ????

Lambert:1995:DPM

[Lam95] D. J. Lambert. Dynamic P-method modeling of one and two dimensional systems using MATLAB. In DeMichele et al. [D^{+}95], pages 72–76. ISBN 0-912053-48-8. ISSN 1046-6770. LCCN TS510.S63 v.2460; TA 654.15 I58 1995. Two volumes.

Lan:2000:TCS

[Lan00]  Tian Lan.  TIA IS-95

CDMA signal generation using MATLAB. Thesis (M.S.), Electrical and Computer Engineering, University of California, Davis, Davis, CA, USA, 2000.

Lathi:1992:LSS

[Lat92] B. P. Lathi. Linear Systems and Signals. Berkeley-Cambridge Press, Carmichael, CA, USA, 1992. ISBN 0-941413-34-9. xiv + 656 pp. LCCN TK5102.5 .L29 1992.

Lathi:2005:LSS

[Lat05] B. P. (Bhagwandas Pannalal) Lathi. Linear systems and signals. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2005. ISBN 0-19-515833-4. xvi + 975 pp. LCCN TK5102.5 L298 2005.

Lawson:1996:LA

[Law96a] Terry Lawson. Linear Algebra. Wiley, New York, NY, USA, 1996. ISBN 0-471-30897-8 (cloth). xvi + 408 pp. LCCN QA184 .L395 1996. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley033/95049808.
                                        html;   http://www.loc.
                                        gov/catdir/toc/onix04/
                                        95049808.html.

Lawson:1996:MLL

[Law96b]  Terry Lawson.  MAT-

LAB Labs for Linear Algebra. Wiley, New York, NY, USA, 1996. ISBN 0-471-14953-5. ???? pp. LCCN ???? URL ftp:/

                                        /ftp.mathworks.com/pub/
                                        books/lawson/. To accom-

pany [LE96].

Law:2005:ISC

[Law05] Alan Law. Introduction To Scientific Computing Using Matlab. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-033369-7. ???? pp. LCCN ????

Lay:1994:LAA

[Lay94] David C. Lay. Linear Algebra and Its Applications. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-52031-1. various pp. LCCN QA184.L397 1994.

Lay:1997:LAA

[Lay97] David C. Lay. Linear algebra and its applications. Addison-Wesley, Reading, MA, USA, second edition, 1997. ISBN 0-201-82478-7, 0-201-82479-5 (Instructor’s edition), 0-201-82477-9 (study guide). various pp. LCCN QA184.L397 1997.

Lay:2003:LAA

[Lay03] David C. Lay. Linear algebra and its applications: Study guide. Addison-Wes-ley, Reading, MA, USA, third edition, 2003. ISBN 0-201-77013-X. x + 300 (est.) pp. LCCN ????

Lay:2006:LAA

[Lay06]  David C. Lay.  Linear

algebra and it’s applications. Pearson/Addison-Wesley, Boston, MA, USA, third edition, 2006. ISBN 0-321-28713-4. ???? pp. LCCN QA184.2 .L39 2006.

Lind:1999:RAS

[LB99] Rick Lind and Marty Brenner. Robust aeroservoelastic stability analysis: flight test applications. Advances in industrial control. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 1-85233-096-1. xv + 204 pp. LCCN TL574.S7 L47 1999.

Ludwig:2000:RCD

[LB00] Reinhold Ludwig and Pavel Bretchko. RF circuit design: theory and applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-095323-7. xiv + 642 pp. LCCN TK6553 .L823 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy032/00269579.html.

Liggett:1998:FMH

[LC98] James A. Liggett and D. A. (David A.) Caughey. Fluid mechanics: an interactive text. ASCE Press, Reston, VA, USA, 1998. ISBN 0-7844-0310-4. LCCN QA901 .L54 1998.

Li:2009:CPD

[LC09] Jichun Li and Yi-Tung Chen. Computational partial differential equations using MATLAB(R). Chapman & Hall/CRC Applied Mathematics and Nonlinear Science Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-8904-8. xiv + 364 pp. With 1 CD-ROM (Windows, Macintosh and UNIX).

Liu:2005:LST

[LCL05] Xinmin Liu, Ben M. Chen, and Zongli Lin. Linear systems toolkit in Matlab: structural decompositions and their applications. Journal of Control Theory and Applications, 3(3):287– 294, 2005. ISSN 1672-6340.

Lourenco:2022:ASL

[LCMCD22] Christopher Lourenco, Jinhao Chen, Erick Moreno-Centeno, and Timothy A. Davis. Algorithm 1021: SPEX left LU, exactly solving sparse linear systems via a sparse left-looking integerpreserving LU factorization. ACM Transactions on Mathematical Software, 48 (2):20:1–20:23, June 2022. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
10.1145/3519024.

Larson:1996:ELA

[LE96] Roland E. Larson and Bruce H. Edwards. Elementary Linear Algebra. D. C. Heath and Company, Lexington, MA, USA, third edition, 1996. ISBN 0-669-39641-9 (student edition), 0-669-39917-5 (international edition). xviii + 568 + 92 pp. LCCN QA184 .L39 1996. See also lab manual [Law96b].

Lurie:2000:CFC

[LE00] B. J. Lurie and Paul J. Enright. Classical feedback control with MATLAB, volume 6 of Control engineering. Marcel Dekker, New York, NY, USA, 2000. ISBN 0-8247-0370-7. xvi + 456 pp. LCCN TJ216 .L865 2000.

Leader:2004:NAS

[Lea04] Jeffery J. Leader. Numerical analysis and scientific computation. Pearson/Addison Wesley, Boston, MA, USA, 2004. ISBN 0-201-73499-0. xiii + 590 pp. LCCN QA297 .L384 2004.

Ledin:2001:SEB

[Led01] Jim Ledin. Simulation engineering: [build better embedded systems faster]. R and D developer series. CMP Books, Lawrence, KS, USA, 2001. ISBN 1-57820-080-6 (paperback). x + 303 pp. LCCN TK7895.E42 L44 2001.

Ledin:2004:ECS

[Led04] Jim Ledin. Embedded control systems in C/C++: an introduction for software developers using MAT-LAB. CMP Books, San Francisco, CA, USA, 2004. ISBN 1-57820-127-6. LCCN QA76.9.S88 L435 2004b.

                                        URL ftp://uiarchive.
                                        cso.uiuc.edu/pub/etext/

gutenberg/;   http://link.
library.utoronto.ca/eir/
EIRdetail.cfm?Resources\
_\_ID=80615&T=resource.

Leis:2002:DSP

[Lei02] John Leis. Digital signal processing: a MATLABbased tutorial approach, volume 20 of Industrial control, computers, and communications series. Research Studies Press, Baldock, Hertfordshire, England, 2002. ISBN 0-86380-276-1. xv + 217 pp. LCCN TK5102.9 .L45 2002.

Leis:2011:DSP

[Lei11] John W. (John William)

Leis. Digital Signal Processing Using MATLAB for Students and Researchers. Wiley, New York, NY, USA, 2011. ISBN 0-470-88091-0. ???? pp. LCCN TK5102.9 .L4525

2011. URL http://www.
loc.gov/catdir/enhancements/
fy1102/2010048285-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1102/2010048285-
t.html.

Leon:1994:LAA

[Leo94] Steven J. Leon. Linear Algebra with Applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 1994. ISBN 0-02-369831-4. xvi + 506 pp. LCCN QA184.L46 1994.

Leon:1998:LAA

[Leo98] Steven J. Leon. Linear Algebra with Applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-849308-1. xiv + 491 pp. LCCN QA184.L46 1998.

Leon:2002:LAA

[Leo02]  Steven J. Leon.  Lin-

ear algebra with applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, sixth edition, 2002. ISBN 0-13-033781-1. xv + 544 pp. LCCN QA184.2 .L46 2002.

Levinson:1992:TEL

[Lev92] D. P. Levinson. Teaching elementary linear algebra with MATLAB to engineering students. In Lum [Lum92], pages 854– 861. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Levy:1996:SEM

[Lev96] Roy Levy. Structural Engineering of Microwave Antennas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-1020-9. xvi + 354 pp. LCCN TK7871.67.M53L48 1996.

Levi:2003:AQM

[Lev03] A. F. J. (Anthony Frederic John) Levi. Applied quantum mechanics. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-81765-X, 0-521-52086-X (paperback). xv + 523 pp. LCCN QC174.12 .L44 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam031/2002073608.
html;   http://www.loc.
gov/catdir/samples/cam041/
2002073608.html;   http:
//www.loc.gov/catdir/toc/
cam031/2002073608.html.

Ljung:1994:MDS

[LG94] Lennart Ljung and Torkel Glad. Modeling of Dynamic Systems. Prentice Hall information and system sciences series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-597097-0. 361 (est.) pp. LCCN QA401 .L58 1994.

Lee:2005:OHF

[LGML05] D.-U. Lee, A. A. Gaffar, O. Mencer, and W. Luk. Optimizing hardware function evaluation. IEEE Transactions on Computers, 54(12):1520–1531, December 2005. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lameed:2013:MAS

[LH13] Nurudeen A. Lameed and Laurie J. Hendren. A modular approach to on-stack replacement in LLVM. ACM SIGPLAN Notices, 48(7): 143–154, July 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Lameed:2014:OMF

[LH14] Nurudeen A. Lameed and Laurie J. Hendren. Optimizing MATLAB feval with dynamic techniques. ACM SIGPLAN Notices, 49(2): 85–96, February 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). DLS ’13 conference proceedings.

Leon:1996:ACE

[LHF96] Steven Leon, Eugene Herman, and Richard Faulkenberry. ATLAST Computer Exercises for Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-270273-8. xxii + 219 pp. LCCN QA185.C65L46 1996. Book consists of a series of MATLAB projects.

Leon:2003:ACE

[LHF03] Steven Leon, Gene Herman, and Richard Faulkenberry. ATLAST Computer Exercises for Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2003.

ISBN 0-13-101121-9. 210 (est.) pp. LCCN ????

Luff:2001:RFF

[LHW01] R. Luff, M. Haeckel, and K. Wallmann. Robust and fast FORTRAN and MAT-LAB(R) libraries to calculate pH distributions in marine systems. Computers and Geosciences, 27(2):157– 169, March 2001. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Liu:2023:MBM

[LHZT23] Guoyang Liu, Janet H.

Hsiao, Weidong Zhou, and

Lan Tian. MartMi-BCI: a

Matlab-based real-time motor imagery brain-computer interface platform. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000675.

Li:1996:IRS

[Li96]  Wei-Hua Li.  Interfac-

ing real-time software with MATLAB: research report. Master of Science, Dept. of Mechanical Engineering, University of California, Berkeley, Berkeley, CA, USA, 1996. various pp.

Li:1999:PRS

[Li99] X. Rong Li. Probability, random signals, and statistics: a textgraph with integrated software for electrical and computer engineers. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-0433-4. 455 (est.) pp. LCCN TK5102.9 .L49 1999.

Li:2020:IFE

[Li20] Gang Li. Introduction to the finite element method and implementation with MAT-LAB. Cambridge University Press, Cambridge, UK, 2020. ISBN 1-108-47168-4 (hardcover), 1-108-66897-6 (e-book), 1-108-55905-0 (e-pub). LCCN TA347.F5 L485 2020.

Lindner:1999:ISS

[Lin99] Douglas K. Lindner. Introduction to signals and systems. WCB/McGraw-Hill, Boston, MA, USA, 1999. ISBN 0-256-25259-9 , 0-07-116489-8. xiv + 969 pp. LCCN TK5102.9 .L55 1999.

Lin:2005:OMM

[Lin05] Michael Lin. Operational Manual for MatLAB tool: “MeasureBrArt”. Thesis (M.D.), UCSD School of Medicine, San Diego, CA, USA, 2005. 73 pp. Independent study project, Class of 2005, [no. 59].

Lipovetsky:2007:BRBa

[Lip07] Stan Lipovetsky. Book review: The Structural Representation of Proximity Matrices with MATLAB by Lawrence Hubert; Phipps Arabie; Jacqueline Meulman. Technometrics, 49(1): 107, February 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471292.

Liu:2023:RMP

[Liu23] Yang Liu. rwl read: a MATLAB program to identify and correct formatting errors in tree-ring measurement files. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022002035.

Li:2023:APA

[LJ23] Yipeng Li and Xiangmin Jiao. ARPIST: Provably accurate and stable numerical integration over spherical triangles. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004204.

Ludtke:1993:SPD

[LJC93] I. Ludtke, M. G. Jayne, and C. Cabral. The simulation of PWM drives using Matlab. In Anonymous [Ano93a], pages 724– 727. ISBN ???? LCCN ???? Two volumes.

Luk:1993:STP

[LJR93] C. K. P. Luk, M. G. Jayne, and D. Rees. Simulation of three-phase pulse width modulated induction motor drives using MATLAB. In Hamza [Ham93], pages 138– 141. ISBN 0-88986-173-0. ISSN 1021-8181. LCCN ????

Ljung:1987:SIT

[Lju87]  Lennart Ljung.  System

Identification: Theory for the User. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-881640-9. xxi + 519 pp. LCCN QA402 .L591 1987.

Ljung:1988:SIT

[Lju88]  Lennart Ljung.  System

Identification Toolbox for use with MATLAB: User’s Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, April 6, 1988. ISBN ???? 38 + 59 pp. LCCN ????

Ljung:1993:SIT

[Lju93]  Lennart Ljung.  System

Identification Toolbox for use with MATLAB: User’s Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1993. ISBN ???? vi + 84 pp. LCCN ????

Ljung:1999:SIT

[Lju99]  Lennart Ljung.  System

identification: theory for the user. Prentice Hall information and system sciences series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-656695-2 (cloth). xxii + 609 pp. LCCN QA402 .L59 1999; QA 402 .L778 1999.

Laub:1986:CST

[LL86] Alan J. Laub and John N.

Little. Control System Toolbox for use with MATLAB: User’s Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, August 27, 1986. ISBN ???? various pp. LCCN ????

Leonard:1992:UMA

[LL92] Naomi Ehrich Leonard and William S. Levine. Using MATLAB to Analyze and Design Control Systems. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1992. ISBN 0-8053-5423-9. 134 pp. LCCN TJ213.L368 1992.

Leonard:1995:UMA

[LL95] Naomi Ehrich Leonard and William S. Levine. Using MATLAB to Analyze and Design Control Systems. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, second edition, 1995. ISBN 0-8053-2193-4. ix + 212 pp. LCCN TJ213 .L368 1995.

Lin:1996:NFS

[LL96] C. T. (Ching Tai) Lin and C. S. G. Lee. Neural fuzzy systems: a neuro-fuzzy synergism to intelligent systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-235169-2. xvi + 797 pp. LCCN TJ217.5.L57 1996.

Larminie:2003:EVT

[LL03] James Larminie and John Lowry. Electric vehicle technology explained. Wiley, New York, NY, USA, 2003. ISBN 0-470-85163-5. xvii + 296 pp. LCCN TL220 .L37 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley039/2003062752.
;   http://www.loc.gov/
catdir/description/wiley039/
2003062752.html;   http:
//www.loc.gov/catdir/toc/
wiley032/2003062752.;
http://www.loc.gov/catdir/
toc/wiley032/2003062752.
html; UCLA.

Lin:2006:MPS

[LLLW06] Xiaoyan Lin, Andong Liu, Yude Li, and Pengju Wu. A MATLAB programming for simulation of X-ray capillaries. Applied Mathematics and Computation, 172 (1):188–197, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:2023:AKM

[LLW^{+}23] Hao Lin, Hongfu Liu, Junjie Wu, Hong Li, and Stephan Günnemann. Algorithm 1038: KCC: a MATLAB package for k-means-based consensus clustering. ACM Transactions on Mathematical Software, 49(4):40:1– 40:??, December 2023. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
10.1145/3616011.

Lee:2018:RAP

[LLZ18] Tsung-Lin Lee, Tien-Yien Li, and Zhonggang Zeng. RankRev: a Matlab package for computing the numerical rank and updating/downdating. Numerical Algorithms, 77(2):559– 576, February 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

https://link.springer.
com/article/10.1007/s11075-
017-0328-7.

Little:1991:MHP

[LM91] John Little and Cleve B.

Moler. MATLAB highperformance numeric computation software: PRO-MATLAB user’s guide: PRO-MATLAB for VAX. MathWorks, South Natick, MA, USA, March 1, 1991. ???? pp. LCCN QA184.M3 L57 1991.

Levanon:2004:RS

[LM04] Nadav Levanon and Eli Mozeson. Radar signals. Wiley, New York, NY, USA, 2004. ISBN 0-471-47378-2 (cloth). xiv + 411 pp. LCCN TK6575 .L478 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley047/2003056882.
                                        html;   http://www.loc.
                                        gov/catdir/description/
                                        wiley041/2003056882.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley041/2003056882.
                                        html.

Laudadio:2024:MPC

[LMAV24] Teresa Laudadio, Nicola Mastronardi, Walter Van Assche, and Paul Van Dooren. A Matlab package computing simultaneous Gaussian quadrature rules for multiple orthogonal polynomials. Journal of Computational and Applied Mathematics, 451(??): ??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042724003583.

Leontaris:2018:ANM

[LMN18] Georgios Leontaris and Oswaldo Morales-Nápoles.

ANDURIL — a MATLAB toolbox for ANalysis and Decisions with UnceRtaInty: Learning from expert judgments. SoftwareX, 7(??):313–317, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711018300608.

Lee:1998:PCM

[LNC98] P. L. Lee, R. B. Newell, and I. T. Cameron. Process control and management. Blackie Academic and Professional, London, UK, 1998. ISBN 0-7514-0457-

8. xvii + 509 pp. LCCN TP155.75 .L44 1998.

Liu:2019:EFL

[LNLB19] Bing Liu, Shiva Nejati, Lucia, and Lionel C. Briand. Effective fault localization of automotive Simulink models: achieving the tradeoff between test oracle effort and fault localization accuracy. Empirical Software Engineering, 24(1):444–490, February 2019. CODEN ES-ENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10664-018-9611-z.

Lougheed:2016:PMO

[LO16] Bryan Lougheed and Stephen Obrochta. MatCal: Open source Bayesian ^{14}C age calibration in Matlab. Journal of Open Research Software, 4(1):e42–??, November 08, 2016. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.130/.

Lowe:2001:CRM

[Löw01]  Arno Löwe.  Chemis-

che Reaktionstechnik mit MATLAB und SIMULINK. (German) [Chemical Reaction Techniques with MAT-LAB and SIMULINK]. Wiley-VCH, Berlin, Germany and New York, NY, USA, 2001. ISBN 3-527-30268-9. xxiii + 503 pp. LCCN ????

Lindfield:1997:UMS

[LP97] G. R. Lindfield and J. E. T. Penny. Using MATLAB for sparse matrices. International journal of mathematical education in science and technology, 28 (3):427–??, ???? 1997. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Landau:2005:FCS

[LP05] Rubin H. Landau and M. J. Páez. A first course in scientific computing: symbolic, graphical, and numeric problem solving using Maple, Java, Mathematica, and Fortran. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-12183-4 (hardcover). xxiv + 481 pp. LCCN Q183.9 .L36 2005.

Landi:2013:NMS

[LP13] Germana Landi and Elena Loli Piccolomini. NPTool: a Matlab software for nonnegative image restoration with Newton projection methods. Numerical Algorithms, 62(3):487– 504, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11075-012-9602-x/
;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=1017-1398&
volume=62&issue=3&spage=
487-504.

Latifis:2017:MVC

[LPD^{+}17] Ioannis Latifis, Karthick Parashar, Grigoris Dimitroulakos, Hans Cappelle, Christakis Lezos, Konstantinos Masselos, and Francky Catthoor. A MATLAB vectorizing compiler targeting applicationspecific instruction set processors. ACM Transactions on Design Automation of Electronic Systems, 22 (2):32:1–32:28, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Latifis:2020:RMC

[LPD^{+}20] Ioannis Latifis, Karthick Parashar, Grigoris Dimitroulakos, Hans Cappelle, Christakis Lezos, Konstantinos Masselos, and Francky Catthoor. A retargetable MATLAB-to-C compiler exploiting custom instructions and data parallelism. ACM Transactions on Embedded Computing Systems, 19(6):50:1– 50:27, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3391898.

Liang:1994:MTR

[LQ94] ?. Liang and ?. Quinn. MATLAB in teaching a robotics course. CoED, 4(1): 52–57, January 1994. CO-DEN CWLJDP. ISSN 0736-8607.

Li:2018:NSD

[LQT18] Zhilin Li, Zhonghua Qiao, and Tao Tang. Numerical solution of differential equations: introduction to finite difference and finite element methods. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-107-16322-6 (hardcover), 1-316-61510-3 (paperback), 1-316-67872-5. ix + 293 pp. LCCN QA371 .L59 2018.

LopezOrozco:1995:MSN

[LRD^{+}95] J. A. Lopez Orozco, P. Ruiperez, J. M. De la Cruz, Aranda, and J. Modeling and simulation of navigation systems:

An INS simulation Matlab toolbox. In Breitenecker and Husinsky [BH95], pages 607–612. ISBN 0-444-82241-0. LCCN QA76.9.C65E966 1995.

Little:1988:SPT

[LS88] John Little and Loren Shure. Signal Processing Toolbox for use with MATLAB: User’s Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, August 29, 1988. ISBN ???? various pp. LCCN CM.0.M3.S5.

Little:1993:SPT

[LS93] John Little and Loren Shure. Signal Processing Toolbox for use with MATLAB: User’s Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1993. ISBN ???? LCCN QA188 .L58 1992.

Lewis:1995:OC

[LS95] Frank L. Lewis and Vassilis L. Syrmos. Optimal control. Wiley, New York, NY, USA, second edition, 1995. ISBN 0-471-03378-2 (cloth). xi + 541 pp. LCCN QA402.3 .L487 1995. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley042/95015649.
html;   http://www.loc.
gov/catdir/description/

                                        wiley034/95015649.html;
                                         http://www.loc.gov/catdir/
                                        toc/onix05/95015649.html.

Lamba:1998:CRM

[LS98] H. Lamba and A. M. Stuart. Convergence results for the MATLAB ODE23 routine. BIT Numerical Mathematics, 38(4):751– 780, 1998. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Lee:2004:OPD

[LS04] H. J. (Hyun Jin) Lee and W. E. Schiesser. Ordinary and partial differential equation routines in C, C++, Fortran, Java^{(}R), Maple^{(}R), and MATLAB^{(}R). Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-423-1. xiii + 519 pp. LCCN QA371.5.D37 L44 2004.

Lemmon:2005:DSS

[LS05a] David R. Lemmon and J. L. (Joseph L.) Schafer. Developing statistical software in Fortran 95. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-28123-1, 0-387-23817-4 (paperback). xv + 323 pp. LCCN QA76.5 .L453

                                        2005. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0662/2004061447-d.html;

http://www.loc.gov/catdir/
enhancements/fy0818/2004061447-
t.html.

Lonngren:2005:FEM

[LS05b] Karl E. Lonngren and Sava V. Savov. Fundamentals of electromagnetics with MATLAB. Scitech Publishing Inc., Raleigh, NC, USA, 2005. ISBN 1-891121-38-

3. xii + 641 pp. LCCN QC760.6 .L55 2005.

Lonngren:2005:IEM

[LS05c] Karl E. (Karl Erik) Lonngren and Sava Vasilev Savov. Introduction to Electromagnetics with Matlab. Scitech Publishing Inc., Raleigh, NC, USA, 2005. ISBN 1-891121-38-3, 1-891121-30-8. xii + 641 pp.

Lu:1995:PPA

[LSH95] Siqing Lu, Edwin Swidenbank, and Brian W. Hogg. Power plant analyser — an applied MATLAB toolbox. In Anonymous [Ano95b], pages 11/1–11/4. ISBN ???? LCCN ????

Lambrette:1995:CMC

[LSPM95] U. Lambrette, B. Schmandt, G. Post, and H. Meyr. COSSAP MATLAB cosimulation. In Anonymous [Ano95s], pages 1789–1793. ISBN ???? LCCN ????

Lorenzo-Seva:2019:MGU

[LSvdV19] Urbano Lorenzo-Seva and Michel van de Velden.

MultipleCar: a graphi-

cal user interface MATLAB toolbox to compute multiple correspondence analysis. Journal of Statistical Software, 90(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v090i04;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v090i04/
v90i04.pdf.

Lorenzo-Seva:2009:CMP

[LSvdVK09] Urbano Lorenzo-Seva, Michel van de Velden, and Henk A. L. Kiers. CAR: a MATLAB package to compute correspondence analysis with rotations. Journal of Statistical Software, 31(8):??, September 2009. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v31/
i08.

Lutovac:2001:FDS

[LTE01] Miroslav D. Lutovac, Dejan V. Tošić, and Brian L. (Brian Lawrence) Evans. Filter design for signal processing using MATLAB and Mathematica. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-201-36130-2. xxvi + 756 pp. LCCN TK7872.F5 L88 2001.

Lu:1996:MNT

[Lu96] Y. Lu. MATLAB NEC toolbox: The cross platform GUI pre- and postprocessing tool for NEC applications. In Anonymous [Ano96t], pages 1142–1149. CODEN CPCEFK. ISBN ???? LCCN ????

Lucarini:2005:KKR

[Luc05] V. (Valerio) Lucarini, editor. Kramers–Krönig relations in optical materials research, volume 110 of Springer series in optical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-23673-2 (hardcover). ix + 160 pp. LCCN QC451 .K8865 2005.

Lum:1991:ICN

[Lum91] Lewis Lum, editor. Proceedings of the Fourth Annual International Conference on Technology in Collegiate Mathematics, Portland State University, Department of Mathematics, Portland, Oregon, November 15–17, 1991. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Lum:1992:PFA

[Lum92] Lewis Lum, editor. Proceedings of the Fifth Annual International Conference on Technology in Collegiate Mathematics: William Rainey Harper College and Northern Illinois University, Rosemont, Illinois, November 12–15, 1992. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Lumkes:2002:CSD

[Lum02] John H. Lumkes. Control strategies for dynamic systems: design and implementation. Mechanical engineering. Marcel Dekker, New York, NY, USA, 2002. ISBN 0-8247-0661-7. vii + 597 pp. LCCN TJ213 .L78 2002. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        toc/fy034/2001059228.html.

Luo:1995:FDF

[Luo95] Xinyue Luo. FIR digital filter design using quadratic and linear programming with MATLAB. Thesis (M.S.), California State University, Northridge, Northridge, CA, USA, 1995. vi + 76 pp.

Luszczek:2009:PPM

[Lus09]  Piotr Luszczek.  Paral-

lel programming in MAT-LAB. The International Journal of High Performance Computing Applications, 23(3):277–283, August 2009. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.

sagepub.com/content/23/
3/277.full.pdf+html.

Lee:2003:SIS

[LV03] Edward A. Lee and P. P.

(Pravin Pratap) Varaiya. Structure and interpretation of signals and systems. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-74551-8. xxi + 647 pp. LCCN TK5102.9 .L43 2003.

Ledoux:2016:MMT

[LV16] Veerle Ledoux and Marnix Van Daele. Matslise 2.0: A Matlab toolbox for Sturm–Liouville computations. ACM Transactions on Mathematical Software, 42(4):29:1–29:18, July 2016. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:

//dl.acm.org/citation.
cfm?id=2839299.

Ledoux:2005:MMP

[LVV05] V. Ledoux, M. Van Daele, and G. Vanden Berghe. MATSLISE: a MATLAB package for the numerical solution of Sturm– Liouville and Schrödinger equations. ACM Transactions on Mathematical Software, 31(4):532–554, December 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Leonard:1994:DNE

[LW94] D. J. Leonard and D. J. Wilcox. Development of a new EMTP using MAT-LAB. In Proceedings of the 29th Universities Power Engineering Conference. Part 2 (of 2), volume 2, pages 641–644. APC, 1994.

Linz:2003:ENM

[LW03] Peter Linz and Richard Wang. Exploring numerical methods: an introduction to scientific computing using MATLAB. Jones and Bartlett Publishers, Boston, MA, USA, 2003. ISBN 0-7637-1499-2. xii + 473 pp. LCCN QA297 . L54 2003.

Laiho:2002:RNP

[LWN02] Jaana Laiho, Achim Wacker, and Tomáš Novosad, editors. Radio network planning and optimisation for UMTS. Wiley, New York, NY, USA, 2002. ISBN 0-471-48653-1. xxv + 484 pp. LCCN TK5103.483 .R34 2002. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley036/2001045566.
                                        html;   http://www.loc.
                                        gov/catdir/toc/wiley021/
                                        2001045566.html.

Lewis:1997:BCS

[LY97] Paul H. Lewis and Chang Yang. Basic Control Systems Engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-597436-4. xi + 450 pp. LCCN TJ213 .L434 1997.

Li:2016:EDP

[LYH^{+}16] Cong Li, Can Yang, Greg Hather, Ray Liu, and Hongyu Zhao. Efficient drug-pathway association analysis via integrative penalized matrix decomposition. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 13(3):531–540, May 2016. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Lynch:2004:DSA

[Lyn04] Stephen Lynch. Dynamical systems with applications using MATLAB. Birkhäuser Boston Inc., Cambridge, MA, USA, 2004. ISBN 0-8176-4321-4, 3-7643-4321-4. xv + 459 pp. LCCN QA614.8 .L97 2004.

Lyshevski:2000:CST

[Lys00a] Sergey Edward Lyshevski.

Control systems theory with engineering applications. Control engineering. Birkhäuser Verlag, Basel, Switzerland, 2000. ISBN 0-8176-4203-X. xi + 416 pp. LCCN TJ213 .L94 2000.

Lyshevski:2000:ESE

[Lys00b] Sergey Edward Lyshevski.

Electromechanical systems, electric machines, and applied mechatronics. Electric power engineering series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-2275-

8. 782 (est.) pp. LCCN TK7881.15 .L97 2000.

Lyshevski:2001:CST

[Lys01] Sergey Edward Lyshevski. Control systems theory with engineering applications. Birkhäuser Verlag, Basel, Switzerland, 2001. ISBN 0-8176-4203-X. xi + 416 pp. LCCN TJ213 .L94 2000.

Lyshevski:2003:ESC

[Lys03] Sergey Edward Lyshevski. Engineering and Scientific Computations Using MAT-LAB. Wiley, New York, NY, USA, 2003. ISBN 0-471-46200-4. x + 227 pp. LCCN TA345 .L97 2003. US$59.95. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley045/2003278438.
html;   http://www.loc.
gov/catdir/description/
wiley041/2003278438.html;
 http://www.loc.gov/catdir/
toc/wiley041/2003278438.
html.

Lyshevski:2005:NME

[Lys05] Sergey Edward Lyshevski. Nano- and micro-electromechanical systems: fundamentals of nano- and microengineering, volume 8 of Nano- and microscience, engineering, technology, and medicine series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2005. ISBN 0-8493-2838-1. 722 (est.) pp. LCCN TK7875 .L96 2005.

Liu:2015:CMP

[LZ15] Xiaohui Liu and Yijun Zuo.

CompPD: A MATLAB package for computing projection depth. Journal of Statistical Software, 65(2): ??, June 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v65/i02.

Li:2017:CSS

[LZ17] Cai Li and Hua Zhou. Code snippet: svt: Singular value thresholding in MAT-LAB. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v081c02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v081c02/
v81c02.pdf.

Murthy:1994:CRI

[M^{+}94] T. K. S. Murthy et al., editors. Computers in railways IV: Fourth International Conference on Computer Aided Design, Manufacture, and Operation in the Railway and Other Mass Transit Systems, COMPRAIL ’94, Madrid, Spain, 7–9 September 1994, volume 4. Computational Mechanics Publications, Southampton; Boston, 1994. ISBN 1-85312-266-1 (Southampton: set), 1-56252-190-X (Boston: set), 1-85312-354-4 (Southampton: v. 1), 1-56252-282-5 (Boston: v. 1), 1-85312-359-5 (Southampton: v. 2), 1-56252-283-3 (Boston: v. 2). LCCN TF507.I55 1994.

Moshirvaziri:1995:MLP

[MA95] Khosrow Moshirvaziri and Mahyar A. Amouzegar. A Matlab linear programming tool for use in global optimization algorithms. Technical report ENG-95-146, UCLA School of Engineering and Applied Science, Los Angeles, CA, USA, 1995. ii + 16 pp.

Mahseredjian:1996:CET

[MA96a] J. Mahseredjian and F. Alvarado. Creating an electromagnetic transients program in MATLAB: MatEMTP. In IEEE [IEE96a], pages 380–388. ISBN ???? ISSN 0885-8977 (print), 1937-4208 (electronic). LCCN ????

Moeini:1996:MTS

[MA96b] A. Moeini and D. P. Atherton. A MATLAB toolkit for the study of oscillations in relay systems. In IEEE [IEE96d], pages 286– 291. ISBN 0-7803-3032-3, 0-7803-3033-1. LCCN TJ212.2.I32495 1996. IEEE catalog number 96TH8136.

Mahseredjian:1997:CETa

[MA97a] J. Mahseredjian and F. Alvarado. Creating an electromagnetic transients program in MATLAB: MatEMTP. IEEE power engineering review, 17(1):59– ??, ???? 1997. ISSN 0272-1724.

Mahseredjian:1997:CETb

[MA97b] J. Mahseredjian and F. Alvarado. Creating an electromagnetic transients program in MATLAB: MatEMTP. IEEE transactions on power delivery: a publication of the Power Engineering Society, 12(1): 380–??, ???? 1997. ISSN 0885-8977 (print), 1937-4208 (electronic).

Magrab:2011:EGM

[MAB^{+}11] Edward B. Magrab, Shapour Azarm, Balakumar Balachandran, James Duncan, Keith E. Herold, and Gregory C. Walsh. An engineer’s guide to MATLAB: with applications from mechanical, aerospace, electrical, civil, and biological systems engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2011. ISBN 0-13-199110-8. xxii + 823 pp. LCCN TA345 .M328 2011.

Maciejowski:1991:MFD

[Mac91] J. M. Maciejowski. Multivariable Frequency Domain Toolbox and its relation to other MATLAB toolboxes. IEE Conference Publication, 1(332):471–475, 1991. CODEN IECPB4. ISSN 0537-9987 (??invalid ISSN checksum??).

MacCluer:2000:IMM

[Mac00] C. R. MacCluer. Industrial mathematics: modeling in industry, science, and government. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-949199-6. xii + 308 pp. LCCN QA401 .M33 2000.

Maciejowski:2002:PCC

[Mac02] Jan Marian Maciejowski. Predictive control: with constraints. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-201-39823-0. xviii + 331 pp. LCCN TJ217.6 .M23 2002.

Mackenroth:2004:RCS

[Mac04] Uwe Mackenroth. Robust control systems: theory and case studies. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20929-8 (hd.bd.). xvi + 519 pp. LCCN TJ217.2 .M33 2004.

MacCluer:2005:CVM

[Mac05] C. R. MacCluer. Calculus of variations: mechanics, control theory, and other applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-142383-5. x + 254 pp. LCCN QA315 .M23 2005.

Macedo:2008:AGM

[MAC08] Raquel S. Macedo, Marcelo F.

Alfradique, and Marcelo Castier. Automatic generation of Matlab functions using Mathematica and Thermath. Computing in Science and Engineering, 10(4):41– 49, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Magrab:2000:EGM

[Mag00] Edward B. Magrab. An engineer’s guide to MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-011335-2 (paperback). xiv + 694 pp. LCCN QA76.73.P22 S42 2000.

Magni:2002:RMC

[Mag02] Jean-François Magni. Robust modal control with a toolbox for use with MAT-LAB. Kluwer Academic… Plenum Publishers, New York, NY, USA, 2002. ISBN 0-306-46773-9. xi + 312 pp. LCCN QA402.3 .M326 2002.

Magrab:2005:EGM

[Mag05] Edward B. Magrab, editor. An engineer’s guide to MATLAB: with applications from mechanical, aerospace, electrical, and civil engineering. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, second edition, 2005. ISBN 0-13-145499-4. ???? pp. LCCN TA345 .A52 2005.

Mahafza:2000:RSA

[Mah00] Bassem R. Mahafza. Radar Systems Analysis and Design Using Matlab. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 1-58488-182-8. 529 (est) pp. LCCN TK6575 .M27 2000. US$94.95, UK£63.99.

Mahafza:2005:RSA

[Mah05] Bassem R. Mahafza. Radar Systems Analysis and Design Using MATLAB. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2005. ISBN 1-58488-532-7. 638 (est.) pp. LCCN TK6575 .M27 2005.

Maillet:2000:TQS

[Mai00] Denis Maillet, editor. Thermal quadrupoles: solving the heat equation through integral transforms. Wiley, New York, NY, USA, 2000.

ISBN 0-471-98320-9. xiii + 370 pp. LCCN QA377 .T48 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley032/00039931.
html;   http://www.loc.
gov/catdir/toc/onix03/
00039931.html.

Makarov:2002:AEM

[Mak02] Sergey N. Makarov. Antenna and EM modeling with Matlab. Wiley, New York, NY, USA, 2002. ISBN 0-471-21876-6 (cloth). xiii + 273 pp. LCCN TK6565.A6 M283 2002. US$99.95. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley044/2002027225.
html;   http://www.loc.
gov/catdir/description/
wiley036/2002027225.html;
 http://www.loc.gov/catdir/
toc/wiley023/2002027225.
html.

Malkinson:1996:CPC

[Mal96] Terrance J. Malkinson, editor. Conference proceedings: 1996 Canadian Conference on Electrical and Computer Engineering, May 26– 29, 1996, the University of Calgary, Alberta, Canada: theme, Glimpse into the 21st century, RF EXPO WEST 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3143-5, 0-7803-3144-3. LCCN TK7801.C36 1996. Two volumes.

Mangiacasale:1996:ACS

[Man96] Luigi Mangiacasale. Airplane Control Systems: Mu-Synthesis with MATLAB. Editrice Levrotto & Bella, Torino, Italy, 1996. ISBN ???? 468 pp. LCCN ????

Manassah:2001:EMC

[Man01] Jamal T. Manassah. Elementary mathematical and computational tools for electrical and computer engineers using MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-1080-6 (paperback). 352 (est.) pp. LCCN TK153 .M362 2001. US$49.95, UK£33.99. URL

                                        http://www.engnetbase.
                                        com/.

Mandal:2003:MSS

[Man03] Mrinal Kr Mandal. Multimedia signals and systems, volume SECS 716 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 2003. ISBN 1-4020-7270-8. xiv + 375 pp. LCCN QA76.575 .M3155 2003.

Marsaglia:1968:RNF

[Mar68] George Marsaglia. Random numbers fall mainly in the planes. Proceedings of the National Academy of Sciences of the United States of America, 61(1): 25–28, September 15, 1968. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). A popularized account of this work appeared as “Are random numbers really random?” [Scientific Research (Philadelphia, PA), 3 (1968), 21–??]. This widelycited paper describes the hyperplane problem that linear congruential generators suffer from, although careful choice of multipliers can minimize its importance: see [CM67, DH97b, DH97a, DH00].

Marcotte:1991:CM

[Mar91] Denis Marcotte. Cokriging with Matlab. Computers and Geosciences, 17(9): 1265–1280, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Marcus:1992:MMT

[Mar92] Marvin Marcus. Matrices and MATLAB: a tutorial. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-562901-2. xvii + 710 pp. LCCN QA188 .M348 1992.

Marchand:1995:GGM

[Mar95a] Patrick R. Marchand. Graphics and GUIs with MAT-LAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-9487-2. vi + 365 pp. LCCN T385 .M3634 1995. US$39.95. Includes two 3.5in floppy disks.

Marlin:1995:PCD

[Mar95b] Thomas E. Marlin. Process Control: Designing Process and Control Systems for Dynamic Performance. McGraw-Hill chemical engineering series. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-040491-7. xxii + 954 pp. LCCN TP155.75.M365 1995.

Marchand:1999:GGM

[Mar99] Patrick Marchand. Graphics and GUIs with MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 1999. ISBN 0-8493-9001-X. 445 pp. LCCN T385.M3634 1999. US$39.95.

Marvasti:2001:NST

[Mar01] Farokh A. Marvasti, editor. Nonuniform sampling: theory and practice. Information technology– transmission, processing, and storage. Kluwer Academic… Plenum Publishers, New York, NY, USA, 2001. ISBN 0-306-46445-4. xxvi + 924 pp. LCCN TK5102.5 .N56857 2001.

MarquesdeSa:2003:ASU

[Mar03] J. P. Marques de Sá. Applied statistics: using SPSS, STATISTICA, and MAT-LAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-01156-0. xxii + 452 pp. LCCN QA276.4 .A575 2003.

MarquesdeSa:2007:ASU

[Mar07] Joaquim P. Marques de Sá. Applied Statistics Using SPSS, STATISTICA, MATLAB and R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-71972-5. xxiv + 506 + 196 pp. LCCN QA276.4 S2232 2007.

Martinez:2011:CSM

[Mar11] Wendy L. Martinez. Computational statistics in MATLAB(R). WIREs Computational Statistics, 3 (1):69–74, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Maronna:2014:BRW

[Mar14] Ricardo Maronna. Book review: Wendy Martínez, Angel R. Martínez, Jeffrey L. Solka: Exploratory data analysis with MATLAB^{(R)}, second edition. Statistical Papers, 55(3):913–914, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s00362-
013-0518-x.

Markussen:2016:PPC

[Mar16] Thor Markussen. Parchar — characterization of suspended particles through image processing in Matlab. Journal of Open Research Software, 4(1): e26–??, July 19, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.114/.

Marzjarani:2019:BRE

[Mar19] Morteza Marzjarani. Book review: Exploratory Data Analysis With MATLAB by Wendy L. Martinez, Angel R. Martinez, and Jeffery L. Solka. Boca Raton, FL: CRC Press, 2017, xxv + 590 pp., $100.00 (Hardback), $46.36 (e-book), ISBN-13: 978-1-4987-7606-6. Technometrics, 61(4):565–566, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

McCalley:1996:SCD

[MASV96] J. D. McCalley, V. Ajjarapu, G. B. Sheble, and V. Vittal. Sophomore course development in power system analysis with interactive Matlab modules. In Cameron et al. [C^{+}96], pages 1101–1104. ISBN 0-7803-3637-2, 0-7803-3636-4, 0-7803-3638-0. LCCN TK3226 .M553 1996. Three volumes. IEEE catalog number: 96CH35995.

Mather:1991:ED

[Mat91] B. G. Mather. Embedding DSP. IEEE Spectrum, 28 (11):52–55, November 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mathews:1992:NMM

[Mat92a] John H. Mathews. Numerical Methods for Mathematics, Science, and Engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1992. ISBN 0-13-624990-6. x + 646 pp. LCCN QA297 .M39 1992.

MathWorks:1992:MUG

[Mat92b] The MathWorks, Inc., Natick, MA, USA. MATLAB User’s Guide, 1992. ???? pp.

Matstoms:1992:SQF

[Mat92c] P. Matstoms. Sparse QR factorization in MATLAB.

Report LiTH-MAT-R-1992-05, Department of Mathematics, Linköping University, Sweden, March 1992. ???? pp.

Mathews:1994:UMO

[Mat94a] John H. Mathews. Using MATLAB to obtain both numerical and graphical solutions to hyperbolic P.D.E.’s. CoED, 4(1):58– 60, January 1994. CODEN CWLJDP. ISSN 0736-8607.

Matstoms:1994:SQF

[Mat94b] Pontus Matstoms. Sparse QR factorization in MAT-LAB. ACM Transactions on Mathematical Software, 20(1):136–159, March 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

MathWorks:1995:MHN

[Mat95a] The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA. MAT-LAB, high-performance numeric computation and visualization software: release notes, version 4.2, 1995. 90 pp.

MathWorks:1995:MQR

[Mat95b] The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA. MAT-LAB, quick reference, January 1995. 15 + 1 pp.

MathWorks:1995:MST

[Mat95c] The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA. MAT-LAB SIMULINK, toolboxes, version 4.2, volume vi edition, 1995. ???? pp. 1 computer laser optical disk, graphics quick reference + 1 MATLAB quick reference.

MathWorks:1997:MIP

[Mat97]  MathWorks, Inc.  Mat-

lab image processing toolbox: user’s guide [for] version 2. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1997. ISBN ???? various pp. LCCN ????

Mitra:1994:PAS

[MB94] S. Mitra and T. Bose. Processing of analytical signals using Matlab. Chemometrics and Intelligent Laboratory Systems, 22(1):3–16, January 1994. CODEN CILSEN. ISSN 0169-7439.

Mullen:2009:HPS

[MBB^{+}09] Julie Mullen, Nadya Bliss, Robert Bond, Jeremy Kepner, Hahn Kim, and Albert Reuther. High-productivity software development with pMatlab. Computing in Science and Engineering, 11(1):75–79, January/ February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Margrave:1995:MNN

[MBBC95] F. Margrave, N. R. Babu, A. Bradshaw, and I. Collins. MATLAB — neural networks toolbox hardware post- processor. In Anonymous [Ano95b], pages 6/1– 6/3. CODEN DCILDN. ISBN ???? ISSN 0963-3308. LCCN ????

Marcus:1993:CM

[MBGV93] Leslie F. Marcus, Elisa Bellos, and Antonio Garcia-Valdecasas. Contributions to Morphometrics. Museo Nacional de Ciencias Naturales, Madrid, Spain, 1993. ISBN 84-00-07353-3. LCCN ????

Mahseredjian:1997:LBE

[MBL^{+}97] J. Mahseredjian, G. Benmouyal, X. Lombard, M. Zouiti, B. Bressac, and L. Gerin-Ljoie. A link between EMTP and MATLAB for user-defined modeling. IEEE power engineering review, 17(8):41–??, ???? 1997. ISSN 0272-1724.

Mikhail:2001:IMP

[MBM01] Edward M. Mikhail, James S. Bethel, and J. Chris Mc-Glone. Introduction to modern photogrammetry. Wiley, New York, NY, USA, 2001. ISBN 0-471-30924-9. ix + 479 pp. LCCN TR693 .M55 2001.

Musch:1995:IAM

[MBMW95] H. E. Musch, G. W. Barton, B. McKay, and M. Willis. Industrial applications of MATLAB based software for advanced data analysis and control. In Anonymous [Ano95d], pages 211– 215. ISBN 0-85825-631-2. ISSN 0313-6922. LCCN ????

Maclean:2021:TEF

[MBR21] John Maclean, J. E. Bunder, and A. J. Roberts. A toolbox of equation-free functions in Matlab /Octave for efficient system level simulation. Numerical Algorithms, 87(4):1729–1748, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11075-020-01027-
z.

Mehl:2015:RTC

[MBS15] Miriam Mehl, Manfred Bischoff, and Michael Schäfer, editors. Recent Trends in Computational Engineering — CE2014: Optimization, Uncertainty, Parallel Algorithms, Coupled and Complex Problems, volume 105 of Lecture Notes in Computational Science and Engineering. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015.

ISBN 3-319-22996-6, 3-319-22997-4 (e-book). 317 (est.) pp. LCCN QA71-90; TA329. URL http:

                                        //www.springerlink.com/
                                        content/978-3-319-22997-
                                        3.

Mayer:1997:MIG

[MC97] M. Mayer and C. Cenker. A MatLab interface for global optimization in pattern recognition. In Burger and Burge [BB97], pages 87–92. ISBN 3-486-24494-9 (München), 3-7029-0436-0 (Wien), 3-85403-103-3 (Österreichischen Computer Gesellschaft). LCCN TA1650.O88 1997.

Melin:2002:MSC

[MC02] Patricia Melin and Oscar Castillo. Modelling, simulation and control of nonlinear dynamical systems: an intelligent approach using soft computing and fractal theory, volume 2 of Numerical insights, 1028-5350. Taylor and Francis, Boca Raton, FL, USA, 2002. ISBN 0-415-27236-X. xi + 249 pp. LCCN QA427 .M45 2002.

Miller:2004:PRP

[MC04] Scott L. Miller and Donald G. Childers. Probability and random processes: with applications to signal processing and communications. Elsevier Academic Press, Amsterdam, The Netherlands, 2004. ISBN 0-12-172651-7 (hardcover). xiii + 536 pp. LCCN TK5102.9 .M556 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els051/2004010367.
html;   http://www.loc.
gov/catdir/toc/els051/
2004010367.html.

McAndrew:2004:IDI

[McA04] Alasdair McAndrew. An introduction to digital image processing with MATLAB. Brooks/Cole, Pacific Grove, CA, USA, 2004. ISBN 0-534-40011-6. ???? pp.

McClellan:1998:CBE

[McC98] James H. McClellan, editor.

Computer-based exercises for signal processing using MATLAB 5. MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-789009-5 (paperback). xii + 404 pp. LCCN TK5102.9 .C567 1998.

Melendez:1996:EOM

[MCd^{+}96] Joaquim Melendez, Joan Colomer, Josep Lluis de la Rosa, Josep Aguilar-Martin, and Josep Vehi. Embedding objects into MATLAB/SIMULINK for process supervision. In IEEE [IEE96d], pages 20–

25. ISBN 0-7803-3032-3, 0-7803-3033-1. LCCN TJ212.2.I32495 1996. IEEE catalog number 96TH8136.

McGillem:1991:TSA

[McG91] Clare D. McGillem. Teaching signal analysis with MATLAB. Frontiers in Education Conference, pages 763–??, January 1991. CO-DEN PFECDR. ISSN 0190-5848. IEEE catalog number 92CH3069-2.

McIlhagga:2016:PMT

[McI16] William McIlhagga. penalized: A MATLAB toolbox for fitting generalized linear models with penalties. Journal of Statistical Software, 72 (??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v072i06;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v072i06/v72i06.pdf.

McMahon:2007:MD

[McM07] David (David M.) McMahon. MATLAB demystified. McGraw-Hill, New York, NY, USA, 2007. ISBN 0-07-148551-1. viii + 326 pp. LCCN QA297 .M427

                                        2007. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0712/2007013608-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0712/2007013608-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0712/2007013608-t.html.

McNelis:2005:NNF

[McN05] Paul D. McNelis. Neural networks in finance: gaining predictive edge in the market. Academic Press advanced finance series. Elsevier Academic Press, Amsterdam, The Netherlands, 2005. ISBN 0-12-485967-4. xv + 243 pp. LCCN HG4012.5 .M38 2005. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip051/2004022859.
html.

Maclay:1995:CDI

[MD95] D. Maclay and R. E. Dorey.

A controller design and implementation environment for the idle speed control of an internal combustion engine. In Anonymous [Ano95b], pages 2–?? ISBN ???? LCCN ????

Mann:2001:MGG

[MDB01] Stephen Mann, Leo Dorst, and Tim Bouma. The making of GABLE: a geometric algebra learning environment in Matlab. In Geometric algebra with applications in science and engineering (Ixtapa–Zihuatanejo, 1999), pages 491–511. Birkhäuser Boston Inc., Cambridge, MA, USA, 2001.

Mahafza:2004:MSR

[ME04] Bassem R. Mahafza and Atef Z. Elsherbeni. MAT-LAB simulations for radar systems design. Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-392-8. 682 (est.) pp. LCCN TK6585 .M34 2004.

Meador:2002:ASP

[Mea02] Don A. Meador. Analog signal processing with Laplace transforms and active filter design. Delmar/ Thomson Learning, Albany, NY, USA, second edition, 2002. ISBN 0-7668-2818-

2. xii + 514 pp. LCCN TK5102.9 .M42 2002.

Meirovitch:2001:FV

[Mei01] Leonard Meirovitch. Fundamentals of vibrations. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-041345-2. xviii + 806 pp. LCCN TA355 .M43 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/mh021/00030494.
html;   http://www.loc.
gov/catdir/toc/mh021/00030494.
html.

Meister:2005:NLG

[Mei05] Andreas Meister. Numerik linearer Gleichungssysteme: eine Einführung in moderne Verfahren: mit MATLAB-Implementierung von C. Vömel. (German) [Numerical linear equation systems: an introduction to modern methods: with a MAT-LAB implementation by C. Vömel]. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 2005. ISBN 3-528-13135-7. ix + 234 pp. LCCN ????

Mendel:1995:LET

[Men95] Jerry M. Mendel. Lessons in Estimation Theory for Signal Processing, Communications, and Control. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-120981-7. xix + 561 pp. LCCN QA276.8.M46 1995.

Mendel:2001:URB

[Men01] Jerry M. Mendel. Uncertain rule-based fuzzy logic systems: introduction and new directions. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-040969-3. xx + 555 pp. LCCN QA9.64 .M46 2001.

Meucci:2005:RAA

[Meu05] Attilio Meucci. Risk and asset allocation, volume ser. is not counted of Springer finance. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-22213-8. ???? pp.

Meurant:2020:PFM

[Meu20] Gérard Meurant. FLOATP toolbox, Matlab software, variable precision floating point arithmetic. Report ????, Commissariat a l’Énergie Atomique (CEA), ????, France, 2020. URL

https://gerard-meurant.
pagesperso-orange.fr/soft_
meurant_n.html.

Mathews:1994:UMP

[MF94] J. H. Mathews and K. D. Fink. Using MATLAB as a programming language for numerical analysis. International journal of mathematical education in science and technology, 25 (4):481–??, ???? 1994. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Mathews:1999:NMU

[MF99] John H. Mathews and Kurtis D. Fink. Numerical methods using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1999. ISBN 0-13-270042-5. viii + 662 pp. LCCN QA297 .M39 1999. Rev. ed. of: Numerical methods for mathematics, science, and engineering. 2nd ed. (1992).

Miranda:2002:ACE

[MF02] Mario J. Miranda and Paul L. Fackler. Applied computational economics and finance. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-63309-4 (paper), 0-262-13420-9. xviii + 510 pp. LCCN HB143.5 .M567 2002.

Mathews:2004:NMU

[MF04] John H. Mathews and Kurtis D. Fink. Numerical methods using MAT-LAB. Pearson, Upper Saddle River, NJ, USA, fourth edition, 2004. ISBN 0-13-065248-2. ix + 680 pp. LCCN QA297 .M39 2004.

Mathew:1995:LTS

[MFO95] R. Mathew, F. Flinders, and W. Oghanna. Locomotive total systems’ simulation using Simulink. In Anonymous [Ano95i], pages 202– 206. ISBN 0-85296-631-8. ISSN 0537-9989. LCCN TF1055 .I57 1995.

Mehra:2013:ASW

[MG13] Mani Mehra and Kavita Goyal. Algorithm 929: a suite on wavelet differentiation algorithms. ACM Transactions on Mathematical Software, 39(4):27:1– 27:28, July 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Mikhael:2018:PMM

[MG18] Shadia Mikhael and Calum

Gray.  Masks2Metrics

(M2M): A Matlab toolbox for gold standard morphometrics. Journal of Open Source Software, 3 (22):436:1–436:3, February 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00436.

Mattsson:1994:IIJ

[MGC94] Sven Erik Mattsson, John O. Gray, and François E. Cellier, editors. IEEE/ IFAC Joint Symposium on Computer-Aided Control System Design: proceedings, March 7–9, 1994, Tucson, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Mulgrew:2003:DSP

[MGT03] Bernard Mulgrew, Peter M. Grant, and John Thompson. Digital signal processing: concepts and applications. Palgrave Macmillan, Basingstoke, UK, second edition, 2003. ISBN 0-333-96356-3 (paperback). xxvi + 380 pp. LCCN TK5102.9 .M85 2003.

Marsolf:1999:UMS

[MGW99] Bret A. Marsolf, Kyle A. Gallivan, and Harry A. G. Wijshoff. The utilization of matrix structure to generate optimized code from MATLAB programs. International Journal of Parallel Programming, 27(2):73– ??, ???? 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Marchand:2003:GGM

[MH03] Patrick Marchand and O. Thomas Holland. Graphics and GUIs with MAT-LAB. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2003. ISBN 1-58488-320-0. 518 (est.) pp. LCCN T385 .M3634 2003. US$64.00.

Hwu:2012:GCG

[mH12] Wen mei Hwu, editor. GPU computing gems. Applications of GPU computing series. Morgan Kaufmann, Boston, MA, jade edition, 2012. ISBN 0-12-385963-8 (hardback). xvi + 541 + 16 pp. LCCN T385 .G6875 2012.

Michalek:1994:LMM

[Mic94a] Jaroslav Michálek. Linear models and MATLAB. In Summer School MAT-LAB 93 Proceedings (Blato, 1993), volume 4 of Folia Fac. Sci. Natur. Univ. Masaryk. Brun. Math., pages 5–24. Masaryk University, Brno, Czech Republic, 1994.

Michalek:1994:SSM

[Mic94b] Jaroslav Michálek, editor.

Summer School MATLAB 93 Proceedings, volume 4 of Folia Facultatis Scientiarium Naturalium Universitatis Masarykianae Brunensis. Mathematica. Masaryk University, Brno, Czech Republic, 1994. ISBN 80-210-1046-0. ii + 78 pp. Held in Blato, August 23–27, 1993.

Michailidis:2000:BRB

[Mic00] George Michailidis. Book review: Essential MAT-LAB for Scientists and Engineers, by B. D. Hahn. Journal of the Royal Statistical Society. Series D (The Statistician), 49(2): 284, ???? 2000. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
2680983.

Middleton:2000:DAE

[Mid00] Gerard V. Middleton. Data analysis in the earth sciences using Matlab. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-393505-1. ix + 260 pp. LCCN QE33.2.S82 M53 2000.

Manolakis:2000:SAS

[MIK00] Dimitris G. Manolakis, Vinay K. Ingle, and Stephen M. Kogan. Statistical and adaptive signal processing: spectral estimation, signal modeling, adaptive filtering, and array processing. McGraw-Hill series in electrical and computer engineering. Computer engineering; McGraw-Hill series in electrical and computer engineering. Communications and signal processing. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-040051-2. xix + 796 pp. LCCN TK5102.9 .M36 2000.

Mikaitis:2023:MMT

[Mik23] Mantas Mikaitis. Monotonicity of multi-term floatingpoint adders. arxiv.org, ?? (??):1–13, April 3, 2023.

URL https://arxiv.org/
pdf/2304.01407.pdf.

Milewski:2020:MSA

[Mil20] Sławomir Milewski. A Matlab software for approximate solution of 2D elliptic problems by means of the meshless Monte Carlo random walk method. Numerical Algorithms, 83(2):565– 591, February 2020. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/content/
pdf/10.1007/s11075-019-
00694-x.pdf.

Mirotznik:1996:TMP

[Mir96] Mark S. Mirotznik. Translating Matlab programs into C code [software reviews]. IEEE Spectrum, 33(2):63– 64, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1997:MCT

[Mir97] Mark S. Mirotznik. Matlabto-C++ time-saver. IEEE Spectrum, 34(4):14–15, April 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mitra:1998:DSP

[Mit98] Sanjit Kumar Mitra. Digital signal processing: a computer-based approach. McGraw-Hill series in electrical engineering. Communications and signal processing. McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-042953-7. xvii + 864 pp. LCCN TK5102.9.M57 1998.

Mitra:1999:DSP

[Mit99] Sanjit Kumar Mitra. Digital signal processing laboratory using MATLAB. WCB/ McGraw-Hill, Boston, MA, USA, 1999. ISBN 0-07-232721-9 (disk), 0-07-232246-2 (manual), 0-07-232876-2 (package). ix + 230 pp. LCCN TK5102.9 .M5723 1999.

Mitra:2006:DSP

[Mit06] Sanjit Kumar Mitra. Digital signal processing: a computer based approach. McGraw-Hill Higher Education, New York, NY, USA, third edition, 2006. ISBN 0-07-286546-6 (text), 0-07-304837-2 (pkg.). xx + 972 pp. LCCN TK5102.9 .M57 2006. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip0420/2004016994.
html.

Milosavljevi:2001:EEA

[MJ01] Igor Z. Milosavljevi and Marwan A. Jabri. Experimental evaluation of automatic array alignment in parallelized Matlab. Journal of Parallel and Distributed Computing, 61(6): 784–809, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
jpdc.2000.1665;   http:
//www.idealibrary.com/
links/doi/10.1006/jpdc.
2000.1665/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/jpdc.
2000.1665/ref.

Meriam:2002:EM

[MKP02] J. L. (James L.) Meriam, L. G. (L. Glenn) Kraige, and William J. (William John) Palm. Engineering mechanics. Wiley, New York, NY, USA, fifth edition, 2002. ISBN 0-471-40646-5 , 0-471-40645-7. ???? pp. LCCN TA350 .M458 2002.

Manurung:2022:YRD

[MKU22] Auralius Manurung, Lisa Kristiana, and Nur Uddin. YADPF: a reusable deterministic dynamic programming implementation in MATLAB. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022000152.

Moler:1990:PMS

[ML90] Cleve B. Moler and John Little, editors. PRO-MATLAB for Sun workstations: Matlab for Sun workstations: user’s guide. MathWorks, South Natick, MA, January 1990. (various) pp. LCCN TA345.5.P7 P75 1991; QA188 .P76 1990.

Moler:2020:HM

[ML20] Cleve Moler and Jack Little. A history of MAT-LAB. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):81:1–81:67, June

                                        2020.  URL https://
                                        dl.acm.org/doi/abs/10.
                                        1145/3386331.

Mondaca:2021:FEM

[MLAS21] S. L. Mondaca, C. A. Leiva, C. A. Acuña, and E. A. Serey. Flow enhancement of mineral pastes to increase water recovery in tailings: a Matlabbased imaging processing tool. Scientific Programming, 2021(1):5607242:1– 5607242:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/5607242.

Moler:1987:PMUb

[MLB87a] C. Moler, J. Little, and S. Bangert. Pro-Matlab User’s Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1987. ISBN ???? LCCN ????

Moler:1987:PMUa

[MLB87b] C. B. Moler, J. N. Little, and S. Bangert. PC-Matlab Users Guide. Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1987.

Moler:1987:PMV

[MLB87c] Cleve Moler, John Little, and Steve Bangert. PRO-MATLAB for VAX/ VMS computer, version 3.1. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1987. ISBN ???? LCCN ????

Martins:2012:PAM

[MLF^{+}12] Pedro Martins, Paulo Lopes, João P. Fernandes, João Saraiva, and João M. P. Cardoso. Program and aspect metrics for MAT-LAB. Lecture Notes in Computer Science, 7336: 217–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.

                                        springer.com/chapter/10.
                                        1007/978-3-642-31128-4_
                                        16/.

Mrozek:1996:MUS

[MM96] Bogumila Mrozek and Zbigniew Mrozek. MATLAB, Uniwersalne srodowisko do obliczen naukowo-technicznych (Polish) [MATLAB, Universal Environment for Scientific and Engineering Computing]. Wydawnictwo PLJ, Warszawa, Poland, 1996. ISBN 83-7101-325-6. ???? pp. LCCN ????

Malek-Madani:1997:AEM

[MM97] Reza Malek-Madani. Advanced engineering mathematics with Mathematica and Matlab. Addison-Wes-ley, Reading, MA, USA, 1997. ISBN 0-201-30822-3. 6 + 3 + 79–705 pp. LCCN ????

Malek-Madani:1998:AEM

[MM98] Reza Malek-Madani. Advanced engineering mathematics with Mathematica and Matlab. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-59881-7 (vol. 1), 0-201-32549-7 (vol. 2). various pp. LCCN TA345 .M34 1998. Two volumes.

Mokhtari:2000:EAM

[MM00] Mohand Mokhtari and Michel Marie. Engineering applications of MAT-LAB 5 and SIMULINK 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-85233-214-X. xiv + 538 pp. LCCN TA345 .M57 2000.

Martinez:2002:CSH

[MM02] Wendy L. Martinez and Angel R. Martinez. Computational Statistics Handbook with MATLAB. Chapman and Hall/CRC, Boca Raton, FL, USA, 2002. ISBN 1-58488-229-8, 1-4200-3563-0 (e-book). xvii + 591 pp. LCCN QA276.4 .M272 2001. US$79.95, UK£53.99.

Martinez:2005:EDA

[MM05] Wendy L. Martinez and Angel R. Martinez. Exploratory data analysis with MAT-LAB. Computer Science and Data Analysis Series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-366-9. xv + 405 pp. LCCN QA278 .M3735 2005.

Martinez:2008:CSH

[MM08] Wendy L. Martinez and Angel R. Martinez. Computational statistics handbook with MATLAB(R). Computer Science and Data Analysis Series. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2008. ISBN 1-58488-566-1. xxiv + 767 pp.

Martaj:2010:MRS

[MM10] Nadia Martaj and Mohand Mokhtari. Matlab r2009, Simulink et Stateflow pour ingénieurs, chercheurs et étudiants. (French) [Matlab r2009, Simulink and Stateflow for engineers, researchers and students]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-11763-5. ???? pp. LCCN ????

Maage:2024:SMT

[MMD24] Ingrid Måge, Josipa Matić, and Katinka Riiser Dankel. SEC2MWD: a MATLAB toolbox for derivation of molecular weight distributions from size exclusion chromatography. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023002984.

Misiti:1995:MMW

[MMOP95] M. Misiti, Y. Misiti, G. Oppenheim, and J.-M. Poggi. MICRONDE: a Matlab Wavelet Toolbox for signals and images. In Antoniadis and Oppenheim [AO95], pages 239–260. ISBN 0-387-94564-4. ISSN 0930-0325. LCCN QA403.3 .W3857 1995.

Mesa-Montoya:2025:IDT

[MMPOQ^{+}25] Carlos Andrés Mesa-Montoya, Néstor Iván Marín Peláez, Kevin David Ortega-Quiñones, German Andrés Holguín-Londoño, Libardo Vicente Vanegas-Useche, Gian Carlo Daraviña-Peña, Edwan Anderson Ariza-Echeverri, and Diego Vergara. Integration of a digital twin framework for trajectory control of a 2RRR planar parallel manipulator using ROS/ Gazebo and MATLAB. Future Internet, 17(4):146, March 26, 2025. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/17/4/146.

Mokhtari:1997:AMM

[MMR97] Mohand Mokhtari, Abdelhalim Mesbah, and Gérard Roucairol. Apprendre et maitriser MAT-LAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62773-1. 728 pp. LCCN ????

Martinez:2011:EDA

[MMS11] Wendy L. Martinez, Angel R. Martinez, and Jeffrey L. Solka. Exploratory data analysis with MAT-LAB. Chapman and Hall/CRC computer science and data analysis. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2011. ISBN 1-4398-1220-9. xix + 508 pp. LCCN QA278 .M3735 2010.

Martinez:2017:EDA

[MMS17] Wendy L. Martinez, Angel R. Martinez, and Jeffrey L. Solka. Exploratory data analysis with MAT-LAB, volume 4 of Computer science and data analysis series. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2017. ISBN 1-4987-7606-X (hardcover), 1-315-33081-4 (Mobi e-book), 1-4987-7607-8 (PDF e-book), 1-315-34984-1 (ePub). xv + 590 pp. LCCN QA278 .M3735 2017.

Melissinos:2003:EMP

[MN03] Adrian C. (Adrian Constantin) Melissinos and Jim Napolitano. Experiments in modern physics. Academic Press, New York, NY, USA, second edition, 2003. ISBN 0-12-489851-3. xviii + 527 pp. LCCN QC33 .M52 2003. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/els031/2002117796.
                                        html;   http://www.loc.
                                        gov/catdir/toc/els031/
                                        2002117796.html.

Matinnejad:2019:TGT

[MNBB19] R. Matinnejad, S. Nejati, L. C. Briand, and T. Bruckmann. Test generation and test prioritization for Simulink models with dynamic behavior. IEEE Transactions on Software Engineering, 45(9):919–944, September 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Marshall:1994:AMC

[MNHH94] M. A. Marshall, W. J.

Nugent, B. C. Hemeon, and A. K. Halder. Applying MATLAB and C to monitor wind and ice loads on a test transmission line. In Baird and El-Hawary [BEH94], pages 159–164. ISBN 0-7803-2416-1, 0-7803-2417-X. LCCN TK 7801 C36 1994.

Moscinski:1995:ACM

[MO95] Jerzy Mościński and Zbigniew Ogonowski, editors. Advanced control with MAT-LAB and SIMULINK: 3rd Summer school — June 1993, Gliwice, Poland. Ellis Horwood, New York, NY, USA, 1995. ISBN 0-13-309667-X. LCCN QA402.3.A33 1995.

Moeller:2004:MCM

[Moe04] Dietmar P. F. Moeller.

Mathematical and computational modeling and simulation: fundamentals and case studies. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-40389-2. xiii + 422 pp. LCCN QA401 .M5396 2004.

Mohan:2001:AED

[Moh01]  Ned Mohan.  Advanced

electric drives: analysis, control and modeling using Simulink. MNPERE, Minneapolis, MN, USA, 2001. ISBN 0-9715292-0-5. ???? pp. LCCN TK4058 .M5778 2001.

Moin:2010:FEN

[Moi10] Parviz Moin. Fundamentals of engineering numerical analysis. Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-88432-2 (hardcover). ???? pp. LCCN TA335 .M65 2010.

Moler:1980:MIM

[Mol80a] Cleve B. Moler. MATLAB — an interactive matrix laboratory. Technical Report 369, University of New Mexico. Dept. of Computer Science, 1980.

Moler:1980:MUG

[Mol80b] Cleve B. Moler. MAT-LAB user’s guide. Technical Report CS81-1, University of New Mexico. Dept. of Computer Science, November 1980. This describes use of Classic Matlab, the prototype for the very-much expanded professional Matlab from The MathWorks. Classic Matlab is no longer available.

Moler:1981:MIG

[Mol81] Cleve B. Moler. MATLAB installation guide. Technical report, University of New Mexico. Dept. of Computer Science, May 1981. This describes installation of Classic Matlab, the prototype for the very-much expanded professional Matlab from The MathWorks. Classic Matlab is no longer available.

Moler:1982:MUG

[Mol82] C. B. Moler. MATLAB users’ guide. Technical Report CS81-1 (Revised), Department of Computer Science, University of New Mexico, 1982.

Moler:1988:MMV

[Mol88]  C. Moler.  MATLAB:

a mathematical visualization laboratory. In Digest of Papers: COMP-CON Spring 88. Thirty-Third IEEE Computer Society International Conference, pages 480–481 (of xvi + 549). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0828-5. LCCN QA75.5 .C58 1988. IEEE Catalog number 88CH2539-5.

Moler:1993:MMM

[Mol93] Cleve B. Moler. MATLAB’s magical mystery tour. The MathWorks Newsletter, 7 (1):8–9, 1993.

Moler:1994:CCB

[Mol94] Cleve B. Moler. Cleve’s corner: Benchmarks — LINPACK and MATLAB: Fame and fortune from megaflops. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Summer/Fall 1994. 3

                                        pp.  URL http://www.
                                        mathworks.com/company/
                                        newsletter/pdf/sumfall94cleve.
                                        pdf.

Moler:1995:CCT

[Mol95a] Cleve B. Moler. Cleve’s corner: a tale of two numbers: With the Pentium, there is a very small chance of making a very large error. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Win-

                                        ter 1995. URL http://www.
                                        mathworks.com/company/
                                        newsletter/pdf/win95cleve.
                                        pdf.

Moler:1995:CCR

[Mol95b] Cleve B. Moler. Cleve’s corner: Random thoughts: 10^{435} years is a very long time. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Fall 1995.

2 pp. URL http://www.
mathworks.com/company/
newsletter/pdf/Cleve.pdf.

Moler:1995:CCQ

[Mol95c] Cleve B. Moler. Cleve’s corner: The QR algorithm: Striving for infallibility. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Summer 1995.

2 pp. URL http://www.
mathworks.com/company/
newsletter/pdf/sum95cleve.
pdf.

Moler:1995:CCW

[Mol95d] Cleve B. Moler. Cleve’s corner: Why there isn’t a parallel MATLAB. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA,

1995. URL http://www.
mathworks.com/company/
newsletter/pdf/spr95cleve.
pdf.

Moler:1996:CCF

[Mol96a] Cleve B. Moler. Cleve’s corner: Floating points: IEEE Standard unifies arithmetic model. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Fall 1996.

3 pp. URL http://www.
mathworks.com/company/
newsletter/pdf/Fall96Cleve.
pdf.

Moler:1996:CCG

[Mol96b] Cleve B. Moler. Cleve’s corner: Golden ODEs: New ordinary differential equation solvers for MATLAB and SIMULINK. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Summer 1996. 3

pp.  URL http://www.
mathworks.com/company/
newsletter/pdf/Sum96.Cleve.
pdf.

Moler:1997:CCW

[Mol97] Cleve B. Moler. Cleve’s corner: Are we there yet? Zero crossing and event handling for differential equations. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, 1997. 2

pp.  URL http://www.
mathworks.com/company/
newsletter/pdf/97slCleve.
pdf.

Moler:1998:CCG

[Mol98a] Cleve B. Moler. Cleve’s corner: Good vibrations: eigshow helps explain eigenvalues and singular values. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 1998.

1 pp. URL http://www.
mathworks.com/company/
newsletter/clevescorner/
win98cleve.shtml.

Moler:1998:CCT

[Mol98b] Cleve B. Moler. Cleve’s corner: Trigonometry is a complex subject: Revisiting inverse, complex, hyperbolic, floating-point trig functions. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Summer 1998.

1 pp. URL http://www.
mathworks.com/company/
newsletter/clevescorner/
sum98cleve.shtml.

Moler:1999:CCOb

[Mol99a] Cleve B. Moler. Cleve’s corner: Objectively speaking: OOPS is not an apology. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter

1999. URL http://www.
mathworks.com/company/
newsletter/clevescorner/
clevecorner.shtml.

Moler:1999:CCOa

[Mol99b] Cleve B. Moler. Cleve’s corner: Optimally speaking: Optimization Toolbox features new methods for large-scale problems. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 1999.

1 pp. URL http://www.
mathworks.com/company/
newsletter/clevescorner/
sm99cleve.shtml.

Moler:2000:CCF

[Mol00a] Cleve B. Moler. Cleve’s corner: Faster finite Fourier transforms: MATLAB 6 incorporates FFTW. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 2000. 1

pp.  URL http://www.
mathworks.com/company/
newsletter/clevescorner/
winter01_cleve.shtml.

Moler:2000:CCM

[Mol00b] Cleve B. Moler. Cleve’s corner: MATLAB incorporates LAPACK: Increasing the speed and capabilities of matrix computation. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 2000.

1 pp. URL http://www.
mathworks.com/company/
newsletter/clevescorner/
winter2000.cleve.shtml.

Moler:2001:CCN

[Mol01] Cleve B. Moler. Cleve’s corner: Normal behavior: Ziggurat algorithm generates normally distributed random numbers. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Spring 2001. 1

pp.  URL http://www.
mathworks.com/company/
newsletter/clevescorner/
spring01_cleve.shtml.

Moler:2002:CCT

[Mol02a] Cleve B. Moler. Cleve’s corner: The tetragamma function and numerical craftsmanship: MATLAB’s special mathematical functions rely on skills from another era. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, February 2002. 1

pp.  URL http://www.
mathworks.com/company/
newsletter/clevescorner/
winter02_cleve.shtml.

Moler:2002:CCW

[Mol02b] Cleve B. Moler. Cleve’s corner: The world’s largest matrix computation: Google’s PageRank is an eigenvector of a matrix of order 2.7 billion. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, October 2002. 1

pp.  URL http://www.
mathworks.com/company/
newsletter/clevescorner/
oct02_cleve.shtml.

Moler:2003:CCS

[Mol03a] Cleve B. Moler. Cleve’s corner: Stiff differential equations: Stiffness is a subtle, difficult, and important concept in the numerical solution of ordinary differential equations. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, May 2003. 1

                                        pp.  URL http://www.
                                        mathworks.com/company/
                                        newsletters/news_notes/
                                        clevescorner/may03_cleve.
                                        html.

Moler:2003:CCM

[Mol03b] Cleve B. Moler. Cleve’s corner: The MathWorks logo is an eigenfunction of the wave equation. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, November 2003. 1

                                        pp.  URL http://www.
                                        mathworks.com/company/
                                        newsletters/news_notes/
                                        clevescorner/win03_cleve.
                                        html.

Moler:2004:CCN

[Mol04a] Cleve B. Moler. Cleve’s corner: Numerical computing with MATLAB: New textbook available in electronic and print editions. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, June 2004. 1

                                        pp.  URL http://www.
                                        mathworks.com/company/
                                        newsletters/news_notes/
                                        clevescorner/june04_cleve.
                                        html.

Moler:2004:CCO

[Mol04b] Cleve B. Moler. Cleve’s corner: The origins of MATLAB. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, December 2004. 1

pp.  URL http://www.
mathworks.com/company/
newsletters/news_notes/
clevescorner/dec04.html.

Moler:2004:NCM

[Mol04c] Cleve B. Moler. Numerical computing with MAT-LAB. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2004. ISBN 0-89871-560-1 (paperback). xi + 336 pp. LCCN QA297 .M625 2004. US$42.50.

Moler:2004:WPT

[Mol04d] Cleve B. Moler. Work in progress — teaching numerical computing with MAT-LAB. In ????, editor, FIE 2004: 34th Annual Conference on Frontiers in Education, 2004, pages T2H/1– T2H/2 (vol. 1). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN ???? LCCN ????

Moler:2006:CCP

[Mol06a] Cleve B. Moler. Cleve’s corner: Professor SVD. The MathWorks News&Notes, pages 26–29, October 2006.

URL http://www.mathworks.
com/mason/tag/proxy.html?
dataid=8193&fileid=35906.

Moler:2006:CCG

[Mol06b] Cleve B. Moler. Cleve’s corner: The growth of MAT-LAB and The MathWorks over two decades. The MathWorks News&Notes, pages 22–24, January 2006.

                                        URL http://www.mathworks.
                                        com/company/newsletters/
                                        news_notes/clevescorner/
                                        jan06.pdf.

Moller:2007:OLC

[Möl07] Karl Dieter Möller. Optics: learning by computing with examples using Math-CAD, Matlab, Mathematica, and Maple. Undergraduate texts in contemporary physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2007. ISBN 0-387-69492-7, 0-387-26168-0. xvi + 453 pp. LCCN QC381 .M66 2007eb.

Moler:2008:NCM

[Mol08] Cleve B. Moler. Numerical computing with MAT-LAB. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2008. ISBN 0-89871-660-8 (hardcover), 0-89871-560-1 (paperback), 0-89871-795-7 (e-book). xi + 336 pp. LCCN QA297 .M625 2008.

Moler:2011:EM

[Mol11] Cleve Moler. Experiments with MATLAB. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 2011. ISBN ???? ???? pp. LCCN

???? URL http://www.
mathworks.com/moler/exm/

.

Moler:2015:TPJ

[Mol15]  Cleve Moler.  Techni-

cal perspective: Not just a matrix laboratory anymore. Communications of the ACM, 58(10):90, October 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.
acm.org/magazines/2015/
10/192374/fulltext.

Moler:2017:CCQ

[Mol17]  Cleve Moler.  Cleve’s

corner: Quadruple precision, 128-bit floating point arithmetic. Math-Works Web site., May 22, 2017. URL https:

//blogs.mathworks.com/
cleve/2017/05/22/quadruple-
precision-128-bit-floating-
point-arithmetic/.

Moler:2018:BHM

[Mol18a] Cleve Moler. A brief history of MATLAB. Math-Works Web site, 2018. URL

https://www.mathworks.
com/company/technical-
articles/a-brief-history-
of-matlab.html.

Moler:2018:HMU

[Mol18b] Cleve Moler. The historic MATLAB Users’ Guide. Web site, February 5,

                                        2018.  URL https://
                                        blogs.mathworks.com/cleve/
                                        2018/02/05/the-historic-
                                        matlab-users-guide/.

Mohammad:2012:CMT

[MON12] Yasser Mohammad, Yoshimasa Ohmoto, and Toyoaki Nishida. CPMD: a Matlab toolbox for change point and constrained motif discovery. Lecture Notes in Computer Science, 7345: 114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-31087-4_
                                        13/.

Moon:1998:ADA

[Moo98] F. C. Moon. Applied dynamics: with applications to multibody and mechatronic systems. Wiley, New York, NY, USA, 1998. ISBN 0-471-13828-2 (cloth). xi + 492 pp. LCCN QA845 .M657 1998; QA 845 .M657 1998X GERSTM.

Moon:2005:ECC

[Moo05] Todd K. Moon. Error correction coding: mathematical methods and algorithms. Wiley-Interscience, New York, NY, USA, 2005. ISBN 0-471-64800-0 (cloth). xlii + 755 pp. LCCN TA331 .M66 2005. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip055/2004031019.
html.

Moore:2007:ME

[Moo07] Holly Moore. MATLAB for engineers. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2007. ISBN 0-13-187244-3. xiii + 608 pp. LCCN TA345 .M585 2007.

Moore:2008:MPI

[Moo08] Holly Moore. Matlab per l’ingegneria. (Italian) [Matlab for engineers]. Pearson education Italia, Milano, Italia, 2008. ISBN 88-7192-447-9. xiv + 606 pp. LCCN ???? Edited by Marco Mauri. Italian translation by Marco Bassetti.

Moore:2009:ME

[Moo09] Holly Moore. MATLAB for engineers. Pearson Education International, New York, NY, USA, second international edition, 2009. ISBN 0-13-136217-8 (paperback). 624 pp. LCCN TA345 .M585 2009b.

Moore:2011:ME

[Moo11] Holly Moore. MATLAB for engineers. Pearson Prentice Hall, Boston, MA, USA, third international edition, 2011. ISBN 0-273-76416-0 (paperback), 0-13-210325-7. xv + 672 pp. LCCN TA345 .M66 2013. International Edition contributions by Sunil Bhat, Visvesvaraya National Institute of Technology, Nagpur, India.

Moore:2015:MEa

[Moo15a] Holly Moore. MATLAB for engineers. Pearson, Boston, MA, USA, fourth edition, 2015. ISBN 0-13-348597-8. xv + 654 pp. LCCN TA345 .M585 2015.

Moore:2015:MEb

[Moo15b] Holly Moore. MATLAB for engineers. Pearson Education Limited, Harlow, UK, fourth global edition, 2015. ISBN 1-292-06053-0 (paperback). 670 pp. LCCN ???? With contributions by Somitra Kumar Sanadhya (Indraprastha Institute of Information Technology, Delhi).

Morrow:1998:PLA

[Mor98]  Greg Morrow.  A par-

allel linear algebra server for MATLAB-like environments. In ACM [ACM98], page ?? ISBN ???? LCCN ???? URL http://

                                        www.supercomp.org/sc98/
                                        papers/.

Morgan:2000:ASM

[Mor00] Byron J. T. Morgan. Applied stochastic modelling.

Arnold texts in statistics. Arnold, London, UK, 2000. ISBN 0-340-74041-8. xxii + 297 pp. LCCN QA274 .M66 2000.

Morel:2018:PGG

[Mor18]  Pierre Morel.  Gramm:

grammar of graphics plotting in Matlab. Journal of Open Source Software, 3(23):568:1–568:4, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00568.

Park:1997:DSE

[mP97] Seok min Park. Derivation of the state equations in symbolic form from bond graphs using MAT-LAB. Thesis (M.S. in Engin.), University of Texas at Austin, Austin, TX, USA, 1997. x + 124 pp.

Menon:1999:CSL

[MP99] Vijay Menon and Keshav Pingali. A case for sourcelevel transformations in MATLAB. In USENIX [USE99], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.
org/publications/library/
proceedings/dsl99/menon.
html.

Menon:2000:CSL

[MP00] Vijay Menon and Keshav Pingali. A case for sourcelevel transformations in MATLAB. ACM SIGPLAN Notices, 35(1):53–65, January 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mehra:2017:ASC

[MP17] Mani Mehra and Kuldip Singh Patel. Algorithm 986: A suite of compact finite difference schemes. ACM Transactions on Mathematical Software, 44(2):23:1– 23:31, October 2017. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3119905.

Montanari:2018:OCC

[MP18] Mattia Montanari and Nik Petrinic. OpenGJK for C, C# and Matlab: Reliable solutions to distance queries between convex bodies in three-dimensional space. SoftwareX, 7(??):352–355, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711018300591.

Mallesham:2025:VIP

[MPVA25] Harish Nagula Mallesham, Kamana Porwal, Jan Valdman, and Sanjib Kumar Acharya. Vectorized implementation of primal hybrid FEM in MAT-LAB. Computers and Mathematics with Applications, 180(??):144–165, February 15, 2025. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122124005625.

Matarasso:1995:AMC

[MR95] Silvano Matarasso and Tommaso Ruggeri. Analisi Matematica 2: Calcolo Differenziale [English: Math Analysis 2: Differential Calculus]. Progetto Leonardo, Bologna, Italy, 1995. ISBN ???? ???? pp. LCCN ????

Monteiro:2011:CRS

[MR11] M. Teresa T. Monteiro and Helena Sofia Rodrigues. Combining the regularization strategy and the SQP to solve MPCC — a MATLAB implementation. Journal of Computational and Applied Mathematics, 235(18):5348–5356, July 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042710002748.

Massei:2020:HTM

[MRK20] Stefano Massei, Leonardo Robol, and Daniel Kress-

ner. hm-toolbox: MAT-

LAB software for HODLR and HSS matrices. SIAM Journal on Scientific Computing, 42(2):C43–C68, ???? 2020. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Mrozek:1995:NMN

[Mro95a]  B. Mrozek.  Nauczaaie

metod numerycznych w srodowisku pakietu MAT-LAB. In Mrozek [Mro95b], pages 101–?? ISBN 83-86547-04-9. LCCN ????

Mrozek:1995:CMA

[Mro95b] Z. Mrozek, editor. Computer and multimedia aided education: MEDIA ’95: Seminar — May 1995, Cracow, Poland. CEC, ????, 1995. ISBN 83-86547-04-9. LCCN ????

Maciejowski:1994:CST

[MS94] J. M. Maciejowski and M. Szymkat. Containers — a step towards objects with Matlab. In Mattsson et al. [MGC94], pages 277– 284 (or 277–283??). ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Moon:2000:MMA

[MS00] Todd K. Moon and Wynn C. Stirling. Mathematical methods and algorithms for signal processing. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-201-36186-8. xxxvi + 937 pp. LCCN TK5102.9 .M63 2000.

Meurant:2014:FVG

[MS14] Gérard Meurant and Alvise Sommariva. Fast variants of the Golub and Welsch algorithm for symmetric weight functions in Matlab. Numerical Algorithms, 67(3):491–506, November 2014. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11075-013-9804-x.

Manuel:2009:MAO

[MSL09] M. Maria Susai Manuel, A. George Maria Selvam, and M. Paul Loganathan. MATLAB applications of the oscillations of forced neutral difference equations with positive and negative coefficients. International Journal of Pure and Applied Mathematics, 54(4): 521–541, 2009. ISSN 1311-8080.

Moura:2019:UJP

[MSS^{+}19] R. A. R. Moura, M. A. O.

Schroeder, S. J. S. Silva, E. G. Nepomuceno, P. H. N. Vieira, and A. C. S. Lima. The usage of Julia programming in grounding grids simulations : An alternative to MATLAB and Python. In 2019 International Symposium on Lightning Protection (XV SIPDA), pages 1– 4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2019.

McClellen:1998:DFM

[MSY98] James H. McClellen, Ronald W. Schafer, and M. A. (Mark A.) Yoder. DSP first: a multimedia approach. Matlab curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-243171-8. xx + 523 pp. LCCN TK5102.9 .M388 1998.

McClellan:2003:SPF

[MSY03] James H. McClellan, Ronald W. Schafer, and M. A. (Mark A.) Yoder. Signal processing first. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-090999-8, 0-13-065562-7 (CD-ROM). xxi + 489 pp. LCCN TK5102.9 .M389 2003.

Marsaglia:1984:FEI

[MT84] George Marsaglia and Wai Wan Tsang. A fast, easily implemented method for sampling from decreasing or symmetric unimodal density functions. SIAM Journal on Scientific and Statistical Computing, 5(2):349– 359, June 1984. CODEN SI-JCD4. ISSN 0196-5204.

Menon:1997:MIM

[MT97] Vijay Menon and Anne E. Trefethen. MultiMAT-LAB: Integrating Matlab with high performance parallel computing. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN

????  URL http://
www.supercomp.org/sc97/
proceedings/TECH/MENON/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Marsaglia:2000:ZMG

[MT00] George Marsaglia and Wai Wan Tsang. The ziggurat method for generating random variables. Journal of Statistical Software, 5(8):1–7, 2000. CO-DEN JSSOBK. ISSN

???? URL http://www.
jstatsoft.org/v05/i08;
http://www.jstatsoft.
org/v05/i08/rnorrexp.c;
http://www.jstatsoft.
org/v05/i08/updates;
http://www.jstatsoft.org/
v05/i08/ziggurat.pdf.

Munro:2012:BRBa

[Mun12] Chris Munro. Book review:

Advanced engineering mathematics with MATLAB (3rd edn.), by D. Duffy. Pp. 1079. £70-99. 2010. ISBN: 978-1-4398-1624-0 (CRC Press). The Mathematical Gazette, 96(536):364–365, July 2012. CODEN MAGAAS. ISSN 0025-5572.

Munro:2013:CPI

[Mun13] Peter R. T. Munro. Computational Photonics: An Introduction With MATLAB, by Marek S. Wartak, Scope: textbook. Level: undergraduate, postgraduate, practising scientists and engineers. Contemporary Physics, 54 (6):297, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moler:2003:NDW

[MV03] Cleve Moler and Charles Van Loan. Nineteen dubious ways to compute the exponential of a matrix, twenty-five years later. SIAM Review, 45(1):3–49, March 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/41801.

Moskovka:2022:FME

[MV22] Alexej Moskovka and Jan Valdman. Fast MATLAB evaluation of nonlinear energies using FEM in 2D and 3D: Nodal elements. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001345.

Minten:1997:BMB

[MVDV97] Willem Minten, Sven Vranckx, Bart De Moor, and Joos Vandewalle. Bondlab, a MATLAB-based GUI for bond graph modeling. Journal A. Revue A. Tijdschrift A. Zeitschrift A, 38(3):11– ??, ???? 1997. CODEN JR-NAAD. ISSN 0771-1107.

Moose:1997:ASN

[MVM97] R. Moose, H. VanLandingham, and N. McLellan. Analysis and simulation of nonlinear systems using students’ version of MAT-LAB. In IEEE [IEE97d], pages 58–62. ISBN 0-7803-3845-6, 0-7803-3844-8, 0-7803-3846-4, 0-7803-3847-2. LCCN TK7801.I33 1997. IEEE Catalog number 97CH36044.

More:1993:OSG

[MW93] Jorge J. Moré and Stephen J.

Wright. Optimization Software Guide, volume 14 of Frontiers in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-322-6. xii + 154 pp. LCCN QA402.5 .M67 1993.

Moore:1998:FCL

[MY98] Hal G. Moore and Adil Yaqub. A first course in linear algebra with applications. Academic Press, New York, NY, USA, third edition, 1998. ISBN 0-12-505760-1. various pp. LCCN QA184 .M464 1998. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/els032/97046320.
                                        html;   http://www.loc.
                                        gov/catdir/toc/els032/
                                        97046320.html.

Money:2005:AEM

[MY05] James H. Money and Qiang Ye. Algorithm 845: EIGIFP: a MATLAB program for solving large symmetric generalized eigenvalue problems. ACM Transactions on Mathematical Software, 31(2):270– 279, June 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Myridis:2017:BRCb

[Myr17] Nikolaos E. Myridis. Book review: Cosmology with MATLAB, by D. Green. Contemporary Physics, 58 (4):357, 2017. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mark:2003:WCN

[MZ03] Jon W. Mark and Weihua Zhuang. Wireless communications and networking. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-040905-7. xii + 356 pp. LCCN TK5103.2 .M37 2003.

Nixon:2002:FEI

[NA02] Mark S. Nixon and Alberto S. Aguado. Feature extraction and image processing. Newnes, Oxford, 2002. ISBN 0-7506-5078-8. xii + 350 pp. LCCN TA1637 .N58 2002.

Nabney:2002:NAP

[Nab02] Ian Nabney. NETLAB: algorithms for pattern recognitions. Advances in pattern recognition. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 1-85233-440-1. xviii + 420 pp. LCCN TA1637 .N33 2002.

Nahin:2000:DIO

[Nah00] Paul J. Nahin. Duelling idiots and other probability puzzlers. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-00979-1 (hardcover), 0-691-10286-4 (paperback). xviii + 269 pp. LCCN QA273 .N29 2000.

Nahin:2001:SRM

[Nah01] Paul J. Nahin. The science of radio: with MATLAB and Electronics Workbench demonstrations. American Institute of Physics, Woodbury, NY, USA, second edition, 2001. ISBN 0-387-95150-4 (softcover). xliv + 466 pp. LCCN TK6550 .N16

Nahin:2002:DIO

[Nah02] Paul J. Nahin. Duelling idiots and other probability puzzlers: with a new preface by the author. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-10286-4 (paperback). xxviii + 269 pp. LCCN QA273 .N29 2002.

Nahin:2008:DDC

[Nah08] Paul J. Nahin. Digital dice: computational solutions to practical probability problems. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-12698-4 (cloth). xi + 263 pp. LCCN QA273.25 .N34 2008.

Nahin:2011:NCT

[Nah11] Paul J. Nahin. Numbercrunching: taming unruly computational problems from mathematical physics to science fiction. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14425-7 (hardcover), 1-4008-3958-0 (e-book). xxvi + 376 pp. LCCN QC20.7.E4 N34 2011. URL http:

                                        //www.jstor.org/stable/
                                        10.2307/j.ctt7rk7v.

Naidu:2003:OCS

[Nai03] D. S. (Desineni S.) Naidu. Optimal control systems. Electrical engineering textbook series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2003. ISBN 0-8493-0892-5. xxvi + 433 pp. LCCN TJ213 .N2655 2003.

Nakamura:1996:NAG

[Nak96] Shoichiro Nakamura. Numerical Analysis and Graphic Visualization with MAT-LAB. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-051518-3. xi + 477 pp. LCCN T385 .N34 1996.

Nakamura:2002:NAG

[Nak02] Shoichiro Nakamura. Numerical analysis and graphic visualization with MAT-LAB. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2002. ISBN 0-13-065489-2. xiii + 519 pp. LCCN T385 .N34 2002.

Nassar:2001:TDS

[Nas01] Carl Nassar. Telecommunications demystified: a streamlined course in digital communications (and some analog) for EE students and practicing engineers. Demystifying technology series. LLH Technology Publishers, Eagle Rock, VA, US, 2001. ISBN 1-878707-55-8. xix + 353 pp. LCCN TK5101 .N257 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els031/00062902.

                                        html;   http://www.loc.
                                        gov/catdir/toc/els031/
                                        00062902.html.

Nasser:2020:PMT

[Nas20] Mohamed M. S. Nasser.

                                        PlgCirMap: a MATLAB

toolbox for computing conformal mappings from polygonal multiply connected domains onto circular domains. SoftwareX, 11(??):Article 100464, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711019303073.

Nataraj:1994:MTQ

[Nat94]  P. S. V. Nataraj.  A

Matlab toolbox for QFTbased synthesis of linear/ nonlinear lumped and linear distributed systems. In Mattsson et al. [MGC94], pages 513–518 (or 513– 517??). ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Noble:1988:ALA

[ND88] Ben Noble and James W. Daniel. Applied Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1988. ISBN 0-13-041260-0. xvi + 521 pp. LCCN QA184 .N61 1988.

Nkwocha:2005:FPM

[NE05] Fidel Nkwocha and Sebastian Elbaum. Fault patterns in Matlab. ACM SIG-SOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Necsulescu:2002:M

[Nec02] D. S. (Dan S.) Necsulescu.

Mechatronics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-201-44491-7. vii + 311 pp. LCCN TJ163.12 .N43 2002.

Nedoma:1995:SCE

[Ned95] P. Nedoma. Solving Control Engineering Problems with MATLAB by K. Ogata. Automatica: the journal of IFAC, the International Federation of Automatic Control, 31(6):924–??, ???? 1995. CODEN ATCAA9. ISSN 0005-1098.

Neidinger:2010:IAD

[Nei10] Richard D. Neidinger. Introduction to automatic differentiation and MAT-LAB object-oriented programming. SIAM Review, 52(3):545–563, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nelson:2017:RCR

[Nel17] Barry L. Nelson. Replicated computations results (RCR) report for “Green Simulation: Reusing the Output of Repeated Experiments”.

ACM Transactions on Modeling and Computer Simulation, 27(4):24:1–24:??, December 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [FS17].

Newnan:1993:AAC

[New93] B. G. Newnan. Approach to algorithmic/conceptual design using MATLAB. In Anonymous [Ano93e], pages 517–522. ISBN ???? LCCN ???? Two volumes.

Newland:1994:IRV

[New94] D. E. Newland. An Introduction to Random Vibrations, Spectral and Wavelet Analysis. Longman Scientific and Technical, Essex, UK, third edition, 1994. ISBN 0-582-21584-6 (Longman), 0-470-22153-4 (Wiley). xxix + 477 pp. LCCN QA935 .N46 1993.

Neirynck:2018:NBA

[NGKM18] Niels Neirynck, Willy Govaerts, Yuri A. Kuznetsov, and Hil G. E. Meijer. Numerical bifurcation analysis of homoclinic orbits embedded in one-dimensional manifolds of maps. ACM Transactions on Mathematical Software, 44(3):25:1– 25:19, April 2018. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3134443.

Ni:2022:BRE

[Ni22] Yang Ni. Book review:

Exploratory Data Analysis with MATLAB, 3rd ed., by Wendy L. Martinez, Angel R. Martinez, and Jeffrey L. Solka. The American Statistician, 76 (1):85–86, 2022. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/
doi/full/10.1080/00031305.
2021.2020000.

Nirmalakhandan:2002:MTE

[Nir02] N. Nirmalakhandan. Modeling tools for environmental engineers and scientists. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2002. ISBN 1-56676-995-7. xi + 312 pp. LCCN GE45.M37 K43 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy034/2001052467.html.

Nise:1995:CSE

[Nis95] Norman Nise. Control Systems Engineering. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, second edition, 1995. ISBN 0-8053-5424-7. xxiv + 853 pp. LCCN TJ213 .N497 1995.

Nise:2004:CSE

[Nis04] Norman S. Nise. Control systems engineering. Wiley, New York, NY, USA, fourth edition, 2004. ISBN 0-471-44577-0. xxii + 983 pp. LCCN TJ213 .N497 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley041/2003278223.
html;   http://www.loc.
gov/catdir/toc/wiley041/
2003278223.html.

Novoselsky:2017:AOM

[NK17] Alexander Novoselsky and Eugene Kagan. Algorithm 974: The OutlierLib — a MATLAB library for outliers’ detection. ACM Transactions on Mathematical Software, 43(4):38:1– 38:3, March 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Nguyen:2018:AGB

[NK18] Duc T. Nguyen and Blair

Kaneshiro. AudExpCreator:

a GUI-based Matlab tool for designing and creating auditory experiments with the Psychophysi’s Toolbox. SoftwareX, 7(??):328– 334, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018301985.

Nekoogar:1999:DCU

[NM99] Farzad Nekoogar and Gene Moriarty. Digital control using digital signal processing. Prentice-Hall information and system sciences series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-089103-7. xiv + 433 pp. LCCN TJ223.M53 N45 1999.

Ntaganda:2006:DHC

[NMS^{+}06] J. Marie Ntaganda, Benjamin Mampassi, Blaise Some, Djibril Seck, and Abdoulaye Samb. Design of the human cardiovascularrespiratory control system using MATLAB. Far East Journal of Applied Mathematics, 22(2):137–153, 2006. ISSN 0972-0960.

Nesterov:1994:IPA

[NN94] Yurii Nesterov and Arkadi Nemirovskii. Interior-Point Polynomial Algorithms in Convex Programming, volume 13 of SIAM studies in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-319-6. ix + 405 pp. LCCN QA402.5 .N4615 1994.

Norgaard:2000:NNM

[Nør00] Magnus Nørgaard, editor.

Neural networks for modelling and control of dynamic systems: a practitioner’s handbook. Advanced textbooks in control and signal processing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-85233-227-1. xiv + 246 pp. LCCN QA76.87 .N4847 2000.

Norton:2004:DMI

[Nor04] Robert L. Norton. Design of machinery: an introduction to the synthesis and analysis of mechanisms and machines. McGraw-Hill series in mechanical engineering. McGraw-Hill, New York, NY, USA, third edition, 2004. ISBN 0-07-286447-8 (set), 0-07-247046-1, 0-07-121496-8 (ISE), 0-07-247049-6 (CD-ROM). xxi + 858 pp. LCCN TJ175 .N58 2004.

Norton:2005:MDI

[Nor05] Robert L. Norton. Machine design: an integrated approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2005. ISBN 0-13-148190-8. ???? pp. LCCN TJ230 .N64 2005.

Nikias:1993:HOS

[NP93] Chrysostomos Nikias and Athina P. Petropulu. Higher Order Spectra Analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-678210-8. xxii + 537 pp. LCCN TK5102.9 .N54 1993. Naraghi-Pour:1996:PTE

[NP96] Mort Naraghi-Pour, editor. Proceedings of the Twentyeighth Southeastern Symposium on System Theory, March 31–April 2, 1996, Baton Rouge, Louisiana, volume 28. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7352-4, 0-8186-7350-4 (??invalid ISBN checksum??). ISSN 0094-2898. LCCN TA 168 S727p 1996.

Nagy:2004:IMI

[NPP04] James G. Nagy, Katrina Palmer, and Lisa Perrone. Iterative methods for image deblurring: a Matlab object-oriented approach. Numerical Algorithms, 36 (1):73–93, May 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp009.kluweronline.
com/IPS/content/ext/x/
J/5058/I/52/A/5/abstract.
htm.

Nedialkov:2015:ADM

[NPT15] Nedialko S. Nedialkov, John D. Pryce, and Guangning Tan. Algorithm 948: DAESA — a Matlab tool for structural analysis of differential-algebraic equations: Software. ACM Transactions on Mathematical Software, 41(2):12:1– 12:14, January 2015. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Nilsson:1996:EC

[NR96a] James William Nilsson and Susan A. Riedel. Electric Circuits. Addison-Wesley, Reading, MA, USA, fifth edition, 1996. ISBN 0-201-55707-X. xxxii + 983 pp. LCCN TK454 .N54 1996.

Nilsson:1996:UCT

[NR96b] James William Nilsson and Susan A. Riedel. Using Computer Tools for Electric Circuits. Addison-Wesley, Reading, MA, USA, fifth edition, 1996. ISBN 0-201-84707-8. ix + 143 pp. LCCN TK454.N542 1996.

Noergaard:1996:NTM

[NRHP96] M. Noergaard, O. Ravn, L. K. Hansen, and N. K. Poulsen. The NNSYSID toolbox — a MATLAB toolbox for system identification with neural networks. In IEEE [IEE96d], pages 374– 379. ISBN 0-7803-3032-3, 0-7803-3033-1. LCCN TJ212.2.I32495 1996. IEEE catalog number 96TH8136.

Noergaard:1996:NCT

[NRPH96] M. Noergaard, O. Ravn, N. K. Poulsen, and L. K. Hansen. NNCTRL — a CANCSD toolkit for MAT-LAB. In IEEE [IEE96d], pages 368–373. ISBN 0-7803-3032-3, 0-7803-3033-1. LCCN TJ212.2.I32495 1996. IEEE catalog number 96TH8136.

Nguyen:2014:CLM

[NSXZ14] Huy Nguyen, Hao Shi, Jie Xu, and Shiwei Zhang. CPMC-Lab: a Matlab package for Constrained Path Monte Carlo calculations. Computer Physics Communications, 185(12):3344– 3357, December 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465514002707.

Nyholm:2008:SAA

[Nyh08] Ken Nyholm. Strategic Asset Allocation in Fixed Income Markets: a Matlab based user’s guide. The Wiley finance series. Wiley, New York, NY, USA, 2008. ISBN 0-470-75362-5. xv + 167 pp.

Ozgun:2011:PMB

[OAKS11] Ozlem Ozgun, Gökhan Apaydin, Mustafa Kuzuoglu, and Levent Sevgi. PETOOL: MATLAB-based one-way and two-way split-step parabolic equation tool for radiowave propagation over variable terrain. Computer Physics Communications, 182(12):2638–2654, December 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465511002669.

Olivier:1993:CAE

[OB93] G. Olivier and R.-P. Bouchard, editors. Computational aspects of electromechanical energy converters and drives: 4th International conference — 1993 Jul: Montréal, Québec, Canada. École Poly Montréal, Montréal, Québec, Canada, 1993. ISBN ???? LCCN ????

OBrien:2013:SBRe

[O’B13] Carl M. O’Brien. Short book review: Statistics for Bioengineering Sciences: With MATLAB and Win-BUGS Support by Brani Vidakovic. International Statistical Review = Revue Internationale de Statistique, 81(3):471–472, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OSullivan:2019:RNL

[OBCG19] S. O’Sullivan, R. E. Bird, W. M. Coombs, and S. Giani. Rapid non-linear finite element analysis of continuous and discontinuous Galerkin methods in MATLAB. Computers and Mathematics with Applications, 78(9):3007–3026, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122119301312.

Oellerich:2023:RSR

[OBD23] Jan Oellerich, Keno Jann Büscher, and Jan Philipp Degel. RooTri: a simple and robust function to approximate the intersection points of a 3D scalar field with an arbitrarily oriented plane in MATLAB. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/
16/9/409.

Otto:2005:IPN

[OD05] Stephen R. Otto and James P. Denier. An introduction to programming and numerical methods in MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-85233-919-5. xvi + 463 pp. LCCN ???? UK£25.00.

Orsak:1995:CSE

[OE95] Geoffrey C. Orsak and Delores M. Etter. Collaborative SP education using the Internet and MAT-LAB. IEEE Signal Processing Magazine, 12(6):23–32, November 1995. CODEN ISPRE6. ISSN 1053-5888 (print), 1558-0792 (electronic).

ODonnell:1992:MIE

[OF92] J. R. O’Donnell and D. K.

Frederick. MATLAB implementation of the Extended Ash Algorithm for finding root loci. In Barker [Bar92], pages 341–346. ISBN 0-08-041269-6. LCCN TJ213 .C57 1992.

Oldaker:1995:IPS

[OG95] R. S. Oldaker and J. M. Gilbert. An integrated package for the simulation and practical evaluation of discrete control systems running within the MATLAB environment. In Anonymous [Ano95b], pages 7/1– 7/3. ISBN ???? LCCN ????

Ogata:1994:DLC

[Oga94a] Katsuhiko Ogata. Designing Linear Control Systems with MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-293226-1. v + 226 pp. LCCN TJ213 .O27 1994.

Ogata:1994:SCE

[Oga94b] Katsuhiko Ogata. Solving Control Engineering Problems with MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-045907-0. v + 359 pp. LCCN TJ213 .O29 1994.

Ogata:1995:DCS

[Oga95] Katsuhiko Ogata. Discrete-Time Control Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1995. ISBN 0-13-034281-5. xi + 745 pp. LCCN QA402 .O4 1995.

Ogata:1997:MCE

[Oga97] Katsuhiko Ogata. Modern Control Engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1997. ISBN 0-13-227307-1. xi + 997 pp. LCCN TJ213.O28 1997.

Ogata:1998:SD

[Oga98] Katsuhiko Ogata. Systems Dynamics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-675745-6. ix + 758 pp. LCCN TA342 .O35 1998.

Ogata:2002:MCE

[Oga02] Katsuhiko Ogata. Modern control engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2002. ISBN 0-13-060907-2. xi + 964 pp. LCCN TJ213 .O28 2002.

Ogata:2004:SD

[Oga04] Katsuhiko Ogata. System dynamics. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2004. ISBN 0-13-142462-9. ix + 768 pp. LCCN TA342 .O35 2004.

Ogata:2008:MCE

[Oga08] Katsuhiko Ogata. MAT-LAB for control engineers. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-615077-2. viii + 433 pp. LCCN ????

Ogata:2010:SNT

[Oga10] Katsuhiko Ogata. Seigyo no tame no mattorabu = Matlab for control engineers. Tōkyō Denki Daigaku Shuppankyoku, Tōkyō, Japan, 2010. ISBN 4-501-32760-X. vi + 447 pp. LCCN ????

Oguni:1995:MUM

[Ogu95] Tsutomu Oguni. MAT-LAB and Its Usage: Modern Applied Mathematics and Computer Graphics. The Science Corporation, ??, Japan, 1995. ISBN 4-7819-0763-6. ???? pp. LCCN ???? In Japanese.

Oner:2017:JCS

[ÖK17] Metin Öner and İpek Deveci Kocakoç. JMASM 49: a compilation of some popular goodness of fit tests for normal distribution: Their algorithms and MATLAB codes (MATLAB). Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/
jmasm/article/view/956.

Ong:1998:DSE

[Ong98] Chee-Mun Ong. Dynamic simulation of electric machinery: using MATLAB/ SIMULINK. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-723785-5. xv + 626 pp. LCCN TK2391 .O533 1998.

Onofri:2001:ECM

[Ono01] Enrico Onofri. Elementary celestial mechanics using Matlab. Computing in Science and Engineering, 3(6):48–53, November/ December 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6048abs.htm;   http:/
/dlib.computer.org/cs/
books/cs2001/pdf/c6048.
pdf.

Orfanidis:1996:ISP

[Orf96] Sophocles J. Orfanidis. Introduction to Signal Processing. Prentice Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-209172-0. xvi + 798 pp. LCCN TK5102.9 .O73 1996.

Olver:2006:ALA

[OS06] Peter J. Olver and Cheri Shakiban. Applied Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-147382-4. 620 (est.) pp. LCCN ????

Oppenheim:1997:SS

[OWN97] Alan V. Oppenheim, Alan S.

Willsky, and Syed Hamid Nawab. Signals and systems. Prentice-Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1997. ISBN 0-13-814757-4. xxx + 957 pp. LCCN QA402 .O63 1997.

Ozbay:2000:IFC

[Özb00] Hitay Özbay. Introduction to feedback control theory. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-1867-X. 217 (est.) pp. LCCN TJ216 .O97 2000.

Ozkurt:2023:PUF

[Ozk23] Nalan Ozkurt. PrepAnnECG: a user friendly MATLAB ECG preprocessing and annotation GUI for health professionals. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002856.

Polking:1999:ODE

[PA99] John C. Polking and David Arnold. Ordinary differential equations using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-011381-6. xii + 224 pp. LCCN QA371.5.D37 P65 1999. Rev. ed. of: MAT-LAB manual, ordinary differential equations (1995).

Polking:2004:ODE

[PA04] John C. Polking and David Arnold. Ordinary differential equations using MAT-LAB. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, third edition, 2004. ISBN 0-13-145679-2. 247 (est.) pp. LCCN QA371.5.D37 P65 2004.

Peng:2011:SSM

[PA11] Jyh-Ying Peng and John A. D. Aston. The state space models toolbox for MATLAB. Journal of Statistical Software, 41(6): ??, May 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v41/i06.

Paarmann:2001:DAA

[Paa01] L. D. Paarmann. Design and analysis of analog filters: a signal processing perspective, volume 617 of SECS. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 2001. ISBN 0-7923-7373-1. xii + 439 pp. LCCN TK5102.9 .P33 2001.

Pace:2000:ATD

[Pac00] Phillip E. Pace. Advanced techniques for digital receivers. Artech House radar library. Artech House Inc., Boston, MA, USA, 2000. ISBN 1-58053-053-2 (paperback). xvii + 430 pp. LCCN TK5103.7 .P334 2000.

Pace:2004:DCL

[Pac04] Phillip E. Pace. Detecting and classifying low probability of intercept radar. Artech House radar library. Artech House Inc., Boston, MA, USA, 2004. ISBN 1-58053-322-1. xxii + 455 pp. LCCN TK6592.L67 P33 2004.

Prasad:2011:ACM

[PAG11] Ashwin Prasad, Jayvant Anantpur, and R. Govindarajan. Automatic compilation of MATLAB programs for synergistic execution on heterogeneous processors. ACM SIG-PLAN Notices, 46(6):152– 163, June 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palm:1998:IME

[Pal98a] William J. (William John) Palm. Introduction to MATLAB for engineers. McGraw-Hill’s BEST–basic engineering series and tools. WCB/McGraw-Hill, Boston, MA, USA, 1998. ISBN 0-07-047328-5. xx + 316 pp. LCCN QA297.P32 1998.

Palm:1998:MAC

[Pal98b] William J. (William John) Palm. Modeling, analysis, and control of dynamic systems. Wiley, New York, NY, USA, second edition, 1998.

ISBN 0-471-07370-9, 0-471-35412-0 (Solutions manual). x + 853 pp. LCCN TJ213.7 .P34 1998.

Palm:2001:IME

[Pal01] William J. (William John)

Palm III. Introduction to MATLAB 6 for engineers. Basic engineering series and tools. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-234983-2 (paperback). xix + 600 pp. LCCN TA345 .P35 2001.

Palm:2005:SD

[Pal05a] William J. (William John)

Palm. System dynamics. McGraw-Hill series in mechanical engineering. McGraw-Hill Higher Education, Boston, MA, USA, 2005. ISBN 0-07-301603-9, 0-256-11449-8. xv + 928 pp. LCCN TJ213 .P228 2005.

Palm:2005:IME

[Pal05b] William J. (William John)

Palm III. Introduction to MATLAB 7 for engineers. McGraw-Hill, New York, NY, USA, second edition, 2005. ISBN 0-07-254818-5, 0-07-292242-7. xiii + 682 pp. LCCN TA345 .P355 2005.

Pang:1989:ESD

[Pan89] Grantham K. H. Pang. Expert system development facility in a MATLAB-derived control environment. In 1989 IEEE Control Systems Society Workshop on Computer Aided Control System Design CACSD (Dec 16 1989: Tampa, FL, USA), pages 132–137. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89TH0270-9.

Pao:1999:EAI

[Pao99] Yen-Ching C. Pao. Engineering Analysis: Interactive Methods and Programs with FORTRAN, Quick-BASIC, MATLAB, and Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-2016-X. 360 (est.) pp. LCCN TA345 .P36 1999. US$79.95.

Pao:2001:EAI

[Pao01] Y. C. Pao. Engineering analysis interactive methods and programs with FORTRAN, QuickBASIC, MATLAB, and Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. Computer data. Title from title screen. Available in ENGnetBASE (CRC Press).

Passino:2004:BOC

[Pas04] Kevin M. Passino. Biomimicry for optimization, control, and automation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 1-85233-804-0. xxxi + 926 pp. LCCN TS156.8 .P245 2004.

Patterson:1994:MCB

[Pat94] B. W. Patterson, editor. Modeling and control in biomedical systems: IFAC Symposium, Galveston, Texas, USA, 27– 30 March 1994. Pergamon Press, New York, NY, USA, 1994. ISBN 0-08-042224-1. LCCN QP39 .M63 1994.

Patzold:2002:MFC

[Pät02] Matthias Pätzold. Mobile fading channels. Wiley, New York, NY, USA, 2002. ISBN 0-471-49549-2. x + 418 pp. LCCN TK6570.M6 P387 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley036/2001056785.
html;   http://www.loc.
gov/catdir/toc/wiley021/
2001056785.html.

Paul:2000:FCA

[Pau00] Clayton R. Paul. Fundamentals of circuit analysis. Wiley, New York, NY, USA, 2000. ISBN 0-471-37195-5. xv + 501 pp. LCCN TK454 .P285 2000.

Paul:2001:FEC

[Pau01] Clayton R. Paul. Fundamentals of electric circuit analysis. Wiley, New York, NY, USA, 2001. ISBN 0-471-37195-5 (cloth). xv + 501 pp. LCCN TK454 .P285 2001; TK 454 .P285 2001X ENGI. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/onix06/00020738.html.

Pave:1993:MSP

[Pav93] A. Pave, editor. Modelling and simulation 1993: Proceedings of the 1993 European Simulation Multiconference, June 7–9, 1993, Ecole Normale Supérieure de Lyon. Society for Computer Simulation, San Diego, CA, USA, 1993. ISBN 1-56555-056-0. LCCN QA76.9.C65 E97 1993.

Poon:2001:COI

[PB01] Ting-Chung Poon and Partha P. Banerjee. Contemporary optical image processing with MATLAB. Elsevier, Amsterdam, The Netherlands, 2001. ISBN 0-08-043788-5 (hardcover). viii + 262 pp. LCCN TA1637 .P65 2001.

Polking:2006:DE

[PBA06] John Polking, Albert Boggess, and David Arnold. Differential Equations. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2006. ISBN 0-13-143738-0. ???? pp. LCCN ????

Psarras:2022:LAM

[PBB22] Christos Psarras, Henrik Barthels, and Paolo Bientinesi. The linear algebra mapping problem. Current state of linear algebra languages and libraries. ACM Transactions on Mathematical Software, 48(3):26:1– 26:??, September 2022. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
10.1145/3549935.

Patiris:2007:TIM

[PBI07] D. L. Patiris, K. Blekas, and K. G. Ioannides. TRIAC II. A Matlab code for track measurements from SSNT detectors. Computer Physics Communications, 177(3):329–338, August 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465507002196.

Poppe:2013:CMO

[PC13] Koen Poppe and Ronald Cools. CHEBINT: a MAT-LAB/Octave toolbox for fast multivariate integration and interpolation based on Chebyshev approximations over hypercubes. ACM Transactions on Mathematical Software, 40(1):2:1–2:13, September 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Peebles:2001:PRV

[Pee01] Peyton Z. Peebles. Probability, random variables, and random signal principles. McGraw-Hill series in electrical and computer engineering. McGraw-Hill, New York, NY, USA, fourth edition, 2001. ISBN 0-07-366007-8. xviii + 462 pp. LCCN TA340 .P43 2001. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/mh021/00034881.
                                        html;   http://www.loc.
                                        gov/catdir/toc/mh021/00034881.
                                        html.

Penney:2004:LAI

[Pen04] Richard C. Penney. Linear algebra: ideas and applications. Wiley-Intersci-ence, New York, NY, USA, second edition, 2004. ISBN 0-471-67620-9 (cloth). xvi + 436 pp. LCCN QA184.2 .P46 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley042/2004048071.
                                        html;   http://www.loc.
                                        gov/catdir/toc/wiley041/
                                        2004048071.html.

Perdikaris:1991:CCS

[Per91] George A. Perdikaris. Computer Controlled Systems: Theory and Applications, volume 8 of International series on microprocessorbased systems engineering. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1991. ISBN 0-7923-1422-0. x + 480 pp. LCCN TJ223.M53 P43 1991.

Persson:1993:MTP

[Per93]  P. Persson.  A Mat-

lab toolbox for PID controller design. In IEEE [IEE93a], pages 230–234. ISBN 0-7803-1223-6. LCCN TJ212.2 .A75 1993.

Part-Enander:1998:AFM

[PES98] Eva Pärt-Enander and Anders Sjöberg. Användarhandledning för MATLAB 5. (Swedish) [User guide for MATLAB 5]. Institutionen för teknisk databehandling, Uppsala, Sweden, 1998. ISBN 91-506-1292-1. 2 + viii + 485 pp. LCCN ????

Part-Enander:1999:MH

[PES99] Eva Pärt-Enander and Anders Sjöberg. The MAT-LAB 5 handbook. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-39845-

1. xv + 559 pp. LCCN QA297.P47 1999. Rev. ed. of: The MATLAB handbook (1996).

Part-Enander:1996:MH

[PESMI96] Eva Pärt-Enander, Anders Sjöberg, Bo Melin, and Pernilla Isaksson. The MATLAB Handbook. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-87757-0. xv + 423 pp. LCCN QA297 .M386 1996.

Petkovska:1996:SNS

[Pet96] M. Petkovska. Simulation of non-linear system dynamics using Matlab with Simulink. Hemijska industrija, 50(2): 65–??, ???? 1996. ISSN 0367-598X.

Popovici:2007:MEC

[PF07] Constantin I. Popovici and Constantin Filipescu. Matlab evaluation of the Γ_{k}^{n}(x) coefficients for PDE solving by wavelet-Galerkin approximation. Libertas Mathematica, 27:85–93, 2007. ISSN 0278-5307.

Pachamanova:2010:SOF

[PF10] Dessislava A. Pachamanova and Frank J. Fabozzi. Simulation and optimization in finance: modeling with MATLAB, @RISK, or VBA, volume 173 of Frank J. Fabozzi series. Wiley, New York, NY, USA, 2010. ISBN 0-470-37189-7. ???? pp. LCCN HG106 .P33 2010. URL http:

                                        //catalogimages.wiley.
                                        com/images/db/jimages/
                                        9780470371893.jpg.

Pfeiffer:1995:BCP

[Pfe95] Paul E. Pfeiffer. Bookware Companion Problems Book:

Basic Probability Topics Using MATLAB. Tom Robbins’ BookWare companion series. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1995. ISBN 0-534-94536-8. xx + 185 pp. LCCN QA273.19.E4P38 1995. US$30.95.

Padulo:2008:RAC

[PFG08] Mattia Padulo, Shaun A.

Forth, and Marin D. Guenov. Robust aircraft conceptual design using automatic differentiation in Matlab. In Bischof et al. [BBH^{+}08], pages 271–280. CODEN LNCSA6. ISBN 3-540-68935-4 (print), 3-540-68942-7 (e-book). ISSN 1439-7358. LCCN QA304 .I58 2008. URL http://

link.springer.com/content/
pdf/10.1007/978-3-540-
68942-3_24.

Paddison:1994:MDS

[PGBG94] J. E. Paddison, R. M.

Goodall, J. Bals, and G. Grubel. Multi-objective design study for a Maglev suspension controller using the database ANDECS– MATLAB environment. In Mattsson et al. [MGC94], pages 239–246. ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Pana:1994:MTF

[PH94] T. Pana and Y. Hori. A MATLAB toolbox for fieldoriented induction motor systems design. In Anonymous [Ano94b], pages 1203– 1210. ISBN ???? LCCN ???? Three volumes.

Phillips:1996:FCS

[PH96] Charles L. Phillips and Royce D. Harbor. Feedback Control Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1996. ISBN 0-13-371691-0. xiv + 683 pp. LCCN TJ216 .P465 1996.

Phillips:2000:FCS

[PH00] Charles L. Phillips and Royce D. Harbor. Feedback control systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2000. ISBN 0-13-949090-6. xiii + 658 pp. LCCN TJ216 .P54 2000.

Phadke:1999:HEE

[Pha99] Arun G. Phadke, editor. Handbook of electrical engineering calculations, volume 106 of Electrical engineering and electronics. Marcel Dekker, New York, NY, USA, 1999. ISBN 0-8247-1955-7. vii + 313 pp. LCCN TS168 .H36 1999.

Phillips:2000:HFE

[Phi00] Chandler A. Phillips. Human factors engineering. Wiley, New York, NY, USA, 2000. ISBN 0-471-24089-3 (cloth). xi + 564 pp. LCCN TA166 .P53 2000.

Pauca:1995:UGP

[PHL95] V. Paúl Pauca, J. Hollingsworth, and K. Liu. User’s guide for the parallel toolbox for MATLAB. Technical report, Wake Forest University, Winston-Salem, NC, USA, 1995. URL ftp:

//ftp.cs.duke.edu/pub/
pauca/papers/pt_manual.
ps.gz.

Peinado:2012:SSD

[PIAH12] Jesus Peinado, Jacinto J.

Ibáñez, Enrique Arias, and Vicente Hernández. Speeding up solving of differential matrix Riccati equations using GPGPU computing and MATLAB. Concurrency and Computation: Practice and Experience, 24(12): 1334–1348, August 25, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierre:1996:SMT

[Pie96a] D. A. Pierre. Supplementary MATLAB tools for systems and control education. In Iskander et al. [I^{+}96], pages 1215–1218. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Pierret:1996:SDF

[Pie96b] Robert F. Pierret. Semiconductor Device Fundamentals. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-54393-1. xxiii + 792 pp. LCCN TK7871.85 .P484 1996.

Pierce:2005:BRI

[Pie05a] John G. Pierce. Book review: Introduction to MAT-LAB with Numerical Preliminaries. Journal of Statistical Software, 14(BR-3): 1–3, August 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v14/b03.

Pietruszka:2005:MIG

[Pie05b] Wolf D. Pietruszka. MAT-LAB in der Ingenieurpraxis. (German) [MAT-LAB in Engineering]. Teubner, Stuttgart, Germany; Leipzig, Germany, 2005. ISBN ???? xi + 320 pp. LCCN ????

Plasil:1997:STP

[PJ97] Frantisek Plasil and Keith (Keith G.) Jeffery, editors. SOFSEM’97: theory and practice of informatics: 24th Seminar on Current Trends in Theory and Practice of Informatics, Milovy, Czech Republic, November 22–29, 1997:proceedings, volume 1338 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63774-5 (softcover). LCCN QA76.751.S62 1997.

Peeva:2004:FRC

[PK04] Ketty Peeva and Yordan Kyosev. Fuzzy relational calculus: theory, applications and software (with CD-ROM), volume 22 of Advances in fuzzy systems. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2004. ISBN 981-256-076-9. ???? pp. LCCN QA248 .P36 2004.

Piiroinen:2008:EDM

[PK08] Petri T. Piiroinen and Yuri A. Kuznetsov. An event-driven method to simulate Filippov systems with accurate computing of sliding motions. ACM Transactions on Mathematical Software, 34(3):13:1–13:24, May 2008. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Penny:1995:NMU

[PL95] John E. T. Penny and George R. Lindfield. Numerical Methods Using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-030966-4. xii + 328 pp. LCCN QA297 .P45 1995. US$40.00.

Penny:2000:NMU

[PL00] J. E. T. (John E. T.) Penny and G. R. (George R.) Lindfield. Numerical methods using Matlab. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2000. ISBN 0-13-012641-1 (paper). xiii + 482 pp. LCCN QA297 .P45 2000.

Pohjolainen:1994:EME

[PMA94] S. Pohjolainen, J. Multisilta, and K. Antchev. Examples of Matlab in engineering education. NATO ASI series. Series F, Computer and system sciences, 132(??):83–92, ???? 1994. CODEN NASFEG. ISSN 0258-1248.

Paprotny:2020:BMT

[PMNWR20] Dominik Paprotny, Oswaldo Morales-Nápoles, Daniël T. H. Worm, and Elisa Ragno. BANSHEE — a MATLAB toolbox for nonparametric Bayesian networks. SoftwareX, 12 (??):Article 100588, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711020303010.

Papadimitriou:2014:MLS

[PMTL14] Stergios Papadimitriou, Seferina Mavroudi, Kostas Theofilatos, and Spiridon Likothanasis. MATLABlike scripting of Java scientific libraries in ScalaLab. Scientific Programming, 22 (3):187–199, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Phillips:1995:DCS

[PN95] Charles L. Phillips and H. Troy Nagle. Digital Control System Analysis and Design. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1995. ISBN 0-13-309832-X. xv + 685 pp. LCCN TJ223.M53 P47 1995.

Poznyak:2000:SLC

[PNGR00] Alexander S. Poznyak, K. Najim, and E. Gomez-Ramirez. Self-learning control of finite Markov chains, volume 4 of Control engineering. Marcel Dekker, New York, NY, USA, 2000. ISBN 0-8247-9429-X. xiii + 298 pp. LCCN QA274.7 .P69 2000.

Pilia:2021:EOS

[PNL^{+}21] Nicolas Pilia, Claudia Nagel, Gustavo Lenis, Silvia Becker, Olaf Dössel, and Axel Loewe. ECGdeli: an open source ECG delineation toolbox for MAT-LAB. SoftwareX, 13(??): Article 100639, January 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711020303526.

Pryce:2015:DMT

[PNT15] John D. Pryce, Nedialko S. Nedialkov, and Guangning Tan. DAESA — a Matlab tool for structural analysis of differential-algebraic equations: Theory. ACM Transactions on Mathematical Software, 41(2):9:1–9:20, January 2015. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Pole:1992:AMS

[Pol92] Stephen J. Pole. Application of Matlab and Simulink to a low-lift sump system. Measurement and control, 25(8):234–239, October 1992. CODEN MEACBX. ISSN 0020-2940.

Polking:1995:MMO

[Pol95] John C. Polking. MAT-LAB Manual for Ordinary Differential Equations. MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-133944-3. xiii + 205 pp. LCCN QA71.15.D37P65 1995.

Porat:1997:CDS

[Por97] Boaz Porat. A course in digital signal processing. Wiley, New York, NY, USA, 1997. ISBN 0-471-14961-6. xxvi + 602 pp. LCCN TK5102.9 .P66 1997.

Portioli:2023:DOB

[Por23] Francesco P. A. Porti-

oli.  DynABlock 2D: an

optimization-based MAT-LAB application for rocking dynamics, nonlinear static and limit analysis of masonry block structures. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023001504.

Pous:1996:UMR

[POVD96] C. Pous, A. Oller, J. Vehi, and J. I. De la Rosa. Using Matlab real-time workshop in teaching control design techniques. In Zalewski [Zal96], pages 153–

158. ISBN 0-8186-7649-3, 0-8186-7651-5. LCCN QA76.54 .R4294 1996. IEEE Computer Society Press order number PR07649. IEEE Order Plan catalog number 96TB1000060.

Pozrikidis:2005:IFS

[Poz05] C. Pozrikidis. Introduction to Finite and Spectral Element Methods Using MATLAB. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-529-7. xxv + 653 pp. LCCN TA347.F5 P7 2005.

Pozrikidis:2014:IFS

[Poz14] Constantine Pozrikidis. Introduction to finite and spectral element methods using MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2014. ISBN 1-4822-0915-2 (hardcover), 1-4822-5074-8 (ePub e-book), 1-4822-0916-0 (PDF e-book), 1-4822-0917-9 (VitalBook ebook). xxv + 804 pp. LCCN TA347.F5 P7 2014.

Phillips:1995:SST

[PP95] Charles L. Phillips and John M. Parr. Signals, Systems and Transforms. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-795253-8. xxi + 708 pp. LCCN TK5102.9.P47 1995.

Pawletta:1995:CPS

[PPD95] S. Pawletta, T. Pawletta, and W. Drewelow. Comparison of parallel simulation techniques — MAT-LAB/PSI. Simulation News Europe, 13:38–39, 1995.

Pantelic:2018:SEP

[PPL^{+}18] Vera Pantelic, Steven Postma, Mark Lawford, Monika Jaskolka, Bennett Mackenzie, Alexandre Korobkine, Marc Bender, Jeff Ong, Gordon Marks, and Alan Wassyng. Software engineering practices and Simulink: bridging the gap. International Journal on Software Tools for Technology Transfer (STTT), 20 (1):95–117, February 2018.

CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL

https://link.springer.
com/article/10.1007/s10009-
017-0450-9.

Phillips:2003:SST

[PPR03] Charles L. Phillips, John M.

Parr, and Eve A. (Eve Ann) Riskin. Signals, systems, and transforms. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2003. ISBN 0-13-041207-4. xx + 765 pp. LCCN TK5102.9 .P47 2003.

Popovici:2006:DMC

[PPS06] Constantin Popovici, Emilia Popovici, and Gheorghe Sufaru. Determination by MATLAB of the coefficients Λ_{k}^{m,n}(x) for solving PDEs by the approximation method wavelet-Galerkin. Stud. Cercet. Ştiinţ. Ser. Mat. Univ. Bacău, 16S:215– 222, 2006. ISSN 1224-2519.

Pachon:2010:PSC

[PPT10] Ricardo Pachón, Rodrigo B.

Platte, and Lloyd N. Trefethen. Piecewise-smooth chebfuns. IMA Journal of Numerical Analysis, 30(4): 898–916, October 2010. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic). URL http:/

/imajna.oxfordjournals.
org/content/30/4/898.full.
pdf+html.

Passmore:2002:LPN

[PR02] Tim Passmore and A. J. Roberts. Low Prandtl number fluid convection modelled using symbolic algebra (REDUCE) and Matlab. The ANZIAM Journal, 44((C)):C590–C626, 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

Poularikas:2006:AFP

[PR06] Alexander D. Poularikas and Zayed M. Ramadan. Adaptive filtering primer with MATLAB. CRC/Taylor and Francis, Boca Raton, FL, USA, 2006. ISBN 0-8493-7043-4. 223 pp. LCCN TK7872.F5 P68 2006 UCB.

Patterson:2014:GIM

[PR14] Michael A. Patterson and Anil V. Rao. GPOPS-II: a MATLAB software for solving multiple-phase optimal control problems using hpadaptive Gaussian quadrature collocation methods and sparse nonlinear programming. ACM Transactions on Mathematical Software, 41(1):1:1–1:37, October 2014. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Pratap:1996:GSM

[Pra96]  Rudra Pratap.  Getting

Started with MATLAB: a Quick Introduction for Scientists and Engineers. Saunders golden sunburst series. Saunders College Publishing, Ft. Worth, TX, USA, 1996. ISBN 0-03-017884-3. iv + 171 pp. LCCN Q183.9.P73 1996.

Pratap:1999:GSM

[Pra99]  Rudra Pratap.  Getting

started with MATLAB 5: a quick introduction for scientists and engineers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512947-4. x + 230 pp. LCCN Q183.9 .P73 1999.

Pratap:2002:GSM

[Pra02]  Rudra Pratap.  Getting

started with MATLAB: a quick introduction for scientists and engineers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-515014-7 (paperback). 245 (est.) pp. LCCN Q183.9 .P74 2002. US$40.11.

Pratap:2006:GSM

[Pra06]  Rudra Pratap.  Getting

started with MATLAB 7/a quick introduction for scientists and engineers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517937-4 (paperback). x + 244 pp. LCCN Q183.9 .P733

Prisman:2000:PDS

[Pri00] Eliezer Z. Prisman. Pricing derivative securities: an interactive, dynamic environment with Maple V and Matlab. Academic Press, New York, NY, USA, 2000. ISBN 0-12-564915-0, 0-12-564916-9 (CD-ROM). xxviii + 754 pp. LCCN HG6024.A3 P74 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els033/99068356.
html;   http://www.loc.
gov/catdir/toc/els033/
99068356.html.

Painter:1996:MSTa

[PS96a] E. Painter and A. Spanias. A MATLAB software tool for the introduction of speech coding fundamentals in a DSP course. In IEEE [IEE96f], pages 1133– 1136. ISBN 0-7803-3193-1, 0-7803-3192-3, 0-7803-3194-X, 0-7803-3195-8. ISSN 0749-8411. LCCN TK 7882 S65 I16 1996. Six volumes.

Painter:1996:MSTb

[PS96b] E. Painter and A. Spanias. A MATLAB software tool for the introduction of speech coding fundamentals in a DSP course. In Iskander et al. [I^{+}96], pages 603–

608. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN T62 .F76 1996. Three volumes. IEEE catalog number: 96CH35946.

Proakis:1998:CCS

[PS98] John G. Proakis and Masoud Salehi. Contemporary communication systems using MATLAB. PWS BookWare companion series. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1998. ISBN 0-534-93804-3 (paperback/disk). xiii + 427 pp. LCCN TK5105.P744 1998.

Proakis:2000:CCS

[PS00] John G. Proakis and Masoud Salehi. Contemporary communication systems using MATLAB. BookWare companion series. Brooks/ Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37173-6. xiii + 428 pp. LCCN TK5105 .P744 2000.

Persson:2004:SMG

[PS04] Per-Olof Persson and Gilbert Strang. A simple mesh generator in MATLAB. SIAM Review, 46(2):329– 345, June 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42912.

Proakis:2004:CCS

[PSB04] John G. Proakis, Masoud Salehi, and Gerhard Bauch. Contemporary communication systems using MATLAB and Simulink. Thomson–Brooks/Cole, Belmont, CA, USA, second edition, 2004. ISBN 0-534-40617-3 (paperback). viii + 487 pp. LCCN TK5105 .P744 2004.

Pukrushpan:2004:CFC

[PSP04] Jay T. Pukrushpan, Anna G. Stefanopoulou, and Huei Peng. Control of fuel cell power systems: principles, modeling, analysis, and feedback design. Advances in industrial control. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 1-85233-816-4. xvii + 161 pp. LCCN TK2931 .P85 2004.

Prusa:2016:DWT

[PSR16] Zdeněk Průsa, Peter L. Søndergaard, and Pavel Rajmic. Discrete wavelet transforms in the large timefrequency analysis toolbox for MATLAB/GNU Octave. ACM Transactions on Mathematical Software, 42(4):32:1–32:23, July 2016. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:

//dl.acm.org/citation.
cfm?id=2839298.

Pagel:1997:CTM

[PSTO97] J. Pagel, Y. Sun, D. Tilbury, and L. Oms. Control tutorials for Matlab on the World Wide web. In Anonymous [Ano97h], pages 3150– 3154. ISBN 0-7803-3833-2, 0-7803-3832-4, 0-7803-3834-0, 0-7803-3835-9. ISSN 0743-1619. LCCN TJ212 .A57 1997. Five volumes.

Polyanin:2008:PDE

[PSZ08] Andrei D. Polyanin, William E.

Schiesser, and Alexei I. Zhurov. Partial differential equation. Scholarpedia, 3(10):4605, 2008. ISSN 1941-6016. URL http:

//www.scholarpedia.org/
article/Partial_differential_
equations.

Petrila:2007:LMP

[PT07] Titus Petrila and Damian Trif. LiScNLE—a Matlab package for some nonlinear partial differential evolution equations. Buletinul Academiei de Ştiinţe a Republicii Moldova. Matematica. Izvestiya Akademii Nauk Respubliki Moldova. Matematika, 3:23–34, 2007. ISSN 1024-7696.

Piazzola:2024:ASG

[PT24] Chiara Piazzola and Lorenzo Tamellini. Algorithm 1040: The Sparse Grids Matlab Kit — a Matlab implementation of sparse grids for high-dimensional function approximation and uncertainty quantification. ACM Transactions on Mathematical Software, 50(1):7:1– 7:22, March 2024. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3630023.

Popovic:1999:MED

[PV99] Dobrivoje Popović and Ljubo Vlacic, editors. Mechatronics in engineering design and product development. Marcel Dekker, New York, NY, USA, 1999. ISBN 0-8247-0226-3. xi + 615 pp. LCCN TJ163.12 .M435 1999.

Palamides:2010:SSL

[PV10] Alex Palamides and Anastasia Veloni. Signals and systems laboratory with MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-3055-X. ???? pp. LCCN TK5102.9 .P35 2010.

Potter:2002:MF

[PWH02] Merle C. Potter, D. C. Wiggert, and Midhat Hondzo. Mechanics of fluids. Brooks/Cole– Thomson Learning, Pacific Grove, CA, USA, third edition, 2002. ISBN 0-534-37996-6. xvi + 863 pp. LCCN TA357 .P725 2002.

Patterson:2013:EOM

[PWR13] Michael A. Patterson, Matthew Weinstein, and Anil V. Rao. An efficient overloaded method for computing derivatives of mathematical functions in MAT-LAB. ACM Transactions on Mathematical Software, 39 (3):17:1–17:36, April 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Passino:1998:FC

[PY98] Kevin M. Passino and Stephen Yurkovich. Fuzzy control. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-18074-X. xviii + 475 pp. LCCN TJ213 .P317 1998.

Papp:2022:AMP

[PY22] Dávid Papp and Sercan Yıldız. Alfonso: Matlab package for nonsymmetric conic optimization. INFORMS Journal on Computing, 34(1):11– 19, Winter 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/
doi/fpi/10.1287/ijoc.2021.
1058.

Qiu:1996:DMT

[QCPG96] L. Qiu, T. Chen, A. Potvin, and J. Gu. Development of MATLAB tools and case study for multirate control design. In IEEE [IEE96e], pages 1278–1283. ISBN 0-7803-3591-0, 0-7803-3590-2, 0-7803-3592-9, 0-7803-3593-

7. ISSN 0191-2216. LCCN TJ 217 I11c 1996. Four volumes.

Quinn:2001:ETF

[QH01] Barry G. Quinn and E. J. (Edward James) Hannan. The estimation and tracking of frequency. Cambridge series in statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80446-9. xi + 266 pp. LCCN TK5102.9 .Q85 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam021/00051944.
html;   http://www.loc.
gov/catdir/toc/cam027/
00051944.html.

Quinn:1998:BGB

[QMS98] M. Quinn, A. Malishevsky, and N. Seelam. Bridging the gap between MATLAB and ScaLAPACK. In IEEE [IEE98], page ?? ISBN 0-8186-8579-4, 0-8186-8581-6 (microfiche). LCCN QA76.9.D5 I157 1998. IEEE Order Plan Catalog Number 98TB100244. IEEE Computer Society Press order number PR08579.

Quarteroni:2003:SCM

[QS03] Alfio Quarteroni and Fausto Saleri. Scientific computing with MATLAB, volume 2 of Texts in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-44363-0 (hard cover). ISSN 1611-0994. ix + 257 pp. LCCN Q183.9 .Q8313 2003.

Quarteroni:2006:SCM

[QS06] Alfio Quarteroni and Fausto Saleri. Scientific computing with MATLAB and Octave, volume 2 of Texts in computational science and engineering. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2006. ISBN 3-540-32612-X. xvi + 318 pp. LCCN Q183.9 .Q8313

2006. URL http://www.
loc.gov/catdir/enhancements/
fy0814/2006928277-d.html;
http://www.loc.gov/catdir/
toc/fy0708/2006928277.
html.

Quarteroni:2014:SCM

[QSG14] Alfio Quarteroni, Fausto Saleri, and Paola Gervasio. Scientific computing with Matlab and Octave, volume 2 of Texts in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-642-45366-X (hard cover). xviii + 450 (est.) pp. LCCN

????  URL http://
link.springer.com/book/
10.1007/978-3-642-45367-
0.

Quatieri:2002:DTS

[Qua02] T. F. (Thomas F.) Quatieri. Discrete-time speech signal processing: principles and practice. Prentice-Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-242942-X. xix + 781 pp. LCCN TK7882.S65 Q38 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy031/2001021821.html.

Quarteroni:2010:SCM

[Qua10] Alfio M. Quarteroni. Scientific computing with Matlab and Octave. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-12429-1. ???? pp. LCCN ????

Qureshi:2001:AAO

[Qur01]  Shehrzad Qureshi.  Al-

gorithm alley: Optimization and fixed-point iteration. Dr. Dobb’s Journal of Software Tools, 26(12):115– 117, December 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/ftp/2001/2001_
12/aa1201.txt;   http:
//www.ddj.com/ftp/2001/
2001_12/aa1201.zip.

Qureshi:2005:EIP

[Qur05] Shehrzad Qureshi. Embedded image processing on the TMS320C6000 DSP: examples in code composer studio and MATLAB. Springer Science+Business Media, New York, NY, USA, 2005. ISBN 0-387-25280-0 , 0-387-25281-9 (e-book). xvi + 433 pp. LCCN TK5102.9 .Q94 2005.

Rodgerson:1992:TSC

[RA92] J. L. Rodgerson and J. H.

Anderson. Teaching systems and control using MAT-LAB. International Journal of Electrical Engineering Education, 29(3):235– 242, July 1992. CODEN IJEEAF. ISSN 0020-7209.

Ranganathan:1995:DHV

[RA95] R. Ranganathan and A. Aia.

Development of heavy vehicle dynamic stability analysis model using MAT-LAB/SIMULINK. In SAE [SAE95a], pages 67–74. ISBN 1-56091-716-4. LCCN TL255.N393 1995.

Rabe:2020:SMG

[Rab20] Martin Rabe. Spectram:

a MATLAB(R) and GNU Octave toolbox for transition model guided deconvolution of dynamic spectroscopic data. Journal of Open Research Software, 8 (1):13–??, June 09, 2020. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.323/.

Radke:1996:MII

[Rad96] Richard Joseph Radke. A Matlab implementation of the Implicitly Restarted Arnoldi Method for solving large-scale eigenvalue problems. Master’s thesis, Rice University, Houston, TX, USA, 1996. 94 pp.

URL http://hdl.handle.
net/1911/14054;   http:
//scholarship.rice.edu/
bitstream/handle/1911/
14054/1379511.PDF.

Rahnamai:1993:FAT

[Rah93] K. Rahnamai. Frequency analysis tools for a first control course using MATLAB. In Kinoglu [Kin93], pages 533–538. ISBN 0-7918-1169-7. LCCN TA345.A86 1993.

Ramirez:1994:PCI

[Ram94] W. Fred Ramirez. Process control and identification. Academic Press, New York, NY, USA, 1994. ISBN 0-12-577240-8. xiii + 424 pp. LCCN TS156.8 .R36 1994. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;

http://www.loc.gov/catdir/
description/els032/93029779.
html;   http://www.loc.
gov/catdir/toc/els032/
93029779.html.

Ramirez:1997:CMP

[Ram97] W. Fred Ramirez. Computational methods for process simulation. Butterworths, London, UK, second edition, 1997. ISBN 0-7506-3541-X. ix + 461 pp. LCCN TP155.7 .R345 1997. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els033/98149714.
html;   http://www.loc.
gov/catdir/toc/els032/
98149714.html.

Randhawa:1996:CIE

[Ran96] S. U. Randhawa, editor.

Computers and industrial engineering: International conference; 19th — March 1996, Miami, FL, volume 31(1) of Computers and Industrial Engineering. Pergamon Press, New York, NY, USA, 1996. ISBN ???? ISSN 0360-8352 (print), 1879-0550 (electronic). LCCN ????

Rao:2002:ANM

[Rao02]  Singiresu S. Rao.  Ap-

plied Numerical Methods for Engineers and Scientists. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-089480-X. xx + 1080 pp. LCCN TA345 .R36 2002.

Rao:2004:MV

[Rao04] S. S. Rao. Mechanical vibrations. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2004. ISBN 0-13-048987-5. xxvi + 1078 pp. LCCN TA355 .R37 2004.

Rao:2011:CPE

[Rao11] T. Subba Rao. Classification, parameter estimation and state estimation — an engineering approach using MATLAB. Journal of Time Series Analysis, 32(2): 194, March 2011. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Ravn:1994:ATU

[Rav94a] O. Ravn. Advanced topics using MATLAB and SIMULINK. In Brdys and Malinowski [BM94], pages 249–262. ISBN 981-02-1391-3. LCCN TJ213 .C568 1994.

Ravn:1994:IEM

[Rav94b]  O. Ravn.  An interac-

tive environment in MAT-LAB. In Brdys and Malinowski [BM94], pages 39– 44. ISBN 981-02-1391-3. LCCN TJ213 .C568 1994.

Reuter:2016:FMG

[RAW^{+}16] Balthasar Reuter, Vadym Aizinger, Manuel Wieland, Florian Frank, and Peter Knabner. FESTUNG: a MATLAB/GNU Octave toolbox for the discontinuous Galerkin method, Part II: Advection operator and slope limiting. Computers and Mathematics with Applications, 72(7):1896–1925, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122116304606.

Rao:1998:WTI

[RB98] Raghuveer M. Rao and Ajit S. Bopardikar. Wavelet transforms: introduction to theory and applications. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-63463-5. xiii + 310 pp. LCCN QA403.3 .R36 1998.

Roy:2004:ACF

[RB04] S. Roy and P. Banerjee.

An algorithm for converting floating-point computations to fixed-point in MATLAB based FPGA design. In Proceedings. 41st Design Automation Conference, June 7–11, 2004, pages 484–

487. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004.

Roy:2005:ATQ

[RB05] Sanghamitra Roy and Prith Banerjee. An algorithm for trading off quantization error with hardware resources for MATLABbased FPGA design. IEEE Transactions on Computers, 54(7):886–896, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=1432671.

Reis:2020:CMC

[RBC20] Luís Reis, João Bispo, and João M. P. Cardoso. Compilation of MATLAB computations to CPU/GPU via C/OpenCL generation. Concurrency and Computation: Practice and Experience, 32(22):e5854:1– e5854:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rao:2010:AGM

[RBD^{+}10] Anil V. Rao, David A. Benson, Christopher Darby, Michael A. Patterson, Camila Francolin, Ilyssa Sanders, and Geoffrey T. Huntington. Algorithm 902: GPOPS, a MATLAB software [sic] for solving multiplephase optimal control problems using the Gauss pseudospectral method. ACM Transactions on Mathematical Software, 37(2):22:1– 22:39, April 2010. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See corrigendum [RBD^{+}11].

Rao:2011:CAG

[RBD^{+}11] Anil V. Rao, David A.

Benson, Christopher Darby, Michael A. Patterson, Camila Francolin, Ilyssa Sanders, and Geoffrey T. Huntington. Corrigendum: Algorithm 902: GPOPS, a MATLAB software for solving multiple-phase optimal control problems using the Gauss pseudospectral method. ACM Transactions on Mathematical Software, 38(1):9:1–9:2, November 2011. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [RBD^{+}10].

Rios-Bolivar:1996:SDA

[RBZ96] M. Rios-Bolivar and A. S. I.

Zinober. Symbolic dynamical adaptive backstepping control design via MAT-LAB. In IEEE [IEE96b], pages 216–221. ISBN 0-7803-3718-2. LCCN TJ216 .I28 1996.

Redfern:1998:MH

[RC98] Darren Redfern and Colin Campbell. The MATLAB 5 handbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-94200-9 (softcover). xi + 488 pp. LCCN QA297 .R395 1998.

Reimer:2013:MBF

[RC13] Ashton S. Reimer and Alexei F. Cheviakov. A Matlab-based finite-difference solver for the Poisson problem with mixed Dirichlet– Neumann boundary conditions. Computer Physics Communications, 184(3): 783–798, March 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465512003293.

See corrigendum [RC16].

Reimer:2016:CMB

[RC16] Ashton S. Reimer and Alexei F. Cheviakov. Corrigendum to: “A Matlabbased finite difference solver for the Poisson problem with mixed Dirichlet– Neumann boundary conditions” [comput. phys. comm. 184(3) (2013) 783– 798]. Computer Physics Communications, 209(??): 200–201, December 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465516302703. See

[RC13].

Ramasawmy:2020:EMT

[RCT20] Danny R. Ramasawmy, Ben T. Cox, and Bradley E.

Treeby. ElasticMatrix:

a MATLAB toolbox for anisotropic elastic wave propagation in layered media. SoftwareX, 11 (??):Article 100397, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711019303048.

Rosjat:2022:DDS

[RD22] Nils Rosjat and Silvia

Daun.  DST (Dynamic
Synchronization Toolbox):

a MATLAB implementation of the dynamic phaselocking pipeline from stimulus transformation into motor action: Dynamic graph analysis reveals a posteriorto- anterior shift in brain network communication of older subjects. Journal of Open Research Software, 10 (1):8–??, August 01, 2022. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.394/.

Rama:2014:SPI

[RDP14] S. Rama, C. Surendra Dilip, and Rajesh Narayana Perumal. A software program to investigate the nucleation kinetics of solution grown crystals using MATLAB platform. Computer Physics Communications, 185(2):661–669, February 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S001046551300338X.

Recktenwald:2000:NMM

[Rec00] Gerald W. Recktenwald. Numerical methods with MATLAB: implementations and applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-030802-1, 0-201-30860-

6. xxviii + 786 pp. LCCN

TA345 .R43 2000.

Reichelt:1993:IFO

[Rei93]  Eric R. Reichelt.  Im-

plementing Fortran ODE solver LSODE using Mat-Lab. Thesis (M.S.), San Diego State University, San Diego, CA, USA, 1993. vii + 65 pp.

Renard:2017:PHO

[Ren17] Philippe Renard. Hytool: an open source MATLAB toolbox for the interpretation of hydraulic tests using analytical solutions. Journal of Open Source Software, 2 (19):441:1–441:3, November 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00441.

Resnick:2019:BRIc

[Res19] Andrew Resnick. Book review: Introductory quantum mechanics with MATLAB for atoms, molecules, clusters and nanocrystals, by J.

R. Chelikowsky, Weinheim, Wiley-VCH Verlag GmbH & Co. KGaA, 2018, 224 pp., £70.00 (paperback), ISBN: 978-3-527-40926-6. Contemporary Physics, 60(4): 322, 2019. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riefstahl:2021:FMP

[RG21] Florian Riefstahl and Felix Gross. FastGAPP — a MATLAB based program supports Earth scientists interpreting geochemical, petrological and sedimentological data. Journal of Open Research Software, 9(1):11–??, May 31, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.333/.

Reis:2022:SCU

[RGM22] Eduardo Reis, Catarina Gralha, and Miguel P. Monteiro. Surveying communities of users of MAT-LAB and clone languages. Journal of Computer Languages, 73(??):1–12, December 2022. CODEN ???? ISSN 2665-9182 (print), 2590-1184 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S2590118422000673.

Rich:1992:ADM

[RH92] Lawrence C. Rich and David R. Hill. Automatic differentiation in MATLAB. Applied Numerical Mathematics: Transactions of IMACS, 9(1):33–43, January 1992. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rao:2000:FFA

[RH00] A. Ramachandra (Adiseshappa Ramachandra) Rao and Khaled H. Hamed. Flood frequency analysis. New directions in civil engineering. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-412-55280-9, 0-8493-0083-5 (cover). xvii + 350 pp. LCCN GB1399.2 .R36 2000.

Ramaswamy:1996:CMP

[RHB96] Shankar Ramaswamy, Eugene W. Hodges, IV, and Prithviraj Banerjee. Compiling MATLAB programs to ScaLAPACK: Exploiting task and data parallelism. In IEEE [IEE96c], pages 613–619. ISBN 0-8186-7255-2, 0-8186-7257-9. LCCN QA 76.58 I56 1996.

Ramsay:2009:FDA

[RHG09] J. O. (James O.) Ramsay, Giles Hooker, and Spencer Graves. Functional data analysis with R and MAT-LAB. Use R! Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-98184-5 (paperback), 0-387-98185-3 (e-book). xi + 207 pp. LCCN QA278 .R365 2009.

Reuter:2021:FOU

[RHR^{+}21] Balthasar Reuter, Hennes Hajduk, Andreas Rupp, Florian Frank, Vadym Aizinger, and Peter Knabner. FESTUNG 1.0: Overview, usage, and example applications of the MATLAB/GNU Octave toolbox for discontinuous Galerkin methods. Computers and Mathematics with Applications, 81(??):3–41, January 1, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122120303254.

Ribbens:2000:SEN

[Rib00] Jack Ribbens. Simultaneous engineering for new product development: manufacturing applications. Wiley, New York, NY, USA, 2000. ISBN 0-471-25265-4 (cloth). xix + 332 pp. LCCN TS170 .R53 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley042/99046112.
html;   http://www.loc.

                                        gov/catdir/description/
                                        wiley033/99046112.html;
                                         http://www.loc.gov/catdir/
                                        toc/onix04/99046112.html.

Ridgel:1995:MIN

[Rid95] Jeffrey A. Ridgel. MAT-LAB implementation of N-dimensional probability density functions (NPDF) for remotely sensed data. Thesis (M.S.), California State University, Chico, Chico, CA, USA, 1995. x + 49 pp.

Rivoire:2001:MSS

[Riv01] Maurice Rivoire. Matlab, Simulink, Stateflow: avec des exercises d’automatique résolus. (French) [Matlab, Simulink, Stateflow: with automatically-solved exercises]. Collection informatique. Editions Technip, Paris, France, 2001. ISBN 2-7108-0789-0. 306 (est.) pp.

Rizzoni:2000:PAE

[Riz00] Giorgio Rizzoni. Principles and applications of electrical engineering. McGraw-Hill, New York, NY, USA, third edition, 2000. ISBN 0-256-26116-4, 0-07-116979-2, 0-07-117727-2 (International edition), 0-07-365467-1 (CD-ROM). xvi + 976 pp. LCCN TK146 .R473 2000.

Rizzardi:2017:ATS

[Riz17] Mariarosaria Rizzardi. Algorithm 981: Talbot Suite DE: Application of modified Talbot’s method to solve differential problems. ACM Transactions on Mathematical Software, 44(2):18:1– 18:23, September 2017. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:

//dl.acm.org/citation.
cfm?id=3089248.

Reyment:1993:AFA

[RJ93] Richard A. Reyment and K. G. Joreskog. Applied Factor Analysis in the Natural Sciences. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-41242-0. xii + 371 pp. LCCN QA278.5.R49 1993.

Ratnanather:2014:ATI

[RKZ^{+}14] J. Tilak Ratnanather, Jung H.

Kim, Sirong Zhang, Anthony M. J. Davis, and Stephen K. Lucas. Algo-

rithm 935: IIPBF, a MATLAB

toolbox for infinite integral of products of two Bessel functions. ACM Transactions on Mathematical Software, 40(2):14:1–14:12, February 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rosenwasser:2000:CCS

[RL00] Efim Rosenwasser and Bernhard P. Lampe. Computer controlled systems: analysis and design with processorientated models. Communications and control engineering, 0178-5354. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-85233-307-

3. xviii + 486 pp. LCCN TJ213 .R653 2000.

Ricci:2003:MSF

[RLH03] Francesco Ricci and Hoang Le-Huy. Modeling and simulation of FPGA-based variable-speed drives using Simulink. Mathematics and Computers in Simulation, 63(3–5):183–195, November 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0378475403000661.

Rupp:2011:SBF

[RLV11] Benjamin Rupp, Howard Lovatt, and Andrea Vezzini. Simulink-based floatingpoint DSP control platform. In Proceedings of the 2011 14th European Conference on Power Electronics and Applications (EPE 2011), pages 1–7. pub-IEEE, pub-IEEE:adr, 2011. URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=
&arnumber=6020132.

Rohrs:1993:LCS

[RMS93] Charles Rohrs, James L. Melsa, and Donald G. Schultz. Linear Control Systems. McGraw-Hill series in electrical and computer engineering. Control theory. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-041525-0. xvi + 555 pp. LCCN QA402.3 .R64 1993.

Robe:1994:SME

[Rob94] Edward D. Robe. SIMULINK modules that emulate digital controllers realized with fixed-point or floating-point arithmetic. Thesis (M.S.), Ohio University, Athens, OH, USA, June 1994. v + 130 pp.

Robbin:1995:MAU

[Rob95] Joel W. Robbin. Matrix algebra using MINImal MATlab. A. K. Peters, Ltd., Wellesley, MA, USA, 1995. ISBN 1-56881-024-5. xvi + 544 pp. LCCN QA188 .R62 1995. US$49.95.

Roberts:1996:UMT

[Rob96] P. D. Roberts. Using MAT-LAB and two-dimensional system theory to compute the stability of an interactive optimal control algorithm. In Anonymous [Ano96m], page 13. CODEN DCILDN. ISBN ???? ISSN 0963-3308. LCCN ????

Robinett:2002:FRD

[Rob02] Rush D. Robinett, editor.

Flexible robot dynamics and controls, volume 19 of IFSR international series on systems science and engineering. Kluwer Academic… Plenum Publishers, New York, NY, USA, 2002. ISBN 0-306-46724-0. xix + 339 pp. LCCN TS191.8 .F59 2002; TS 191.8 .F59 2002X ENGI.

Roberts:2004:SSA

[Rob04] Michael J. Roberts. Signals and systems: analysis using transform methods and MATLAB. McGraw-Hill, New York, NY, USA, 2004. ISBN 0-07-249942-7. xvii + 1054 pp. LCCN TK5102.9 .R63 2004 UCB.

Rodgerson:1992:TST

[Rod92] J. L. Rodgerson. Teaching systems theory using MAT-LAB. In Van Wyk [Van92b], pages 562–565. ISBN 0-7803-0835-2, 0-7803-0836-0, 0-7803-0837-9. LCCN TK5 .A47 1992.

Rogers:2000:PSO

[Rog00] Graham Rogers. Power system oscillations, volume SECS 539. Power electronics and power systems of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 2000. ISBN 0-7923-7712-5. xi + 328 pp. LCCN TK1010 .R64 2000.

Rogers:2003:AMI

[Rog03] Robert M. Rogers. Applied mathematics in integrated navigation systems. AIAA education series. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024–2518, second edition, 2003. ISBN 1-56347-656-8 (Hardcover). xvi + 330 pp. LCCN TL695 .R64 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip048/2003019359.
html.

Romanowicz:1997:MIT

[Rom97] R. Romanowicz. A Matlab implementation of Top-Model. Hydrological processes, 11(9):1115–??, ???? 1997. ISSN 0885-6087.

Roque:2013:SNA

[Roq13] C. M. C. Roque. Symbolic and numerical analysis of plates in bending using Matlab. Journal of Symbolic Computation, 61–62 (??):3–11, ???? 2013. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0747717113001247.

Rossiter:2003:MBP

[Ros03] J. A. Rossiter. Model-based predictive control: a practical approach. Control series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2003. ISBN 0-8493-1291-4. 318 (est.) pp. LCCN TJ217.6 .R67 2003.

Roughgarden:1998:PET

[Rou98]  Jonathan  Roughgarden.

Primer of ecological theory. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-442062-4. xiv + 456 pp. LCCN QH541.15.M3 R68 1998.

Rovati:1990:VIG

[Rov90] G. Enrico Rovati. A versatile implementation of the Gauss–Newton minimization algorithm using MAT-LAB for Macintosh microcomputers. Computer Methods and Programs in Biomedicine, 32(2):161–167, June 1990. CODEN CMP-BEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Rovenski:2010:MCS

[Rov10a] Vladimir Rovenski. Modeling of curves and surfaces with Matlab. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-71277-1. ???? pp. LCCN ????

Rovenskii:2010:MCS

[Rov10b] Vladimir Y. Rovenskii. Modeling of curves and surfaces with MATLAB. Springer undergraduate texts in mathematics and technology. Springer Science+Business Media, New York, NY, USA, 2010. ISBN 0-387-71278-X, 0-387-71277-

1. xv + 452 pp. LCCN QA448.D38 R68 2010.

Rowe:2003:MBS

[Row03] Daniel B. Rowe. Multivariate Bayesian statistics: models for source separation and signal unmixing. Chapman and Hall/CRC, Boca Raton, FL, USA, 2003. ISBN 1-58488-318-9. xx + 329 pp. LCCN QA279.5 .R68 2003.

Roblin:2001:HSH

[RR01] Patrick Roblin and Hans Rohdin. High-speed heterostructure devices. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-78152-3. xxxiv + 688 pp. LCCN TK7871.85 .R56 2001.

Roblin:2002:HSH

[RR02] Patrick Roblin and Hans Rohdin. High-speed heterostructure devices: from device concepts to circuit modeling. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-78152-3. xxxiv + 688 pp. LCCN TK7871.85 .R56 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam021/00066719.
html;   http://www.loc.

                                        gov/catdir/toc/cam027/
                                        00066719.html.

Ramsay:2002:AFD

[RS02] J. O. (James O.) Ramsay and B. W. Silverman. Applied functional data analysis: methods and case studies. Springer series in statistics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95414-7 (paperback). x + 190 pp. LCCN QA278 .R35 2002.

Ratcliffe:2008:GMT

[RS08] Sarah J. Ratcliffe and Justine Shults. GEE-QBOX: a MATLAB toolbox for generalized estimating equations and quasileast squares. Journal of Statistical Software, 25(14): 1–14, May 2008. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v25/i14.

Ryabtsev:2009:TVS

[RS09] Michael Ryabtsev and Ofer Strichman. Translation validation: From Simulink to C. In Bouajjani and Maler [BM09], pages 696– 701. ISBN 3-642-02657-5 (paperback), 3-642-02658-3. LCCN QA76.76.V47 .C38 2009.

Rogel-Salazar:2015:OHP

[RS15a]  J. Rogel-Salazar.  One

Hundred Physics Visualisations Using Matlab, (With DVD-ROM), by Dan Green. Scope: Review. Level: Undergraduate. Contemporary Physics, 56(1):100, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2015:EMO

[RS15b] Jesus Rogel-Salazar. Essential MATLAB and Octave. Taylor and Francis, CRC Press, Boca Raton, FL, USA, 2015. ISBN 1-4822-3463-7 (paperback), 1-4822-3466-1 (ePub e-book), 1-4822-3464-5 (PDF e-book), 1-4822-3465-3 (VitalBook ebook). xxv + 261 pp. LCCN QA76.95 .R59 2015.

Rojas:2008:ALM

[RSS08] Marielba Rojas, Sandra A.

Santos, and Danny C. Sorensen. Algorithm 873: LSTRS: MATLAB software for large-scale trust-region subproblems and regularization. ACM Transactions on Mathematical Software, 34 (2):11:1–11:28, March 2008. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ryeznik:2015:RMS

[RSW15] Yevgen Ryeznik, Oleksandr Sverdlov, and Weng Kee

Wong.  RARtool: A

MATLAB software package for designing responseadaptive randomized clinical trials with time-to-event outcomes. Journal of Statistical Software, 66 (1):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v066i01;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v066i01/v66i01.pdf.

Reznik:1996:EMA

[RT96] L. Reznik and B. Tavron. Employment of MATLAB in the analysis of radionuclide deposition and leakage. In Anonymous [Ano96g], page 5.2. ISBN ???? LCCN ????

Rule:2002:OAC

[Rul02] Keith Rule. Oscilloscope Analysis and Connectivity Made Easy. Tektronix Inc., M/S 50-PTLD, Beaverton, OR 97077-0001, USA, October 28, 2002. URL http:

                                        //www.tek.com/site/mn/
                                        mnfinder_detail/1,1096,
                                        ,00.html?id=6241&pn=071104601.

Rump:1995:IIL

[Rum95] Siegfried M. Rump. INT-LAB: Interval laboratory, a MATLAB toolbox for interval arithmetic. World-Wide Web document., 1995. URL http://

                                        www.ti3.tu-harburg.de/
                                        rump/intlab;   https:/
                                        /www.tuhh.de/ti3/rump/
                                        intlab/.

Rump:2001:FVA

[Rum01] Siegfried M. Rump. Fast verification algorithms in Matlab. In Symbolic algebraic methods and verification methods (Dagstuhl, 1999), pages 209–226. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Rump:2017:IPK

[Rum17] Siegfried M. Rump. IEEE754 precision-k base-β arithmetic inherited by precisionm base-β arithmetic for k < m. ACM Transactions on Mathematical Software, 43(3):20:1–20:15, January 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=2785965.

Rump:2023:IPP

[Rum23] Siegfried M. Rump. IEEE-754 precision-p base-β arithmetic implemented in binary. ACM Transactions on Mathematical Software, 49(4):32:1–32:??, December 2023. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596218;   https://www.
tuhh.de/ti3/paper/rump/
Ru23b.pdf.

Rust:1993:GMC

[Rus93] J. Rust. GAUSS and MAT-LAB: a comparison. Journal of Applied Econometrics, 8(3):307–324, July 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Russo:2008:IPR

[Rus08] Ralph P. Russo. Intuitive probability and random processes using MAT-LAB. The American Statistician, 62(2):181–182, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rahman:2013:FMA

[RV13] Talal Rahman and Jan Valdman. Fast MAT-LAB assembly of FEM matrices in 2D and 3D: Nodal elements. Applied Mathematics and Computation, 219(13):7151–7158, March 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300311010836.

Renes:1992:MGC

[RVV^{+}92] W. A. Renes, M. Vanbegin, P. Van Dooren, J. W. J. Beckers, and J. The MAT-LAB gateway compiler. A tool for automatic linking of FORTRAN routines to MATLAB. In Barker [Bar92], pages 95– 100. ISBN 0-08-041269-6. LCCN TJ213 .C57 1992.

Redivo-Zaglia:2012:SMT

[RZR12] Michela Redivo-Zaglia and Giuseppe Rodriguez. smt: a Matlab toolbox for structured matrices. Numerical Algorithms, 59(4):639– 659, April 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1017-1398&volume=
59&issue=4&spage=639.

Singh:2001:SDT

[SA01] Krishna Kumari Singh and Gayatri Agnihotri. System design through MAT-LAB, Control Toolbox and SIMULINK. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 1-85233-337-5 (paperback). xxiv + 487 pp. LCCN QA76.9.S88 S5673 2001.

Saadat:1993:CAC

[Saa93]  Hadi Saadat.  Compu-

tational Aids in Control Systems Using MATLAB. McGraw-Hill series in electrical and computer engineering. Control theory. Mc-Graw-Hill, New York, NY, USA, 1993. ISBN 0-07-911358-3. xii + 141 pp. LCCN TA165.A1 S2 1993.

SAE:1995:NDA

[SAE95a] SAE, editor. New developments in axle, steering, suspension, and classic technology: International truck and bus meeting, November 1995, Winston-Salem, NC, USA, volume SP-1128 of Papers — Society of Automotive Engineers New York. Society of Automotive Engineers, Warrendale, PA, USA, 1995. ISBN 1-56091-716-4. LCCN TL255.N393 1995.

SAE:1995:VCA

[SAE95b] SAE, editor. Vehicle computer applications: vehicle systems and driving simulation: Session, International congress — February 27 – March 2, 1995, Detroit, MI, number SAE/SP-95/1080 in Papers — Society of Automotive Engineers New York 1995; SP-1080. Society of Automotive Engineers, Warrendale, PA, USA, 1995. ISBN 1-56091-630-3. LCCN TL240.V4237 1995.

Sharifzadeh:2015:MBM

[SAKG15] Mohsen Sharifzadeh, Hosein Afarideh, Hosein Khalafi, and Reza Gholipour. A Matlab-based Monte Carlo algorithm for transport of gamma-rays in matter. Monte Carlo Methods and Applications, 21(1):77–??, March 2015. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://

www.degruyter.com/view/
j/mcma.2015.21.issue-1/
mcma-2014-0011/mcma-2014-
0011.xml.

Sarra:2010:AMP

[Sar10] Scott A. Sarra. Algorithm 899: The Matlab postprocessing toolkit. ACM Transactions on Mathematical Software, 37(1):10:1–10:15, January 2010. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sarra:2017:MRB

[Sar17] Scott A. Sarra. The Matlab Radial Basis Function Toolbox. Journal of Open Research Software, 5 (1):8–??, March 27, 2017. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.131/.

Sas:1996:NVE

[Sas96]  P. Sas, editor.  Noise

and vibration engineering: International conference — September 1996, Leuven, Belgium, PRO-CEEDINGS OF THE IN-TERNATIONAL SEMI-NAR ON MODAL ANAL-YSIS CONF 21//V2. Departement Werktuigkunde, Katholieke Universiteit Leuven, Leuven, Belgium, 1996.

ISBN 90-73802-60-1. LCCN

????

Sayed:2003:FAF

[Say03] Ali H. Sayed. Fundamentals of adaptive filtering. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-471-46126-1. xxxvii + 1125 pp. LCCN TK7872.F5 S29 2003. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley039/2003273186.
                                        ;   http://www.loc.gov/
                                        catdir/description/wiley039/
                                        2003273186.html; UCLA.

Scharf:1990:FCE

[SB90] Louis L. Scharf and Richard T. Behrens. A First Course in Electrical and Computer Engineering with MATLAB Programs and Experiments. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-53472-X. xv + 269 pp. LCCN TK168 .S34 1990.

Stonick:1996:LSS

[SB96] Virginia L. Stonick and Kevin Bradley. Labs for signals and systems using MATLAB. The PWS BookWare companion series. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1996. ISBN 0-534-93808-6. xvi + 143 pp. LCCN TK5102.9 .S76 1996.

Strang:1997:LAG

[SB97] Gilbert Strang and K. (Kai)

Borre. Linear algebra, geodesy, and GPS. Wellesley-Cambridge Press, Wellesley, MA, USA, 1997. ISBN 0-9614088-6-3 (hardcover). xvi + 624 pp. LCCN TA347.L5 S87 1997.

Solomon:2010:FDI

[SB10] Chris Solomon and Toby Breckon. Fundamentals of digital image processing: a practical approach with examples in Matlab. Wiley, New York, NY, USA, 2010. ISBN 0-470-84472-8. ???? pp. LCCN TA1637 .S65154 2010. URL http:

//catalogimages.wiley.
com/images/db/jimages/
9780470844724.jpg.

Switkes:2004:MTR

[SBC04] Jennifer Switkes, Robert L.

Borrelli, and Courtney S. Coleman. Matlab technology resource manual [for] Differential equations: a modeling perspective. Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-48387-7 (paperback). vi + 53 pp. US$16.95.

Santana:2010:MMP

[SBL^{+}10] Roberto Santana, Concha Bielza, Pedro Larrañaga, Jose A. Lozano, Carlos Echegoyen, Alexander Mendiburu, Rubén Armañanzas, and Siddartha Shakya. Mateda-2.0: a MATLAB package for the implementation and analysis of estimation of distribution algorithms. Journal of Statistical Software, 35(7):??, July 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v35/i07.

Schmeisser:2017:PER

[SBL^{+}17] Andre Schmeißer, Daniel Burkhart, Dominik Linn, Johannes Schnebele, Manuel Ettmüller, Simone Gramsch, and Walter Arne.

Ensight4Matlab:  read,

process, and write files in EnSight(R) gold format from C++ or MAT-LAB(R). Journal of Open Source Software, 2 (20):217:1, December 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00217.

Smyl:2019:OMP

[SBL19] Danny Smyl, Sven Bossuyt, and Dong Liu. Open-QSEI: a MATLAB package for quasi static elasticity imaging. SoftwareX, 9(??):73–76, January/June 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018300256.

Sorensen:1997:DSP

[SC97] Henrik V. Sorensen and Jianping Chen. A Digital Signal Processing Laboratory Using the TMS320C30. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-741828-0. viii + 215 pp. LCCN TK5102.9.S67 1997. Includes a set of functions for the MATLAB Signal Processing Toolbox.

Sotack:1999:HSC

[SCB99] Robert A. Sotack, Rajiv S.

Chowdhry, and Carey S. Buttrill. High speed civil transport aircraft simulation: reference-H cycle 1: MATLAB implementation. Washington, DC, USA, 1999. ???? pp. Shipping list no.: 2000-0596-M.

Scokaert:1995:SPA

[SCC95] P. O. M. Scokaert, C. M.

Chow, and D. W. Clarke. Simulink predictive adaptive control environment (space) user’s guide. Technical Report 2057, University of Oxford. Dept. of Engineering Science, Oxford, UK, ???? 1995. ???? pp.

Schussler:1994:DSA

[Sch94] H. W. Schüssler. Digital Signalverarbeitung I: Analyse diskreter Signale und System [English: Digital Signal Processing, Volume I: Analysis of Discrete Signals and Systems]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1994. ISBN 3-540-57428-X. LCCN ????

Scheinerman:1996:IDS

[Sch96] Edward R. Scheinerman. Invitation to Dynamical Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-185000-8. xvii + 373 pp. LCCN QA614.8.S34 1996.

Schaufelberger:1997:SEM

[Sch97] W. Schaufelberger. The student edition of MAT-LAB. Automatica: the journal of IFAC, the International Federation of Automatic Control, 33(5):1005– ??, ???? 1997. CODEN AT-CAA9. ISSN 0005-1098.

Schmidt:1998:IAF

[Sch98] Louis V. Schmidt. Introduction to aircraft flight dynamics. AIAA education series. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024–2518, 1998. ISBN 1-56347-226-0. xii + 397 pp. LCCN TL570 .S33 1998.

Schleher:1999:EWI

[Sch99] D. Curtis Schleher. Electronic warfare in the information age. Artech House radar library. Artech House Inc., Boston, MA, USA, 1999. ISBN 0-89006-526-8. xiv + 605 pp. LCCN UG485 .S3497 1999.

Schreppers:2004:TMM

[Sch04] Walter Schreppers. Towards a multiprecision MAT-LAB environment. World-Wide Web slides presentation., November 17,

2004. URL http://www.
wogsymposium.ugent.be/
Schreppers.pdf.  WOG

symposium, Universiteit Gent, Belgium.

Schott:2005:IMG

[Sch05]  Dieter Schott.  Inge-

nieurmathematik mit MAT-LAB. (German) [Engineering Mathematics with MATLAB]. Carl Hanser, München, Germany, 2005. ISBN ???? 519 pp. LCCN ????

Schilling:2011:FDS

[Sch11] Robert J. Schilling. Fundamentals of digital signal processing using Matlab. Cengage Learning, Mason, OH, USA, second edition, 2011. ISBN 0-8400-6909-X. ???? pp. LCCN

???? URL http://www.
loc.gov/catdir/enhancements/
fy1105/2010938460-b.html;
 http://www.loc.gov/catdir/
enhancements/fy1105/2010938460-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1105/2010938460-t.html.

Schiesser:2012:PDE

[Sch12] W. E. Schiesser. Partial differential equation analysis in biomedical engineering: case studies with MAT-LAB. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-02280-0 (hardback). ???? pp. LCCN ????

Schiesser:2014:ICM

[Sch14] William E. Schiesser. An introductory comparison of Matlab and R: Applications to ordinary and partial differential equations. Report, Lehigh University, Bethlehem, PA 18105, USA, January 15, 2014. 48

                                        pp.  URL http://www.
                                        lehigh.edu/~wes1/R_PDE/
                                        R-Matlab_compare.pdf.

SotoPrieto:1995:ALC

[SCL95] Manuel Jesus Soto Prieto, Vicente Cordoba, and Jose Luis, aut. Algebra lineal: con Matlab y Maple. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-526815-X. viii + 301 pp. LCCN ????

Sjostrand:2018:SMT

[SCL^{+}18] Karl Sjöstrand, Line Harder Clemmensen, Rasmus Larsen, Gudmundur Einarsson, and Bjarne Ersbøll. SpaSM: A MATLAB toolbox for sparse statistical modeling. Journal of Statistical Software, 84(??):??, ????

2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v084i10;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v084i10/
v84i10.pdf.

Smith:1991:OFK

[SD91] P. R. Smith and C. Dailly.

Optimal filtering and Kalman filtering using MATLAB. IEE Conference Publication, 2(332):734–737, 1991. CODEN IECPB4. ISSN 0537-9987 (??invalid ISSN checksum??).

Stearns:1996:SPA

[SD96] Samuel D. Stearns and Ruth A. David. Signal processing algorithms in MAT-LAB. Prentice-Hall signal processing series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-045154-1. xvi + 372 pp. LCCN TK5102.9 .S72 1996.

Sigmon:2002:MP

[SD02] Kermit Sigmon and Timothy A. Davis. MAT-LAB primer. Chapman and Hall/CRC, Boca Raton, FL, USA, sixth edition, 2002. ISBN 1-58488-294-8 (paperback). xii + 163 pp. LCCN QA297 .S4787 2002. US$16.95, UK£11.99.

Sanchez:2004:JMB

[SDPM04] J. Sanchez, S. Dormido, R. Pastor, and F. Morilla. A Java/ Matlab-based environment for remote control system laboratories: Illustrated with an inverted pendulum. IEEE Transactions on Education, 47(3): 321–329, 2004. CODEN IEEDAB. ISSN 0018-9359.

Seborg:2004:PDC

[SEM04a] Dale E. Seborg, Thomas F. Edgar, and Duncan A. Mellichamp. Process dynamics and control. Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-00077-9. xv + 713 pp. LCCN TP155.75 .S43 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        toc/wiley032/2003049712.
                                        html.

Semmlow:2004:BBI

[Sem04b] John L. Semmlow. Biosignal and biomedical image processing: MATLAB-based applications, volume 22 of Signal processing series. Marcel Dekker, New York, NY, USA, 2004. ISBN 0-8247-4803-4. xviii + 423 pp. LCCN R857.O6 S463 2004.

Semmlow:2005:CSS

[Sem05] John L. Semmlow. Circuits, systems, and signals for bioengineers: a MATLAB-based introduction. Academic Press series in biomedical engineering. Elsevier Academic Press, Amsterdam, The Netherlands, 2005. ISBN 0-12-088493-3. xiv + 446 pp. LCCN R856 .S453 2005.

Sethares:2005:TTS

[Set05] William A. Sethares. Tuning, timbre, spectrum, scale. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2005. ISBN 1-85233-797-4. xviii + 426 pp. LCCN QC225.7 .S48 2005.

Seibold:2014:SSO

[SF14] Benjamin Seibold and Martin Frank. StaRMAP — a second order staggered Grid method for spherical harmonics moment equations of radiative transfer. ACM Transactions on Mathematical Software, 41(1):4:1–4:28, October 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Saddique:1991:FQC

[SFK91] S. M. Saddique, M. Farsi, and K. Z. Karam. Finetuned QFT control in a Matlab environment. IEE Conference Publication, 2 (332):1257–1260, 1991. CO-DEN IECPB4. ISSN 0537-9987 (??invalid ISSN checksum??).

Senini:1994:AMT

[SFPO94] S. Senini, F. Flinders, S. Pang, and W. Oghanna. Asynchronous motor traction drive simulation using Simulink dynamic systems analysis package. In Murthy et al. [M^{+}94], pages 445– 454. ISBN 1-85312-266-1 (Southampton: set), 1-56252-190-X (Boston: set), 1-85312-354-4 (Southampton: v. 1), 1-56252-282-5 (Boston: v. 1), 1-85312-359-5 (Southampton: v. 2), 1-56252-283-3 (Boston: v. 2). LCCN TF507.I55 1994.

Schiesser:2009:CPD

[SG09] W. E. Schiesser and Graham W. Griffiths. A compendium of partial differential equation models: method of lines analysis with Matlab. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-51986-1 (hardback). xiii + 474 + 2 pp. LCCN QA377 .S3538 2009. URL http://

                                        www.loc.gov/catdir/toc/
                                        fy0905/2008045816.html.

Sorensen:1995:KSU

[SGA95] O. B. Sorensen, A. Goucem, and D. P. Atherton. Kit software for use with MAT-LAB. In Anonymous [Ano95b], pages 8/1–8/7. CODEN DCILDN. ISBN ???? ISSN 0963-3308. LCCN ????

Serrano-Gotarredona:1998:ART

[SGLBA98] Teresa Serrano-Gotarredona, Bernabé Linares-Barranco, and Andreas G. Andreou. Adaptive resonance theory microchips, volume SECS 456 of Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1998. ISBN 0-7923-8231-5. xxi + 234 pp. LCCN TK7874.75 .S47 1998.

Shampine:2003:SOM

[SGT03] Lawrence F. Shampine, I. Gladwell, and S. Thompson. Solving ODEs with MATLAB. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-82404-4, 0-521-53094-6 (paperback). viii + 263 pp. LCCN QA371.5.D37 S43 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam031/2003041958.
html;   http://www.loc.
gov/catdir/samples/cam034/
2003041958.html;   http:
//www.loc.gov/catdir/toc/
cam031/2003041958.html.

Shahian:1993:CSD

[SH93] Bahram Shahian and Michael Hassul. Control System Design Using Matlab. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-174061-X. xx + 503 pp. LCCN TJ213 .S424 1993.

Schilling:2000:ANM

[SH00] Robert J. (Robert Joseph) Schilling and Sandra L. Harris. Applied numerical methods for engineers using MATLAB and C. Brooks/ Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37014-4. xx + 715 pp. LCCN TA345 .S34 2000.

Schilling:2005:FDS

[SH05] Robert J. (Robert Joseph) Schilling and Sandra L. Harris. Fundamentals of digital signal processing using MATLAB. Thomson, Southbank, Victoria, Australia, 2005. ISBN 0-534-39150-8. xiii + 786 pp.

Shonkwiler:2009:MB

[SH09] Ronald W. Shonkwiler and J. V. Herod. Mathematical biology: an introduction with Maple and Matlab. Undergraduate texts in mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 0-387-70984-3 (paperback), 0-387-70983-5 (hardcover). ISSN 0172-6056. xiii + 551 pp. LCCN QH323.5 .S56 2009.

Stearns:2011:DSP

[SH11] Samuel D. Stearns and Donald R. Hush. Digital signal processing with examples in MATLAB. Electrical engineering and applied signal processing series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2011. ISBN 1-4398-3782-1. ???? pp. LCCN TK5102.9 .S719 2011.

Schmidt:2018:WMP

[SH18] Burkhard Schmidt and Carsten Hartmann. WavePacket: A Matlab package for numerical quantum dynamics. II: Open quantum systems, optimal control, and model reduction. Computer Physics Communications, 228(??):229– 244, July 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465518300614.

Shampine:2002:SM

[Sha02] L. F. Shampine. Solving 0 = F (t, y(t), y′(t)) in Matlab. J. Numer. Math., 10(4): 291–310, 2002. ISSN 1570-2820.

Shames:2003:MF

[Sha03] Irving Herman Shames. Mechanics of fluids. McGraw-Hill series in mechanical engineering. McGraw-Hill, New York, NY, USA, fourth edition, 2003. ISBN 0-07-247210-3, 0-07-119889-X. various pp. LCCN TA357 .S44 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/mh024/2002071906.
html;   http://www.loc.
gov/catdir/toc/mh023/2002071906.
html.

Shacham:2004:BRB

[Sha04] Mordechai Shacham. Book review: Scientific Computing with MATLAB: A. Quarteroni, F. Saleri, Springer-Verlag, Berlin, 2003, ISBN 3-540-44363-0. Computer Physics Communications, 161(3):183–185, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S001046550400147X.

Shampine:2005:SHP

[Sha05] L. F. Shampine. Solving hyperbolic PDEs in MAT-LAB. ANACM. Applied Numerical Analysis and Computational Mathematics, 2 (3):346–358, 2005. ISSN 1611-8170.

Shampine:2007:AND

[Sha07] L. F. Shampine. Accurate numerical derivatives in MATLAB. ACM Transactions on Mathematical Software, 33(4):26:1–26:17, August 2007. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Shampine:2008:M

[Sha08a] L. F. Shampine. Matlab. Applied Mathematics and Computation, 202(1): 266–274, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shampine:2008:MPQ

[Sha08b] L. F. Shampine. MATLAB program for quadrature in 2D. Applied Mathematics and Computation, 202(1): 266–274, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shampine:2008:VAQ

[Sha08c]  L. F. Shampine.  Vec-

torized adaptive quadrature in Matlab. Journal of Computational and Applied Mathematics, 211(2): 131–140, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Shampine:2009:VSO

[Sha09] L. F. Shampine. Vectorized solution of ODEs in Matlab. Scalable Computing: Practice and Experience, 10(4):337–345, December 2009. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.
org/vols/vol10/no4/SCPE_
10_4_01.pdf;   http://
www.scpe.org/vols/vol10/
no4/SCPE_10_4_01.zip.

Shampine:2011:IOF

[Sha11] L. F. Shampine. Integrating oscillatory functions in MATLAB. International Journal of Computer Mathematics, 88(??):2348–2358, ???? 2011. CODEN IJC-MAT. ISSN 0020-7160. See also Part II [Sha12].

Shampine:2012:IOF

[Sha12] L. F. Shampine. Integrating Oscillatory Functions in Matlab, II. Electronic Transactions on Numerical Analysis (ETNA), 39: 403–413, 2012. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://

                                        etna.mcs.kent.edu/vol.
                                        39.2012/pp403-413.dir/
                                        pp403-413.pdf;   http:/
                                        /etna.mcs.kent.edu/volumes/
                                        2011-2020/vol39/abstract.
                                        php?vol=39&pages=403-413.

See also Part I [Sha11].

Sherrick:2004:CSS

[She04] John D. Sherrick. Concepts in systems and signals. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2004. ISBN 0-13-178271-1. ix + 447 pp. LCCN TK5102.9 .S5325 2004.

Sheng:2008:MCM

[She08a] Yanyan Sheng. Markov chain Monte Carlo estimation of normal Ogive IRT models in MATLAB. Journal of Statistical Software, 25(8):1–15, April 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v25/i08.

Sheng:2008:MPM

[She08b] Yanyan Sheng. A MAT-LAB package for Markov chain Monte Carlo with a multi-unidimensional IRT model. Journal of Statistical Software, 28(10): ??, November 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v28/i10.

Sheng:2010:BEM

[She10] Yanyan Sheng. Bayesian estimation of MIRT models with general and specific latent traits in MATLAB. Journal of Statistical Software, 34(3):??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v34/
i03.

Shinners:1992:MCS

[Shi92] Stanley M. Shinners. Modern Control System Theory and Design. Wiley, New York, NY, USA, 1992. ISBN 0-471-55008-6. xvi + 855 pp. LCCN TJ213.S15 1992.

Shinners:1998:AMC

[Shi98a] Stanley M. Shinners. Advanced modern control system theory and design. Wiley, New York, NY, USA, 1998. ISBN 0-471-31857-4. xvi + 607 pp. LCCN TJ213 .S45543 1998. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley042/98030680.
html;   http://www.loc.
gov/catdir/description/
wiley033/98030680.html;
 http://www.loc.gov/catdir/
toc/onix05/98030680.html.

Shinners:1998:MCS

[Shi98b] Stanley M. Shinners. Modern control system theory and design. Wiley, New York, NY, USA, second edition, 1998. ISBN 0-471-24906-8 (cloth). xviii + 720 pp. LCCN TJ213 .S455443 1998. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley041/97034538.
html;   http://www.loc.
gov/catdir/description/
wiley031/97034538.html;
 http://www.loc.gov/catdir/
toc/onix01/97034538.html.

Shiavi:1999:IAS

[Shi99] Richard Shiavi. Introduction to applied statistical signal analysis. Academic Press series in biomedical engineering. Academic Press, New York, NY, USA, second edition, 1999. ISBN 0-12-640010-5. xx + 390 pp. LCCN TK5102.5 .S474 1999. URL ftp://

uiarchive.cso.uiuc.edu/

pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els033/98022148.
html;   http://www.loc.
gov/catdir/toc/els032/
98022148.html.

Shu:2001:BRS

[Shu01]  Chi-Wang Shu.  Book

review: Spectral methods in Matlab. Mathematics of Computation, 70(235):1337–1338, July 2001. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/mcom/2001-
70-235/S0025-5718-01-01388-
6/bookrev-S0025-5718-01-
01388-6.html;   http:/
/www.ams.org/mcom/2001-
70-235/S0025-5718-01-01388-
6/S0025-5718-01-01388-
6.dvi;   http://www.ams.
org/mcom/2001-70-235/S0025-
5718-01-01388-6/S0025-
5718-01-01388-6.pdf;
http://www.ams.org/mcom/
2001-70-235/S0025-5718-
01-01388-6/S0025-5718-
01-01388-6.ps;   http:/
/www.ams.org/mcom/2001-
70-235/S0025-5718-01-01388-
6/S0025-5718-01-01388-
6.tex.

Shubbar:2017:UMI

[Shu17]  Safa Shubbar.  Ultra-

sound medical imaging systems using telemedicine and blockchain for remote monitoring of responses to neoadjuvant chemotherapy in women’s breast cancer: Concept and implementation. M.S., Kent State University, Kent, OH, USA,

2017.  133 pp.  URL

                                        http://search.proquest.
                                        com/pqdtglobal/docview/
                                        2059846207.

Saito:1996:SCS

[SHX96] Osami Saito, Hideki Hino, and Li Xu. Symbolic CAD system for algebraic approach of MATLAB — symbolic control toolbox. In Anonymous [Ano96v], pages 1338–1343. ISBN 0-85296-668-7, 0-85296-666-0. ISSN 0537-9989. LCCN TJ212.2 .U32 1996; TK5.I4 no.427. Two volumes.

Schetzen:2000:DSL

[SI00] Martin Schetzen and Vinay K. Ingle. Discrete systems laboratory using MATLAB. BookWare companion series. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37463-8. xv + 119 pp. LCCN TK7887.6 .S346 2000.

Sigmon:1992:MP

[Sig92] Kermit Sigmon. MATLAB primer. Technical report, Department of Mathematics, University of Florida, Gainesville, 1992. ???? pp. Available in TEX source code form by email from

                                        sigmon@math.ufl.edu.

Sigmon:1994:MP

[Sig94]  Kermit Sigmon.  MAT-

LAB Primer. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, fourth edition, 1994. ISBN 0-8493-9440-6. 101 pp. LCCN QA297.S4787 1994. US$12.95.

Silvester:1996:SEE

[Sil96] P. P. Silvester, editor. Software for electrical engineering analysis and design: Third International Conference on Software for Electrical Engineering Analysis and Design, Electrosoft ’96. Computational Mechanics Publications, Southampton, 1996. ISBN 1-85312-395-1, 1-85312-385-1 (??invalid ISBN checksum??). LCCN TK5.I59 1996.

Simonsen:1999:EFM

[Sim99] Leif Simonsen. Elektriske filter med Matlab, PSPICE og LabVIEW. (Norwegian) [Electrical Fields with Matlab, PSPICE and Lab-VIEW]. Høgskolen i Narvik, Narvik, Norway, 1999. ISBN ???? 223 pp. LCCN ????

Sincovec:1993:PSS

[Sin93] Richard F. Sincovec, editor. Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, Norfolk, VA, March, 1993. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3. LCCN QA76.58.S55 1993. Two volumes.

Strum:1994:CLS

[SK94] Robert D. Strum and Donald E. Kirk. Contemporary linear systems using MATLAB. Tom Robbins’ BookWare companion series. PWS Pub. Co., Boston, 1994. ISBN 0-534-93273-8 (recycled paper). xxx + 673 pp. LCCN QA402 .S858 1994.

Strum:1995:CLS

[SK95] Robert D. Strum and Donald E. Kirk. Contemporary linear systems using MATLAB 4.0. The PWS BookWare companion series. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1995. ISBN 0-534-94710-7. xxx + 673 pp. LCCN QA402 .S858 1996. US$45.95.

Strum:2000:CLS

[SK00] Robert D. Strum and Donald E. Kirk. Contemporary linear systems using MAT-LAB. BookWare companion series. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37172-8. xxx + 673 pp. LCCN QA402 .S858 2000.

Schmidt:2019:WMP

[SKA19] Burkhard Schmidt, Rupert Klein, and Leonardo Cancissu Araujo. WavePacket: a Matlab package for numerical quantum dynamics. III. Quantum-classical simulations and surface hopping trajectories. Journal of Computational Chemistry, 40(30):2677–2688, November 15, 2019. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shampine:2005:UAS

[SKF05] L. F. Shampine, Robert Ketzscher, and Shaun A. Forth. Using AD to solve BVPs in Matlab. ACM Transactions on Mathematical Software, 31(1):79–94, March 2005. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Shearer:1997:DMC

[SKG97] J. Lowen Shearer, Bohdan T. Kulakowski, and John F. (John Francis) Gardner. Dynamic Modeling and Control of Engineering Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1997. ISBN 0-13-356403-7 (hardcover). xx + 375 pp. LCCN TA342 .S54

Stevens:2003:ACS

[SL03] Brian L. Stevens and Frank L. Lewis. Aircraft control and simulation. Wiley, New York, NY, USA, second edition, 2003. ISBN 0-471-37145-9 (cloth). xvi + 664 pp. LCCN TL678 .S74 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley046/2003043250.
html;   http://www.loc.
gov/catdir/description/
wiley039/2003043250.html;
 http://www.loc.gov/catdir/
toc/wiley032/2003043250.
html.

Schmidt:2017:WMP

[SL17] Burkhard Schmidt and Ulf Lorenz. WavePacket: a Matlab package for numerical quantum dynamics. I: Closed quantum systems and discrete variable representations. Computer Physics Communications, 213(??):223–234, April 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465516303861.

Soutas-Little:1999:EMD

[SLI99a] Robert W. Soutas-Little and D. J. Inman. Engineering mechanics: Dynamics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999.

ISBN 0-13-278409-2. xvii + 702 pp. LCCN TA352 .S615 1999.

Soutas-Little:1999:EMS

[SLI99b] Robert W. Soutas-Little and D. J. Inman. Engineering mechanics: Statics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-769001-0. xii + 564 pp. LCCN TA351 .S58 1999.

Szatkowski:2023:LGB

[SLM23] Mateusz Szatkowski, Przemysław Litwin, and Jan Masajada. LBSA: a GUIbased Matlab software for the advanced laser beam shaping with spatial light modulators. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711023000389.

Szymkat:1994:TDT

[SM94] M. Szymkat and J. M. Maciejowski. Time Delay Toolbox for Matlab. In Mattsson et al. [MGC94], pages 505– 512 (or 505–511??). ISBN 0-7803-1800-5, 0-7803-1801-3. LCCN TJ 212.2 I3256 1994.

Singer:1995:ADA

[SM95] G. Singer and Y. Meashio.

Analysis of a double actuator electrohydraulic system for structural testing. In Anonymous [Ano95b], pages 4–?? ISBN ???? LCCN ????

Stoica:1997:ISA

[SM97] Petre Stoica and Randolph L. Moses. Introduction to spectral analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-258419-0. xviii + 319 pp. LCCN QA320 .S86 1997.

Stoica:2005:SAS

[SM05] Petre Stoica and Randolph L. Moses. Spectral analysis of signals. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-113956-8. xxii + 452 pp. LCCN QA320 .S864 2005.

Stubben:2007:EAD

[SM07] Chris Stubben and Brook Milligan. Estimating and analyzing demographic models using the popbio package in R. Journal of Statistical Software, 22(11): 1–23, September 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v22/i11.

Sharma:2009:MLP

[SM09] Gaurav Sharma and Jos Martin. MATLAB^{(}R): a language for parallel computing. International Journal of Parallel Programming, 37(1):3– 36, February 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=1&spage=3.

Silva:2014:OST

[SM14] Ikaro Silva and George Moody. An open-source toolbox for analysing and processing PhysioNet databases in MATLAB and Octave. Journal of Open Research Software, 2(1):e27– ??, September 24, 2014. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.bi/.

Sizemore:2015:MD

[SM15] Jim Sizemore and John Mueller. MATLAB for dummies. For dummies. For Dummies, Hoboken, NJ, USA, 2015. ISBN 1-118-82010-X (paperback), 1-118-82003-7 (ebook), 1-118-82434-2 (ebook). xiv + 406 pp. LCCN QA297 .S52 2015.

Shigley:2004:MED

[SMB04] Joseph Edward Shigley, Charles R. Mischke, and Richard G. (Richard Gordon) Budynas. Mechanical engineering design. McGraw-Hill series in mechanical engineering. Mc-Graw-Hill, New York, NY, USA, seventh edition, 2004. ISBN 0-07-252036-1. xxv + 1030 pp. LCCN TJ230 .S5 2004.

Smith:1997:MPB

[Smi97] Rick L. Smith. The MAT-LAB project book for linear algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-521337-1. 243 (est.) pp. LCCN QA185.C65 S65 1997.

Saifuddin:1995:AMS

[SMS95] A. B. Saifuddin, J. M. Maciejowski, and M. Szymkat. Analysis of multivariable Smith predictors using MATLAB containers. In Anonymous [Ano95b], pages 1/1–1/4. ISBN ???? LCCN ????

Saifuddin:1996:CCC

[SMS96] A. B. Saifuddin, J. M. Maciejowski, and M. Szymkat. Computational chains for CACSD using MAT-LAB containers. In IEEE [IEE96d], pages 392– 397. ISBN 0-7803-3032-3, 0-7803-3033-1. LCCN TJ212.2.I32495 1996. IEEE catalog number 96TH8136.

Strang:1996:WFB

[SN96] Gilbert Strang and Truong Nguyen. Wavelets and Filter Banks. Wellesley-Cambridge Press, Wellesley, MA, USA, 1996. ISBN 0-9614088-7-1. xxi + 490 pp. LCCN TK7872.F5 S79 1996. See [Stu96].

Schneider:2001:AAM

[SN01] Tapio Schneider and Arnold Neumaier. Algorithm 808: ARfit—a Matlab package for the estimation of parameters and eigenmodes of multivariate autoregressive models. ACM Transactions on Mathematical Software, 27(1):58–65, March 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Senini:1993:NIM

[SO93] S. Senini and W. Oghanna.

A nonlinear inductor model for implementation with “SIMULINK” power electronics laboratory. In Anonymous [Ano93d], pages 28–34. ISBN ???? LCCN ????

Some:2007:UNM

[Som07] Longin Some. Using the new Matlab-based method of lines (MATMOL) toolbox to solve three test problems by moving-grid techniques. Far East Journal of Applied Mathematics, 26(1): 121–138, 2007. ISSN 0972-0960.

Soumekh:1994:FAI

[Sou94] Mehrdad Soumekh. Fourier Array Imaging. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-063769-6. xv + 559 pp. LCCN TK5102.9 .S67 1994.

Soumekh:1999:SAR

[Sou99] Mehrdad Soumekh. Synthetic aperture radar signal processing with MAT-LAB algorithms. Wiley, New York, NY, USA, 1999. ISBN 0-471-29706-2 (cloth). xxvii + 616 pp. LCCN TK6592.S95 S68 1999.

Schoukens:1991:ILS

[SP91] J. (Johan) Schoukens and R. (Rik) Pintelon. Identification of Linear Systems: a Practical Guideline to Accurate Modeling. Pergamon Press, New York, NY, USA, 1991. ISBN 0-08-040734-X. xxi + 332 pp. LCCN QA402.3 .S345 1991; QA 402.3 .S345 1991. US$110.00.

Skogestad:1996:MFC

[SP96] Sigurd Skogestad and Ian Postlethwaite. Multivariable feedback control: analysis and design. Wiley, New York, NY, USA, 1996. ISBN 0-471-94330-4 (paperback), 0-471-94277-4. xi + 559 pp. LCCN TJ216 .S47 1996. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley033/96203203.
html;   http://www.loc.
gov/catdir/toc/onix04/
96203203.html.

Stripinis:2022:DND

[SP22] Linas Stripinis and Remigijus Paulavicius. DI-RECTGO: a new DIRECTtype MATLAB toolbox for derivative-free global optimization. ACM Transactions on Mathematical Software, 48(4):41:1–41:??, December 2022. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3559755.

Spall:2003:ISS

[Spa03] James C. Spall. Introduction to stochastic search and optimization: estimation, simulation, and control. Wiley-Interscience series in discrete mathematics and optimization. Wiley, New York, NY, USA, 2003. ISBN 0-471-33052-3 (cloth). xx + 595 pp. LCCN QA274 .S63 2003. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley034/2002038049.
html;   http://www.loc.
gov/catdir/toc/wiley031/
2002038049.html;   http:
//www3.interscience.wiley.
com/cgi-bin/bookhome/109868994.

Speciale:1995:CGV

[Spe95] Ross A. Speciale. Computation and graphic visualization of plane-wave Kspace spectra and far-field patterns with MATLAB 4.0. In IEEE [IEE95a], pages 1090–?? CODEN IAPSBG. ISBN 0-7803-2720-9. ISSN 0272-4693. LCCN TK 7871.6 A2 1995. Four volumes. IEEE catalog number: 95CH35814.

Speciale:1996:CGV

[Spe96] Ross A. Speciale. Computation and graphic visualization of plane-wave k-space spectra and farfield patterns with MAT-LAB 4.0. In Anonymous [Ano96t], pages 1150–1157. CODEN CPCEFK. ISBN ???? LCCN ????

Spooner:2002:SAC

[Spo02] Jeffrey Spooner, editor. Stable adaptive control and estimation for nonlinear systems: neural and fuzzy approximator techniques. Adaptive and learning systems for signal processing, communications, and control. Wiley, New York, NY, USA, 2002. ISBN 0-471-41546-4. xvii + 545 pp. LCCN QA402.35 .S73 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley043/2002512362.
html;   http://www.loc.
gov/catdir/description/
wiley035/2002512362.html;
 http://www.loc.gov/catdir/
toc/wiley031/2002512362.
html.

Smits:1990:LAC

[SR90] J. G. M. M. Smits and J. J. M. Rijpkema. A linear algebra course with PC-Matlab: Some experiences. Technical report, Preprint, Department of Mathematics and Computer Science, Eindhoven University of Technology, The Netherlands, ????, August 1990. ???? pp.

Scott:1994:MCC

[SR94] P. H. Scott and J. P. Russell. Modelling of a combined cycle conversion using Simulink. In Anonymous [Ano94d], pages 368– 373. ISBN ???? ISSN 0537-9989. LCCN ???? Two volumes.

Shampine:1995:SMO

[SR95] Lawrence F. Shampine and Mark W. Reichelt. The Matlab ODE suite. Manuscript, 1995.

Shampine:1997:MOS

[SR97] Lawrence F. Shampine and Mark W. Reichelt. The MATLAB ODE suite. SIAM Journal on Scientific Computing, 18(1):1– 22, ???? 1997. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). Dedicated to C. William Gear on the occasion of his 60th birthday.

Shampine:2002:MOS

[SR02] Lawrence F. Shampine and Mark W. Reichelt. The Matlab ODE suite. World-Wide Web document, February 27, 2002.

URL http://www.cs.berkeley.
edu/~wkahan/Math128/ODEsuite.

pdf. Lecture notes for Math 128.

Shults:2009:AML

[SR09] Justine Shults and Sarah J. Ratcliffe. Analysis of multilevel correlated data in the framework of generalized estimating equations via xtmultcorr procedures in Stata and qls functions in Matlab. Statistics and its Interface, 2(2):187–196, 2009. ISSN 1938-7989.

Shampine:1999:SID

[SRK99] Lawrence F. Shampine, Mark W. Reichelt, and Jacek A. Kierzenka. Solving index-1 DAEs in MATLAB and Simulink. SIAM Review, 41(3):538–552, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/33425.

Sankowski:1996:SNM

[SS96a] D. Sankowski and P. Szaruga. Simulation of neural models of electroheat systems using neural network toolbox for Matlab. In Javor et al. [JLM96], pages 681– 685. ISBN 1-56555-097-8. LCCN ????

Sciavicco:1996:ISM

[SS96b] Lorenzo Sciavicco and Bruno Siciliano. Instructor’s Solutions Manual for Modeling and Control of Robot Manipulators. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-057218-6. ???? pp. LCCN ???? This manual is for the book [SS96c].

Sciavicco:1996:MCR

[SS96c] Lorenzo Sciavicco and Bruno Siciliano. Modeling and Control of Robot Manipulators. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-057217-8. xvii + 358 pp. LCCN TJ210.3.M63 1996. An instructor’s manual is available [SS96b].

Sciavicco:2000:MCR

[SS00] L. (Lorenzo) Sciavicco and Bruno Siciliano. Modelling and control of robot manipulators. Advanced textbooks in control and signal processing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-85233-221-2. xxiii + 377 pp. LCCN TJ211.35. S43 2000.

Sumathi:2010:CIP

[SS10] S. Sumathi and P. Surekha.

Computational intelligence paradigms: theory and applications using MATLAB.

CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-0903-8 (electronic book), 1-4398-0902-X (hardback). xxiii + 829 pp. LCCN Q342 .S94 2010.

SouzadeCursi:2015:UQS

[SS15] Eduardo Souza de Cursi and Rubens Sampaio. Uncertainty Quantification and Stochastic Modeling with Matlab. ISTE Press Ltd, London, UK, 2015. ISBN 0-08-100471-0 (e-book), 1-78548-005-7. LCCN QA274.2.

URL http://alltitles.
ebrary.com/Doc?id=11040161;
 http://lib.myilibrary.
com?id=762946;   http:/
/public.eblib.com/choice/
PublicFullRecord.aspx?
p=2007484;   http://www.
sciencedirect.com/science/
book/9781785480058.

Saavedra:2018:OMI

[SS18] Pablo Saavedra and Clemens Simmer. An Octave/MATLAB(R) interface for rapid processing of SMOS L1C full polarization brightness temperature. Journal of Open Research Software, 6(1): 2–??, January 08, 2018. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.165/.

Santina:1994:DCS

[SSH94a] Mohammed S. Santina, Allen R. Stubberud, and Gene H. Hostetter. Digital Control System Design. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 1994. ISBN 0-03-076012-

7. xvii + 797 pp. LCCN TJ223.M53S26 1994. An instructor’s manual is available [SSH94b].

Santina:1994:ISM

[SSH94b] Mohammed S. Santina, Allen R. Stubberud, and Gene H. Hostetter. Instructor’s Solutions Manual for Digital Control System Design. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 1994. ISBN 0-03-000692-9. ???? pp. LCCN ???? This manual is for the book [SSH94a].

Saff:2003:FCA

[SSTD03] Edward B. Saff, Arthur David Snider, Lloyd N. (Lloyd Nicholas) Trefethen, and Tobin A. (Tobin Allen) Driscoll. Fundamentals of complex analysis: with applications to engineering and science. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2003. ISBN 0-13-907874-6, 0-13-017968-X (International ed.). xi + 511 pp. LCCN QA300 .S18 2003.

Snorek:1995:MSE

[SSV95] M. Snorek, M. Sujansky, and A. Verbraeck, editors. Modelling and simulation 1995: European simulation multiconference — June 1995, Prague, Czech Republic, Modelling and Simulation — Society for Computer Simulation International Conference. Society for Computer Simulation, San Diego, CA, USA, 1995. ISBN 1-56555-080-3. LCCN ????

Saucez:2009:MIM

[SSW09] P. Saucez, L. Some, and A. Vande Wouwer. Matlab implementation of a moving grid method based on the equidistribution principle. Applied Mathematics and Computation, 215(5):1821–1829, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006821.

Sigmon:1998:MP

[ST98] Kermit Sigmon and The MathWorks Staff. MAT-LAB Primer. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, fifth edition, 1998. ISBN 0-8493-1305-8. 130 pp. LCCN QA297.S4787 1998. US$14.95.

Shampine:2001:SDM

[ST01] L. F. Shampine and S. Thompson. Solving DDEs in MAT-LAB. Applied Numerical Mathematics: Transactions of IMACS, 37(4):441– 458, 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sadkane:2012:ASM

[ST12] Miloud Sadkane and Ahmed Touhami. Algorithm 918:

specdicho: a MATLAB

program for the spectral dichotomy of regular matrix pencils. ACM Transactions on Mathematical Software, 38(3):21:1–21:13, April 2012. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sanfelici:2024:FMV

[ST24] Simona Sanfelici and Giacomo Toscano. The Fourier–Malliavin Volatility (FMVol) MATLAB(R) library. Mathematics and Computers in Simulation, 226:338–353, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0378475424002568.

Stanley:2003:NAA

[Sta03a] William D. Stanley. Network analysis with applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2003. ISBN 0-13-060246-9. x + 662 pp. LCCN TK454.2 .S72 2003.

Stanley:2003:TCA

[Sta03b] William D. Stanley. Transform circuit analysis for engineering and technology. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fifth edition, 2003. ISBN 0-13-060259-0. 487 pp. LCCN TK454 .S7 2003. Stanoyevitch:2005:IMN

[Sta05a] Alexander Stanoyevitch. Introduction to MATLAB^{(}R) with numerical preliminaries. Wiley-Interscience, New York, NY, USA, 2005. ISBN 0-471-69737-0. x + 331 pp.

Stanoyevitch:2005:INO

[Sta05b] Alexander Stanoyevitch. Introduction to numerical ordinary and partial differential equations using MATLAB^{(}R). Pure and Applied Mathematics (New York). Wiley-Interscience, New York, NY, USA, 2005. ISBN 0-471-69738-9. xiv + 813 pp.

Shawe-Taylor:2004:KMP

[STC04] John Shawe-Taylor and Nello Cristianini. Kernel methods for pattern analysis. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-81397-2 (hardcover). xiv + 462 pp. LCCN Q325.5 .S475 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam051/2003069590.
html;   http://www.loc.
gov/catdir/toc/cam051/
2003069590.html.

Stenger:1996:BRS

[Ste96] Frank Stenger. Book reviews: Solving Problems in Scientific Computing Using MAPLE and MATLAB, by Walter Gander and Jírí Hrebícek. Mathematics of Computation, 65(214):880– 882, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:

//www.ams.org/jourcgi/
jour-pbprocess?fn=110&
arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Stefani:2002:DFC

[Ste02] Raymond T. Stefani, editor. Design of feedback control systems. The Oxford series in electrical and computer engineering. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, fourth edition, 2002. ISBN 0-19-514249-7. xvi + 848 pp. LCCN TJ216 .D417 2001.

Stearns:2003:DSP

[Ste03] Samuel D. Stearns. Digital signal processing with examples in MATLAB. Electrical engineering and applied signal processing series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2003. ISBN 0-8493-1091-1. 336 (est.) pp. LCCN TK5102.9 .S719 2003.

Stein:2008:CWC

[Ste08] William A. Stein. Can we create a viable free open source alternative to Magma, Maple, Mathematica and Matlab? In Jeffrey [Jef08], pages 5–6. ISBN 1-59593-904-0. LCCN ????

Stewart:2009:FMP

[Ste09]  G. W. Stewart.  Flap:

a Matlab package for adjustable precision floatingpoint arithmetic. Report, Department of Computer Science, University of Maryland, College Park, MD, USA, 2009. URL

                                        http://www.cs.umd.edu/~
                                        stewart/flap/flap.html.

Stenger:2010:HSN

[Ste10] Frank Stenger. Handbook of Sinc Numerical Methods. Chapman and Hall/CRC numerical analysis and scientific computation series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-2158-5 (hardback), 1-4398-2159-3 (e-book). xx + 463 pp. LCCN QA372 .S8195

                                        2010. URL http://www.

crcpress.com/product/isbn/
9781439821589.

Stein:2013:SCV

[Ste13] William Stein. Sage: creating a viable free open source alternative to Magma, Maple, Mathematica, and MATLAB. In Felipe Cucker, Teresa Krick, Allan Pinkus, and Agnes Szanto, editors, Foundations of computational mathematics, Budapest 2011, volume 403 of London Mathematical Society lecture note series, pages 230–238. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-60407-9 (paperback), 1-139-09540-4 (e-book), 1-139-61690-0. LCCN QA297 .F635 2011.

Son:2001:RDR

[STF01] Jae Sok Son, Gabriel Thomas, and Benjamin C. Flores. Range-Doppler radar imaging and motion compensation. Artech House radar library. Artech House Inc., Boston, MA, USA, 2001. ISBN 1-58053-102-4. xi + 238 pp. LCCN TK6592.D6 S66 2001.

Sticklen:2004:TPS

[Sti04] Jon H. Sticklen. Technical problem solving: with Matlab and Excel. Great Lakes Press, Wildwood, MO, USA, 2004. ISBN 1-881018-80-6. ???? pp.

Stone:2013:BRT

[Sto13] James V. Stone. Bayes’ rule: a tutorial introduction to Bayesian analysis. Sebtel Press, Lexington, KY, USA, 2013. ISBN 0-9563728-4-8 (paperback). 170 pp. LCCN QA279.5 .S766 2013.

Strang:1993:ILA

[Str93]  Gilbert Strang.  Intro-

duction to Linear Algebra. Wellesley-Cambridge Press, Wellesley, MA, USA, 1993. ISBN 0-9614088-5-5. viii + 472 pp. LCCN QA184 .S78 1993.

Strang:2003:ILA

[Str03] Gilbert Strang. Introduction to linear algebra. Wellesley-Cambridge Press, Wellesley, MA, USA, third edition, 2003. ISBN 0-9614088-9-8. viii + 568 pp. LCCN QA184 .S77 2003.

Strang:2006:LAA

[Str06]  Gilbert Strang.  Lin-

ear algebra and its applications. Thomson– Brooks/Cole, Belmont, CA, fourth edition, 2006. ISBN 0-03-010567-6 (student edition). viii + 487 pp. LCCN QA184 .S8 2006.

Studt:1996:WTO

[Stu96] Tim Studt. Wavelet technology offers designers alternative to Fourier analysis. Research & Development, 38(6):51, May 1996.

CODEN REDEEA. ISSN 0746-9179. Discusses MAT-LAB’s Wavelet Toolbox and Strang and Nguyen’s book [SN96].

Sturm:1999:USM

[Stu99] Jos F. Sturm. Using Se-DuMi 1.02, a MATLAB toolbox for optimization over symmetric cones. Optimization Methods and Software, 11/12(1-4):625–653, 1999. CODEN OMSOE2. ISSN 1055-6788. Interior point methods.

Swiercz:1993:TAW

[SU93] M. Swiercz and B. Ulanicki.

Transient analysis in water distribution networks using the SIMULINK package. In Coulbeck [Cou93], pages 151–166. ISBN 0-86380-154-4. LCCN TD353.I524 1993.

Szymkat:1994:MBE

[SU94] M. Szymkat and T. Uhl.

MATLAB based environmental for mechatronic robot modelling. In Anonymous [Ano94a], pages 817– 829. ISBN ???? LCCN ???? Three volumes.

Su:1996:AF

[Su96] Kendall Su. Analog filters.

Chapman and Hall, Ltd., London, UK, 1996. ISBN 0-412-63840-1. xii + 366 pp. LCCN TK7872.F5 S798

Sutton:2017:VEM

[Sut17] Oliver J. Sutton. The virtual element method in 50 lines of MATLAB. Numerical Algorithms, 75(4):1141– 1159, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/s11075-016-0235-
3.pdf.

Stark:1994:PRP

[SW94] Henry Stark and John W. (John William) Woods. Probability, Random Processes, and Estimation Theory for Engineers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1994. ISBN 0-13-728791-7. xiv + 618 pp. LCCN TA340 .S71 1994. US$48.00.

Stone:1995:DDF

[SW95] D. J. Stone and C. R. Walters. DSP demonstration facility using a sound card and Matlab. In Anonymous [Ano95t], pages 4/1– 4/4. ISBN ???? LCCN ????

Stavrakos:1994:ICC

[SWG^{+}94] Nicholas J. Stavrakos, Bruce C. Wheeler, Sherald H. Gordon, Donald T. Wicklow, Richard V. Greene, and Robert B. Schudy. Identification of contaminated corn from FTIR spectra:

A MATLAB tool for pattern recognition and neural network studies. In Dagli [Dag94], pages 779–786 (or 779–784??). ISBN 0-7918-0045-8. LCCN ???? Three volumes.

Stark:2002:PRP

[SWrpSetfe02] Henry Stark, John W.

(John William) Woods, random processes Stark, Henry, Probability, and estimation theory for engineers. Probability and random processes with applications to signal processing. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2002. ISBN 0-13-020071-9. xv + 689 pp. LCCN TK5102.9 .S717 2002.

StGeorge:1997:SCP

[SWS97] B. St. George, E. Wooten, and L. Sellami. Speech coding and phoneme classification using MATLAB and NeuralWorks. In EP Innovations [EP 97], pages 12– ?? ISBN 0-7803-4087-6, 0-7803-4086-8, 0-7803-4088-4, 0-7803-4089-2. ISSN 0190-5848. LCCN T62 .F76 1997. Three volumes.

Semaan:2020:SSC

[SY20] Richard Semaan and Vikas Yadav. SCOUT: Signal Correction and Uncertainty Quantification Toolbox in MATLAB. SoftwareX, 11 (??):Article 100474, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711019303711.

Sydow:1995:SAS

[Syd95] A. Sydow, editor. Systems analysis and simulation: Symposium — June 1995, Berlin, Germany, volume 18/19 of Systems Analysis Modelling Simulation. Gordon and Breach, ????, 1995. CODEN SAMSEC. ISBN 2-88449-064-7. ISSN 0232-9298. LCCN ????

Stork:2004:CMM

[SYTD04] David G. Stork, Elad Yom-Tov, and Richard O. Duda. Computer manual in MATLAB to accompany Pattern Classification. Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-42977-5. ix + 134 pp. LCCN Q327 .S76 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley047/2004044014.
                                        html;   http://www.loc.
                                        gov/catdir/description/
                                        wiley041/2004044014.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley041/2004044014.
                                        html.

Su:2024:HSA

[SYW^{+}24] Zhuo Su, Zehong Yu, Dongyan Wang, Yixiao Yang, Rui Wang, Wanli Chang, Aiguo Cui, and Yu Jiang. HSTCG: Stateaware simulink model test case generation with heuristic strategy. IEEE Transactions on Software Engineering, 50(12):3088–3103, December 2024. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Szabo:2004:DUI

[Sza04] Thomas L. Szabo. Diagnostic ultrasound imaging: inside out. Academic Press series in biomedical engineering. Elsevier Academic Press, Amsterdam, The Netherlands, 2004. ISBN 0-12-680145-2. xxii + 549 pp. LCCN RC78.7.U4 S98 2004.

Stripinis:2021:MEA

[SZCP21] Linas Stripinis, Julius Zilinskas, Leocadio G. Casado, and Remigijus Paulavicius. On MATLAB experience in accelerating DIRECT–GLce algorithm for constrained global optimization through dynamic data structures and parallelization. Applied Mathematics and Computation, 390(??):Article 125596, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305518.

Smiljanic:2014:MBP

[SZM^{+}14] J. Smiljanić, M. Zezelj, V. Milanović, J. Radovanović, and I. Stanković. MATLABbased program for optimization of quantum cascade laser active region parameters and calculation of output characteristics in magnetic field. Computer Physics Communications, 185(3):998–1006, March 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465513003706.

Szymkat:1993:KWW

[Szy93] Maciej Szymkat. Komputerowe wspomaganie w projektowaniu ukladów regulacji [English: Computer-Aided Control Systems Design]. Wydawnictwa Naukowo-Techniczne, Warszawa, Poland, 1993. ISBN 83-204-1655-8. LCCN ????

Thierry:2015:IDO

[TACA15] Bertrand Thierry, Xavier Antoine, Chokri Chniti, and Hasan Alzubaidi. μ-diff: an open-source Matlab toolbox for computing multiple scattering problems by disks. Computer Physics Communications, 192(??): 348–362, July 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465515001125.

Taylor:1994:PSS

[Tay94] Fred Taylor. Principles of Signals and Systems. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-063197-2, 0-07-911171-

8. xx + 562 pp. LCCN TK5102.5.T36 1994.

Taylor:1995:RHS

[Tay95] James H. Taylor. Rigorous handling of state events in MATLAB. In IEEE [IEE95b], pages 156–

161. ISBN 0-7803-2551-6, 0-7803-2550-8. LCCN TJ212.2.I3247 1995.

Taylor:1999:BRC

[Tay99] Peter R. Taylor. Book review: Y. C. Pao, Engineering analysis: Interactive methods and programs with FORTRAN, QuickBasic, MATLAB, and Mathematica (1999) CRC Press,Bristol 0-8493-2016-X. Computer Physics Communications, 120(2–3):271– 272, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465599002374.

Teolis:1995:WPW

[TB95] Anthony Teolis and John S.

Baras. Wavelet processing workstation: an interactive MATLAB-based computational tool for wavelet processing. Proceedings of the SPIE — The International Society for Optical Engineering, 2491/1: 592–603, 1995. CODEN PSISDG. ISBN 0-8194-1844-7. ISSN 0277-786X (print), 1996-756X (electronic).

Trefethen:1997:NLA

[TB97] Lloyd N. (Lloyd Nicholas) Trefethen and David Bau. Numerical linear algebra. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-361-7. xii + 361 pp. LCCN QA184 .T74 1997.

Takrouni:2021:SID

[TBH21] Manel Takrouni, Rim Bouhouch, and Salem Hasnaoui. Simulink implementation of the data distribution service for vehicular controllers on top of GBE and AFDX. The Computer Journal, 64(6):860– 879, June 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/64/6/860/6124919.

Tuschak:1994:NCC

[TBHS94] R. Tuschak, R. Bars, M. Habermayer, and B. Szucs. A new curriculum in control education supported by MATLAB. In Ichikawa and Furuta [IF94], pages 185– 188. ISBN 0-08-042230-6. LCCN TJ212.2 .A393 1995.

Taylor:1997:EMT

[TC97a] J. H. Taylor and C. Chan.

Enhanced MATLAB tools for linear and nonlinear system stability. In Boullart et al. [BLM97], pages 293– 298. ISBN 0-08-042383-3. LCCN TJ212.2 .C33 1997.

Taylor:1997:MTL

[TC97b] J. H. Taylor and C. Chan.

MATLAB tools for linear and nonlinear system stability theorem implementation. In IEEE [IEE97e], pages 42–47. ISBN 0-7803-3877-4, 0-7803-3876-6, 0-7803-3878-2. ISSN 1085-1992. LCCN TJ212.2.C565 1997. IEEE catalog number: 97CH36055.

Chau:2004:CBW

[tC04] Foo tim Chau, editor.

Chemometrics: from basics to wavelet transform, volume 164 of Chemical analysis. Wiley, New York, NY, USA, 2004. ISBN 0-471-20242-8. xiv + 316 pp. LCCN QD79.I5 C44 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley046/2003002429.
html;   http://www.loc.
gov/catdir/description/
wiley0310/2003002429.html;
http://www.loc.gov/catdir/
toc/wiley032/2003002429.
html.

Tisan:2013:SNN

[TC13] A. Tisan and M. Cirstea. SOM neural network design – a new Simulink library based approach targeting FPGA implementation. Mathematics and Computers in Simulation, 91:134–149, May 2013. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0378475412001139.

Tong:2022:GAI

[TCD^{+}22] Xin Tong, Sou-Cheng T. Choi, Yuhan Ding, Fred J. Hickernell, Lan Jiang, Lluís Antoni Jiménez Rugama, Jagadeeswaran Rathinavel, Kan Zhang, Yizhi Zhang, and Xuan Zhou. Guaranteed

Automatic Integration Library

(GAIL): an open-source MATLAB library for function approximation, optimization, and integration. Journal of Open Research Software, 10(1):7– ??, July 29, 2022. CO-DEN ???? ISSN 2049-9647. URL https:/

/openresearchsoftware.
metajnl.com/articles/10.
5334/jors.381/.

Terven:2016:KKT

[TCE16] Juan R. Terven and Diana M. Córdova-Esparza. Kin2. A Kinect 2 toolbox for MATLAB. Science of Computer Programming, 130(??):97–106, November 15, 2016. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642316300569.

Terven:2021:KAK

[TCE21] Juan R. Terven and Diana M. Córdova-Esparza. KinZ: an Azure Kinect toolkit for Python and Matlab. Science of Computer Programming, 211(??):??, November 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167642321000952.

Thomson:1998:TVA

[TD98] William Tyrrell Thomson and Marie Dillon Dahleh. Theory of vibration with applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fifth edition, 1998. ISBN 0-13-651068-X. xiii + 524 pp. LCCN TA355 .T47 1998.

Teolis:1998:CSP

[Teo98]  Anthony Teolis.  Com-

putational signal processing with wavelets. Applied and numerical harmonic analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 1998. ISBN 0-8176-3909-8 (Boston), 3-7643-3909-8 (Basel). xxiv + 324 pp. LCCN TK5102.9 .T44 1998.

Terzuoli:1994:CPA

[Ter94] Andy Terzuoli, editor. Conference proceedings: 10th Annual Review of Progress in Applied Computational Electromagnetics; at the Doubletree Hotel & Convention Center, Monterey, CA, March 21–26, 1994, volume 2. University of Alabama, Tuscaloosa, AL, USA, 1994. ISBN ???? LCCN ???? Two volumes.

Tewari:2002:MCD

[Tew02] Ashish Tewari. Modern control design with MATLAB and SIMULINK. Wiley, New York, NY, USA, 2002. ISBN 0-471-49679-0. xiii + 503 pp. LCCN QA402.3 .T44 2002. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/wiley037/2002284777.
                                        html;   http://www.loc.
                                        gov/catdir/toc/wiley023/
                                        2002284777.html.

Takakura:2002:CUC

[TF02] T. (Tadashi) Takakura and Wei Fang. Climate under cover. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, second edition, 2002. ISBN 1-4020-0845-7 (hardcover), 1-4020-0846-5 (paperback). xi + 190 pp. LCCN SB416 .T34 2002; SB416 .T34 2002X ESCI.

Trauth:2010:MRE

[TG10] Martin Trauth and Marwan Gebber. Matlab recipes for earth sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2010. ISBN 3-642-12761-4. ???? pp. LCCN ????

Trauth:2006:MRE

[TGM06] Martin H. Trauth, Robin Gebbers, and Norbert Marwan. MATLAB recipes for earth sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-27983-0 (hardback). xi + 237 pp. LCCN QE33.2.S82 T7 2006. Includes CD-ROM.

Tamplin:2001:ACW

[TH01] M. R. Tamplin and A. Hamilton. Ant circuit world: An ant algorithm MATLAB^{T M} toolbox for the design, visualisation and analysis of analogue circuits. Lecture Notes in Computer Science, 2210:151–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2210/22100151.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2210/22100151.
pdf.

Taylor:2009:CCT

[TH09] Alan Taylor and Desmond J. Higham. CONTEST: a controllable test matrix toolbox for MATLAB. ACM Transactions on Mathematical Software, 35(4):26:1–26:17, February 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

MathWorks:1990:PMS

[The90] The MathWorks, Inc. PRO-MATLAB for Sun Workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, January 31, 1990. ISBN ???? various pp. LCCN CM.0.M3.P7.

MathWorks:1992:MHPb

[The92a] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: External Interface Guide: for UNIX workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1992. ISBN ???? LCCN QA188 .M384 1993.

MathWorks:1992:MHPd

[The92b] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: External Interface Guide: for UNIX workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1992. ISBN ???? LCCN QA188 .M384 1993. MathWorks:1992:MHPc

[The92c] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: Reference Guide. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1992. ISBN ???? iv + 548 pp. LCCN QA188 .M387 1992. MathWorks:1992:MHPa

[The92d] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: User’s Guide: for UNIX workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1992. ISBN ???? LCCN QA188 .M389 1992.

MathWorks:1992:SEMb

[The92e] The MathWorks, Inc. The Student Edition of Matlab for Macintosh Computers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-855990-2. xiv + 494 pp. LCCN ????

MathWorks:1992:SEMc

[The92f] The MathWorks, Inc. The Student Edition of Matlab for MS-DOS Personal Computers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-855974-0, 0-13-855982-1. xiv + 494 pp. LCCN QA297 .S8433 1992. System requirements for disk: IBM PC/XT, PC/AT, PS/2, or compatible MS-DOS Adapter (CGA), Enhanced Graphics Adapter (EGA), Hercules Monochrome Graphics Card (HGC), Video Graphics Array (VGA), AT and T 6300 Graphics Card (ATT), or compatible; color graphics are available on systems with a VGA or EGA (EGA must have 256 KB video memory); Epson 9-pin, HP Laserjet, or fully compatible printer.

MathWorks:1992:SEMa

[The92g] The MathWorks, Inc. The Student Edition of MAT-LAB: Student User Guide. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-856006-4. xiv + 494 pp. LCCN QA297 .S8434 1992.

Therrien:1992:DRS

[The92h] Charles W. Therrien. Discrete Random Signals and Statistical Signal Processing. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-852112-3. xx + 727 pp. LCCN TK5102.5 .T48 1992.

MathWorks:1993:MHPa

[The93a] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: Building a Graphical User Interface. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1993. ISBN ???? LCCN QA188 .M383 1993.

MathWorks:1993:MHPb

[The93b] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: New Features Guide, version 4.0. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1993. ISBN ???? 25 pp. LCCN QA188 .M384 1993.

MathWorks:1993:MHPc

[The93c] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: Release Notes, Version 4.1: for UNIX Workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1993. ISBN ???? 59 pp. LCCN QA188 .M388 1993.

Mathworks:1996:SES

[The96] The MathWorks, Inc. The Student Edition of SIMULINK User’s Guide. The Matlab curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-452435-7, 0-13-452427-6 (with Windows software), 0-13-452310-5 (with Macintosh and Power Macintosh software). ix + 240 pp. LCCN QA76.76.C65S88 1996.

MathWorks:1997:SEM

[The97] The MathWorks, Inc. The Student Edition of Matlab: High-Performance Numeric Computation and Visualization Software: Version 5 User’s Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN ???? ???? pp. LCCN ????

MathWorks:1998:SES

[The98] The MathWorks, Inc. The student edition of Simulink: dynamic system simulation for MATLAB: user’s guide. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-659673-8 (Windows 95/NT CD-ROM), 0-13-659699-1 (Macintosh/Power Macintosh CD-ROM). xiii + 224 pp. LCCN QA76.9.C65S78 1998.

Thinh:1995:CMM

[Thi95] N. X. Thinh. A collection of MATLAB and MATHCAD simulation for environmentally relevant systems. In Sydow [Syd95], pages 841– 844. CODEN SAMSEC.

ISBN 2-88449-064-7. ISSN 0232-9298. LCCN ????

Hart:2019:UAM

[tHLMN19] Cornelis Marcel Pieter ’t Hart, Georgios Leontaris, and Oswaldo Morales-Nápoles. Update (1.1) to ANDURIL — a MAT-LAB toolbox for ANalysis and Decisions with UnceRtaInty: Learning from expert judgments: AN-DURYL. SoftwareX, 10 (??):Article 100295, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711019302419.

Thomaseth:1994:MAP

[Tho94] K. Thomaseth. Modelling and analysis of physiological systems with MATLAB: Software development for Macintosh computer. In Patterson [Pat94], pages 431–432. ISBN 0-08-042224-

1. LCCN QP39 .M63 1994.

Thomson:2000:ITP

[Tho00] William J. Thomson. Introduction to transport phenomena. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-454828-0. xv + 509 pp. LCCN TP156.T7 T48 2000.

Thompson:2004:IFE

[Tho04] Erik G. Thompson. An introduction to the finite element method: theory, programming, and applications. Wiley, New York, NY, USA, 2004. ISBN 0-471-26753-8 (ClOTH/CD-ROM). xiii + 343 pp. LCCN TA347.F5 .T47 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley041/2003063425.
html;   http://www.loc.
gov/catdir/toc/wiley041/
2003063425.html.

Thyagarajan:2010:SIV

[Thy10] K. S. Thyagarajan. Still image and video compression with MATLAB. Wiley, New York, NY, USA, 2010. ISBN 0-470-48416-0. ???? pp. LCCN TA1638 .T48 2010.

Tibaldi:1993:NIM

[Tib93] Marco Tibaldi. Note Introduttive a MATLAB e Control System Toolbox [English: Introductory Notes for MATLAB and the Control System Toolbox]. Progetto Leonardo, Bologna, Italy, 1993. ISBN ???? LCCN ????

Tang:1992:IMF

[TJML92] R. Tang, N. A. Jalel, A. R. Mirzai, and J. R. Leigh. Identification and modelling of fermentation process using MATLAB: a case study. IFAC Symposia Series, 10:331–334, 1992. CO-DEN ISYSEK. ISBN 0-08-041710-8. ISSN 0962-9505.

Taylor:1997:MSH

[TK97] J. H. Taylor and D. Kebede.

Modeling and simulation of hybrid systems in MAT-LAB. In Gertler et al. [GCP97], pages 275–280. ISBN 0-08-042926-2 (plenary vol.), 0-08-042909-2 (vol. A), 0-08-042910-6 (vol. B), 0-08-042911-4 (vol. C), 0-08-042912-2 (vol. D), 0-08-042913-0 (vol. E), 0-08-042914-9 (vol. F), 0-08-042915-7 (vol. G), 0-08-042916-5 (vol. H), 0-08-042917-3 (vol. I), 0-08-042918-1 (vol. J), 0-08-042919-X (vol. K), 0-08-042920-3 (vol. L), 0-08-042921-1 (vol. M), 0-08-042922-X (vol. N), 0-08-042923-8 (vol. O), 0-08-042924-6 (vol. P), 0-08-042925-4 (vol. Q). LCCN ????

Tsang:2000:SEW

[TKD00] Leung Tsang, Jin Au Kong, and Kung-Hau Ding. Scattering of electromagnetic waves. $p Theories and applications. Wiley series in remote sensing. Wiley, New York, NY, USA, 2000. ISBN 0-471-38799-1 (cloth). xiv + 426 pp. LCCN QC665.S3 T73 2000. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley037/00040863.
html;   http://www.loc.

                                        gov/catdir/toc/onix06/
                                        00040863.html.

Turjan:2007:CIC

[TKD07] Alexandru Turjan, Bart Kienhuis, and Ed Deprettere. Classifying interprocess communication in process network representation of nested-loop programs. ACM Transactions on Embedded Computing Systems, 6(2):13:1–13:??, May 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tilbury:1997:DIW

[TM97] D. Tilbury and W. Messner. Development and integration of Web-based software tutorials for an undergraduate curriculum: Control tutorials for Matlab. In EP Innovations [EP 97], pages 1070–1075. ISBN 0-7803-4087-6, 0-7803-4086-8, 0-7803-4088-4, 0-7803-4089-2. ISSN 0190-5848. LCCN T62 .F76 1997. Three volumes.

Trefethen:1999:MMM

[TMC^{+}99] Anne E. Trefethen, Vijay S. Menon, Chi-Chao Chang, Grzegorz J. Czajkowski, Chris Myers, and Lloyd N. Trefethen. MultiMATLAB: MATLAB on multiple processors. Technical report, Cornell Theory Center and Computer Science Department, Cornell University, Ithaca, NY, August 19, 1999. URL

http://www.cs.cornell.
edu/Info/People/lnt/multimatlab.
html.

Tolic-Norrelykke:2004:MPP

[TNBSF04] Iva Marija Tolić-Nørrelykke, Kirstine Berg-Sørensen, and Henrik Flyvbjerg. Matlab program for precision calibration of optical tweezers. Computer Physics Communications, 159(3):225–240, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465504001043.

Tobin:2011:DSA

[Tob11] Stephen M. Tobin. DC servos: application and design with MATLAB. Taylor and Francis, Boca Raton, FL, USA, 2011. ISBN 1-4200-8003-2. ???? pp. LCCN TJ214 .T63 2011.

Torkelsson:2002:CP

[Tor02]  Björn Torkelsson.  The

CONLAB Project. Technical report, Department of Computing Science, Umeå universitet, S-901 87 Umeå, Sweden, 2002.

URL http://www.cs.umu.
se/research/conlab/;
mailto:conlab@cs.umu.se.

From the Web site: “CON-LAB (CONcurrent LABoratory) is an environment for developing algorithms for parallel computer architectures. It is an interactive environment in which one can simulate MIMD architectures with distributed memory and communication with message passing, as well as MIMD architectures with shared memory.

CONLAB is an extension of MATLAB with control structures for expressing parallel execution of programs and primitives for message passing, use of shared memory and synchronization. The language used in CONLAB is a language that is close to the informal algorithm specification languages that many algorithm designers use.

… CONLAB Compiler: A prototype compiler, CLC, has been developed for CONLAB. The compiler translates its input files into C-language files to which it applies a suitable C compiler. In order to achieve high performance the computational kernels of executable code are built on level 1-2-3 BLAS (Basic Linear Algebra Subprograms) and LAPACK (Linerar Algebra Package). Communication via explicit message passing is performed by PVM.

Programs generated by CLC has successfully been executed on heterogeneous workstations. Work is undergoing to port the compiler to IBM SP/2.”.

Toth:2008:QVP

[Tót08] Géza Tóth. QUBIT4MATLAB V3.0: a program package for quantum information science and quantum optics for Matlab. Computer Physics Communications, 179(6):430–437, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465508001367.

Townsend:2016:ROP

[Tow16] Alex Townsend. A review of Orthogonal Polynomials in MATLAB: Exercises and Solutions by Gautschi. Web document., November

5, 2016. URL http://pi.
math.cornell.edu/~ajt/
papers/Gautschi.pdf.

Talole:2003:GLD

[TP03] S. E. Talole and S. B.

Phadke. Generating L^{A}TEX documents through Matlab. TUGboat, 24(2):245– 248, 2003. ISSN 0896-3207.

Torres:1996:OPF

[TQ96] G. L. Torres and V. H.

Quintana. Optimal power flow via interior point methods: An educational tool in Matlab. In Malkinson [Mal96], pages 996– 999. ISBN 0-7803-3143-5, 0-7803-3144-3. LCCN TK7801.C36 1996. Two volumes.

Thomas:2004:ADL

[TR04] Roland E. Thomas and Albert J. Rosa. The analysis and design of linear circuits: Laplace early. Wiley, New York, NY, USA, fourth edition, 2004. ISBN 0-471-27213-2; 0-471-43299-7. xviii + 788 + 38 pp. LCCN TK454 .T466 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley038/2003270625.
html;   http://www.loc.
gov/catdir/toc/wiley032/
2003270625.html;   http:
//www.loc.gov/catdir/toc/
wiley041/2004268992.html.

Trappenberg:2002:FCN

[Tra02] Thomas P. Trappenberg. Fundamentals of computational neuroscience. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-851582-0 (hardcover), 0-19-851583-9 (paperback). xvi + 338 pp. LCCN QP357.5 .T746 2002; QP 357.5 .T746 2002X GER-STM. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/fy031/2002070356.html.

Tranter:2004:PCS

[Tra04] William H. Tranter, editor. Principles of communication systems simulation with wireless applications, volume 16 of Prentice Hall communications engineering and emerging technologies series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-494790-8. xxii + 778 pp. LCCN TK5102.5 .P673 2004.

Trappenberg:2010:FCN

[Tra10] Thomas P. Trappenberg.

Fundamentals of computational neuroscience. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2010. ISBN 0-19-956841-3 (paperback). xxv + 390 pp. LCCN QP357.5 .T746 2010.

Taylor:2011:CAS

[TRD11] Ken Taylor, Scott Rickard, and Konstantinos Drakakis. Costas arrays: Survey, standardization, and MATLAB toolbox. ACM Transactions on Mathematical Software, 37(4):41:1–41:31, February 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Trefethen:1986:MPC

[Tre86] Lloyd N. Trefethen. MAT-LAB programs for CF approximation. In C. K. Chui, L. L. Schumaker, and J. D.

Ward, editors, Approximation Theory V, pages ??–?? (of xviii + 654). Academic Press, New York, NY, USA, 1986. ISBN 0-12-174581-3. LCCN QA221 .I561 1986.

Trefethen:2000:SMM

[Tre00] Lloyd N. Trefethen. Spectral Methods in MAT-LAB. Software, environments, tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-465-6 (paperback), 0-89871-959-3 (e-book). xvi + 165 pp. LCCN QA377 .T65 2000.

Trefethen:2010:HTQ

[Tre10] Lloyd N. Trefethen. Householder triangularization of a quasimatrix. IMA Journal of Numerical Analysis, 30(4):887–897, October 2010. CODEN IJ-NADH. ISSN 0272-4979 (print), 1464-3642 (electronic). URL http://

imajna.oxfordjournals.
org/content/30/4/887.full.
pdf+html.

Trefethen:2015:CNF

[Tre15] Lloyd N. Trefethen. Computing numerically with functions instead of numbers. Communications of the ACM, 58(10):91–97, October 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.
acm.org/magazines/2015/
10/192390/fulltext.

Trif:2008:MPS

[Tri08] Damian Trif. Matlab package for the Schrödinger equation. Journal of Mathematical Chemistry, 43(3):1163–1176, March 2008. CODEN JM-CHEG. ISSN 0259-9791 (print), 1572-8897 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10910-007-9266-2.

Thielicke:2014:PPT

[TS14] William Thielicke and Eize Stamhuis. PIVlab — towards user-friendly, affordable and accurate digital particle image velocimetry in MATLAB. Journal of Open Research Software, 2(1):e30–??, October 16, 2014. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.bl/.

Thielicke:2021:PIV

[TS21] William Thielicke and René Sonntag. Particle image velocimetry for MATLAB: Accuracy and enhanced algorithms in PIVlab. Journal of Open Research Software, 9(1):12–??, May 31, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.

                                        metajnl.com/articles/10.
                                        5334/jors.334/.

Teimouri:2021:PFE

[TSA21] Mehdi Teimouri, Zahra Seyedghorban, and Fatemeh

                                        Amirjani. Fragments-Expert:

a graphical user interface MATLAB toolbox for classification of file fragments. Concurrency and Computation: Practice and Experience, 33(9):e6154:1– e6154:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tripakis:2005:TDT

[TSCC05] Stavros Tripakis, Christos Sofronis, Paul Caspi, and Adrian Curic. Translating discrete-time Simulink to Lustre. ACM Transactions on Embedded Computing Systems, 4(4):779– 818, November 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tian:1993:CAM

[TSM93] Daquan Tian, Soroosh Sorooshian, and Donald E. Myers. Correspondence analysis with Matlab. Computers and Geosciences, 19(7):1007–1022, August 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Trefethen:1996:MMM

[TT^{+}96] L. N. Trefethen, A. E. Trefethen, et al. MultiMAT-LAB: MATLAB on multiple processors. Technical Report CTC96TR293, Cornell Theory Center, Cornell University, Ithaca, NY, USA, ???? 1996.

Turin:2004:PAM

[TtsT04] William Turin and William Digital transmission systems Turin. Performance analysis and modeling of digital transmission systems. Information technology– transmission, processing, and storage. Kluwer Academic… Plenum Publishers, New York, NY, USA, 2004. ISBN 0-306-48191-X. xxi + 441 pp. LCCN TK5103.7 .T8675 2004.

Toh:1999:SMS

[TTT99] K. C. Toh, M. J. Todd, and R. H. Tütüncü. SDPT3—a MATLAB software package for semidefinite programming, version 1.3. Optimization Methods and Software, 11/12(1-4):545–581, 1999. CODEN OMSOE2. ISSN 1055-6788. Interior point methods.

Tsoukalas:1997:FNA

[TU97] Lefteri H. Tsoukalas and Robert E. Uhrig. Fuzzy and neural approaches in engineering. Adaptive and learning systems for signal processing, communications, and control. Wiley, New York, NY, USA, 1997. ISBN 0-471-16003-2 (cloth). xix + 587 pp. LCCN QA76.87 .T76 1997. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley041/96014102.
html;   http://www.loc.
gov/catdir/description/
wiley031/96014102.html;
 http://www.loc.gov/catdir/
toc/wiley022/96014102.
html.

Turner:2000:GSC

[Tur00] P. R. (Peter R.) Turner. Guide to scientific computing. Macmillan mathematical guides. MacMillan Publishing Company, New York, NY, USA, second edition, 2000. ISBN 0-333-79450-8. ix + 301 pp. LCCN QA297 .T88 2000.

Turner:2001:GSC

[Tur01] P. R. (Peter R.) Turner. Guide to scientific computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2001. ISBN 0-8493-1242-6 , 0-333-79450-8. ix + 301 pp. LCCN QA297 .T88 2001.

Turcotte:1998:CAM

[TW98] Louis H. Turcotte and H. B. (Howard B.) Wilson. Computer applications in mechanics of materials using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-749060-7. xi + 321 pp. LCCN TA405 .T787 1998.

Trappe:2002:ICC

[TW02] Wade Trappe and Lawrence C.

Washington. Introduction to cryptography: with coding theory. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-061814-4. xiii + 490 pp. LCCN QA268 .T73 2002.

Trappe:2006:ICC

[TW06] Wade Trappe and Lawrence C.

Washington. Introduction to Cryptography: with Coding Theory. Pearson Prentice Hall, Upper Saddle River, NJ, USA, second edition, 2006. ISBN 0-13-186239-1, 0-13-198199-4 (paperback). xiv + 577 pp. LCCN QA268 .T73 2006.

Truskey:2004:TPB

[TYK04] George A. Truskey, Fan Yuan, Ph D., and David F. Katz. Transport phenomena in biological systems. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-042204-5. xxi + 793 pp. LCCN QH509 .T78 2004.

Teng:2016:EGP

[TYL^{+}16] Ben Teng, Can Yang, Jiming Liu, Zhipeng Cai, and Xiang Wan. Exploring the genetic patterns of complex diseases via the integrative genome-wide approach. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 13(3):557–564, May 2016. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Uhl:1995:CAD

[UB95] T. Uhl and M. Betemps. Computer aided design of robotic systems with pneumatic actuator—SIMULINK LIBRARY. In Sydow [Syd95], pages 551–554. CO-DEN SAMSEC. ISBN 2-88449-064-7. ISSN 0232-9298. LCCN ????

Ucar:2010:MPI

[UÇA10] Bora Uçar, Ümit V. Çatalyürek, and Cevdet Aykanat. A matrix partitioning interface to PaToH in MATLAB. Parallel Computing, 36(5–6):254– 272, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

UCLA:2004:EAU

[UCL04] UCLA. Engineering analysis using Matlab/Simulink: Engineering 851.88, Reg. No. Q6965, June 15–July 29, 2004. Short course program, University of California, University Extension, Dept. of Engineering, Information Systems and Technical Management, Short Course Program, Los Angeles, CA, USA, 2004. UCLA Extension presents the onsite short course at Boeing Huntington Beach, CA.

Umez-Eronini:1999:SDC

[UE99] Eronini Umez-Eronini. System dynamics and control. PWS Publishing Company, Pacific Grove, 1999. ISBN 0-534-94451-5. xiv + 993 pp. LCCN TJ213 .U453 1999.

Uhlig:2002:TLA

[Uhl02] Frank Uhlig. Transform linear algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-041535-9. xviii + 503 pp. LCCN QA184.2 .U48 2002.

Uberhuber:2005:M

[ÜKP05] Christoph W. Überhuber, Stefan Katzenbeisser, and Dirk Praetorius. MATLAB 7. Springer, Wien / New York, 2005. ISBN 3-211-21137-3. ix + 309 pp. LCCN ????

USENIX:1994:PUS

[USE94] USENIX Association, editor. Proceedings of the USENIX Symposium on Very High Level Languages (VHLL): October 26–28, 1994, Santa Fe, New Mexico, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

USENIX:1999:PCD

[USE99] USENIX, editor. Proceedings of the 2nd Conference on Domain-Specific Languages (DSL ’99), October 3–5, 1999, Austin, Texas, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.
org/publications/library/
proceedings/dsl99/.

Umansky:2012:NAM

[UW12] Moti Umansky and Daphne Weihs. Novel algorithm and MATLAB-based program for automated power law analysis of single particle, time-dependent meansquare displacement. Computer Physics Communications, 183(8):1783–1792, August 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465512001063.

Vassiliadis:1994:TSD

[VA94] C. A. Vassiliadis and D. Akos. Twin spiral distinction via backpropagation networks using the MATLAB Neural Network Toolbox. In Chen et al. [CAC94], pages 643–648. ISBN 2-907669-30-3. LCCN ????

Vaccaro:1995:DCS

[Vac95] Richard J. Vaccaro. Digital Control: a State-Space Approach. McGraw-Hill series in electrical and computer engineering. Control theory. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-066781-0. xvii + 455 pp. LCCN TJ223.M53 V32 1995.

VanLoan:1992:CFF

[Van92a] Charles F. Van Loan. Computational Frameworks for the Fast Fourier Transform. Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-285-8 (paperback), 1-61197-099-7 (e-book). xiii + 273 pp. LCCN QA403.5 .V35 1992.

VanWyk:1992:APA

[Van92b] D. Van Wyk, editor.

AFRICON ’92: proceedings, Third Africon conference, 22–24 September, 1992, Convention Centre, Royal Swazi Sun, Ezulwini Valley, Swaziland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0835-2, 0-7803-0836-0, 0-7803-0837-9. LCCN TK5 .A47 1992.

VanLoan:1997:ISC

[Van97] Charles F. Van Loan. Introduction to scientific computing: a matrix-vector approach using MATLAB. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-125444-8. xi + 347 pp. LCCN QA76.9.M35V375 1997.

VanLoan:2000:ISC

[Van00a] Charles F. Van Loan. Introduction to scientific computing: a matrix–vector approach using MATLAB. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2000. ISBN 0-13-949157-0. xi + 367 pp. LCCN QA76.9.M35 V375 2000.

Vancura:2000:MTD

[Van00b] Tobias Vancura. Matlab — a tool for doing numerics. Linux Journal, 70:??, February 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

VanderSluis:2001:TPS

[Van01a] Lou Van der Sluis. Transients in power systems. Wiley, New York, NY, USA, 2001. ISBN 0-471-48639-6. xiv + 207 pp. LCCN TK3226 .V23 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;

http://www.loc.gov/catdir/
description/wiley036/2001026198.
html;   http://www.loc.
gov/catdir/toc/wiley021/
2001026198.html.

VanTrees:2001:DEM

[Van01b] Harry L. Van Trees. Detection, estimation, and modulation theory. Wiley, New York, NY, USA, 2001. ISBN 0-471-09517-6 (v. 1), 0-471-44678-5 (v. 2), 0-471-09390-4 (v. 4), 0-471-10793-X. ???? pp. LCCN TK5102.5 .V3 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/wiley022/2002276850.
html.

VanTrees:2002:OAP

[Van02] Harry L. Van Trees. Optimum array processing, volume 4 of Detection, estimation and modulation theory. Wiley-Interscience, New York, NY, USA, 2002. ISBN 0-471-09390-4 (hardcover). 1443 pp.

Vandev:2004:ISD

[Van04] D. L. Vandev. Interactive stepwise discriminant analysis in MATLAB. Pliska, Studia mathematica Bulgarica, 16:291–298, 2004. ISSN 0204-9805.

Varga:2004:GHC

[Var04] Richard S. Varga. Geršgorin and His Circles, volume 36 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21100-4 (hardcover), 3-642-17798-0 (e-book). ISSN 0179-3632. x + 226 pp. LCCN QA193 .V37

2004. URL http://www.
loc.gov/catdir/enhancements/
fy0815/2004104814-d.html.

Vazquez:2016:NDI

[Váz16] R. Vázquez. A new design for the implementation of isogeometric analysis in Octave and Matlab: GeoPDEs 3.0. Computers and Mathematics with Applications, 72(3):523–554, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122116302681.

Villaverde:2018:PTI

[VBB18] Alejandro F. Villaverde, Kolja Becker, and Julio R. Banga. PREMER: a tool to infer biological networks. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 15(4): 1193–1202, July 2018. CO-DEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

VanDeun:2006:ACI

[VC06a] Joris Van Deun and Ronald Cools. Algorithm 858: Computing infinite range integrals of an arbitrary product of Bessel functions. ACM Transactions on Mathematical Software, 32(4):580–596, December 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

VanDeun:2006:MIA

[VC06b] Joris Van Deun and Ronald Cools. A Matlab implementation of an algorithm for computing integrals of products of Bessel functions. In Mathematical software—ICMS 2006, volume 4151 of Lecture Notes in Comput. Sci., pages 284–295. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Volakis:1998:FEM

[VCK98] John Leonidas Volakis, A. (Arindam) Chatterjee, and Leo C. Kempel. Finite element method for electromagnetics: antennas, microwave circuits, and scattering applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-3425-6. xvi + 344 pp. LCCN TK7867.2 .V65 1998. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        bios/wiley043/97048768.
                                        html;   http://www.loc.
                                        gov/catdir/description/
                                        wiley036/97048768.html;
                                         http://www.loc.gov/catdir/
                                        toc/onix07/97048768.html.

vanderHeijden:2005:CPE

[vdH05] Ferdinand van der Heijden. Classification, Parameter Estimation and State Estimation: an Engineering Approach Using Matlab. Wiley, New York, NY, USA, 2005. ISBN 0-470-09015-4. ???? pp. LCCN ????

vanOverschee:1996:SIL

[vdM96] Peter van Overschee and Bart L. R. de Moor. Subspace identification for linear systems: theory, implementation, applications. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1996. ISBN 0-7923-9717-7. xiv + 254 pp. LCCN QA402 .O94 1996.

Vehkalahti:2007:SBRa

[Veh07] Kimmo Vehkalahti. Short book review: Applied Statistics Using SPSS, STATIS-TICA, MATLAB and R, 2nd Edition, by Joaquim P. Marques de Sá. International Statistical Review = Revue Internationale de Statistique, 75(3):421– 422, December 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
41509892.

Venkataraman:2002:AOM

[Ven02] P. Venkataraman. Applied optimization with MATLAB programming. Wiley, New York, NY, USA, 2002. ISBN 0-471-34958-5 (cloth). xvii + 398 pp. LCCN QA402.5 .V42 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley044/2001026938.
html;   http://www.loc.
gov/catdir/description/
wiley037/2001026938.html;
 http://www.loc.gov/catdir/
toc/onix05/2001026938.
html.

Verzani:2009:CSH

[Ver09]  John Verzani.  Compu-

tational statistics handbook with MATLAB. Second edition [book review of MR2404036]. SIAM Review, 51(3):654–656, 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vesely:1994:MOL

[Ves94] Vítězslav Veselý. MAT-LAB operators of left/right division and generalized inverse. In Summer School MATLAB 93 Proceedings (Blato, 1993), volume 4 of Folia Fac. Sci. Natur. Univ. Masaryk. Brun. Math., pages 25–40. Masaryk Univ., Brno, 1994.

Vesely:1998:FAF

[Ves98] Vítězslav Veselý. Fast algorithms of Fourier and Hartley transform and their implementation in MAT-LAB. In Summer School DATASTAT 97, Proceedings (Hodůnka), volume 7 of Folia Fac. Sci. Natur. Univ. Masaryk. Brun. Math., pages 185–232. Masaryk Univ., Brno, 1998.

VanLoan:2010:ITC

[VF10] Charles F. Van Loan and K.-Y. Daisy Fan. Insight through computing: a MAT-LAB introduction to computational science and engineering. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2010. ISBN 0-89871-691-8. xviii + 434 pp. LCCN QA297 .V25

                                        2010. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy1007/2009030277-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1007/2009030277-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1007/2009030277-t.html.

Vaz:2004:SSI

[VFG04] A. Ismael F. Vaz, Edite M. G. P. Fernandes, and M. Paula S. F. Gomes.

SIPAMPL: Semi-infinite programming with AMPL. ACM Transactions on Mathematical Software, 30(1):47– 61, March 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Varmaz:2013:RAW

[VFV13] Armin Varmaz, Christian Fieberg, and Andreas Var-

wig.  RMatlab-app2web:

Web deployment of R/ MATLAB applications. Journal of Statistical Software, 54(5):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v54/
i05.

Verboven:2004:MSR

[VH04] Sabine Verboven and Mia Hubert. MATLAB software for robust statistical methods. In COMPSTAT 2004—Proceedings in Computational Statistics, pages 1941–1946. Physica, Heidelberg, 2004.

Verboven:2010:MLL

[VH10] Sabine Verboven and Mia Hubert. MATLAB library LIBRA. WIREs Computational Statistics, 2(4):509– 515, July/August 2010. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Versypt:2017:PAO

[VHM17] Ashlee N. Ford Versypt, Grace K. Harrell, and Alexandra N. McPeak.

ACEInhibPKPD: An open-

source MATLAB app for a pharmacokinetic/pharmacodynamic model of ACE inhibition. Journal of Open Source Software, 2(17):340:1–340:2, September 2017. CO-DEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/
10.21105/joss.00340.

Vidakovic:2011:SBS

[Vid11] Brani Vidakovic. Statistics for bioengineering sciences: with MATLAB and Win-BUGS support. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4614-0393-6 (hardcover), 1-4614-0394-4 (ebook). ISSN 1431-875X. xvi + 753 pp. LCCN R857.M34 V53 2011.

Viniotis:1998:PRP

[Vin98]  Yannis Viniotis.  Prob-

ability and random processes for electrical engineers. McGraw-Hill series in electrical and computer engineering. Communications and signal processing. WCB/McGraw-Hill, Boston, MA, USA, 1998. ISBN 0-07-067491-4. xviii + 676 pp. LCCN QA273 .V55 1998. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/mh023/97037609.
html;   http://www.loc.
gov/catdir/toc/mh022/97037609.
html.

Vittoria:2011:MDW

[Vit11] C. Vittoria. Magnetics, dielectrics, and wave propagation and MATLAB codes. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-4199-3. ???? pp. LCCN QC760.4.M37 V58 2011.

Voyiadjis:2005:MCM

[VK05] G. Z. Voyiadjis and Peter Issa Kattan. Mechanics of composite materials with MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-24353-4 (hardcover). xi + 336 pp. LCCN TA418.9.C6 V69 2005. Includes CD-ROM.

Vuorinen:2016:DGT

[VK16] V. Vuorinen and K. Keskinen. DNSLab: a gateway to turbulent flow simulation in Matlab. Computer Physics Communications, 203(??): 278–289, June 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465516300388.

VanRooyen:2000:STP

[VLV00] Pieter Van Rooyen, M. N. Lötter, and Danie Van Wyk. Space-time processing for CDMA mobile communications, volume SECS 544 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 2000. ISBN 0-7923-7759-1. xx + 308 pp. LCCN TK5103.45 .V36 2000.

Vogel:2002:CMI

[Vog02] Curtis R. Vogel. Computational methods for inverse problems. Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2002. ISBN 0-89871-507-5. xvi + 183 pp. LCCN QA377 .V575 2002.

vonGrunigen:1993:DSG

[von93] Daniel Ch. von Grunigen. Digitale Signalverarbeitung, Grundlagen und Anwendungen Beispiele und Ubungen mit MATLAB [English: Digital Signal Processing: Fundamentals and Applications; Examples and Exercises with MAT-LAB]. Aufbau Taschenbuch Verlag, Berlin, Germany, 1993. ISBN 3-8007-1971-1, 3-905214-16-4.

LCCN ???? URL ftp:/

/ftp.mathworks.com/pub/
books/vonGrunigen/.

Vergara-Perez:2016:MMP

[VPM16] Sandra Vergara-Perez and Marcelo Marucho. MP-BEC, a Matlab Program for Biomolecular Electrostatic Calculations. Computer Physics Communications, 198(??):179–194, January 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465515003288.

Villada:2023:AMB

[VRVAC23] Sebastian Montoya Villada, Erick Reyes-Vera, and Mauricio Arias-Correa. AnIMAGE: a MATLAB-based tool for generating microstrip antennas with complex shapes. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S235271102300198X.

Vich:2000:CF

[VS00] Robert Vích and Zdeněk Smékal. Číslicové filtry, volume roč. 105, 2000 čís. spisu 482 of Česká matice technická. Academia, Prague, Czechoslovakia, 2000. ISBN 80-200-0761-X. 218 (est.) pp. LCCN T4 .C42 (QA297).

Vandebril:2008:MCS

[VvBM08] Raf Vandebril, Marc van Barel, and Nicola Mastronardi. Matrix Computations and Semiseparable Matrices. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-8714-3 (vol. 1, hardcover), 0-8018-9052-7 (vol. 2). xviii + 575 (vol. 1), xvi + 498 (vol.

2) pp. LCCN QA188 .V36

                                        2008. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0740/2007030657-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0740/2007030657-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/ecip0723/
                                        2007030657.html.

Vasiliu:1993:USA

[VVM93] N. Vasiliu, D. Vasiliu, and J. C. Mare. Using Simulink and ACSL for the simulation of an hydraulic power system. In Pave [Pav93], pages 185–192. ISBN 1-56555-056-0. LCCN QA76.9.C65 E97 1993.

Veres:1996:UGM

[VWHK96] S. M. Veres, D. S. Wall, S. Hermsmeyer, and A. V. Kuntsevich. Using GBT for MATLAB, version 5.1, in identification and control. In Anonymous [Ano96v], pages 216–221. ISBN 0-85296-668-7, 0-85296-666-

0. ISSN 0537-9989. LCCN TJ212.2 .U32 1996; TK5.I4 no.427. Two volumes.

Velasco:2025:SUF

[VZDCM25] Manel Velasco, Isiah Zaplana, Arnau Dòria-Cerezo, and Pau Martí. Symbolic and user-friendly geometric algebra routines (SUGAR) for computations in Matlab. ACM Transactions on Mathematical Software, 51 (2):11:1–11:??, June 2025. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Wicks:1997:PIM

[W^{+}97] Alfred L. Wicks et al., editors. Proceedings of the 15th International Modal Analysis Conference: IMAC: February 3–6, 1997, Sheraton World Resort, Orlando, Florida, volume 3089 of Proceedings of the SPIE — The International Society for Optical Engineering. Society for Experimental Mechanics Inc, Bethel, CT, USA, 1997. ISBN 0-912053-53-4. ISSN 1046-6770. LCCN TS510.S63 v.3089. Two volumes.

Welfert:1996:ANM

[WA96] Bruno D. Welfert and Ricardo Aguilar. Applied numerical methods and graphical visualization. Computer Applications in Engineering Education, 4(2): 127–143, 1996. CODEN CAPEED. ISSN 1061-3773.

Walnut:2002:IWA

[Wal02] David F. Walnut. An introduction to wavelet analysis. Applied and numerical harmonic analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 2002. ISBN 0-8176-3962-4. xvii + 449 pp. LCCN QA403.3 .W335 2002.

Walker:2018:FMC

[Wal18] Shawn W. Walker. FELIC-ITY: a Matlab/C++ toolbox for developing finite element methods and simulation modeling. SIAM Journal on Scientific Computing, 40(2):C234–C257, ???? 2018. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Wang:2015:MPM

[Wan15] Chun-Chao Wang. A MAT-LAB package for multivariate normality test. Journal of Statistical Computation and Simulation, 85(1): 166–188, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wartak:2013:CPI

[War13] Marek S. Wartak. Computational photonics: an introduction with MAT-LAB. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-00552-3 (hardcover). xiii + 452 pp. LCCN TK8304 .W37 2013. URL http:

//assets.cambridge.org/
97811070/05525/cover/9781107005525.
jpg.

Watts:1993:PPP

[Wat93] J. W. Watts. Phase plane portraits of second order systems using Simulink. In Grayson [Gra93], page ?? ISBN 0-7803-1482-4, 0-7803-1483-2, 0-7803-1484-

0. ISSN 0190-5848. LCCN ????

Watkins:2002:FMC

[Wat02] David S. Watkins. Fundamentals of matrix computations. Pure and applied mathematics. Wiley-Interscience, New York, NY, USA, second edition, 2002. ISBN 0-471-21394-2. xiii + 618 pp. LCCN QA188 .W38 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley044/2002071332.
html;   http://www.loc.
gov/catdir/description/
wiley036/2002071332.html;
http://www.loc.gov/catdir/
toc/wiley023/2002071332.
html.

Wilkes:1999:FMC

[WB99] James O. Wilkes and Stacy G. Bike. Fluid mechanics for chemical engineers. Prentice Hall international series in the physical and chemical engineering sciences. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-739897-2. xiv + 599 pp. LCCN TP155.7 .W55 1999.

Willkomm:2012:IDD

[WB12] Johannes Willkomm and Christian H. Bischof. The impact of dynamic data reshaping on adjoint code generation for weakly-typed languages such as Matlab. In Forth et al. [FHP^{+}12], pages 127–138. CODEN LNCSA6. ISBN 3-642-30022-7 (print), 3-642-30023-5 (e-book). ISSN 1439-7358. LCCN ????

URL http://link.springer.
com/content/pdf/10.1007/
978-3-642-30023-3_12. Pro-

ceedings of the Sixth International Conference on Automatic Differentiation (AD2012) held July 23–27, 2012, in Fort Collins, Colorado, USA.

Wiik:2016:CBV

[WB16] Jonatan Wiik and Pontus Boström. Contract-based verification of MATLABstyle matrix programs. Formal Aspects of Computing, 28(1):79–107, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL

http://link.springer.com/
article/10.1007/s00165-
015-0353-z.

Wilson:1992:UMA

[WCS92] H. B. Wilson, G. S. Chang, and D. Singh. Using MAT-LAB to analyze cable dynamics and related numerical problems. American Society of Mechanical Engineers, Petroleum Division (Publication) PD, 44:87–95, 1992. CODEN ASMPEX.

Watts:1996:SRS

[WD96] J. W. Watts and T. E.

Dwan. Sensitivity and robustness simulations using MATLAB. In Hamza [Ham96], pages 105–107. ISBN 0-88986-201-X. LCCN QA76.9.C65I28 1996.

Wicks:1998:PIM

[WD98] Alfred L. Wicks and Dominick J. DeMichele, editors. Proceedings of the 16th International Modal Analysis Conference: February 2– 5, 1998, Fess Parker’s Doubletree Resort, Santa Barbara, Calif, volume 3243. Society for Experimental Mechanics Inc, Bethel, CT, USA, 1998. ISBN 0-912053-59-3. ISSN 1046-6770. LCCN TA654.15 .I57 1998. Two volumes.

Wolkenhauer:1996:FST

[WE96] O. Wolkenhauer and J. M.

Edmunds. A fuzzy systems toolbox for use with MATLAB. In Anonymous [Ano96f], pages 9–?? ISBN ???? ISSN 0963-3308. LCCN ????

Weaver:1997:PTM

[Wea97] Mark Weaver. Programmer’s toolchest — modeling with Matlab designing and implementing complicated systems is a complex process. Mark describes how he uses MatLab. Dr. Dobb’s Journal of Software Tools, 22(11):80–83, 94–95, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Webb:1997:LER

[Web97] Peter Webb. Letter to the Editor: Response to Wilson: Teach science and software engineering with Matlab. IEEE Computational Science & Engineering, 4(2): 4–5, April/June 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1997/pdf/c2004.
pdf.

Weiss:2012:FRR

[Wei12] Andreas Weiss. A flexible and reliable radar simulator in Matlab OOP for optimizing tracking algorithms. Lecture Notes in Computer Science, 6927: 472–476, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/978-3-642-27549-
4_60.

Wellstead:1993:ACS

[Wel93] P. E. Wellstead. Adaptive control and signal processing enhancements to SIMULINK. In Anonymous [Ano93c], pages 4–?? ISBN ???? LCCN ????

Wallack:1998:MMM

[WEM98] Aaron Wallack, Ioannis Z.

Emiris, and Dinesh Manocha. MARS: a MAPLE/MATLAB/ C resultant-based solver. In Gloor [Glo98], pages 244– 251. ISBN 1-58113-002-3. LCCN ???? URL http://

www.acm.org/pubs/citations/
proceedings/issac/281508/
p244-wallack/.

Wentworth:2005:FEE

[Wen05] Stuart M. Wentworth. Fundamentals of electromagnetics with engineering applications. Wiley, New York, NY, USA, 2005. ISBN 0-471-26355-9. xx + 588 pp. LCCN TK146 .W435 2005. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/wiley041/2004049216.
html.

Werner:2003:DSM

[Wer03]  Martin Werner.  Digi-

tale Signalverarbeitung mit MATLAB. (German) [Digital Signal Processing with MATLAB]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 2003.

ISBN 3-528-13930-7. x +

305 pp. LCCN ????

Wette:2020:SPO

[Wet20] Karl Wette. SWIGLAL: Python and Octave interfaces to the LALSuite gravitational-wave data analysis libraries. SoftwareX, 12 (??):Article 100634, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711020303472.

Whidborne:1995:MMT

[WGP95] J. F. Whidborne, D.-W. Gu, and I. Postlethwaite. MODCONS — a MAT-LAB toolbox for multiobjective control system design. In Anonymous [Ano95b], pages 9/1–9/4. CODEN DCILDN. ISBN ???? ISSN 0963-3308. LCCN ????

White:1991:EAA

[Whi91] B. A. White. Eigenstructure assignment for aerospace applications. MATLAB implementation. IEE Conference Publication, 2(332): 738–743, 1991. CODEN IECPB4. ISSN 0537-9987 (??invalid ISSN checksum??).

Whitaker:2000:SMA

[Whi00] Jerry C. Whitaker, editor. Signal measurement, analysis, and testing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-0048-7. 313 (est.) pp. LCCN TK5102.9 .S537 2000.

White:2004:CMM

[Whi04] R. E. (Robert E.) White.

Computational Mathematics: Models, Methods, and Analysis with MATLAB and MPI. Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-364-2. xvi + 385 pp. LCCN QA297 .W495 2004.

White:2007:EMM

[Whi07] Robert E. White. Elements of matrix modeling and computing with MAT-LAB(R). Chapman and Hall/CRC, Boca Raton, FL, USA, 2007. ISBN 1-58488-627-7. xvi + 402 pp.

Wilson:2002:AMM

[WHT02] H. B. Wilson, David Halpern, and Louis H. Turcotte. Advanced mathematics and mechanics applications using MATLAB. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2002. xiii + 678 pp. LCCN TA345 .W55 2003.

Wieseman:1994:WER

[Wie94] C. D. Wieseman, editor. Workshop entitled “The role of computers in LARC R and D” — 1994 Jun: Hampton; VA, number 10159 in NASA Conference Publication. National Aeronautics and Space Administration, Washington, DC, USA, 1994. ISBN ???? ISSN 0191-7811. LCCN ????

Wikle:2004:BRB

[Wik04] Christopher K. Wikle. Book review: Computational Statistics Handbook with MATLAB, by Wendy Martinez; Angel R. Martinez. Journal of the American Statistical Association, 99 (466):566, April 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/miranda.asa.catchword.
org/vl=1515603/cl=66/nw=
1/rpsv/cw/asa/01621459/
v99n466/s36/p566;   http:
//www.jstor.org/stable/
27590422.

Williams:1996:LAA

[Wil96a] Gareth Williams. Linear Algebra with Applications. Wm. C. Brown Publishers, Dubuque, IA, USA, 1996. ISBN 0-697-26849-7. xix + 512 + 40 + 4 pp. LCCN ????

Wilson:1996:FSO

[Wil96b] Raymond G. Wilson. Fourier Series and Optical Transform Techniques in Contemporary Optics: An Introduction. Wiley, New York, NY, USA, 1996. ISBN 0-471-30357-7. xvii + 325 pp. LCCN QC454.F7 W55 1995.

Wilson:1999:SDA

[Wil99] Hugh R. (Hugh Reid) Wilson. Spikes, decisions, and actions: the dynamical foundations of neuroscience. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-852431-5 (hardcover), 0-19-852430-7 (paperback). ix + 307 pp. LCCN QP357.5 .W54 1999.

Williams:2003:PSR

[Wil03] Richard H. Williams. Probability, statistics, and random processes for engineers. Thomson, Brooks/Cole, Pacific Grove, CA, 2003. ISBN 0-534-36888-3. xii + 372 pp. LCCN TA340 .W52 2003.

Witkovsky:2004:MAT

[Wit04] Viktor Witkovský. Matlab algorithm TDIST: the distribution of a linear combination of Student’s T random variables. In COMP-STAT 2004—Proceedings in Computational Statistics, pages 1995–2002. Physica, Heidelberg, 2004.

Wilkie:2002:CEI

[WJK02] Jacqueline Wilkie, Michael Johnson, and Reza Katebi. Control engineering: an introductory course. Palgrave, Basingstoke, 2002. ISBN 0-333-77129-X. xv + 750 pp.

Wrixon:1995:MTC

[WK95] A. Wrixon and M. P. Kennedy. A MATLAB tool for calculating Lyapunov exponents from a chaotic time series. In Kennedy [Ken95], pages 205–208. ISBN 1-898473-24-2. LCCN ????

Waldron:1999:KDD

[WK99] Kenneth J. Waldron and Gary L. Kinzel. Kinematics, dynamics, and design of machinery. Wiley, New York, NY, USA, 1999. ISBN 0-471-58399-5 (cloth). xiv + 640 pp. LCCN TJ175 .W35 1999. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley041/98025607.
html;   http://www.loc.
gov/catdir/description/
wiley032/98025607.html;
 http://www.loc.gov/catdir/
toc/onix03/98025607.html.

Westwick:2003:INP

[WK03] David T. Westwick and Robert E. Kearney. Identification of nonlinear physiological systems. IEEE Press series on biomedical engineering. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-471-27456-9 (cloth). xii + 261 pp. LCCN QP33.6.M36 W475 2003. URL ftp://

uiarchive.cso.uiuc.edu/

pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley044/2003043255.
html;   http://www.loc.
gov/catdir/description/
wiley037/2003043255.html;
http://www.loc.gov/catdir/
toc/wiley032/2003043255.
html.

Wells:1994:EEE

[WL94] D. P. Wells and J. Lebaric.

EMAG 2.0 — enhanced 2D electrostatic and magnetostatic solver in MATLAB. In Terzuoli [Ter94], pages 522–527. ISBN ???? LCCN ???? Two volumes.

Woods:1997:MSD

[WL97] Robert L. Woods and Kent L. Lawrence. Modeling and Simulation of Dynamic Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-337379-1. xix + 521 pp. LCCN TA342 .W66 1997.

Wallisch:2009:MN

[WLB^{+}09] Pascal Wallisch, Michael Lusignan, Marc Benayoun, Tanya I. Baker, Adam S. Dickey, and Nicholas G. Hatsopoulos. Matlab(R) for neuroscientists. Elsevier/Academic Press, Amsterdam, 2009. ISBN 0-12-374551-9. xiv + 394 pp. An introduction to scientific computing in Matlab.

Weeks:1995:AEM

[WM95] R. W. Weeks and J. J. Moskwa. Automotive engine modeling for realtime control using MAT-LAB/SIMULINK. In SAE [SAE95b], pages 123–138. ISBN 1-56091-630-3. LCCN TL240.V4237 1995.

Weeks:1996:AEM

[WM96] R. W. Weeks and J. J. Moskwa. Automotive engine modeling for real-time control using MATLAB/ SIMULINK. S.A.E. transactions, 104(5):295–??, ???? 1996. ISSN 0096-736X.

Wang:2002:VPC

[WOZ02] Yao Wang, Jörn Ostermann, and Ya-Qin Zhang. Video processing and communications. Prentice-Hall signal processing series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-017547-1. xxvii + 595 pp. LCCN TK5105.2 .W36 2002.

Wette:2018:POL

[WPK^{+}18] Karl Wette, Reinhard Prix, David Keitel, Matthew Pitkin, Christoph Dreissigacker, John T. Whelan, and Paola Leaci. OctApps: a library of Octave functions for continuous gravitationalwave data analysis. Journal of Open Source Software, 3 (26):707:1–707:3, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00707.

Weideman:2000:MDM

[WR00] J. A. C. Weideman and S. C.

Reddy. A MATLAB differentiation matrix suite. ACM Transactions on Mathematical Software, 26(4):465–519, December 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Westland:2004:CCS

[WR04] Stephen Westland and Caterina Ripamonti. Computational colour science using MATLAB. Wiley, New York, NY, USA, 2004. ISBN 0-470-84562-7 (cloth). x + 207 pp. LCCN QC495.8 .W47 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley041/2003021483.
;   http://www.loc.gov/
catdir/description/wiley041/
2003021483.html;   http:
//www.loc.gov/catdir/toc/
wiley041/2003021483.;
http://www.loc.gov/catdir/
toc/wiley041/2003021483.
html; UCLA.

Weinstein:2016:STO

[WR16] Matthew J. Weinstein and Anil V. Rao. A source transformation via operator overloading method for the automatic differentiation of mathematical functions in MATLAB. ACM Transactions on Mathematical Software, 42(2):11:1– 11:42, June 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Weinstein:2017:AAT

[WR17] Matthew J. Weinstein and Anil V. Rao. Algorithm 984: ADiGator, a toolbox for the algorithmic differentiation of mathematical functions in MATLAB using source transformation via operator overloading. ACM Transactions on Mathematical Software, 44(2):21:1– 21:25, September 2017. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:

//dl.acm.org/citation.
cfm?id=3104990.

Wright:1997:UMM

[Wri97] M. Wright. The use of Matlab in model based processing. In Anonymous [Ano97l], pages 4–?? ISBN ???? ISSN 0963-3308. LCCN ????

Whittle:2004:AIK

[WS04] Jon Whittle and Johann Schumann. Automating the implementation of Kalman filter algorithms. ACM Transactions on Mathematical Software, 30(4):434–453, December 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Wouwer:2005:MIU

[WSST05] A. Vande Wouwer, P. Saucez, W. E. Schiesser, and S. Thompson. A MATLAB implementation of upwind finite differences and adaptive grids in the method of lines. Journal of Computational and Applied Mathematics, 183(2):245–258, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Wilson:1994:AMM

[WT94] Howard B. Wilson and Louis H. Turcotte. Advanced Mathematics and Mechanics Applications Using MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1994. ISBN 0-8493-2482-3. 405 pp. LCCN TA345 .W55 1994. US$61.95.

Wilson:1997:AMM

[WT97] Howard B. Wilson, Jr. and Louis H. Turcotte. Advanced Mathematics and Mechanics Applications Using Matlab. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 1997. ISBN 0-8493-1686-3. 632 pp. LCCN TA345.W55 1997. US$69.95. URL http://

www.crcpress.com.

Wilson:2003:AMM

[WTH03] Howard B. Wilson, Louis H. Turcotte, and David Halpern. Advanced mathematics and mechanics applications using MATLAB. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2003. ISBN 1-58488-262-X. xiv + 678 pp. LCCN TA345 .W55 2003.

Wagenpfeil:2000:MBS

[WTL00] Stefan Wagenpfeil, Uwe Treiber, and Antonie Lehmer. A MATLAB-based software tool for changepoint detection and nonlinear regression in dose-response relationships. Lecture Notes in Computer Science, 1933: 190–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1933/19330190.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1933/19330190.
                                        pdf.

Wunsch:2005:CVA

[Wun05] A. David Wunsch. Complex variables with applications. Pearson Addison-Wesley, Boston, third edition, 2005. ISBN 0-201-75609-9. xv + 677 pp. LCCN QA331.7 .W86 2005.

Wheless:1994:FDA

[WW94] W. P. Wheless and C. S.

Wheless. Finite difference analysis with MAT-LAB and VMAP in undergraduate instruction. In Terzuoli [Ter94], pages 536– 544. ISBN ???? LCCN ???? Two volumes.

Webb:1999:MSL

[WW99] Peter Webb and Gregory V.

Wilson. Matlab as a scripting language: a simple way to do powerful things. Dr. Dobb’s Journal of Software Tools, 24(1):94, 96– 99, January 1999. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ddj/ftp/
1999/1999_01/matlab.txt.

Welch:2006:RTD

[WWM06] Thad B. Welch, Cameron H. G. Wright, and Michael G. Morrow. Real-time digital signal processing from MATLAB to C with the TMS320Cx DSK. CRC/Taylor and Francis, Boca Raton, FL, USA, 2006. ISBN 0-8493-7382-4. xxxiii + 363 pp. LCCN TK5102.9 .W44 2006.

Xue:2007:LFC

[XCA07] Dingyü Xue, YangQuan Chen, and Derek P. Atherton. Linear feedback control: analysis and design with MATLAB. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2007. ISBN 0-89871-638-1. xii + 356 pp. LCCN TA345 .Z84 2007.

Xu:2020:MSM

[XSS20] Qimen Xu, Abhiraj Sharma, and Phanish Suryanarayana. M-SPARC: Matlab-simulation package for ab-initio realspace calculations. SoftwareX, 11(??):Article 100423, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711019303966.

Xu:2022:UGC

[XWZ^{+}22] Xiong Xu, Shuling Wang, Bohua Zhan, Xiangyu Jin, Jean-Pierre Talpin, and Naijun Zhan. Unified graphical co-modeling, analysis and verification of cyberphysical systems by combining AADL and Simulink/ Stateflow. Theoretical Computer Science, 903(??):1–25, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0304397521006691.

Xu:2023:UGC

[XWZ^{+}23] Xiong Xu, Shuling Wang, Bohua Zhan, Xiangyu Jin, Naijun Zhan, and Jean-Pierre Talpin. Unified graphical co-modelling, analysis and verification of cyber-physical systems by combining AADL and Simulink/Stateflow. ACM SIGADA Ada Letters, 43 (1):46–49, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3631483.3631487.

Xiong:2024:GGP

[XZL24] Hongqiang Xiong, Zhiyu Zhang, and Jing Li. GPRlab: a ground penetrating radar data processing and analysis software based on MATLAB. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711024000918.

Yaz:1995:UMTa

[YA95a] E. E. Yaz and A. Azemi.

Utilizing MATLAB in two graduate electrical engineering courses. In Bergadaa [Ber95], pages 2C6.1–2C6.4. CODEN PFECDR. ISBN 2-911209-00-1. ISSN 0190-5848. LCCN ????

Yaz:1995:UMTb

[YA95b] E. E. Yaz and A. Azemi.

Utilizing MATLAB in two graduate electrical engineering courses. In Budny [Bud95], pages 2C6.1– 2C6.4. ISBN 0-7803-3023-4, 0-7803-3022-6, 0-7803-3024-2, 0-7803-3025-0. ISSN 0190-5848. LCCN T62 .F76 1995. Two volumes.

Yaniv:1999:QFD

[Yan99]  Oded Yaniv.  Quantita-

tive feedback design of linear and nonlinear control systems, volume [SECS 509] of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1999. ISBN 0-7923-8529-2. xxiii + 369 pp. LCCN TJ216 .Y36 1999.

Yang:2005:SSS

[Yan05]  Bingen Yang.  Stress,

Strain, and Structural Dynamics: An Interactive Handbook of Formulas, Solutions, and MATLAB Toolboxes. Elsevier Academic Press, Amsterdam, The Netherlands, 2005. ISBN 0-12-787767-3. xvii + 942 pp. LCCN TA648.3 .Y36 2005. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip051/2004022861.
html.

Yang:2017:CMI

[Yan17] Yaguang Yang. CurveLP — a MATLAB implementation of an infeasible interiorpoint algorithm for linear programming. Numerical Algorithms, 74(4):967– 996, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11075-016-0180-1.

Yang:2005:ANM

[YCCM05] Won Young Yang, Wenwu Cao, Tae-Sang Chung, and John Morris. Applied numerical methods using MATLAB. Wiley, New York, NY, USA, 2005. ISBN 0-471-69833-4. xiv + 509 pp. LCCN QA297 .Y36 2005. UK£50.50. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip0418/2004013108.
html.

Yates:1999:PSP

[YG99] Roy D. Yates and David J.

Goodman. Probability and stochastic processes: a friendly introduction for electrical and computer engineers. Wiley, New York, NY, USA, 1999. ISBN 0-471-17837-3 (cloth). xvii + 454 pp. LCCN QA273 .Y384 1999.

Yang:2022:AMB

[YHC^{+}22] Tao Yang, Lizhu Hu, Anja Chen, Maroš Tunák, Shangyong Zhang, Deyou Yu, Xiaodong Tan, Michal Petrů, Ivan Mašín, and Sundaramoorthy Palanisamy. AFDeter: a MATLABbased tool for simple and rapid determination of the structural parameters and the airflow-related properties of fibrous materials. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711022001315.

Yu:2024:SSC

[YJ24] Bo Yu and Ruijiang Jing. SCTBEM: a scaled coordinate transformation boundary element method with 99-line MATLAB code for solving Poisson’s equation. Computer Physics Communications, 300(??): ??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0010465524001085.

Yoon:1994:NSO

[YKS94] T.-S. Yoon, Seung Woo Kim, and W. Shin. New software for old diffractometer: A Matlab based single crystal data collection softwate system. In Anonymous [Ano94e], page PM01. ISBN ???? ISSN 0569-4221. LCCN ????

Yen:1999:FLI

[YL99] John Yen and Reza Langari. Fuzzy logic: intelligence, control, and information. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-525817-0. xxiv + 548 pp. LCCN QA9.64 .Y46 1999.

Yang:2009:KMT

[YMLW09] Chao Yang, Juan C. Meza, Byounghak Lee, and Lin-Wang Wang. KSSOLV — a MATLAB toolbox for solving the Kohn–Sham equations. ACM Transactions on Mathematical Software, 36 (2):10:1–10:35, March 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Yang:2019:BMT

[YR19] Jingjing Yang and Peng Ren. BFDA: a MATLAB toolbox for Bayesian functional data analysis. Journal of Statistical Software, 89(??):??, ???? 2019. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v089i02;
https://www.jstatsoft.
org/index.php/jss/article/
view/v089i02/v89i02.pdf.

Yuksel:1996:MIM

[Yuk96] Ibrahim Yuksel. MATLAB ile Mühendislik Sistemlerinin Analizi ve Cözümü (Turkish) [Analysis and Solution of Engineering Systems with MATLAB]. Uludag University, Bursa, Turkey, 1996. ISBN 975-564-049-5. ???? pp. LCCN ????

Zolzer:2002:DDA

[ZA02] Udo Zölzer and Xavier Amatriain, editors. DAFX: digital audio effects. Wiley, New York, NY, USA, 2002. ISBN 0-471-49078-4. xix + 533 pp. LCCN TK5105.8863 D34 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley036/2002280336.
html;   http://www.loc.
gov/catdir/toc/wiley022/
2002280336.html.

Zaghloul:2011:ACF

[ZA11] Mofreh R. Zaghloul and Ahmed N. Ali. Algorithm 916: Computing the Faddeyeva and Voigt functions. ACM Transactions on Mathematical Software, 38(2):15:1–15:22, December 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Zag16].

Zaghloul:2016:RAC

[Zag16] Mofreh R. Zaghloul. Remark on “Algorithm 916: Computing the Faddeyeva and Voigt Functions”: Efficiency improvements and Fortran translation. ACM Transactions on Mathematical Software, 42(3):26:1– 26:9, May 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [ZA11].

Zalewski:1996:RSE

[Zal96] J. Zalewski, editor. Realtime systems education. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7649-3, 0-8186-7651-5. LCCN QA76.54 .R4294 1996. IEEE Computer Society Press order number PR07649. IEEE Order Plan catalog number 96TB1000060.

Zarchan:1997:TSM

[Zar97] Paul Zarchan. Tactical and strategic missile guidance, volume 176 of Progress in astronautics and aeronautics. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024–2518, third edition, 1997. ISBN 1-56347-254-6. xxv + 611 pp. LCCN TL507 .P75 vol. 176.

Zhan:2008:SBF

[ZC08] Yuan Zhan and John A.

Clark. A search-based framework for automatic testing of MATLAB/Simulink. The Journal of Systems and Software, 81(2):262– 285, February 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:1998:ERC

[ZD98] Kemin Zhou and John Comstock Doyle. Essentials of robust control. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-525833-2. xvii + 411 pp. LCCN QA402.3 .Z475 1998.

Zielke:2003:GLL

[ZD03] Gerhard Zielke and Volker Drygalla. Genaue Lösung linearer Gleichungssysteme. (German) [Exact solution of linear systems of equations]. Mitteilungen der Gesellschaft für Angewandte Mathematik und Mechanik, 26(??):7–107, ???? 2003. ISSN 0936-7195.

URL http://www.wiley-
vch.de/publish/en/journals/
alphabeticIndex/2250/.

Zietz:1995:MAS

[ZE95] Stanley Zietz and Craig Elicker. Modeling and analysis of sonar backscatter intensity for ocean bottom classification. Oceans Conference Record (IEEE), 3: 1836–1839, 1995. CODEN OCNSDK. ISSN 0197-7385. IEEE catalog number 95CB35870.

Zekollari:2017:TMF

[Zek17] Harry Zekollari. TopoZeko: a MATLAB function for 3-D and 4-D topographical visualization in geosciences. SoftwareX, 6 (??):278–284, ???? 2017. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711017300535.

Zempleni:1997:SAU

[Zem97] Andras Zempleni, editor.

Statistics at universities: its impact for society Proceedings of the workshop on statistics at universities: its impact for society: Workshop — May 1997, Budapest. Eötvös University Press, Budapest, Hungary, 1997. ISBN 963-463-082-0. LCCN ????

Zendah:1997:SFW

[Zen97] Sami R. Zendah. Simulation of four wheel steering vehicles in MATLAB. Thesis (M.S. Eng.), Wright State University, Dayton, OH, USA, 1997. x + 114 pp.

Zeng:2004:AMM

[Zen04a] Zhonggang Zeng. Algorithm 835: MultRoot—a Matlab package for computing polynomial roots and multiplicities. ACM Transactions on Mathematical Software, 30(2):218–236, June 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Zeng:2004:MPC

[Zen04b] Zhonggang Zeng. A Matlab package computing polynomial roots and multiplicities. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38 (1):28–29, March 2004. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zhang:1998:SLS

[Zha98] Yin Zhang. Solving largescale linear programs by interior-point methods under the MATLAB environment. Optimization Methods and Software, 10(1):1– 31, 1998. CODEN OM-SOE2. ISSN 1055-6788.

Zhu:2001:MSI

[Zhu01] Yucai Zhu. Multivariable system identification for process control. Pergamon Press, New York, NY, USA, 2001. ISBN 0-08-043985-

3. xxi + 349 pp. LCCN TS156.8 .Z495 2001.

Zoubir:2004:BTS

[ZI04] Abdelhak M. Zoubir and D. Robert Iskander. Bootstrap techniques for signal processing. Cambridge, Cambridge, [England], 2004. ISBN 0-521-83127-X (HB). xiv + 217 pp. LCCN TK5102.9 .Z68 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam041/2004040407.
html;   http://www.loc.
gov/catdir/toc/cam041/
2004040407.html.

Ziemer:1997:EEP

[Zie97] Rodger E. Ziemer. Elements of engineering probability and statistics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-02-431620-2. xiv + 322 pp. LCCN TK153 .Z525 1997.

Zinober:1993:MDS

[Zin93] Alan S. I. Zinober. Matlab design of sliding mode multivariable control systems. In IEEE [IEE93b], pages 2471–2476. ISBN 0-7803-1299-6, 0-7803-1298-8, 0-7803-1300-3. ISSN 0888-3610. LCCN TJ 217 I11c. Four volumes.

Zilouchian:2001:ICS

[ZJ01] Ali Zilouchian and Mohammad Jamshidi, editors. Intelligent control systems using soft computing methodologies. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-1875-0. 472 (est.) pp. LCCN TJ217.5 .I5435 2001.

Zhang:2023:VMS

[ZJKS23] Boqin Zhang, Xin Jing, Shashikant Kumar, and Phanish Suryanarayana. Version 2.0.0 — M-SPARC: Matlab-Simulation Package for Ab-initio Real-space Calculations. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022002138.

Zelyk:2004:CTI

[ZL04] Ya. I. Zelyk and M. M. Lychak. Computer technology of interval-set analysis in MATLAB. Kibernet. Sistem. Anal., 1(1):122–138, 190, 2004. ISSN 0023-1274.

Zeng:2013:NMT

[ZL13] Zhonggang Zeng and Tien-Yien Li. NAClab: a Matlab toolbox for numerical algebraic computation. ACM Communications in Computer Algebra, 47(3–4):170– 173, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zhao:2017:TXB

[ZL17] Xueqian Zhao and Zhonghai Lu. A tool for xMASbased modeling and analysis of communication fabrics in Simulink. ACM Transactions on Modeling and Computer Simulation, 27(3): 16:1–16:??, September 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zlateva:2017:MBL

[Zla17] Ivelina Zlateva. Matlabbased length/weight relationship analysis of commercial fishery samples taken from the Black Sea (Bulgaria). Journal of FisheriesSciences.com, 11(4): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/matlabbased-
lengthweight-relationship-
analysis-of-commercial-
fishery-samples-taken-
from-the-black-sea-bulgaria.
php?aid=20582.

Zou:2023:MMT

[ZLLT23] Jiaqi Zou, Zonghao Li, Xuanying Liu, and Haonan Tong. MSCPDPLab: a MAT-LAB toolbox for transfer learning based multisource cross-project defect prediction. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022002047.

Zhao:2023:NCP

[ZLMQ23] Lingxiao Zhao, Zhiyang Li, Yue Ma, and Leilei Qu. A novel cryptocurrency price time series hybrid prediction model via machine learning with MATLAB/Simulink. The Journal of Supercomputing, 79(14):15358–15389, September 2023. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
023-05242-y.

Zhan:2019:UGC

[ZLW^{+}19] Haolan Zhan, Qianqian Lin, Shuling Wang, Jean-Pierre Talpin, Xiong Xu, and Naijun Zhan. Unified graphical co-modelling of cyber-physical systems using AADL and Simulink/ Stateflow. In de Oliveira Salazar) Ribeiro and Sampaio [dOSRS19], pages 109– 129. ISBN 3-030-31037-X (paperback), 3-030-31038-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .U86

                                        2019. URL http://link.
                                        springer.com/.

Zarchan:2000:FKF

[ZM00] Paul Zarchan and Howard Musoff. Fundamentals of Kalman filtering: a practical approach, volume 190 of Progress in astronautics and aeronautics. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024–2518, 2000. ISBN 1-56347-455-7. xx + 664 pp. LCCN TL507 .P75 vol. 190.

Zhang:2017:UVM

[ZSW^{+}17] Hanqing Zhang, Tim Stangner, Krister Wiklund, Alvaro Rodriguez, and Magnus Andersson. UmUTracker: a versatile MATLAB program for automated particle tracking of 2D light microscopy or 3D digital holography data. Computer Physics Communications, 219(??):390–399, October 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465517301820.

Ziemer:2002:PCS

[ZT02] Rodger E. Ziemer and William H. Tranter. Principles of communication: systems, modulation, and noise. Wiley, New York, NY, USA, fifth edition, 2002. ISBN 0-471-39253-7 (cloth). ix + 637 pp. LCCN TK5105 .Z54 2002. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wiley036/2001045389.
html;   http://www.loc.
gov/catdir/toc/onix07/
2001045389.html.

Ziemer:1998:SSC

[ZTF98] Rodger E. Ziemer, William H.

Tranter, and D. Ronald Fannin. Signals and systems: continuous and discrete. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 1998. ISBN 0-13-496456-X. xvii + 622 pp. LCCN QA402 .Z53 1998.

Zinober:1993:VMT

[ZW93a] A. S. I. Zinober and C. A.

Woodham. VSC MATLAB Toolbox for multivariable variable structure control systems. Pure Math. Appl., 4(1):83–97, 1993. ISSN 1218-4586.

Zinober:1993:DVS

[ZW93b] Alan S. I. Zinober and C. A. Woodham. Design of variable structure control systems using PC-MATLAB. Elektrotehniski Vestnik, 60(2-3):82–90, February–March 1993. CODEN ELVEA2. ISSN 0013-5852.

Zhang:2008:GBN

[ZZC^{+}08] Yu Nong Zhang, Yu Heng Zhang, Ke Chen, Bing Huang Cai, and Wei Mu Ma. Gradient-based neural network for solving linear matrix equations and its MAT-LAB simulative verification. Acta Scientiarum Naturalium Universitatis Sunyatseni. Zhongshan Daxue Xuebao. Ziran Kexue Ban, 47(3):26–32, 2008. CODEN CHTHAJ. ISSN 0529-6579.

Zheng:2014:IMS

[ZZG^{+}14] Liang Zheng, Huai Zhang, Taras Gerya, Matthew Knepley, David A. Yuen, and Yaolin Shi. Implementation of a multigrid solver on a GPU for Stokes equations with strongly variable viscosity based on Matlab and CUDA. The International Journal of High Performance Computing Applications, 28(1):50–60, February 2014. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/28/
1/50.full.pdf+html.